1

A Bibliography of Publications in Parallel Computing

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

03 June 2024 Version 1.100

Title word cross-reference

(k) [HRR08a]. (ν) [Kam87]. (r_{∞}, n_{1/2}, s_{1/2}) [Hoc85b]. (r_{∞}, n_{1/2}) [Tem89]. 0 [LLW^{+}15]. 0/1 [CJ92, MS99b]. 1 [LLW^{+}15, Tak03, Ume01, YL86]. 2 [AW13, BSE88, BAS13, Cap88, DZD01, EY86, Lee97, Mar92, MM96b, RVGG01, SA16, TT00, WSR14, YL86]. 2, 576, 980, 370, 000 [Tak10]. 2k [DT95]. 3 [AAS13, AW13, BFG^{+}07, BCMG^{+}07, Bri95, BAS13, CYDJ21, CDM06, CM95, Chr98, EGTD99, ER95b, EK98, FMS^{+}06, GV01, GJS93a, HKL01, HSS07, HB84, JHD^{+}22, KT97, KTN^{+}14, KD13, KYLH01, Lee97, Lip99b, LCLL00, LPMD01, MW94, MDC^{+}08, OCE^{+}07, RVGG01, SHPA05, SSS99, SC19, TKG97, TA14, VvBv90, Wan09, WTLW23, YE94, YR18, ZLM98]. 4 [ER95b]. ^{+} [Lin94]. ^{2} [CDZ^{+}21]. ^{3} [MMP^{+}21, PB11]. _{//} [C^{+}01a]. _{c} [KSM^{+}94]. A [AND05]. α_{critical} [MOW95]. AX + XB = C [EG94a]. C [WCC^{+}22, BCYB11]. d [Tro00]. d − D [BDK98]. D^{3} [Evr01]. E^{2} [KKŽ05]. E^{3} [KKŽ05]. ℓ_{1} [BE93c]. ϵ [BG97]. f_{1/2} [HC89]. g [MS89]. I [CC98]. ijk [Ort88, OR88]. k [BY21, BDSd94, Bro88a, CCP98, Kon02, LS97, LS92, PER17, She92, TT00, YTCS95]. kd [WTLW23]. L_{1}(L_{∞}) [Jeo91b]. L_{i} [RSSS94]. LU [BETR17, CHQORS18b, TOO20, DHW97, BFR93]. m [CGMM99, Zho93]. μ [SM13]. N [Gor97, LBWR90, Lip99b, Bro88a, CCP98, GS04, GKS98, HWYL89, YTCS95, SSB^{+}91]. O(1) [Wan09]. O(2^{n/2}) [Che08]. P [EW93, GLMBMPMV03, Koç97]. π [Tak18, Tak10]. QR [AD14, AHW14, BTLK18, CTW14, KKW14, TOO20, AB95, BHM^{+}95, CG89b, GK92, Hen93, KC93b, Mat95, OW90, OHZ98, SOH94, Ste87, SNO99, WE94, Wri91, Zhu93]. s [CG89a, CS96]. S_{n} [ARW93]. \sqrt{N} [HK90b]. τ [RGDM15]. W Z [YE95, KS87].

-adic [Koç97]. -algorithm [Ste87]. -ary [Bro88a, CCP98, TT00, YTCS95]. -bit [Ume01]. -body [GKS98, LBWR90, Lip99b]. -Cholesky [TOO20]. -communication [MS99b]. -Cube [GS04, Bro88a, HWYL89]. -cubes [CCP98, YTCS95]. -cyclic [EW93]. -D [Cap88, EK98, SSS99, Tak03, BSE88, Bri95, CDM06, DZD01, HSS07, HB84, KTN^{+}14, LCLL00, MW94, MM96b, YL86, ZLM98]. -decomposition [AHW14]. -dimensional [LS92]. -factorisation [KS87]. -factorization [OHZ98]. -graphs [Gor97]. -inversion [MS89]. -Lattice [GAW96]. -Lop [RGDM15, WCC^{+}22]. -Machine [Evr01]. -means [BY21, BCYB11]. -median [GLMBMPMV03]. -meshes [Tro00]. -order [AND05]. -parallel [Kat03]. -processor [BDSd94]. -relaxation [BG97]. -selection [She92]. -SMP [PB11]. -square [TT00]. -step [CG89a, CS96, CGMM99]. -steps [Mar92]. -towards [SSL03]. -tree [WTLW23]. -way [EY86, PER17, Zho93].

/hypersonic [SOS97].

0 [TYKA95]. ’02 [AGPS03].

1 [ADE84, BNK15, RSRM96, Tem88, Tem89, WSL88]. 10 [LLVM21a, MRJ89]. 10-P [MRJ89]. 100/200 [TKI85]. 100k [SC19]. 164 [Hoc89]. 1991 [BH92]. 1M [HL88].

2 [BHM94, BRH90, But92, CYXL18, DCG90, DD87, FSY88, Hoc94, KN88, MSS^{+}05]. 2-CPU [Hoc85b]. 2.0 [LC17, LT09]. 2.0para [LT09]. 200 [HL88, TKI85]. 2000 [NU05]. 2006 [MTW07]. 2007 [KB08]. 2012 [SN14]. 2016 [AGV^{+}17, Ano18k, DLO17, MSS19, VBC19]. 2017 [FYH^{+}18, GT19, LCE^{+}18, LLX^{+}18, LHL18]. 2019 [Ano19a, Ano19l, TH20]. 2020 [Ano20a, Ano20b, Ano20c, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano21n, Ano22m]. 2021 [Ano21a, Ano21b, Ano21o, Ano21l, Ano21m, Ano21p, Ano21q, Ano21r]. 2022 [Ano22a, Ano22h, Ano22i, Ano22j, Ano22k]. 2023 [Ano23e, Ano23f, Ano23g, Ano23i]. 2024 [Ano24c, Ano24d]. 205 [LN87, Riz85, Won88]. 25-processor [FBMV88]. 256 [BBDN21]. 26th [Ano16k]. 2D [HM97, MKK03, DT95, GJS93b, LSS88, RSSS94]. 2nd [Ano88a]. 2S [SV94].

3 [FCS^{+}19, HCH^{+}96]. 3090

[Car88a, CK90, GSZ88]. 3090/VF [CK90]. 3090E [Web90].

4 [NU05, OKSY92, SKY93, SKS^{+}95]. 4.0 [BPJ22, DWS^{+}21]. 4.5 [DFP^{+}19]. 40 [WSL88]. 48 [Meu87, Nag88].

5 [BL94a, SNS^{+}97, TM94]. 500 [Heg96]. 5E [BEK95b].

6000 [AL93b, NMW93]. 6240 [KW90].

6400 [Dun91]. 65064a [HE88].

76 [HL88]. 77 [Sne88].

8 [WSL88]. 860 [Dun91, GKV94, HC92a]. 88j [HE88]. 8th [ABH18].

90 [CCL04]. 9076 [Bri95]. 93 [Ano94].

AADL [DD89]. ab-initio} [SKE^{+}22]. ABCLib DRSSED [KKHY06a]. ABCLibScript [KKHY06b]. abstractions [CCS03]. accelerate [LXD^{+}23]. Accelerated [LHZ^{+}20, ADF93, BBB^{+}22, BGWR21, BWV^{+}17, CYDJ21, DN16, FKM^{+}22, JNWJ18, JNC^{+}19, MKJC21a, MKJC21b, MAA^{+}21, NDN20, PTGF20, SRK^{+}21, SSV^{+}16, ŚDJ^{+}22, TDB10, TLC23, ZAA^{+}21, CTZ^{+}18]. accelerates [HBC19]. Accelerating [ADEQO19, CZJS12, CFW^{+}22, GTD18, GB14, PVBR23, RSD16, SGP22, TND10, XLS^{+}17, DBI^{+}17]. Acceleration [SWW99, YMJT10, BOV09, CR23, DZLK20, GPW^{+}08, KS16, LPNV20, LTS^{+}24, OYS08, WWZ^{+}18, XLY^{+}20]. accelerator [BNS^{+}07, HHGA15, LXL^{+}22, NSF^{+}22, WYBdJ^{+}21, YLW^{+}13]. accelerator-based [BNS^{+}07]. accelerators [ABB^{+}16, cFM07, KP12, Luc01, WSR14]. access [BDS88, CZJS12, CDK^{+}03, EHHS89, FD18, Sch91b, SLH^{+}18, YZWcF14]. accesses [BS96, TGL02]. accidents [TAB^{+}19]. accounting [YX07]. Accuracy [MAM^{+}09, AÖ22, Coo19, DW00, Sun95, Zha91]. Accurate [AFD13, WCC^{+}22, Moo04, YORO08, ZEC^{+}17]. AcHEe [KKK^{+}18]. achievable [HMS^{+}19]. achieve [SHN12]. achieved [NHS^{+}95]. achievements [Gil94a]. Achieving [Taf01, WYBdJ^{+}21, CHLO85, NW96]. Acknowledgement [Ano08, Ano09a, Ano18l]. Acknowledgment [Ano10a, Ano23h]. ACM [Ano15a]. across [MD04]. action [WL13a]. Active [AKSS07, AGL08, CB17, CC00c, HK98, PCN10, WS98, TM94]. actor [SSBT19]. actual [LAS90]. Ad [MFGEL17, CWK09, SSS^{+}05]. Ad-hoc [MFGEL17]. Ada [BKL^{+}19, CCS87]. adaptable [PAD02]. adaptation [OBG00, WRS12]. Adapting [ADPV03, CS89a, KLP11, MRSB94]. Adaptive [AKK99, CZJS12, CTW14, CD00, GP96, HM97, Las02, LFL11, MSE07, MB00, NC97, PSS01, RE98, RC15, YTCS95, ADK22, BN01, BCCS07, BE88, BLCR21a, BLCR21b, Bur90, CS89c, DLM97, DLR94, DR95, DPSW00, DG05, GCH21, GV06, GZ99, HM89a, HSÇ12, HAK^{+}21, HM86, IDS16, Kd88, KU93, LLP00, LDK16, LR16, LOKM99, MQ89, MT00, MC09, Mic90, OIK21, QCC02, SM13, SFB^{+}97, THG98, TMCC02, TGE92, TCL92, THK14, TSJB00, WZS^{+}14, WY11, XCR17, MD04]. Adaptively [YLT^{+}23]. adaptivity [BFVRC14]. addition [VTmL03]. Additive [AO89, EG94a, LG09, PHD16]. Address [NZHY11, MS20, PEO16, SVC07, SSOB03, Vio04]. Addressing [RG92, Pan01]. ADI [AIIV98, RSK99, Wak04]. adic [Koç97]. adjoint [TW19]. adjust [BP97]. Adjusting [SWR^{+}13, WC94]. adjustment [da 90]. Admission [GE11]. Adoptability [SC92]. Advanced [DDdS02, DDdSL02, HHGA15, KBGZ88, DSCP88, SXBD97]. Advances [MS23, NLG99, AC00]. aerodynamic [ND17]. aerodynamics [AKK16]. aerospace [Man01a, MZ01, ND17]. AES [FLYL16, FLYL21]. affine [BBC^{+}11, BPK12, DR96, LL98]. affinity [CJLS14]. after [KC93b]. AGE [EC91, Eva91]. Agent [SA18, Y^{+}02, CD98, MCB05]. agent-based [CD98, MCB05]. agents [BGP^{+}97, LWG06]. Aggregate [LEH14]. aggregation [Beb97, CGH^{+}19, SHN12]. aggregation/disaggregation [Beb97]. aging [BBDN21]. agreement [CC93]. ahead [CHQORS18b, TRSC^{+}19]. aid [BDS88]. aided [IJCL96]. AIOC [CDZ^{+}21]. air [BWZ95, DM97, OZ02, Zla88, LPNV20]. aircraft [A^{+}01b, Ger94]. airplane [SYAU07]. Aitken [Eva89, LG09, MBC92]. Aitken-additive [LG09]. ALC [YA24]. ALC-PSO [YA24]. algebra [BLRR02, BLKD09, CHQORS18a, CWB92, CIO^{+}17, CGG04, DGP^{+}16, EL95, ES06, HH89, LN87, NE01, ORM^{+}10, PWM00, RS88, TKC^{+}14, TDB10]. Algebraic [CMP92, AB16, BR90, BBQOQO00, DQRR00, KS01, MW88, PHD16, Slo91]. Algorithm [CZJS12, ER18, MWH95, Sen91, YFK03, Ale94, Als01, ARW93, AM93, AB95, AT98, AL88, Aub11, BSH88, BBB^{+}14, BAMK07, BFY06, BKS06, BE93a, Bar97, BW92, Bas94, BOV02, BO03, BOVG10, BE87, BY21, BT99, BEK^{+}11, BE88, BDK95, BC91, BWW89, BF06, Blo03, BSD11, BE93c, BL93a, Bon91, BGM03, BCYB11, Bur04, CDRV97, CGG01, Cha99, CF93, CCS94, CC02b, CC02a, Che08, CJ92, CC95, CLSM98, CST02, CH92, CG92, CSW01, CFH89, DVP90, De 95, DEH^{+}11, Don05, DUG^{+}06, DZD01, EK92, EHE92, ES89a, EMV^{+}18, EPMPU02, EY86, EB88, Eva90, ED91, EY94, FLW86, FF21, Far96, FHL87, FLYL21, FJPA17, FIMF99, GKT^{+}15, GS89, GWLS05, GZHX17, GCH21, GLG06, GV14, Gir02, GMMT16, GKAS91, GÁVRRL18, GH98, GT90, GM89, GM03, GE93, HJ05, HMTX93]. algorithm [Han98, HJ97b, HP94, HS89, HT94, HM89c, HSSM17, HHH92, HC94, HB99, HP91, HM97, HGM20, Hur93, Hwa04, HSN89, IB01, IOH05, IT08, JLY18, Jeo91a, Jou89, KGS08, Kal92, Kan97, KK03, Ken99, KJA15, KL90a, KY94, KB20, KTN^{+}14, KJA05, KN09, KCN99, KŠR04a, KŠR04b, Kra84, KKSS90, KS91b, KP93, KK94, KKB92, KSS06, Lec89, LSA^{+}95, LE91, LM00, LMG09, LW11, LLW^{+}15, LLD19, LHG^{+}23, LO92, LW15, LC97, MS98, MG95, ME84, ME93b, MM00, MP11, MMT07, MMS90, Miš98, MHH97, MR85, Möl99, Moo04, MSB91, MKC92, MA22, NSH^{+}21, NTHY22, NM01, Nat90, NW96, NOS92, NSS15, O’L87, ORM^{+}10, OV06, OW91, OWZ91, ÖS94, OHZ98, OW97, PG93, Par91, PW84, PYLE21a, PYLE21b, PYZC11, Pin91, Pis92, PGBF^{+}07, PS06, PKR00, QJ06, Ram07]. algorithm [RW96, RH99, RE92, RMACG10, SOH94, SMK91, SL03, SM91, SK87, She95, SKC90, SLY90, ST21, Shy90, SML^{+}14, ST95, SGP22, Son94, SE93, SS87, Ste87, SMM90, SM88, Str08, SNO99, SW00, Sun95, SB92b, Tak03, TKK^{+}05, TTO^{+}22, TSCS14, Tem88, Ter13, TY91a, TCL92, TJJ93, TL94, VR95, VvBv90, VIHD90, WP88, Wak04, WWR05, Wan08, WZL09, Wan09, WL13b, WE92, WWT01, WLZ^{+}23, WTLW23, WSR14, Yal97, YORO08, YA24, YLE95, YHZ21, ZM88a, Zer90, ZK92, ZMMW90, Zha91, ZHC^{+}23, ZG16, ZG88, ZM88b, Zub90, dV94, dGR95, JTS^{+}11, Ter10]. Algorithm-level [CZJS12]. Algorithmic [DA06, ME89, AFPG12, DL04]. Algorithms [AGPS03, ASSS11, AGKS15, AGKS16, BDG^{+}10, BFR93, CGNR06, FC18, FYEHP89, HFL^{+}10, KTAB^{+}19, MAH^{+}19, PMMAM10, Tap84a, ABB^{+}21, AAGS18, AÖ22, ALNT04, AAB^{+}06, AD14, Arb92, AJ89, AV15, AU88, AGGG06, AB16, BM08, BCVC05, BO91, BHK89, BS97, BEDdC16, BP95, BE93b, BGSS14, BAR98, BGMT00, BD18, BA95, BWL00, Bog92, BSB^{+}22, BFVRC14, BTLK18, BDSV98, Bro89, BLKD09, CMT04a, Cal96, Car88a, ÇFG^{+}12, CZ91, CDBL08, CJBK93, CCJ90, CC91, CL92, CL02, CDW95, C^{+}01b, CGM^{+}92, CR90, CKM93, DLM97, DDP90a, DDP90b, DY91, DN16, DM04, DQRR00, DSS86, DFH^{+}13, DAD11, DR02, Eck90, Ekl04, EAR93, El 93, ESV10, ES88, ER95a, ER95b, EG92a, EM85, EM88, EA95, Fie96, FOH87, FJS85, FB91, GK03, Gen84, GV11, GJMM89, GAR15, GGL21, GAMR03, Gor97]. algorithms [GM07, GPS^{+}08, GK04, HKSK97, HA05, HCJ03, HM89b, HW91, HI92, HCK94, Hot89, HWW92, HK90a, HK90b, HC92b, HK91, IR93, IKK15, JW09, JG93, JWS13, JCC^{+}24, Jor86, KB85, KN88, KH95, Kd88, KKW14, KKU16, Kas85, KBBC88, Kat86, KLP11, KH92, Kim90, KC91, Kim98, Kon02, KST02, KU93, KKB93, LSS05, Lah00, LD99, LPM11, LAHM14, LNS03, LR04, LF89, LBWR90, LFL11, Lin91b, Lin01, Lip99b, MKS91, MR89, McB88b, MF16, MCG98, MNS87, MRSB94, MS90, MOF04, MRJ89, MGSK88, MWR95, Mun99, MB00, MO99, NA08, NZ92, Num05, OW90, Obe85, OIK21, Ols95, OW94, PS89a, PCL23, Pet84, PVP08, PD11, PB94, Qui88, RR89, Reg01, Rib84, RV96, Rod85, RRS^{+}00, Ron84, RS87, RR14, SN88, SS02, SSY02]. algorithms [San99, SST09, SMT^{+}04, SB96, SAGV21, SB94b, SAWH88, Sha06, SKC90, SLG^{+}22, SC92, Sim91, SPMB23, SPK18, SSN04, SE87, Sto89, Stp93, Str87, Swa84, SWC99, TS02, TTH95, TNZM20, TPK^{+}13, TOO20, Thu92, TSJB00, TCT00, TCS04, Trä95, Trä18, TL96, TO89, Tyr90, VP95, VTmL03, VB92, VL05, WC00, WRB97, Wat00, Wen95, WC90, WSB90, Wri91, WS94, WWL^{+}22, YK06, YK08, dCC89, tV96, ML00]. algorithms-description [CR90]. alignment [DEH^{+}11, OGC^{+}15, Ste13, TS09, ZHC^{+}23]. all-FPGA [SDMS12]. All-pairs [SA13]. all-reduce [PGW16]. all-row [GYH94]. Alleviation [FTY^{+}20]. Allgather [KTAB^{+}19]. Allgatherv [KTAB^{+}19]. Alliant [WSL88]. allocation [AL93a, BEDdC16, BCRSR11, CLZ01, DHL07, EEH^{+}19, EG90, GRP22, GPC88, GV08, HM02, HMH^{+}13, JSC97, KSP97, LLL10, NL98, SS94a, SSS^{+}09, SYY^{+}22, ZYZ^{+}16]. allocator [IYV^{+}22]. allocators [RDS13]. allowing [CRGR^{+}13]. allreduce [JOL^{+}21, RBB17]. allscale [JGT^{+}20]. almost [PG91]. alpha [HSN89]. alpha/beta [HSN89]. Alphaserver [WBS06]. Alternate [HW94, DWX^{+}12]. alternating [BE92, EC91, Mei86, ZL94]. Alternative [GGFF93, Trä12, CGMS94, KJA05]. Alternatives [Rot95]. ambulance [GLS01]. AMC [HHGA15]. AMD [BBL^{+}19]. Amdahl [MOW96, YMG14]. AMG [BDP20, Ema10, MX07]. AMG-preconditioners [Ema10]. among [CHLO85]. AMR [WAT20]. AMT [Hil92]. analog [PB98]. analyses [CI98, FA11]. Analysis [ADGS10, DZM^{+}13, DFP^{+}19, EO91, FLYL21, HSSM17, KD97, LOKM99, MRT93, PCA01, San99, SNK06, SDv98, VJ12, van87a, AGL08, ALD01, AZ98, AKSS07, AKK99, ARR19, AW13, BM08, BP93, Bjø87, BFH23, BOS^{+}09, BDS88, Bro86, BOG15, CMT04b, CC96b, CMPM^{+}15, CC92, DWB99, DLPS92, FHK^{+}15, Feo88, FGHB94, GLP08, GP90, Gir02, GKAS91, GLP98, GT07, GT90, GZTR07, GMS04, GT92, HMTX93, HKW13, HW91, Hil92, HSFS14, IR00, IHM^{+}12, JAA06, JPK^{+}15, JCE96, Kat01, KCG08, LR04, LyHW^{+}22, LSA^{+}95, LR99, LWZL24, LGD^{+}15, Lin08, Liu98, MSE95, NMAB19, NS02, NA08, NL98, NMI^{+}24, OAJ^{+}16, PSW18, Pis92, PVP08, Psa02, QCC02, RRS^{+}00, SNS^{+}97, SBM^{+}98, SHN12, SHCS86, SFB^{+}97, SMJ15, SV97, Stp93, TPK^{+}14, TFS91, TA14, WE97, WOKH96, YNO97, YLCT18]. analysis [ZAA^{+}21]. analytic [GSC^{+}22, LND^{+}19, MWR95]. Analytical [TDW^{+}04, GPZ08, JMLBL98]. Analytics [NOG^{+}22, CC16, CLJ14, IYV^{+}22, KMB^{+}18, KG21, XLS^{+}17]. Analyzing [BEDdC16, CGHBS18, Coo19, HHOMR06, HG12, LGC97, OKF16, RDS13, YS96, YMYH24, AKP^{+}98, MM96b]. ancestor [LO92]. AND/OR [VKH99]. aneurysms [PIG^{+}16]. Angela [Hol11]. angular [FSY88, GICM18]. animation [CDV08]. anisotropic [MSV15, Oli96, PMLT03]. ANN [AFT23, KK94]. annealing [AC89, BL93a, Dod90, DW00, JK91, LR16, MG95, WWR05, YLE95]. annihilation [Meg90a]. annotations [AAA16]. anomaly [Lin91a]. ant [KŠR04a, KŠR04b, SSBT19]. ant-colony [KŠR04a, KŠR04b]. anti [CDRV97]. APE [Rap99]. APE100 [GLS99]. APEmille [Tri99]. aperture [FFLV96]. APGAS [HXCZ14]. API [IMQO^{+}18]. apid [FD02]. Application [EMV^{+}18, Hem94, KP12, LMH08, PY00, PMAL14, Sun95, WZV^{+}16, AAO13, ADMV05, ADR^{+}05, ADE84, BM93, BBB^{+}22, BMS01a, BOS^{+}09, CB17, CLC08, CRGM16, DUG^{+}06, DL06, Gao87, GPW^{+}08, GIRT19, KNTG08, LFS^{+}19, LLW^{+}24, LHZ^{+}22, LV92, LF88, MAJD17, MAM^{+}09, NMW93, ORS90, OCSBY01, OILZ17, OCPT97, PAD02, SSS99, SFSV13, SF91, STT94, TB03, VR95, VCK^{+}11, WBPM97, WBN^{+}17, YSBM97, MPZS13]. Application-aware [PMAL14]. application-centric [SFSV13]. application-level [CRGM16]. Application-specific [MPZS13]. Applications [ABH18, AGPS03, ASSS11, AGKS15, AGKS16, BDG^{+}10, CGNR06, DMWW88, FG01, OSZ93, PS20, Sod02, TFV16, AAB^{+}16, AAGS18, AQ93, AK89b, ANG^{+}20, AFD13, AMMV98, BDPV99, BH14, BMS01b, BFL^{+}01, BRWL09, BDNP11, BCRSR11, BOV09, BFVRC14, BTLK18, BCC^{+}97a, BCC^{+}97b, BK07, CMT04b, CTS02, CMSL06, CP18, CAA98, CHR00a, CSB00, CMPM^{+}15, CL92, DDdSL02, DCD97, Dec00, DPFT19, EHR^{+}98, Emm84, DST15, FD18, FGBN19, FAS02, LHR^{+}19, FJPA17, FBAB13, FA96, FMM^{+}02, GRP22, GWWM09, GJMM18, GW08, Gra91, GPS^{+}08, GAPZ00, HL97, HA05, HCT16, HCH^{+}96, HAJK01, HST87, HKN89, HTB01, IMQO^{+}18, JNWJ18, JNC^{+}19, Kan97, KST02, KKSS90, kLH95, Li89, LTV96, LWW^{+}21, MKK^{+}19, MD04, MKL^{+}01, MCB16, MSMC15a, MZ01, MHL06, MCG^{+}12, MW88, NAC^{+}14, NS02, ND17]. applications [NSM23, OK22, OID^{+}12, PK05, PSS94, PGK^{+}18, PMV^{+}20, PRR03, RDS13, R^{+}00, RGGP^{+}18, RHM^{+}88, SKST08, SAGV21, SDMS12, SHN12, SHB19, SMJ15, Sol88, ST88, SC05, SR10, SRS^{+}19, TS02, TDG^{+}18, TB22, TTH09, Tor91, Tou02, TF98, THK^{+}99, Van02, VV07, WH97, WYT^{+}16, WAT20, WLN^{+}96, WOKH96, YXQ^{+}21, YMYH24, ZAA^{+}21, ZEgT04, dSS09, dNdRRL^{+}21]. applicative [OM90]. Applied [Ano87, BCG00, BP02, Dod89, GG10, HWB92, ML00, RGBC11]. Applying [LQ92, SMC16, Bur04]. approach [ADGS10, Bek95a, BC97, BR99, BCS15, CCG^{+}17, CDZ^{+}21, CA92, CRGM16, CDM^{+}10, DF91, DFRZ02, DL05, EGTD99, FHK^{+}11, FM15, GO88, GK19, HAK^{+}21, JMLBL98, KPL^{+}12, Kot97, LCLL00, LXD^{+}23, MGSK87, MWR95, NA03, NC97, PAF^{+}97, PB98, SHPA05, SB94a, SLC^{+}16, THG98, TRSHB04, VV07, VBS^{+}15, VFG12, Vég18, WA03, WJK98, Web90, Wol89, YA24, YWP11, ZLJ93, ZGG92, dNdRRL^{+}21]. Approaches [Kut02, AKB^{+}19, BN01, DCY^{+}22, GJMMS97, GLP98, JCC^{+}24, Lau93, SER90, Tyr90]. Approaching [MSOCG^{+}16]. Approximate [AHBD18, CC93, Daḡ07, GCH21, Gur88, HRR08a, KKK^{+}18, Mil01, TLS16]. approximately [SSH97]. approximating [RH13]. Approximation [Lah00, Mun99, AÖ22, BG91a, BE93c, CLHL23, Cor00, HGM20, KS91b, MKC92]. approximations [GGL21, SSL19]. April [Hol11, WBD99a, Ano20a, Ano21a]. AR-modified [GGFF93]. arallel [FD02]. arbitrarily [BPK12, Wag89]. Arbitrary [DZ90, LPNV20, MJ94, RV96, SSB^{+}91]. arc [CSEK03, KH95]. Architectural [VJ11, Web90, MSE19]. Architecture [Ano16k, ALTZ02, CGG04, DZ90, Ger04, KR10, KU93, LF00, Wat87, AB02, AK89b, CLC^{+}18, Cho92, DFR92, DWH^{+}08, DEH^{+}11, DSSD18, DAG^{+}09, DBI^{+}17, EBSS94, Ekl04, Evr01, FFLV96, AKK16, FYH^{+}18, GP86, GWWM09, HISS92, IM88, JHSV02, JGT^{+}20, Jou89, KSS06, KAD09, LNLK13, LLW^{+}24, Lil94, LHL18, LWJ^{+}17, MO89, MMSS22, MF16, MP95, NFG^{+}13, QH19, RGT17, SKS^{+}95, SKG02, SCD92, Sod02, SSGF00, Tre85, Tre88, WMCU97, WE93, WE94, WRS12, Y^{+}99, HZY^{+}19, BKS19, BKL^{+}19, BBL^{+}19, LYC^{+}19, WHL19]. Architecture-independent [LF00].

Architectures [BWL00, HFL^{+}10, SYP13, ARCH05, ACMT20, Ale91, ABB^{+}16, ADPV03, BJ18, BZ99, Bjø87, Bog92, BCS15, BCMSW03, BCMSW04, BLKD09, CFF17, CHQORS18a, ÇFG^{+}12, CDZ^{+}98, CCJ90, CLSM98, CS19, CS89b, CG92, CGH^{+}19, CDGI15, CRT89, CIO^{+}17, DD89, DR94, DFO93, DEGS95, DTR18, DSCP88, DvdG92, Erc88, FLS^{+}21, FD18, FR95, Gao90, Gil94b, HHK95, KK11, KZV97, Kut02, Kuz98, LD99, LLX^{+}18, LGD^{+}15, LTV96, LBGO23, MS00, MSE19, Mat95, McB88b, MD88, NS11, NZ92, ORM^{+}10, PMS^{+}13, PS89a, PYLE21a, PYLE21b, Pea19, PB11, RP85, SS89a, SSB12, SKH^{+}12, Sel95, SS04, Sol88, SWR^{+}18, TB07, TS21, Uhl96, VKS^{+}15, WYBdJ^{+}21, WSR14, YMJT10, ZMMW90, vCd90]. architectures/algorithms [LD99]. Area [RKS92, FGG^{+}98, FCM03, KBG^{+}01, MSM98, NFG^{+}13, SB97]. area- [NFG^{+}13]. Argonne [BRH90]. Argonne/GMD [BRH90]. argument [GF13]. argument-based [GF13]. Argus [ZP16]. ARIAL [FD02]. arising [GKT^{+}15]. arithmetic [DS93, Sch87, SS87, Tak10]. ARM [ZJD^{+}23]. Arnoldi [HRT07, IS18]. Array [Lin91a, ARB94, ARB95, Bek95a, BR89, Bol86, CCCP92, CCL04, CC99b, CMM^{+}88b, EB88, Eva92, FF92, FYEHP89, GS06, GNY00, GE92, Hil92, Kak88, Kal92, KS07, KLS^{+}88, LC02, LL90, Lin91b, McC89, PS89b, PRW88, Ris90, SS91a, SB95, Son92, TTH95, Tor91, WW95, WE93, WE94, YL86, Num05]. array-based [McC89]. array-processor [Str87]. Arrays [SA18, ABGC^{+}14, ACC^{+}88, BE93a, Bar95, BR90, BP90, BK97, CS89a, CJBK93, CC91, CMP92, DY91, DS93, DS92, EM86, EM87, EG92b, EG94b, FMB98, GE93, HI92, HCL05, JWS13, KL00, KC93a, Lin90, Lin91a, Lin94, LT90, MM96a, ME89, ME90c, MBK12, MS89, ONOK13, RF90, SOS97, SSH97, TH89, WB88, WWL^{+}22, ZLJ93]. Arria [LLVM21b, LLVM21a]. art [LR88, SKH^{+}12]. ARTE [MPZS13]. articulation [KH95]. Artificial [Loh96, BJSN04, CH92, Fah85, Mar95, Sin90]. ary [Bro88a, CCP98, TT00, YTCS95]. ASA [XCR17]. ASA-FTL [XCR17]. Asian [SÖB07]. ASIC [vCd90]. Aspects [GWC^{+}99, BWZ95, CW97, CEM^{+}99, DL04, DNL15, EV89, Hol95, JG93, KPS90, LBL95, kLH95, LTV96, Par87b, Ron84, STKA96, WRB97]. assembly [EL16, SPMB23]. Assessing [HBH^{+}16]. Assessment [KR10, MT95, PTK04]. asset [CDM^{+}10]. asset-liability [CDM^{+}10]. assets [ZCPT00]. Assigning [MT91]. Assignment [DN16, LW15, BTZ06, FYEHP89, GHM97, KY98, LC02, LX00, LZCT15, MT95, Man01b, MT89, ME90b, NDN20, SS91b, Tai91, TN13, WS94]. assimilation [Ran97, TL96]. assist [PB16, Van02, VV07]. assisted [ADK22, NS11]. associated [SSD01]. association [VA14]. Associative [SA18, ARZ97, DM90, DS92, MP92, Mar95, PPZ93]. assumed [BFY06]. astronomical [WYX^{+}22, dCC89]. astrophysical [CCP^{+}21, Sug99]. asymmetric [ACH^{+}18, CHQORS18b, GBH98]. asymmetry [CIO^{+}17]. asymptotic [WZCG14]. Async [MGCB^{+}10]. asynchronization [FSG19]. Asynchronous [BÁJG^{+}20, El 93, FB91, HM89a, MQ89, OIK21, TLC^{+}21, WE91a, ATT89, BM93, BCVC05, Bai99, BP93, BP95, BG97, BC91, CZ91, CGM^{+}92, CGOP03, CB09, DCR^{+}16, ED91, GAR15, HP97, HWB92, OILZ17, PHM^{+}22, RRS^{+}00, RSC^{+}19, SB96, SK87, Son94, ÜD89, Wei93, WP19, YCBD19, ZDE95, RHWF23]. ATAPE [PW22]. ATCOR [BRS^{+}08]. Athread [LHZ^{+}22]. ATLAS [WPD01]. ATM [Sab97]. Atmospheric [Adr99, AFG^{+}97, DS96, HLDS95, KNTG08, MSW98, MSV15, PFS^{+}04, PODd16, SB97, WME^{+}95, YNY^{+}17]. Atomic [BCA08, SSV^{+}16, BF92, DS13, LWG06]. atoms [SKE^{+}22]. attacking [BF06]. attainable [Coo19]. ATTEMPT [TYKA95]. ATTEMPT-0 [TYKA95]. attention [GH98]. attribute [CRGR^{+}13]. Auction [AGGG06, BC91, PB94, SSB12]. auction-based [SSB12]. augmented [HCTH05, MLX07]. August [Hol11, Ano20b]. Aurora [TSEE21]. Austin/Texas [SWR^{+}18]. Author [Ano84a, Ano01, Ano02a]. auto [KKHY06b, KKHY06a, CCGG14]. Auto-tuned [CCGG14]. auto-tuning [KKHY06b, KKHY06a]. AutoDock [SVTSM^{+}22]. Automata [ATT92, ATT89, AvHL^{+}97, BMS01b, BMS01a, BM01a, BMA06, CDR^{+}95, CCEJ01, CFMM97, FMS^{+}06, Hee97, KVW97b, MHK97, M^{+}00, MCM01, OCSBY01, SBZ04, SC19, Sut97, Wei97, Wei01, Yan04, KVW97a]. Automated [CQ97, WPD01, TLC23]. Automatic [AKNS91, EJLC97, EJLC00, Fis90, GAPZ00, JIC96, LF98, MS99a, RW89, VFG12, APBcF16, AvHL^{+}97, AART13, AGK98, CYDJ21, CMT04b, CMSL06, CDGM96, GAMR03, GF13, Has01, HM01, KBGZ88, LMJC96, LCD91, LLFZ23, PB23, PHM^{+}22, SM13, SFB^{+}97, ZBG88]. automatically [CGG04, Glo95a, Her00, MFGEL17]. Automating [Xue94]. automaton [Ume01, WNES01]. automotive [Gin99, vWWM^{+}19]. autonomic [CDZ^{+}21]. Autonomous [DMT06]. Autotasking [EO91, Nag90]. autotuning [BWV^{+}17]. availability [BDF^{+}00]. average [Tap84b]. averaging [CGG01]. avoidance [CZTS99, CSH^{+}11]. Avoiding [WLYJ13, IS18]. AVX512 [AAS20]. Aware [DGKF19, HVA^{+}16, ZWJ^{+}19, AFT23, CDA^{+}16, DCN^{+}15, FCC16, GIRT19, HCT16, HG12, JWS13, KBG^{+}01, LFS14, LCLA19, LBGO23, LWW^{+}21, MSMC15a, NB12, PMAL14, PVK^{+}22, RR16, SYY^{+}22, TDC19, XCR17, YXQ^{+}21, ZP16, ZYZ^{+}16]. awareness [BHR09, PCLM18, SS04, VGS14]. axis [Wan09]. axon [WS04].

back [CS89b, MMC97]. back-propagation [MMC97]. backplane [SP94]. backprojection [RW98]. backpropagation [WZ90, ZMMW90]. Bag [dSS09]. Bag-of-Tasks [dSS09]. Balance [EY90, CHQORS18a, ORM^{+}10]. balanced [BF95, BM01b, CG92, Gor97, GH92, LZ00, SSH97, TRLD13]. Balancing [SG16, BMP15, BCPB05, BD18, BP97, BS05, Bog92, CE94, CS00, CFF17, CG87, DLM97, DR93, DLR94, DR95, Dec00, DG90, DFM99, DPSW00, EB93, EB94, FMD98, FG96, GGV04, GPS03, HL97, Hor93b, HB99, JH97, KNP97, Koh95, MD04, MG19, MMT07, MCG^{+}12, MSE07, MZ95, NC97, OBG00, PRR03, RHWF23, RCAP11, San99, SSLK00, SGS95, Son94, TSJB00, VR95, YEC97, YSBM97, Zam99]. band [CS87, MMS90, TRSC^{+}19, ZVWS88]. banded [Con89, DG95a, DS84, DJ87, EM88, HS90, Lan96a, Mei85, MP11, PS06, RLH19, VB95, ZM94]. bandsymmetric [Arb92]. bandwidth [CD00, LNLK13, ML23, SST09, ZYZ^{+}16]. Banerjee [CC00a]. banks [GGL^{+}05]. banyan [MO89]. Barnes [GKS98]. barrier [AJ89, FK98, Nic95]. barriers [Axe86]. base [DFRC94]. based [AK89a, AKB^{+}19, ALH^{+}14, ARZ97, ACLN03, ALTZ02, AKP^{+}98, AATK10, BS00a, BN01, BBL^{+}16, BM01a, BVP^{+}89, Bar12, BWV^{+}17, BY21, BCB02, BFG^{+}07, BDP20, BKK24, BNS^{+}07, BFH23, BDH95, BÁJG^{+}20, BFVRC14, BI07, BJSN04, BDK06, Bra88, BRS^{+}08, Bri95, CYDJ21, CRGR^{+}13, CDDG93, CCEJ01, CFF17, CAC^{+}09, CD98, CCCP92, Cv98, CL05, CGH^{+}19, CV02, CKS03, DTV21, DCR^{+}16, DP00, DFRZ02, DAG^{+}09, DGP^{+}16, EJLC00, FD18, FLYL16, FHK^{+}11, FLPG18, FA11, FGHB94, FMS^{+}06, Ger98, GMMT16, GF13, GHH^{+}03, GZ99, GLD08, GVH08, GSC^{+}22, GK04, GPZ08, Haa98, HJS23, HSÇ12, HAK^{+}21, HJ97a, Her00, HST05, HM86, Hot89, HF20, IR93, IS18, Jou89, JMA^{+}13, KYS23, KK03, KG21, KPL^{+}12, Kog85, KM89, KS01, LJD93, LAHM14, LSA^{+}95, LS90, Li89, LWZL24, LLL10, LLFZ23]. based [LKP24, LTS^{+}24, LZCT15, Lop93, LPAZ97, LBH07, MKK^{+}19, MP92, MPZS13, MCB05, McC89, MMT07, MAB^{+}21, Miš98, MRSB94, MCG^{+}12, MKL^{+}88, NMMG13, Num05, OAJ^{+}16, OIK21, ÖW10, OGC^{+}15, PB98, PK05, Pel97, PGGL17, POHS14, PHCR05, PGG11, PSF^{+}15, PGW16, RHWF23, RMS^{+}18b, RBS10, RVD10, STP^{+}19, SSSG03, SSB12, SKE^{+}22, SSN^{+}21, Sch92, SSS99, She95, SLG^{+}22, SCL^{+}23, ST21, SML^{+}14, SSS^{+}09, SSN04, SSBT19, SIH14, SW03, Tak06, TW19, TKK^{+}05, TLS^{+}08, TDC19, TNZM20, TSCS14, TB07, TND10, TE92, VP95, VFG12, VRGÁ15, VKS^{+}15, VK92, WWR05, Wan09, WL13b, WCC^{+}22, WAT20, WL14, WYBdJ^{+}21, WWL^{+}22, WTLW23, XXQ^{+}15, XRW^{+}19, YMJT10, Y^{+}99, YTSI13, YTCS97, Y^{+}02, YHZ21, YLT^{+}23, YZB^{+}23, ZLT19, de 96, Lin91b]. Basic [DT95, DLdS06, LN87, HM89a]. basins [WFW^{+}11]. basis [Bak00, GPSK09, KS91b, LSS05, Uhl96, WYBdJ^{+}21]. Basker [BETR17]. batch [BM11]. Batched [BTLK18, ADFQO19]. Batching [JSC97]. Bayesian [LTG23, SYAU07, YMJT10]. BBN [PBTC89]. BDSC [KJA15]. be [FLW87, Par86, PTS^{+}12]. beam [LPCA98, OIH10, SCCPM14, SKH^{+}12]. beamforming [GMF00]. bearing [TB03]. bee [BW01, GÁVRRL18]. BeeGFS [BPC21a, BPC21b]. behavior [AAO13, AFD13, Beb97, CBP^{+}07, MAJD17, TCS04]. Behavioral [FKPB15, FF21]. Behaviour [HK98]. BEM [BCS15]. Bench [NL91]. benchmark [BCL91, BOS^{+}09, GSE^{+}15, KK92, LXC^{+}18, MKL^{+}01, Oed92, PB23, ZCT^{+}20, van91]. Benchmarking [AGH^{+}94, FFZ^{+}18, MHL06, NL91, SVTSM^{+}22, Gin99, HL88, Hoc91]. Benchmarks [Xu07, But92, Dix91, Hey91, Wei91]. Benders [NZ97]. benefit [PGK^{+}18]. Beowulf [DWM^{+}01]. Bernoulli [ME90a]. Bernstein [MSE93]. Best [Ano15a, SN14, Ano18k, GT19, TO89, Uhl96]. best-bound-first [TO89]. beta [God00, HSN89, SHCS86]. beta-networks [SHCS86]. Bethe [PMV^{+}20]. better [SG91]. between [BL93b, BK97, CR23, CHQORS18a, DW99, LKGD16, MSP93]. Beyond [DM90, SC19]. bi [BSD11, MMT06, VP94, BP02]. Bi-CGSTAB [BP02]. bi-diagonalization [VP94]. bi-objective [BSD11, MMT06]. bias [NTHY22]. biased [SR97, Zam99]. BiCG [ADEQO19]. BiCGStab [Coo19, CV17]. bidiagonal [GTD18, GL99, Kuz98, Lan96a, LPNJR96, TND10, VP94, Yal97]. bidirectional [MP92]. Big [LK14, STE23, BHB21, PFTB15, RBB^{+}22, XLS^{+}17]. big-data [PFTB15]. biharmonic [LV92, MS04b]. binary [AJ95, AND05, AFS14, Cho92, LH02, SGS95, Tou02, Ume01]. bindings [VGRS16]. binomial [DP00, Ger04]. bio [GHC04, HR04, RD07, RBS10]. bio-computing [GHC04, HR04]. bio-inspired [RBS10]. bio-molecules [RD07]. bioinformatic [JBC^{+}04]. bioinformatics [DEH^{+}11, OGC^{+}15, SKST08, SWYM17, VRGÁ15]. Bioinspired [HFL^{+}10]. biological [ORS90, ZHC^{+}23]. biologically [HSR^{+}14]. biology [BA08]. biomedical [Rob99, SHCR98, TRSHB04]. Biomolecular [BWT^{+}08, AAV08]. biomolecules [EMV^{+}18, SSS92]. bipartite [AB16, LPM11, LAHM14, SKC90]. bipartitioning [KB20]. bisection [VR95]. Bit [TLS16, PS89b, SAOKM01, SR97, SWSG92, Ume01]. bit-level [PS89b]. Bit-parallel [TLS16]. bit-reversal [SAOKM01]. black [BP02, MCY^{+}24, SS88, SY21]. black-box [MCY^{+}24]. BLAS [DMWW88, Lip99b, LN87, Phi91]. BLAST [HMS^{+}07, YZWcF14]. BLAZE [MV87b]. Block [DZLK20, ER95a, IS18, ME93b, SY21, ADFQO19, BOV02, BO03, BOVG10, BDK95, BF06, BVWH14, BP02, BM90, CTW14, DR18, DNL15, DAD11, DS05, DR02, EY90, Fie96, GS89, HS90, HAK^{+}21, HRR08a, Hwa04, KGS08, LC02, LKHL03, Li97, LHG^{+}23, MW11, Meh93, MAD^{+}16, O’L87, OV06, PG91, RM97, RR16, Sch91b, SZ02, SE98, Ter13, WB88, WE94, Wan09, WAT20, van87b]. block-algorithms [Fie96]. block-based [Wan09, WAT20]. block-compressed [BVWH14]. block-cyclic [Hwa04, LC02]. block-diagonal [KGS08]. block-ILU [HRR08a]. block-Jacobi [BOVG10, OV06, ADFQO19, BOV02, BO03]. block-level [MAD^{+}16]. block-oriented [LKHL03]. block-preconditioners [DNL15]. block-sparse [RR16]. block-structured [HAK^{+}21]. block-tridiagonal [DR18, SE98, Ter13]. blocked [AHW14]. Blocking [NSS15, BO03, BCS15, HGLR07, KS01, PGGL17, PHCR05, STP^{+}19]. blocks [DBZ^{+}19, HRR08b, Mar14]. blockwise [Bai99]. blood [GGFF93]. Blue [WZV^{+}16, BJV^{+}16, MV17]. Bluestein [SSB^{+}91]. Board [Ano18e, Ano18f, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano06, Ano09b, Ano10b, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano18a]. Board [Ano18b, Ano18c, Ano18d, Ano18g, Ano18h, Ano18i, Ano18j, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano22g, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23a, Ano23b, Ano23c, Ano23d, Ano24a, Ano24b]. Body [ER18, LCE^{+}18, CLC^{+}18, CDV08, DSSD18, FYH^{+}18, GKS98, LLX^{+}18, LBWR90, Lip99b, LHL18, MBB^{+}18, Sug99, SWR^{+}18]. Boeing [FMS01, MKL^{+}01]. boid [LKP24]. Boltzmann [AK89a, CGK^{+}16, DF91, DFR92, DFO93, FHK^{+}11, FHK^{+}15, GPLW17, HAK^{+}21, HWB92, OKTR13, Sch92, SP93, XT11]. BondMachine [MMSS22]. bone [TA14]. Boolean [HWYL89, Sen91, WP88]. Boosting [MC09]. bootstrapping [GAA^{+}13]. bordered [DS05, VP94]. bordering [Eva92, LE91]. both [Hsi06, LF00]. bottleneck [PJW^{+}22, RG92]. bottlenecks [DSG17, HC89, JKHK08]. Bound [GMMT16, BCG00, CC92, Chr00b, DM03, DCY^{+}22, DFRC90, DUG^{+}06, FF21, GK92, MD04, MT90, MMT07, SDMS12, TYLL22, TO89, ADMV05, Lau93]. bound-constrained [DM03]. boundary [DL92, Dav88, Gol88, GW87, Kie91, Lop93, MPJ03, PG93, Pir93, Pir96, RW96]. Bounding [CRD02]. bounds [LMKH97]. box [MCY^{+}24, SS88]. BPC [LY93]. BPC-permutations [LY93]. BPMN [STE23]. brain [PJW^{+}22, VLSPL19]. Branch [BCG00, GMMT16, DCY^{+}22, DFRC90, DUG^{+}06, FF21, MD04, MMT07, Ral03, TO89, ADMV05, Lau93]. Branch-and-Bound [GMMT16, BCG00, DCY^{+}22, DUG^{+}06, MD04, TO89]. branching [PCL23]. bricked [MSE07]. bridges [HHH92, SS89b, ZM88b]. bridging [NMMG13]. Bringing [Col04]. Broadband [BAMK07, BAK09, KR10, PMMAM10, SKST08, CD00]. broadcast [AMC^{+}19, BT99, GT92, HT92, HK90a, Jia09, LS98, SST09, Sim91, THK^{+}99]. Broadcasting [BMT92, CL92, CHH^{+}01, GP96, ST02, ZM96]. broadening [HK91]. Broadwell [DSSD18, FYH^{+}18, LCE^{+}18, LHL18]. Brokering [BG07]. Bruijn [FJPA17, Bar97]. BSC [SLH^{+}18]. BSP [BJvOR03, BI07, DHS00, GS01, Ger15, HMS^{+}98, IB01, JHSV02, EB98]. BSPlib [HMS^{+}98]. bubble [MP87]. Buchberger [Sen91]. bucket [JON08]. buffer [YHZ21]. buffered [Lin93]. Buffering [Sor84, BL99, MC09]. Buffers [ZS21]. build [DFRZ02, EL95]. Building [BGWW97, CMT04a, RBB^{+}22, HRR08b]. built [DK01]. bulge [KKW14]. bulge-chasing [KKW14]. Bulk [KS99, PM03, SRS^{+}19]. bulk-loading [PM03]. Burst [ZS21]. Burst-Buffers [ZS21]. bursty [SP97, WL98]. Bus [ZKCL04, BV04, FHKZ88, Kat01, Tak06, THK^{+}99, WH94, VP95]. bus-based [Tak06, VP95]. bus-connected [BV04]. buses [LPH00]. butterfly [Bro87, TTH95, Tou02, PBTC89]. BVP [GH89]. bypass [SP97, WL98]. bytecode [A^{+}01a].

C [ATA^{+}16, Hol11, FSCL06, Son86, WSL88]. C-1 [WSL88]. C2 [SM13]. C90 [Oed92].

CA [BW01, FGS97]. CA-like [FGS97]. CAA [IS18]. Cabana [HJS23]. Cabana-based [HJS23]. Cache [GM07, ZWJ^{+}19, APPG15, BGR00, CD00, DG94, HCT16, LFS14, MSS^{+}05, PSS01, PJW^{+}22, RH13, SIH14, YB11]. cache-aware [HCT16]. Cache-efficient [GM07]. cache-oblivious [YB11]. cached [LBH07]. caches [AFD13, RB14]. caching [AKSS07]. CAD [CAA98]. Calculating [Fre89]. calculation [ALH^{+}14, BWT^{+}08, DL04, FSY88, HWL^{+}22, LCLL00, NHS^{+}95, SMC16, Tak10]. calculations [ABB^{+}11, BF92, Ber00, DD87, EEH^{+}19, FGHB94, GSMY^{+}07, HH91, JNWJ18, JNC^{+}19, PGK^{+}00, RKMJ10, SB94a, Sil88]. calculators [Hee97]. calculus [HL99, VP92]. Calendar [Ano84b, Ano97]. calibration [God00, WWT01]. Callback [SSN^{+}21]. Callback-based [SSN^{+}21]. CAMELOT [DS03]. can [FLW87, Par86]. cancellation [LHS13]. Cannon [ML20]. Cannon-type [ML20]. capability [HISS92]. capacitated [GPS03, Ral03]. capacity [JSC97, MLM^{+}00, ZLPF16]. CAPDYN [Pry97]. CAPTools [IJCL96]. car [CLL99]. carbon [PODd16]. cardinality [AB16]. Carlo [vS92, ACC^{+}88, BSH88, BLFT84, DP95, EEH^{+}19, FHL87, FHJ^{+}84, GPW^{+}08, HA97, HSFS14, Ken99, LM00, mM94, MD88, MAM^{+}09, MB88, ÖW10, ZLM98]. CARP [GG10]. carried [BCR00]. Carrying [SR14]. CARS [LCLA19]. Cartesian [Gro19]. case [AFY^{+}16, AMMV98, AGK98, Ber00, BF06, CDM06, Chr98, CLZ01, GJS93b, GCC19, JLY18, Kas84, LyHW^{+}22, Liu98, LF88, MMT06, MAM^{+}09, MST94, NAC^{+}14, Oli96, RSK99, SEAH^{+}08, SSLK00, SWYM17, Sod02, THMH21, Wal01, WLZ^{+}23]. cases [SMP11]. cationic [STKA96]. causal [PLY02]. cause [SMJ15]. cavity [Ste88]. Cayley [BT99, GHM97, LJD93]. CC [HK02b]. CC-NUMA [HK02b]. CCF [AAS20]. ccNUMA [SSGF00]. CDC [LN87]. CELIP [Has90]. Cell [BMA06, BL94a, Hat06, HHSM88, LF88, MII^{+}11, SLS^{+}21, Ume01, KAD09, BAMK07, BAK09, KR10, PMMAM10, PTS^{+}12, SKST08, WRS12]. Cell-centric [BMA06]. Cell/BE [PTS^{+}12]. cells [AIIV98, GE92]. Cellular [ATT92, BMS01b, BMS01a, KVW97a, OCSBY01, SBZ04, Wei97, ATT89, AvHL^{+}97, BM01a, BMA06, CDR^{+}95, CCEJ01, CCCP92, CFMM97, DS03, FMS^{+}06, Hee97, Kat86, KVW97b, MHK97, MZ04, MCM01, Seu85, SC19, Sut97, Ume01, WNES01, Wei01, Yan04, Has90]. Cenju [HCH^{+}96]. Cenju-3 [HCH^{+}96]. centers [PVBR23, ZCT^{+}20]. centrality [SSKÇ15]. centralized [PCL23]. Centre [Kau94]. Centric [NDW^{+}19, BMA06, CHZ^{+}19, CFD^{+}16, IYV^{+}22, RH12, SFSV13]. Centroid [PJV00]. centroidal [JDG02]. CEPROL [Seu85]. ceramics [LVC16]. cerebrovascular [TAB^{+}19]. CFD [A^{+}01b, DK04, Ema10, EK98, FMS01, FR95, GK17, GKKS01, HAJK01, IUY^{+}04, JJAP^{+}20, Taf01]. CFD-codes [Ema10]. CFD’92 [Str92]. CFD’93 [Heb93]. CG [ABB^{+}16, BRS97, GG10]. CG3DR [WCKM11]. CGM [TNZM20]. CGM-based [TNZM20]. CGSTAB [BP02]. chain [Bur90]. chaining [TE92]. chains [Beb97, Reg01]. challenge [AGV^{+}17, CDK^{+}03, Hoc94, WBN^{+}17]. Challenges [Che98b, BJK^{+}17, GScFM13, Gin99, KCG08, MP08b, Sch95, WLK^{+}18]. Chamfer [ER95b]. Chan [SPW^{+}15]. change [CS00, TL94]. Changes [Ree07, TKG97]. changing [PCLM18]. channel [FT93, HT92, JHD^{+}22, SF91, WLZ^{+}23, Xu07]. channels [BKK24, CWK09, FMA17, N^{+}00]. Characterising [BS96]. characteristic [SK91, YLCT18]. characteristics [AAV08, PGAA16]. Characterization [FSKF06, FCM03, HMP^{+}16, HWYL89, YHE^{+}19]. characterizations [BCL91]. characterize [HC89]. Characterizing [FLPG18, GScFM13, LBGO23, PGK^{+}18, TYSF13]. chasing [KKW14]. Chebyshev [GR02]. check [ARCH05]. checking [BSK03]. Checkpoint [AKB^{+}19, NSM23]. Checkpoint/restart [AKB^{+}19, NSM23]. checkpointing [CQ01, GAR15, HW02, PY00]. checkpointing-recovery [CQ01]. Chemical [BGL^{+}88, DS96, DF00, NMW93]. chemistry [AC00, BS00a, Les00, LXD^{+}23, NRN00, R^{+}00, SSGF00]. Cheyette [God00]. Cheyette-beta [God00]. Chick [YHE^{+}19]. Chinese [KC93a]. Chip [SYP13, AFY^{+}16, LNLK13, NMMG13, PMS^{+}13, Tak06, GP86, AAS13]. chips [WLYJ13]. Chiral [EHN99]. Cholesky [TOO20, BO91, CGH^{+}19, Con89, DG95a, GP90, GHNL87, GLN89, Haa98, LKHL03, LC99a, Liu86, Ng93, RSD16, Stp92, YPL12, ZE92, ZG88]. Cholla [SRK^{+}21]. choose [CHZ^{+}19]. chopping [PG93]. Chromium [LBH07]. chromodynamics [BH99]. chromosome [BMCA98]. Chunked [ONOK13]. Chunks [RR14, ARR19, RR16]. Cilk [AV15]. Cimmino [Slo91]. Cinema [OAJ^{+}16]. cinematic [SLS^{+}21]. circuit [SM91]. circuits [Tor91]. Circulant [DZLK20, Far96]. circular [KT97, KH95, SKY93]. circular-arc [KH95]. circulation [JKH95, MSW98, WME^{+}95]. City [ER95a]. class [Bai99, BE89, BLKD09, DP90a, ED91, KN88, KC91, TCL92, Thu92, VL05]. classical [CT88]. Classification [MAJD17, BMA06, BKK24, EG92a, Gao87, GIRT19, Li91, MCM01, PPZ93, SSN04]. Client [TDC19, BA02]. client-server [BA02]. Client-side [TDC19]. climate [CDK^{+}03, DFM^{+}95, DF95a, DF95b, HRW^{+}95]. climatology [Adr99]. clique [Shy90, YL19]. CLJP [MX07]. clocks [CZTS99, DB03, LWZL24]. cloning [Hat06]. closed [HMTX93, KR10]. closed-loop [KR10]. closeness [SSKÇ15]. closet [Col04]. cloth [TB07]. Cloud [BY21, PCC19, BEDdC16, DST15, KRS^{+}21, LSC^{+}15, SWYM17, ST21, STE23, SLH^{+}13b, dNdRRL^{+}21, RMRN05]. clouds [CLY^{+}19]. ClustalW [DL06]. Cluster [AGV^{+}17, Bar12, BKS19, BRS^{+}08, FYH^{+}18, GGL^{+}05, HZY^{+}19, HG17, LCE^{+}18, LYC^{+}19, LHL18, NGM19, SER90, TBM16a, WHL19, AL93b, ABMN02, BCPB05, BEDdC16, BWL00, BKL^{+}19, BBL^{+}19, CSEK03, CLC^{+}18, CKS03, DWM^{+}01, DSSD18, GSMY^{+}07, GHS02, GW08, GP98, Hat06, HH17, IMQO^{+}18, JLLC22, JAA06, JMA^{+}13, KEEF01, KI05, LLX^{+}18, MBB^{+}18, MHH97, Moo04, NHS^{+}95, OIK21, PWY03, PL00, RBB^{+}22, SNS^{+}97, SKN04, SVC07, SDMS12, SSOB03, SS04, SSS^{+}09, SWR^{+}18, Tak18, TSLL17, WBS06, Wat00, WBN^{+}17, XT11, RMRN05]. Cluster-based [Bar12, CKS03, JMA^{+}13, OIK21]. Clustered [DFO93, Ara03, CE94, KBG^{+}01, NS11]. Clustering [DZ90, DL05, LF89, LSF^{+}15, MS98, Ols95, PRW88, SLG^{+}22, Tak06, TMCC02, ZRPI89, ZEC^{+}17]. Clusters [DST01, HVA^{+}16, ADPV03, AS00, ADMV05, BCVC05, BM11, BL94b, CGK^{+}16, CdCM07, CC00c, DGBP05, DL04, DHS00, DFH^{+}13, FKPB15, FWL03, FHK^{+}11, FHK^{+}15, HL97, HCL05, JS13, JKHK08, KNTG08, MCY^{+}24, MS03b, MKJC21a, MKJC21b, MSV15, NMW93, BHB21, OKTR13, PRS^{+}14, PM95, PSF^{+}15, RHWF23, SLG^{+}22, SMW^{+}05, VCK^{+}11, WZL09, WZS^{+}14, WY11, XLS^{+}17, ZCT^{+}20, ZKCL04]. CM [BP89, BEK95b, But92, SNS^{+}97, TM94]. CM-2 [But92]. CM-5 [SNS^{+}97, TM94]. CM-5E [BEK95b]. CM2 [SS94b]. CM5 [She96, SS94b]. CMMD [TM94]. CMP [KRN^{+}11, LLL10]. CMP-based [LLL10]. CNN [NSF^{+}22]. co [CCG^{+}17, DTV21, Kim98, LC17, Num05].

Co-Array [Num05]. co-design [LC17]. co-designed [CCG^{+}17]. co-minima [Kim98]. co-simulations [DTV21]. coalescing [MCP^{+}14]. Coarray [CFF17, FD18, FGBN19]. Coarray-based [CFF17, FD18]. coarrays [SC19]. Coarse [BBC^{+}11, CDBL08, BMP15, CS95, DG90, GPSK09, HHK95, KY98, KHHC20, Lil94, MAFC14, OHZ98, TTO^{+}22, dV94]. coarse-grain [KY98]. Coarse-grained [BBC^{+}11, HHK95, Lil94, MAFC14, OHZ98, TTO^{+}22, dV94]. coarse/fine [BMP15]. coarsening [MX07]. Cobra [BG01]. code [AG00, BDI^{+}95, BWV^{+}17, BDSV98, Bri95, Bur90, Cap88, CEM^{+}99, CDH^{+}03, DF91, DLR94, Dow90, DAG^{+}09, EJLC97, FMS01, FR95, GJS93a, GDAK06, GHH^{+}03, GAW96, Hat06, HC05, IJM^{+}05, IUY^{+}04, JIC96, Kie91, KSM^{+}94, KPL^{+}12, KRS^{+}21, LMJC96, LR89, mM94, MWI^{+}15, MB88, MPJ03, OCE^{+}07, PATC99, POHS14, PJW^{+}22, RKMJ10, SSP^{+}98, SLC^{+}16, STKA96, Taf01, TX00, TLC23, TBM^{+}16b, ZGZS20]. codebook [AB02]. codes [ARCH05, ALD01, BSE88, BF92, BC97, CGK^{+}16, Cha88, ES88, Ema10, EJLC00, FDTZ04, GKS07, GK17, GH89, IJCL96, MJ99, PHCR05, SMTT96]. coding [Kut02, Riz85, SKN04, SM15]. Coefficient [JNC^{+}19]. coefficients [AIIV98]. coffee [CCEJ01]. coherence [CD00, WC94]. coherent [MSS^{+}05]. collaboration [KYS23]. collaborative [LHZ^{+}22]. collapse [OCSBY01]. collection [Bev89, DH84]. collections [Wil85]. collective [BPJ22, CC96a, GGA19, HGLR07, HMS^{+}19, HG12, IZS^{+}20, KBG^{+}01, NZHY11, PGBF^{+}07, Trä12, THMH21, WBPM97]. Collectives [ZGZS20, Trä18]. collision [Wan08]. collision-free [Wan08]. collisions [CT88]. collocation [BP02]. Colloquium [Ano88a]. colony [KŠR04a, KŠR04b, SSBT19]. color [SPW^{+}15, SM15]. color-coding [SM15]. color-entropy [SPW^{+}15]. coloring [BSB^{+}22, ÇFG^{+}12, GK17, ZK92]. colouring [BFM90, Zer90]. column [AC94a, ST95]. columns [KC93b, PLP97]. combination [PHCR05]. combinations [TL90]. Combinatorial [Hag91, BAMK07, MGSK88, Rib84]. Combined [BNK15, Ala89, GV08]. Combining [HRR08b, MFS^{+}19, Lil94, MMTU18]. combustion [CYXL18, DG95b]. come [Rau98]. Comments [Eis89, FLW87, HE88, RT88, SSY02, Sim91, SE87]. commercial [CTS02]. commit [Yua94]. committor [BWT^{+}08]. commodity [B^{+}00, DK01]. common [GÁVRRL18, KS16, LO92, Lin94, TNZM20, TTO^{+}22, WFW^{+}11]. communicating [DC15, Ram99, Wil85]. Communication [Ber89, DCN^{+}15, ER18, GLN89, GS96, HHK95, HHQ06, HL99, LW11, SAOKM01, Ste90, VB92, Ann89, Arv92, ABMN02, AFY^{+}16, BFY06, BDK98, BAB^{+}02, BE87, Bek95a, BMG07, Cal96, CAHT17, CC95, CWC00, Cho92, CFW^{+}22, CNK93, CV17, Coo19, De 95, DVGG98, DT96, DWS^{+}21, DZD01, FMA17, FKK^{+}06, FBAB13, FA96, GP90, GZH^{+}23, GMS04, GNY00, GE92, HK02a, HC89, Hoc89, HC92a, Hoc94, HT92, HEB96, HK90a, HHJP16, Hwa04, IS18, IZS^{+}20, JWS13, JOL^{+}21, Joh93, KH95, KLP11, KBG^{+}01, KK92, KG03, KC93b, LC02, LFL11, LC99a, LX00, LBGO23, LBH07, MS99b, MFGEL17, Mun99, NZHY11, OKF16, Oed92, OID^{+}12, PVK^{+}22, PM95, PP98, RM97, RBM^{+}16, SS89a, SS95, SKS^{+}95, SSY02, SAOKZ06, SW91, SV94, TG09, Trä12, THMH21, Ume01, VP95, VLL90, Ver99]. communication [WBPM97, WZL09, WCC^{+}22, WWL^{+}22, YCBD19, YG93, YMG14, ZK92, ZL94, ZGG92, BD95]. communication-avoiding [IS18]. Communication-aware [DCN^{+}15, JWS13]. Communication-efficient [GS96]. communication-hiding [CV17]. Communication-optimal [HHQ06]. communication-related [Arv92]. communication/computation [DVGG98]. communications [ADLL03, BKT91, CC96a, CMT96, EPS98, EJLC97, GGA19, Has01, HRT07, HC92a, Lah00, PS84, VJ11]. communicators [GJMM18]. community [DFM^{+}95, GGH^{+}22, HRW^{+}95, LHK15, TMD^{+}97, ZY16]. Compact [QP16, GGL21, PW84]. compaction [GF03, SC92]. Comparative [FRC^{+}95, GT92, BEK95b, FB19, LCD91, SB94a, SH88, SS94b]. compared [HL88]. Comparing [AJ89, Bai88, BD18, LHR^{+}19, MTV08]. Comparison [GP85, HC92a, KHHC20, PER17, Ahm97, ABMT99, BM93, BP89, BDOS95, But92, CWB92, CKLM14, GLP98, GE93, HJ97b, Hur93, KN88, MSP93, MOF04, PVP08, RW98, Sha06, SH01, TSJB00, Trä95, YTSI13]. comparisons [BE89, LR89]. compatible [BDP20, EL16, OIH10]. compensation [HWL^{+}22]. Competition [AGV^{+}17, BKS19, FYH^{+}18, HZY^{+}19, HG17, LCE^{+}18, LYC^{+}19, LHL18, NGM19, WHL19, BKL^{+}19, BBL^{+}19, CLC^{+}18, DSSD18, HH17, LLX^{+}18, MBB^{+}18, SWR^{+}18, TSLL17, WBN^{+}17]. competitive [SMT^{+}04]. Compilation [GPPS99, KCRB98, CAS09, JPK^{+}15, KY98, MFS^{+}19, MAFC14]. compile [HJ05]. compile-time [HJ05]. Compiler [ADK22, AKP^{+}98, BWV^{+}17, NS11, AART13, FAS02, FDTZ04, HT00, SM13, SNK06, SSO^{+}14, THH^{+}05]. Compiler-assisted [ADK22, NS11]. Compiler-based [AKP^{+}98, BWV^{+}17]. compilers [DFRR91, LCD91, PB23, YTCS97]. Compiling [BZ99, Lee95, WMCU97, A^{+}01a]. complement [DO88, XKL^{+}22]. complete [A^{+}01b, CC95, CHH^{+}01, LH04b, PRS^{+}14, TT00]. completeness [KB20]. completion [KKU16, NYFK06, PHM^{+}22, SSN^{+}21, SPK18]. complex [APRP97, AV15, BMS01a, Heg96, HGS10, IS18, LS93, Reu99, TO99, YR18]. complex-shaped [YR18]. Complexity [HBH^{+}16, BKT91, CS95, ČSR05, Dar00, DB03, Feo88, GK92, HJ05, LAS90, NMARD10, RG09, YXQ^{+}21]. complexity-effective [RG09]. compliant [BK07]. Component [CGG01, BPP10, CDM06, DFRZ02, DAG^{+}09, FMM^{+}02, HSÇ12, HLP10, IKK15, Rob00]. component-based [DFRZ02, DAG^{+}09, HSÇ12]. Components [BGL01, B^{+}00, CCS^{+}18, DK01, LDK16, NBB^{+}02]. composite [Thu91]. composition [LDK16, OILZ17, SKN04]. compositional [CCH^{+}21]. compositions [GS06]. compounds [BM01a]. Compressed [YPL12, BVWH14, GAR15, HA05, LB23, SMW^{+}05]. compressible [CLL99, Ema10, HM01, Lan96b, OJ90, SB94a, TW19, TMD^{+}97, YLCT18]. compression [AHL02, BTLK18, De 95, KNTG08, SS91a]. Comput [HE88, SVS02]. Computation [Tak18, AGO97, AHBLR12, ABH^{+}10, BM09, BG02, BL93b, BFVRC14, Bro86, CM95, CR23, CARW91, CKM94, CCRR91, Cor99, CJ95, DS96, DVGG98, DM90, ES89c, Fis90, FP98, FPT91, FCM03, GL90, Got89, JTS^{+}11, KT97, KK01, KKB92, LHK^{+}96, Li90, MAB17, MS00, McC89, MP02, dTNOR^{+}04, ORM^{+}10, OA91, PAF^{+}97, RLH19, RD07, RBS10, SOS97, SK91, SK19, SMK20, SHH^{+}97, TTH95, TO99, TLC23, WC90, XT11, XXQ^{+}15, YD97, ZM96]. Computational [Ano88b, BBDN21, CEM^{+}99, HRW^{+}95, Les00, Liu86, MLM^{+}00, NRN00, R^{+}00, She96, A^{+}02, BA08, BMS01a, BG07, BR84, CDR^{+}95, CdCM07, CW06, CL05, DM03, DGNP88, DST01, DNL15, FGBN19, AKK16, Feo88, Gen84, GWC^{+}99, GK92, GV06, GIF^{+}10, GM04b, GH98, HBH^{+}16, JMLBL98, MMT06, NLG99, PATC99, QWD^{+}21a, QWD^{+}21b, Sýk84, TW19, TKG97, WRB97, ZEgT04]. computational-graph [QWD^{+}21a, QWD^{+}21b]. Computations [AK89a, ZVWS88, APRP97, Bar95, BI07, BGR00, CNK93, CDC^{+}87, CB09, EM90, GF89, GVH08, HSS07, Her00, HHR87, JS13, KCRB98, KD13, LBL95, Lan99, Lin08, Lip99b, ME93a, MRRP11, ND17, OSZ93, OZ02, PTGF20, PVK^{+}22, RCAP11, SA13, SBZ04, SVTSM^{+}22, Ste90, TS09, WL13b, WL14, Zom95]. compute [SB92b]. computed [SSS99]. Computer [Ano16k, IJCL96, AMP92, BHM^{+}95, BJP^{+}89, BCYB11, BJSN04, CCCP92, CCRR91, CMRT88, DD89, DSCP88, Dow95, Eck90, ETV91, EY90, EL90, FLS^{+}21, Fre89, GR89, GF89, GK92, HH89, HKSK97, Han85, Hio96, Hoc87, Hoc89, HST87, HHSM88, IU87, IHM^{+}12, Jeo91a, Jeo91b, KBBC88, KTN^{+}14, KBD93, LQ92, mM94, MMSS22, MR89, MJ94, Maw99, MMS90, MSW91, ME92c, NMARD10, PJV00, PAD02, Rea90, RBS10, SW91, Sch91a, SHRN98, Sch87, Sho17, SB92a, SS92, Sug99, Tre85, Uka99, WN91, WBPM97, WZ90, SWSG92]. computer/array [Str87]. computers [AL93a, AO89, APG92, AAZ96, AL88, Bai88, BMS97, BE89, BE92, BE93b, BBQOQO00, BRRV11, Ber00, BP00, BFM90, BEK95b, BP86, BT88, Bru91, CF91, CDW95, CT88, CMM^{+}88a, DL92, DDP90a, DDP90b, DS03, DR93, DLR94, Deá90, DJH05, DG15, EL16, ER89, GMBZ90, Gen84, GMF00, Ger98, GHH^{+}03, Güs99, HISS92, Hat06, HW91, Hot89, Ken99, KC91, KSP97, KKŽ05, Koh95, KLS^{+}88, Kuz98, LPM11, LR07, LPCA98, Lec89, LC02, LE91, Lop93, Mak94a, NS94, NM01, ND95, ORS90, Ort88, PG91, Par87a, PW84, Par87b, PZE94, PJW^{+}22, Rap99, RS87, RV95, SK97a, Sch91b, Ste88, SP93, Swa84, TS99, Wal94a, Wal94b, WLCG02, WS12, WAB89, Wri90, ZRPI89, ZM96, dV94, de 96, van86, van87b]. Computing [Ano87, Ano94, Ano15a, BAK09, BVC16, BVC19, BMYK98, Bis89, BDO17, CHR18, CP17, CBV13, DHS89, HK92a, Hor93a, KŠR04b, Lun98, NDW^{+}19, RST11, RC17, TBM16a, WBD99a, AC00, AU88, AS00, ABMN02, AMMV98, BGWR21, BCPB05, BG01, BY21, BMCA98, BHK00, BJK^{+}17, Blo03, BWL00, BDRV99, BH99, B^{+}00, CIS99, CTS02, CS93, CLS^{+}16, CD98, CCG^{+}17, CLC08, CP18, CKRZ98, CHR00a, Cla03, CC99a, CMM03, CQ97, DDdS02, DT97, DST01, DF95a, DF95b, ESV10, EHF^{+}97, FWL03, cFM07, FG01, FGP23, FMM10, GKT^{+}15, GK03, GGV04, Ger15, GYL00, GHC04, Hac89, HP05, HP97, HK00, HP94, Hoc85a, HK06, HOSS91, HR04, HMH^{+}13, JJM^{+}11, KLR05, KH95, KCG08, Kog85, KG03, KNR00, Kon02, KKK^{+}18, KS99, LLPV06, LTG23, LD99, LM03]. computing [Las02, LG01, LM00, LLFZ23, MD04, Man01a, MKL^{+}01, MT97, MSZM14, MT00, MSK14, MTK03a, MTK03b, MSE95, MOW96, MSOCG^{+}16, MHL06, NHS^{+}95, ONB11, OGC^{+}15, PY00, Pel97, PWY03, PCC19, QH19, RMRN05, Rob99, RBP^{+}17, Sab97, SEAH^{+}08, SÖB07, SVC07, SRK^{+}21, SP94, SGS95, SR14, SS94a, SC03, ST21, SVC19, SBM^{+}22, SHCR98, SDMS99, SDMS05, SMTT96, SGDM94, Sun97b, SG99, TKK^{+}05, TLS^{+}08, TND10, TAB^{+}19, TF01, UZ02, VRL03, WL13a, WA03, Wan08, WFW^{+}11, WJK98, WS12, W^{+}99, YPG03, YCS07, YDTS01, YD07, ZLT19, Zen99, ZJD^{+}23, ZYZ^{+}16, Ano16k, BBD^{+}12, Jou97, Ben24, MSS19]. concatenated [SMM90]. concatenation [ASH92, LyHW^{+}22]. concentrations [PODd16]. Concept [Qua92, GRP22, PAD02, BDM99]. Concepts [Ale91, DHW97, EM88, Jor88, SAWH88, Tol02]. conceptual [IJCL96]. Concurrency [Ghe85, PSW18, Van90, WLK97].

Concurrent [DG95a, MAB17, ATT89, Arv92, BCRSR11, CDW95, DSS86, GW08, HISS92, JDH96, LOST16, MT89, PS84, SS94a, Son86, SGDM94, Wal94a, Wal94b, WW90, WCC^{+}22, ZLJ93]. condensation [MV99, RV95]. condensed [ACH^{+}18, DvdG92, WCKM11]. Conditional [Fu07, KLL^{+}09, MJ95]. conditions [BP95, GW87, RW96, ÜD89]. cone [LPCA98, OIH10, SKH^{+}12]. cone-beam [LPCA98, SKH^{+}12]. CONeP [YNY^{+}17]. Conference [Ano87, Ano88b, BH92, BR84, Ano18k, Ano15a]. conferences [Ano97]. configurable [GGL^{+}05, ZSI02]. configuration [DHS89, Far96, HOSS91, KRN^{+}11, KI05]. configurations [ADEQO19, CG93, JJAP^{+}20, Wil85]. confinement [TTT^{+}92]. Conflict [CL92, LY93]. Conflict-free [CL92, LY93]. conflicting [SG16]. conflicts [LX00]. Congestion [HP97, CDZ^{+}21, TT00]. congruent [Box92, SLY90]. congruential [Mas98, MC04]. Conjugate [VJ12, BM90, CG89a, CGMM99, Daḡ07, ETV91, EL90, EG94a, GZHX17, GV14, HGLR07, Kru97, Meu87, O’L87, RR89, Sea86, Str08, TGE92, van86]. connected [AAZ96, BV04, CC95, Che98a, CCS^{+}18, HI92, Hor93a, HST87, HHH92, IKK15, Jeo91a, Jeo91b, MSS^{+}05, Sch91a, Son92, WN91, YD01, ZM88b, Li94]. connected-component [IKK15]. Connecting [Sch88a]. Connection [DLPS92, McB88a, Sin90, AIO95, CHH^{+}01, CGOP03, SS94b, BL94a, JK92, LSP92]. connectionless [GRBA15]. connections [WSB90]. connectivity [FJPA17, Ume01, WSB90]. conquer [Arb92, Bon91, ER95a, ER95b, GTD18, Gor97, HvNJB12, LPAZ97, LSF^{+}15, Meh93, RM91, WP19]. conservation [DG05, GL90, TDG^{+}18]. conservative [BT01]. conserving [BTZ06]. considerations [EG90, JKH95]. considering [WLZ^{+}23]. consistencies [LG03]. consistency [DPFT19, OGC^{+}15, PWY03, RRA11]. consistency-based [OGC^{+}15]. consistent [SSP^{+}98]. Constant [Kim98, AIIV98, BT99, Hee97, Mar92, OSZ93]. Constant-time [Kim98, OSZ93]. constants [BM93]. constrained [BP95, BHR09, BE93c, BDK96, CAS09, DM03, DFO93, ECLV12, KJA15, Kha12, KAM^{+}20, MKK^{+}19, NFG^{+}13, QWD^{+}21a, QWD^{+}21b, QJ06, SWCBQ19, TTO^{+}22, TSCS14]. constraint [Her00, Kon11]. constraints [BMP15, BGMT00, GO88, TNZM20, WL13b]. Constructing [CG93, CST02, Far96, Jeo91a, WTLW23]. Construction [EM87, Hio96, CRS88, Hwa02, KS07, KSS06, MGCB^{+}10, YTCW07, ZGL^{+}19]. consuming [DST15]. consumption [ACM^{+}15, LLFZ23, RMS^{+}18b, TYLL22, WC15]. container [BGM03, BCG00]. containment [SSLK00]. Contention [BHR09, GNY00, CGGG03, DSG17, HK90b, LCLA19, SS04, WCC^{+}22, WH94]. contention-aware [LCLA19]. contention-based [WCC^{+}22]. Contention-free [GNY00]. Context [LWW^{+}21, BCVC05, CC96b, DR18, JC94, LGM09]. context-free [CC96b, JC94]. continuation [SWSG92]. Continuations [SSN^{+}21]. Continuous [AAB^{+}16, SVS01, SVS02, ZSI02, ALNT04, Beb97, GM98, RBL97]. continuous-time [Beb97]. continuum [SG16]. continuum-Lagrangian [SG16]. contour [ALH^{+}14, PS89b, SPW^{+}15, YTSI13]. contraction [AED21, BM93, Rag97]. Contractive [BM93]. Contribution [FP98, CMP92]. Control [Tre85, Ala89, Amm89, ABH^{+}10, CDZ^{+}21, DCG90, FJS85, GE11, HP97, KR10, MPJ03, PSF^{+}15, Qua92, Tap84a, WLK97, Wor97, Wri90].

Control-driven [Tre85]. controlled [CZJS12, OKF16, PY00]. Controller [HHGA15]. Controlling [BBDN21, DP95, MCB16]. convection [CEM^{+}99, CKLM14, FB19, GG10, LPMD01]. convection-diffusion [LPMD01]. conventional [CIO^{+}17]. Convergence [BP93, DW99, ESK88, ÜD89]. conversion [BF92, FCC16, Gol86, HA14, SLS^{+}21]. convex [CC02a, CLHL23, DZD01, El 93, EM85, HI92, LSS05, SE87, Xue96, ZM88a, WSL88]. Convey [BNK15]. convolution [AFK01, EG94b, JHD^{+}22, ZSH97]. convolutional [LXL^{+}22]. convolvers [Ume89]. Cooley [JK92]. cooperating [LL90]. cooperation [Jia09]. cooperative [GM03, LH04a, MMT06, RB14, RR07, TCT00, TCS04]. cooperativity [Yan04]. Coordinate [BDOS95, BBL^{+}16, DRST03, NC97]. coordinate-based [BBL^{+}16]. coordinated [LSC^{+}15]. Coordination [ACH98, WCKM11, Cv98, DF98, EK98, PY00]. coordination-based [Cv98]. coprocessor [HXW^{+}13]. coprocessors [GVH08]. CORDIC [BAZ93]. core [AGL08, AAV08, ABB^{+}16, BJ18, BNS^{+}07, BB21, CFF17, CU04, ÇFG^{+}12, CLA15, CJLS14, CDGI15, CN98, CKS03, CIO^{+}17, DTR18, DEH^{+}11, DHW97, DFH^{+}13, FD18, AKK16, Ger15, GB14, JHD^{+}22, JJM^{+}11, KCRB98, KD10, LXC^{+}18, LXL^{+}22, MMP^{+}21, NMAB19, PB11, PTS^{+}12, RGT17, SBP12, SA13, SWCBQ19, SFSV13, SPW^{+}15, TYLL22, VJ11, VA14, WZS^{+}14, WYX^{+}22, WY11, WPS^{+}08, ZJD^{+}23]. core-level [TYLL22]. CoreLib [WPS^{+}08]. cores [GIF^{+}10, JMA^{+}13, MPZS13, SC19]. corner [Wan09]. coroutine [WH22]. correction [HWL^{+}22]. corrections [BMP15]. Correctness [DFP^{+}19, vS92]. corrector [RM97]. correlated [Ber00, Lin08]. Correlation [JNC^{+}19, CYDJ21, CCRR91, Cor99, DLPS92, GICM18]. correlations [DP95, dP90b]. Corrigendum [ZK92]. cortical [YMJT10]. Cost [LSC^{+}15, SLH^{+}13b, ATA^{+}16, BG97, BR99, CLY^{+}19, Che08, CC97b, CC00c, CWDG07, CJ95, DHS00, FIMF99, HBH^{+}16, KV99, LLW^{+}15, LC90, LO92, LC99a, LWW^{+}21, Mar92, SA16, SSL03, SB92b, YNY^{+}17]. cost-effective [CWDG07, YNY^{+}17]. Cost-efficient [LSC^{+}15, SLH^{+}13b]. cost-optimal [LLW^{+}15, LC90, LO92, SB92b]. cost-rejection [CLY^{+}19]. costs [Gre89, LFS^{+}19, MMP^{+}21, NDN20, TN13]. Couillard [MAFC14]. count [SWR^{+}13]. counteract [KP96]. counters [IZS^{+}20]. counting [Bev89, Kim90, Ski02]. coupled [CC99b, Eva90, FBMV88, Gro87, HSS07, PL00, Qui88, Wat00, WE91b]. coupled-cluster [PL00, Wat00]. Coupling [GIF^{+}10, Wei01, TL96]. covariance [GGFF93]. Cover [Ano02f, SL90a]. coverage [ZSI02]. CP [A^{+}99, KKU16, NNB^{+}99, Uka99]. CP-PACS [A^{+}99, Uka99, NNB^{+}99]. CPMD [AC00, HC05]. CPU [BEK^{+}11, CTW14, CKLM14, FLYL21, FHK^{+}11, FHK^{+}15, GK19, HM02, Hoc85b, LLW^{+}15, LFL11, LGD^{+}15, SPW^{+}15, TPK^{+}13, WZCG14, YWP11]. CPU/GPU [ADEQO19, GK19]. CPUs [KD10, Nag88, SFSV13]. Crank [ZL94]. CRAY [Hoc85b, ARW93, Cal85, DCG90, DD87, DH91, EO91, FSY88, Gur88, HL88, HKN89, Meu87, MSW91, Nag88, Nag90, Pin91, SW91, Tem88, Tem89, But92, CC96a, GYL00, HEB96, KN88, KLN90, OL86, Oed92, Sea86, WLN^{+}96, ZCPT00]. CRAY-1 [Tem88, Tem89]. CRAY-1M [HL88]. CRAY-2 [DCG90, DD87, FSY88, But92, KN88]. CRAY-X [HL88]. CRAY-X/MP [HL88]. credit [PSF^{+}15]. credit-based [PSF^{+}15]. critical [Kha12]. Criticality [NB12]. critique [Yue97]. cross [GSC^{+}22]. cross-platform [GSC^{+}22]. crossbar [AB93, Qua92]. crossed [EBSS94]. cryptanalytic [OK22]. cryptography [SBZ04]. cryptosystems [BF06]. Crystal [LHZ^{+}22, Kam87]. Crystals [LHZ^{+}20]. CS [BCM94, BHM94, DTV21, Hoc94]. CS-2 [BHM94, Hoc94, BCM94]. CSD [HL99]. CT [FA11, SKH^{+}12]. Cube [GS04, Bro88a, EBSS94, Hor93a, HWYL89, LC17, LH00, PMAL14, RBB^{+}22, Tak06, Li94]. cube-connected [Hor93a, Li94]. cube-shaped [PMAL14]. cube-type [Tak06]. cubes [CCP98, HC04, HCTH05, MLX07, YTCS95, LH04b]. CUDA [DS13, DR18, DWL^{+}12, GScFM13, HVA^{+}16, HLP10, KD13, NSM23, OKTR13, OIH10, PVK^{+}22, PRS^{+}14, SBM^{+}22, YA24, YMYH24]. CUDA-Aware [HVA^{+}16, PVK^{+}22]. CUDA-enabled [DS13]. CUDA-sharing [PRS^{+}14]. CUDA-to-OpenCL [GScFM13]. cumulants [MS00]. Current [FK94, DDdSL02]. curve [CARW91]. curves [Ban97, GZ99, Kon11]. curvilinear [NC97]. curvilinear-coordinate [NC97]. custom [LC17, JNC^{+}19]. Cut [FC18, Ral03, ZG16]. Cut-off [FC18]. Cuts [MBC97]. cutting [HAK^{+}21]. Cyber [LN87, PGGP19, HL88, Riz85, Won88]. cyber-physical [PGGP19]. CYBERPLUS [KRW88]. cycle [CC92, CKM93, Fu03, Hsi06, RS92, YX07]. cycles [AFJG06, KF95, KL00, Li94]. Cyclic [CTZ^{+}18, Fie96, AM93, BE93b, Chr00b, EW93, GS89, Hwa02, Hwa04, KS91a, LC02, NHS^{+}95, Sch91b, SWSG92]. cylinders [KT97]. cylindrical [MO89].

D [AAS13, AW13, CYDJ21, Cap88, EK98, GV01, HKL01, RVGG01, SSS99, Tak03, VvBv90, WSR14, AW13, BSE88, BFG^{+}07, BCMG^{+}07, Bri95, BAS13, CDM06, CM95, Chr98, DZD01, EGTD99, FMS^{+}06, GJS93a, HSS07, HB84, JHD^{+}22, KT97, KTN^{+}14, KD13, KYLH01, Lee97, LCLL00, LPMD01, MW94, MM96b, MDC^{+}08, OCE^{+}07, RVGG01, SHPA05, SA16, SC19, TKG97, TT00, TA14, Wan09, WTLW23, YR18, YL86, ZLM98]. D/ [RVGG01]. DACCOSIM [DTV21]. DAG [BBD^{+}12, LKHL03, YZB^{+}23]. DAGs [Ara03]. DAGuE [BBD^{+}12]. damaging [FC05]. dance [WBPM97]. DAP [AMP92, BP89, Car88b, Dav88, EL91, LL88, PR90, Wai88, WC90]. DAP510 [EG92a]. DAPSYS [HKK97b]. DaSH [GSE^{+}15]. Data [ACMT20, A^{+}02, AZ98, ABG^{+}03, CHLS90, CSH^{+}11, FAS02, FMD98, FTY^{+}20, HS03, KA19, KP93, KS95, LK14, LR07, MV17, Mv88, NDW^{+}19, NOG^{+}22, NZ92, OW97, RM97, RW98, RZ95, ROZ01, RC17, RH12, SS89a, TPK^{+}14, Van90, WXL^{+}18, ZPAT99, ZM96, AL93a, ABB^{+}16, ATA^{+}16, ANG^{+}20, AKSS07, AFT23, AGK98, BSH88, BG91a, BKS06, BF95, BEDdC16, BD18, B^{+}02, BM02, BDOS95, BETR17, BDK06, BGR00, BCMSW03, BCMSW04, BJV^{+}16, CTS02, CH98, CFS01, CWK09, CZJS12, CS19, CDK^{+}03, CC16, C^{+}01b, CRD02, ČSR05, CMM03, CV02, CDA^{+}16, DGNP88, De 95, DR93, Dec00, DF00, DAD11, DMG^{+}04, EG95, Evr01, EG90, GWLS05, GWC^{+}99, GMMT16, GM04a, GM04b, GLD08, GSC^{+}22, Haa98, HLvHA11, HMTX93, HCJ03, HCR01, Hwa04, IOH05, IYV^{+}22, KMB^{+}18, KN88, Kan93]. data [Kan97, KMLM97, Kas85, KY98, KG21, LLP00, LyHW^{+}22, LC89, LWJ^{+}17, Mah96, MSE19, MJ94, MIA^{+}07, MC09, MMM13, MQ97, MS20, BHB21, NL98, NJ02, NMI^{+}24, ONOK13, PGGP19, PM03, PVBR23, PFTB15, PB16, PS91, RBB^{+}22, Ran97, SNK06, SB94a, SSH97, SSS99, SC03, STE23, SV97, SPMB23, SSN04, TS02, Thu92, TS09, Tre85, TL96, THK^{+}99, Vio04, WYX^{+}22, WJK98, WCM^{+}14, WE97, WE91b, WLK^{+}18, XLS^{+}17, Y^{+}02, YZWcF14, YEC97, YM05, ZLD15, ZJDW18, ZCT^{+}20, ZGL^{+}19, BCC^{+}97b, MAFC14]. data-centric [IYV^{+}22]. data-dependence [NMI^{+}24]. data-dependent [BM02]. data-driven [Evr01, MSE19, Tre85, ZLD15]. data-flow [EG95, EG90, MAFC14]. Data-Intensive [RC17, ANG^{+}20, TS02]. data-localization [KY98]. data-oriented [PB16]. Data-parallel [FMD98, HS03, RW98, ZPAT99, ADEQO19, BDOS95, MIA^{+}07, SB94a, SV97, ZGL^{+}19]. Data-structure [BCC^{+}97b]. Data-task [ROZ01]. database [Bak00, BF95, BV96, MC09, NS02, Reu99, SCM^{+}98, SJJ85, TDW^{+}04, ZHC^{+}23]. databases [LTS^{+}24, PRW88, PWM00]. DataCutter [B^{+}01]. DataExchange [EPS98]. dataflow [ACM20, GSE^{+}15, Gao90, HSÇ12, MMTU18, NLG99, OKSY92, RKS92, TB22, WMCU97, WvR16]. dataflow-graph [RKS92]. DataFrame [GSC^{+}22]. DataFrame-based [GSC^{+}22]. datagrams [GRBA15]. dataparallel [OJ90]. datarepresentation [SJJ85]. dataset [LyHW^{+}22]. datasets [B^{+}01, LMC05, NR20, SMW^{+}05]. datatypes [CAHT17]. DBT [DZ90]. DBT-Transformed [DZ90]. DC [XZYQ21]. DDM [BDK06]. DDS [LLFZ23, WL13b]. DDS-based [LLFZ23]. deadlines [Ver99]. Deadlock [Fei91, Ann89, LWZL24, WL13b]. deadlocks [Ara03, WLYJ13]. Debugging [KGV97, CSH^{+}11, KCD^{+}97, RGdS^{+}13, WOKH96]. DEC [KW90]. December [Ano19a, Ano20c, Ano21b, Ano22a]. decentralized [AGGG06, LMG09, MTV08, TKK^{+}05, Yua94]. decimal [Tak10]. decision [GE11, KNR00, QP16]. decision-making [KNR00]. decision-trees [QP16]. declustered [SSH97]. decoding [ARCH05, AFS14, De 95]. decomposable [NDN20, TN13]. decomposed [EEH^{+}19]. decomposition [AC94b, AHW14, AIV95, AW13, BG91a, BMP15, BO91, BZ96, Bis89, BM01b, BFR93, CS95, CUSR88, CMV^{+}06, DBVS01, GK03, GHP10, GK92, GW87, HJ97a, HP94, HJ97c, HZ93, HK06, HSFS14, IR93, IR00, KS91a, KN09, Kon11, KC93b, Kon02, MW11, NZ97, PAF^{+}97, SMC16, SG16, SB92a, TOO20, Ter13, TRSC^{+}19, WE94, WA03, Wri91, WTLW23, YPL12, YLCT18]. decompositions [CC97a, FRC^{+}95, Ste94]. decoupled [Evr01, RG09]. Decoupling [SE93]. dedicated [ABMT99, DFR92, GGA19, GW08, RSC^{+}19, Sug99]. Deep [DZLK20, AMC^{+}19, CDZ^{+}21, DBZ^{+}19, GZH^{+}23, JAH^{+}18, KLP11, LWW^{+}21]. Defeasible [WBD99b, WBD99a]. defect [MPJ03]. defined [KYS23]. definite [HRR02, NYFK06, ZVWS88]. definiteness [EV89]. definitions [Kut85]. deflection [AFY^{+}16]. deformable [SHPA05]. degradable [WWL^{+}22]. Degree [VTmL03, BT99, BDK96, LAS90]. degree-constrained [BDK96]. Delaunay [FCC16, KKŽ05, NR20]. delay [CC92, SAOKM01, WLZ^{+}23]. delayed [BLT^{+}22]. delays [BDK98, BP93, BP95, DT96, GMS04, HK02a, MS99b, Mun99, VLL90, Ver99, YG93]. deleting [KC93b]. Delft [de 87]. delivery [CGGG03]. Delta [HC92a, Pir96]. demand [CS97b, JSC97, SWR^{+}13, Tre85]. demand-driven [Tre85]. demography [MSOCG^{+}16]. Dempster [WH93]. Dempster-Shafer [WH93]. Denelcor [DH84, HLM84, Hir86]. Dense [BLRR02, RRS88, ASA16, BBB^{+}14, BP90, CTW14, CG89b, CTT89, CIO^{+}17, DGP^{+}16, EB98, FKK^{+}06, Kal92, KMB^{+}18, Kuz98, LR07, MS03a, Mal02, MRBQO14, Mel87, ONOK13, SZS22, TDB10]. densely [MHE19]. density [ARCH05, MCM01, NHS^{+}95, PTGF20, WYBdJ^{+}21]. dependence [AZ98, CDRV97, JCE96, LC89, NMI^{+}24]. dependencies [Bra88, BCR00, GF13]. dependency [LKHL03, MT91, MT95, Man01b]. dependent [BM02, RSK99, SMW^{+}05, WLYJ13]. deployed [ZS21]. deployment [BK07, NOG^{+}22]. depot [BCG00]. depth [SSS99]. Derivation [DQRR00, WL14, Cv98]. derivatives [LM00]. derived [BOS^{+}09, CAHT17, FF21]. Deriving [vCd90, Tyr90]. descent [BG02, KL90b, MA22]. describing [Zla88]. description [BEK95b, CR90, DFRR91]. descriptions [SBP12]. Design [BCPB05, BAZ93, BGRD00, CCH^{+}21, DFM^{+}95, FD18, GM98, GV08, LGD^{+}15, MGT^{+}13, MS20, SKY93, SCM^{+}98, SR10, VGRS16, ZGZS20, AB02, BAMK07, BHM94, Cv98, CLHL23, DAA94, Eva89, Eva92, EG90, FSCL06, FWL03, GGA19, GLP98, Gor97, GVH08, GM89, HRW^{+}95, Kas85, LT94, LC17, ML00, MGCB^{+}10, MCC04, ME90b, MAS06, NOG^{+}22, PSF^{+}15, RBB^{+}22, RG09, RSC^{+}19, SSD^{+}94, SS91b, TL90, UN87, Ume89, Wal94a, Wal94b, BWL00]. design-pattern [MAS06]. Design-time [CCH^{+}21]. designed [CCG^{+}17, CTA22]. Designing [MTV08, PP98, KJA05, Lin01, ZLJ93]. Designs [HVA^{+}16, ALD01, EMB89, EM90, ME90a, MBC92]. desktop [CdCM07]. detach [PHM^{+}22]. detail [SSV^{+}16]. detailed [Wei91]. detect [DPFT19]. detecting [ZLJ93]. detection [Arv92, BJSN04, BNK15, CCS^{+}18, CRD02, DIR97, DCN^{+}15, Fei91, GGH^{+}22, GGL99, GF13, HISS92, LHS13, LSA^{+}95, LHK15, MMTU18, MSE93, SPW^{+}15, TDW03, WL13b, XXQ^{+}15, ZY16]. detection-based [WL13b]. Determination [Bra88, Ban97, GM04a, GM04b, KSS06]. determine [MFGEL17]. determining [BFY06, BW92]. determinism [vS92]. deterministic [CCP98, KJA05].

Deutschland [SK97a]. Deutschland-Modell [SK97a]. Developing [AMMV98, BSH88, PRR03, CSH^{+}11, DS87, Kas84, SN88]. Development [Oya99, WWR05, WH91, CLC08, Gil88, JKH95, KCD^{+}97, MMP^{+}21, OKSY92, SEAH^{+}08, Sho17, YCS07, Ano87, R^{+}00]. Developments [DvdV99]. device [LS90, MW94]. devices [DWH^{+}08, LK10, PZE94, YZB^{+}23, ZJDW18]. DF [GSC^{+}22]. DGST [ZGL^{+}19]. diagnosable [Cho92]. diagnosing [GWWM09]. diagnosis [BAR98, Don05, Li90, SYAU07, TJ97]. diagnostic [BS97, BMYK98, CDDG93, GGFF93]. diagonal [DS05, KGS08, PG91, Sun95, YD97, Zha91]. diagonalisation [RH99, TY91a]. diagonalization [VP94]. diagonalizing [KG87]. diagonals [SM87]. diagram [Jeo91a, Jeo91b]. diagrams [ES89c]. dial [ACGL04]. dial-a-ride [ACGL04]. dialects [CWB92]. dialkyl [STKA96]. diameter [GDMS97]. diatomic [MT97]. Dichotomy [Ter10]. dictionary [DG90]. difference [EL90, GF89, MWI^{+}15, MSW98, Ram07, She96, Thu90, Thu91, WME^{+}95, Won88]. different [DF00, JCC^{+}24, NAC^{+}14, VRGÁ15]. Differential [SKN04, AAHF97, BP02, CLS93, EM87, GR89, Gol88, GW87, HM89a, HK92b, MQ89, RSSS94, RMS^{+}18b, SHPA05, SN88, SS88, DST15]. differentiation [CDGM96, Fis90, HM01]. diffusion [BCMSW03, BCMSW04, CKLM14, FBMV88, HSR^{+}14, HSS07, Hor93b, HB99, LRH97, LPMD01, MD88, MMS09, Wei97]. digit [SB94b, Tak18]. digit-reversal [SB94b]. digital [CYDJ21, EG94b, EA95, PS89b, Wel89, WS98, Y^{+}02]. digits [Tak10]. Dimemas [LGC97]. Dimension [HT98, CCRR91, Cor99]. Dimension-exchange [HT98]. dimensional [AL93b, AIIV98, Bas94, BE87, B^{+}02, BP86, CFMM97, CC00a, CCW01, DE09, DG99, DDF^{+}10, EM85, EJLC00, Far94, GGL21, Hee97, HHJP16, HF20, JWS13, KR97, LPCA98, LS92, MM96a, MSA09, MCM01, OJ90, PSW18, QCC02, Ram07, RSK99, Sim91, SBO^{+}91, Tro00, WBS06, YB11]. dimensions [Eva91, GHW08]. dimeric [PB98]. diodes [SAWH88]. dioxide [PODd16]. Dirac [RSK99]. direct [ADPV03, Bjø87, CL02, EW93, GW95, GZTR07, HRR02, PGGL17, PER17, RW98, SZS22, VKS^{+}15, Zam99]. direct-neighbour [Zam99]. directed [MQ97, SCD92]. Direction [CC01]. directions [FK94, UAR^{+}99]. directive [GHH^{+}03, HST05, KKHY06b]. directive-based [GHH^{+}03, HST05]. directives [BC19b, CLZ01, GB14, VGS14]. directory [HK02b]. disaggregation [Beb97]. discontinuous [HBC19]. discovering [MMO07]. Discovery [BDO17, AED21, BOG15, CRGR^{+}13, CC16, LLFZ23, TKK^{+}05, WCM^{+}14, YBM05]. discrepancy [Lec89]. discrete [BAK09, CLS93, DF98, EK92, GJMMS97, Mil01, Not95, PM89, Wri90, CT94, YR18]. discrete-time [CLS93, Wri90]. discretization [BP02]. Discretizations [ABB^{+}21, Lep99]. DISCS [RC14]. DISCS-2013 [RC14]. disjoint [GS04, KSP97, SHCD00, Tou02]. disk [AGL08, CH97, CN98, MQ97, San02, SSH97]. disk-directed [MQ97]. disks [BGV97, DBZ^{+}19, TTH05]. dispatch [RG09]. dispersed [NC02]. displacement [YLCT18]. displays [CKS03, CWDG07, LBH07]. dissection [Con90]. distance [BL93b, CR23, CC95, DCY^{+}22, ER95a, ER95b, FIMF99, GS04, HCR01, LHK^{+}96, ST02]. distance-insensitive [ST02]. distant [BCVC05]. Distributed [ABGC^{+}14, BP95, B^{+}01, CJLS14, DHL07, DIR97, GS10, HKK97a, HKK97b, HCJ03, HK90a, HC92b, ISH03, JON08, KYS23, Pap98, RCAP11, SSBT19, Tol02, TDW03, WE97, ABGR21, AGH^{+}94, AJ95, AC89, ARZ97, AKSS07, AFT23, AKK99, A^{+}01a, ADMV05, AB16, BBB^{+}14, BG91a, Bai88, BEY00, BGWW97, BKS06, BZ96, BCA08, BY21, BAR98, BZ99, BGP^{+}97, Bev89, BP00, B^{+}02, Bis89, BSK03, BSB^{+}22, BVWH14, BÁJG^{+}20, BBD^{+}12, BE97, BDK06, BCS15, Bro89, Bru91, BBB19, CMT04a, CE94, Cal96, CS00, CS93, CPdM^{+}97, CKRZ98, CTZ^{+}18, CPR^{+}18, CSH^{+}11, CDW95, CL05, CBP^{+}07, CSW01, CB09, CRT89, CJ95, DDdS02, DDdSL02, DD89, DL92, DR94, DR93, DLR94, DvdG92, DUG^{+}06, DZG94, EZ99, EPS98, Ekl04, ES06, FFLV96, Fie96, FBAB13, FMM10, GKS07, GZH^{+}23, GE11]. distributed [GLG06, GPC88, GGH^{+}22, GD90, GMF00, GHH^{+}03, Güs99, Hey91, Hil92, HJ97c, HMH^{+}13, IOH05, IT08, IKK15, IHM^{+}12, JLY18, JIC96, Kan97, KJA15, KG03, KA96, Kuz98, LPM11, Lee95, LC02, LH88, LWG06, LLD19, LSL02, LLFZ23, LS93, LGMdRA^{+}19, LF00, LMKH97, MMC97, MM00, MCC04, MW11, MW94, MHH97, MBC97, MOF04, ME92c, NOS92, NA08, ORM^{+}10, OIK21, PY00, PLY02, PK05, PRW88, RBB^{+}22, RMRN05, RGBCS09, Rot95, RV95, RBL97, Sab97, SYAU07, SAOKZ06, SSKÇ15, SNK06, SK97a, SBM^{+}98, Sch88b, SS94a, SSP^{+}98, SC03, SS01, SB95, Sol88, Son92, Son94, SB92a, Sun97a, SW03, SG99, TX00, TS88, TMCC02, TB07, Trä95, Trä18, Van02, VV07, WW95, Wal94a, Wal94b, WC94, WZL09, WYT^{+}16, WWZ^{+}18, WME^{+}95, WHW^{+}11, Wil88, WvR16, WP19, YX07, YPG03, YHZ21]. distributed [YLT^{+}23, YSBM97, ZLT19, ZM88a, ZEC^{+}17, ZGL^{+}19, dGR95, di 97]. distributed-heterogeneous [BCS15]. Distributed-memory [Tol02, ABGR21, AB16, BZ99, CTZ^{+}18, CPR^{+}18, CJ95, FFLV96, Fie96, FBAB13, GGH^{+}22, GD90, IKK15, LF00, MW94, Rot95, SNK06, Sun97a, Wil88]. distributed-shared [ADMV05]. Distributed-sum [TDW03]. distributed-termination [Son92]. distributing [TY91b]. Distribution [PMMAM10, BV04, CAA98, CGM05, DTV21, GWC^{+}99, Haa98, HT98, Hwa04, LR07, LC02, PB16]. distributions [RvG05, RZ95]. disturbed [MMS09]. Divide [Arb92, Bon91, Meh93, RM91, ER95a, ER95b, GTD18, Gor97, HvNJB12, LPAZ97, LSF^{+}15]. Divide-and-conquer [RM91, ER95a, ER95b, Gor97, HvNJB12, LSF^{+}15]. Divisible [BDM99, BLR03, BD10, BD95, DG99]. division [NSF^{+}22]. DLP [HXW^{+}13]. DMR [IMQO^{+}18, XXQ^{+}15]. DNA [BMCA98, TS09]. DNAml [CDZ^{+}98]. DNNs [QWD^{+}21a, QWD^{+}21b]. DNS [KRS^{+}21, Xu07]. DOACROSS [LK00]. docking [SVTSM^{+}22, VBS^{+}15]. document [PW87]. dog [LK14]. Domain [CMV^{+}06, DRST03, GW87, Hat06, YLCT18, AIV95, AW13, BG91a, BMP15, BZ96, BCS15, CC97a, CS95, DWB99, EEH^{+}19, FRC^{+}95, GHP10, HAK^{+}21, HJ97a, HJ97c, HZ93, HSFS14, Kon11, PAF^{+}97, PRV87, Ram07, SMC16, SGP22, Ter13, WA03, WAT20]. domains [LSS88, RSSS94]. dominance [Kim90]. dominant [Sun95, Zha91]. dominated [GG10]. dominating [GHC04]. dominating-set [GHC04]. dominators [Hor93a]. Doppler [TB03]. Doppler-bearing [TB03]. dot [YORO08]. DotDFS [PGG11]. double [BE92, LE91, SNO99, WRS12]. double-bordering [LE91]. double-precision [WRS12]. doubling [RBB17]. downdating [BHM^{+}95, YK08]. Downloading [PS91]. drawing [ME92b]. drift [BP95]. driven [Evr01, FSG19, JMA^{+}13, KRN^{+}11, Lin93, MSE19, MBK12, Tre85, WRS12, ZLD15, dNdRRL^{+}21]. drives [XCR17]. Driving [LBH07]. drug [PB98]. DSDADI [LRH97]. DSM [KKH04, Lef97]. DSP [KG03]. Dual [DK04, HC05, NYFK06]. Dual-level [DK04, HC05]. DuctTeip [ZLT19]. Dunn [BHB21]. Duplicate [TFS91]. duplication [Ara03, HK02a, KK03, Man01b]. duplications [LC89]. during [CRD02, FC05]. DVFS [GIRT19, TYLL22]. DVFS-aware [GIRT19]. DynaGrid [BK07]. Dynamic [BM11, BFL^{+}01, BOV02, CJLS14, DWH^{+}08, ECG93, EB93, HL97, Han85, KS05, Koh95, KAM^{+}20, LHS13, LOST16, TRSC^{+}19, AC03, ACGL04, ASH92, BM01a, BCPB05, BGG^{+}99, BO03, BNS^{+}07, BSK03, Bog92, BDK06, BK07, CLS93, CZTS99, CRGR^{+}13, CC97a, CCJ90, CF04, CNK93, CGM05, DCG90, DHL07, DIR97, EG95, FKPB15, FG96, GLS01, GGA19, HM02, Hor93b, HB99, JSC97, JDH96, KK11, LPNV20, LZCT15, LP92, MSMC15a, MMO07, MCP^{+}14, M^{+}00, MAD^{+}16, NC97, OKF16, OBG00, OID^{+}12, PAG^{+}05, PSF^{+}15, RCAP11, RMACG10, RGBCS09, RR14, SG02a, SSLK00, SYY^{+}22, TSJB00, TE92, VR95, WYT^{+}16, Wol89, WH94, ZCPT00]. dynamical [GLS99, KLN90]. dynamically [Kao08, PCLM18, SSS99, ZSI02]. Dynamicity [VV07]. Dynamics [ER18, AFPG12, BL94a, BCA08, FP98, Gen84, GCC19, GIF^{+}10, GVH08, HC05, Kas84, Ken90, KHN01, LBL95, LVC16, LBWR90, LHZ^{+}20, LR89, MSOCG^{+}16, MO99, NC97, N^{+}00, PJV00, SKE^{+}22, SSS92, STKA96, TJJ93, XRW^{+}19, Zom95].

E2SC [AMN16]. ear [IR93]. Early [CC99b, CH97, XH96, HHK^{+}19, Oed92]. earth [AC03, ABG^{+}03, SCM^{+}98, Sod02, CFN03, HUYK04, IUY^{+}04, Nak05, Sat04, SZ04, YS04]. EARTH-MANNA [Sod02].

Earthquake [BKL^{+}19, WHL19, BKS19, BBL^{+}19, HZY^{+}19, LYC^{+}19]. earthquakes [FC05]. ease [TKI85]. easily [Cho92]. easily-diagnosable [Cho92]. easy [GHS02, MK97]. easy-to-use [MK97]. echocardiography [CDM06]. ECMWF [Kau94, GJS93b]. economics [KNR00]. Eddy [WH15, SML^{+}14]. Eden [HHOMR06]. edge [HC04, HWW92, MLX07, MDC^{+}08, Rag97, TT00, YZB^{+}23, ZG16]. edge-fault-tolerant [MLX07]. edge-preserving [MDC^{+}08]. Edges [Tou02, Far94, Hsi06, KRS^{+}21, LS97]. Edges-disjoint [Tou02]. edit [DCY^{+}22]. Editor [RC14, SYP13, GT19, VBC19]. Editorial [AMN16, Ano18e, Ano18f, Ben24, DF95a, Hol10, Hol17, Jou02, Jou04, KP12, KC06, MV11, MS23, RST05, PS20, SAGV21, TS02, UZ02, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano06, Ano09b, Ano10b, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g]. Editorial [Ano16h, Ano16i, Ano16j, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano18a, Ano18b, Ano18c, Ano18d, Ano18g, Ano18h, Ano18i, Ano18j, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano22g, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23a, Ano23b, Ano23c, Ano23d, Ano24a, Ano24b]. Editors [BH14]. Edwards [RW96]. EEGs [EG92a]. EF5 [OWW16]. Effect [AB93, Man01b, HEB96, LFS^{+}19, Wor92, YMG14]. effective [BC19b, CWDG07, EJLC00, HC94, JH97, RG09, YNY^{+}17]. effectiveness [PB23, SC92]. Effectivity [Str87]. Effects [Axe86, MCB16, MBC97, PW89]. efficacy [CBP^{+}07, GScFM13]. efficiencies [DZM^{+}13, FRC^{+}95]. Efficiency [Bro89, AQO18, CHZ^{+}19, CBP^{+}07, CMM^{+}88a, FLW87, FLYL21, GIRT19, HZ93, KKK^{+}18, LT19, LS93, Ori10, Par86, SP93, SYY^{+}22, ZCT^{+}20, tV96]. Efficient [ABB^{+}21, AAS13, AAB^{+}06, ADTV01, BO91, BFVRC14, CAS09, CS93, CC00c, CCRR91, DCY^{+}22, DFM99, ED04, FK98, FCZY17, FCS^{+}19, GL99, HT00, HA05, Hee97, HXW^{+}13, IR93, JW09, JC94, JWS13, KD13, Lan99, ML23, Mar14, MKK03, MS04b, MSW98, MNP93, MO99, NSM23, OV06, PJW^{+}22, RR07, RGBC11, RSC^{+}19, SCCPM14, SMT^{+}04, SV94, SHH^{+}97, SR97, TNZM20, TPK^{+}13, YK06, YD01, ZGL^{+}19, ZM88b, Ale91, AAS20, AK89b, ABMN02, AFY^{+}16, AG00, BBB^{+}14, BFG^{+}07, Ber89, Bev89, BSD11, CGG01, CC02b, CF91, CC95, Che98a, CST02, CS19, CP08, CNK93, CG89a, CT88, CGM^{+}92, DBZ^{+}19, DWS^{+}23, EAR93, EL95, FLW86, FSCL06, FWL03, FMA17, FLYL16, FG89, FA11, FBG^{+}12, GLG06, GG10, GHS02, GM07, GS96, HW02, HRR08a, HCL05, HGM20, HKR^{+}10, Hwa04, IB01, KK01]. efficient [KA19, KC97, LMJC96, LSC^{+}15, LCLA19, LWJ^{+}17, MCY^{+}24, MS98, MGCB^{+}10, MSMC15b, MMT07, MMP^{+}21, MS20, NOS92, Not95, OW91, PLY02, PRS^{+}14, RH99, RBM^{+}16, Sab97, SKN04, SOH94, SM91, SWCBQ19, SP97, She95, SWYM17, SLG^{+}22, SGP22, SLH^{+}13b, TKC^{+}14, TX00, WYX^{+}22, WE92, WL98, YCS07, YLW^{+}13, YL19, YZB^{+}23, Yua94, ZLT19, ZHC^{+}23]. efficiently [KLP11, WW92]. effort [WPS^{+}08]. EFFT [AV15]. eigenproblems [ML20]. eigensolution [Pin91]. eigensolver [ABG^{+}06, BPS01, IS18, KKHY06a, PB11]. eigensolvers [DR18, HRT07, KMK^{+}19, YTSI13]. eigenvalue [ALH^{+}14, Arb92, ABB^{+}11, AHW14, Bol86, CRS88, CJ95, DvdG92, EM90, Gut88, Kal90, Kal92, KK01, Lan99, LQ92, MV99, Nat90, PMV^{+}20, RLH19, RKMJ10, RV95, SK19, SMK20, THM^{+}95, ZM94]. eigenvalue-eigenvector [EM90]. eigenvalues [BW92, GKT^{+}15, GD90, WC90]. eigenvector [EM90, SK19, SMK20]. eigenvectors [GD90, WC90]. eight [TTH95]. eight-neighbor [TTH95]. ejection [Reg01]. Elan [BHM94]. Elan-Elite [BHM94]. Elastic [WS12]. elasticity [dNdRRL^{+}21]. elasticity-driven [dNdRRL^{+}21]. elastohydrodynamic [ADTV01]. electric [SHRN98]. electroencephalograms [EG92a]. electromagnetic [IHM^{+}12]. electronic [ABB^{+}11, Ber00, MP95, SKE^{+}22, SLS^{+}21]. Electrostatic [RD07, HSS09]. Element [YR18, AKK99, BFL^{+}01, BM09, DL92, Dav88, DDF^{+}10, EZ99, EL95, ETV91, FKM^{+}22, FA11, FR95, FHMS93, GAR15, HJ97c, Kra84, LLP00, LR99, MW11, Nak05, NP05, PZE94, RWT97, SHH^{+}97, Sil88, SPMB23, TLC23, TA14, Wai88, YNO97, YLCT18]. element-by-element [ETV91]. elementary [de 87]. elements [GPSK09, LL88, SM88]. Eliminating [DSG17]. Elimination [BKT91, Gir02, ADFQO19, BR90, CG87, CMRT88, CTT89, DGP^{+}16, GT92, KKB92, Leu89, Liu89, MKS91, MR89, Mel87, MMS90, PW84, Ris90, RDB^{+}90]. Elite [BHM94]. elliptic [Eva91, GS89, GW87, HJ97a, Mil01, Pet97, Pir93, Won88]. ELLR [VFG12]. ELLR-T [VFG12]. ELPA [KMK^{+}19]. EM-4 [OKSY92, SKY93, SKS^{+}95]. EM-ML [GH98]. embedded [ADK22, DCD97, HK91, JLLC22, JON08, JMA^{+}13, MCB16, OWW16, RMACG10, SVC07, SSO^{+}14, YLW^{+}13]. Embedding [KLL^{+}09, TT00, Fu03, Hsi06, KF95, KL00]. embeddings [GH92]. emergent [CBP^{+}07]. Emerging [ONB11, BOV09, MII^{+}11, WOV^{+}09]. emission [Möl99]. Empirical [LD98, MJ99, PCC19, BHK89, BGG^{+}99, Dod90, WPD01]. employing [LNLK13]. Emu [YHE^{+}19]. emulate [FGP23]. Emulating [Wel89]. emulation [Bar97]. enabled [BWT^{+}08, DS13, DG15, DUG^{+}06, DMG^{+}04, GM04a, LNA06]. Enabling [APBcF16, BGWR21, HCT16, ZEC^{+}17, FBAB13]. enclosing [JCL92]. encoder [BFG^{+}07]. encoding [AALK01, PGBF^{+}07]. encountered [Hir86]. End [BVC16, BVC19, BDO17, CBV13, AAV08, MK12, PSF^{+}15, SVC19]. end-to-end [PSF^{+}15]. Energy [CHQORS18a, MSK14, SLG^{+}22, AQO18, AFY^{+}16, FLYL21, GPSK09, GIRT19, KKK^{+}18, LyHW^{+}22, LKGD16, LT19, MKK^{+}19, NB12, OKF16, RMS^{+}18b, SLS^{+}21, SWCBQ19, SYY^{+}22, TKC^{+}14, dNdRRL^{+}21]. energy-constrained [MKK^{+}19]. Energy-efficient [SLG^{+}22, AFY^{+}16]. engine [BBD^{+}12, CLL99, CRS88, DCR^{+}16, PR90, SML^{+}14, Wal01, Yaş01, BAMK07, BAK09, KR10, PMMAM10, SKST08]. engineering [DMG^{+}04, GSZ88, KSM^{+}94, Pap98, RMS^{+}18a, YPG03]. engines [DBI^{+}17, GGA19, HXW^{+}13, MGCB^{+}10, SYAU07, TFS91]. enhance [GHP10, RBB^{+}22]. Enhanced [NLP^{+}15, FP92, GSMY^{+}07, HRT07, HXW^{+}13, TSCS14]. enhancement [BDK06]. Enhancing [HMS^{+}19, MSMC15a, MQ97]. Enormous [AV15]. entrainment [KRS^{+}21]. entropy [SPW^{+}15]. enumeration [ME84, SDv98, YL19]. Environment [FTY^{+}20, AS86, AKK99, BDPV99, BCC^{+}97a, Bro87, Bro88a, BGRD00, CDR^{+}95, CDM06, CLA15, CDJ08, DS87, ES06, FMM^{+}02, HST05, IJM^{+}05, Jor86, KCD^{+}97, KHC92, KGV97, KSS06, MKK^{+}19, MSW91, MOF04, MKL^{+}88, NL91, NJ02, NRN00, PW87, PCC19, Pry97, RSSS94, SSP^{+}98, ST21, SWC99, TSCS14, Van02, ZLM98, vv91]. Environments [HKK97a, Jou97, A^{+}02, AKSS07, AMMV98, B^{+}02, BSB^{+}22, BGK^{+}98, CS93, CHBS18, CSW01, DDdS02, DDdSL02, DCG^{+}07, DF98, DT97, DBI^{+}17, EHR^{+}98, JBWE14, LGMdRA^{+}19, MSP93, McB94a, NOG^{+}22, TRSHB04, WA03, YCS07]. enzyme [Yan04]. EPEX [DGNP88]. EPEX/FORTRAN [DGNP88]. epidemiological [MSMC15b]. EpiGraph [MSMC15b]. EPS [MTPE90, de 87]. EPYC [BBL^{+}19]. Equalizing [MD04]. equation [Bjø87, BCS15, CMM^{+}88a, CKLM14, DWB99, FBMV88, HJ97b, HWB92, LRH97, MS04b, RSK99, SHvAP13, Sch92, SP93]. equational [GJMMS97, OTI^{+}89]. equations [ADF93, AAHF97, AIIV98, BD89a, BG91b, Bas94, BE92, BE93b, BBQOQO00, BEK^{+}11, BWW89, Bon91, BP02, Cap88, DH84, EM87, ER89, EL90, ED91, GR89, GS89, GYH94, GV01, Gol88, GW87, HSS07, HM89a, HK92b, Jou89, KR97, Koç97, KS16, LWS02, LSS88, Lop93, MQ89, ME90c, Mei85, Mil01, Oli96, OJ90, PATC99, Pis92, RF90, Reu88, RSSS94, RMS^{+}18b, RO88, SHPA05, SN88, SS88, SSL19, Sch87, SK87, Ter10, Ter13, Won88]. equilibrium [DF00, XLY^{+}20]. equivalent [TH89]. era [CP18, SR14]. EREW [HK06, She97]. Erlangen [DSSD18]. Erratum [SVS02, Wal94b, WBD99a]. Error [Pis92, Stp93, FGS97, HISS92, XXQ^{+}15, ZLJ93]. error-detecting [ZLJ93]. errors [Arv92, CHR18, DPFT19, TYSF13]. eruptions [OCE^{+}07]. ESA [ZHC^{+}23]. Esprit [Hey94]. essentials [ST88]. estimates [MW88, SH91]. estimating [YK06]. Estimation [PMMAM10, HDH97, KI05, Kol94, Mal02, MS99a, TDW^{+}04, VJ12]. estimations [KM88]. estimator [GGFF93, Güs99, YK08]. ETA [MRJ89]. Ethernet [Gog11]. Euclidean [CC95, Gur88, LHK^{+}96]. EULAG [WSR14]. Euler [Ger94, Cap88, DLR94, DR95, PEO16, di 97]. Euler/Navier [DLR94]. Eulerian [DAD11]. EuroBen [van91]. EuroMPI [Ano18k, Ano22m, GT19, TH20]. EuroMPI/USA [Ano22m, GT19]. Europe [BRS^{+}17, Duf84]. European [Kau94]. EuroPVM [MTW07]. EuroPVM/MPI [MTW07]. Evaluating [BLCR21a, BLCR21b, CRT89, FL21, KKK^{+}18, LXC^{+}18, RHM^{+}88, SSH97, Beb97, CH97, GScFM13, TG09]. Evaluation [BCVC05, DF00, FR95, IKK15, PEO16, SKH^{+}12, SWSG92, vWWM^{+}19, AE93, BAR98, BDH95, BMG07, CC96a, CCP98, CRSS99, DWS^{+}21, DBI^{+}17, FKK^{+}06, FB19, GV08, GICM18, HB90, HHK^{+}19, Her88, HHOMR06, HHR87, HK02b, HL13, JMLBL98, KK92, KSS98, KP93, LAS90, MP92, MS20, Obe85, Ori10, OW94, OW97, PB23, PCN10, PKR00, Rib84, SNS^{+}97, SAOKZ06, SPCB22, SFSV13, SCM^{+}98, SS04, Sol88, TYKA95, Y^{+}99]. evaluation-interpolation [KP93]. Eve [OW97]. even [BE93b, BR89, MP87]. Event [KRN^{+}11, BRWL09, DF98, FSG19, GJMMS97, MK12, PM89]. Event-driven [KRN^{+}11, FSG19]. events [Ano97, HHK^{+}19]. Everglades [HW02]. Evolution [MGSK88, SGDM94, AIV95, MGSK87, SSD^{+}94, YLE95, DST15]. Evolutionary [GM04a, BFVRC14, CMT04a, DS07, dTNOR^{+}04, RMACG10, YHZ21]. evolutive [MOF04]. Evolving [MCM01, PYZC11]. Exact [Bog92, BG21a, BG21b, BBQOQO00, DCY^{+}22, DR02, DGP^{+}16, FBAB13, KB20, LHK^{+}96, LG03, Obe85]. examine [LFS^{+}19]. example [CR90, DFRR91]. Exascale [ABB^{+}21, NDW^{+}19, PS20, CHBS18, CCG^{+}17, MAB^{+}21]. exceptions [LTG23]. exchange [CS89a, CC00b, HT98, Kas85, Kat01, MMM13, SBM^{+}22]. exclusion [TNZM20]. executable [OTI^{+}89]. executing [MKK^{+}19, SLH^{+}13b, WW92]. execution [ATA^{+}16, A^{+}01a, BPEL05, BV96, CCGG14, Eck90, Fer97, FM88, GW08, GB14, KD97, KI05, LGM09, Mah96, MFS^{+}19, MAD^{+}16, Psa02, RvG05, Reu99, TDW03, WW90, WL14, WvR16]. execution-time [KI05]. Existence [Rao97, HE88, KS87]. exogenous [Mal02]. expansion [LH02]. expected [CAHT17, CHR18]. experience [BC19b, FPT91, HKN89, KS99, MCC04, NHS^{+}95, Zha95]. Experiences [HL88, HLM84, KSM^{+}94, NOG^{+}22, BBB^{+}22, BWZ95, CW97, CH97, FLS^{+}21, Hol95, HB94, LBL95, kLH95, Nag88, NBB^{+}02, Pei88, SGDM94, STKA96, WRB97, XH96]. experiment [CC99b, Son92]. Experimental [DBI^{+}17, HISS92, CFS01, Lef97, MTPE90, PVP08, TFS91, Trä95]. experiments [DCG90, Kam87, YD97, HAJK01]. expert [CDDG93, Li89, TE92]. experts [EPMPU02]. Explicit [GVH08, BE95, BE89, BE92, EY90, EC91, KR97, ME90c, NA08, OJ90, SBP12, Thu90, Thu91]. Explicitly [SWC99, Vég18]. exploit [CIO^{+}17, RRP03]. Exploitation [JCE96, Lee97]. Exploiting [ABMQO11, ABB^{+}16, DL97, FMA17, Gao90, GW95, JAH^{+}18, Nag90, Zam99, DVGG98, HXW^{+}13, MR86, Rob00, THH^{+}05]. EXPLORER [LK00]. Exploring [CS19, DZLK20, GKS07, GSMY^{+}07, LGMdRA^{+}19, Pea19, PTK04, RMS^{+}18b, SKST08, BJ18, LC17]. explosive [OCE^{+}07, SSLK00]. Exponential [SSL19]. exponentials [ABH^{+}10]. Express [Ahm97, FK94]. Expressing [GJMM89, CWB92]. Expression [CR90, Bur04, HL99]. expression-rewriting [HL99]. expressions [SS87]. expressive [Trä12]. extended [LBH07]. extendible [ONOK13]. Extending [AAA16, BM02, BPP10, Ger15, JCC^{+}24]. Extensible [DAG^{+}09]. extension [CC98, JH97, MZ95, SS04, Son86]. Extensions [HVA^{+}16, Hoe12, Wit98, BCC^{+}97b, HP05, ZCBD22]. External [BBB^{+}94, CSW01, MRT93]. extracted [Kao08]. extraction [BDSV98, MS03a]. extrapolation [EM87, Eva89, HK92b, SY87]. Extreme [BKL^{+}19, GAA^{+}13, HMP^{+}16, MSZM14, RBP^{+}17]. extreme-scale [MSZM14]. extremes [Cla03].

F [Ano03k]. fabric [KRN^{+}11, WLZ^{+}23]. Face [GK17, SPW^{+}15]. facilitate [LMJC96]. facilitating [HLvHA11, WCM^{+}14]. facility [KKHY06b, KKHY06a, SK97b]. FACOM [TKI85]. factor [BO03, ESK88, God00, Tem88]. factorial [ALD01]. factorisation [HE88, KS87, YE95]. Factorization [CGH^{+}19, AB95, AD14, AC94b, ARR19, BETR17, BB21, CU04, CHQORS18b, CTW14, CG89b, Con89, DG95a, DSS86, DHW97, DFH^{+}13, DM02, GP90, GMBZ90, GHNL87, GLN89, GH90, Hen93, HRR08a, KKB92, LR07, LKHL03, Leu89, LC99a, Liu86, Mat95, Mil01, Ng93, OW90, Ort88, OR88, OHZ98, Rao97, RSD16, Stp92, ZE92, Zhu93, ZG88]. factorizations [DW99, MRBQO14, mMvdV01]. factors [AGL08]. Failure [SRS^{+}19, LHR^{+}19, FCZY17]. failure-prone [LHR^{+}19]. Fair [MCY^{+}24]. fairness [LFS14]. Falk [NSS15]. false [CAC^{+}09]. family [GDMS97]. fan [VP95]. fan-in [VP95]. farm [GV06]. farms [Bus93, FCZY17]. FAS [SS90]. Fast [AV15, AW13, BT88, EL16, Gol86, GHC04, HWL^{+}22, JK91, MKJC21a, MKJC21b, NMI^{+}24, RR97, RKMJ10, RBS10, SB94b, Wen95, BG02, BWW89, BNK15, Cha88, CTZ^{+}18, Dod90, DZG94, GS89, GR01, GF03, Heg96, HJ97b, Hol95, LO92, LCLL00, Pet97, Pir93, Pir96, SKH^{+}12, SPCB22, SWW99, Ter13, WYT^{+}16, AFS14]. faster [BR89, OWZ91, SCCP08]. fat [ACM^{+}15]. fat-tree [ACM^{+}15]. Fault [ASDOK01, Fu03, GAR15, Hsi06, HCTH05, KH97, LH04b, NFG^{+}13, SYAU07, SHCS86, XCS05, BS97, BHR09, CC02a, CS97a, Cho92, FBD01, FSKF06, Far96, Fu07, GLP98, GP98, HXCZ14, HG12, KL00, Leu93, Li90, Loh96, MLX07, ME89, PY00, QJ06, SS01, SR17, SW00, Tou02, Yua94, ZL04]. fault- [CS97a]. Fault-tolerance [SHCS86]. Fault-tolerant [ASDOK01, Fu03, GAR15, Hsi06, KH97, LH04b, XCS05, BHR09, CC02a, Cho92, Far96, Fu07, KL00, Loh96, PY00, QJ06, SR17, SW00, Yua94, ZL04]. faults [CC02a, CS97a, HC04, KLL^{+}09, Shi01]. faulty [EAR93, Far94, GP96, Hsi06, LH02]. FEAST [GKT^{+}15]. feature [WHW91]. features [BPJ22, Kao08, Oed92]. February [Ano23e, Ano24c]. Federated [BKK24]. Feedback [CZJS12, DTV21, MM95]. Feedback-controlled [CZJS12]. feedforward [Han98]. FEM [DPSW00, GSMY^{+}07]. FEniCS [TLC23]. fermions [EHN99, GLS99, KLN90]. FETI [KHHC20, RBM^{+}16]. FFT [AL88, AGGM90, Cal96, DWM^{+}01, GGL21, HP91, IB01, JK92, MKJC21a, MKJC21b, MKK03, ND95, Swa84, SSB^{+}91, Tak03, TTH95, Tem88]. FFTs [BP86, CN98, Swa87, SH01, SBO^{+}91, WBS06, AAZ96]. FGCS [UAR^{+}99]. Fibonacci [Mak94a, MS04a]. fidelity [DCG^{+}07]. FIDISOL [SS88]. field [Gup99, IHM^{+}12, Qua92, SMC16, WAT20]. fields [BCCS07, SHRN98]. Filaments [LF00]. file [BOS^{+}09, CPdM^{+}97, CJLS14, FLYL16, GKS07, LZCT15, MK97, MQ97, NK97, PCN10, PGG11, TDC19]. filestore [PR90]. fill [Ala89]. fill-in [Ala89]. filled [BM01a]. Filling [Kon11, GZ99]. filter [DOD23, Hot89]. filtered [RW98]. filtering [BDK06, CS89c, EA95, MSW98, TB03]. filters [EG94b, TGE92]. finance [KNR00, Zen99]. financial [BGRD00, PGK^{+}00, PŚ00]. find [SS89b]. Finding [Box92, CF90, GD90, LTG23, LS97, CKM93, FB91, HRR08a, HM89c, HHH92, HWW92, Jan06, MBC92, MS89, RVD10, SLY90, Shy90, ZM88b]. Fine [KKW14, PGGP19, PTS^{+}12, Ste13, XDZL10, BMP15, DFRC90, Gao90, HS89, Lil94, MG19, MQ97, ZGG92]. Fine-grain [PTS^{+}12, Gao90]. Fine-grained [KKW14, PGGP19, Ste13, XDZL10, DFRC90, Lil94, MG19, MQ97]. Finite [LL88, PZE94, SB96, AKK99, BFL^{+}01, BM09, DDF^{+}10, EZ99, EL95, FF92, FA11, FHMS93, GF89, GAR15, HJ97c, Kra84, LLP00, LR99, Lin93, MWI^{+}15, MW11, MSW98, Nak05, Ram07, RWT97, Riz85, SHH^{+}97, She96, Sil88, SPMB23, TW19, TLC23, TA14, Wai88, WME^{+}95, YNO97, YLCT18]. finite-buffered [Lin93]. finite-difference [GF89, MSW98, Ram07]. finite-element [DDF^{+}10, GAR15, Kra84, Nak05]. finite-volume [Riz85]. Fire [HW02]. First [GJS93a, CKS03, GV08, Kre13, ME90c, SlWC89, TO89]. first-fit [GV08]. first-party [Kre13]. Fischer [Sut97]. fit [GV08]. fitness [Hur93]. fitting [CH98, C^{+}01b, SM87]. five [GP85]. Fixed [PMS^{+}13, CHLS90, DDP90b, PER17, Wei93, ZRPI89]. fixed-size [CHLS90, DDP90b]. fixing [SS94b]. flash [XCR17, DAG^{+}09, CCP^{+}21]. Flash-X [CCP^{+}21]. fleet [BCG00]. flexible [DFRZ02, DBI^{+}17, EMV^{+}18, FHK^{+}11, PYZC11, SPMB23, KL00]. floating [Kam88, LTG23, LHS13, ZG16]. floating-point [Kam88, LTG23, LHS13, ZG16]. flow [BSH88, BG97, CW97, DF00, El 93, EG95, EG90, FMS^{+}06, GPLW17, GGFF93, HM01, JS13, JJAP^{+}20, KYLH01, LS93, Lun98, NA03, NP05, OCE^{+}07, PAF^{+}97, PCA01, PSF^{+}15, SB94a, SHH^{+}97, TW19, TO99, TCL92, TSJB00, WNES01, XT11, Xu07, MMT06, MAFC14]. Flow-Shop [MMT06]. flows [BM09, CLL99, DL04, Ema10, KT97, Lan96b, MT97, PMLT03, RWT97, SOS97, WH15, YD97]. fluid [ATL^{+}12, Gen84, GIF^{+}10, Kas84, LR89, MRRP11, PIG^{+}16, RWT97, YNO97]. fluid-particulate [ATL^{+}12]. fluid-structure [PIG^{+}16, RWT97]. FMI [DTV21]. FMI-CS-based [DTV21]. focus [GH98]. focus-of-attention [GH98]. fog [KYS23]. folding [BE87, CLSM98, EG92b, OCSBY01, Rob00]. followed [HBH^{+}16]. following [BG01]. footprint [WH15]. force [GVH08, RD07]. forecast [GJS93b]. forecasting [Sel95]. Forecasts [Kau94]. forefronts [BR84]. foreign [CC00b]. forest [DDP90b, HAK^{+}21]. forest-of-octrees-based [HAK^{+}21]. Foreword [Ano18k, BVC19, CHD09]. Fork [FC18]. form [BDK95, DvdG92, GL99, JW09, KK11, Kuz98, Lan96a, ML20, NW96, TRSC^{+}19]. formal [Cv98]. format [AAS20, KKU16, LB23, Mar14, SPMB23]. formation [Yan04]. forms [ACH^{+}18, DS05, Ort88, OR88, TND10]. formula [Eva89]. formulating [BEDdC16]. Formulation [CC96b, AB16, Bek95a, Bro86, MMC97, RWT97]. formulations [GKS98, SPK18]. forthcoming [Ano97]. FORTRAN [BRH90, BDS88, CK90, DGNP88, DS87, GSZ88, LMJC96, RRS88, Sne88, WMCU97, van86, BZ99, BCC^{+}97a, BCC^{+}97b, BC19b, CCL04, CWB92, CI98, FGBN19, KY98, KT00, MVZ98, MZ01, Num05, PFS^{+}04, SC19, ZBC94]. forward [CDM^{+}10, MRW^{+}08, TE92, YE94]. four [Che98a, DE09]. four-connected [Che98a]. four-dimensional [DE09]. Fourier [AV15, AW13, CT94, Cha88, DZG94, Heg96, HJ97b, RW98, SCCP08, Sur10]. FP [Lin91b]. FP-based [Lin91b]. FPGA [BCB02, GPW^{+}08, GVH08, KSS06, LLVM21a, LLVM21b, LTS^{+}24, OYS08, SDMS12, XDZL10]. FPGA-based [GVH08, LTS^{+}24]. FPGAs [HMS^{+}07, LLVM21a, LLVM21b, MC09, YPL12]. FPS [CFH89, Hoc89, Pop88]. FPS-164 [Hoc89]. Framework [Bro86, SLH^{+}13a, BG21a, BG21b, BCB02, BCCS07, BG98, BK07, CKRZ98, CLHL23, CCP^{+}21, DG05, EZ99, GAA^{+}13, GF13, HL99, JON08, JGT^{+}20, JMA^{+}13, KCG08, LLVM21a, LLVM21b, LPAZ97, MPZS13, MGT^{+}13, MW88, SSLK00, SR10, Vio04, WFW^{+}11, WH22, WS12, YZWcF14, SLC^{+}16]. frameworks [ACMT20, HCJ03, KG21, MAB^{+}02]. Fredholm [ADF93, BD89a, ER89]. Free [BPK12, FC18, Wil88, Ann89, CC96b, CL92, CS97a, GNY00, HGS10, JC94, LY93, Wan08]. Free-Lagrange [Wil88]. FreeTIV [PAD02]. frequency [ACMT20, CHQORS18a]. frequent [Ski02, XZYQ21]. Friedrich [DSSD18]. Friedrich-Alexander [DSSD18]. friendliness [TLC23]. Front [Ano02f, HK91, KLP11]. frontal [EGTD99]. Frontiers [Ano15a]. FT [HXCZ14]. FTL [XCR17]. FUJITSU [HL88, Heg96, IU87, NRN00, R^{+}00, SE98]. Full [XRW^{+}19, CDV08, HM86, SST09]. full-multigrid [HM86]. Full-neighbor-list [XRW^{+}19]. fully [Gen87, HHH92, Lev90, LOKM99, RV95, TKK^{+}05, ZLM98]. fully-pipelined [HHH92]. Function [Kog85, Fis90, GICM18, MSB91, TH89, YLE95]. Function-based [Kog85]. functional [BHS99, Cha99, EZ99, Mah96, PTGF20, WYBdJ^{+}21]. functionally [Kas85]. functions [Bus93, CDGM96, Hur93, KS91b]. Fundamental [THK^{+}99, CKM93]. fundamental-cycle [CKM93]. Further [Tem89]. fused [HBC19]. fusion [Dar00, GZH^{+}23, TBM^{+}16b]. future [Cla03, FK94, MSE19, MVZ98, SR14, UAR^{+}99]. Fuzzy [DS93, DZ90, FGS97, TMCC02, WC15, ZRPI89]. FX [WSL88]. FX/8 [WSL88].

Gaining [AGH^{+}94]. galactic [SRK^{+}21]. Galerkin [HBC19]. Galley [NK97]. GAMMA [CC00c]. ganglia [MCC04]. garbage [Bev89]. gas [MT97, SR97]. gather [Trä18]. gating [N^{+}00]. Gauge [GAW96, BP89, Hio96, SS94b]. Gauss [ADFQO19, BP02, CTT89, Mel87]. GAUSSIAN [GYL00, BKT91, BR90, CH98, CG87, CMRT88, DGP^{+}16, Leu89, MR89, MMS90, PW84, Ris90, RDB^{+}90, WYBdJ^{+}21]. GCG [EL90]. gene [Bur04, CFS01, BJV^{+}16, MV17, WZV^{+}16]. Gene/Q [BJV^{+}16, MV17, WZV^{+}16]. GeneaLog [PGGP19]. General [Gup99, MV99, ADLL03, AHBLR12, BE95, BCB02, CHH^{+}01, CPR^{+}18, DZG94, FG96, Hac89, JKH95, KKT00, LPNJR96, LSS88, LK00, MT90, MHT06, MP11, MSW98, Mun99, PJW^{+}22, RSSS94, Sil88, WYT^{+}16, WME^{+}95, XKL^{+}22, YK06, dV94]. general-purpose [Hac89]. Generalisation [RBB17]. Generalised [CKM94, EL90]. Generalized [LWS02, NC02, Bol86, CC01, HWYL89, KS16, KK94, ML20, Nat90, NSS15, RLH19, RV95, ZM94, LW15]. generate [ME93b, LEH14]. Generate-Test-Aggregate [LEH14]. generated [DLPS92]. Generating [HvNJB12, IJM^{+}05, LC02, TX00, TH89, EK92, GLG06, Lec89, LT94, TL90, TL94]. Generation [BDO17, AND05, AKNS91, BCGS93, BGWW97, BWV^{+}17, BDSV98, BRS^{+}08, CCS94, EAR93, EYP^{+}20, EJLC97, Glo95a, GDAK06, Hoe12, Hwa04, JIC96, KPL^{+}12, Kra84, LMJC96, LLW^{+}24, Lin90, MS99a, PVBR23, SR97, SLC^{+}16, Sun97b, TLC23, YK87]. generations [HL99]. Generative [DF98, MAS06]. generator [APG92, Glo95b]. generators [DP90a, Deá90, DLS09, DWS^{+}23, Mak94a, MM95, Mas98, MC04, MS04a, SMC03, dP90b]. generic [BBD^{+}12, BAS13, Gog11, HLvHA11]. Genesis [Hey91, GHS02, Kau94, McB94b, STT94, Hey94]. Genetic [WSB90, AD96, ALNT04, CA92, DM04, Ekl04, Hur93, IT08, LW15, MG95, Moo04, Müh90, MSB91, ORM^{+}10, OIK21, OCPT97, POHS14, SBM^{+}98, SVS01, SVS02, Wal01, WWR05, WWT01, ML00]. genetic-based [POHS14]. genlocking [CWDG07]. GenoMap [SBM^{+}98]. genome [CS19]. Genomic [AGV^{+}17, GGL^{+}05, LH04a]. genomics [JNWJ18, JNC^{+}19]. genotyping [SBM^{+}98]. Geocomputation [Cla03]. geographic [HCJ03]. geographical [CMM03]. geographically [SSP^{+}98]. geological [DS07, DDF^{+}10]. Geometric [GO88, AK89b, BWV^{+}17, BBB19, KS95, LNS03]. Geometrical [FM15]. geometries [CLL99, ZEC^{+}17]. geometry [BE87, SM90, Sýk84]. geophysical [BPC21a, BPC21b]. Germain [MC04]. Germany [BRS^{+}17]. GESIMA [AFG^{+}97]. GF [PW84]. GF11 [KBD93, WZ90]. GFLOP [Taf01]. GFLOP/s [Taf01]. gigabit [EHR^{+}98]. Ginkgo [CTA22]. Givens [HK06, Kon02, VKS^{+}15]. Glacier [CCG^{+}17]. Global [CMPM^{+}15, HHJP16, NZHY11, SS92, TCT00, BK94, BE88, BOS^{+}09, BT01, ES89a, FWL03, GV14, HRT07, HM97, HCL05, Kat03, MS20, PEO16, PODd16, RE98, SWR^{+}13]. GMD [BRH90]. GMRES [Li97, VvW98]. Goals [Gil94a]. gossiping [ST02]. GPGPU [MVdLvN16, TYSF13, TS09]. GPGPUs [GIRT19, DWX^{+}12]. GPU [ABB^{+}21, ABGR21, ADEQO19, ACM20, BGWR21, BWV^{+}17, BG21a, BG21b, BEK^{+}11, BSB^{+}22, BGB10, CYDJ21, CGK^{+}16, CR23, CTW14, CFW^{+}22, CKLM14, DN16, DTR18, DG15, DZLK20, DWL^{+}12, ESV10, FFZ^{+}18, FMA17, FLYL21, FHK^{+}11, FHK^{+}15, FKM^{+}22, GZHX17, GMMT16, GSMY^{+}07, GK19, HSR^{+}14, HJS23, HVA^{+}16, HWL^{+}22, HF20, JS13, JHD^{+}22, KPL^{+}12, KS16, LLW^{+}15, LGD^{+}15, LKP24, MCY^{+}24, MAB^{+}21, MSV15, MAA^{+}21, NDN20, OKTR13, OKF16, PTS^{+}12, RBS10, SHvAP13, SPW^{+}15, SCL^{+}23, SY21, SGP22, SVTSM^{+}22, SSV^{+}16, SMW^{+}05, SIH14, SYY^{+}22, ŚDJ^{+}22, TPK^{+}13, TND10, TDB10, TS21, TLS16, TLC23, VKS^{+}15, WZCG14, WL14, WSR14, XT11, XLY^{+}20, YMYH24, YWP11, ZAA^{+}21]. GPU-accelerated [BWV^{+}17, DN16, FKM^{+}22, MAA^{+}21, NDN20, SSV^{+}16, ŚDJ^{+}22, TLC23, ZAA^{+}21]. GPU-based [GMMT16, HF20, LKP24, MAB^{+}21, RBS10, SIH14, TND10]. GPU-controlled [OKF16]. GPU-enabled [DG15]. GPU-enhanced [GSMY^{+}07]. GPU-parallel [HWL^{+}22, SGP22]. GPUs [AÖ22, AFS14, AD14, AGD^{+}17, BDP20, BTLK18, BC19b, CCS^{+}18, DS13, GCH21, GTD18, GS22, HLP10, OK22, OIH10, RSD16, RA20, SPCB22, VFG12, VJ12, WC15, YPL12, AHBLR12]. Gradient [VJ12, BM90, CG89a, CGMM99, Daḡ07, El 93, ETV91, EL90, EG94a, Fis90, GZHX17, GV14, Han98, HGLR07, KL90b, Kru97, Meu87, MZ95, MA22, O’L87, RR89, Sea86, Str08, TGE92]. gradient-like [RR89]. gradients [van86]. Graham [Chr00b]. grain [BSH88, Gao90, GW95, HS89, KY98, PTS^{+}12]. grained [BBC^{+}11, CDBL08, DG90, DFRC90, HHK95, KKW14, Lil94, MAFC14, MG19, MQ97, OHZ98, PGGP19, Ste13, TTO^{+}22, XDZL10, ZGG92, dV94]. Gram [BO91, BLT^{+}22, OW90, Str08, WC91]. granularity [Cal85, Mah96, MB88, MRSB94, SCL^{+}23, YMY^{+}19]. GRAPE [Sug99]. Graph [BHK00, BOG15, ÇFG^{+}12, Dod89, HK00, HWYL89, MMS09, NSH^{+}21, NTHY22, Pel97, SKC90, ABGR21, BM08, BVP^{+}89, BSB^{+}22, BFH23, BGL^{+}88, CDBL08, CLA15, CL92, CWC00, CP08, CKM93, DCY^{+}22, DCR^{+}16, DDP90a, Far96, GPC88, HLP10, HHH92, Hwa02, IR93, JDH96, KH95, LS98, LPH00, LC89, Lin01, LSF^{+}15, ME84, Mar92, Miš98, MPD00, OTI^{+}89, PD11, PER17, QWD^{+}21a, QWD^{+}21b, RM91, RKS92, SS89b, Shy90, TW19, ZY16, Zer90, ZK92, Zho93, ZM88b]. graph-based [BFH23, DCR^{+}16]. graph-partitioning [MPD00]. graph-theoretic [BM08]. graphical [DRJSW97, KCD^{+}97]. Graphics [CM95, ADFQO19, CJ97, CWDG07, DOD23, GM07, HSS09, MIA^{+}07, MAM^{+}09, RSB05, RB03, SRH07, SA13, SCL^{+}23, Sur10]. graphs [AB16, BDK98, BT99, CHR18, Che96, DT96, DUG^{+}06, EG95, FKPB15, Fei91, FJPA17, Fu07, Gor97, Got89, Hwa02, KKT00, KF95, Kru97, LJD93, LZ00, MT91, MT95, Man01b, YL19, WCKM11]. GRASP [ABR03, RR07]. gray [LCLL00, Cha88]. greater [FLW87, Par86]. Greedy [Kon02]. GREMLINS [CDJ08]. Grid [DMG^{+}04, JBC^{+}04, Laf02, PGG11, Sol88, A^{+}02, BDK98, BCMG^{+}07, BFVRC14, BDK06, BRS^{+}08, CS95, CDK^{+}03, CW06, CGQ93, CL05, CDJ08, DR95, DM04, DUG^{+}06, EK98, LHR^{+}19, GGV04, GM04a, GM04b, HST05, KTN^{+}14, LLPV06, LG09, MMO07, MMT06, MMT07, MST94, PTK04, SSS^{+}05, SG16, ST88, ŚDJ^{+}22, TLS^{+}08, WVHR16, YCS07, YBM05, ZMMW90, DWH^{+}08, FMM^{+}02, LNA06, TKK^{+}05, TRSHB04, WA03]. Grid-based [PGG11, BDK06, MMT07]. Grid-enabled [DMG^{+}04, DUG^{+}06, GM04a, LNA06]. grid-shaped [LHR^{+}19]. gridding [WYX^{+}22]. grids [BCCS07, BG07, CRGR^{+}13, CdCM07, Cér07, ČSR05, DA06, DL04, DST01, GCP11, GV06, HJ97c, JBC^{+}07, JCWP07, LM06, LGM09, NP05, Pop88, PMLT03, Ran00, Sim91, Thu91, MTV08, MHL06, Ram07]. GridSAT [CW06]. Ground [LG01, FC05]. group [BE89, BE92, EC91, KR97, Kan97, ME90c, NA08, OA08, van91]. group-ware [Kan97]. grouping [GTW^{+}20]. groups [EB94, LJD93]. growth [WS04]. GS [XXQ^{+}15]. GS-DMR [XXQ^{+}15]. guarantee [FHL87]. guards [YMY^{+}19]. Guest [DF95a, PS20, RC14, SAGV21, SYP13, TS02, UZ02, BH14, GT19, VBC19]. GUI [VGS14]. GUI-awareness [VGS14]. guided [CYDJ21, GBH98, LG21, NB12]. GWAS [PFTB15]. gyration [Rob00]. Gyrokinetic [MII^{+}11, MWI^{+}15].

Hadamard [AFS14]. Hadoop

[CJLS14, RBB^{+}22]. half [BBDN21]. halo [SBM^{+}22]. Hamiltonian [Che96, CST02, KL00]. hamiltonicity [Fu07, HCTH05]. hand [CLA14, KMB^{+}18, Ter10, WP19]. handling [BE97, DG94, Fer97]. Hardware [GJMM18, HUYK04, Sab97, BCB02, BBDN21, BNK15, CCG^{+}17, CT94, Gog11, GM07, HBC19, IZS^{+}20, LTS^{+}24, NMMG13, PMS^{+}13, PGAA16, SKH^{+}12, NRN00, Oya99]. hardware-software [BNK15]. hardware/software [NMMG13]. Hari [NSS15]. harmonic [IHM^{+}12]. harness [BE95, FBD01]. Harry [Ano03k]. hash [QP16]. hashing [GZ99]. Haswell [PGAA16]. Hausdorff [BL93b]. Hazel [BRS^{+}17]. HBPFP [XZYQ21]. HBPFP-DC [XZYQ21]. HC [BNK15]. HC-1 [BNK15]. HCW [MSS19]. HDF5 [LyHW^{+}22]. head [SHRN98, Wor97]. headed [Laf02]. health [PTK04]. heap [MNS87]. heaps [DP00, LL90]. heating [TTT^{+}92]. heavyweight [RGdS^{+}13]. height [Ver00]. helix [Meg90b]. Helmholtz [HJ97b]. HEP [DH84, HLM84, Hir86]. Hermite [BP02, JW09]. Hermitian [WP19]. Hertz [ADTV01]. Hessenberg [BDK95, KK11, SK91, SNO99, TND10]. Hessian [Fis90]. Heterogeneity [Erc88, FMA17, KKK^{+}18, LR07, PRV08, SS94a]. Heterogeneous [CP17, KLR05, LB23, LFS14, MOW96, Sun97b, SG99, AAS13, ALNT04, AAO13, AFK01, AMMV98, BDPV99, BCVC05, BM11, BLRR02, BLR03, BHR09, BDNP11, BRRV11, BSD11, BDRV99, BCS15, CFF17, CLA15, CP18, CYXL18, CC99a, CGM05, FLS^{+}21, FLYL21, FHK^{+}11, FHK^{+}15, FGP23, GGH^{+}22, HJ05, HJS23, JAA06, JCC^{+}24, Kha12, KI05, LTG23, Lan90, Las02, LR04, LDK16, LV15, LXD^{+}23, LBGO23, MKK^{+}19, MGT^{+}13, OS21, Pea19, PB16, PVP08, PP98, PCC19, QJ06, RCAP11, SBP12, SSO^{+}14, SLG^{+}22, ST21, SS04, SBM^{+}22, YA24, ZJDW18]. Heterogeneous-aware [LFS14]. heterogeneously [JLLC22]. HeteroPar [MSS19]. Heuristic [WS94, AG00, CNK93, DT96, GLS01, GPS03, NDN20, ST21, TN13]. heuristic-based [ST21]. heuristical [Zer90, ZK92]. Heuristics [BD10, BDK96, CGM05, ALT04, ABMT99, ACGL04, BMA06, CA92, Liu98, LHK15, MMT06, RR07]. HEVC [JLLC22]. hexadecimal [Tak18]. hidden [DVP90]. Hiding [GV14, CV17, Coo19, MS03b]. Hierarchical [ATT92, BV96, DFH^{+}13, MSZM14, TY91b, vOS^{+}98, AFY^{+}16, ADMV05, BETR17, BTLK18, CLA15, CPR^{+}18, DL05, EPMPU02, GJMM18, HA05, Kol94, KJ07, MTV08, Ols95, PW22, RG92, SHN12, SB92a, WZS^{+}14, YMJT10]. hierarchically [GAR15]. hierarchies [KLP11]. hierarchy [CG89a, FDTZ04]. High [Ano16k, BAMK07, BA08, BVC16, BVC19, BY21, BDO17, CIS99, CLL99, CCS03, CBV13, CDK^{+}03, CMM03, CV02, DDF^{+}10, EHF^{+}97, cFM07, GJMMS97, Ger98, GS22, Gog11, GYL00, GLDS96, GKKS01, HVA^{+}16, HR04, JJM^{+}11, MD04, MS23, MVZ98, MDC^{+}08, ND17, NU05, OIH10, OGC^{+}15, PFTB15, PMV^{+}20, SH90, SHCR98, Zen99, AAV08, ACMT20, A^{+}02, ABMN02, BJ18, BGRD00, CZJS12, CJLS14, CD00, Cla03, DS03, DCG^{+}07, DWB99, DK04, EPS98, FKK^{+}06, FGP23, GKS07, GMF00, GGL21, GL97, HJ05, HRR02, HHJP16, HMH^{+}13, HF20, KRW88, KT00, KCG08, KA19, KG03, LyHW^{+}22, LG01, LVC16, LLD19, MKL^{+}01, MAH^{+}19, MFS^{+}19, MSK14, MK97, MK12, MHL06, MGT^{+}13, NS94, OILZ17, PVK^{+}22, PGG11, QH19, RGBCS09, RW89, SAWH88, She93, SCM^{+}98, SVC19, SDMS99, SDMS05, SNO99, Tak06, TKC^{+}14, TPK^{+}14]. high [TRSHB04, TAB^{+}19, WJK98, Wel89, WS98, WvR16, YDTS01, ZLM98, ZSI02, BZ99, BBD^{+}12, BCC^{+}97a, BCC^{+}97b, DHS00, MZ01, PFS^{+}04, KSM^{+}94]. high-dimensional [HHJP16]. High-End [BVC16, BVC19, BDO17, CBV13, AAV08, MK12, SVC19]. high-fidelity [DCG^{+}07]. high-frequency [ACMT20]. High-level [CCS03, Ger98, BJ18, MGT^{+}13, RW89]. High-order [SH90, DK04, GGL21, PVK^{+}22]. High-performance [BA08, BY21, CDK^{+}03, CV02, DDF^{+}10, EHF^{+}97, cFM07, Gog11, GLDS96, GKKS01, HR04, MD04, ND17, OIH10, Zen99, A^{+}02, ABMN02, CZJS12, CJLS14, FGP23, GMF00, GL97, HRR02, KCG08, KG03, LVC16, MAH^{+}19, MFS^{+}19, MK97, OILZ17, QH19, RGBCS09, SCM^{+}98, SDMS99, Tak06, TRSHB04, TAB^{+}19, WS98, ZLM98]. high-speed [FKK^{+}06]. high-T [KSM^{+}94]. high-throughput [PGG11, TPK^{+}14]. Higher [PS84, KHC92, MS00]. Higher-order [PS84]. highlights [McB88b].

Highly [BN01, SC95, BF95, BP86, EMV^{+}18, Ori10, PW84, PW87, RBM^{+}16]. highly-flexible [EMV^{+}18]. hints [HK02b]. HIRLAM [SB97]. historical [FC05]. histories [CRD02]. History [MVZ98]. Hitachi [Tak03]. HMC [LC17]. HMC-Sim-2.0 [LC17]. HMMer [OYS08]. hoc [CWK09, MFGEL17, SSS^{+}05]. Hogs [ORM^{+}10]. HoL [PGGL17]. HoL-blocking [PGGL17]. Holistic [Sto89]. holographic [EMB89]. HOM4PS [LT09]. HOM4PS-2.0 [LT09]. HOM4PS-2.0para [LT09]. homogeneous [CGM05, KRN^{+}11, MCG^{+}12, PVP08]. homology [BNK15, SNS^{+}97]. homotopy [CARW91]. honey [GÁVRRL18]. honey-bee [GÁVRRL18]. Honeycomb [TTH05]. Hopfield [SH90]. host [EB94, NTHY22, Str87]. host-computer [Str87]. host-computer/array-processor [Str87]. hot [AB93]. hot-spots [AB93]. Hough [vCd90]. householder [RH99, CUSR88, KCN99, OW90]. householder-QL [RH99]. hp [GPSK09, LLP00]. HPC [AAB^{+}16, ACMT20, ANG^{+}20, BDPV99, BRS^{+}17, BOS^{+}09, EYP^{+}20, ECLV12, LHR^{+}19, Gin99, KP12, KAM^{+}20, LCLA19, MSE19, MKJC21a, MKJC21b, NOG^{+}22, Pan01, PRS^{+}14, PB16, PGK^{+}18, PSF^{+}15, RGGP^{+}18, SPK18, WL13b, XLS^{+}17, YXQ^{+}21, ZS21]. HPCToolkit [ZAA^{+}21]. HPD [ML20]. HPF [BCC^{+}97b, DRST03, EGKU02, SNK06, SDv98]. HPFIT [BCC^{+}97a, BCC^{+}97b]. HTA [FBG^{+}12]. hub [CSEK03]. Hughes [EG90]. hull [DZD01, EM85, HI92, SE87]. human [BM08, SHRN98, VLSPL19]. humanoid [CDV08]. hundreds [SKE^{+}22]. Hungarian [DN16]. Hut [GKS98]. HVH [FT93]. Hybrid [AGLP06, EMV^{+}18, Hor91, IDS16, LWJ^{+}17, PS20, TBM^{+}16b, WL13a, ADR^{+}05, BG21a, BG21b, BEK^{+}11, CLY^{+}19, Cha99, CLC08, CTW14, GSE^{+}15, GPS03, GHW08, GÁVRRL18, GK19, IUY^{+}04, KCN99, LSC^{+}15, MRRP11, MA22, Nak05, Nat90, PGK^{+}18, POHS14, PS06, RB14, RBM^{+}16, ST21, SSL03, TPK^{+}13, TND10, TDB10, VK17, WW92, WY11, WS12, WSR14, YWP11, ZWJ^{+}19, ZGZS20, Ken99]. hybrid-memory [RB14]. hybrid-parallel [MA22]. Hybrid-view [TBM^{+}16b]. hydrodynamic [LKP24]. hydrodynamics [BSE88, FD18, Wil88]. hydrological [WFW^{+}11]. hydrophobic [Rob00]. Hyper [AIO95, Lip99b, LPPS01, BFG^{+}07, THK^{+}99]. hyper-bus [THK^{+}99]. Hyper-ring [AIO95]. Hyper-systolic [Lip99b, LPPS01]. hyper-threading [BFG^{+}07]. hyperbolic [DG05, EL90, ME90c]. hypercomputing [CC97a]. Hypercube [BF92, Nel93, AL93a, AM93, Bro88b, CARW91, CC93, DDP90a, DDP90b, DG90, DFRC90, DFRC94, Don05, FOH87, Fu03, GP90, GMBZ90, GLN89, GF03, HCK94, Hsi06, KSP97, LD99, LS92, MV87a, PW89, PW22, Pop88, SER90, SKC90, SSB^{+}91, SH01, TFS91, VP95, ZRPI89, ZMMW90, ZM96, KL00]. hypercubes [ASDOK01, BE88, Ber89, BD95, Cha88, CHLS90, CS97a, DVGG98, Dun91, EAR93, GP98, GS96, HC94, KLL^{+}09, LBL95, LY93, LC99b, LF88, PS91, RDB^{+}90, SAOKM01, She92, She95, SlWC89, SBO^{+}91, TY91b, VB92, VB95, XCS05, YTCW07, Zhu93]. hypercubic [dG98]. hypergraph [FBAB13]. Hyperion [A^{+}01a]. hyperledger [WLZ^{+}23]. hypersonic [SOS97]. hyperspectral [PVP08]. hypre [FLS^{+}21]. HySet [BG21a, BG21b].

I-Cluster [RMRN05]. I/O

[Bak00, BPC21a, BPC21b, CDZ^{+}21,

DBI^{+}17, HMP^{+}16, MV17, MQ97, PUSS97, SDMS12, TDC19, TF98, WG97, YM05, ZS21, ZJDW18, ZYZ^{+}16, SC03]. I/O-aware [ZYZ^{+}16]. IBM [AL93b, BBB^{+}94, Bri95, Car88a, CK90, GSZ88, GYL00, NMW93, Pet91, RRS88, Reu88, WZV^{+}16, Web90]. ICAP [CC99b]. ICCG [RG92]. ICENI [FMM^{+}02]. ICL [Car88b, Dav88, PRW88, WC90]. icosahedral [KTN^{+}14, WCKM11]. ideas [HM89a, HJ97a]. identical [DC15]. identification [EAR93, Leu93, VBS^{+}15]. Identifying [CFS01, LSS05]. idle [RMRN05]. IFC [Ano02f]. IFIP [KVW97a]. IFS [BDI^{+}95]. ignited [Yaş01]. ignition [Yaş01]. II [ABMN02, BCC^{+}97b, Eis89, ER95b, OR88, PWY03, Pir96]. III [Ano88b, SSGF00, Tol02]. illumination [Kat03]. illustrated [CT88]. illustrative [CR90]. ILP [HXW^{+}13]. ILU [HRR08a, PLP97, SZ02, VvW98]. ILU-preconditioned [VvW98]. ILUPACK [ABB^{+}16, ADEQO19]. Image [Has90, YEC97, AFK01, AATK10, BZ96, BCB02, BP00, BCYB11, Che98a, Chr98, DCDV98, EA95, FCC16, GBH98, GWC^{+}99, HL97, HA97, IOH05, Kao08, Kut02, LSS05, LPCA98, kLH95, Lee97, LS90, MP08a, MP08b, MCG98, NJ02, OAJ^{+}16, OA08, PPZ93, SKN04, SKG02, SG02b, TMCC02, TPK^{+}14, Tor91, UZ02, WE97, dCC89]. image-based [AATK10, OAJ^{+}16]. image-guided [GBH98]. image-to-mesh [FCC16]. images [BL93b, FA11, LCLL00, Möl99, SHPA05, Ume01]. imaging [CDM06, MDC^{+}08, SSS^{+}09, SHCR98, TAB^{+}19, WJK98]. imbalanced [MG19]. imbedding [WAB89]. Immortal [TS21, JBWE14]. immune [BJSN04]. Impact [ADLL03, BKT91, GLU03, KNTG08, BRS^{+}17, CB17, HH91, KV99, PRV87, PTK04, PRV08, RIB10, TB22, YMYH24]. impacts [Sat04]. imperative [OM90, SR14]. implement [Gro19]. Implementation [Ale94, BHM^{+}95, BMG07, Cal96, CHLO85, DL92, DSS86, DWS^{+}21, EG92b, HLDS95, HK02b, KW90, KSS98, mM94, MP87, PCN10, Ram99, RBM^{+}16, Sen91, SB97, Ste88, Tem88, TM90, ZLM98, AJ95, Ale91, AC89, AAHF97, ADLL03, ARZ97, AIV95, AW13, BCPB05, Beb97, BG97, BFG^{+}07, Bjø87, BS05, BDK96, BRH90, BP02, Bri95, BAS13, CCEJ01, Car98, CDZ^{+}98, CG89a, CDGM96, CT94, CH97, CGOP03, Daḡ07, DAA94, DF00, Dod89, Duf86, DOD23, El 93, EY90, FG89, FGG^{+}98, FCS^{+}19, GM98, God00, GLDS96, GL97, GT07, Güs99, HJS23, HW02, HCL05, HKR^{+}10, KK01, Kie91, KKSS90, KHHC20, LNS03, Li89, LMG09, LLD19, LS92, MRW^{+}08, ME92a, Mar95, MCC04, MKK03, Mic97, MW94, MNP93, Möl99, NR01, O’L87, OJ90, ÖS94, PPZ93, PCA01, Pet97, Pir96]. implementation [PAD02, Pry97, RGT17, Ram07, SKY93, SKS^{+}95, SCCP08, SCCPM14, SSB12, SAWH88, SP97, ST95, Sol88, Ste87, Tak10, TGE92, TSCS14, TMD^{+}97, VvBv90, VIHD90, VGRS16, WH91, WZ90, WL98, YA24, YL19, ZCPT00, di 97]. Implementations [DFP^{+}19, Sin90, BBK90, BC91, CAC^{+}09, CZ91, CKLM14, GGFF93, JDG02, LG03, LLW^{+}15, NAC^{+}14, RWT97, Rib84, RR07, SPCB22, Ste13, TF01, WLK^{+}18, Zha95, van86]. implemented [ARB94, ARB95, ETV91, FBMV88, HHSM88, KC91, PW84]. Implementing [AD14, BE93c, Bus93, CC97b, HTB01, Ris90, RVGG01, LHZ97]. implications [GEvGCP09, OCSBY01]. implicit [ES88, GKKS01, MPJ03]. impractical [Par87a]. improve [BBL^{+}16, GIRT19, ZWJ^{+}19]. Improved [EG92a, Lep99, MWI^{+}15, RV96, She92, ZS21, BKS06, CC02a, CJ92, HB99, Jeo91a, KG87, KB20, Nag90, SYY^{+}22, VR95, SG02a]. improvement [ACM20, Gin99, MPD00]. improvements [ASH92, OGC^{+}15]. Improving [APPG15, ASA16, AAO13, BPC21a, BPC21b, BJV^{+}16, HSÇ12, IMQO^{+}18, LFS14, Mah96, OK22, SIH14, dSS09, Coo19, KJ07, MSE19]. impure [BHS99]. ImRP [FTY^{+}20]. in-network [BCRSR11]. including [MT00, OCSBY01, Ter13]. Incomplete [AHBD18, BB21, DW99, Haa98, HCK94, HC94, SHCD00, TY91b, mMvdV01]. incompressible [BM09, GV01, JS13, KYLH01, NP05, XT11, YD97]. Incorporating [EG95]. increase [CPdM^{+}97, OID^{+}12]. increasing [JSC97, YMY^{+}19]. Incremental [SSKÇ15, HM02]. indefinite [BBB^{+}14, LWS02, SS02, SZS22]. Independent [Leu89, PLP97, BSE88, BDSd94, FMB98, Ger04, JHSV02, Lin91a, LF00, MV17, MT90, RvG05, YTCW07]. Index [Ano99a, Ano00, Ano02a, BHB21, Ano84a, Ano01, SRK^{+}21]. indications [Laf02]. indices [YS96]. indirect [Bro88a, NTHY22]. inductance [MS03a]. induction [MRW^{+}08, OCPT97]. Industrial [SMTT96, TS99]. industry [EHF^{+}97]. inelastic [CT88, DDF^{+}10]. inequality [CB97]. inexact [BMP15]. inference [LH04a]. InfiniBand [PK05]. InfiniBand-based [PK05]. infinity [CC00a]. infix [SS87]. influence [HHOMR06]. Influences [Gin99]. information [BGP^{+}97, Gro19, JCWP07, JCE96, LMJC96, MTV08, PSS01, Zam99]. Infrastructure [Ano13i, HKW13, LC17, SSO^{+}14]. inherently [TDG^{+}18]. inhibition [Yan04]. initial [HK91, KH92, Kie94, MGT^{+}13, Pir96, PER17]. initialization [BCA08]. Initiative [GHH18, HG17, NGM19, vOS^{+}98]. initio [BS00a, BCA08, HC05, SSGF00, SKE^{+}22]. injured [CS97a]. Inner [Rod85, GKT^{+}15]. Inner/outer [Rod85]. innermost [TM09]. input [AS86, EG95, WL14]. input/output [AS86, EG95]. inputs [LTG23]. insensitive [ST02]. insertion [APPG15]. insight [AGH^{+}94]. inspector [ATA^{+}16]. inspector-execution [ATA^{+}16]. inspired [RBS10, ZEgT04]. instances [BPK12]. Institute [KB08]. instruction [KLS^{+}88, SS91a]. instructions [CC92]. Instrumentation [NS02, CZTS99]. insurance [CDM^{+}10]. Integer [GMMT16, FF21, HA14, KC93a, MHH97, NW96]. integral [ADF93, ALH^{+}14, BD89a, DWB99, ER89, WY11, YTSI13, ZSH97]. integrals [FSY88]. Integrated [KMLM97, BCC^{+}97a, BCC^{+}97b, AKK16, JMA^{+}13, SV97, WE97]. Integrating [LMJC96, LTS^{+}24, OYS08, STP^{+}19]. integration [BGWR21, Bur90, DLM97, EM86, ES89b, HLvHA11, PATC99, RMS^{+}18a, SR17, TJJ93, dG98]. integrative [TRSHB04]. integrators [SSL19]. Intel [BKS19, BKL^{+}19, HZY^{+}19, LYC^{+}19, WHL19, AV15, BRH90, CLC^{+}18, DR15, DSSD18, Dun91, FYH^{+}18, GAW96, GKV94, HC92a, Hoc94, HSSM17, LCE^{+}18, LLX^{+}18, LHL18, Pir96, SNS^{+}97, SWR^{+}18, Tak18]. intelligence [Fah85, Loh96]. intelligent [YBM05]. intense [RMRN05]. Intensive [ANG^{+}20, RC17, CTS02, FAS02, SC03, TS02, TS09]. Inter [MSE95, BFY06, LFL11, Ume01]. inter-cell [Ume01]. inter-communication [BFY06]. inter-node [LFL11]. Inter-procedural [MSE95]. interaction [BM08, BJP^{+}89, CDH^{+}03, DRST03, HOSS91, PIG^{+}16]. interactions [LKP24, MM96b, RWT97]. Interactive [LMC05, WOKH96, CSB00, DGBP05, GBH98, IJM^{+}05, ISH03, Li89, RBL97]. interconnect [BHM94, KRN^{+}11, PMS^{+}13]. Interconnection [KV99, ACM^{+}10, ACM^{+}15, Bro87, CWC00, Dzw91, FSKF06, FHKZ88, GHM97, HCK94, HW94, KJ07, LJD93, Leu93, LH00, Miš98, NSH^{+}21, NMARD10, She93, SS01, Y^{+}99]. interconnections [JK91]. interconnects [FKK^{+}06, RBP^{+}17]. Interface [BC19a, GLDS96, UÇA10, ACLN03, EYP^{+}20, GPLW17, HHK^{+}19, Nag90, OILZ17, Pie94, RMS^{+}18a, Wal94a, Wal94b, BBB^{+}94, FGG^{+}98]. interface-resolved [GPLW17]. interfaces [BGL01, FCS^{+}19, Hem94, Trä12, THMH21]. interference [OL86, ZJDW18]. Intergroup [KTAB^{+}19]. interior [DM03, DR03, NYFK06, RWT97]. interior-point [DR03, NYFK06]. Interlock [CZTS99]. interlocking [GMBZ90]. intermediate [BVP^{+}89]. internal [Car88a, MRJ89, RS87]. International [ABH18, Ano88b, Ano88a, Ano16k, BH92, Emm84, Jou97, SHB19, HJB^{+}21, Ano87]. Interoperability [EGTD99, WPS^{+}08]. interplay [LKGD16, RMS^{+}18b]. interpolation [BAS13, Jan06, KP93, MKC92, SMK91, WA03]. Interprocedural [CI98, NL98, JCE96]. interprocessor [SKS^{+}95]. interprocessor-communication [SKS^{+}95]. interruptions [BRRV11]. intersection [Ban97]. interval [GYH94, LR16, Sch87, Shy90, UN87, Ver99]. interval-ordered [Ver99]. intervals [SC95]. intra [RRP03]. intra-tile [RRP03]. intracranial [PIG^{+}16]. intrinsics [CCL04]. Introducing [Vég18]. Introduction [DF95b, KVW97b, SYP13, VBC19, BH14, Cro97, GT19, KP12, RC14, WG97]. intrusion [BJSN04]. Invariant [WAB89, HA14]. invariants [MS89]. Inverse [VBS^{+}15, ARR19, Daḡ07, FPT91, GCH21, MR85, WE93]. Inverses [AHBD18]. Inversion [KL90b, CKM94, CGQ93, FF92, JLY18, KP93, KK94, LE91, LT90, MS89, NM01]. invert [ALH^{+}14]. Investigating [MM00, WBN^{+}17]. Investigation [MSE93, SDMS12, BHK89, Dod90]. inviscid [PEO16]. IO [TGL02]. ion [N^{+}00]. ionization [XLY^{+}20]. iPregel [CHZ^{+}19]. IPS [GE92]. iPSC [BRH90, Dun91, GKV94, HC92a]. iPSC/2 [BRH90]. iPSC/860 [Dun91, GKV94, HC92a]. IR [ZJDW18]. Irregular [TFV16, BCC^{+}97b, CHR00a, Dec00, GAPZ00, GPZ08, HT00, HMP^{+}16, Her00, KS05, LBGO23, PHCR05, PAG^{+}05, SVTSM^{+}22, TPK^{+}13, TB22, Trä18, ZPAT99]. irregularities [LLP00]. irregularly [SW03]. Ising [AL93b, HB90]. Island [RIB10]. isogeometric [PSW18]. Isopycnic [BDOS95]. isostatic [HWL^{+}22]. isotropic [VB92, VB95]. Issue [BVC16, CP17, GHH18, HG17, MSS19, NGM19, PS20, RST11, RC17, SYP13, TBM16a, TFV16, AMN16, AAGS18, Ano18k, Ano22m, ASSS11, AGKS15, AGKS16, BH14, BVC19, BC19a, BDG^{+}10, BOG15, CP18, CBV13, DF95b, GT19, KST02, KVW97b, RC14, SAGV21, TH20]. Issues [Ano87, BPEL05, MP08b, BLRR02, DM03, DWB99, DSCP88, HLDS95, Hir86, NBB^{+}02, Pan01, VV07, WLN^{+}96]. itemset [XZYQ21]. Iteration [WWZ^{+}18, GR02, HWB92, Xue96, BBC^{+}11]. iterations [BM93, BP93, ÜD89, Wei93, ZDE95]. Iterative [HI92, LPNV20, ABMQO11, ARB94, ARB95, BCVC05, BOVG10, BD89b, BB21, CGG01, CS96, Eis89, ESV10, ES06, Eva84, EY90, FHMS93, FG96, GKT^{+}15, HEB96, JP94, KMB^{+}18, KA96, MS03a, MS03b, MNP93, Mis87, MS90, Nak05, PS89a, PGW16, RP85, RDS13, RT88, Rod85, SB96, Son94, ZGG92]. iteratively [ZG16]. ixed [FD02].

Jacobi [BOVG10, OV06, ADFQO19, BOV02, BO03, DL97, KG87, ME93b, ME93a, RVGG01, TY91a]. Jacobi-sets [ME93b, ME93a]. Japan [FC05, Ohb85, Oya99]. Japanese [YK87]. Java [A^{+}01a, FWL03, VGRS16, VGS14]. Javelin [BB21]. jet [JJAP^{+}20, SML^{+}14]. Job [GF03, ABR03, ECLV12, GCP11, MSW91, TRLD13]. jobs [BM11, BD95, FCZY17, KEEF01]. Join [FC18, BG21a, BG21b, She95]. joining [DL06]. joint [LWW^{+}21]. Jordan [Ano03k, ADFQO19, CTT89, Mel87]. Josephus [LH00, LH04b]. Journal [MSS19]. JPEG [NU05]. JStar [UWC14]. July [Ano21l, Ano21m, Ano22h, Ano23f]. June [Ano21n, Ano21o, Ano24d]. just [FK94].

Kaczmarz [ESV10]. Kalman [Hot89]. KAMM [Adr99]. Kaps [HE88]. Karlsruhe [Adr99]. Kepler [TLS16]. kernel [AAO13, HBH^{+}16, LG21]. kernels [BLRR02, EL95, KKW14, WMCU97]. Key [DHW97, ZKCL04, NE01, SBZ04, WZL09]. Kilauea [Kat03]. kind [ADF93]. kinetic [CDH^{+}03]. Kirchhoff [SSS99]. knapsack [CCS94, CJ92, CFH89, LLW^{+}15, LC97, SSY02]. Knights [HSSM17]. knowledge [Bra88, DFRC94, LH04a, LMJC96, YTCS97]. knowledge-base [DFRC94]. knowledge-based [Bra88, YTCS97]. Knuth [OW97]. Knuth-Eve [OW97]. Kohonen [TMCC02]. Kronecker [AAHF97]. Krylov [AGD^{+}17, BLT^{+}22, LR99, MSZM14, de 96]. KSR [RSRM96]. KSR-1 [RSRM96]. KSR2 [BEK95b]. Kutta [ES89b, MPJ03].

labeling [IKK15]. labelling [HLP10]. laboratories [EPS98]. laboratory [TLS^{+}08]. laden [GPLW17]. Lagged [Mak94a, MS04a]. Lagged-Fibonacci [Mak94a, MS04a]. Lagrange [Rob00, Wil88]. Lagrangian [DAD11, NDN20, SG16, TN13]. Lambda [CC00a]. Lambiotte [Kra84]. LAN [ALT04, AAB^{+}06, DG15]. LAN/WAN [AAB^{+}06, ALT04]. Lana [AT98]. Lana-Match [AT98]. Lanczos [ALH^{+}14, BCA08, BF06, CSW01, KC91, SWC99, YTSI13]. Lanczos-like [KC91]. Landing [HSSM17]. landmark [KSS06]. Landscape [HW02, MVdLvN16]. landslide [DS03]. Langemeyer [NSS15]. Language [Has90, KT99, BN01, BVP^{+}89, BE97, CMM^{+}88b, FAS02, HXCZ14, Jor88, MIA^{+}07, MV87b, OTI^{+}89, Seu85, Sne88, Son86, TBM^{+}16b, UWC14, VP92, WHW^{+}11]. language-based [BN01]. languages [ANG^{+}20, ACH98, Ghe85, JC94, LOST16, SR14, Vio04]. LAPACK [Phi91]. LAPT [CDA^{+}16]. Large [Cér07, CDC^{+}87, IHM^{+}12, MZ04, ORS90, OA08, SMW^{+}05, TA14, VRL03, YNO97, van87b, ASA16, ALD01, AMC^{+}19, BSH88, BBB^{+}14, BW92, B^{+}01, B^{+}02, BCMG^{+}07, Bjø87, BG98, BPC21a, BPC21b, BR84, CLS93, CGK^{+}16, CGG01, CS97b, CBP^{+}07, CV17, CDH^{+}03, CDJ08, DM03, DLS09, DWS^{+}23, GR89, GMW96, GWLS05, GW95, GGL^{+}05, HCJ03, HT94, HF20, IU87, JBC^{+}07, JCWP07, LMC05, LLFZ23, MHE19, MSMC15b, MP02, NHS^{+}95, NFG^{+}13, NR20, ONB11, OID^{+}12, PD11, Qua92, RS08, RJ97, SSSG03, SZ02, SML^{+}14, SBM^{+}22, Ste90, SLH^{+}13b, WYT^{+}16, WC90, WZ90, Won88, WP19, YX07, YXQ^{+}21, YM05, ZZ03, WH15]. large-grain [BSH88]. large-message [AMC^{+}19]. Large-scale [CDC^{+}87, IHM^{+}12, MZ04, ORS90, OA08, YNO97, ASA16, B^{+}02, BG98, BR84, CS97b, HCJ03, LLFZ23, MHE19, NFG^{+}13, NR20, ONB11, OID^{+}12, PD11, Qua92, WYT^{+}16, YX07, YXQ^{+}21, YM05]. larger [XC02]. largest [CIG^{+}21, SlWC89]. laser [CDH^{+}03]. last [APPG15, Zen99]. latencies [MS03b]. Latency [SC05, ACM20, CC97b, GV14, KG21, LW11, LGM09, NSH^{+}21, NTHY22, PMS^{+}13, SA16, Tap84b, VCK^{+}11]. latforms [FD02]. Lattice [GLS99, GAW96, Uka99, A^{+}99, BP89, CIS99, CGK^{+}16, EHN99, FHK^{+}11, FHK^{+}15, GPLW17, HAK^{+}21, Hio96, Kat86, Lep99, Lip99a, OKTR13, SR97, XT11]. lattice-Boltzmann [CGK^{+}16]. law [MOW96, YMG14]. laws [DG05, GL90]. layer [Dod89, FT93, LHG^{+}23, Müh90, PFS^{+}04, RBM^{+}16, SVC07, SM90, XCR17]. layered [Kak88, WM97]. layout [AGK98, GLD08]. layouts [BETR17, SSH97]. LC [LC99b]. LCAP [Hoc89]. LDPC [AFS14, GKS07]. leading [BRS^{+}17]. LEAN [BVP^{+}89]. learned [Gil94a, Sho17]. Learning [SA18, AMC^{+}19, BKK24, BSK03, CDZ^{+}21, DBZ^{+}19, GZH^{+}23, HT92, HT94, Ken90, LG21, LWW^{+}21, MMC97, SMT^{+}04, WWZ^{+}18, WZ90, da 90]. leasing [LSC^{+}15]. least [KCN99, LWS02, MS90, NP05, OHZ98, PGW16, Str08, Sun97a, Wei93, YK08, Zhu93, ZG16]. least-squares [KCN99, LWS02, MS90, NP05]. Lees [RW96]. Lees-Edwards [RW96]. legacy [CMPM^{+}15, SR14]. Lemur [CHLO85]. length [MT90, PW89]. lengths [KLL^{+}09]. less [ZRS13]. Lessons [Sho17, Gil94a]. level [APPG15, ABMQO11, BJ18, BS97, CCS03, CZJS12, CL05, CRGM16, DK04, FFZ^{+}18, FCC16, GRP22, Ger98, Hor93b, HRR08b, HC05, JS13, LHG^{+}23, Lip99b, LCLL00, LND^{+}19, ML00, MAD^{+}16, MGT^{+}13, NR20, PS89b, RGT17, RW89, SNK06, SG02a, SZ02, SK97b, TW19, TYLL22, XLS^{+}17, DMWW88]. level- [Lip99b]. levels [BDK06, GHP10]. Levenshtein [CR23]. Leveraging [MRBQO14, BPJ22, LFL11]. lexical [Hil92]. lexicographic [LT94, TL90]. liabilities [ZCPT00]. liability [CDM^{+}10]. LIBI [GAA^{+}13]. Libmonitor [Kre13]. libraries [DWS^{+}21, FB19, FMB98, LG21, Lip99b, SH88, Y^{+}02]. library [AV15, ABMN02, BGWR21, BCPB05, BJvOR03, BVWH14, CHHW94, CTA22, CGG04, DAA94, DMWW88, HMS^{+}98, KT99, KMK^{+}19, LEH14, Pry97, SEAH^{+}08, SBM^{+}22, WS98, WPS^{+}08]. LICRA [Kan97]. life [CDM^{+}10]. Lifeline [RHWF23]. Lifeline-based [RHWF23]. Lifetime [ZWJ^{+}19]. lifting [ALTZ02]. ligand [ZEC^{+}17]. lightweight [MCB16, PCL23, SK97b]. like [Ale91, CS89a, FGS97, NMARD10, RR89, TSEE21, VGS14, KC91]. likelihood [Mal02]. limit [JCC^{+}24]. Limitations [Müh90, Sch95]. limited [BDF^{+}00, BD10, Lah00, MSM98, SB97, ZS21]. limited-size [ZS21]. LINDA [MSP93, CGMS94, Car98]. line [AALK01, BBQOQO00, CLJ14, Gol86, IT08, Leu93, LDK16, ME92b, RBMO20, ZSH97]. Linear [AvHL^{+}97, HH89, Jou89, SSD01, Sut97, ŚDJ^{+}22, Ume01, ABMQO11, AGLP06, AB95, ARB94, ARB95, AIV95, BBB^{+}14, BG91b, BBK90, Bar95, BLRR02, BP93, BG97, BGMT00, BWW89, Bjø87, Bog92, Bol86, Bon91, BP90, Bru91, BLKD09, CHQORS18a, CTZ^{+}18, CHLS90, CL02, CLA14, CPR^{+}18, CS96, CB97, CWB92, CV17, CMM^{+}88a, CTT89, CIO^{+}17, CDJ08, CGG04, DY91, DR18, Don05, DH84, DvdV99, DGP^{+}16, DR03, ESV10, EL95, ES06, ESK88, EB88, EG92b, EG94b, EB98, FF21, GKT^{+}15, GR89, GMW96, GYH94, Gao87, GW95, GH90, GG10, GPSK09, GE93, HS90, JK91, JP94, KMB^{+}18, KS89, KL00, KL90a, Koç97, KPS90, LWS02, LL90, LN87, MS03a, Mal02, Mas98, MC04, MG19, Mei85, Mel87, Mel88, MNP93, NHS^{+}95, NZ97, NE01, Pis92, PGW16, RRS88, RR89, RS08]. linear [RP85, Reu88, RS88, SSSG03, SS02, SK87, SZ02, Slo91, SE93, Stp93, SE98, Sun97a, TKC^{+}14, Ter10, Ter13, TDB10, THK14, Wag89, WE93, Wan08, WP19, YK06, ZM94, dP90b, van87a, van87b, DN16]. Linear-time [Ume01, Wan08]. linearized [HWB92]. lines [WAB89]. link [KSP97, KLL^{+}09]. link-disjoint [KSP97]. link-faults [KLL^{+}09]. linkage [SBM^{+}98]. linked [SSP^{+}98]. links [BMT92, CF04]. LINPACK [KK92, DZM^{+}13]. liquid [PBTC89]. LISSOM [CC02b]. List [KKT00, SS04, YG93, Che08, CFH89, LC97, SSY10, ST21, XRW^{+}19]. lists [KS95]. live [TKK^{+}05]. Livermore [Feo88]. LIW [CCRR91]. llc [DLdS06]. Load [BP97, BS05, BM01b, DR93, EB94, GGV04, GH92, LX00, MG19, MCG^{+}12, SGS95, YSBM97, BCPB05, BV04, BF95, BD18, Bog92, CE94, CS00, CFF17, CG87, DLM97, DLR94, DR95, Dec00, DG90, DFM99, DPSW00, EB93, FMD98, FG96, HL97, Hor93b, HB99, JH97, KNP97, KEEF01, Koh95, MD04, MMT07, MSE07, MZ95, NC97, OBG00, PRR03, RHWF23, RCAP11, San99, SSLK00, Son94, TSJB00, VR95, YEC97, Zam99]. Load-balancing [GGV04, MG19, BD18, HL97, MZ95, TSJB00]. load-imbalanced [MG19]. Load-skewing [LX00]. loading [BGM03, PM03]. loads [BD10, DG99, San99]. Local [BBL^{+}16, BT01, DC15, EGKU02, ER18, TJ97, ACM20, BL99, BFM90, CKLM14, Fre89, GHNL87, MPD00, Ng93, PHM^{+}22, RS88, SDv98]. local-memory [GHNL87, Ng93]. Locality [RR16, CDA^{+}16, FCC16, GPZ08, HCR01, Mah96, PCLM18, SIH14, Vio04]. Locality-aware [RR16, CDA^{+}16, FCC16]. locality-awareness [PCLM18]. locality-based [GPZ08]. localization [ATA^{+}16, KY98]. locally [SSP^{+}98]. location [BCG00, CSEK03, GPS03]. Log [GLP08]. Log-map [GLP08]. logarithmic [BO91, GDMS97]. LOGFLOW [KNP97]. logging [PLY02]. logic [CF91, Her00, UAR^{+}99, VKH99, Wel89]. logical [CZTS99, LWZL24]. logistics [LM03]. LogP [KKT00, LZ00, Ver00]. LogP-machines [LZ00]. Long [dP90b, CR23, LL88, OWW16, ZCBD22, Zla88]. Long-range [dP90b, Zla88]. long-track [OWW16]. longest [Lin94, TNZM20, TTO^{+}22]. look [BRS^{+}08, CHQORS18b, TRSC^{+}19, Wei91]. look-ahead [CHQORS18b, TRSC^{+}19]. lookahead [BT01]. Loop [BDSV98, BBC^{+}11, BCR00, CDRV97, DR94, Dar00, DWX^{+}12, DR96, GGL99, KR10, MJ95, MSE93, NL98, WMCU97, Xue94, YTCS97, YCS07]. loop-carried [BCR00]. loops [BPK12, BCR00, CK90, DFRR91, Feo88, GGL99, GDMS97, LK00, SM13, TM09, WW92, Xue96, Xue97]. loosely [CC99b, Gro87]. loosely-coupled [Gro87]. Lop [RGDM15, WCC^{+}22]. loss [Lin93]. Low [BLT^{+}22, LLFZ23, RGT17, XXQ^{+}15, ARCH05, CFD^{+}16, CHQORS18a, CC97b, CC00c, DHS00, GPSK09, HJ05, KG21, LAS90, Lec89, NSH^{+}21, NTHY22, N^{+}00, SSL03, Trä18, XKL^{+}22, YA24]. low-cost [CC00c]. low-density [ARCH05]. low-latency [NSH^{+}21, NTHY22]. Low-level [RGT17]. Low-overhead [XXQ^{+}15]. low-power [CFD^{+}16, CHQORS18a, YA24]. low-rank [XKL^{+}22]. Low-synch [BLT^{+}22]. Lower [TYLL22, GK92]. lowest [LO92]. Lowry [WWT01]. LPN [TSCS14]. LR [JCC^{+}24]. LR-TDDFT [JCC^{+}24]. LU [AC94b, BB21, MW11, SB92a]. Lunch [FC18]. Lustre [CDZ^{+}21]. LVQ [HKSK97]. Lyapunov [BEK^{+}11, KS16]. LZ2 [De 95].

M [HE88, FD02, DG94]. M-cache [DG94]. M3Set [BE97]. MAC [CWK09]. Machine [BL94a, DLPS92, McB88a, Sin90, SA18, AK89a, Amm89, AGGG06, AGGM90, BDK96, Car98, DFR92, DG95b, DZG94, EHE92, EG90, FHMS93, GKAS91, Hor93a, LG21, LKTZ88, MA22, OTI^{+}89, OKSY92, SB94b, Sor84, SJJ85, WP88, ZLPF16, DF91, Evr01, JK92, LSP92]. machines [AIO95, BKS06, BRS97, BMYK98, BD89b, Bro99, BWZ95, Cal96, CdCM07, DG90, ED04, GMS04, JK91, JCL92, KA96, Lan99, Lan90, Li91, LS93, LZ00, LF00, MAB17, MW94, MOW96, NH11, SN88, SNK06, SS94b, TX00, TPK^{+}13, Wor97, ZZ03, ZGG92, dGR95, di 97, DFO93]. macro [MSW91]. macro-tasking [MSW91]. macros [BRH90]. macroscopic [Wei01]. macrotasking [DO88]. MAD [KGV97]. Madeleine [ABMN02]. magnetic [MRW^{+}08, SHRN98]. magnetohydrodynamic [KT00]. magnetohydrodynamics [CCP^{+}21]. Maintenance [WE91b]. major [WLK^{+}18]. makespan [CHR18, ST21]. making [KNR00]. MALLBA [AAB^{+}06]. malleable [IMQO^{+}18, MSMC15a]. management [A^{+}02, ACM^{+}15, Bak00, BCG00, BLCR21a, BLCR21b, BGL01, CHBS18, CDM^{+}10, DCR^{+}16, DWH^{+}08, GJMM18, Kan97, KKH04, KAM^{+}20, LLP00, LF98, LCLA19, LMH08, MKK^{+}19, PŚ00, RMRN05, RGBC11, SVC07, TPK^{+}14, WXL^{+}18, Y^{+}02, YD01, YBM05, ZCPT00, vOS^{+}98, MPZS13]. manager [BÁJG^{+}20]. managers [RMACG10]. managing [LLP00]. maneuvering [TB03]. manifesto [Col04]. manifold [Han85, Pap98]. Manila [MHE19]. manipulating [MHT06]. manipulation [AL93a]. MANNA [Sod02]. manufacturer [SH91]. many [ABB^{+}16, BB21, Cal85, CFF17, CDGI15, DTR18, DEH^{+}11, FD18, AKK16, GB14, JMA^{+}13, LXC^{+}18, MMP^{+}21, NMAB19, SBP12, SA13, Sug99, Vég18, ZJD^{+}23, RHWF23]. many-body [Sug99]. many-core [ABB^{+}16, BB21, CFF17, CDGI15, DTR18, DEH^{+}11, FD18, GB14, LXC^{+}18, MMP^{+}21, NMAB19, SBP12, SA13, ZJD^{+}23]. many-cores [JMA^{+}13]. many-processor [Cal85, Vég18]. Many-Task [RHWF23]. manycore [AAO13, CS19, MII^{+}11, MCB16, TDB10]. Manycores [ABH18, BH14, CFD^{+}16, SHB19]. map [GLP08, KSS06, TN13, WHW91]. map-reduce [TN13]. Mapping [BDNP11, DR94, DR96, FFLV96, DST15, HKSK97, HHSM88, KM89, LBL95, MM96a, Pop88, SSS^{+}05, TMCC02, Wol89, ARCH05, AAO13, ANG^{+}20, BBL^{+}16, BE93a, CHLS90, CS19, CA92, CDA^{+}16, DCN^{+}15, GAMR03, HCT16, Hen93, Hur93, KNP97, KTN^{+}14, MGSK87, RV96, RH12, SER90, SNK06, SM91, YXQ^{+}21, YL86, ZLD15]. mappings [BM93, RKS92]. MapReduce [BD18, FCZY17, KdBMA^{+}13, LWJ^{+}17, MMM13, PD11, RDS13, SWYM17, XLY^{+}20]. March [Ano20l, Ano22i]. marching [Vaj84]. market [CC00b, CL05]. marketplace [SDMS99, SDMS05]. Markov [Beb97]. masks [APRP97]. mass [ND95]. massive [DM90, Ger94, KTN^{+}14, PCL23, Xue94]. Massively [B^{+}00, CGK^{+}16, DFO93, GL90, TMD^{+}97, W^{+}99, AK89a, AMP92, BRS97, BHM^{+}95, BEK95b, BWZ95, BK97, ÇFG^{+}12, CR90, DFR92, DEGS95, DFRC94, DAG^{+}09, Ekl04, GWWM09, Ger98, GHC04, HLDS95, HISS92, KK01, KCN99, Li97, NNB^{+}99, NZ92, Rea90, RR97, RKMJ10, SHPA05, SSB12, TJ97, WRB97, WOKH96, YNO97, ZJD^{+}23]. massively-parallel [HISS92]. Master [CMSL06, BPP10, HP05, JMLBL98, PRV08, dSS09]. master-slave [HP05, PRV08, dSS09]. master-worker [BPP10]. Master/Worker [CMSL06]. masters [MV99]. match [HGM20, AT98]. matching [AB16, BDP20, BGL^{+}88, CDBL08, Dod89, FLPG18, GGA19, HHQ06, KBBC88, LPM11, LAHM14, LFS^{+}19, LSF^{+}15, MPD00, SHPA05, SSB12, SKC90, TLS16]. matchings [OA91]. materials [AC00, YD97]. math [CTA22]. Mathematical [ES89a, LS90, Zla88]. mating [GÁVRRL18]. MATLAB [UÇA10, EL16]. Matrices [DZLK20, BW92, BRS97, CU04, Con89, DR18, EB98, GD90, GKAS91, KGS08, KKSS90, Kuz98, Lan96a, LE91, Liu89, ML23, MAH^{+}19, MJ94, MW11, Mel88, MP02, MMS90, Pet84, Pin91, SK91, SPMB23, SGP22, SMM90, THM^{+}95, VP94, VB95, Wai88, WC90, XKL^{+}22, ZVWS88, GV11]. Matrix [AGPS03, ASSS11, AGKS15, AGKS16, ABH^{+}10, BDG^{+}10, CGNR06, DS13, FOH87, GMMT16, MV87a, UÇA10, ASA16, AAGS18, Ala95, AMP92, AC94b, AB16, Bar95, Bar97, BG01, BGSS14, BR89, Ber89, BDK95, BF06, BVWH14, BTLK18, CKM94, CS87, CDW95, CG89b, CGQ93, CMT96, DD87, DEGS95, DTR18, DSS86, Dow95, DB03, EM88, EMB89, EM90, EL91, Eva92, EW93, FAF16, FG89, Fis90, FPT91, FG96, GR01, GPS^{+}08, GS96, GKV94, HH91, HCR01, HRR08b, Joh93, Kak88, KG87, KB20, KST02, KP93, KK94, KD10, KAD09, LB23, Leu89, Lin91a, LPPS01, LV15, LT90, Mal02, ML20, MRBQO14, Mar14, MAH^{+}19, MWH95, ME89, Meg90b, MS89, NMAB19, NYFK06, NM01, Nel93, NW96, OZ02, Pet91, RH99, Ris90, RZ95, RR16, SAGV21, SA16, Sil88, SPMB23, Ste90, TY91a, TM90]. matrix [VTmL03, VFG12, WP88, WE93, WL14, WOV^{+}09, WRS12, YE95, YB11, ZJD^{+}23, ZDE95]. matrix-algebraic [AB16]. matrix-based [WL14]. matrix-matrix [NMAB19, Ris90, RR16]. matrix-squaring [EL91]. Matrix-Vector [DS13, AMP92, CMT96, DEGS95, FG89, GR01, HCR01, KD10, Lin91a, LV15, Mar14, WOV^{+}09, YB11]. maturity [FMM10]. max [SB92b]. max-min [SB92b]. maximal [AB16, CS97a, CC92, Coo19, GS04, HC04, SM88, YL19]. maximize [ZLPF16]. Maximizing [LL98]. maximum [BM93, CMM^{+}88a, HM89c, LAHM14, Mal02, NA03, SKC90, Shy90, Wen95]. Maxwell [ABG^{+}06, DE09, HKL01]. May [Ano20m, Ano21p, Ano22j]. MBSP [Ger15]. MCD [MMP^{+}21]. MCSPARSE [GMW96]. means [BY21, BCYB11, GÁVRRL18]. measure [OWZ91, Zub90]. Measurement [ZAA^{+}21, CB17, EHHS89, OL86]. measurements [Hoc85b, PSS94, RH13, Tem89]. measures [HCR01]. mechanic [SWW99]. mechanical [CT88, KR00]. mechanism [DG94, GZH^{+}23, MSS^{+}05, PSF^{+}15, RGBCS09, YMY^{+}19]. mechanisms [CPdM^{+}97, PGGL17]. media [DP02, DDF^{+}10, GM98, JSC97, RBL97, WB97]. medial [Wan09]. median [GLMBMPMV03]. mediaprocessors [MKK03]. medical [BMYK98, Chr98, DCDV98, SHPA05, WJK98]. Medium [Kau94]. Medium-range [Kau94]. meeting [DA06]. Megacell [LKTZ88]. MEGAFLOW [A^{+}01b]. Megathrust [BKL^{+}19, BBL^{+}19, WHL19, BKS19, HZY^{+}19, LYC^{+}19]. Meiko [BCM94, BHM94, Hoc94]. melts [OCSBY01]. membranes [SSV^{+}16]. Memoriam [Hol11]. memories [ARZ97, PPZ93]. Memory [BS00b, CAA98, DCD97, DGBP05, KKH04, Sch91b, ABGR21, AGH^{+}94, AGL08, AFT23, ADMV05, AB16, BG91a, Bai88, BKS06, BZ96, BZ99, BD10, BC04, BFM90, BM01b, BD89b, BDS88, Bro87, Bro88b, BS96, Bru91, BCMSW03, BMG07, BFR93, Cal96, CHZ^{+}19, CTZ^{+}18, CPR^{+}18, CSH^{+}11, CDW95, CG89a, CG89b, CSW01, CRT89, CJ95, DL92, DR94, DR93, DLR94, DH91, DPFT19, DvdG92, DZG94, DM02, ECG93, EHHS89, FFLV96, FFZ^{+}18, Fie96, FBMV88, FA11, FHMS93, FBAB13, FGP23, FDTZ04, FM88, Fre89, GSE^{+}15, GGH^{+}22, GD90, GHNL87, Ger98, GK92, GHH^{+}03, GB14, GL97, GLU03, HH91, HM02, HCT16, Hey91, HC89, HK90b, IKK15, IYV^{+}22, IHM^{+}12, JLY18, JIC96, Jor86, KW90, KK03, KK11, KJA15, KV99, KKŽ05, KZV97, KA96, Kuz98, LSL14, Lan99, LPM11, LR07]. memory [Lee95, LC02, LC17, LLD19, LSL02, LXD^{+}23, LS93, LF00, MP92, MMC97, MAB17, Mar95, MM00, MMTU18, Mat95, MSSC10, MCP^{+}14, MW94, MBC97, Ng93, NA08, OL86, OILZ17, PLY02, PK05, PRV87, PB16, PGK^{+}18, PGAA16, Phi91, QWD^{+}21a, QWD^{+}21b, RB14, RDS13, RGDM15, RMACG10, RS88, RG92, Rot95, RV95, SSKÇ15, SNK06, SK97a, SG02a, SB94b, SLH^{+}18, SCL^{+}23, SPMB23, Sol88, SB92a, Sun97a, TX00, TSCS14, TB07, Tol02, VA14, WW95, Wal94a, Wal94b, WC94, WME^{+}95, WY11, WH15, WE92, Wil88, Wri91, YL19, YSBM97, ZE92, ZWJ^{+}19, dGR95, di 97, DR15, HL13]. memory-constrained [QWD^{+}21a, QWD^{+}21b, TSCS14]. memory-coupled [FBMV88]. merge [AHBLR12, EY86]. mergesort [BGV97]. Merging [GS01, GS06, CG92, Eva90, HK90b, LWW^{+}21, VIHD90, WE91a]. Mesh [HST87, AAZ96, BN01, BCYB11, BAS13, CC95, CD00, DHS89, DPSW00, DG05, DG99, EJLC00, FK98, Far94, FCC16, Fie96, Glo95b, GV08, HHK95, HAK^{+}21, HI92, HT98, IDS16, IJCL96, Jan06, Jeo91a, Jeo91b, Kak88, KH97, KŠR04a, KŠR04b, Lan96b, Lee97, MM96b, MGT^{+}13, OBG00, RVGG01, Sch91a, Son92, VTmL03, WC00, WN91, Wan08, YD01, YSBM97, dG98, de 96]. mesh-adaptive [DG05]. mesh-based [de 96]. Mesh-connected [HST87, CC95, HI92, Jeo91a, Jeo91b, Son92, WN91, YD01]. mesh-of-processors [DHS89]. mesh-particle [BAS13]. mesh-partitioning [KŠR04a, KŠR04b]. meshes [Bar97, BL99, BMT92, CC02a, CST02, DAD11, Glo95a, HB94, HM97, MB00, NC02, OSZ93, RV96, ST02, SW00, TT00, Tro00, VB92, ZL04]. Mesoscale [Adr99, Mic97, AFG^{+}97, TMD^{+}97]. Message [BC19a, BCM94, BL94a, Fer97, GDAK06, PW89, SSOB03, AB02, ARCH05, AMC^{+}19, AE93, BRWL09, BDOS95, BWL00, CHHW94, CGMS94, CDZ^{+}98, CCS03, DG94, DM02, FK94, GGA19, Gog11, HP05, HM86, HGM20, ISO^{+}03, Jia09, JOL^{+}21, JIC96, KM89, KNTG08, LFS^{+}19, MRBQO14, McB94a, MRT93, MBK12, MKL^{+}88, NC02, Pie94, Pry97, SK97a, Ste90, TO91, Wak04, Wal94a, Wal94b, WLYJ13, WY11, FGG^{+}98, GLDS96, WZL09]. message-based [HM86, KM89, MKL^{+}88]. message-dependent [WLYJ13]. message-driven [MBK12]. message-handling [DG94]. Message-passing [BL94a, GDAK06, ARCH05, AE93, BRWL09, BDOS95, BWL00, CHHW94, CGMS94, CDZ^{+}98, CCS03, Gog11, JOL^{+}21, MRBQO14, MRT93, Pry97, SK97a, TO91]. messages [BL94b, Gir02, TM94]. Messaging [ZKCL04]. meta [MT00, MMT06, SK97b]. meta-heuristics [MMT06]. meta-level [SK97b]. meta-systems [MT00]. metabolic [JTS^{+}11]. Metacomputing [EHR^{+}98, vOS^{+}98, BGK^{+}98, BG98, DFRZ02]. metadata [DCR^{+}16]. metagenomic [FJPA17]. Metall [IYV^{+}22]. metaprogramming [TSEE21]. meteorological [Car88b, MSM98]. Method [DVGG98, FTY^{+}20, MKJC21a, MKJC21b, YR18, AL93a, AGL08, ALH^{+}14, AHBLR12, BSE88, BG91a, Beb97, BE92, BBQOQO00, BG97, Ber00, BFM90, BP02, BBB19, CH98, CLS^{+}16, CAC^{+}09, Car88b, Cha99, Cv98, CL02, CV17, Coo19, CGMM99, CKLM14, DM03, Daḡ07, DL92, DL97, Dav88, DR03, ES88, ETV91, ESK88, ES89b, EC91, EL91, Eva91, Eva92, EG94a, Gao87, GPLW17, Gol88, GdCCS03, GTW^{+}20, GKS98, Has01, HAK^{+}21, HK92b, Hor93b, HHSM88, HWL^{+}22, ISO^{+}03, KR97, KS89, Kan93, KG87, Kas85, Koç97, Kru97, LRH97, Li97, ME90a, MW11, MBC92, Mei85, Meu87, Mv88, Mil01, MS91, NYFK06, NC02, OJ90, PLP97, PJ84, PJV00, PL00, Pir96, QWD^{+}21a, QWD^{+}21b, RSK99, RV95, RVGG01, SCCP08, SCCPM14, SS89b, Sch87, SY87, Slo91, Sur10, TW19, TGE92, Ter13]. method [VBS^{+}15, VJ12, VA14, Wak04, WL13a, WWR05, WYX^{+}22, WBS06, WAT20, WY11, WAB89, WH93, WP19, XT11, XRW^{+}19, YE94, YE95, YTSI13, YXQ^{+}21, YWP11, YL19, dF99, van87a, Ger94]. Methodologies [OKSY92]. methodology [BE93a, CMT04b, DSCP88, GF03]. Methods [Lop93, Reu99, TO99, ADF93, AÖ22, AFPG12, Bai99, BRS97, BE89, BFH23, BD89b, BM90, CS95, CG89a, CS96, DF00, DL04, DK04, Eis89, Eva84, EY90, EL90, EK98, FBMV88, GT92, HA05, HM89a, HZ93, HM86, HHR87, JDG02, KW90, KHC92, LR99, LP92, MS03a, MRW^{+}08, MQ89, MSZM14, ME90c, Meh93, Mis87, Nak05, ÖW10, Oli96, Ort88, OR88, PS89a, Pel97, PŚ00, PER17, RM97, RT88, Rod85, SGS95, She96, Str08, SP93, TKG97, Thu90, Thu91, TLC^{+}21, WB88, Wat00, YM05, YSBM97, Zha95, de 96]. METIS [TTT^{+}92]. metric [FBAB13, Jeo91b, SG91]. Metrics [Ori10, PMAL14, Wor91]. Mfold [MSSC10]. MGR [Kam87]. Miami [BDOS95]. MICOM [BDOS95]. Micro [EO91, BY21, FA11, LXC^{+}18, NR01, TA14]. Micro- [EO91]. micro-benchmark [LXC^{+}18]. micro-finite [FA11]. micro-services [BY21]. micro-simulation [NR01]. microbenchmark [YHE^{+}19]. microcode [SC92]. microprocessor [RG09]. Microscopic [NS94, Wei01]. microservices [dNdRRL^{+}21]. Microtasking [DO88]. Microwave [TAB^{+}19]. Middleware [BAB^{+}02, DFRZ02, GE11, HSÇ12]. migration [BTZ06, BK07, CRGM16, GV08, HB84, LSL02, RIB10, SSS99]. mildly [Pir93, Pir96]. military [LG01]. millions [SKE^{+}22]. MILU [SY21]. MIMD [AB02, AGGM90, ADE84, BBK90, BE89, BE92, BE93b, Bog92, BT88, Chr98, CMRT88, CF90, CJ95, ER95a, ER95b, FFLV96, GKAS91, Hoc85a, Jor86, Kas85, Koh95, Kut02, LPCA98, LE91, MR89, MW94, MMS90, NM01, SN88, Sen91, SSL03, Sor84, Ste88, Uhl96, ZBG88]. MIMD/SIMD [MW94, ZBG88]. min [SB92b, ZG16, MSS^{+}05]. min-cut [ZG16]. minima [Kim98]. Minimal [BHK89, BR90, CF04, MA22, Tap84b, TL94, XC02, ZL04]. minimize [LX00]. Minimizing [HK02a, IZS^{+}20, Joh93, MMP^{+}21, LL98, WL14]. Minimum [BDK96, LC99a, ACM20, BG97, Cal96, CG93, CMT96, DDP90b, HWW92, LS97, Ume89, WL13a]. Minimum-weight [BDK96]. mining [C^{+}01b, CV02, GM04a, GK04, HLvHA11, SSN04, VA14, XZYQ21]. minisupercomputers [WSL88]. MIP [APPG15]. Miscellaneous [Ano97]. miss [LD98, RH13]. Mixed [OM90, BEK^{+}11, HSS07, HHOMR06, PWY03, RC15, SK87, YLCT18]. mixed-mode [HSS07]. mixed-precision [BEK^{+}11]. mixing [CdCM07, GJMMS97, KRS^{+}21]. mixture [EPMPU02]. MkIIa [ADE84]. ML [GH98]. MLEM [Möl99]. MLP [Taf01]. MM5 [Mic97]. MM90 [Mic97]. mobile [APPG15]. mobility [DMT06]. Möbius [HC04]. mode [GWC^{+}99, HSS07]. Model [BDOS95, WRS12, AL93b, ARR19, AFG^{+}97, BJ18, BMS01a, BDI^{+}95, BFVRC14, BI07, BJSN04, CE94, CCEJ01, CK90, CC96b, CC00b, CH92, CGM^{+}92, CH97, DM97, DGNP88, DS03, DFM^{+}95, FK98, FGHB94, FMS^{+}06, FBAB13, FG96, FT93, GZHX17, GE11, GLS01, GS01, Ger15, GCP11, God00, GPZ08, HRW^{+}95, HLDS95, HLvHA11, HB90, HT92, HB84, IB01, JAA06, JKH95, JMA^{+}13, Kas84, Kat86, KD97, KSS98, KA96, KK94, KS99, LH88, Lin01, MRW^{+}08, MSM98, MS99a, MM96b, MZ95, NLG99, Nak05, OCPT97, PFS^{+}04, PEO16, RRS^{+}00, RR14, RR16, SHN12, SKC90, SPW^{+}15, SB97, SSBT19, SW03, TMD^{+}97, Tol02, Van02, Ver00, WNES01, WM97, WAT20, WME^{+}95, WvR16, WLK^{+}18, WWT01, YK06, YK08, Yaş01, YHZ21, ZLT19, ZCPT00, de 96, Adr99, HW02, Mic97]. Model [RIB10]. Model-driven [WRS12, JMA^{+}13]. Modeling [AE93, CMSL06, CC99a, GHH^{+}03, HCR01, LGM09, LF88, MRSB94, RGDM15, TYSF13, AATK10, BMS97, BMS01b, CC96a, CLA15, CTW14, CCH^{+}21, DS96, DF95a, DF95b, ES89a, FHK^{+}15, HGS10, Kot97, KNTG08, LG01, Lin08, LND^{+}19, MVdLvN16, MJ99, NMAB19, NS94, PCC19, SXBD97, SLS^{+}21, WN91, WFW^{+}11, WCC^{+}22, vWWM^{+}19]. Modell [SK97a]. modelled [CM95]. Modelling [EHHS89, ME92c, OL86, PWM00, Ano99b, BR99, BS96, BGRD00, DS07, DRJSW97, MHE19, MST94, MSV15, ORS90, OZ02, PTK04, SHRN98]. Models [ABH18, BVC16, BVC19, ZDE95, AGL08, BH14, BK94, BG07, BÁJG^{+}20, BPP10, BWZ95, BPC21a, BPC21b, CBV13, CFW^{+}22, FF21, FRS^{+}88, GSE^{+}15, GK03, Gil94b, HK00, Hoe12, JMLBL98, KLS^{+}88, LD98, Liu86, LP92, MM96a, Mal02, MPZS13, MCB05, MTV08, MMTU18, MLM^{+}00, MSW98, NZHY11, RE98, RP85, Rob00, STP^{+}19, San02, SA16, SHB19, SVC19, TL96, Xu07, YMJT10, YNY^{+}17, Zla88]. moderate [Uhl96]. modern [BMYK98, CGH^{+}19, EYP^{+}20, GS22, Gil88, LT19, PCN10, TF98]. modified [CH92, CKLM14, OW90, WN91, YE94, YLCT18, GGFF93]. modifying [KJ07]. ModSAF [BG98]. MODTRAN [WLCG02, BRS^{+}08]. MODTRAN-based [BRS^{+}08]. Modular [Bar95, TO91]. modulation [BKK24]. moduli [Mas98, MC04]. moldable [LHR^{+}19, MMSS22]. Molecular [GM04b, N^{+}00, AFPG12, BL94a, BCA08, CCCP92, DG15, DMG^{+}04, GCC19, GM04a, GVH08, GHC04, HC05, KHN01, LBL95, LVC16, LBWR90, LHZ^{+}20, MT97, MO99, NC97, PJV00, PBTC89, SKE^{+}22, SSS92, SWW99, SVTSM^{+}22, STKA96, TJJ93, XRW^{+}19]. molecules [RD07]. moment [IS18]. moment-based [IS18]. moments [LCLL00]. monadic [BI07]. Mondriaan [BF06]. monitoring [AAB^{+}16, JCWP07, KNP97, Kre13, MCC04]. Monitors [BL94b]. mono [MPJ03]. mono-implicit [MPJ03]. monodirectional [BMT92]. monotonic [BPEL05]. Monte [Ken99, vS92, ACC^{+}88, BSH88, BLFT84, DP95, EEH^{+}19, FHL87, FHJ^{+}84, GPW^{+}08, HA97, HSFS14, LM00, mM94, MD88, MAM^{+}09, MB88, ÖW10, ZLM98]. morphology [LS90, VLSPL19]. most [HWW92, LS97]. motif [AED21]. motifs [FM15]. motions [FC05]. motivations [Gup99]. movement [BJV^{+}16, MV17, MS20, THK^{+}99]. moves [DW00]. moving [DAD11]. MP [Oed92, ARW93, Cal85, DH91, FSY88, Gur88, HL88, Hoc85b, HKN89, KN88, MSW91, Nag90, OL86, Pin91, RR97, SW91, Sea86, Tem89]. MP-1 [RR97]. MP/48 [Meu87, Nag88]. MP2 [Ber00]. MPAS [OAJ^{+}16]. MPAS-ocean [OAJ^{+}16]. MPC [BPJ22, Las02]. MPCD [HJS23]. MPDATA [WSR14]. MPEG [AALK01, BC04, NU05]. MPEG-2 [AALK01]. MPEG-4 [NU05]. MPI [AKB^{+}19, ADLL03, ADR^{+}05, AMC^{+}19, BPJ22, CAHT17, CRGM16, DSG17, DR18, DPFT19, DWS^{+}21, EYP^{+}20, FBD01, FLPG18, FL21, FCS^{+}19, GJMM18, GLDS96, GL97, GT07, Gro19, HMS^{+}19, HJB^{+}21, HG12, IDS16, JJM^{+}11, KTAB^{+}19, LK14, LFS^{+}19, LWZL24, LFL11, LGMdRA^{+}19, MV17, MSMC15a, MAS06, MRRP11, MMM13, MTW07, Nak05, PVK^{+}22, PGBF^{+}07, PD11, PHM^{+}22, Ram07, RMS^{+}18a, RGDM15, RGGP^{+}18, RSC^{+}19, STP^{+}19, SSN^{+}21, SC19, SBM^{+}22, SRS^{+}19, TSCS14, TGL02, TG09, Trä12, THMH21, VLSPL19, VGRS16, Wal01, WCC^{+}22, ZJDW18, ZCBD22, ZGZS20]. MPI-3 [FCS^{+}19]. MPI-based [TSCS14]. MPI-CUDA [DR18]. MPI-IO [TGL02]. MPI-thread [IDS16]. MPI/OpenMP [ADR^{+}05]. MPI T [HHK^{+}19]. MPICH2 [BMG07]. mpiecho [RGdS^{+}13]. MPP [Hoc94, SSP^{+}98, SWC99, XH96]. MPPs [MS03b]. MPSoC [MFS^{+}19]. MPSoCs [SSO^{+}14]. MR [HE88, PB11]. MRO [MMM13]. MRO-MPI [MMM13]. MRSG [KdBMA^{+}13]. MSPAI [HKR^{+}10]. Multi [BKL^{+}19, DGKF19, HHGA15, JS13, LCE^{+}18, LHG^{+}23, SHvAP13, SA18, WHL19, XT11, ZP16, ZL04, ABGR21, AAV08, ADEQO19, ACM20, ACGL04, BJ18, Bar95, BL99, BFL^{+}01, BKS19, BL94a, BD10, B^{+}02, BNS^{+}07, BNK15, BBL^{+}19, CRGR^{+}13, ÇFG^{+}12, CLA15, CC00a, CCW01, CLC^{+}18, CSB00, CDGI15, CIO^{+}17, DR95, Dod89, DFH^{+}13, DSSD18, DWL^{+}12, EMV^{+}18, EJLC00, FWL03, FMA17, Far94, FYH^{+}18, FT93, GZHX17, GCH21, Ger15, HSR^{+}14, HZY^{+}19, Hor93b, HRR08b, HF20, IM88, JLLC22, JAA06, JJM^{+}11, Kat01, KKH04, Kie91, KD10, LLX^{+}18, LYC^{+}19, LHL18, LXL^{+}22, LNA06, MCY^{+}24, MII^{+}11, MWI^{+}15, MBB^{+}18, MPZS13, MSW91, Müh90, NMAB19, NMMG13, PFS^{+}04, Pea19, POHS14, PB11, PL00, PTS^{+}12, RGT17, SWCBQ19, SC03, SFSV13, SPW^{+}15, SM90, SWR^{+}18, VLSPL19, VJ11, VK17, VA14, WZS^{+}14, WYX^{+}22, WAT20, WY11, Wor97]. multi [WLZ^{+}23, YLW^{+}13, vv91, LGD^{+}15]. multi- [CDGI15, MII^{+}11]. Multi-accelerator [HHGA15]. Multi-Agent [SA18]. multi-attribute [CRGR^{+}13]. Multi-Body [LCE^{+}18, CLC^{+}18, DSSD18, FYH^{+}18, LLX^{+}18, LHL18, MBB^{+}18, SWR^{+}18]. multi-cell [BL94a]. multi-channel [WLZ^{+}23]. multi-chip [NMMG13]. multi-cluster [JAA06]. multi-core [AAV08, BJ18, BNS^{+}07, ÇFG^{+}12, CLA15, CIO^{+}17, DFH^{+}13, Ger15, JJM^{+}11, KD10, LXL^{+}22, NMAB19, PB11, PTS^{+}12, RGT17, SWCBQ19, SFSV13, SPW^{+}15, VJ11, VA14, WZS^{+}14, WYX^{+}22, WY11]. multi-cores [MPZS13]. multi-CPU [ADEQO19, LGD^{+}15]. multi-CPU/GPU [ADEQO19]. multi-dimensional [B^{+}02, CC00a, CCW01, EJLC00, Far94]. Multi-GPU [SHvAP13, XT11, ABGR21, FMA17, GZHX17, HSR^{+}14, LGD^{+}15]. multi-GPUs [GCH21]. multi-grid [DR95]. multi-head [Wor97]. multi-improvement [ACM20]. multi-job [MSW91]. multi-layer [FT93, LHG^{+}23, Müh90, PFS^{+}04, SM90].

Multi-level

[JS13, LHG^{+}23, Hor93b, HRR08b].

multi-linear [Bar95]. multi-morphology [VLSPL19]. Multi-Objective [DGKF19, HF20, LNA06, VK17]. multi-parallel [BNK15]. multi-partitioning [BFL^{+}01]. Multi-phase [ZL04]. multi-phase-field [WAT20]. Multi-Physics [BKL^{+}19, WHL19, BKS19, BBL^{+}19, HZY^{+}19, LYC^{+}19, Pea19]. multi-platform [DWL^{+}12]. multi-point [POHS14]. multi-point-boundary-value [Kie91]. multi-processor [Kat01, YLW^{+}13]. multi-reference [PL00]. multi-round [BD10]. multi-scale [MWI^{+}15]. multi-scattering [BL99]. multi-storage [SC03]. Multi-tenancy [ZP16]. multi-tenant [MCY^{+}24]. multi-threaded [CSB00, KKH04]. multi-threading [FWL03]. multi-tree [EMV^{+}18]. multi-user [vv91]. multi-vector [IM88]. multi-vehicle [ACGL04]. multi-view [JLLC22]. multiblock [DLR94, Ran00, di 97]. multibody [FP98, GIF^{+}10]. MultiBoson [dF99]. Multicast [CWC00, ACLN03, GLG06, KEEF01, LBH07, Miš98, NLP^{+}15]. multicasting [XCS05]. MultiCL [APBcF16]. multicommodity [GPS03]. multicomputer [BF95, BR99, CGM^{+}92, CF90, Don05, DG94, Kam87, MO89, MRT93, NA08, RVGG01, TTO^{+}22]. multicomputers [AC94a, AC94b, CDR^{+}95, DHL07, GPC88, Lee95, Lee97, MM96b, SH01, WW95, YD01]. multiconfigurable [Tap84b]. multicore [ACH^{+}18, AART13, BM08, BLKD09, CHQORS18a, CHQORS18b, CS19, EL16, LLW^{+}15, LT19, WOV^{+}09, YMJT10, YMG14]. Multicores [ABH18, ADK22, BH14, OK22, SHB19]. multidimensional [AKSS07, DLM97, FA96, NDN20, SS92, TN13]. multifrontal [AGL08, ADGS10, Duf86, GS22, GLU03, LSL14, YWP11]. multigrain [MS03b]. Multigrid [LS93, Ale91, Ale94, BWV^{+}17, DL04, FBMV88, FA11, GAR15, GZ99, HKL01, HJ97b, HM86, HK92b, IHM^{+}12, KS01, LSS88, LPMD01, Oli96, PMLT03, RBMO20, RSSS94, SS90, TCL92, di 97]. multigrid-extrapolation [HK92b]. Multilevel [HSS09, ABG^{+}06, CLA14, HS90, HM89a, MQ89, MPD00, PER17, WC00, XKL^{+}22]. multimedia [AFT23, Che98b, DCD97, HXW^{+}13]. multimicroprocessor [Kas85]. multiobjective [DST15, dTNOR^{+}04]. multioutput [NAC^{+}14]. multipath [BKK24]. multiphase [OCE^{+}07]. multiphysics [DAG^{+}09, PIG^{+}16]. multipipelines [Cho91]. multiple [ALH^{+}14, BCRSR11, BDK06, BJV^{+}16, CWK09, DGNP88, DLS09, DWS^{+}23, Dod90, GHP10, Gol88, GDMS97, Hoc89, Jia09, KMB^{+}18, Kie94, Nag88, OGC^{+}15, Qua92, SSP^{+}98, SSS^{+}05, Tak10, TNZM20, Ter10, VTmL03, VJ12, VK92, WP19]. multiple-precision [Tak10]. multiplexers [Lin93]. multiplexing [NSF^{+}22]. Multiplication [DS13, MJ94, ASA16, AMP92, BR89, Ber89, BVWH14, CS87, DEGS95, DTR18, FAF16, FOH87, GR01, GS96, GKV94, HRR08b, Joh93, Kak88, KAD09, LB23, Lin91a, LPPS01, LV15, ML20, Mar14, Nel93, Pet91, Ris90, RZ95, RR16, SMM90, TM90, VTmL03, WOV^{+}09, WRS12, YB11]. multiplications [MAH^{+}19]. multiplicative [MS04a]. Multipole [MKJC21a, MKJC21b, SSSG03, SWW99]. Multipole-based [SSSG03]. Multiprocessing [VGS14, GBH98, GO88, KI05, SG02a]. Multiprocessor [AD96, KRW88, Swa87, VLL90, AALK01, ABMT99, AGB97, AB93, ADMV05, ADE84, Axe86, AE93, BG91a, BA95, BC04, BDSd94, BS96, BFR93, CG93, CG87, CG89b, CG92, Cor00, DFRC94, Dzw91, ECG93, Eis89, ES88, EO91, FHKZ88, GP85, GD90, GSZ88, GHNL87, GV08, Hag91, Hat06, Her88, KW90, KD97, KHC92, KLN90, KM89, LH88, Lil94, Mah96, MRW^{+}08, MM00, MF16, MSS^{+}05, Mis87, MNS87, Nat90, Ng93, PJ84, PRV87, PZE94, Phi91, RRS88, RR89, RT88, RS88, RG92, She93, SJJ85, Sto89, TYKA95, Wor92, Wri91, WHW91, Y^{+}99, ZE92, ZG88]. multiprocessor-systems [Hag91]. multiprocessors [APRP97, BS97, BM01b, CG89a, CHLO85, CJ95, DFRC90, DT96, Eva90, Fie96, FM88, GYH94, Gre89, HM89a, HM86, HK02b, HB84, Joh93, Jor88, KS89, KK03, KV99, MMC97, MQ89, MKL^{+}88, PSS01, Pir93, Qui88, Rot95, SM91, Stp92, Sun97a, TO91, WW90, WW92, WE91b, WE92, Wol89]. multiprogramming [NL91]. multiresolution [GWLS05, NC97]. Multiscale [SLS^{+}21]. Multiscattering [Li94]. multiselection [Als01, OW91, She97]. multisplitting [Bai99, PS89a, ZDE95]. multistage [Blo03, HW94, Leu93, LMH08, Y^{+}99]. multitask [DST15]. Multitasking [FSY88, HKN89, Meu87, CCS87, DCG90, Nag88, Nag90]. Multithreaded [DTR18, AFD13, A^{+}01a, ÇFG^{+}12, DSG17, Evr01, LLL10, Mar14, NB12, Sod02, TG09, TDW03]. multivariable [KS91b]. Multivariate [LND^{+}19, dG98]. Munich [DSSD18]. MUPPET [MKL^{+}88]. MVAPICH [RMS^{+}18a]. MX [Gog11].

N [GZTR07]. N3S [CLL99]. N3S-NATUR [CLL99]. NAG [DMWW88]. nanoclients [Cla03]. nanoelectronics [LK10, WRB97]. Nanyang [WHL19, LHL18]. NASA [BJK^{+}17]. national [YK87]. NATUR [CLL99]. nature [ZEgT04]. nature-inspired [ZEgT04]. Navier [GV01, VvBv90, AIIV98, Bas94, DLR94, FKM^{+}22, LSS88, OJ90, PATC99, PCA01]. NCAR [Mic97, HRW^{+}95]. NCCL [AMC^{+}19]. NCCL2 [AMC^{+}19]. Ncube [Dun91, SV94]. near [GZH^{+}23, TDG^{+}18]. near-optimal [GZH^{+}23]. near-threshold [TDG^{+}18]. Nearest [Li91, BJP^{+}89, CAC^{+}09, DFM99, San99]. nearest-neighbour [BJP^{+}89]. nearly [VB95]. NEC [HCH^{+}96, Wat87]. negating [RH13]. neighbor [DFM99, DL06, EY86, Li91, San99, TTH95, XRW^{+}19]. neighbor-joining [DL06]. Neighborhood [OID^{+}12]. neighbors [CAC^{+}09]. neighbour [BJP^{+}89, Zam99]. NekRS [FKM^{+}22]. Nemesis [BMG07]. nephrons [BJP^{+}89]. nervous [WS04]. nested [BPK12, BS05, CCGG14, Con90, HCT16, MCB16, MSZM14, THH^{+}05, WW92, Xue96, Xue97]. NestedMP [HCT16]. nesting [YNY^{+}17]. nests [DR94, DR96, GGL99]. net [Mak94b, MS89, WN91, WH94]. nets [CFC99]. NetSolve [CD98]. Network [ACLN03, EG90, KBG^{+}01, Ann89, AFK01, AED21, BM08, Bar97, BCRSR11, BP97, BWT^{+}08, CB17, CZ91, CC02b, CF04, CH92, CD00, DT95, Dzw91, El 93, EC91, EB94, FRS^{+}88, FA96, FHKZ88, FCM03, GS04, GDMS97, HWYL89, KYS23, Kal90, KS05, KRN^{+}11, KV99, KTN^{+}14, Kol94, KJ07, kLH95, LX00, LLL10, LH00, MM96a, MS03b, NTHY22, NZ92, PMS^{+}13, PYLE21a, PYLE21b, PVBR23, PJW^{+}22, RGBC11, SS95, SKY93, SOS97, SVC07, She93, SKC90, SS01, Sun97b, Tak06, TMCC02, Tou02, THK^{+}99, WC91, WH91, WH93, WS94, WLZ^{+}23, YMJT10, Y^{+}99, YLT^{+}23, ZM88a]. network-based [KYS23]. network-embedded [SVC07]. Networks [DZLK20, AK89a, AQ93, ACM^{+}10, ACM^{+}15, ACLN03, AFJG06, BG91b, BGL^{+}88, CWC00, CHH^{+}01, CFS01, CWK09, CB97, CB09, CMM^{+}88b, DG95a, DJH05, Dod89, EHR^{+}98, EAKT90, FK98, FSKF06, HMTX93, Han98, HP97, HCK94, HGS10, HK90a, HP91, HW94, JAH^{+}18, Ken90, KL90b, LJD93, Lan90, Las02, LS98, Leu93, LK00, LXL^{+}22, MT91, ME92a, Mar95, MS03b, Miš98, Müh90, NSH^{+}21, NMARD10, ORS90, PGGL17, PW22, PVP08, RD07, RGBCS09, RGBC11, Sab97, SYAU07, SH90, SAOKZ06, SVS01, SVS02, Sha06, SHCS86, SF91, Shi01, Sin90, SS90, SHCD00, SM87, TY91a, TO91, VP95, WBPM97, WLYJ13, WC15, Wel89, WSB90, dV94, da 90, dG98, AAS13]. Networks-on-Chip [AAS13]. networks-on-chips [WLYJ13]. networks-steps [Müh90]. Neumann [GW87, PJW^{+}22]. Neural [CFC99, CB97, DZLK20, FRS^{+}88, Ken90, CC02b, CH92, DJH05, EPMPU02, EAKT90, Han98, JAH^{+}18, KRN^{+}11, KL90b, Lan90, LXL^{+}22, MM96a, ME92a, Mar95, Müh90, ORS90, PMS^{+}13, PYLE21a, PYLE21b, PYZC11, PJW^{+}22, SF91, Sin90, TO91, WC15, WSB90, dV94]. neuromimetic [NFG^{+}13]. neutral [CDM^{+}10]. neutron [FBMV88, MD88, TS21]. Newton [BBQOQO00, CH98, TLC^{+}21]. Next [BDO17, EYP^{+}20, Hoe12, PVBR23, Sun97b, Zen99]. Next-Generation [BDO17, Hoe12, PVBR23]. Nial [GJMM89]. Nicolson [ZL94]. no [HE88, FC18]. NoC [JMA^{+}13, NMMG13]. NoC-based [NMMG13]. Node [GP98, Reg01, SHCD00, BB21, GS04, Gro19, LFL11, LBGO23, PS91, SS04]. node-aware [LBGO23]. Node-disjoint [SHCD00]. Node-ejection [Reg01]. Node-to-set [GP98]. nodes [BBB^{+}22, FMA17, Fig06, GP96, GGL^{+}05, HJS23, PTGF20]. noise [SML^{+}14]. Non [CC97a, AFS14, AIV95, BPEL05, CLHL23, DF00, GW08, GDAK06, Haa98, HGLR07, JLY18, LSS05, Lan96b, MSE07, NAC^{+}14, STP^{+}19, SK19, SMK20, Ver99, WP19, XLY^{+}20, Xue94, Xue96, Xue97]. non-binary [AFS14]. non-blocking [HGLR07, STP^{+}19]. non-convex [CLHL23, LSS05, Xue96]. non-dedicated [GW08]. non-equilibrium [DF00, XLY^{+}20]. non-Hermitian [WP19]. non-linear [AIV95]. non-monotonic [BPEL05]. non-overlapping [Haa98, Lan96b]. non-perfectly [Xue97]. non-rectangular [GDAK06]. non-symmetric [JLY18, SK19, SMK20]. non-trivial [NAC^{+}14]. Non-uniform [CC97a, Ver99]. non-uniformly [MSE07]. non-unimodular [Xue94]. non-zero [Ver99]. Noncommittal [Nic95]. noncontiguous [TGL02]. nondeterminism [MMTU18]. Nonlinear [Amm89, MTK03a, BE92, CZ91, DDF^{+}10, ED91, GYH94, Gol88, HJ97a, IS18, LR88, Pir93, Pir96, Sch95, Sch87, SK87, dP90b]. nonparaxial [SCCPM14]. nonrigid [IOH05]. nonsymmetric [BDK95, GMW96, KS89, RR89]. Noor [Kra84]. Noor-Lambiotte [Kra84]. normal [JW09, NW96]. normalized [YS96]. norms [BM93, CT88]. northeastern [BBL^{+}19, MBB^{+}18]. NoSQL [ZP16]. notation [Num05]. notational [Bek95a]. Note [Ano19m, CARW91, AE86, Con89, DG90, GP90, Jan87, Jou04, SS91b, WMCU97]. notification [SSN^{+}21]. Notified [FD18]. Novel [SYP13, VA14, AQ93, AART13, BCPB05, DL05, GTW^{+}20, LNLK13, LH00, MHK97, QJ06, RBB^{+}22, ST21, VL05, Wan09, YZWcF14]. November [Ano19l, Ano20n, Ano23g]. NOWs [MT00]. NP [KB20]. NP-completeness [KB20]. NPDP [PB23]. NRL [WM97]. NTHU [LYC^{+}19, CLC^{+}18]. nuclear [HSFS14, TBM^{+}16b]. Nullspace [JTS^{+}11]. NUMA [ABB^{+}16, CCGG14, Car98, HK02b, MCB16]. number [APG92, CRD02, DP90a, Deá90, Lah00, Mak94a, SMC03, Ume89, dP90b]. numbers [AGH^{+}94, MM95]. Numerical [AFPG12, CDGI15, DG95b, DNL15, DR03, Gen84, HB94, JCC^{+}24, Luc01, LK10, Zha95, AAHF97, Ano99b, ADTV01, Bro89, Bur90, DAA94, DM90, ER89, GM07, Gup99, HJ97b, HB84, HHJP16, KKHY06b, LP92, MTK03b, Num05, PATC99, Pir93, Pir96, RRA11, Rod85, TKC^{+}14, XRW^{+}19, YLCT18, ZPAT99, YD97]. Nürnberg [DSSD18]. NVIDIA [LLX^{+}18, MBB^{+}18, SRK^{+}21, SWR^{+}18, CLC^{+}18]. NVM [ZWJ^{+}19]. NWChemEx [WYBdJ^{+}21]. NX [Pie94].

O [Bak00, BPC21a, BPC21b, CDZ^{+}21, DBI^{+}17, HMP^{+}16, MV17, MQ97, PUSS97, SDMS12, SC03, TDC19, TF98, WG97, YM05, ZS21, ZJDW18]. O-aware [ZYZ^{+}16]. OASys [VKH99]. Obituary [Ano03k]. object [AAA16, BV96, CKRZ98, CLZ01, FWL03, Fer97, RW89, SW03, TDC19]. object-based [TDC19]. object-oriented [BV96, Fer97, RW89]. Objective [DGKF19, BSD11, HF20, LNA06, MMT06, VK17]. objects [BE97, HK98, Jes88]. Oblivious [LC99b, YB11]. observation [AC03]. Observations [LNA06, SSY10]. observed [CAHT17]. OCamlP3l [CMV^{+}06]. OCCAM [Jes88]. occluded [IR00]. Ocean [BDOS95, KSS98, OAJ^{+}16, WM97]. October [Ano20o, Ano21q, Ano22k]. Octopus [GSC^{+}22]. Octopus-DF [GSC^{+}22]. octrees [HA05, HAK^{+}21]. odd [BE93b, MP87]. odd-even [BE93b, MP87]. ODE [GH89, Kie91, PG93, RM97]. ODEs [MPJ03]. OF-WFBP [GZH^{+}23]. off [AALK01, CLJ14, DW99, FC18, LDK16]. off-line [AALK01, CLJ14, LDK16]. Offloading [DFP^{+}19, TSEE21]. oil [BBK90, YLCT18, Zha95]. oil-water [YLCT18]. OLAP [RBB^{+}22]. old [ES88, LK14]. omega [SKY93, GS10]. OmpSs [YA24]. on- [CLJ14]. On-Chip [SYP13, AFY^{+}16, LNLK13, PMS^{+}13, Tak06]. on-demand [JSC97]. On-line [Leu93]. on-node [BB21]. One [BT99, CFMM97, CC92, DPFT19, Don05, Hee97, RVGG01, ST02, Ver00, YCBD19]. one-dimensional [CFMM97]. one-port [ST02]. one-sided [DPFT19, RVGG01, YCBD19]. one-step [Don05]. One-to-all [BT99]. ones [Che96]. Online [CLY^{+}19, SMJ15, DCN^{+}15, WNES01, YNY^{+}17]. onto [DR94, DT96, MM96a, MT91, Pop88, RV96, SER90, SM91, TMCC02, WMCU97]. open [IR93, Gog11, VGRS16]. Open-MX [Gog11]. OpenACC [LLVM21a, LLVM21b, LHZ^{+}20, LHZ^{+}22, WLK^{+}18]. OpenACC-to-FPGA [LLVM21a, LLVM21b]. OpenCL [APBcF16, BAS13, DWL^{+}12, GScFM13, SFSV13, SBM^{+}22, TSEE21]. OpenCL-like [TSEE21]. openFabrics [FCS^{+}19]. OpenFPGA [WPS^{+}08]. OpenGL [LHZ97]. OpenGR [HST05]. OpenMP [ADK22, ATL^{+}12, ADMV05, ADR^{+}05, BBB^{+}22, BS05, BC19b, CM05, DFP^{+}19, FSG19, HP05, HAJK01, HCL05, IJM^{+}05, JKHK08, JJM^{+}11, MCB05, MRRP11, Nak05, NU05, PK05, THH^{+}05, VLSPL19, VGS14, SSGF00]. OpenMP-like [VGS14]. OpenMP/MPI [Nak05]. OpenSHMEM [HVA^{+}16]. operating [GHS02, JBWE14, Sch88b]. operation [Sho17]. Operational [NOG^{+}22, BRS^{+}08, EAR93]. operations [BMG07, CIO^{+}17, DS13, EG95, GGV04, HHK95, HGLR07, HMS^{+}19, Hwa02, IDS16, IZS^{+}20, KS95, LC17, MV87a, ME89, ORM^{+}10, PGW16, TKC^{+}14, THK^{+}99]. operator [FHMS93]. operators [BM02]. Opportunities [BJK^{+}17, Sch95]. Optical [EM88, RBP^{+}17, Kat01, LPH00, PVBR23]. Optimal [BDK98, Dow90, Fig06, HK90b, IR00, KH95, MR89, Mar92, NM01, OS21, PW22, RRP03, ST02, She97, Als01, BE93a, CS95, Che08, CL05, CMP92, DM04, DZD01, Dzw91, FIMF99, GZH^{+}23, GAMR03, HT94, HHQ06, KR10, LLW^{+}15, LC90, LO92, MMS90, MRSB94, MB00, OWZ91, RE92, SSY10, SM88, SB92b, SH01, TT00, TF98, TL94, WL14, Wri90, WLZ^{+}23, YTCW07, Zub90]. Optimisation [FMM^{+}02, Dod89, GS10, JOL^{+}21, KHC92, dTNOR^{+}04, WC00]. Optimising [Sil88]. Optimization [CLSM98, CDH^{+}03, EEH^{+}19, FBG^{+}12, HW91, KRS^{+}21, LLVM21a, LLVM21b, LM06, WW95, WOV^{+}09, ARCH05, ALNT04, ALT04, AAB^{+}06, BP95, BCMG^{+}07, CZ91, CCL04, DF91, DFO93, DCD97, DGKF19, ES89a, GZHX17, GÁVRRL18, HF20, LLPV06, LTG23, LW11, LLW^{+}24, LHZ^{+}22, LR88, LNA06, MII^{+}11, MAH^{+}19, MSSC10, MTK03a, MTK03b, MGSK88, MA22, NMAB19, NS11, NSH^{+}21, NTHY22, NBB^{+}02, PB98, PYLE21a, PYLE21b, PGK^{+}00, PŚ00, PHCR05, RKS92, SYAU07, SH90, Sch95, SLH^{+}13a, She96, ST21, SMP11, STE23, SPK18, SWR^{+}13, SSBT19, SS92, ŚDJ^{+}22, TCT00, WWR05, WS12, XLY^{+}20, ZM88a, VK17]. Optimizations [KMK^{+}19, LKGD16, LXC^{+}18, MV17, WPD01]. optimize [AKSS07, ORM^{+}10, RMACG10, WZL09]. Optimized [AFS14, AMC^{+}19, PM95, Bar12, DPSW00, HBC19, LDK16, MMM13, ND95, SB92a, YCBD19]. optimizer [MSB91, YLE95]. Optimizing [CFN03, DJH05, EGKU02, HGLR07, JHD^{+}22, KI05, KAD09, LXL^{+}22, SM87, TGL02, TS09, VCK^{+}11, ZJD^{+}23, BLCR21a, BLCR21b, DEH^{+}11, GF89, GK19, PB23, Str87, WSB90]. optimum [ML00, UN87, Ume89]. Option [FMM10, Ger04, PGK^{+}00, SÖB07, Sur10]. options [Pea19]. OR-parallel [MBC97]. Orchestrating [CCS^{+}18]. Order [SS95, AND05, DCD97, DLS09, DWS^{+}23, DK04, GGL21, HA14, KD97, KHC92, LT94, MS00, ME90c, PVK^{+}22, PS84, SH90, TL90, TL94]. order-invariant [HA14]. ordered [CB09, OIK21, TS88, Ver99]. Ordering [DW99, BOV02, BMCA98, CP08, GM03, Rag97]. orderings [Leu89]. orderly [WL98]. orders [MS99b]. ordinary [EM87, GR89, Gol88]. Organisational [Par87b]. organization [HH91, KV99, SY87, da 90]. organizations [CS89a]. organizing [ORS90, WHW91]. orientation [Sch91a]. Oriented [DWH^{+}08, AAA16, BV96, Fer97, GBH98, GJMMS97, KAM^{+}20, LKHL03, PB16, RW89, VvBv90, HL13]. Origin2000 [WBS06]. Orthogonal [Kuz98, ACH^{+}18, EK92, Hot89, ME90b, Par91, SS91b, Stp92]. orthogonalisation [WC91]. orthogonality [THMH21]. orthogonalization [CSW01]. OT [PUSS97]. other [Ale94, ESV10, KLS^{+}88]. OTIS [Jan06]. OTIS-mesh [Jan06]. Ouragan [C^{+}01a]. Out-of-core [CKS03, AGL08, CU04, CN98, DHW97, Ger15, KCRB98]. out-of-order [KD97]. out-tree [HK02a]. outer [EM88, Rod85]. Output [MMTU18, AS86, BBDN21, CDRV97, EG95]. outtrees [Ver00]. over-relaxation [PJ84]. overcome [DW99]. Overcoming [FC18, JKHK08]. OVERFLOW [Taf01]. OVERFLOW-MLP [Taf01]. Overhead [DFP^{+}19, Cal96, CRGM16, KP96, MW91, Trä18, XXQ^{+}15]. overheads [Hoc89]. overlap [DVGG98, KGS08]. overlapped [BMP15, EJLC97, KH97, mMvdV01]. Overlapping [WB88, CWK09, Haa98, Lan96b, MMM13]. Overlay [TF98]. overrelaxation [ESK88]. oversubscribed [MKK^{+}19]. Overview [Alm85, Ohb85, HHK^{+}19, Hey94, IJCL96, McB94a, SG02b, Tre88]. OWL [QH19].

P [FD02, MRJ89]. P100 [CLC^{+}18]. P2P [CRGR^{+}13, MTV08]. P2P-based [CRGR^{+}13]. P2S2 [BVC19, SVC19, VBC19]. P3M [YSBM97]. p4 [BL94b]. package [LF00, RS88, Ste88, XKL^{+}22, YSBM97]. packages [WLN^{+}96]. packet [Ann89, ALTZ02, Lin93, Uhl96]. PACS [A^{+}99, NNB^{+}99, Uka99]. PAD [SN14]. Padé [MKC92]. Paderborn [BJvOR03]. page [CDA^{+}16]. pair [HGM20]. pairs [GS04, GT90, SA13]. pairwise [DEH^{+}11]. Pajé [CSB00]. pancake [AQ93, KF95]. Panconnectivity [MLX07]. Pancyclicity [HC04, MLX07]. paper [FLW87, HE88]. Papers [Ano22m, Ano15a, Ano18k, GT19, MTW07, TH20]. PAPS [WH97]. PAR-Bench [NL91]. parabolic [AAHF97, AIV95, BE92, HK92b, KR97, Lop93, Pir96]. paradigm [AD96, BA02, BM02, BPP10, JON08, MNS87, OID^{+}12, PWY03, Sto89, ZLD15, da 90]. paradigms [ONB11, ZEgT04]. ParadisEO [CMT04a]. parafoil [TKG97]. Paragon [BEK95b, DG95b, GAW96, Hoc94, SNS^{+}97, XH96]. Paragon/XP [SNS^{+}97]. Paralldroid [AAA16]. Parallel [AB02, Adr99, ADF93, AND05, ABR03, AFPG12, ALH^{+}14, Ala89, Ala95, ALNT04, ALT04, AAHF97, Ano87, Ano88b, AQO18, AHBD18, AATK10, AS00, ACGL04, ABB^{+}11, AIV95, AW13, BD89a, BVC16, BVC19, BM01a, Bar97, Beb97, BM09, BE93b, BG02, BDG^{+}10, BGSS14, BG97, BGMT00, BPS01, BC91, BP00, BMCA98, BC04, BSK03, BSB^{+}22, BJP^{+}89, BETR17, BL93b, BCYB11, BH99, BP02, CLS93, CTS02, CGGG03, CDZ^{+}98, CJ97, CF93, CS95, CHR00a, CL02, CDW95, CS96, CW97, CGQ93, CMT96, Con90, Cor99, CLZ01, CGMM99, CMRT88, CKM93, DS07, DM03, DS96, DDP90a, DP00, DCG^{+}07, DFRC90, DBVS01, DL04, DFRR91, DF95a, DAD11, Duf86, ER18, ECLV12, Eva84, EW93, Fah85, FF21, Fer97, FMS01, FHMS93, FPT91, FC05, GYH94, GMBZ90, GK03, GH90]. Parallel [Gil94b, GHW08, God00, GGL21, Gra91, GZ99, GPS^{+}08, GPSK09, GK04, Haa98, HKL01, HKK97b, HE88, HLP10, Heb93, HK92a, HJ97a, Hen93, HRT07, HFL^{+}10, Hil92, HZ93, HT92, HM86, HCK94, HM01, HWW92, HK91, JLLC22, Jeo91b, JCL92, Jou97, JNWJ18, JNC^{+}19, KR97, KB85, KT97, KK11, KKU16, KBBC88, KH92, Kie94, Kim90, Koç97, KS91a, KKŽ05, KHN01, KC93b, KNR00, KŠR04b, KS01, KPH97, KD10, KKB93, LSS05, LRH97, Lan96a, LPM11, Lan96b, LNS03, LM03, LPCA98, kLH95, LHK^{+}96, Lef97, LF89, Li89, LHZ97, LM00, LLW^{+}24, Lin90, LL90, Lin93, LSS88, Lip99a, LP92, LR16, LHK15, MS03a, MG95, MKS91, Mal02, Man01a, MKL^{+}01, MAFC14, Mas98, MC04, Mat95, MT97, MHT06, MS03b, MKJC21a, MKJC21b, MT00, MMT06, Mel87, Mel88, MP08a, MP08b, MP02]. Parallel [Mic90, MTK03b, Mil01, MCG98, MSW91, M^{+}00, MS90, MZ95, NR01, Nak05, BHB21, dTNOR^{+}04, NBB^{+}02, Num05, O’L87, OW90, Obe85, OBG00, Oli96, Ols95, OR88, OA91, ÖS94, OCPT97, OZ02, PPZ93, PAF^{+}97, PM03, PM89, PS89a, Par86, PUSS97, PSS94, Pet84, Pet97, PB94, PHD16, PGW16, Qui88, RR89, RSB05, Rag97, Ral03, RS08, RWT97, RP85, RB03, RA20, RLH19, RST11, RO88, RG92, SÖB07, SS02, SRH07, SXBD97, SOS97, SK97a, Sch91a, SSS92, SAWH88, SVS02, SCD92, SZ02, SYP13, SFB^{+}97, SB95, SVC19, Ski02, SM15, Stp92, Str08, Str92, Sun97a, Sur10, Tak10, TS02, TTH95, TY91a, TKG97, TE92, Tre88, TL96, TJJ93, UZ02, Uml94, Vaj84, VP94, VIHD90, Wag89, WC00, WC91, Wat00, WBD99a, Wei93, WS04]. Parallel [WG97, Wri91, YK08, YPG03, YTCW07, YD97, Yue97, ZRPI89, ZE92, ZSH97, Zom95, ZEgT04, de 87, dGR95, di 97, mMvdV01, AK89a, ASA16, ABGR21, AO89, AGH^{+}94, AJ95, AGL08, AAGS18, AHL02, APBcF16, ARCH05, AAB^{+}06, ADEQO19, Alm85, Als01, APG92, ATA^{+}16, ADPV03, AGLP06, ADGS10, AM93, AB95, AMP92, AKSS07, AC00, AKK99, ACH98, ABG^{+}06, AT98, ADTV01, AFK01, AKP^{+}98, AKNS91, AV15, AGGG06, ABH^{+}10, AHW14, ASH92, AGGM90, AIIV98, ADMV05, AE93, BSH88, BM08, BE95, Bai99, BMS97, BS00a, Bak00, BBB^{+}94, BCGS93, BPEL05, BEY00, BDF^{+}00, BN01, BKT91, BMS01a, BGWW97, Ban97, BKS06, BO91, BM11, Bar12, BHK89, BDI^{+}95, BGV97, BW92, BRS97, BFL^{+}01, BV96, Bas94, BOV02, BO03, BOVG10, BE92, BA02, BHM^{+}95, BAR98, BBQOQO00, BMT92, Ber00, BE88]. parallel [BDK95, B^{+}02, BM02, BF06, BHK00, Bjø87, BGL^{+}04, BDOS95, BFM90, Blo03, BE93c, BHS99, BL93a, BDK96, BM01b, BFH23, BRH90, Bon91, BEK95b, BDH95, BGM03, BT01, Box92, BGR00, BP86, B^{+}00, Bri95, Bro89, Bro99, BNK15, BWZ95, Bro86, Bru91, BK97, BS00b, BL94b, BGL01, BLKD09, CMT04a, CZTS99, CYDJ21, CE94, CGK^{+}16, Cal96, CMT04b, CDDG93, CS00, CDR^{+}95, CS93, Cap88, CCEJ01, CH98, CS89a, CU04, CLS^{+}16, CDM06, CLL99, CPdM^{+}97, CM95, CCCP92, CGG01, Cha99, CCS03, CDBL08, CC96b, CUSR88, CCS94, CC02b, CCL04, CSB00, CC00b, Cv98, CMPM^{+}15, CJ92, Che98a, CLSM98, CFN03, CLA14, CCS^{+}18, CLHL23, CP08, CC97b, CC00c, CSH^{+}11, CFW^{+}22, CNK93, C^{+}01b, Chr98, CC99a, CQ97, CDC^{+}87, CWB92, Col04, CDGI15, CGM^{+}92, CDGM96]. parallel [Con89, CV17, CV02, CH97, CN98, Cor00, CKS03, CGOP03, CDM^{+}10, CQ01, CWDG07, CDH^{+}03, CR90, CRSS99, Cro97, CGM05, CCH^{+}21, DDdS02, DDdSL02, DCY^{+}22, DY91, DR18, De 95, DFR92, DFO93, DCD97, DR93, DLR94, DR95, DP90a, DP95, Deá90, DEGS95, DFRC94, DRJSW97, DAA94, DLS09, DJH05, DF00, DPSW00, DWM^{+}01, DF98, DH84, DS84, DS87, DJ87, DHW97, DT97, DFM^{+}95, DF95b, DG05, DL06, DZG94, DAG^{+}09, DvdV99, DS05, DR02, DW00, DR03, DGKF19, DZD01, Eck90, EGTD99, ED04, EK92, Eis89, Ekl04, EHF^{+}97, Ema10, EG92a, ES89a, EPMPU02, EM85, EY86, ESK88, ER89, ES89c, ES89b, EY90, Eva90, EL90, ED91, Eva91, EG94a, FSCL06, FHL87, FLYL16, FLYL21, FCC16, Feo88, FAS02, FM15, Fig06, Fis90, FP98, FK89, FGHB94, FJPA17, FG01]. parallel [FR95, FMS^{+}06, FMD98, FBAB13, FP92, FJS85, Fre89, FIMF99, FT93, GKT^{+}15, GS89, GP90, GWLS05, GZHX17, GL90, GV01, GGNY11, GGL99, GWWM09, GLS01, GPS03, GSZ88, GLN89, Ger04, Ger98, Ger94, GR01, GWC^{+}99, GJMM89, Glo95a, Glo95b, GKAS91, GdCCS03, GTW^{+}20, GÁVRRL18, GEvGCP09, GYL00, GG10, GW08, GGFF93, GKS98, Gro87, GKKS01, GL99, GT90, GM03, GHC04, GPPS99, GLS99, Güs99, GKV94, GPZ08, Hac89, HS90, HKSK97, HLDS95, HSS07, HT00, HLvHA11, HISS92, HMTX93, Han98, HA05, HKK97a, HCJ03, HW02, HCT16, HP97, HCH^{+}96, HM89b, HK00, HRR02, HJ97b, HDH97, HW91, HP94, Hio96, Hir86, HJ97c, Hoe12, HS03, HS89, HK98, HT94, HM89c, HK92b, HST87, HTB01, HK90b, HP91, HR04, HHQ06, HWL^{+}22, HKR^{+}10, HRR08b, HF20, Hwa04, HSN89, IR93]. parallel [IR00, IS18, IB01, IOH05, IHM^{+}12, JLY18, JC94, JG93, JHSV02, JK91, Jeo91a, JJM^{+}11, JIC96, JKH95, JON08, Jor86, JDG02, KCD^{+}97, KNP97, Kal92, KCRB98, Kao08, KMLM97, KKW14, KG87, Kas85, KY98, KK01, KKHY06a, Kat03, Ken99, KL90a, KC91, KY94, Kim97, KTN^{+}14, KG03, Kol94, KR00, KSS98, KCN99, KST02, KC97, Kot97, KP93, KS99, KS07, KKB92, KBD93, KS95, KYLH01, KSS06, KLS^{+}88, Kut85, LD99, Lan90, LAHM14, Las02, LR04, LLP00, Lau93, LAS90, LC02, LyHW^{+}22, LF98, LSA^{+}95, LMJC96, Leu89, LBWR90, LLS^{+}93, Li97, LMG09, LW11, LLW^{+}15, LHG^{+}23, LMC05, LC90, LO92, LC99a, Lin01, Lin08, Lip99b, Liu86, Liu89, Liu98, LEH14, LW15, LLFZ23, LTV96, LPMD01, LZCT15, LS92, LR88, LGMdRA^{+}19, Lop93, LC97, LG09, LQ92, LMH08, mM94, MAB^{+}02, MKK^{+}19, MV99]. parallel [MP92, MMC97, MD04, MS98, MAB17, ME84, Mak94a, Mak94b, ML23, MS00, ME93a, MGCB^{+}10, Mar92, MS04a, MJ94, MSP93, MWH95, MW11, MV87a, MIA^{+}07, MV87b, Mei85, Mei86, Mic97, MS99a, MJ99, MTK03a, MS04b, MK97, MBK12, MRRP11, MSW98, Mis87, MX07, MSE95, MS99b, MK12, MOW95, MSOCG^{+}16, Moo04, MTPE90, ME92c, MGT^{+}13, MGSK87, MSB91, MPJ03, MW91, MS89, MO99, MA22, NHS^{+}95, NL91, NS94, NC97, NYFK06, NNB^{+}99, Nat90, NAC^{+}14, NS02, NW96, NA08, NJ02, NZ92, NZ97, NK97, NRN00, NP05, Not95, NE01, NH11, ORS90, Ohb85, OTI^{+}89, OV06, OW91, OCE^{+}07, Ori10, OHZ98, OW94, OW97, PLP97, Pap98, PG91, PG93, Par87a, Par91, PW84, Par87b, PTGF20, PATC99, PJV00, PYZC11, PŚ00, PCN10, Pin91]. parallel [Pir96, Pis92, PVP08, PW87, PS06, PKR00, PP98, PMLT03, Psa02, PWM00, QH19, RM91, Ram07, Ran97, RM97, RW96, Rea90, RH99, Reg01, RvG05, RJ97, RBMO20, RR97, R^{+}00, Reu99, RE92, RR07, RBM^{+}16, RMACG10, RSSS94, RT88, RHM^{+}88, RRA11, RRS^{+}00, RW98, RKMJ10, Ron84, RV95, RBL97, RR16, RW89, SS89a, SER90, SNS^{+}97, SHPA05, SSY02, SCCP08, SCCPM14, San02, SKN04, SMT^{+}04, SAOKZ06, SSB12, SB94a, SAGV21, SV94, Sch95, SHN12, SSL19, SP94, SMK20, SKG02, Sel95, SR14, SG02b, SHH^{+}97, SK87, She95, She97, SZS22, SLY90, SMP11, Shy90, SMJ15, SSL03, SV97, SB97, SPMB23, SR97, ST95, SH88, Sne88, SGP22, SR17, SS87, SMC03, SB92a, Ste13, Ste87, SS90, Ste94, SE87, SM88, Stp93, SS92, SP93, SMTT96]. parallel [SR10, Sug99, SE98, SG91, Sun95, SG99, SB92b, STKA96, Tak03, THG98, TDC19, TNZM20, TTO^{+}22, Ter13, TO99, TS99, TMD^{+}97, Thu92, TTH09, TDW^{+}04, TLC^{+}21, THK14, TSJB00, TCT00, TLS16, TF98, TJ97, TO89, TL94, TBM^{+}16b, Tyr90, UAR^{+}99, Uhl96, Van02, VBS^{+}15, VL05, VP92, Vio04, VK92, VKH99, VA14, WH97, Wak04, WL13a, WN91, WRB97, WLK97, WLCG02, WWR05, WFW^{+}11, WB97, Wen95, WC90, WE91a, WE92, WHW^{+}11, Wil88, WvR16, WOKH96, WZ90, Wit98, Wol89, WAB89, W^{+}99, Wor97, Wri90, WP19, WTLW23, XDZL10, XRW^{+}19, XKL^{+}22, XZYQ21, YNO97, YORO08, YTSI13, YS96, YR18, YXQ^{+}21, Yan04, YCS07, YD07, YA24, YZWcF14, YEC97, YM05, YL19, YLT^{+}23, ZLT19, ZLM98, ZCPT00, ZZ03, ZPAT99, ZY16, Zer90, ZK92, Zha91, ZM94, ZJDW18, ZJD^{+}23, ZDE95, ZG16]. parallel [ZGL^{+}19, ZM96, ZVWS88, ZG88, ZGG92, dV94, de 96, tV96, van87a, van87b, Ano94, AGPS03, ASSS11, AGKS15, AGKS16, A^{+}01b, BG01, CGNR06, DC15, GRP22, GV11, SE93, Ter10, TF01, vS92, FLW87, YFK03, MBC97, Ben24, MSS19]. Parallel-architecture-directed [SCD92]. parallel-in-time [BFH23, SSL19, SR17]. parallel-independent [RvG05]. parallel/distributed [BGWW97]. Parallelisation [DWB99, TB03, EJLC97, EJLC00, GH89, Has01, IJCL96, LMJC96]. parallelise [PFS^{+}04]. Parallelism [HGS10, VRGÁ15, VvW98, vv91, ABMQO11, ABB^{+}16, ATL^{+}12, BNS^{+}07, BS05, BETR17, BDSV98, CCGG14, Cha99, CS19, CC99b, DCDV98, DW99, DK04, DM90, Gao90, GW95, GHP10, HK91, HC05, JS13, Kog85, LSL14, Lee97, Lil94, LL98, LF00, MCB16, MRBQO14, McC89, NR20, OKF16, PTS^{+}12, RRP03, ROZ01, RVD10, SSOB03, THH^{+}05, Xue94]. parallelism-based [VRGÁ15]. parallelism=simplicity=performance [Wel89]. Parallelizable [DWS^{+}23, SC95]. Parallelization [AL93b, Ano99b, ARR19, AFG^{+}97, AED21, BSE88, DE09, GLMBMPMV03, HB90, JTS^{+}11, KPS90, LT09, MRW^{+}08, MSM98, MSSC10, PL00, SPW^{+}15, WSR14, ARW93, AART13, AFT23, BK94, BBC^{+}11, BDSV98, BCG00, Bra88, BCC^{+}97a, BCC^{+}97b, CDRV97, Che08, DTV21, DL97, Dow90, DL05, DGP^{+}16, EMV^{+}18, FHK^{+}11, GHH^{+}03, GAPZ00, HM02, HOSS91, IDS16, IJM^{+}05, Kd88, KN09, KBGZ88, LK00, LPAZ97, ML00, MCB05, MB88, MSE93, NM01, NSS15, PCA01, PCL23, POHS14, PAG^{+}05, RE98, Rau98, RR14, SM13, SOH94, SWW99, SC05, SNO99, TCL92, WY11, YTCS97, ZBG88, vWWM^{+}19]. parallelizations [JW09, SSY10]. parallelize [EK98, MS91]. Parallelized [MM95, GJS93a, Koh95, PJ84, SL03, WWT01]. Parallelizing [DWX^{+}12, DEH^{+}11, GJS93b, Her00, Hwa02, KJA15, KLN90, RGdS^{+}13, Sch92, Sea86, WHW91, Zha95, BC97, DFRR91, KSM^{+}94, MJ95, WH93, YTCS97, dV94, Hol95]. Parameter [BCMG^{+}07, HC89]. Parameterization [ÖW10]. parameterized [BPK12]. Parameterizing [MS04a]. parameters [GE11, Hoc91, WL14]. Parametrization [Hoc87]. parareal [Aub11]. Parawell [Eck90]. paraxial [HSS07]. parcel [NC02]. ParConnect [AGV^{+}17, BCE^{+}17, HSSM17, HH17, TSLL17, WBN^{+}17, YLF^{+}17]. PARDISO [SG02a]. parenthesis [HHQ06]. parGeMSLR [XKL^{+}22]. parity [ARCH05, CLHL23, SSH97]. parity-check [ARCH05]. parity-declustered [SSH97]. PARMACS [CHHW94]. PARMESH [Glo95b]. ParPar [KEEF01]. parsing [BC04, CC96b, Hil92, SS87]. Part [ER95a, ER95b, BCC^{+}97a, BCC^{+}97b]. partial [AAHF97, ABB^{+}11, BCR00, BP02, CG87, GW87, HM89a, HK92b, KY98, MQ89, Meg90b, Pin91, RSSS94, RMS^{+}18b, SN88, SS88, XCS05]. partial-static [KY98]. Partially [CB09, CDGM96, IR00, OS21, Son94, TS88]. Particle [SSN04, BAS13, DBVS01, DOD23, GPLW17, Hat06, HHSM88, HF20, Koh95, LW11, LMC05, LF88, Luc01, MII^{+}11, SYAU07, Sch92, SAWH88, SG16, TB03, WVHR16, YSBM97, VK17]. particle-in-cell [Hat06, HHSM88, LF88, MII^{+}11]. particle-in-grid [WVHR16]. particle-laden [GPLW17]. particle-mesh [BAS13, YSBM97]. particle-particle [YSBM97]. particles [YR18]. particulate [ATL^{+}12]. Partition [FTY^{+}20, AFT23, BFY06, FP92, Mei85, Mv88, PMAL14, SL90b]. partitionable [PW22]. partitioned [DWS^{+}21, MS20, PEO16, NZHY11]. partitioner [ABGR21]. Partitioning [Che96, HHR87, KP96, Ran00, UÇA10, Wai88, AGV^{+}17, AFK01, AG00, Bak00, BFL^{+}01, BF06, BHK00, BBB19, CLA15, DPSW00, EB94, Fig06, Got89, HAK^{+}21, HK00, HB94, KGS08, Kd88, KŠR04a, KŠR04b, Kru97, Lan96b, LFS14, LWZL24, MS98, MMS09, MPD00, MOF04, PRV87, Pel97, PB16, PER17, QWD^{+}21a, QWD^{+}21b, SER90, SA16, Thu90, Thu91, Thu92, WC00, WW90, WAT20, Yal97, ZY16, Zho93]. partitions [LL98]. party [Kre13, WBPM97, WLN^{+}96]. ParVoro [WTLW23]. pass [Gir02, HMS^{+}07, KY94]. Passing [BC19a, FGG^{+}98, GLDS96, AB02, ARCH05, AE93, BCM94, BL94a, BRWL09, BDOS95, BWL00, CHHW94, CGMS94, CDZ^{+}98, CCS03, DM02, FK94, Gog11, GDAK06, HP05, HGM20, JOL^{+}21, JIC96, MRBQO14, McB94a, MRT93, NC02, Pie94, Pry97, SK97a, SSOB03, Ste90, TO91, Wal94a, Wal94b, WY11]. passing/shared [WY11]. passive [HKW13]. PaStiX [HRR02]. Patch [FHK^{+}11]. Patch-based [FHK^{+}11]. Path [CMP92, KJA05, ABR03, AÖ22, BG01, BR90, BGB10, CCS87, CG93, DQRR00, EMV^{+}18, GT90, HTB01, HC92b, LNS03, PB94, SB95, Trä95, Wan08, XCS05]. path-relinking [ABR03]. paths [BJV^{+}16, CST02, EHE92, GTW^{+}20, GS04, Kha12, KJA05, KLL^{+}09, SHCD00, TF01]. pathways [JTS^{+}11]. PaToH [UÇA10]. Pattern [Yan04, AKNS91, BCGS93, MAS06, TLS16, vWWM^{+}19]. patterns [BDS88, DRST03, FSKF06, FKK^{+}06, GÁVRRL18, LGMdRA^{+}19, MAB^{+}02, MFGEL17, PM95, SLH^{+}18]. PAX [HST87, HHSM88]. PC [BS00a, SKN04]. PC-based [BS00a]. PCG [PLP97]. PCs [DL04]. PDAC [HMTX93]. PDE [GZ99, HHR87, McB88a, Not95]. PDEs [GG10, Gro87, HHJP16, MSV15, NBB^{+}02]. PEAB [YHZ21]. PEACE [Sch88a, SP94, Sch88b]. Pearson [RvG05]. peeling [Wan09]. Peer [CdCM07, MMO07, AFJG06, MTV08]. Peer-to-peer [CdCM07, MMO07, AFJG06]. PEI [VP92]. Peking [FYH^{+}18]. Penn [Mic97]. perceptron [CH92, HT92, HT94, KK94, Müh90, SM90]. perceptrons [Dod89]. Percipient [NDW^{+}19]. percolation [CCEJ01]. Perfect [MP95]. perfectly [Xue97]. perform [MCM01]. Performance [AÖ22, AAV08, Ano16k, A^{+}99, ASH92, ADR^{+}05, BAR98, BZ99, BBD^{+}12, BCC^{+}97a, BCC^{+}97b, CC96a, CPdM^{+}97, CLA15, CCP98, CRSS99, DM97, DY91, Dun91, EGKU02, Ema10, FKK^{+}06, FHK^{+}15, FK89, FGHB94, GICM18, GEvGCP09, GLD08, GZTR07, GKV94, HVA^{+}16, HMP^{+}16, Her88, Hoc91, HL13, KN88, Kat01, KCG08, KK92, KM88, Kol94, LG21, Lin08, LMKH97, MP92, MS23, MZ01, MST94, MW88, NMAB19, NMW93, NA08, OW94, PFS^{+}04, Pan01, PGK^{+}00, PHCR05, PKR00, QCC02, Rib84, SNS^{+}97, SAOKZ06, SH91, Sha06, SE98, WME^{+}95, XLY^{+}20, Y^{+}99, YTSI13, YCBD19, APPG15, ASA16, Ahm97, Ala95, A^{+}02, ADLL03, AB93, ABMN02, ADE84, Axe86, BG91b, BAMK07, BA08, BKS19, BE89, BY21, BGL^{+}04, BWL00, BHS99, BFH23, BP90, BDH95, BOS^{+}09, BPJ22, BLCR21a, BLCR21b]. performance [BJV^{+}16, BBL^{+}19, BGRD00, CIS99, CMT04b, CHZ^{+}19, CK90, CU04, CAHT17, CLL99, CB17, CFD^{+}16, CGHBS18, CMSL06, CFN03, CZJS12, CDK^{+}03, CJLS14, Cla03, CC99a, CMM03, CQ97, CT88, CV02, CCP^{+}21, CCH^{+}21, DS03, DK01, DH91, DHS00, DSCP88, DFM^{+}95, DWL^{+}12, DDF^{+}10, Eck90, EPS98, EHF^{+}97, FAF16, FD18, cFM07, LHR^{+}19, FGP23, FDTZ04, GKS07, GMF00, Ger94, GS22, Gin99, GHP10, Gog11, GV08, GYL00, Gra91, GPSK09, GLDS96, GL97, GKKS01, HJ05, HH91, HLDS95, HMTX93, HSÇ12, HZY^{+}19, HB90, HRR02, HKW13, HBH^{+}16, HDH97, HW91, HHOMR06, HWB92, Hoc87, HAJK01, HMS^{+}19, HHR87, HEB96, HR04, HG12, HMH^{+}13, HC05, IJCL96, JG93, JHSV02, JAA06, JMLBL98, JKHK08, JJM^{+}11, KMLM97, KRW88, KT00, KBG^{+}01, KV99, KG03, KSS98, KJ07, LR04, LKGD16, LG01]. performance [LFS^{+}19, LVC16, LLW^{+}15, LLD19, LXC^{+}18, LYC^{+}19, LBGO23, LR89, LF88, MVdLvN16, MD04, MKL^{+}01, MSM98, MSMC15a, MAH^{+}19, MFS^{+}19, MSE19, McB88b, McB94c, MQ89, MSK14, MAM^{+}09, MSS^{+}05, MS99a, MJ99, MK97, MAB^{+}21, MGT^{+}13, MDC^{+}08, NS94, NS02, ND17, NU05, OKF16, OIH10, OGC^{+}15, OILZ17, PRV87, PSS94, PFTB15, PGK^{+}18, PMV^{+}20, Phi91, QH19, RBB^{+}22, RMS^{+}18a, RB14, RDS13, R^{+}00, RGDM15, RGBCS09, RH12, SLH^{+}13a, SLH^{+}18, SHCS86, She93, SCM^{+}98, SMJ15, SV97, SB97, SVTSM^{+}22, Sor84, SHCR98, SDMS99, SDMS05, SNO99, SR10, SG91, Tak06, TKC^{+}14, TYKA95, TG09, TRSHB04, TAB^{+}19, TLC23, VJ12, WH97, WN91, WBS06, WJK98, WSL88, Wat87, Web90, WYBdJ^{+}21, WS98, WvR16, WOKH96, Wor91, WHL19, XT11, XH96, YS96, YDTS01, ZLM98, Zen99, ZGZS20, dNdRRL^{+}21, de 96]. performance [van86, MVZ98]. performance-aware [KBG^{+}01, MSMC15a]. performance-centric [CFD^{+}16]. performance-portable [DWL^{+}12, FAF16, MAB^{+}21]. performances [DL04]. Performing [CN98]. perimeter [BW01]. perimeter-time [BW01]. Periodic [TM09, TYLL22]. permutation [LJD93, She95, Wit98]. permutation-based [She95]. permutations [LT94, Lin90, LY93, LC99b, TL94]. persistent [BE97, HMS^{+}19, IYV^{+}22]. perspective [BJK^{+}17, PODd16, SL03, W^{+}99]. Perspectives [Her00, McB94c]. pessimistic [Don05]. PET [Che98a, GH98]. peta [LK10]. peta-scale [LK10]. Petascale [CYXL18, MSV15, SML^{+}14, BOV09, SWR^{+}13, ZYZ^{+}16]. Petri [WN91, CFC99, MS89]. PETSc [CDM06, HM01, MAB^{+}21]. PFASST [SR17]. PGAS [PEO16, TBM^{+}16b]. phase [ADGS10, DE09, LS92, WAT20, ZL04]. phases [DZD01]. PhD [Hol11]. phenomena [KRS^{+}21, PIG^{+}16]. phenomenon [YR18]. Phi [HSSM17, Tak18, TLS16]. philosophy [UWC14]. photon [BLFT84]. photosynthetic [SLS^{+}21, SSV^{+}16]. Physical [SJJ85, PGGP19, RBB^{+}22]. Physically [TB07]. Physics [BKL^{+}19, WHL19, BKS19, BBL^{+}19, Gup99, HZY^{+}19, LyHW^{+}22, LYC^{+}19, Pea19, Rap99]. PI [PUSS97]. PI/OT [PUSS97]. PIC [DBVS01, Hat06]. pickup [CGGG03]. pilot [GCP11]. pilot-job [GCP11]. Pin [LC89]. Pipeline [CCJ90, HK06, SWSG92, CT88, FHKZ88, GAMR03, MR86, MCG^{+}12, NAC^{+}14, RVD10, Tap84a, UN87, Wil85]. pipeline-interval-optimum [UN87]. Pipelined [BD89b, BE89, BAZ93, CC92, CV17, Coo19, Gao87, HHH92, LPH00, MRT93, RCAP11, Sor84]. pipelines [MG19, Tap84b]. pipelining [CWK09]. pivoting [Ala89, AC94a, CG87, Meg90b, Pet84, RDB^{+}90]. pixel [JON08]. placement [Has01, MSA09]. planar [HI92, Kim90, KSS06, SE87]. plane [RBMO20, SM88]. Planning [HMS^{+}19, MLM^{+}00, SB95]. Plasma [TTT^{+}92, CDH^{+}03, DBVS01, MWI^{+}15, RKMJ10]. platform [AGB97, CC97b, CTA22, DTV21, DWL^{+}12, GSC^{+}22, KR10, NMMG13, SBP12, ZLM98]. Platforms [LCE^{+}18, AAS20, BLRR02, BLR03, BEDdC16, BEK^{+}11, BDNP11, BDRV99, EYP^{+}20, FLYL21, LHR^{+}19, MII^{+}11, MAA^{+}21, SWCBQ19, WOV^{+}09, ZGL^{+}19, dSS09]. Plugging [CDRV97]. Plus [AV15]. PMAA [AGPS03]. PMAA’04 [CGNR06]. PMAA’10 [ASSS11]. PMAA’14 [AGKS15, AGKS16]. PMAA’16 [AAGS18]. PMAA’18 [SAGV21]. PMAM [SHB19]. PMAM’17 [ABH18]. PMC [MHE19]. PMC-space [MHE19]. PMIRKDC [MPJ03]. PMIx [CHBS18]. PML [Ram07]. PnetCDF [BPC21a, BPC21b]. pNJTree [DL06]. point [ADLL03, CGOP03, CB09, DM03, DNL15, DR03, GGA19, GICM18, Kam88, Kie91, LTG23, LHS13, LBGO23, NYFK06, PJ84, POHS14, ZG16]. point-to-point [ADLL03, CGOP03, CB09, GGA19, LBGO23]. pointer [Her00]. pointer-based [Her00]. pointerchain [GCC19]. pointers [AZ98, GCC19]. points [KH95, Wei93]. Poisson [BT88, LV92, SBO^{+}91, Vaj84]. polar [HP94]. Polder [vOS^{+}98]. policies [APPG15, LGC97, PGAA16, Zam99]. Policy [ZWJ^{+}19, MMM13, RBL97]. pollutants [Zla88]. pollution [BWZ95, OZ02]. polyacenes [NHS^{+}95]. polycrystalline [LVC16]. Polydisperse [ER18]. polygon [AU88]. polygonal [AU88]. polygonization [HS03]. polygons [Kim97, LSS05]. Polyhedral [SLC^{+}16]. polymer [Bis84, OCSBY01]. polymerization [EGTD99]. Polynomial [Jan06, AO89, But92, CF90, Fre89, FB91, KP93, LT09, LWS02, PW89, SMK91]. Polynomials [Sen91, EK92, LAS90, OW97, SK91]. pool [YHZ21]. pool-based [YHZ21]. popular [Wei91]. populated [MHE19]. population [LP92, MSOCG^{+}16]. poroelasticity [TA14]. port [ST02]. portability [CHLO85, CTA22, CCP^{+}21, DM97, LG21, WYBdJ^{+}21, tV96]. Portable [CHHW94, JHSV02, SEAH^{+}08, YX07, ABMN02, BRH90, BAS13, DAA94, DFRZ02, DS87, DWL^{+}12, FAF16, GLDS96, KA96, MAB^{+}21, Van02, WS12]. portfolio [CLHL23, GS10, LMH08]. Porting [BBB^{+}22, FLS^{+}21, MAA^{+}21, PMMAM10, WLN^{+}96]. ports [CC00c]. Position [GV11]. Positive [EV89, HRR02, NYFK06, ZVWS88]. positron [Möl99, ZLM98]. possible [FLW86, GK03]. possibly [CLA14]. Potential [LCE^{+}18, CLC^{+}18, DSSD18, FYH^{+}18, LLX^{+}18, LHL18, MBB^{+}18, PTK04, SWR^{+}18]. potentials [HSS09]. Power [ACM^{+}10, ACM^{+}15, BLCR21a, BLCR21b, CFD^{+}16, CHQORS18a, DZM^{+}13, ECLV12, EL91, KAM^{+}20, LKGD16, MHE19, Mak94b, NS11, NFG^{+}13, OW94, ŚDJ^{+}22, TKC^{+}14, TDG^{+}18, TYLL22, WZV^{+}16, WC15, YA24, YZB^{+}23, ZLPF16, Bri95]. power- [NFG^{+}13]. power-constrained [KAM^{+}20]. power-efficient [YZB^{+}23]. powering [CKM94]. powers [EMB89, EM90, SB92b]. pp [WBD99a]. PPS [MCY^{+}24]. Practical [BWZ95, CW97, Hol95, JKH95, LBL95, kLH95, STKA96, Trä18, WRB97, KT99, KA96, KS99, Par87a]. Practice [TFV16, CAA98, NOG^{+}22, ZGZS20]. pragmatic [Col04]. PRAM [CC96b, HK06, KT99, She97]. pre [BOVG10, KEEF01, OV06, Wak04]. pre-load [KEEF01]. pre-processing [BOVG10, OV06]. pre-propagation [Wak04]. precedence [BHR09, MKS91, QJ06]. precision [BEK^{+}11, JAH^{+}18, LPNV20, Tak10, WRS12, ZGG92]. precompiler [GF89]. Preconditioned [AGD^{+}17, BRS97, ABB^{+}16, ABG^{+}06, BM90, CG89a, ETV91, GYH94, GZHX17, GV14, TGE92, VvW98, van86]. preconditioner [BP02, Daḡ07, EG94a, GPSK09, HJ97c, HKR^{+}10, PLP97, PHD16, RMS^{+}18b, SY21]. preconditioners [AO89, CGMM99, DNL15, EV89, Ema10, GHW08, Haa98, LWS02, SSSG03, Str08]. Preconditioning [AHBD18, ADFQO19, BPS01, GCH21, Lip99a, SZ02, Wai88, XKL^{+}22, YD97]. predict [CK90, FDTZ04]. predictable [MFS^{+}19]. Predicting [BGL^{+}04, RvG05, WC15]. Prediction [CB17, AFD13, ADR^{+}05, ADE84, BG07, CQ97, GJS93a, HJ97b, ISO^{+}03, LZCT15, MJ99, SXBD97, WH97, XDZL10, XH96]. prediction-based [LZCT15]. predictions [AÖ22]. Predictive [FTY^{+}20, BLCR21a, BLCR21b]. predictor [RM97]. predictor-corrector [RM97]. preemptable [BDF^{+}00]. preemption [GF03]. Preface [Ano13i, Ano16k, CC16, Pet92, Thu10]. prefetcher [CZJS12]. prefetching [PSS01, RG09, YEC97]. prefiltering [BNK15]. prefix [ABH^{+}10, ZM96]. Preliminary [Ger94, WLK^{+}18]. PREM [MFS^{+}19]. preprocessed [SPW^{+}15]. preprototype [Pei88]. prescribed [TKG97]. presence [CHR18, MW91, SAOKM01, TYSF13]. preserving [MDC^{+}08, SS95]. price [Ger04]. pricing [FMM10, SÖB07, Sur10]. primal [NYFK06]. primal-dual [NYFK06]. prime [Mas98, Tem88]. primitives [STP^{+}19]. principles [Her88]. prior [NMI^{+}24]. priorities [SS04]. Priority [Bro99, Bur04, DP00, WH94]. prism [HWL^{+}22]. Probabilistic [BS97, JDG02, YMY^{+}19, ZS21]. probabilities [EB93]. probability [ASDOK01, BWT^{+}08]. probe [KG21]. probe-based [KG21]. Problem [CMP92, MW91, AC89, Als01, ACM20, Arb92, ADTV01, ACGL04, ABH^{+}10, AHW14, BW01, BR90, BG97, Bev89, BDK96, Bol86, BGM03, BCG00, CGGG03, CCS87, CCS94, CJ92, CDK^{+}03, CA92, CW97, CWB92, Cor00, CGOP03, CFH89, DCY^{+}22, DG90, DG95b, DQRR00, DvdG92, EM85, Eva91, GLMBMPMV03, GPS03, Gol88, GT90, GM03, GHC04, Gur88, Gut88, HI92, HC92b, JK91, Kal90, Kal92, KŠR04a, KŠR04b, LLW^{+}15, LS97, Lin91a, LO92, Lin94, LV92, LC97, ME90b, MMT06, MP95, MOF04, MGSK87, Nag88, NA03, Nat90, OWZ91, Par91, PMV^{+}20, Reg01, SSY02, SK19, SMK20, SF91, SL90a, SL90b, SE87, SM88, SS91b, Tai91, TNZM20, TTO^{+}22, TSCS14, Thu92, TRSHB04, WBPM97, Wen95, Zhu93, Zub90, Bur04, DN16, DUG^{+}06, LW15, MNS87]. Problem-heap [MNS87]. problem-size [Lin91a]. Problems [DZ90, AK89b, ABB^{+}11, AIV95, BP95, BCC^{+}97b, BGB10, BR84, CSEK03, CGG01, CW06, DM03, DFO93, El 93, EK98, FYEHP89, GSZ88, HJ97a, HF20, HK91, KH95, KR10, KH92, Kie91, Kie94, Kim98, KR00, LLPV06, LPM11, LPH00, LPMD01, LQ92, LNA06, MV99, MWH95, MST94, MR86, MS90, NDN20, NP05, OIK21, OHZ98, PG93, PMV^{+}20, Pir93, Pir96, RLH19, RS88, RV95, SMK91, Sch91b, SSD01, SC95, Sug99, Sun97a, SW03, ŚDJ^{+}22, Sýk84, TN13, TO99, TCL92, VV07, Wri90, ZPAT99, ZM88a, ZM94, ZLD15, ZG16]. procedural [MSE95]. procedure [CF91, IS18, LS92, Mei86, Riz85, YNY^{+}17]. procedures [SS94b]. Proceedings [HKK97b, Jou97, KVW97a]. Process [CHBS18, Jia09, BO91, BGL01, DCN^{+}15, GJMMS97, HG12, PWM00, Qua92, SK97b, SWR^{+}13, WH91, YXQ^{+}21]. processes [ATT89, DS07, EGTD99, HSR^{+}14, MZ04, Ram99, SLS^{+}21, SlWC89, WFW^{+}11, Wil85, vS92]. Processing [B^{+}02, Adr99, AHL02, ACMT20, Alm85, ASH92, BM01a, BOVG10, BCB02, BCRSR11, B^{+}01, BWT^{+}08, Bro87, Bro88a, CLS93, CS89a, CC00c, Chr98, CMM^{+}88b, DP02, DCDV98, DOD23, EM88, Fah85, FMS01, HL97, HLvHA11, HSS09, HCJ03, Hir86, JON08, KN88, Kan93, KY98, KK01, KA19, Kot97, Kut85, kLH95, LS90, LXD^{+}23, LWJ^{+}17, MP08a, MP08b, MHH97, MSW91, dTNOR^{+}04, NJ02, Ohb85, OV06, Ori10, SMT^{+}04, SKG02, SSS99, SG02b, SS01, SC92, SY87, SH88, Sur10, TMCC02, Tor91, UZ02, WLK97, WH22, WB97, YD97, Zom95, dCC89, Has90]. processing-in-memory [LXD^{+}23]. processor [AFJG06, BAMK07, BFY06, BDSd94, Bol86, Bro87, BAZ93, Bus93, Cal85, CC91, DHL07, DJ87, DS92, EEH^{+}19, EB88, EG92b, FFLV96, FBMV88, FYEHP89, GIF^{+}10, GB14, GE92, Hil92, IM88, JWS13, Kas84, Kat01, KSP97, KAD09, LXC^{+}18, MM96a, Mak94b, MT91, MR86, NNB^{+}99, PRW88, Qua92, RV96, SV94, SS94a, SB95, Str87, TTH95, VvBv90, Vég18, WYX^{+}22, Wil88, WS94, WRS12, YD01, YLW^{+}13, ZJD^{+}23, ZVWS88, de 87]. processor-memory [Bro87]. Processors [Ano88b, AAV08, ACH^{+}18, ABMT99, Ann89, ADFQO19, AKNS91, BDF^{+}00, BV04, BGG^{+}99, Bis89, CAS09, CHQORS18b, CZJS12, CC92, CC99b, DC15, DHS89, DG99, FK89, FMR90, Gro87, HK90b, Li97, LL90, LLL10, LV15, MV87a, McB88a, MIA^{+}07, Meg90a, MAM^{+}09, MSW98, MOW95, NMAB19, NB12, NRN00, R^{+}00, RHM^{+}88, SA13, Son92, Tak18, VJ11, WXL^{+}18, YNO97, YL86, Zla88]. Procrustes [Par91]. Producing [HAJK01]. product [AAHF97, Bar97, CMT96, EM88, FG89, HCR01, KD10, VFG12, YORO08]. production [AAB^{+}16, BDI^{+}95, GCP11, LGM09, NOG^{+}22, Taf01, TRLD13]. productive [SLH^{+}13a]. productively [SDMS12]. productivity [IMQO^{+}18]. products [KKSS90, KS91b, NMAB19, Obe85]. profile [MSK14, RVD10]. profile-based [RVD10]. profiler [MAD^{+}16, PB16]. profiles [MAJD17]. Profiling [KC97, EYP^{+}20, MK12, NMI^{+}24, WZV^{+}16, WZCG14]. Program [GPC88, LKGD16, LWZL24, Psa02, SS91a, CZTS99, CDZ^{+}98, DGNP88, DL06, GJS93b, KMLM97, ROZ01, RH12, SNS^{+}97, SCD92, WvR16, YS96]. programmability [CHZ^{+}19]. programmable [LS90]. Programming [ANG^{+}20, ABH18, BVC16, BVC19, DSCP88, Jor88, MS23, WLN^{+}96, YFK03, AQO18, ACH98, ARR19, BE95, BH14, Blo03, Bog92, BRH90, BÁJG^{+}20, BL94b, CLS93, CHHW94, CDM06, CCS03, CHLS90, CCJ90, CBV13, CFW^{+}22, CMV^{+}06, Col04, CV02, CGM05, DM03, DCG90, DWL^{+}12, ECG93, FF21, FSG19, GSE^{+}15, GGNY11, Ger98, Ghe85, Gil94b, GEvGCP09, Hem94, Her00, HMS^{+}98, HST05, Hoe12, IUY^{+}04, JMA^{+}13, KT99, KA96, Laf02, Lef97, LEH14, MAFC14, MMTU18, MSP93, MIA^{+}07, MV87b, M^{+}00, MKL^{+}88, Nak05, NZ97, NZHY11, NRN00, ONB11, OCPT97, PS84, RRA11, RR14, RR16, STP^{+}19, SBP12, SW91, Seu85, SK97b, SHB19, SVC19, Son86, Sto89, TBM^{+}16b, UAR^{+}99, Van02, VP92, VKH99, Wal01, WYT^{+}16, Yue97, ZLT19, de 87]. programs [ATA^{+}16, AZ98, A^{+}99, AKP^{+}98, Arv92, Bek95a, BGMT00, BGL^{+}04, BHS99, BDS88, BS96, BS00b, BGL01, CCL04, CSH^{+}11, CI98, CCH^{+}21, DR93, DS87, DR03, DGKF19, EGTD99, EO91, Fer97, FGHB94, FC18, FBG^{+}12, GTW^{+}20, HvNJB12, Hwa02, Hwa04, JPK^{+}15, KCD^{+}97, KY94, Lee95, LF98, LMH08, MAB^{+}02, Mah96, MS98, MS99a, MBK12, MHH97, NL91, NYFK06, OM90, PAG^{+}05, PS91, Psa02, RvG05, RVD10, RW89, SER90, SM91, SS94a, SV97, Str87, SLH^{+}13b, THH^{+}05, WW90, Wol89, YMY^{+}19, ZZ03, ZJDW18]. programs/data [PS91]. progress [MCM01, RSC^{+}19]. progressive [RR97]. project [C^{+}01a, Hey94, KNP97, SG02b, Tro94, UAR^{+}99, WPD01, YK87, AAB^{+}06]. projection [ADF93, ALH^{+}14, GK04, KS89, KW90, Slo91]. projector [KHHC20]. projects [McB94b]. Prolog [KNP97, MBC97, VK92]. promise [Hac89]. prone [LHR^{+}19]. proof [GRP22, HE88, KS87, KB20]. propagation [AATK10, CFD^{+}16, DDF^{+}10, FGS97, MMC97, SCCPM14, SGP22, TPK^{+}13, Wak04]. propagator [LSP92]. properties [BM01a, BA95, EBSS94, EAKT90, ME92a, TCT00, YLT^{+}23]. Proposal [MSS19]. proposals [RS88]. propositional [BSK03, CF91]. protecting [FLYL16]. protein [BM08, GÁVRRL18, KHN01, OCSBY01, Rob00]. protein-interaction [BM08]. proteins [FM15, VBS^{+}15]. Proteus [JAH^{+}18]. protocol [BDK06, CWK09, LLFZ23, RGT17, YLW^{+}13, Yua94]. protocols [ACLN03]. provenance [PGGP19]. proving [CF91]. provisioning [GCP11]. Proximity [Bur04, BP97]. Pruning [LDK16]. PSelInv [JLY18]. Pseudo [FHJ^{+}84, FHL87, mMvdV01]. pseudo-overlapped [mMvdV01]. Pseudo-random [FHJ^{+}84, FHL87]. pseudorandom [MM95]. pseudospectra [BG02, MP02]. pseudospectral [Bri95, MRRP11]. pseudospectrum [BG01]. PSFGA [dTNOR^{+}04]. psi [Tak06]. psi-cube [Tak06]. PSO [YA24]. PT [CP08]. PT-Scotch [CP08]. PUB [BJvOR03]. Publisher [Ano19m]. Purcell [CL02]. Purdue [BKL^{+}19]. pure [BHS99]. purpose [AHBLR12, CT94, DZG94, Hac89, PJW^{+}22]. push [LAHM14]. push-relabel [LAHM14]. PVM [Ahm97, ADMV05, CDGM96, DG95a, JH97, LHZ97, MSP93, SGS95, SSS99, SGDM94]. PyCUDA [KPL^{+}12]. PyOpenCL [KPL^{+}12]. pyramidical [EMB89].

Q [BJV^{+}16, CDZ^{+}21, MV17, WZV^{+}16]. Q-learning [CDZ^{+}21]. QCD [A^{+}99, CIS99, GLS99, KLN90, Lep99, Lip99a, Uka99]. QCDSP [Maw99]. QIF [EB88]. QL [RH99]. QMPI [EYP^{+}20]. QoS [JMA^{+}13, NMMG13, SYY^{+}22]. QoS-aware [SYY^{+}22]. QR [BOVG10, BM01b, DFH^{+}13, DM02, GK03, HK06, Kon02]. QSDF [Got89]. QTMS [YXQ^{+}21]. quadrant [GMBZ90]. quadratic [DM03, DR03, Tai91, YXQ^{+}21, ZZ03]. quadrature [BE88]. quadrillionth [Tak18]. quadtree [PGBF^{+}07, WA03]. Quaff [FSCL06]. Quality [MD04, MPD00, DM97, OIK21]. quantification [MHK97]. quantify [TDG^{+}18]. quantization [AB02]. Quantized [BV04]. Quantum [MAM^{+}09, ABH^{+}10, BS00a, BJK^{+}17, BH99, GPW^{+}08, Gup99, JPK^{+}15, KR00, LXD^{+}23, NMW93, PJV00, SSGF00]. quantum-mechanical [KR00]. Quark [LSP92]. Quasi [LWG06, TLC^{+}21, Beb97, CH98, KS16, LM00, ÖW10]. quasi- [KS16]. Quasi-atomic [LWG06]. quasi-Monte [LM00, ÖW10]. Quasi-Newton [TLC^{+}21, CH98]. quasi-stationary [Beb97]. queen [BW01]. Quenched [Rap99]. queries [CRGR^{+}13, HM02, LND^{+}19, Reu99]. query [BV96, RBB^{+}22]. question [BC97]. queue [CAS09, SP97, WL98]. queueing [HMTX93]. queues [Bro99, DP00, RG09]. queuing [MPZS13, TB22]. quickest [HC92b]. quicksort [Lev90, FP92].

R [FD02]. race [CSH^{+}11, CRD02]. rack [MP87, TM90]. Racoon [DG05]. radar [FFLV96]. radio [AATK10, WYX^{+}22]. radio-wave [AATK10]. radiosity [APRP97, RR97]. radius [Rob00, Wor92]. radix [BAZ93, TTH95]. Radon [FF92]. RAIL [FMB98]. RAMA [MK97]. ran [HSSM17]. random [APG92, DP90a, Deá90, FHL87, FHJ^{+}84, Mak94a, San99, SR97, SMC03, TS21, dP90b]. randomization [FMM10]. randomized [BBB^{+}14, BGV97, CA92, ÖW10]. range [CRGR^{+}13, GWC^{+}99, GVH08, Kau94, Kim98, She95, Zla88, dP90b]. range-join [She95]. Rank [Meg90a, CKM94, DT95, KTN^{+}14, Kon02, XKL^{+}22]. rank- [DT95, Kon02]. rank-mapping [KTN^{+}14]. rapid [AAB^{+}16, FHMS93]. rarefied [MT97]. Rate [CS00, CLY^{+}19]. rates [LD98]. rational [MKC92, SSL19]. ray [AATK10, DGBP05, Lee97, RJ97, SGS95, SSV^{+}16, TS21, YEC97]. raycasting [MSE07]. rays [TS21]. RC dag [PWY03]. RDMA [GRBA15]. re [GGL^{+}05, KSM^{+}94, ZSI02]. re-configurable [GGL^{+}05, ZSI02]. re-engineering [KSM^{+}94]. Reaching [CC93, HC05]. reacting [SHH^{+}97]. reaction [HSR^{+}14, Wei97]. reactive [AFPG12, Ram99, WH15]. reactor [HSFS14]. read [CS19]. Real [AGB97, GMF00, Heg96, HCH^{+}96, BW92, BEDdC16, BDH95, BG98, CDV08, GLS01, HA14, Ken90, MF16, NA03, QJ06, RHM^{+}88, SWCBQ19, SMTT96, TYLL22, WH91, YZB^{+}23, vWWM^{+}19]. Real-time [AGB97, GMF00, BG98, CDV08, GLS01, MF16, NA03, QJ06, RHM^{+}88, SWCBQ19, TYLL22, WH91, YZB^{+}23, vWWM^{+}19]. real-to-integer [HA14]. realisations [WVHR16]. realism [San02]. Realistic [HDH97, SHRN98]. reality [BGWW97]. realization [KSS06]. Realizing [McC89]. rearrangement [HRT07]. reasoning [BPEL05, QH19]. Receiving [ISO^{+}03]. recognition [CC96b, JC94, Ume01]. recombinative [MG95]. Reconciling [San02]. Reconfigurable [Cho91, DP02, BCYB11, DT95, Kas85, KCG08, LPH00, MO89, MRT93, MB00, OSZ93, PW22, SEAH^{+}08, VTmL03, Wan08, KB08]. Reconfiguration [WWL^{+}22, JWS13, MSMC15a]. Reconfiguring [LH02, RGBCS09]. reconstruction [BMCA98, BCMG^{+}07, BBB19, Che98a, DL06, GH98, KBBC88, LPCA98, Möl99, OIH10, RW98, SKH^{+}12, WCKM11]. records [BBDN21]. recoverable [PLY02]. recovery [CQ01, FCZY17, LH88, LWG06, MSK14, SRS^{+}19]. rectangle [JCL92]. rectangular [GDAK06, TTH05]. Rectangulating [Kim97]. rectilinear [Kim97]. recurrence [DQRR00, RF90, Stp93]. recurrences [Gao87]. Recursion [DGP^{+}16]. recursions [AE86]. Recursive [KCN99, SE93, Bus93, CHH^{+}01, CS89c, DLS09, DWS^{+}23, EZ99, ME93b, NM01, RBB17, SHCD00, dV94, Mar14]. red [BP02, SY21]. red-black [BP02, SY21]. redistribution [GNY00, HBH^{+}16, MQ97, Van90, WW95, BLRR02]. reduce [ATA^{+}16, CCGG14, PGW16, TN13]. Reduced [SKS^{+}95, ZM96]. Reducing [CRGM16, NMARD10, SA16, TB22, WH15, DB03]. Reduction [DvdG92, Ala89, AM93, BE93b, BDK95, CDGI15, DCD97, EY94, GS89, GTD18, GL99, HM97, KK11, Kuz98, Lan96a, LC89, ML20, ML23, OTI^{+}89, PGGL17, RM91, SST09, Sch91b, TDG^{+}18, TRSC^{+}19, TND10, YE94, ZM96, CTZ^{+}18]. reductions [ACH^{+}18, GPZ08, HT00, ZCBD22]. redundancy [ZLJ93, ZRS13]. Redundant [FMB98, Gir02, MKS91]. reference [Bev89, PL00]. REFINE [BA95]. refined [PSW18]. refinement [BN01, DLR94, DR95, HAK^{+}21, LPNV20, PGW16, VP92]. refining [Zha95]. Region [TPK^{+}14, AU88, DGKF19]. region-Aware [DGKF19]. Regional [BMS97, Adr99, SXBD97, YNY^{+}17]. regions [Box92, KH97, LFL11, SLY90]. register [MM95, TM09, Ume89]. register-number-minimum [Ume89]. registration [Chr98, IOH05, WJK98]. regression [GK03, YK08]. regular [ACM^{+}10, LLS^{+}93, Lin91b, SS01, Tro00]. regularization [BE93a, MDC^{+}08]. regularized [Zhu93]. regulatory [CFS01]. reinforcement [Ken90]. rejection [CLY^{+}19]. relabel [LAHM14]. related [Arv92, DW99, KH95, SMK91]. Relation [BI07, KLS^{+}88]. Relation-based [BI07]. relational [BF95, LTS^{+}24, TDW^{+}04]. relations [BK97, MKS91]. Relative [BE89]. relaxation [Bai99, BG97, CZ91, El 93, LG09, PJ84, RBMO20, SY21]. Relaxed [MX07, ZDE95]. relevant [HSR^{+}14]. reliability [CYDJ21, HWYL89, TYSF13]. reliability-guided [CYDJ21]. Reliable [ZRS13, DHS00]. relinking [ABR03]. relocation [GLS01]. remaindering [KC93a]. remapping [CNK93]. remarks [Tro94]. remote [AGO97, CDK^{+}03]. removal [CDRV97, DVP90, TFS91]. Removing [ZJDW18]. renderer [Kat03]. Rendering [RJ97, CM95, CFN03, CKS03, Cro97, DCG^{+}07, FB19, GWLS05, ISH03, KPH97, SL03, SIH14, Wit98]. Reordering [Liu89, GLU03, LC99a, PHCR05]. reorthogonalization [BLT^{+}22, KC93b]. replay [CSH^{+}11]. replicated [Kan97]. replicated-data [Kan97]. replication [ČSR05, CBP^{+}07, ZJDW18]. report [DLO17, MCM01, TSLL17, YLF^{+}17]. reporting [CF93]. represent [Wil85]. representation [Ken90, MT89, TPK^{+}14]. Reprint [FJPA17]. Reproducibility [DLO17, GHH18, HG17, NGM19, AGV^{+}17, CDGI15, FHL87, TSLL17, WBN^{+}17, XRW^{+}19, YLF^{+}17]. reproducible [LHG^{+}23]. Reproducing [BCE^{+}17, BKS19, BKL^{+}19, BBL^{+}19, CLC^{+}18, DSSD18, FYH^{+}18, HZY^{+}19, LCE^{+}18, LLX^{+}18, LYC^{+}19, LHL18, MBB^{+}18, SWR^{+}18, WHL19]. request [WLYJ13]. request-request [WLYJ13]. requests [CLY^{+}19]. Requirements [GT07, BMYK98, BS00b, GPS03, SG16]. rerouting [CF04]. resampling [DOD23, KHC92, LR16]. Research [Ano87, Ohb85, SHCR98]. reservation [TKK^{+}05, ZP16]. reservoir [BBK90, Mic90]. resilience [AQO18, CHQORS18a, RMS^{+}18b]. Resilient [FGBN19, TDG^{+}18]. resizable [SR10]. resolution [Evr01, HW94, VK92]. resolved [GPLW17]. resolving [Ara03]. resonant [Ste88]. Resource [BCRSR11, CGH^{+}19, GRP22, MSA09, BK07, CRGR^{+}13, CLY^{+}19, CL05, FL21, GCP11, HMH^{+}13, KJA15, LCLA19, MKK^{+}19, SVC07, SS94a, STE23, SSS^{+}09, SYY^{+}22, TKK^{+}05, YBM05, ZP16, ZRS13, vOS^{+}98]. resource-constrained [KJA15]. Resources [BFVRC14, ASH92, BGK^{+}98, LSC^{+}15, MMO07, PVBR23, RMRN05, RSC^{+}19, WCM^{+}14]. respect [HWW92, LGM09]. response [TDW^{+}04]. responsibilities [SH91]. responsive [GE11]. Restart [SMP11, AKB^{+}19, NSM23]. restarted [SWC99]. restoration [BZ96, BP00, IT08]. Restricted [DR15, Mal02]. Restructuring [GP86, MJ95]. Results [HH17, Car88a, DFRR91, GJS93a, GLN89, Ger94, HISS92, HBH^{+}16, MQ89, MRJ89, Oed92, RS87, Uka99, VIHD90, WBN^{+}17, WLN^{+}96]. Rete [AT98]. Rete-Match [AT98]. Rethinking [TS21]. retrieval [AFT23, BGP^{+}97, Kao08, PW87]. reusability [ES06]. reusable [CMT04a, NE01]. reversal [SAOKM01, SB94b]. Review [LPNJR96, AK89a, Kog85, ŚDJ^{+}22]. Reviewer [Ano18l, Ano23h]. Reviewers [Ano10a, Ano08, Ano09a]. revisited [GR02, Uml94]. Revisiting [CIO^{+}17, PCLM18, RS92, dNdRRL^{+}21].

Revolutionary [BOV09]. reweighted [ZG16]. rewriting [BVP^{+}89, CMPM^{+}15, HL99]. Reynolds [ADTV01]. Reynolds-Hertz [ADTV01]. RI [Ber00]. RI-MP2 [Ber00]. Riccati [BBQOQO00]. rich [DCR^{+}16]. ride [ACGL04]. right [CLA14, KMB^{+}18, Pan01, Ter10, WP19]. right-hand [CLA14, KMB^{+}18, Ter10, WP19]. Rigid [ER18, LHR^{+}19]. Rijkswaterstaat [VRL03]. Ring [LD99, AIO95, CS97a, FMR90, KJ07, Meg90a, KJ07]. rings [AFY^{+}16, BGG^{+}99, CMT96, KL00, Pop88]. RISC [AL93b, NMW93]. rise [XC02]. risk [CLHL23, CDM^{+}10]. risk-neutral [CDM^{+}10]. river [WFW^{+}11]. RMA [FCS^{+}19]. RMESH [Kim98]. RNA [CLSM98, TLS^{+}08, XDZL10]. RNAVLab [TLS^{+}08]. robot [DHS89, Zom95]. Robust [RB14, SMK20, SSS^{+}09, Tai91, GG10, HSFS14, SN88]. robustness [BGMT00]. roles [Sat04]. rollback [MSK14]. rollback-recovery [MSK14]. rollout [GM03]. Romberg [EM86]. room [ZLPF16]. root [Jan06, MBC92, SMJ15]. root-cause [SMJ15]. roots [CF90, GCC19, MR85]. rotating [BE87, CEM^{+}99, Lun98, SSL19]. rotation [Jou89]. rotationally [CT88]. rotations [VKS^{+}15]. rototyping [FD02]. round [BD10]. rounding [Obe85]. routed [CWC00, FK98, HM97, MM96b, Miš98, RGBC11, SW00]. router [FT93, LNLK13, LNLK13]. routine [GKV94]. routines [CHQORS18a, DT95, DH84, DGP^{+}16]. Routing [PJW^{+}22, AAS13, AQ93, ASDOK01, AFY^{+}16, BL99, CC02a, CCP98, GP98, HCK94, HC94, KH97, KJA05, LY93, LC99b, Loh96, LH04b, LOKM99, NLP^{+}15, PGGL17, QCC02, Ral03, Reg01, RE92, ST02, SVC07, Sha06, SF91, Shi01, SW00, Tor91, XCS05, YTCS95, ZL04]. row [AC94a, BVWH14, GYH94, LB23]. row/column [AC94a]. rowwise [PJ84]. RS0 [MX07]. rubber [BM01a]. rule [CFMM97, Li89, TE92, VA14]. rule-based [Li89, TE92]. Run [DLR94, DR95, MPZS13, Rau98, GE11, HT00, KPL^{+}12, LK00, RRP03, SM13]. Run-Time [MPZS13, DLR94, DR95, Rau98, GE11, HT00, KPL^{+}12, LK00]. Runge [ES89b, MPJ03]. Running [BWZ95, dSS09]. runs [Dod90]. Runtime [BNS^{+}07, AKB^{+}19, ADK22, BÁJG^{+}20, HXCZ14, MFGEL17, YLT^{+}23]. runtimes [OK22, RHWF23].

S [ADE84, BEK95b, WH94, Taf01]. S-1 [ADE84]. S/NET [WH94]. S2SS [MSS^{+}05]. S2SS-MIN [MSS^{+}05]. SaC [GS06]. saddle [DNL15]. safety [GT07]. SAGE [NDW^{+}19]. salesman [AC89, Gur88, JK91]. Salesperson [Bur04]. Salpeter [PMV^{+}20]. samples [ED04]. sampling [LLS^{+}93]. SAN [GLD08]. SAN-based [GLD08]. Sandia [W^{+}99]. satisfiability [BSK03, CW06]. saturation [TM09]. saving [ACM^{+}10]. SBAC [SN14]. SC16 [BCE^{+}17, HG17, HH17, WBN^{+}17]. SC17 [DSSD18]. SC’18 [NGM19]. ScaffCC [JPK^{+}15]. Scalability [ALD01, Ber00, DLM97, IUY^{+}04, MWR95, RSRM96, ARR19, AW13, BMP15, GHW08, GSMY^{+}07, HRT07, JJAP^{+}20, KNTG08, LTV96, OID^{+}12, RSK99, SG02a, SHN12, VTmL03, VLSPL19, WLN^{+}96, dSS09]. Scalable [BRWL09, C^{+}01b, DLS09, DWM^{+}01, Gor97, GKS98, GRBA15, HA97, JPK^{+}15, JP94, KTAB^{+}19, KTN^{+}14, KS07, MK12, NZ97, OKTR13, PVK^{+}22, RBMO20, RC17, SK19, WRB97, AALK01, Ano13i, ARB94, ARB95, BBB^{+}94, BY21, BB21, Bri95, CQ97, DWH^{+}08, DFM^{+}95, EPMPU02, FA11, GWWM09, GPLW17, GM98, HKW13, HAJK01, KMB^{+}18, KKW14, KG21, LW15, LHK15, Mic97, MRRP11, BHB21, PSS01, PYLE21a, PYLE21b, SS01, Trä12, THM^{+}95, VTmL03, Wak04, WCM^{+}14, WvR16, WTLW23, Xu07, YZWcF14, ZEC^{+}17, ZGL^{+}19, GHH^{+}03]. scalar [AE86, BCCS07, CDC^{+}87, KD10, Phi91]. Scale [BKL^{+}19, ASA16, B^{+}02, BCMG^{+}07, Bjø87, BG98, BR84, CLS93, Cér07, CS97b, CDC^{+}87, CLJ14, GAA^{+}13, HCJ03, HMP^{+}16, HGS10, IU87, IHM^{+}12, JBC^{+}07, JCWP07, LLFZ23, LK10, MWI^{+}15, MHE19, MSMC15b, MSZM14, MZ04, NHS^{+}95, NFG^{+}13, NR20, ORS90, OA08, ONB11, OID^{+}12, PD11, Qua92, RBP^{+}17, TA14, VRL03, WYT^{+}16, WP19, YNO97, YX07, YXQ^{+}21, YM05]. scale-free [HGS10]. scales [HSR^{+}14, SLS^{+}21]. Scaling [JG93, CHQORS18a, GGH^{+}22, LFL11, MWI^{+}15, MAJD17, MP95, MW91, SDMS12, Wor92]. scan [Gol86, SST09]. scan-line [Gol86]. scanning [GGL^{+}05]. scatter [GLMBMPMV03, GGV04, Trä18]. scattered [BG91a, CUSR88]. Scattering [BL99, FMR90, SSP^{+}98]. Scavenger [SW91]. SCC [DLO17]. SCC’17 [GHH18]. scenes [CM95, Lee97, RJ97]. SCFGs [XDZL10]. schedule [BGR00, DC15, MT90, Moo04]. scheduler [KG21, LCLA19, TS88, TDC19]. schedulers [CGHBS18, CIO^{+}17, KAM^{+}20]. schedules [BDK98]. Scheduling [Aub11, BDF^{+}00, BLR03, BDSd94, BCR00, CS97b, DG99, GB14, Kha12, LZ00, Mis87, MS99b, Ver99, Ver00, AD96, ABR03, APBcF16, AGLP06, ABMT99, AMP92, AGGG06, BEY00, BKS06, BM11, BGG^{+}99, BHR09, BD10, BPK12, BNS^{+}07, BDM99, BR99, BSD11, CLY^{+}19, CHQORS18b, CC92, Chr00b, CL05, Cor00, DM04, DT96, Eis89, ECLV12, GO88, GMS04, GNY00, HJ05, HK02a, KKT00, KK03, KK11, KS05, KJA15, KP96, KY94, LGC97, Lah00, LLP00, LKHL03, LSC^{+}15, LGD^{+}15, Liu86, Liu98, LND^{+}19, LWW^{+}21, MCY^{+}24, Mar92, MFS^{+}19, MCP^{+}14, Mun99, NL98, OS21, PVBR23, PW22, PRV08, PKR00, QJ06, RRP03, RC15, RT88, RBL97, SM13, SG02a, SWCBQ19, SLG^{+}22, SCL^{+}23, ST21, SS04, SLH^{+}13b, SR10, TRLD13, TYLL22, TE92, VLL90, VK17, WL13b, WZS^{+}14, WZCG14, WS94, WLZ^{+}23, YG93]. scheduling [YCS07, YZB^{+}23, ZS21, ZLPF16, ZRS13, BD95]. scheme [AALK01, AQ93, Ara03, CHH^{+}01, CSH^{+}11, CNK93, Cor00, CQ01, CGM05, FG89, KYS23, KY98, MC09, MRRP11, PLY02, SKN04, SS90, SNO99, WW92, WC94, WZS^{+}14, WH94, XXQ^{+}15, ZL94]. schemes [BSE88, BFM90, DFM99, Duf86, GLD08, HWB92, HHSM88, HHJP16, MT95, RHWF23, RV96, SDv98, YD01, ZY16, ZGG92]. Schlegl [HE88]. Schmidt [BO91, BLT^{+}22, OW90, Str08, WC91]. Schur [XKL^{+}22]. Schwarz [ESK88, LG09, Mei86, PHD16, Rod85, YCBD19]. Science [Ano88b, ABG^{+}03, AC00, BRS^{+}17, CDR^{+}95, SCM^{+}98, WCM^{+}14]. Scientific [BCL91, BDO17, Jou97, BC97, BOS^{+}09, BOG15, CCGG14, CTS02, CD98, CC16, DT97, DST01, Eck90, FDTZ04, GRP22, HST87, LTV96, MV87b, NS02, ONB11, ONOK13, OID^{+}12, Pel97, SC03, SC05, VK17, YPG03]. scientific-application [BOS^{+}09]. Scilab [C^{+}01a, C^{+}01a]. Scotch [CP08]. Scout [MIA^{+}07]. scramjet [CYXL18]. scratches [IT08]. Screen [Ban97]. Screen-parallel [Ban97]. screening [JBC^{+}07]. scripting [KPL^{+}12, WHW^{+}11]. scripting-based [KPL^{+}12]. SCS [WSL88]. SCS-40 [WSL88]. SDAFT [YZWcF14]. SDPARA [YFK03]. Search [Wal01, AFT23, ACM20, ABGC^{+}14, ACGL04, BBL^{+}16, BBQOQO00, BGM03, CGGG03, Dzw91, EMV^{+}18, FM15, GLMBMPMV03, GLS01, Loh96, MGCB^{+}10, PR90, PKR00, RRA11, SKC90, Tai91, THG98, TCT00, TCS04, TO89, ZHC^{+}23, DC15]. searches [MC09]. Searching [EAKT90, GÁVRRL18, PRW88, CCS94, HSN89, ME92a, Wal01]. second [ADF93, BGWW97]. secondary [Bjø87, TLS^{+}08, XDZL10]. secret [SBZ04]. secure [CLJ14, FLYL16, JCWP07]. Security [BGP^{+}97]. seed [CYDJ21]. seepage [YLCT18]. SegFAUlt [DLO17]. Segment [YZB^{+}23, ZL94, CF93, CRD02]. segmentation [AGB97, BCYB11, GWC^{+}99, LSA^{+}95, MCG98]. segmented [AL88, LV15, SPCB22]. Seidel [BP02]. Seismic [CFD^{+}16, DDF^{+}10, HB84, KN88, SSS99]. Selected [Ano22m, MTW07, PŚ00, TH20, JLY18, KHHC20]. Selecting [LSL02]. Selection [BFVRC14, SlWC89, CYDJ21, FMA17, GF03, KJA05, LG21, PMAL14, PGBF^{+}07, RA20, She92, Uhl96]. selective [CSW01, DCG^{+}07]. selectsort [HK90a]. Self [BFVRC14, DR15, GV06, NH11, SM13, ORS90, SSP^{+}98, TS88, WC94, WHW91, YCS07, ZSI02, da 90]. Self-adaptive [GV06]. Self-adaptivity [BFVRC14]. self-adjusting [WC94]. self-consistent [SSP^{+}98]. self-organization [da 90]. self-organizing [ORS90, WHW91]. self-scheduler [TS88]. self-scheduling [YCS07]. Self-similarity [NH11]. self-testing [ZSI02]. Self-tuning [DR15]. Semantic [Bak00, AKSS07, Vio04]. semi [DLPS92, PCL23, ZBG88]. semi-automatic [ZBG88]. semi-centralized [PCL23]. semi-systolic [DLPS92]. semiconductor [MW94, PZE94]. semidefinite [NYFK06, YFK03]. sensing [OA08, YPL12]. Sensitivity [GMS04]. sensors [OA08]. separable [CDGM96]. separation [XCR17]. September [Ano20p, Ano21r, Ano22l, Ano23i]. sequence [DEH^{+}11, EAR93, GK04, MC09, OGC^{+}15, TS09, ZHC^{+}23]. sequences [AvHL^{+}97, BMCA98, GÁVRRL18, HK06, Lec89]. sequential [AS86, FLW86, GW08, GM03, MA22, RVD10, TTO^{+}22, Wil85, Reg01]. serial [BF92, Hot89, Lan90, SMP11, Y^{+}99]. serialization [PGAA16]. series [CCRR91, Cor99, DLPS92, MS99b, OW94, Pop88]. server [BA02, CS97b, DFRC94, GLD08, JSC97]. servers [Che98b, GM98, GLP98, RBL97]. Service [SVC07, BK07, JCWP07, KG03]. services [BY21, BGK^{+}98]. set [BG21a, BG21b, BCC^{+}97a, BCC^{+}97b, CKM93, GP98, GHC04, HT94, HL99, Hwa04, Leu89, MSE93, SL90a, Ski02, THMH21, WYBdJ^{+}21, XCS05]. set-to-set [GP98]. set-up [XCS05]. sets [BE97, EAKT90, JON08, LC02, ME92a, ME93b, ME93a]. Several [AFK01, But92, CLA14]. SGD [MA22]. SGPM [WH22]. SHA [BBDN21]. SHA-256 [BBDN21]. SHA3 [FLYL16]. shadows [ISH03]. Shafer [WH93]. shallow [PAF^{+}97, SSL19]. shallow-water [SSL19]. Shape [DPSW00, HKSK97, ML00, MJ94, TKG97]. Shape-optimized [DPSW00]. shaped [LHR^{+}19, PMAL14, YR18]. Sharded [LNLK13]. sharding [LNLK13]. shared [APPG15, ATA^{+}16, AFD13, ADMV05, Bai88, BC04, BM01b, BD89b, Bro88b, BS96, BMG07, BFR93, CSH^{+}11, CG89b, DM02, ECG93, FHMS93, FM88, GSE^{+}15, Ger98, GK92, GHH^{+}03, GB14, GL97, HCT16, Jor86, KW90, Kan93, KK03, KK11, KJA15, KV99, KKŽ05, KZV97, LSL14, Lan99, LFS14, LSL02, LF00, MAB17, MMTU18, Mat95, MSSC10, MBK12, OILZ17, PLY02, PK05, PRV87, Phi91, RGDM15, SG02a, SB94b, SSOB03, SPMB23, VA14, WC94, WY11, WE91b, WE92, Wri91, YXQ^{+}21, ZE92]. shared- [LF00]. shared-data [ATA^{+}16]. shared-memory [BC04, BMG07, CG89b, ECG93, GB14, GL97, KK11, LSL14, PK05, SB94b]. sharing [AFJG06, DGBP05, Fig06, PRS^{+}14, SCL^{+}23, YX07]. sharper [GT90]. shear [PCA01, SL03]. shear-warp [SL03]. sheared [FB19]. shells [CEM^{+}99, Wan09]. Shift [Mak94b, ALH^{+}14, MM95, SNO99]. shift-invert [ALH^{+}14]. Shift-net [Mak94b]. shooting [Gol88, KH92, Kie94]. shop [ABR03, MMT06]. Short [Bek95a, BD95, CC95, Cho92, CNK93, De 95, GE92, HC92a, KH95, KK92, KC93b, Oed92, SSY02, SW91, VP95, ZK92, ZL94, FCZY17, GVH08, HE88, KS87, KAD09, ZGG92]. short-range [GVH08]. short-vector [KAD09]. shortest [AÖ22, CCS87, EHE92, GS04, GT90, HTB01, LNS03, PB94, Trä95, TF01]. shortest-path [AÖ22]. shower [MB88]. shrinking [RS92]. shuffle [CS89a]. shuffle-exchange [CS89a]. SI [AMN16, VBC19]. side [MBC97, TDC19]. side-effects [MBC97]. sided [ACH^{+}18, DPFT19, RVGG01, YCBD19]. sides [CLA14, KMB^{+}18, Ter10, WP19]. SIEMENS [EHHS89]. SIFT [CYDJ21]. signal [FFLV96]. signaling [MZ04]. signatures [PW87]. silent [CHR18]. Silicon [LHZ^{+}20, LHZ^{+}22]. Silicon-Crystal [LHZ^{+}22]. SilkRoad [PWY03]. Sim [LC17]. SIMD [BE89, BDK96, CC95, Chr98, DEGS95, DJH05, GK92, HXW^{+}13, JK91, JCL92, Kas85, KCN99, KAD09, Li91, MD88, MP95, MW94, SSL03, WZ90, ZRPI89, ZBG88]. SIMD/MIMD [Kas85]. SimGrid [KdBMA^{+}13]. similarity [BG21a, BG21b, JNWJ18, MC09, NH11]. Simple [BGV97, LLP00, Lau93, RBL97, BC19b, IB01, JH97, SS89b, Y^{+}99, GP86]. simplex [MM00]. simplicity [San02]. Simplified [Rob00, CK90, GO88]. SIMULA [PM89]. simulate [EGTD99, PBTC89]. simulated [AC89, BL93a, DW00, JK91, LR16, MG95, OWW16, WWR05]. Simulating [TO91, ZLPF16, MSOCG^{+}16]. Simulation [CCCP92, DS03, DS92, HSR^{+}14, Lan90, WLK97, Amm89, ACC^{+}88, AMMV98, ADR^{+}05, BBK90, BM01a, BJP^{+}89, BDK06, BG98, CE94, CYXL18, CFC99, CDK^{+}03, CW97, CIG^{+}21, CT88, CLZ01, CQ01, CCP^{+}21, DRJSW97, DG95b, DBVS01, DF98, DLPS92, DAG^{+}09, ED04, FLPG18, FMS^{+}06, FC05, GJMMS97, GPLW17, GAW96, HM01, HGS10, KD97, KHN01, LLPV06, LG01, LFS^{+}19, LHZ^{+}20, Lin93, LXD^{+}23, Luc01, MD88, Mic90, MB88, NR01, OL86, OCE^{+}07, OCSBY01, PM89, PZE94, POHS14, PCC19, PJW^{+}22, RBB17, SSLK00, SWW99, SAWH88, SM91, SML^{+}14, SP93, STKA96, TKG97, TS99, TB07, TCT00, TBM^{+}16b, WNES01, WS04, ZLM98, Zha95]. simulation-generated [DLPS92]. Simulations [BKL^{+}19, WHL19, AAV08, BKS19, Bis84, BT01, BBL^{+}19, CLL99, CFD^{+}16, DTV21, DWM^{+}01, DDF^{+}10, EZ99, FHK^{+}15, GCC19, GKS98, GVH08, Gup99, HZY^{+}19, HBC19, KT00, Koh95, LLP00, LVC16, LBWR90, LYC^{+}19, LS93, LK10, MSMC15b, MZ04, MO99, OAJ^{+}16, ORS90, Par87a, PCA01, Pea19, SHvAP13, SKE^{+}22, Sch92, SR97, VLSPL19, XLY^{+}20, XRW^{+}19, YR18, Yaş01, WH15]. simulator [ADE84, KdBMA^{+}13, LKP24, MW94, TTT^{+}92, CFN03, HUYK04, IUY^{+}04, Nak05, Sat04, SZ04, YS04]. simultaneous [DW00, ED91, Kat01]. Single [HMS^{+}07, CCS87, DGNP88, HT92, Trä95, THMH21]. single-channel [HT92]. single-program-multiple-data [DGNP88]. single-source [Trä95]. singular [Bis89, KS91a, KN09, Lan99, TRSC^{+}19]. SISC [RGT17]. situ [OAJ^{+}16]. Six [BKL^{+}19]. Size [DZ90, ADFQO19, CAS09, CS95, CHLS90, CTW14, CMP92, DDP90b, DCD97, HSR^{+}14, Lin91a, MW91, ZRPI89, ZS21]. size-optimal [CMP92]. skeletal [Col04, GV06]. skeleton [CMV^{+}06, CV02]. skeleton-based [CV02]. Skeletons [SG02b, BCB02, Col04, DA06, DMT06, DLdS06, FSCL06, GS06, HHOMR06, MHT06]. Skew [FTY^{+}20, PMV^{+}20]. skew-symmetric [PMV^{+}20]. skewed [BF95]. skewing [LX00]. skews [BD18]. skid [RH13]. SkIE [BDPV99]. SKIPPER [SG02b]. SKYHI [JKH95]. Skylake [BKS19, BKL^{+}19, HZY^{+}19, LYC^{+}19, WHL19, CLC^{+}18, LCE^{+}18, LLX^{+}18, SWR^{+}18]. slackers [ORM^{+}10]. slave [HP05, JMLBL98, PRV08, dSS09]. SLEPc [DR18]. slicing [LOST16, BBC^{+}11]. Slow [Dod90]. small [JHD^{+}22, MAH^{+}19]. smart [OA08]. Smith [JW09, NW96, Ste13]. SMO [MA22]. smoothing [PS89b]. SMP [Hat06, JKHK08, PB11, RDS13, SSOB03]. SMPs [ADPV03, NU05]. SMPSs [MRBQO14]. SNAIL [MSS^{+}05, Y^{+}99]. SNAIL-2 [MSS^{+}05]. socket [Gro19]. Sodan [Hol11]. soft [EMB89, ISH03, MF16, TYSF13, XXQ^{+}15]. soft-systolic [EMB89]. Software [BVC16, BVC19, PS20, RH13, SYP13, YS04, BPP10, BNK15, CCGG14, CD98, CCG^{+}17, CBV13, CWDG07, ES06, FGP23, GAA^{+}13, Gut88, HH89, HA05, JH97, KYS23, KKHY06b, KRW88, KKH04, LSL02, Nag90, NMMG13, NE01, Oya99, PSS01, PK05, R^{+}00, SP94, SKG02, SVC19, STT94, WPD01, tV96]. solid [KH97, XCR17]. Solution [Cap88, FBMV88, Wri90, ADF93, Ala95, ABMQO11, AAHF97, AGLP06, ADGS10, ABB^{+}11, AIIV98, BD89a, Bas94, BEK^{+}11, Bev89, Bon91, CL02, CW97, CL05, CV17, DWL^{+}12, DvdV99, DR03, EB88, ER89, Eva92, EW93, EY94, GKT^{+}15, GS89, Gao87, GHC04, Gur88, HJ97b, HHJP16, JP94, Kie94, KPS90, KKB93, LRH97, LLP00, LSS88, McB88a, Mel87, Mel88, MLM^{+}00, MS90, PRS^{+}14, PMV^{+}20, Pir93, Pis92, RS08, RP85, RSSS94, RO88, Sch87, SE98, Sun97a, THK14, VL05, Wag89, WCM^{+}14, XKL^{+}22, van87a]. solutions [CWB92, Far96, Lan96b, MNP93, MGSK87, PFTB15, RS88, Sab97, SHH^{+}97, SM90]. solve [ACM20, BGMT00, KR10, LQ92, RG92, TSCS14, WP19]. solved [LMH08]. solver [AJ95, ABB^{+}16, ADPV03, ADTV01, AIIV98, Bjø87, BB21, Bru91, CLL99, CTZ^{+}18, CLA14, CPR^{+}18, CCP^{+}21, CDJ08, DE09, DR95, DF00, EL95, FKM^{+}22, FA11, GYH94, GV01, GW95, GG10, GZ99, GPSK09, GLU03, HS90, HRR02, HJ97c, HGLR07, IHM^{+}12, KMB^{+}18, KYLH01, LSL14, LLW^{+}24, LC90, LPMD01, NP05, Not95, PCA01, Pet97, PS06, PHD16, PMLT03, PEO16, Rea90, RBMO20, SS88, SZS22, TW19, TSJB00, THM^{+}95, VKS^{+}15, ZZ03, di 97]. solvers [ADLL03, AGD^{+}17, ARB94, ARB95, BBK90, BLT^{+}22, BT88, Cv98, DR18, DS84, DS05, EB98, FHMS93, GS22, GHP10, HKL01, HJ97a, HEB96, KS16, KA96, LPNJR96, LV92, MS03b, MS04b, MS91, MSV15, Nak05, OKTR13, PGW16, RRS88, RBM^{+}16, SG16, SBO^{+}91, ŚDJ^{+}22, Vaj84, ZM94]. Solving [BBQOQO00, Bol86, BGB10, CSEK03, CMM^{+}88a, DZ90, DJ87, GR89, GMW96, Gro87, HSS07, KR00, KŠR04a, KŠR04b, LLPV06, LPH00, PG91, Reu88, SL90a, SW03, Won88, APRP97, BBB^{+}14, BG91b, BE93b, BWW89, BGM03, CW06, CB97, CKLM14, DM03, EM87, ED91, FF21, Fie96, HM89a, HK91, Jou89, KS89, KL90a, Koç97, LT09, Lop93, LNA06, MWH95, MQ89, ME90c, Mei85, MP11, Mil01, MHH97, MOF04, Nag88, PW89, PMV^{+}20, RR89, RM97, Rot95, SN88, SZ02, SE93, SM88, Stp93, Ter10, Ter13, TRSHB04, WB88]. SOM [HKSK97]. Some [BK97, Hir86, LTV96, LR89, MR86, Par87a, Tro94, BP97, Bro89, BM90, Car88a, DS84, DSS86, DNL15, KH95, Laf02, MRJ89, NBB^{+}02, RS87, Sýk84, TSJB00, Wei91, Zla88, ZKCL04]. SOME-Bus [ZKCL04]. sonar [GMF00]. Sophie [MC04]. SOR [BFM90, CKLM14, DWX^{+}12, Eva84, Gen87, Kam87, Mei86]. sort [CKS03, EY86, MP87]. sort-first [CKS03]. sorted [SC95]. sorters [Hor91]. Sorting [CJBK93, AS86, Car88a, Eva90, GK19, HK90a, HK90b, LLS^{+}93, LL90, LS92, MRT93, MRJ89, Qui88, RS87, SPCB22, Uml94, WE92, ZM96]. sorting-merging [Eva90]. source [CCS87, RGBC11, Trä95]. sources [TB03]. SP2 [GYL00, XH96]. Space [BBC^{+}11, BCMSW04, CFMM97, DE09, Dec00, DHS89, EM85, FWL03, Fig06, GZ99, GPSK09, KHHC20, MHE19, MS20, PEO16, SSOB03, Sur10, Kon11, NZHY11]. space-filling [GZ99]. space-sharing [Fig06]. spaces [Xue96]. Spacetime [BR90]. Spacetime-minimal [BR90]. spacetrees [WVHR16]. Spanning [FA96, BHK89, BDK96, DDP90b, DY91, HWW92, LS97, ME84, TY91b, Tou02, YTCW07]. Spark [FTY^{+}20, SLG^{+}22, Yaş01, CCP^{+}21, XZYQ21]. spark-ignited [Yaş01]. Sparse [AHBD18, BVWH14, DS13, DD87, Pet91, ASA16, Ala95, ABMQO11, ADPV03, ADLL03, AMP92, BRS97, BPS01, BF06, Bjø87, BETR17, BB21, BJV^{+}16, CGG01, CCL04, CPR^{+}18, CDJ08, DEGS95, DTR18, DS05, EL16, EK98, FG89, GMW96, GP90, GCH21, GLN89, GW95, GR01, GS22, GH90, HRR02, HCR01, JP94, KB20, KD10, KKB93, LB23, LKHL03, LLW^{+}24, LC99a, Liu86, Liu89, LV15, MM00, Mar14, Mat95, Mel88, MP02, MNP93, MS90, NMAB19, ORM^{+}10, OHZ98, Pet84, Pin91, RR89, RP85, RSD16, RZ95, Rot95, RR16, SSSG03, SA16, SZ02, SPMB23, SGP22, Sun97a, THK14, VFG12, WP88, Wag89, WE94, WOV^{+}09, XKL^{+}22, YL19, YB11, ZE92, ZJD^{+}23, ZG88, SLC^{+}16]. sparse-grid [EK98]. Spatial [AFT23, NSF^{+}22, FSKF06, PM03, WA03]. Spatial- [NSF^{+}22]. Spatial-aware [AFT23]. SPEC [Dix91]. Special [AMN16, AAGS18, Ano22m, ASSS11, AGKS15, AGKS16, BH14, BVC16, BC19a, BDG^{+}10, BOG15, CP17, CP18, CBV13, CT94, GHH18, GT19, HG17, KST02, MSS19, NGM19, PS20, RST11, RC17, SYP13, TBM16a, TH20, TFV16, Ano18k, BVC19, DF95b, DR03, KVW97b, RC14, SAGV21]. specialized [BDNP11]. specific [MPZS13]. specification [GO88, HK98, KKHY06b, LOST16]. specifications [RW89]. specified [CLZ01]. Specifying [BGK^{+}98, DD89]. SPECT [MDC^{+}08]. Spectral [LSF^{+}15, ALH^{+}14, BK94, Bas94, Car88b, Cha99, DWM^{+}01, FKM^{+}22, FR95, GGFF93, GPSK09, Kru97, MWI^{+}15, NP05, VR95]. spectral/finite [MWI^{+}15]. spectral/hp [GPSK09]. speculation [FCZY17, NB12, RGT17]. Speculative [LV15, GRP22, NB12]. Speed [ARB94, ARB95, DWB99, DW00, FKK^{+}06, GJMMS97, Hee97]. Speed-up [ARB94, ARB95, Hee97]. Speedup [OIK21, FLW86, HOSS91, Jan87, NHS^{+}95, YR18, Zub90, ZGG92]. speedups [CRT89, Fis91]. sphere [KPH97, SSL19]. spherical [CEM^{+}99]. Sphynx [ABGR21]. SPIKE [MP11, PS06, SML^{+}14]. spiking [PMS^{+}13, PJW^{+}22]. SpiNNaker [NFG^{+}13, NLP^{+}15]. Splicing [GRP22]. spline [BZ96]. Split [EAR93, AHBLR12, SCCP08, WBS06]. split-and-merge [AHBLR12]. split-step [SCCP08]. SPMD [Has01, MCP^{+}14, PRR03, SV97]. SpMV [AAS20]. spots [AB93]. spray [NC02]. SQL [HM02]. square [CUSR88, MR85, TT00]. square-roots [MR85]. squares [KCN99, LWS02, MS90, NP05, OHZ98, PGW16, Str08, Sun97a, YK08, Zhu93, ZG16]. squaring [EL91]. SR8000 [Tak03]. SS [IS18, RSSS94]. SS-CAA [IS18]. SSOR [Lip99a]. Stability [Ron84, Yal97, Gao87, YE94]. Stabilized [DS05]. stack [UN87]. stage [BGMT00, GTD18, KK11, Zho93]. stages [NAC^{+}14, SRS^{+}19]. staging [KA19]. stamping [DPFT19]. Standard [Sne88, ML20, Wal94a, Wal94b, GLDS96]. staples [Hio96]. star [AQ93, AED21, CWC00, Fu07, LS98, Miš98, SS95]. StarFish [FMB98]. starvation [Ann89]. State [LR88, Mic97, SWR^{+}18, LRH97, SKH^{+}12, XCR17]. State-of-the-art [LR88, SKH^{+}12]. State/NCAR [Mic97]. statement [BPK12]. statements [HvNJB12, LC02]. states [GWWM09, NHS^{+}95]. Static [BRRV11, BDRV99, CHQORS18b, SS94a, ČSR05, Gir02, HAK^{+}21, KY98, MJ95, MFS^{+}19, NMI^{+}24, YEC97]. stationary [Beb97]. Statistical [DLPS92, CTW14]. statistics [FL21, POHS14]. status [DDdSL02, MVZ98, Pei88]. steady [LRH97]. stealing [LNLK13, YMY^{+}19]. Steiner [DUG^{+}06]. stencil [KD13, PVK^{+}22, XXQ^{+}15]. stencil-based [XXQ^{+}15]. stencils [GGL21]. step [BSE88, BKS06, CG89a, CS96, CGMM99, Don05, SCCP08, WE97]. stepping [Sur10, WBD99b, WBD99a]. steps [ALTZ02, Mar92, Müh90]. stiff [DF00]. stiffness [Kra84]. Stochastic [Güs99, BP93, BP95, BGMT00, BG07, Blo03, Dod89, LMH08, MA22, NZ97, OK22, SHPA05, TLC^{+}21, ZCPT00]. Stokes [AIIV98, Bas94, DLR94, FKM^{+}22, GV01, LSS88, OJ90, PATC99, PCA01, VvBv90]. Storage [AGL08, NDW^{+}19, AAS20, BAB^{+}02, Bjø87, BWT^{+}08, Che98b, DBZ^{+}19, DCD97, DWH^{+}08, FG89, FMB98, GKS07, LF98, LFS14, LCLA19, LWJ^{+}17, ND95, ONOK13, PCN10, RGGP^{+}18, SC03, SDv98, WL13b, XLS^{+}17, ZJDW18, DWH^{+}08]. store [ZP16]. storing [KS91a]. straggler [TDC19]. straggler-aware [TDC19]. Straightforward [GGNY11, Thu91]. strategem [CUSR88]. Strategies [CT88, FLS^{+}21, Ran97, AFK01, BRRV11, BG07, BCG00, BLCR21a, BLCR21b, DW99, Fer97, FMD98, Gin99, GV08, KM89, KHHC20, Lan96b, LLP00, LDK16, Lil94, LGD^{+}15, LBGO23, Loh96, LK10, MSK14, PHD16, PP98, Ran00, SMP11]. strategy [AC94a, Buz86, DF00, HW94, KJA05, LZCT15, ML00, MMT07, MX07, NC02, PTGF20, PCA01, PCL23, SOH94, Thu90]. Stratix [LLVM21b, LLVM21a]. stream [ACMT20, BCRSR11, JSC97, KA19, LGMdRA^{+}19, LBH07]. stream-processing [BCRSR11]. Streaming [FTY^{+}20, ACMT20, HMS^{+}07, PGGP19, SDMS12, TB22]. streams [CS97b]. stride [YE94]. strides [EY94]. string [GE93]. strings [CR23, IR00]. striped [Mel88]. striping [TF98]. strong [FC05, MWI^{+}15, Sch91a]. strongly [CCS^{+}18, GG10, KKW14]. structural [Bjø87, CMPM^{+}15, LR99, LVC16]. Structure [THK14, ABB^{+}11, BF92, Ber00, BCC^{+}97b, GMMT16, GEvGCP09, GM04a, GM04b, LSA^{+}95, MP95, N^{+}00, PIG^{+}16, RWT97, SKE^{+}22, XDZL10, da 90]. Structure-adaptive [THK14]. structure-based [SKE^{+}22]. Structured [DEGS95, WW90, BN01, CKRZ98, CTZ^{+}18, CV02, DAD11, EJLC00, HAK^{+}21, IJCL96, PMLT03, Ran00, RBMO20, SW03, WE94]. structures [ATT92, Bro86, Hag91, Han85, LH02, NZ92, TLS^{+}08, WE91b]. Structuring [Jor86]. Student [CLC^{+}18, HG17, MBB^{+}18, NGM19, SWR^{+}18, TSLL17, DSSD18, HH17, WBN^{+}17, AGV^{+}17, BKS19, BKL^{+}19, BBL^{+}19, FYH^{+}18, HZY^{+}19, LCE^{+}18, LLX^{+}18, LYC^{+}19, LHL18, WHL19]. studies [Cal85, GGFF93, TTT^{+}92, WLK97]. Study [PK05, PGAA16, ANG^{+}20, AMMV98, AGK98, BGG^{+}99, Ber00, BF06, CDM06, Chr98, CEM^{+}99, CLZ01, GCC19, GHW08, GW08, GH98, GLD08, Kas84, LSL14, LyHW^{+}22, Lef97, LCD91, LBWR90, LF88, MMT06, MAM^{+}09, MST94, MOF04, NAC^{+}14, N^{+}00, RSRM96, RSK99, SEAH^{+}08, SB94a, SSLK00, SWYM17, SH88, Sod02, SS94b, TCT00, TO89, Wal01, WLK^{+}18, YM05, ZY16]. studying [TLS^{+}08]. Sub [DM04, ES88]. sub-implicit [ES88]. Subcube [AC94b, CS97a, CGQ93, GF03, KSP97]. subcube-grid [CGQ93]. subcube-ring [CS97a]. subcubes [EAR93]. subdomain [KS01, PHD16]. subdomains [mMvdV01]. subgraphs [FA96]. subject [MS99b, Ver99]. submatrix [CG92]. suboptimal [CKM93]. subprograms [LN87]. subroutine [DZG94]. subsequence [Lin94, TNZM20, TTO^{+}22]. subset [GK03]. subspace [LR99, de 96]. substitution [CS89b]. substitutions [LC89]. substrate [AFY^{+}16]. substring [TNZM20, TTO^{+}22]. substring-exclusion [TNZM20]. subsystem [BMG07]. subtasking [HS89]. subtasks [SSS^{+}05]. subtraction [ST95]. subword [DCDV98]. Successive [EL90, CKM94, EMB89, PJ84]. Sufficient [ÜD89]. suffix [ABGC^{+}14, KS07, ZGL^{+}19]. suitable [Blo03, LG09, MAS06, PJV00]. suite [DFRR91, GSE^{+}15, LXC^{+}18, MKL^{+}01, PB23, TG09]. sum [HA14, LV15, TDW03, Wen95]. Sumatra [BKS19, BKL^{+}19, BBL^{+}19, HZY^{+}19, LYC^{+}19, WHL19]. Summary [Kau94, FL21]. summation [HSS09, LHG^{+}23]. Summer [KB08]. Summit [BC19b]. sums [Obe85]. SUNDIALS [BGWR21]. Sunway [JCC^{+}24, LLW^{+}24, LHZ^{+}22, ZHC^{+}23]. super [MTV08, NHS^{+}95, SOS97]. super- [SOS97]. super-/hypersonic [SOS97]. super-linear [NHS^{+}95]. super-peer [MTV08]. SUPERB [ZBG88, ZBC94]. supercell [OWW16]. supercomputer [BJV^{+}16, CYXL18, CDC^{+}87, DK01, Emm84, Erc88, Gil88, Gil94b, Gil94a, GL97, GAW96, JCC^{+}24, SL90a, Wat87]. Supercomputers [Duf84, BS00a, BWV^{+}17, BLCR21a, BLCR21b, DZM^{+}13, EHHS89, Gra91, HL88, Hoc91, Hos88, MRSB94, Oya99, RD07, SH91, TKI85, TRLD13, WLCG02, ZEC^{+}17].

Supercomputing [PB98, YK87, BH92]. Supercomputing-based [PB98]. superconductivity [KSM^{+}94]. supercore [vWWM^{+}19]. supercube [LH02]. Superlinear [FLW86, YR18, Fis91, Jan87]. Supernodal [Ng93]. supernodes [HRR08a]. superscalar [SC92]. supersonic [JJAP^{+}20]. Support [CCL04, SYP13, ADK22, AART13, BPJ22, DLR94, Evr01, FAS02, HT00, JMA^{+}13, KKHY06b, KT99, MCB16, MSE19, MA22, PTK04, SBP12, VJ11, ZZ03, ZPAT99]. Supporting [FSG19, DF98, FK98, SK97b, TDW03, LK00]. supremum [SB92b]. SUPRENUM [Ano88a, Ale91, Ale94, Gil88, Gil94a, Gut88, Hem94, Kam88, KK92, KM88, LH88, McB94b, McB94c, MST94, PSS94, Pei88, Sch88b, SAWH88, ST88, Tro94]. SUPRENUM-like [Ale91]. surface [C^{+}01b, DVP90]. surfaces [DG15]. surfactants [STKA96]. surgery [GBH98]. Survey [GLP98, FGP23, Ghe85, HJB^{+}21, HMH^{+}13, LJD93, TKC^{+}14, WXL^{+}18]. sustainable [TRLD13]. SVD [BOV02, BO03, BOVG10, BTLK18, GTD18, LLD19, NSS15, OV06]. SVM [MA22]. SVM-SMO-SGD [MA22]. SW26010 [LXC^{+}18]. SWAMP [Ste13]. Swarm [VK17, HF20, LW11, SYAU07, SSN04]. swarm-based [SSN04]. swarms [HF20]. Swift [WHW^{+}11]. switch [CB17, Kol94, LWW^{+}21, Qua92, WWL^{+}22]. switched [Sha06]. switches [Jes88]. switching [Ann89]. SX [TSEE21, Wat87]. SX-Aurora [TSEE21]. SYCL [LG21]. Symbolic [GHNL87, GH90, JCE96, Kan93, KKB92, Ng93, ZG88]. symmetric [ABB^{+}11, AHW14, BBB^{+}14, BW92, BC04, Eva92, GR01, Gut88, Hat06, HRR02, JLY18, Kal90, Kal92, KG87, KD10, LWS02, LQ92, MRW^{+}08, ML23, Mar14, Nat90, PMV^{+}20, Pet84, PB11, Pin91, RS08, SK19, SMK20, SZS22, THM^{+}95, WC90, ZVWS88]. symmetric-multiprocessor [Hat06]. Symmetry [LJD93, DL97, NTHY22]. Symposium [Ano16k, Emm84]. Sync [MGCB^{+}10]. Sync/Async [MGCB^{+}10]. synch [BLT^{+}22]. Synchronised [FM88]. Synchronization [ATT89, FJS85, Gre89, Hoc89, Axe86, BRWL09, BMG07, FK98, GV14, GEvGCP09, GF13, HKW13, HvNJB12, LL98, Nic95, YMG14, YLW^{+}13]. synchronized [Y^{+}99]. synchronizing [CWDG07]. Synchronous [CZ91, BC91, Bol86, EHE92, Kan97, KS99, LY93, MT89, SRS^{+}19, WWL^{+}22]. Synergy [HM89b]. syntax [Num05]. Synthesis [CC91, CMP92, DTV21, GM89, FF21, HA97, Lin91b, MS00, WE97, da 90]. Synthesizing [RF90]. synthetic [FFLV96]. System [Eck90, Oed92, PS20, ADK22, AAB^{+}16, AALK01, AC03, A^{+}01a, AGO97, AE93, BCGS93, BF95, BS97, BV96, Bis89, BDSd94, BJP^{+}89, Bon91, BOS^{+}09, Bri95, BL94b, BGL01, CDDG93, CD98, Cv98, Che98a, CS89c, CW06, CDV08, DBZ^{+}19, DPFT19, DS84, Eva92, DST15, FLYL16, FBMV88, FK94, FMB98, GW95, GHS02, GSC^{+}22, HUYK04, HKSK97, HCH^{+}96, Hoc89, IM88, JH97, JLLC22, JBWE14, Jou89, KMB^{+}18, KRW88, KRN^{+}11, KL90a, KLN90, KTN^{+}14, Kol94, KKB93, KAM^{+}20, Lef97, LBWR90, LC90, LBH07, mM94, MP92, MCC04, MTV08, MAS06, MSSC10, MST94, MK97, MTPE90, NL91, NS02, NK97, PK05, PGK^{+}18, Pir96, Pis92, PS06, PGG11, SKN04, SXBD97, SBM^{+}98, Sch88b, SC03, SSL03, SSS^{+}09, ST88, Str87, SE98, SGDM94, TMCC02, Ter10, UAR^{+}99, VKH99, WC94, WYT^{+}16, Wat87]. system [WB97, WS04, YXQ^{+}21, YS04, ZM94, ZWJ^{+}19, de 87, AL93b, CJLS14, GHH^{+}03, NMW93]. system-level [BS97]. System/6000 [AL93b, NMW93]. Systematic [MS00, ANG^{+}20, Kot97, LEH14, ZLJ93]. Systemic [TCS04, RBS10]. Systems [BVC16, BVC19, HKK97b, KB08, PS20, PTS^{+}12, RC17, SA18, ASA16, ABGR21, AGH^{+}94, AAB^{+}16, AAV08, AAO13, ACMT20, Ale94, ABMQO11, AGLP06, AB95, ATL^{+}12, AART13, AB93, ASH92, AE93, BBB^{+}14, BG91b, Bak00, BBB^{+}94, BPEL05, BEY00, BMS01a, BGWW97, BZ96, BAR98, BHR09, BGP^{+}97, BNS^{+}07, BSD11, BP90, BDH95, BJSN04, Bro89, Bru91, CE94, CCGG14, CS00, CPdM^{+}97, CGMS94, CG93, CLC08, CTZ^{+}18, CL02, CBV13, CTW14, CLA14, CPR^{+}18, CS87, CSH^{+}11, CFW^{+}22, CS96, CB97, CBP^{+}07, CC99a, CGM^{+}92, CV17, CN98, CMM^{+}88a, CTT89, CDH^{+}03, CRSS99, CGM05, DY91, DRJSW97, DIR97, Don05, DJ87, DNL15, DG94, DvdV99, Eis89, ESV10, EO91, ECLV12, ESK88, EB88, EW93, EY94, Fie96, FGP23, FHKZ88, GP85, GKT^{+}15, GR89, GMW96, GJMMS97, GGH^{+}22, GF03, GH90]. systems [GAA^{+}13, GV08, GYL00, GG10, GPPS99, GLS99, Hag91, HJ05, HS90, HKK97a, HRR02, Her88, HMH^{+}13, IKK15, JAA06, JJM^{+}11, JOL^{+}21, JIC96, JP94, KS89, Kas85, KD97, KKH04, KLP11, KJA15, Kha12, KBG^{+}01, Koç97, KG03, KM88, KR00, KCN99, Kot97, KM89, KPS90, LLPV06, Li89, LT09, LDK16, LLD19, LSL02, LCLA19, Lin01, Liu98, LT19, LZCT15, LPAZ97, LQ92, LMKH97, MD04, MS03a, MKS91, MRW^{+}08, MHE19, Mak94b, MT89, MMO07, Meh93, Mei85, MT00, Mel87, Mel88, MP11, MAB^{+}21, MNP93, Mis87, MK12, MOF04, ME92c, MGT^{+}13, MGSK87, NMMG13, OS21, OR88, PGGP19, Pap98, PG91, PY00, PLY02, PW89, PCN10, PP98, QJ06, Qua92, RRS88, RR89, RS08, Rea90, RP85, Reu88, RCAP11, RMACG10, RT88, RO88, Rot95, RBP^{+}17, SEAH^{+}08, SSSG03, SS02]. systems [SBP12, SAOKZ06, SG02a, Sch87, SS94a, She93, SZ02, SZS22, SVC19, SSD01, SPMB23, Slo91, SE93, SBM^{+}22, Ste90, Stp93, Sýk84, Tak06, TDC19, TYLL22, Ter13, TDW^{+}04, TDB10, THK14, TE92, TJ97, VL05, VA14, Wag89, WB88, WE93, WE94, WLK97, WL13b, WWZ^{+}18, Wei97, WHW91, WP19, Yal97, YX07, YMG14, YLW^{+}13, YSBM97, YK87, ZSI02, ZYZ^{+}16, vWWM^{+}19, van87a, van87b]. Systolic [AS86, AU88, CTT89, DZ90, DLPS92, EMB89, EM90, EA95, KC93a, KKSS90, KS91b, Li90, ME90a, ME92a, ME90c, MBC92, MS89, MKC92, SMK91, SK91, TGE92, BE87, Bek95a, BR89, BR90, CS89a, DVP90, DS93, DQRR00, EM86, EM87, Eva89, Eva92, EG94b, FF92, GM89, GE93, Hor91, HHH92, Jou89, Kat86, KLS^{+}88, LT94, LS90, Lin90, Lin91a, Lin94, Lip99b, LPPS01, LT90, MO89, Mar95, ME90b, Meg90b, ME92b, MF16, Möl99, MRSB94, PS89b, RF90, Ris90, SS89b, SS91a, SS91b, SM87, Tor91, TH89, TL90, UN87, Ume89, WE93, WE94, ZLJ93, ZM88b].

T [KSM^{+}94, Pop88, TM90, VFG12]. T-RACK [TM90]. T-Series [Pop88]. T20 [CFH89]. T3D [CC96a, HEB96, WLN^{+}96, XH96]. T3E [GYL00, ZCPT00]. table [CDA^{+}16]. tables [BRS^{+}08, DM90, EM87, QP16]. taboo [Tai91]. tabu [BGM03, CGGG03, GLS01, PKR00, THG98, ACGL04]. TaihuLight [LHZ^{+}20, LHZ^{+}22, ZHC^{+}23]. Talbot [dGR95]. tangent [CARW91]. Tank [SH90]. TaPIA [TFV16]. target [GTW^{+}20, HKW13]. Targeting [BC19b, TLC23, ZG16]. targets [VBS^{+}15]. Task [BFH23, BTZ06, Cal85, GBH98, GHM97, LKHL03, SZS22, APBcF16, ARCH05, ABB^{+}16, ADEQO19, Ara03, BBL^{+}16, BKS06, BM11, BDM99, BR99, BÁJG^{+}20, CHR18, CGHBS18, CBP^{+}07, CGH^{+}19, CIO^{+}17, DT96, EB93, GRP22, GV08, GV06, GF13, HJ05, JDH96, KKT00, KK03, Kd88, KY98, KS05, LX00, Liu86, LWW^{+}21, LZ00, MKS91, Man01b, Mar92, MRBQO14, NAC^{+}14, NJ02, OS21, PW22, PKR00, RRP03, RMS^{+}18b, ROZ01, STP^{+}19, SS94a, SLG^{+}22, SLH^{+}13b, TSLL17, VB95, WZS^{+}14, WvR16, WS94, ZLT19, RHWF23]. task-based [BÁJG^{+}20, CGH^{+}19, RMS^{+}18b, STP^{+}19, ZLT19]. task-graphs [LZ00]. task-level [GRP22]. Task-oriented [GBH98]. Task-parallel [SZS22, APBcF16]. task-parallelism [MRBQO14]. task-transfer [EB93]. task-uncoordinated [WvR16]. task/data [ADEQO19]. task/data-parallel [ADEQO19]. Taskgraph [Hur93]. tasking [MSW91, VLSPL19]. TaskInsight [CGHBS18]. Tasks [RR14, ABMT99, ARR19, Aub11, BDF^{+}00, BHR09, BDSd94, DC15, DRST03, Fig06, MT90, Moo04, QJ06, SSN04, TS88, TYLL22, VB92, Ver99, YZB^{+}23, RR16, dSS09]. TAU [RMS^{+}18a]. Tausch [SBM^{+}22]. Taxonomy [Kut85, Wor91]. TC [SM13]. TC/SL [SM13]. TCASM [OILZ17]. TDDFT [JCC^{+}24]. Team [BKS19, HZY^{+}19, WHL19, BKL^{+}19, BBL^{+}19, CLC^{+}18, CGOP03, LLX^{+}18, LYC^{+}19, LHL18, MBB^{+}18, DLO17, DSSD18, FYH^{+}18, LCE^{+}18, SWR^{+}18]. Technical [DSSD18]. technique [AAHF97, BZ96, CCS94, CHLO85, Glo95a, HOSS91, MFGEL17, Sil88, TB03, TE92]. Techniques [Lee95, AAS13, AFPG12, Ano99b, AATK10, AGK98, CDRV97, DR95, DBVS01, DW99, FGP23, FBG^{+}12, GLP98, GPPS99, Güs99, JSC97, KCRB98, KBGZ88, Loh96, Lop93, MAH^{+}19, MSW98, Pet91, PHCR05, Psa02, RE98, RH12, RH13, SWW99, SZ02, TKC^{+}14, YS96, YTCS97, CAA98]. Technological [LHL18, WHL19]. technologies [BOV09, CDK^{+}03, LNA06, VRGÁ15]. Technology [BKS19, BFG^{+}07, BRS^{+}17, IU87, TLS^{+}08]. telecommunication [BJSN04]. tempering [LMG09]. templates [PUSS97, TPK^{+}14].

Temporal [TF01, YLT^{+}23]. tenancy [ZP16]. tenant [MCY^{+}24]. Tensor [HBC19, GZH^{+}23, JHD^{+}22, KKU16, KKSS90, KS91b, Num05, SPK18]. Tensor-optimized [HBC19]. tera [Xu07]. tera-scalable [Xu07]. teraflop [HC05]. Teraflops [Maw99]. TERM [OTI^{+}89]. Termination [GGL99, DIR97, SB96, Son92, TDW03]. terms [OKF16]. Tersoff [CLC^{+}18, DSSD18, FYH^{+}18, LCE^{+}18, LLX^{+}18, LHL18, MBB^{+}18, SWR^{+}18]. tertiary [Che98b, FMB98]. tessellations [JDG02, WTLW23]. Test [LEH14, TG09, AKNS91, BCGS93, CC98, CC00a, CC01, CCW01, DFRR91, HISS92]. testbed [TYKA95, WH97]. Testing [SMC03, OKSY92, ZSI02]. Tests [MSE93]. tetrahedral [Glo95a, OBG00]. Texas [SWR^{+}18]. text [ABGC^{+}14, PW87]. their [BRH90, CL92, GCC19, Gil94b, Güs99, JDG02]. theorem [CF91, FC18]. theoretic [BM08, HWYL89]. Theoretical [BA95, BMS01a]. theory [BP89, Gup99, Hio96, LPH00, M^{+}00, PTGF20, WYBdJ^{+}21, GAW96, TFV16]. Thermal [SWCBQ19, FB19, HSS07, ZCT^{+}20]. Thermal-constrained [SWCBQ19]. ThermoBench [ZCT^{+}20]. thin [TOO20]. thinning [HW91, HS89, NOS92]. third [WLN^{+}96]. third-party [WLN^{+}96]. Thread [CAC^{+}09, GT07, LLL10, MCP^{+}14, AKB^{+}19, ABMQO11, CDA^{+}16, DCN^{+}15, GCH21, HCT16, IDS16, MAJD17, RGT17, VJ11]. thread-adaptive [GCH21]. Thread-based [CAC^{+}09, AKB^{+}19]. thread-level [ABMQO11, RGT17]. Thread-safety [GT07]. threaded [CSB00, KKH04, MP11]. threading [BFG^{+}07, FWL03]. threads [LSL02]. Three [Ram07, ABMT99, Bas94, BP86, DG99, DDF^{+}10, FB19, GHW08, GICM18, GGL21, HL88, JWS13, LPCA98, LS92, MSA09, OJ90, PSW18, SBO^{+}91, WBS06]. three- [LPCA98]. three-dimensional [Bas94, BP86, DG99, DDF^{+}10, GGL21, JWS13, MSA09, OJ90, PSW18, SBO^{+}91, WBS06]. three-point [GICM18]. threshold [TDG^{+}18]. throughput [GK19, KA19, MC09, MHL06, PGG11, TPK^{+}14, VCK^{+}11, ZLPF16]. thunderstorm [OWW16]. Tianhe [CYXL18]. Tianhe-2 [CYXL18]. Tight [TYLL22]. tightly [Eva90, Qui88, WE91b]. tile [RRP03]. tiled [BLKD09, CKS03, LBH07, SZS22, TX00]. tiling [BDRV99, DWX^{+}12, GDAK06, XC02]. Time [BCS15, MPZS13, XC02, AGB97, BSE88, BGWR21, BO91, BE93a, Beb97, BW01, BP95, BCCS07, BFH23, BDH95, BG98, CLS93, CCGG14, CDV08, CCRR91, Cor99, CCH^{+}21, DLR94, DR95, DPFT19, DWB99, Don05, DLPS92, DST15, GE11, GLS01, GMF00, HJ05, HSR^{+}14, HT00, HK92b, IHM^{+}12, Joh93, Kim98, KI05, KPL^{+}12, LK00, LT19, MF16, NA03, NSF^{+}22, OSZ93, QJ06, Ram07, RRP03, RSK99, Rau98, RHM^{+}88, SM13, SSL19, SWCBQ19, SR17, SC05, SMW^{+}05, Sur10, TKK^{+}05, TYLL22, TDW^{+}04, Ume89, Ume01, Wan08, Wan09, WBD99b, WBD99a, WL14, WH91, Wri90, YXQ^{+}21, YZB^{+}23, vWWM^{+}19, CQ01]. time- [NSF^{+}22]. time-consuming [DST15]. time-dependent [RSK99, SMW^{+}05]. Time-domain [BCS15, Ram07]. time-harmonic [IHM^{+}12]. Time-minimal [XC02]. time-optimal [BE93a]. time-optimum [Ume89]. time-parallel [HK92b]. time-stamping [DPFT19]. time-stepping [Sur10, WBD99b, WBD99a]. time-to-live [TKK^{+}05]. time-varying [BCCS07]. times [RvG05, SSS99]. timestamp [BRWL09]. Timing [Car88a, GKAS91, MRJ89, RS87]. TMC [SNS^{+}97]. today [Pan01]. Toeplitz [BE93b, CLA14, EY94, GKAS91, KL90a, LPNJR96, Pis92]. tokamak [TTT^{+}92]. token [HT98]. Tokyo [FC05]. tolerance [HXCZ14, ME89, NFG^{+}13, SHCS86, SC05, Tou02]. tolerant [ASDOK01, BHR09, CC02a, Cho92, FBD01, Far96, Fu03, Fu07, GAR15, GLP98, GP98, Hsi06, HG12, KL00, KH97, LW11, Loh96, LH04b, MLX07, PY00, QJ06, SS01, SR17, SW00, XCS05, Yua94, ZL04]. Tolerating [CS97a, Far94]. tomographic [BBB19, RW98, TAB^{+}19]. tomography [LPCA98, MRW^{+}08, MF16, Möl99, WH91, ZLM98]. tool [BDH95, Bra88, CSB00, CP08, EJLC97, FDTZ04, GWWM09, JBC^{+}04, JJAP^{+}20, Kre13, Lin91b, RMS^{+}18a, RVD10, SV97, ZBG88]. Tools [AGK98, BDS88, Jou97, KNP97, Ano13i, AKP^{+}98, BCC^{+}97a, BCC^{+}97b, CBV13, DT97, EHR^{+}98, EO91, HKK97a, IJCL96, KBGZ88, LMJC96, MVdLvN16, OGC^{+}15, Pan01, Pel97, RGdS^{+}13, SLH^{+}18]. toolset [DRJSW97]. top [CGH^{+}19, Lef97]. TOP500 [DZM^{+}13]. Topics [CP17]. Topological [EBSS94, MAB17]. topologies [BP97, Gro19, GDMS97, GHM97, KS05, PMAL14, PCLM18]. Topology [da 90, Fie96, GJMM18, Gor97, RIB10, YXQ^{+}21]. topology-aware [YXQ^{+}21]. tori [LOKM99, MSA09]. tornado [OWW16]. toroidal [CMP92]. torso [SHRN98]. torus [DT95, FA96, Hen93, HT98, KTN^{+}14, LD99, QCC02, RD07, Shi01, SH01, KJ07]. torus-wrap [Hen93]. total [RBM^{+}16]. Touchstone [Pir96, HC92a]. Trace [MK12, BDH95, ED04, FLPG18, Lin93, YS96]. trace-based [FLPG18]. trace-driven [Lin93]. traces [BRWL09]. Tracing [GCC19, AATK10, DGBP05, Lee97, MK12, RJ97, SGS95, SSV^{+}16, YEC97]. track [CGGG03, OWW16]. tracking [CARW91, CDV08, TB03]. TracSim [ZLPF16]. trade [DW99]. trade-off [DW99]. Tradeoffs [MB88]. Trading [DW00, OCPT97]. traffic [CW97, CLZ01, NS94, PJW^{+}22, SAOKM01, SP97, WNES01, WL98]. training [EPMPU02, Han98, HT94, QWD^{+}21a, QWD^{+}21b, ZZ03]. Trajectory [GRP22]. Traleika [CCG^{+}17]. transaction [RC15, WLK97, WH22]. Transactional [DR15, HL13, PGAA16, RC15]. transactions [WLZ^{+}23]. transcoding [Bar12]. transfer [A^{+}02, EB93, PGG11]. transform [ALTZ02, BFG^{+}07, Cha99, CC95, DZG94, FF92, HJ97b, Hol95, LHK^{+}96, Wan09, vCd90, AFS14, AW13, CT94]. transformation [ER95a, ER95b, MCG98, RW89, SCD92, TH89, Wil85, YS96]. Transformations [CFMM97, Xue96, BBC^{+}11, CMPM^{+}15, DQRR00, Dow90, EG92b, GDAK06, Hot89, Xue94, Xue97]. Transformed [DZ90]. transforming [Psa02]. Transforms [AV15, BAK09, Cha88, FIMF99, Heg96]. TRANSIMS [NR01]. translation [HCL05, XCR17]. translator [GScFM13]. transmission [CWK09]. Transparent [HXCZ14, CZTS99, GHS02, LFL11, SKG02]. transport [ARW93, BSH88, BLFT84, FHL87, HWB92, Mv88, Oli96, PFS^{+}04, TS21, Zla88]. transportation [FG01, HTB01]. transpose [CDW95, WBS06]. transposed [Mar14]. Transposing [Dow95]. Transposition [VB95, MP87]. Transputer [ARZ97, ACC^{+}88, BG91b, BP90, BGL^{+}88, CDDG93, CCCP92, CJBK93, CMM^{+}88b, Dod89, EC91, HP91, Kal90, Kal92, LBWR90, LS92, MP92, MRSB94, SS90, TY91a, WB88, Wel89, WH91, WH93, SOS97]. Transputer-based [ARZ97, CDDG93, CCCP92, MP92]. Transputers [Jes88, ARB94, ARB95, Bur90, DY91, MP87, WC91]. TransTOOL [BCC^{+}97a]. transversal [TGE92]. trapped [ZLPF16]. travel [SSS99]. travel-times [SSS99]. traveling [Gur88, JK91, Bur04]. travelling [AC89]. traversal [DCR^{+}16]. traversals [CL92, KB85]. traversed [CF04]. Treatment [Zla88, LP92]. Tree [HAK^{+}21, AJ95, ACM^{+}15, BV04, BDK96, Cho92, DY91, DL06, EHE92, EMV^{+}18, FHL87, Ger04, GMS04, GH92, GK04, HKSK97, HK02a, HWW92, HSN89, KB85, KKB92, LS97, LH02, NMARD10, PYZC11, SST09, SGS95, WTLW23, ZGL^{+}19]. tree-like [NMARD10]. tree-projection-based [GK04]. trees [AND05, BHK89, Bar97, BL99, FHJ^{+}84, GE11, GLG06, ME84, MHT06, Mun99, OA91, QP16, ST02, TY91b, Tou02, TT00, YTCW07]. Trends [IU87, YD07, BEY00, DvdV99, SDMS05, SGDM94, YDTS01]. trendsetter [Gil88]. tri [VP94]. tri- [VP94]. Triangular [LT90, Cv98, Fie96, KS16, KKB93, LLW^{+}24, ML20, ME89, NM01, RO88, RG92, Rot95, THK14]. triangularisation [Meg90b]. triangularizations [DR02]. Triangulating [DG15]. triangulation [AU88, KKŽ05, NR20]. tricks [LK14]. tridiagonal [AJ95, AB95, BW92, BE93b, BWW89, Bon91, Bru91, DR18, EY94, HS90, Kal90, KL90a, KPS90, LC90, Lin01, LPAZ97, LQ92, MWH95, Meh93, MS91, PB11, Pis92, RS08, Rea90, Reu88, SE93, SE98, Ter10, Ter13, TND10, THM^{+}95, VKS^{+}15, VL05, WB88, van87a, van87b]. tridiagonalization [CUSR88]. trigger [LTG23]. triggered [IZS^{+}20]. trivial [NAC^{+}14]. truncated [SML^{+}14]. truncating [Ram07]. Tsinghua [HZY^{+}19, LLX^{+}18]. TSUBASA [TSEE21]. Tsukuba [NNB^{+}99]. Tsunamigenic [BKS19, BKL^{+}19, BBL^{+}19, HZY^{+}19, WHL19, LYC^{+}19]. Tukey [JK92]. tuned [CCGG14, CGG04]. Tuning [NZHY11, POHS14, TTH09, CMSL06, CSB00, DR15, JHSV02, KKHY06b, KKHY06a, Kra84, LDK16, SWYM17, VFG12]. tunneling [GdCCS03]. turbulence [CIG^{+}21, MWI^{+}15, MRRP11, RKMJ10]. turbulent [CEM^{+}99, Xu07]. Turing [Wor97]. turn [FK98]. turning [IMQO^{+}18]. Two [ACH^{+}18, AIIV98, BP86, DDP90b, EM85, FCC16, Kon11, Mei86, SST09, SG02a, Tro00, WVHR16, WC90, YB11, Zho93, AL93b, ADLL03, BKS06, BE87, BGMT00, Che08, CL05, CWB92, CFH89, Eva91, GTD18, Gir02, God00, GMS04, GLS99, HAJK01, JCC^{+}24, KR97, Kak88, KW90, KK11, KY94, Li91, LND^{+}19, LC97, MM96a, MKL^{+}01, ML00, MSP93, MCM01, NR20, QCC02, RSK99, SSY10, SNK06, SZ02, Sim91, SE87, Stp93, TW19, Trä95, XLS^{+}17]. Two- [BP86]. Two-constraint [Kon11]. Two-dimensional [AIIV98, Tro00, YB11, AL93b, BE87, KR97, MM96a, MCM01, QCC02, RSK99, Sim91]. two-factor [God00]. two-layered [Kak88]. Two-level [FCC16, SG02a, LND^{+}19, ML00, NR20, SNK06, TW19, XLS^{+}17]. two-list [Che08, LC97, SSY10]. two-pass [Gir02, KY94]. Two-sided [ACH^{+}18]. Two-stage [Zho93, BGMT00, KK11]. two-step [BKS06]. Two-tree [SST09]. type [ML20, Slo91, Tak06, WLYJ13]. types [BDNP11, Li91].

U.S.A. [Hoc85a]. Ultracomputer [Bis84]. ultrascale [DBI^{+}17]. ultrasonic [GGFF93]. ultrasound [BMYK98, WH91]. unbalanced [Par91]. unbounded [LH02]. uncertainties [LMKH97]. uncertainty [MLM^{+}00, PŚ00]. unconstrained [CLS93]. uncoordinated [WvR16]. underrelaxation [EL90]. Understanding [SLH^{+}18, AAB^{+}16]. undirected [HHH92, SS89b]. Unicast [Miš98]. Unicast-based [Miš98]. unifications [VK92]. Unified [GSC^{+}22, LPAZ97, ATA^{+}16, CLJ14, GGA19]. unified-parallel-C [ATA^{+}16]. Uniform [Deá90, BDSd94, CC97a, DR94, VP95, Ver99]. uniformly [MSE07, Trä12]. unifying [AC94b, Lin01, SBM^{+}98]. Unimodular [Xue97, Dow90, Xue94]. uniqueness [Rao97]. unit [CARW91, DOD23, Kam88]. unite [WP19]. units [HSS09, Sur10, Wor97]. unity [Par86, FLW87]. univariate [OW97]. universal [She92]. University [BKS19, BJvOR03, BKL^{+}19, DSSD18, HZY^{+}19, LCE^{+}18, SWR^{+}18, WHL19, BBL^{+}19, FYH^{+}18, LLX^{+}18, LHL18, MBB^{+}18, NNB^{+}99, SWR^{+}18]. UNIX [Sch88a, DY91]. unknown [KSS06]. unrecoverable [BRRV11]. unreliable [GRBA15]. unsaturated [FMS^{+}06]. unsteady [DL04, GV01, KT97, KYLH01, TW19]. Unstructured [AFJG06, AKK16, BCCS07, BGR00, CGG01, CFN03, DR95, DL04, GK17, HJ97c, HB94, LLW^{+}24, LWJ^{+}17, MGT^{+}13, NC02, NP05, SHH^{+}97]. unsymmetric [CS96, CV17, EY94, GD90, YWP11]. untransposed [Mar14]. up- [BHM^{+}95]. update [DT95]. updates [Ree07]. Updating [Ste94, AD14, GLG06, Kon02]. Uphill [DOD23]. upload [CJLS14]. upper [BDK95, TND10]. upper-Hessenberg [BDK95]. URV [Ste94]. USA [Ano22m, GT19]. usage [FL21, IZS^{+}20, PJW^{+}22]. Use [EL95, GSZ88, PW87, BAB^{+}02, DCDV98, GHS02, MK97, SB94a, TKI85]. used [PBTC89]. User [BBB^{+}94, CLZ01, LMJC96, SH91, SKG02, TLC23, vv91]. user-friendliness [TLC23]. user-specified [CLZ01]. users [HJB^{+}21]. uses [Cha99]. Using [ATA^{+}16, CLS^{+}16, CD98, EK98, GHP10, Gro19, HJS23, KEEF01, KT00, LFS^{+}19, MHE19, MBK12, Nag88, ORM^{+}10, PFS^{+}04, SBP12, TDG^{+}18, YTCS97, ZM94, ZCBD22, ATT89, AALK01, ASDOK01, ALD01, ACM20, ABGC^{+}14, ARR19, AED21, ASH92, AIV95, AIIV98, AMMV98, AW13, BSH88, BSE88, BM08, BN01, Bar97, BE87, BCA08, BG21a, BG21b, BCB02, BBQOQO00, BG07, BOS^{+}09, BDK06, BCC^{+}97a, BCC^{+}97b, B^{+}00, BGL^{+}88, BPC21a, BPC21b, BJV^{+}16, BMG07, But92, CZTS99, CLL99, CUSR88, CCS94, Cv98, CS97a, CW06, CDGM96, CKLM14, CWDG07, DG95a, DS13, DFRC94, DWX^{+}12, DM04, DCN^{+}15, DZLK20, DM90, Dow90, EM86, EM88, EM90, EB93, EB94, FGBN19, FKK^{+}06, FHL87, cFM07, GKS07, GMW96, GE11, GRP22, GTD18, GK03, GD90, GPLW17, GMMT16, Gol86, GM07, GM04a, HSR^{+}14, HSS09, HW02]. using [Hen93, HCR01, HM01, HK90b, HWL^{+}22, Hur93, IJM^{+}05, IS18, IB01, IZS^{+}20, IHM^{+}12, JHSV02, JKHK08, JJM^{+}11, KK11, KY98, Kha12, KCN99, KS91b, KK94, KD13, LGC97, LL88, Lan96b, LKHL03, LPNV20, LG01, LPH00, LHZ^{+}20, LR16, LF00, LF88, LNA06, MRBQO14, MSMC15a, MC09, MHH97, MMM13, MSW91, MMP^{+}21, MA22, Nak05, NTHY22, NYFK06, NR20, NU05, OIH10, PSS01, Par87b, PVK^{+}22, PTK04, PVP08, PRW88, PTS^{+}12, Rag97, RvG05, RJ97, Ron84, RBS10, RR16, SYAU07, SNS^{+}97, SHvAP13, SSN^{+}21, SGS95, SSS99, SLH^{+}18, Sil88, SSGF00, SBM^{+}22, SA18, Str08, TTH95, TAB^{+}19, VGS14, Wal01, WCKM11, WAT20, Wel89, WC90, WH91, WH93, XDZL10, XCS05, XLS^{+}17, XZYQ21, YS96, YA24, YEC97, ZLJ93, ZG16]. Utah [LCE^{+}18]. Utility [MKK^{+}19, CL05]. Utility-based [MKK^{+}19, CL05]. utilization [CB17, CT88, Han85, Pea19, SYY^{+}22]. utilize [KLP11]. utilizing [BETR17].

V100 [MBB^{+}18, SWR^{+}18]. VAIL

[ZWJ^{+}19]. validation [BHB21]. valuations [Ger04]. value [Bis89, Gol88, HM89c, HK91, KH92, Kie91, Kie94, KS91a, KN09, KAM^{+}20, Lan99, Lop93, MPJ03, PG93, Pir93, Pir96, TRSC^{+}19]. value-oriented [KAM^{+}20]. valuing [LM00]. variabilities [LMKH97]. variability [JAH^{+}18]. Variable [ADFQO19, AIIV98, BO03, Evr01, KLL^{+}09, LC89, ZGG92]. Variable-size [ADFQO19]. variables [RH12]. variant [EL91, Gen87, Li97, NSS15, Zer90, ZK92]. variants [Mei86]. variation [CGHBS18]. variational [Ran97, TL96]. VARMAX [ÖS94]. varying [BCCS07]. VAX [KW90]. VBARMS [CLS^{+}16]. Vector [Ano87, Ano88b, CC01, DS13, GH89, GMMT16, Hos88, Ort88, Riz85, AB02, Amm89, AMP92, AL88, BE89, BWW89, Bis89, Bjø87, Bro87, Bro88a, CK90, CARW91, CF91, CDC^{+}87, CT88, CMT96, CMM^{+}88a, DEGS95, Dow95, ETV91, FG89, FP98, GR89, GF89, Gen84, GSZ88, GR01, GL97, HH89, HCR01, Her88, Hot89, HB84, IM88, JNWJ18, Kam88, Kan93, Kas84, KD97, Kie91, KD10, KAD09, LB23, LPNJR96, Lec89, Lin91a, LV15, LXL^{+}22, Mar14, MV87a, MA22, ND95, NRN00, Pet91, RRS88, RR89, R^{+}00, Rib84, RZ95, RS87, Sch87, Sch91b, SY87, Sor84, Swa84, VTmL03, VFG12, WP88, WYX^{+}22, Web90, WOV^{+}09, YB11, ZZ03, ZCBD22, ZVWS88, Zla88, van86, van87b]. Vector- [GH89, Her88]. vector-parallel [NRN00, R^{+}00]. Vector-supercomputers [Hos88]. Vectorisation [Gol88, Gol86, Sil88]. vectorizable [EV89, Gen87]. Vectorization [BM90, DCG90, LCE^{+}18, AE86, Buz86, CLC^{+}18, DSSD18, FYH^{+}18, KPS90, LLX^{+}18, LHL18, MBB^{+}18, MCP^{+}14, SL90b, SWR^{+}18, Wak04]. Vectorized [BLFT84, Kra84, dCC89, Lev90]. Vectorizing [DBZ^{+}19, LCD91]. vectors [ASDOK01, LL88]. vehicle [ACGL04, Bis84, LG01, Ral03, Reg01]. vehicles [LG01]. velocity [Lun98]. verification [BDM99, DS92, YLT^{+}23]. versatile [BR99]. versatility [LLS^{+}93]. version [Adr99, ADEQO19, AM93, AT98, CC00a, CCW01, GYL00, RH99, SK97a, SB97, YFK03]. versions [BDOS95, SSS^{+}05, YMYH24]. versus [Ahm97, BT01, BRS^{+}08, But92, Dod90, TLS16]. Vertex [CHZ^{+}19, GPSK09]. Vertex-centric [CHZ^{+}19]. vertical [IT08]. vertices [GS04, Hsi06, PER17]. very [B^{+}01, DHS00, MAH^{+}19, Wel89]. Vese [SPW^{+}15]. VF [CK90, Pet91, Reu88]. via [BMCA98, CB17, CMPM^{+}15, CTW14, CLZ01, DBZ^{+}19, DR03, FLPG18, HRT07, HCL05, LTG23, LB23, MAFC14, SHPA05, TDG^{+}18, ZJDW18]. viability [SKST08]. ViC* [CH97]. Victim [ZWJ^{+}19]. Victim-Aware [ZWJ^{+}19]. Video [AHL02, AALK01, AGB97, Bar12, BC04, CS97b, GLD08, Kut02, MP08a, MP08b, UZ02]. video-on-demand [CS97b]. Vienna [ZBC94]. View [LG03, AC94b, DDdSL02, JLLC22, PCLM18, Sch95, TBM^{+}16b, VRGÁ15, HL13]. View-Oriented [HL13]. Virtual [BCMSW03, Dec00, JBC^{+}07, SAGV21, BGWW97, DCG^{+}07, Ger98, JBWE14, MOF04, TLS^{+}08, YX07]. virtualization [TDG^{+}18]. virus [WCKM11]. viscous [PMLT03]. Visibility [APRP97, AU88, CF93]. vision [KBBC88, PAD02]. vision-application [PAD02]. Visions [CLC08]. visualisation [CJ97, KC97, RB03]. Visualization [CC16, OWW16, Rob99, BCC^{+}97b, BLCR21a, BLCR21b, CSB00, FB19, GWLS05, KMLM97, KZV97, KRS^{+}21, LMC05, MMP^{+}21, OAJ^{+}16, RSB05, SRH07, SHN12, SLS^{+}21, SSV^{+}16, SMW^{+}05, WE97, YM05]. Visualizations [LVC16]. Visualizing [CIG^{+}21, PIG^{+}16, BCCS07, SRK^{+}21]. vital [HWW92, LS97]. Vlasiator [SHvAP13]. Vlasov [DE09, SHvAP13]. VLIW [NS11]. VLSI [AK89b, ASH92, BE93a, Bar95, CS87, CS89c, CS89b, Cho91, DZ90, EG92b, GE92, LV92, MNP93, SM91, Sýk84, Tor91]. VLSI/WSI [Cho91]. VMEC [ROZ01]. VOD [GLP98]. voice [Lin93]. volcanic [OCE^{+}07]. Volta [LLX^{+}18]. voltage [CHQORS18a, SAWH88, TDG^{+}18]. voltage-frequency [CHQORS18a]. Volume [KZV97, Ano84a, Ano01, CYDJ21, CFN03, FB19, FBAB13, GWLS05, HK02a, KC97, Riz85, SL03, SMW^{+}05, SIH14, TW19, WE97, Wit98]. volumes [MSE07]. Voronoi [JDG02, WTLW23]. Voronoĭ [ES89c, Jeo91a, Jeo91b]. Vorton [Kas84]. VP [EHHS89, HL88, TKI85]. VP-100 [TKI85]. VP-100/200 [TKI85]. VP-200 [HL88]. VPP [Heg96]. VPP500 [SE98]. VPP700 [KYLH01]. vs [BBC^{+}11, CHZ^{+}19, Chr98, DLPS92, OIK21, SC19]. vs. [DT95]. VTC [NU05]. VVP [IM88].

wait [Fei91, GWWM09]. wait-for [Fei91]. waiting [GF03]. Walsh [AFS14]. WAN [ALT04, AAB^{+}06]. Wang [Mv88]. ware [Kan97]. warehouse [RBB^{+}22]. warehousing [DMG^{+}04]. warp [FFZ^{+}18, SL03, CQ01]. warping [Wit98]. WarpX [MAA^{+}21]. Warsaw [BKS19]. water [PAF^{+}97, SSL19, YLCT18]. Waterman [Ste13]. watershed [MAB17, MCG98]. wave [AATK10, BCS15, CFD^{+}16, DDF^{+}10, HSS07, KLP11, Tol02]. wave-front [KLP11]. waveform [LG09]. wavefront [EB88, TPK^{+}13]. wavefunction [BCA08]. Wavelet [Uhl96, ALTZ02, BFG^{+}07, Hol95, LSA^{+}95, WC15]. wavenumber [PATC99]. WAVEWATCH [Tol02]. way [CCGG14, EY86, PER17, Zho93]. weak [GSMY^{+}07]. weakly [Ema10]. Weather [Kau94, Sel95, BMS97, BK94, BDI^{+}95, DF95a, DF95b, GJS93a, GJS93b, HJ97b, MJ99, RE98, SXBD97]. Web [MGCB^{+}10, Wal01]. webs [ABG^{+}03]. Weight [GV11, BDK96, N^{+}00, Shy90, da 90]. weighted [BDP20, BSD11, FIMF99, SSB12]. weights [ZG16]. WFBP [GZH^{+}23]. where [Laf02]. Which [BC97, Cha99]. while [GGL99]. whole [AAB^{+}16]. whole-system [AAB^{+}16]. Wide [FGG^{+}98, Con89, FCM03, KBG^{+}01]. Wide-area [FGG^{+}98, KBG^{+}01]. Wilson [GLS99]. wind [SMC16, FMS01]. windows [RGGP^{+}18]. winds [SRK^{+}21]. WinSGL [CWDG07]. wire [Tor91]. wire-routing [Tor91]. within [FMM^{+}02, MHE19, OWW16]. without [BL99, Fei91, HK90b, RSC^{+}19, YG93]. WK [SHCD00]. WK-recursive [SHCD00]. work [BTZ06, CS89a, CGM05, Kau94, MCM01, YMY^{+}19]. work-conserving [BTZ06]. work-stealing [YMY^{+}19]. Worker [CMSL06, BPP10]. Workflow [YBM05, BDNP11, DBI^{+}17, STE23, VK17, WL13b, ZRS13]. workflows [VCK^{+}11]. workload [Kot97, LSL02, LMKH97, ZP16]. workload-aware [ZP16]. workloads [APBcF16, AMC^{+}19, BLR03, LSC^{+}15, MCP^{+}14, RC15]. worksharing [BRRV11]. Workshop [ABH18, DT97, HKK97b, Jou97, KVW97a, SHB19, BVC19]. Workshops [MSS19]. workstation [CS93, DG95a, HP97, NHS^{+}95, NMW93, PM95, SNS^{+}97, SOS97]. workstations [AL93b, ACLN03, AFK01, AS00, CSEK03, HSR^{+}14, HL97, kLH95, LX00, LK00, MHH97, PVP08]. world [CIG^{+}21, DZM^{+}13, Ken90]. Wormhole [LS98, Shi01, ACLN03, CWC00, CCP98, FK98, HM97, KH97, LOKM99, MM96b, Miš98, QCC02, Sha06, SW00, YTCS95, ZL04]. wormhole-based [ACLN03]. wormhole-routed [CWC00, FK98, HM97, MM96b, Miš98]. wormhole-switched [Sha06]. Worst [Liu98, WLZ^{+}23]. Worst-case [Liu98, WLZ^{+}23]. wrap [Hen93]. wraparound [BMT92, VB92]. wrapped [Tou02]. Writing [FAF16]. Wroclaw [BKS19]. WSI [Cho91]. WSRF [BK07]. WSRF-compliant [BK07]. WZ [HE88, Rao97]. WZ-factorisation [HE88].

X [Cal85, CCP^{+}21, DH91, FSY88, Gur88, Hoc85b, HKN89, KN88, Meu87, Nag88, OL86, SW91, Sea86, Tem89]. X-MP [Cal85, DH91, FSY88, Gur88, Hoc85b, HKN89, KN88, Meu87, Nag88, OL86, SW91, Sea86, Tem89]. X-MP/48 [Meu87, Nag88]. X/MP [HL88]. X10 [HXCZ14]. X10-FT [HXCZ14]. XcalableMP [TBM^{+}16b]. Xeon [HSSM17, Tak18, TLS16]. XOR [PGGL17]. XOR-based [PGGL17]. XP [BEK95b, SNS^{+}97]. XP/S [BEK95b].

Y-MP [Oed92, DH91, MSW91, Nag90, Pin91]. Yada [GGNY11]. YARN [SCL^{+}23]. years [Zen99].

zero [FB91, Ver99, XC02]. zero-finding [FB91]. zeros [Fre89]. zerotree [Kut02]. Zimmermann [NSS15]. Zipcode [SSD^{+}94]. Zipfian [BD18]. Zolo [LLD19]. Zolo-SVD [LLD19].

References

Aoki:1999:PLQ

[A^{+}99] S. Aoki et al. Performance of lattice QCD programs on CP-PACS. Parallel Computing, 25(10– 11):1243–1255, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/26/20/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/26/20/
article.pdf.

Antoniu:2001:HSC

[A^{+}01a] Gabriel Antoniu et al. The Hyperion system: Compiling multithreaded Java bytecode for distributed execution. Parallel Computing, 27(10):1279–1297, September 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/40/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/40/27/
article.pdf.

Aumann:2001:MPC

[A^{+}01b] P. Aumann et al. MEGAFLOW: Parallel complete aircraft CFD. Parallel Computing, 27(4):415–440, March 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/28/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/28/27/
article.pdf.

Allcock:2002:DMT

[A^{+}02] Bill Allcock et al. Data management and transfer in high-performance computational grid environments. Parallel Computing, 28(5):749–771, May 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/57/31/
abstract.html.

Acosta:2016:EPO

[AAA16] Alejandro Acosta, Sergio Afonso, and Francisco Almeida. Extending Paralldroid with object oriented annotations. Parallel Computing, 57(??):25–36, September 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300126.

Alba:2006:EPL

[AAB^{+}06] E. Alba, F. Almeida, M. Blesa, C. Cotta, M. Díaz, I. Dorta, J. Gabarró, C. León, G. Luque, J. Petit, C. Rodríguez, A. Rojas, and F. Xhafa. Efficient parallel LAN/WAN algorithms for optimization. the MALLBA Project. Parallel Computing, 32 (5–6):415–440, June 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Agelastos:2016:CWS

[AAB^{+}16] Anthony Agelastos, Benjamin Allan, Jim Brandt, Ann Gentile, Sophia Lefantzi, Steve Monk, Jeff Ogden, Mahesh Rajan, and Joel Stevenson. Continuous whole-system monitoring toward rapid understanding of production HPC applications and systems. Parallel Computing, 58(??):90–106, October 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300394.

Agullo:2018:SIP

[AAGS18] Emmanuel Agullo, Peter Arbenz, Luc Giraud, and Olaf Schenk. Special issue on parallel matrix algorithms and applications (PMAA’16). Parallel Computing, 74(??):1– 2, ???? 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300115.

Amer:1997:PIK

[AAHF97] M. A. Amer, B. A. Abdel-Hamida, and D. Fausett. Parallel implementation of the Kronecker product technique for numerical solution of parabolic partial differential equations. Parallel Computing, 22(14):1997–2005, March 24, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/

cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=14&aid=1140.

Ahmad:2001:SLM

[AALK01] Ishfaq Ahmad, Shahriar M.

Akramullah, Ming L. Liou, and Muhammad Kafil. A scalable off-line MPEG-2 video encoding scheme using a multiprocessor system. Parallel Computing, 27(6):823–846, May 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/30/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/30/27/
article.pdf.

Albayrak:2013:IAB

[AAO13] Omer Erdil Albayrak, Ismail Akturk, and Ozcan Ozturk. Improving application behavior on heterogeneous manycore systems through kernel mapping. Parallel Computing, 39(12):867–878, December 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001038.

Andion:2013:NCS

[AART13] José M. Andión, Manuel Arenaz, Gabriel Rodríguez, and Juan Touriño. A novel compiler support for automatic parallelization on multicore systems. Parallel Computing, 39(9):442–460, September 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000471.

Agyeman:2013:ERT

[AAS13] Michael Opoku Agyeman, Ali Ahmadinia, and Alireza Shahrabi. Efficient routing techniques in heterogeneous 3D Networks-on-Chip. Parallel Computing, 39(9):389– 407, September 2013. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819113000537.

Almasri:2020:CES

[AAS20] Mohammad Almasri and Walid Abu-Sufah. CCF: an efficient SpMV storage format for AVX512 platforms. Parallel Computing, 100(??):Article 102710, December 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300958.

Athanaileas:2010:PRW

[AATK10] T. E. Athanaileas, G. E.

Athanasiadou, G. V. Tsoulos, and D. I. Kaklamani. Parallel radiowave propagation modeling with image-based ray tracing techniques. Parallel Computing, 36(12): 679–695, December 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Alam:2008:PCB

[AAV08] Sadaf R. Alam, Pratul K.

Agarwal, and Jeffrey S. Vetter. Performance characteristics of biomolecular simulations on highend systems with multicore processors. Parallel Computing, 34(11):640– 651, November 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Arguello:1996:FMC

[AAZ96] Francisco Argüello, Margarita Amor, and Emilio L. Zapata. FFTs on mesh connected computers. Parallel Computing, 22(1):19– 38, February 20, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cgi-bin/cas/tree/
store/parco/cas_sub/
browse/browse.cgi?year=

1996&volume=22&issue=
1&aid=1041.

Atiquzzaman:1993:EHS

[AB93] M. Atiquzzaman and M. M. Banat. Effect of hot-spots on the performance of crossbar multiprocessor systems. Parallel Computing, 19(4):455– 461, April 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Amodio:1995:PFA

[AB95] Pierluigi Amodio and Luigi Brugnano. The parallel QR factorization algorithm for tridiagonal linear systems. Parallel Computing, 21(7):1097–1110, July 11, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=7&aid=990.

Abbas:2002:PCD

[AB02] Hazem M. Abbas and Mohamed M. Bayoumi. Parallel codebook design for vector quantization on a message passing MIMD architecture. Parallel Computing, 28(7–8):1079–1093, August 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/58/34/
abstract.html.

Azad:2016:MAF

[AB16] Ariful Azad and Aydin Buluç. A matrixalgebraic formulation of distributed-memory maximal cardinality matching algorithms in bipartite graphs. Parallel Computing, 58(??):117–130, October 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300370.

Auckenthaler:2011:PSP

[ABB^{+}11] T. Auckenthaler, V. Blum, H.-J. Bungartz, T. Huckle, R. Johanni, L. Krämer, B. Lang, H. Lederer, and P. R. Willems. Parallel solution of partial symmetric eigenvalue problems from electronic structure calculations. Parallel Computing, 37(12):783–794, December 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000494.

Aliaga:2016:ETD

[ABB^{+}16] José I. Aliaga, Rosa M.

Badia, Maria Barreda, Matthias Bollhöfer, Ernesto Dufrechou, Pablo Ezzatti, and Enrique S. Quintana-Ortí. Exploiting task and data parallelism in ILU-PACK’s preconditioned CG solver on NUMA architectures and manycore accelerators. Parallel Computing, 54(??):97– 107, May 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001581.

Abdelfattah:2021:GAE

[ABB^{+}21] Ahmad Abdelfattah, Valeria Barra, Natalie Beams, Ryan Bleile, Jed Brown, Jean-Sylvain Camier, Robert Carson, Noel Chalmers, Veselin Dobrev, Yohann Dudouit, Paul Fischer, Ali Karakus, Stefan Kerkemeier, Tzanio Kolev, Yu-Hsiang Lan, Elia Merzari, Misun Min, Malachi Phillips, Thilina Rathnayake, Robert Rieben, Thomas Stitt, Ananias Tomboulides, Stanimire Tomov, Vladimir Tomov, Arturo Vargas, Tim Warburton, and Kenneth Weiss. GPU algorithms for efficient exascale discretizations. Parallel Computing, 108(??):??, December 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819121000879.

Ananthanarayan:2003:DWE

[ABG^{+}03] Asvin Ananthanarayan, Rajiv Balachandran, Robert Grossman, Yunhong Gu, Xinwei Hong, Jorge Levera, and Marco Mazzucco. Data webs for earth science data. Parallel Computing, 29(10):1363–1379, October 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Arbenz:2006:PMP

[ABG^{+}06] Peter Arbenz, Martin Bečka, Roman Geus, Ulrich Hetmaniuk, and Tiziano Mengotti. On a parallel multilevel preconditioned Maxwell eigensolver. Parallel Computing, 32(2):157–165, February 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Arroyuelo:2014:DTS

[ABGC^{+}14] Diego Arroyuelo, Carolina Bonacic, Veronica Gil-Costa, Mauricio Marin, and Gonzalo Navarro. Distributed text search using suffix arrays. Parallel Computing, 40(9):471–495, October 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000805.

Acer:2021:SPM

[ABGR21] Seher Acer, Erik G. Boman, Christian A. Glusa, and Sivasankaran Rajamanickam. Sphynx: a parallel multi-GPU graph partitioner for distributedmemory systems. Parallel Computing, 106(??):??, September 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000272.

Auckenthaler:2010:MEP

[ABH^{+}10] T. Auckenthaler, M. Bader, T. Huckle, A. Spörl, and K. Waldherr. Matrix exponentials and parallel prefix computation in a quantum control problem. Parallel Computing, 36(5–6):359– 369, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Amer:2018:IWP

[ABH18] Abdelhalim Amer, Pavan Balaji, and Zhiyi Huang. 8th International Workshop on Programming Models and Applications for Multicores and Manycores (PMAM’17). Parallel Computing, 78(??):84, October 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301145.

Aumage:2002:MIP

[ABMN02] Olivier Aumage, Luc Bougé, Jean-François Méhaut, and Raymond Namyst. Madeleine II: a portable and efficient communication library for highperformance cluster computing. Parallel Computing, 28(4):607–626, April 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/39/29/
abstract.html.

Aliaga:2011:ETL

[ABMQO11] José I. Aliaga, Matthias Bollhöfer, Alberto F. Martín, and Enrique S. Quintana-Orti’. Exploiting thread-level parallelism in the iterative solution of sparse linear systems. Parallel Computing, 37(3):183–202, March 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819110001432.

Amoura:1999:CHS

[ABMT99] A. K. Amoura, E. Bampis, Y. Manoussakis, and Zs. Tuza. A comparison of heuristics for scheduling multiprocessor tasks on three dedicated processors. Parallel Computing, 25(1):49–61, January 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1999/
                                        25/1/1378.pdf.

Aiex:2003:PGP

[ABR03] R. M. Aiex, S. Binato, and M. G. C. Resende. Parallel GRASP with path-relinking for job shop scheduling. Parallel Computing, 29(4):393–430, April 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Allwright:1989:DIS

[AC89] James R. A. Allwright and D. B. Carpenter. A distributed implementation of simulated annealing for the travelling salesman problem. Parallel Computing, 10(3):335–338, May 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Angelaccio:1994:RCP

[AC94a] M. Angelaccio and M. Colajanni. The row/ column pivoting strategy on multicomputers. Parallel Computing, 20(2): 197–213, February 24, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=2&aid=821.

Angelaccio:1994:SMD

[AC94b] M. Angelaccio and M. Colajanni. Subcube matrix decomposition: a unifying view for LU factorization on multicomputers. Parallel Computing, 20(2):257–270, February 24, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=2&aid=825.

Andreoni:2000:NAC

[AC00] Wanda Andreoni and Alessandro Curioni. New advances in chemistry and materials science with CPMD and parallel computing. Parallel Computing, 26(7–8):819–842, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/42/29/24/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/29/24/
                                        article.pdf.

Aloisio:2003:DEO

[AC03] Giovanni Aloisio and Massimo Cafaro. A dynamic earth observation system. Parallel Computing, 29(10):1357–1362, October 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Askew:1988:MCS

[ACC^{+}88] C. R. Askew, D. B. Carpenter, J. T. Chalker, A. J. G. Hey, M. Moore, D. A. Nicole, and D. J. Pritchard. Monte Carlo simulation on transputer arrays. Parallel Computing, 6(2):247–258, February 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Attanasio:2004:PTS

[ACGL04] Andrea Attanasio, Jean-François Cordeau, Gianpaolo Ghiani, and Gilbert Laporte. Parallel Tabu search heuristics for the dynamic multi-vehicle diala-ride problem. Parallel Computing, 30(3):377–387, March 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Arbab:1998:CLP

[ACH98] F. Arbab, P. Ciancarini, and C. Hankin. Coordination languages for parallel programming. Parallel Computing, 24 (7):989–1004, July 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/7/1316.pdf.

Alonso:2018:TSO

[ACH^{+}18] Pedro Alonso, Sandra Catalán, José R. Herrero, Enrique S. Quintana-Ortí, and Rafael Rodríguez-Sánchez. Two-sided orthogonal reductions to condensed forms on asymmetric multicore processors. Parallel Computing, 78(??):85–100, October 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300784.

Anglano:2003:NIM

[ACLN03] Cosimo Anglano, Claudio Casetti, Emilio Leonardi, and Fabio Neri. Network interface multicast protocols for wormhole-based networks of workstations. Parallel Computing, 29(2): 255–283, February 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Alonso:2010:PSR

[ACM^{+}10] Marina Alonso, Salvador Coll, Juan-Miguel Martínez, Vicente Santonja, Pedro López, and José Duato. Power saving in regular interconnection networks. Parallel Computing, 36 (12):696–712, December 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Alonso:2015:PCM

[ACM^{+}15] M. Alonso, S. Coll, J. M. Martínez, V. Santonja, and P. López. Power consumption management in fat-tree interconnection networks. Parallel Computing, 48(??):59–80, October 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000599.

Araujo:2020:MIL

[ACM20] Rodolfo Pereira Araujo, Igor Machado Coelho, and Leandro Augusto Justen Marzulo. A multiimprovement local search using dataflow and GPU to solve the minimum latency problem. Parallel Computing, 98(??):Article 102661, October 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300545.

Aldinucci:2020:DSP

[ACMT20] Marco Aldinucci, Valeria Cardellini, Gabriele Mencagli, and Massimo Torquati. Data stream processing in HPC systems: New frameworks and architectures for highfrequency streaming. Parallel Computing, 98(??): Article 102694, October 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300843.

Ahmad:1996:MSG

[AD96] Imtiaz Ahmad and Muhammad K. Dhodhi. Multiprocessor scheduling in a genetic paradigm. Parallel Computing, 22(3):395–406, April 29, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/

                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=3&aid=1047.

Andrew:2014:IFU

[AD14] Robert Andrew and Nicholas Dingle. Implementing QR factorization updating algorithms on GPUs. Parallel Computing, 40(7):161– 172, July 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000337.

Axelrod:1984:SMP

[ADE84] T. Axelrod, P. Dubois, and P. Eltgroth. A simulator for MIMD performance prediction: application to the S-1 MkIIa multiprocessor. Parallel Computing, 1(3–4):237–274, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Aliaga:2019:ATD

[ADEQO19] José I. Aliaga, Ernesto Dufrechou, Pablo Ezzatti, and Enrique S. Quintana-Ortí. Accelerating the task/data-parallel version of ILUPACK’s BiCG in multi-CPU/GPU configurations. Parallel Computing, 85(??):79–87, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301777.

Aggarwal:1993:PSF

[ADF93] Rajesh Aggarwal, David R.

Dellwo, and Morton B. Friedman. Parallel solution of Fredholm integral equations of the second kind by accelerated projection methods. Parallel Computing, 19(10): 1105–1115, October 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anzt:2019:VSB

[ADFQO19] Hartwig Anzt, Jack Dongarra, Goran Flegar, and Enrique S. Quintana-Ortí. Variable-size batched Gauss–Jordan elimination for block-Jacobi preconditioning on graphics processors. Parallel Computing, 81(??):131–146, January 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117302107.

Amestoy:2010:ASP

[ADGS10] P. Amestoy, I. S. Duff, A. Guermouche, and Tz. Slavova. Analysis of the solution phase of a parallel multifrontal approach.

Parallel Computing, 36(1): 3–15, January 2010. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Agathos:2022:CAA

[ADK22] Spiros N. Agathos, Vassilios V. Dimakopoulos, and Ilias K. Kasmeridis. Compiler-assisted, adaptive runtime system for the support of OpenMP in embedded multicores. Parallel Computing, 110(??):??, May 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819122000035.

Amestoy:2003:IIM

[ADLL03] Patrick R. Amestoy, Iain S. Duff, Jean-Yves L’Excellent, and Xiaoye S. Li. Impact of the implementation of MPI point-to-point communications on the performance of two general sparse solvers. Parallel Computing, 29(7):833–849, July 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Aversa:2005:HDS

[ADMV05] Rocco Aversa, Beniamino Di Martino, Nicola Mazzocca, and Salvatore Venticinque. A hierarchical distributed-shared memory parallel Branch & Bound application with PVM and OpenMP for multiprocessor clusters. Parallel Computing, 31 (10–12):1034–1047, October/December 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Amestoy:2003:APS

[ADPV03] Patrick R. Amestoy, Iain S.

Duff, Stéphane Pralet, and Christof Vömel. Adapting a parallel sparse direct solver to architectures with clusters of SMPs. Parallel Computing, 29(11– 12):1645–1668, November/ December 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Adrian:1999:PPR

[Adr99] Gerhard Adrian. Parallel processing in regional climatology: The parallel version of the “Karlsruhe Atmospheric Mesoscale Model’’ (KAMM). Parallel Computing, 25(7):777–787, August 13, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1999&volume=
25&issue=7&aid=1409;

                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/32/
                                        23/17/abstract.html;
                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/32/
                                        23/17/article.pdf.

Aversa:2005:PPT

[ADR^{+}05] Rocco Aversa, Beniamino Di Martino, Massimiliano Rak, Salvatore Venticinque, and Umberto Villano. Performance prediction through simulation of a hybrid MPI/ OpenMP application. Parallel Computing, 31(10– 12):1013–1033, October/ December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Arenaz:2001:EPN

[ADTV01] M. Arenaz, R. Doallo, J. Touriño, and C. Vázquez. Efficient parallel numerical solver for the elastohydrodynamic Reynolds-Hertz problem. Parallel Computing, 27(13):1743–1765, December 1, 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/47/43/31/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/43/31/
                                        article.pdf.

Axelsson:1986:NVS

[AE86] O. Axelsson and V. Eijkhout. A note on the vectorization of scalar recursions. Parallel Computing, 3(1):73–83, March 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Azaria:1993:MEN

[AE93] H. Azaria and Y. Elovici.

Modeling and evaluation of a new message-passing system for parallel multiprocessor systems. Parallel Computing, 19(6):633–649, June 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ateskan:2021:PNM

[AED21] Esra Ruzgar Ateskan, Kayhan Erciyes, and Mehmet Emin Dalkilic. Parallelization of network motif discovery using star contraction. Parallel Computing, 101(??): Article 102734, April 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120301174.

Andrade:2013:APB

[AFD13] Diego Andrade, Basilio B.

Fraguela, and Ramón Doallo. Accurate prediction of the behavior of multithreaded applications in shared caches. Parallel Computing, 39(1):36– 57, January 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000919.

Ashworth:1997:PGM

[AFG^{+}97] M. Ashworth, F. Foelkel, V. Gülzow, K. Kleese, D. P. Eppel, H. Kapitza, and S. Unger. Parallelization of the GESIMA mesoscale atmospheric model. Parallel Computing, 23(14):2201–2213, December 17, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=14&aid=1267.

Awan:2006:UPP

[AFJG06] Asad Awan, Ronaldo A. Ferreira, Suresh Jagannathan, and Ananth Grama. Unstructured peer-to-peer networks for sharing processor cycles. Parallel Computing, 32(2):115–135, February 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Aritsugi:2001:SPS

[AFK01] Masayoshi Aritsugi, Hiroki Fukatsu, and Yoshinari Kanamori. Several partitioning strategies for parallel image convolution in a network of heterogeneous workstations. Parallel Computing, 27(3):269–293, February 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/22/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/22/26/
article.pdf.

Aktulga:2012:PRM

[AFPG12] H. M. Aktulga, J. C. Fogarty, S. A. Pandit, and A. Y. Grama. Parallel reactive molecular dynamics: Numerical methods and algorithmic techniques. Parallel Computing, 38(4–5):245–259, April/May 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001074.

Andrade:2014:OFW

[AFS14] Joao Andrade, Gabriel Falcao, and Vitor Silva. Optimized Fast Walsh– Hadamard Transform on GPUs for non-binary LDPC decoding. Parallel Computing, 40(9):449–453, October 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000817.

Andrade:2023:SAD

[AFT23] Guilherme Andrade, Renato Ferreira, and George Teodoro. Spatial-aware data partition for distributed memory parallelization of ANN search in multimedia retrieval. Parallel Computing, 115(??): ??, February 2023. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819122000813.

Ausavarungnirun:2016:CHR

[AFY^{+}16] Rachata Ausavarungnirun, Chris Fallin, Xiangyao Yu, Kevin Kai-Wei Chang, Greg Nazario, Reetuparna Das, Gabriel H. Loh, and Onur Mutlu. A case for hierarchical rings with deflection routing: an energy-efficient onchip communication substrate. Parallel Computing, 54(??):29–45, May 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000399.

Ayed:2000:EHC

[AG00] Moez Ayed and Jean-Luc Gaudiot. An efficient heuristic for code partitioning. Parallel Computing, 26(4):399–426, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/26/24/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/26/24/
article.pdf.

Arapis:1997:RTS

[AGB97] Constantin Arapis, Simon Gibbs, and Christian Breiteneder. Real-time segmentation of video on a multiprocessor platform. Parallel Computing, 23 (12):1777–1792, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=12&aid=1245.

Anzt:2017:PKS

[AGD^{+}17] Hartwig Anzt, Mark Gates, Jack Dongarra, Moritz Kreutzer, Gerhard Wellein, and Martin Köhler. Preconditioned Krylov solvers on GPUs. Parallel Computing, 68(??):32–44, October 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117300777.

Attanasio:2006:AAD

[AGGG06] Andrea Attanasio, Gianpaolo Ghiani, Lucio Grandinetti, and Francesca Guerriero. Auction algorithms for decentralized parallel machine scheduling. Parallel Computing, 32(9):701–709, October 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Averbuch:1990:PFM

[AGGM90] Amir Averbuch, Eran Gabber, Boaz Gordissky, and Yoav Medan. A parallel FFT on an MIMD machine. Parallel Computing, 15(1–3):61–74, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Addison:1994:BDM

[AGH^{+}94] C. A. Addison, V. S. Getov, A. J. G. Hey, R. W. Hockney, and I. C. Wolton. Benchmarking for distributed memory parallel systems: Gaining insight from numbers. Parallel Computing, 20(10– 11):1653–1668, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=920.

Ayguade:1998:TTA

[AGK98] Eduard Ayguadé, Jordi Garcia, and Ulrich Kremer. Tools and techniques for automatic data layout: a case study. Parallel Computing, 24(3–4):557– 578, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/3-4/1302.pdf.

Arbenz:2015:SIP

[AGKS15] Peter Arbenz, Laura Grigori, Rolf Krause, and Olaf Schenk. Special issue on Parallel Matrix Algorithms and Applications (PMAA’14). Parallel Computing, 49(??):99–100, November 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001234.

Arbenz:2016:SIP

[AGKS16] Peter Arbenz, Laura Grigori, Rolf Krause, and Olaf Schenk. Special issue on Parallel Matrix Algorithms and Applications (PMAA’14). Parallel Computing, 57 (??):135–136, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300709.

Agullo:2008:PCM

[AGL08] Emmanuel Agullo, Abdou Guermouche, and Jean-Yves L’Excellent. A parallel out-of-core multifrontal method: Storage of factors on disk and analysis of models for an out-of-core active memory. Parallel Computing, 34(6–8): 296–317, July 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Amestoy:2006:HSP

[AGLP06] Patrick R. Amestoy, Abdou Guermouche, Jean-Yves L’Excellent, and Stéphane Pralet. Hybrid scheduling for the parallel solution of linear systems. Parallel Computing, 32(2): 136–156, February 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Arbenz:1997:RCS

[AGO97] P. Arbenz, W. Gander, and M. Oettli. The remote computation system. Parallel Computing, 23 (10):1421–1428, November 7, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=10&aid=1219.

Arbenz:2003:PMA

[AGPS03] Peter Arbenz, Efstratios Gallopoulos, Bernard Philippe, and Yousef Saad. Parallel Matrix Algorithms and Applications (PMAA ’02). Parallel Computing, 29(9):1117–1119, September 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ababao:2017:SCC

[AGV^{+}17] Rainier Ababao, Joe A.

Garcia, Joseph Voss, W. Cyrus Proctor, and R. Todd Evans. Student Cluster Competition 2016 reproducibility challenge: Genomic partitioning with ParConnect. Parallel Computing, 70(??):5–10, December 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300947.

Anzt:2018:ISA

[AHBD18] Hartwig Anzt, Thomas K. Huckle, Jürgen Bräckle, and Jack Dongarra. Incomplete sparse approximate inverses for parallel preconditioning. Parallel Computing, 71(??):1– 22, January 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911730176X.

Arguello:2012:SMM

[AHBLR12] F. Argüello, D. B. Heras, M. Bóo, and J. Lamas-Rodríguez. The splitand-merge method in general purpose computation on GPUs. Parallel Computing, 38(6–7):277–288, June/July 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000208.

Ahmad:2002:VCP

[AHL02] Ishfaq Ahmad, Yong He, and Ming L. Liou. Video compression with parallel processing. Parallel Computing, 28(7–8):1039–1078, August 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/58/33/
abstract.html.

Ahmad:1997:EVP

[Ahm97]  Ishfaq Ahmad.  Ex-

press versus PVM: a performance comparison. Parallel Computing, 23 (6):783–812, June 20, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=6&aid=1138.

Auckenthaler:2014:BDP

[AHW14] T. Auckenthaler, T. Huckle, and R. Wittmann. A blocked QR-decomposition for the parallel symmetric eigenvalue problem. Parallel Computing, 40(7):186– 194, July 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000404.

Averbuch:1998:TDP

[AIIV98] A. Averbuch, L. Ioffe, M. Israeli, and L. Vozovoi. Two-dimensional parallel solver for the solution of Navier–Stokes equations with constant and variable coefficients using ADI on cells. Parallel Computing, 24(5–6):673–699, June 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/5-6/1310.pdf.

Altman:1995:HRC

[AIO95] Tom Altman, Yoshihide Igarashi, and Koji Obokata. Hyper-ring connection machines. Parallel Computing, 21(8):1327–1338, August 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=8&aid=1001.

Averbuch:1995:PIN

[AIV95] A. Averbuch, M. Israeli, and L. Vozovoi. Parallel implementation of non-linear evolution problems using parabolic domain decomposition. Parallel Computing, 21 (7):1151–1183, July 11, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/

cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=7&aid=991.

Arenstorf:1989:CBA

[AJ89] N. S. Arenstorf and H. F.

Jordan. Comparing barrier algorithms. Parallel Computing, 12(2):157–170, November 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Agui:1995:BTI

[AJ95] Juan C. Agüí and Javier Jiménez. A binary tree implementation of a parallel distributed tridiagonal solver. Parallel Computing, 21(2):233–241, February 17, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=2&aid=940.

Aarts:1989:CMP

[AK89a] Emile H. L. Aarts and Jan H. M. Korst. Computations in massively parallel networks based on the Boltzmann machine: a review. Parallel Computing, 9(2):129–145, January 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). PARLE: Conference on Parallel Architectures and Languages— Europe (Eindhoven, 1987).

Alnuweiri:1989:EVA

[AK89b] Hussein M. Alnuweiri and V. K. Prasanna Kumar. An efficient VLSI architecture with applications to geometric problems. Parallel Computing, 12(1):71– 93, October 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Adam:2019:CRA

[AKB^{+}19] Julien Adam, Maxime Kermarquer, Jean-Baptiste Besnard, Leonardo Bautista-Gomez, Marc Pérache, Patrick Carribault, Julien Jaeger, Allen D. Malony, and Sameer Shende. Checkpoint/restart approaches for a threadbased MPI runtime. Parallel Computing, 85(??):204– 219, July 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118303247.

Annamalai:1999:AFE

[AKK99] V. Annamalai, C. S. Krishnamoorthy, and V. Kamakoti. Adaptive finite element analysis on a parallel and distributed environment. Parallel Computing, 25(12):1413–1434, November 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/35/23/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/35/23/
article.pdf.

Farhan:2016:UCA

[AKK16] Mohammed A. Al Farhan, Dinesh K. Kaushik, and David E. Keyes. Unstructured computational aerodynamics on many integrated core architecture. Parallel Computing, 59(??):97–118, November 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300564.

Arvindam:1991:ATP

[AKNS91] S. Arvindam, V. Kumar, V. Nageshwara Rao, and V. Singh. Automatic test pattern generation on parallel processors. Parallel Computing, 17(12):1323– 1342, December 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Armstrong:1998:CBT

[AKP^{+}98] Brian Armstrong, Seon Wook Kim, Insung Park, Michael Voss, and Rudolf Eigenmann. Compiler-based tools for analyzing parallel programs. Parallel Computing, 24(3–4):401– 420, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/3-4/1296.pdf.

Andrade:2007:ASC

[AKSS07] Henrique Andrade, Tahsin Kurc, Alan Sussman, and Joel Saltz. Active semantic caching to optimize multidimensional data analysis in parallel and distributed environments. Parallel Computing, 33(7–8):497– 520, August 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ashworth:1988:SFA

[AL88] Mike Ashworth and Andrew G. Lyne. A segmented FFT algorithm for vector computers. Parallel Computing, 6(2):217–224, February 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Aboelaze:1993:MDA

[AL93a] Mokhtar Aboelaze and De-Lei L. Lee. A method for data allocation and manipulation in hypercube computers. Parallel Computing, 19(5):497–510, May 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Altevogt:1993:PTD

[AL93b] P. Altevogt and A. Linke.

Parallelization of the twodimensional Ising model on a cluster of IBM RISC System/6000 workstations. Parallel Computing, 19 (9):1041–1052, September 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Alaghband:1989:PPC

[Ala89] Gita Alaghband. Parallel pivoting combined with parallel reduction and fill-in control. Parallel Computing, 11(2):201–221, ???? 4, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Alaghband:1995:PSM

[Ala95] Gita Alaghband. Parallel sparse matrix solution and performance. Parallel Computing, 21 (9):1407–1430, September 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.

cgi?year=1995&volume=
21&issue=9&aid=1008.

Alabdulkareem:2001:SAL

[ALD01] M. Alabdulkareem, S. Lakshmivarahan, and S. K. Dhall. Scalability analysis of large codes using factorial designs. Parallel Computing, 27(9):1145–1171, August 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/35/23/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/35/23/
article.pdf.

Alef:1991:CEM

[Ale91] M. Alef. Concepts for efficient multigrid implementation on SUPRENUMlike architectures. Parallel Computing, 17(1):1–16, April 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Alef:1994:IMA

[Ale94] Manfred Alef. Implementation of a multigrid algorithm on SUPRENUM and other systems. Parallel Computing, 20(10– 11):1547–1557, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=912.

Aktulga:2014:PEC

[ALH^{+}14] Hasan Metin Aktulga, Lin Lin, Christopher Haine, Esmond G. Ng, and Chao Yang. Parallel eigenvalue calculation based on multiple shift-invert Lanczos and contour integral based spectral projection method. Parallel Computing, 40(7):195–212, July 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000325.

Almasi:1985:OPP

[Alm85] G. S. Almasi. Overview of parallel processing. Parallel Computing, 2(3):191– 203, November 1985. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Alba:2004:PHG

[ALNT04] E. Alba, F. Luna, A. J.

Nebro, and J. M. Troya. Parallel heterogeneous genetic algorithms for continuous optimization. Parallel Computing, 30(5–6): 699–719, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Alsuwaiyel:2001:OPA

[Als01] Muhammad H. Alsuwaiyel. An optimal parallel algorithm for the multiselection problem. Parallel Computing, 27(6):861– 865, May 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/30/29/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/30/29/
                                        article.pdf.

Alba:2004:PLW

[ALT04] E. Alba, G. Luque, and J. M. Troya. Parallel LAN/WAN heuristics for optimization. Parallel Computing, 30(5–6): 611–628, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Arguello:2002:AWP

[ALTZ02] Francisco Argüello, Juan López, María A. Trenas, and Emilio L. Zapata. Architecture for wavelet packet transform based on lifting steps. Parallel Computing, 28(7–8):1023–1037, August 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/58/32/
abstract.html.

Amodio:1993:PVC

[AM93] P. Amodio and N. Mastronardi. A parallel version of the cyclic reduction algorithm on a hypercube. Parallel Computing, 19(11):1273–1281, November 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Awan:2019:OLM

[AMC^{+}19] Ammar Ahmad Awan, Karthik Vadambacheri Manian, Ching-Hsiang Chu, Hari Subramoni, and Dhabaleswar K. Panda. Optimized large-message broadcast for deep learning workloads: MPI, MPI + NCCL, or NCCL2? Parallel Computing, 85(??):141– 152, July 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303284.

Amman:1989:NCS

[Amm89] H. M. Amman. Nonlinear control simulation on a vector machine. Parallel Computing, 10(1):123– 127, March 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Aversa:1998:DAH

[AMMV98] R. Aversa, A. Mazzeo, N. Mazzocca, and U. Villano. Developing applications for heterogeneous computing environments using simulation: a case study. Parallel Computing, 24(5–6):741–761, June 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/5-6/1308.pdf.

Abhinav:2016:ESI

[AMN16] Vishnu Abhinav, Andres Marquez, and Dimitris Nikolopoulos. Editorial of the special issue: SI: E2SC. Parallel Computing, 57(??):107, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300680.

Andersen:1992:SSM

[AMP92] J. Andersen, G. Mitra, and D. Parkinson. The scheduling of sparse matrix-vector multiplication on a massively parallel DAP computer. Parallel Computing, 18(6):675–697, June 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ahrabian:2005:PGB

[AND05] Hayedeh Ahrabian and Abbas Nowzari-Dalini. Parallel generation of binary trees in A-order. Parallel Computing, 31 (8–9):948–955, August/ September 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Amaral:2020:PLD

[ANG^{+}20] Vasco Amaral, Beatriz Norberto, Miguel Goulão, Marco Aldinucci, Siegfried Benkner, Andrea Bracciali, Paulo Carreira, Edgars Celms, Luís Correia, Clemens Grelck, Helen Karatza, Christoph Kessler, Peter Kilpatrick, Hugo Martiniano, Ilias Mavridis, Sabri Pllana, Ana Respício, José Simão, Luís Veiga, and Ari Visa. Programming languages for data-intensive HPC applications: a systematic mapping study. Parallel Computing, 91(??): Article 102584, March 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301759.

Annot:1989:DFS

[Ann89] J. K. Annot. A deadlock free and starvation free network of packet switching communication processors. Parallel Computing, 9 (2):147–162, January 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:1984:AIV

[Ano84a] Anonymous. Author index to volume 1 (1984). Parallel Computing, 1 (3–4):345–346, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819184903570.

Anonymous:1984:C

[Ano84b] Anonymous. Calendar. Parallel Computing, 1 (3–4):343–344, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819184903417.

Anonymous:1987:ICV

[Ano87] Anonymous. International Conference on Vector and Parallel Computing — Issues in Applied Research and Development. Parallel Computing, 5(1–2):??, July 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:1988:ISC

[Ano88a] Anonymous. 2nd International SUPRENUM Colloquium. Parallel Computing, 7(3):??, September 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:1988:ICV

[Ano88b] Anonymous. International conference on vector and parallel processors in computational science III. Parallel Computing, 8(1–3):??, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:1994:PC

[Ano94]  Anonymous.  Parallel

Computing 93. Parallel Computing, 20(3):409–??, March 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:1997:MCF

[Ano97]  Anonymous.  Miscella-

neous: Calendar of forthcoming conferences and events. Parallel Computing, 23(6):813–??, June 20, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:1999:I

[Ano99a] Anonymous. Index. Parallel Computing, 25(13– 14):2189–2196, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/36/49/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/49/
article.pdf.

Anonymous:1999:PTN

[Ano99b] Anonymous. Parallelization techniques for numerical modelling. Parallel Computing, 25(7):775–776, August 13, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/23/35/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/23/35/
article.pdf.

Anonymous:2000:I

[Ano00] Anonymous. Index. Parallel Computing, 26(13– 14):1955–1962, December 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/34/35/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/34/35/
article.pdf.

Anonymous:2001:AIV

[Ano01] Anonymous. Author index to volume 27. Parallel Computing, 27(14): 1937–1944, December 31, 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/44/32/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/44/32/
article.pdf.

Anonymous:2002:AI

[Ano02a] Anonymous. Author index. Parallel Computing, 28(12):1833–1839, December 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2002:EBa

[Ano02b]  Anonymous.  Editorial

Board. Parallel Computing, 28(9):CO2, September ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/59/27/
abstract.html.

Anonymous:2002:EBb

[Ano02c]  Anonymous.  Editorial

Board. Parallel Computing, 28(10):CO2, October ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/60/27/
                                        abstract.html.

Anonymous:2002:EBc

[Ano02d]  Anonymous.  Editorial

Board. Parallel Computing, 28(11):CO2, November ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/61/25/
                                        abstract.html.

Anonymous:2002:EBd

[Ano02e]  Anonymous.  Editorial

Board. Parallel Computing, 28(12):CO2, December 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2002:IIF

[Ano02f] Anonymous. IFC — inside front cover (Editorial Board). Parallel Computing, 28(7–8):CO2, August 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/58/27/
                                        abstract.html.

Anonymous:2003:EBa

[Ano03a]  Anonymous.  Editorial

Board. Parallel Computing, 29(1):CO2, January 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2003:EBb

[Ano03b]  Anonymous.  Editorial

Board. Parallel Computing, 29(2):CO2, February 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2003:EBc

[Ano03c]  Anonymous.  Editorial

Board. Parallel Computing, 29(3):CO2, March 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2003:EBd

[Ano03d]  Anonymous.  Editorial

Board. Parallel Computing, 29(4):CO2, April 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2003:EBe

[Ano03e]  Anonymous.  Editorial

Board. Parallel Computing, 29(5):CO2, May 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2003:EBf

[Ano03f]  Anonymous.  Editorial

Board. Parallel Computing, 29(6):CO2, June 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2003:EBg

[Ano03g]  Anonymous.  Editorial

Board. Parallel Computing, 29(7):CO2, July 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2003:EBh

[Ano03h]  Anonymous.  Editorial

Board. Parallel Computing, 29(8):CO2, August 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2003:EBi

[Ano03i]  Anonymous.  Editorial

Board. Parallel Computing, 29(9):CO2, September 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2003:EBj

[Ano03j]  Anonymous.  Editorial

Board. Parallel Computing, 29(10):CO2, October 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2003:OHF

[Ano03k] Anonymous. Obituary: Harry F. Jordan. Parallel Computing, 29(4):iii, April 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2004:EBa

[Ano04a]  Anonymous.  Editorial

Board. Parallel Computing, 30(1):CO2, January 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2004:EBb

[Ano04b]  Anonymous.  Editorial

Board. Parallel Computing, 30(2):CO2, February 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2004:EBc

[Ano04c]  Anonymous.  Editorial

Board. Parallel Computing, 30(3):CO2, March 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2004:EBd

[Ano04d]  Anonymous.  Editorial

Board. Parallel Computing, 30(4):CO2, April 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2004:EBe

[Ano04e]  Anonymous.  Editorial

Board. Parallel Computing, 30(5–6):CO2, May/ June 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2004:EBf

[Ano04f]  Anonymous.  Editorial

Board. Parallel Computing, 30(7):CO2, July 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2004:EBg

[Ano04g]  Anonymous.  Editorial

Board. Parallel Computing, 30(8):CO2, August 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2004:EBh

[Ano04h]  Anonymous.  Edito-

rial Board. Parallel Computing, 30(9–10):CO2, September/October 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2006:EB

[Ano06]  Anonymous.  Editorial

Board. Parallel Computing, 32(1):iv–vi, January 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2008:AR

[Ano08]  Anonymous.  Acknowl-

edgement to reviewers. Parallel Computing, 34(1): 66–68, January 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2009:AR

[Ano09a]  Anonymous.  Acknowl-

edgement to reviewers. Parallel Computing, 35(1): 54–55, January 2009. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2009:EB

[Ano09b]  Anonymous.  Editorial

Board. Parallel Computing, ??(??):??, January 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2010:AR

[Ano10a] Anonymous. Acknowledgment to reviewers. Parallel Computing, 36(1):68– 69, January 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2010:EB

[Ano10b]  Anonymous.  Editorial

Board. Parallel Computing, 36(1):??, January 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2011:EBa

[Ano11a]  Anonymous.  Editorial

Board. Parallel Computing, 37(2):ifc, February 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911100007X.

Anonymous:2011:EBb

[Ano11b]  Anonymous.  Editorial

Board. Parallel Computing, 37(3):??, March 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000275.

Anonymous:2011:EBc

[Ano11c]  Anonymous.  Editorial

Board. Parallel Computing, 37(4–5):??, April/ May 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000469.

Anonymous:2011:EBd

[Ano11d]  Anonymous.  Editorial

Board. Parallel Computing, 37(6–7):??, June/ July 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000561.

Anonymous:2011:EBe

[Ano11e]  Anonymous.  Editorial

Board. Parallel Computing, 37(8):??, August 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000731.

Anonymous:2011:EBf

[Ano11f]  Anonymous.  Editorial

Board. Parallel Computing, 37(9):??, September 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000937.

Anonymous:2011:EBg

[Ano11g]  Anonymous.  Edito-

rial Board. Parallel Computing, 37(10–11): ??, October/November 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001128.

Anonymous:2011:EBh

[Ano11h]  Anonymous.  Editorial

Board. Parallel Computing, 37(12):??, December 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111001463.

Anonymous:2012:EBa

[Ano12a]  Anonymous.  Editorial

Board. Parallel Computing, 38(1–2):??, January/ February 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111001803.

Anonymous:2012:EBb

[Ano12b]  Anonymous.  Editorial

Board. Parallel Computing, 38(3):??, March 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781911200004X.

Anonymous:2012:EBc

[Ano12c]  Anonymous.  Editorial

Board. Parallel Computing, 38(4–5):??, April/ May 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819112000142.

Anonymous:2012:EBd

[Ano12d]  Anonymous.  Editorial

Board. Parallel Computing, 38(6–7):??, June/ July 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000257.

Anonymous:2012:EBe

[Ano12e]  Anonymous.  Editorial

Board. Parallel Computing, 38(8):??, August 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911200035X.

Anonymous:2012:EBf

[Ano12f]  Anonymous.  Editorial

Board. Parallel Computing, 38(9):??, September 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000506.

Anonymous:2012:EBg

[Ano12g]  Anonymous.  Edito-

rial Board. Parallel Computing, 38(10–11): ??, October/November 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819112000701.

Anonymous:2012:EBh

[Ano12h]  Anonymous.  Editorial

Board. Parallel Computing, 38(12):??, December 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819112000828.

Anonymous:2013:EBa

[Ano13a]  Anonymous.  Editorial

Board. Parallel Computing, 39(1):??, January 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000045.

Anonymous:2013:EBb

[Ano13b]  Anonymous.  Editorial

Board. Parallel Computing, 39(2):??, February 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000124.

Anonymous:2013:EBc

[Ano13c]  Anonymous.  Editorial

Board. Parallel Computing, 39(3):??, March 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000240.

Anonymous:2013:EBd

[Ano13d]  Anonymous.  Editorial

Board. Parallel Computing, 39(4–5):??, April/ May 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000380.

Anonymous:2013:EBe

[Ano13e]  Anonymous.  Editorial

Board. Parallel Computing, 39(6–7):??, June/ July 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000598.

Anonymous:2013:EBf

[Ano13f]  Anonymous.  Editorial

Board. Parallel Computing, 39(8):??, August 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000756.

Anonymous:2013:EBg

[Ano13g]  Anonymous.  Editorial

Board. Parallel Computing, 39(9):??, September 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000860.

Anonymous:2013:EBh

[Ano13h]  Anonymous.  Editorial

Board. Parallel Computing, 39(10):??, October 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113001099.

Anonymous:2013:PIS

[Ano13i] Anonymous. Preface: Infrastructure for scalable tools. Parallel Computing, 39(3):113, March 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000227.

Anonymous:2014:EBa

[Ano14a]  Anonymous.  Editorial

Board. Parallel Computing, 40(5–6):IFC, May 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000556.

Anonymous:2014:EBb

[Ano14b]  Anonymous.  Editorial

Board. Parallel Computing, 40(7):IFC, July 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000702.

Anonymous:2014:EBc

[Ano14c]  Anonymous.  Editorial

Board. Parallel Computing, 40(8):IFC, August 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000866.

Anonymous:2014:EBd

[Ano14d]  Anonymous.  Editorial

Board. Parallel Computing, 40(9):IFC, October 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000982.

Anonymous:2014:EBe

[Ano14e]  Anonymous.  Editorial

Board. Parallel Computing, 40(10):IFC, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114001318.

Anonymous:2015:BPA

[Ano15a] Anonymous. Best papers from ACM Computing Frontiers 2014 Conference. Parallel Computing, 45 (??):1, June 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000654.

Anonymous:2015:EBa

[Ano15b]  Anonymous.  Editorial

Board. Parallel Computing, 41(??):IFC, January 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000034.

Anonymous:2015:EBb

[Ano15c]  Anonymous.  Editorial

Board. Parallel Computing, 42(??):IFC, February 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000228.

Anonymous:2015:EBc

[Ano15d]  Anonymous.  Editorial

Board. Parallel Computing, 43(??):ifc, March 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000356.

Anonymous:2015:EBd

[Ano15e]  Anonymous.  Editorial

Board. Parallel Computing, 44(??):ifc, May 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000526.

Anonymous:2015:EBe

[Ano15f]  Anonymous.  Editorial

Board. Parallel Computing, 45(??):ifc, June 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000678.

Anonymous:2015:EBf

[Ano15g]  Anonymous.  Editorial

Board. Parallel Computing, 46(??):ifc, July 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000824.

Anonymous:2015:EBg

[Ano15h]  Anonymous.  Editorial

Board. Parallel Computing, 47(??):ifc, August 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000952.

Anonymous:2015:EBh

[Ano15i]  Anonymous.  Editorial

Board. Parallel Computing, 48(??):ifc, October 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115001106.

Anonymous:2015:EBi

[Ano15j]  Anonymous.  Editorial

Board. Parallel Computing, 49(??):ifc, November 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001271.

Anonymous:2015:EBj

[Ano15k]  Anonymous.  Editorial

Board. Parallel Computing, 50(??):ifc, December 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001507.

Anonymous:2016:EBa

[Ano16a]  Anonymous.  Editorial

Board. Parallel Computing, 51(??):ifc, January 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000260.

Anonymous:2016:EBb

[Ano16b]  Anonymous.  Editorial

Board. Parallel Computing, 52(??):ifc, February 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000466.

Anonymous:2016:EBc

[Ano16c]  Anonymous.  Editorial

Board. Parallel Computing, 53(??):ifc, April 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300047.

Anonymous:2016:EBd

[Ano16d]  Anonymous.  Editorial

Board. Parallel Computing, 54(??):ifc, May 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300163.

Anonymous:2016:EBe

[Ano16e]  Anonymous.  Editorial

Board. Parallel Computing, 55(??):ifc, July 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300266.

Anonymous:2016:EBf

[Ano16f]  Anonymous.  Editorial

Board. Parallel Computing, 56(??):ifc, August 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300503.

Anonymous:2016:EBg

[Ano16g]  Anonymous.  Editorial

Board. Parallel Computing, 57(??):ifc, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300679.

Anonymous:2016:EBh

[Ano16h]  Anonymous.  Editorial

Board. Parallel Computing, 58(??):ifc, October 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300850.

Anonymous:2016:EBi

[Ano16i]  Anonymous.  Editorial

Board. Parallel Computing, 59(??):ifc, November 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116301107.

Anonymous:2016:EBj

[Ano16j]  Anonymous.  Editorial

Board. Parallel Computing, 60(??):ifc, December 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116301211.

Anonymous:2016:PIS

[Ano16k]  Anonymous.  Preface:

26th International Symposium on Computer Architecture and High Performance Computing. Parallel Computing, 54(??):1, May 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911630014X.

Anonymous:2017:EBk

[Ano17a]  Anonymous.  Editorial

Board. Parallel Computing, 61(??):ifc, January 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300030.

Anonymous:2017:EBl

[Ano17b]  Anonymous.  Editorial

Board. Parallel Computing, 62(??):ifc, February 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300212.

Anonymous:2017:EBm

[Ano17c]  Anonymous.  Editorial

Board. Parallel Computing, 63(??):ifc, April 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300303.

Anonymous:2017:EBn

[Ano17d]  Anonymous.  Editorial

Board. Parallel Computing, 64(??):ifc, May 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300431.

Anonymous:2017:EBo

[Ano17e]  Anonymous.  Editorial

Board. Parallel Computing, 65(??):ifc, July 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300649.

Anonymous:2017:EBp

[Ano17f]  Anonymous.  Editorial

Board. Parallel Computing, 66(??):ifc, August 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300790.

Anonymous:2017:EBq

[Ano17g]  Anonymous.  Editorial

Board. Parallel Computing, 67(??):ifc, September 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301102.

Anonymous:2017:EBr

[Ano17h]  Anonymous.  Editorial

Board. Parallel Computing, 68(??):ifc, October 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301023.

Anonymous:2017:EBs

[Ano17i]  Anonymous.  Editorial

Board. Parallel Computing, 69(??):ifc, November 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301576.

Anonymous:2017:EBt

[Ano17j]  Anonymous.  Editorial

Board. Parallel Computing, 70(??):ifc, December 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301667.

Anonymous:2018:EBa

[Ano18a]  Anonymous.  Editorial

Board. Parallel Computing, 71(??):ii, January 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301989.

Anonymous:2018:EBb

[Ano18b]  Anonymous.  Editorial

Board. Parallel Computing, 72(??):ii, February 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300206.

Anonymous:2018:EBc

[Ano18c]  Anonymous.  Editorial

Board. Parallel Computing, 74(??):ii, ???? 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118300383.

Anonymous:2018:EBd

[Ano18d]  Anonymous.  Editorial

Board. Parallel Computing, 74(??):ii, ???? 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118300449.

Anonymous:2018:EBe

[Ano18e]  Anonymous.  Editorial

board. Parallel Computing, 75(??):ii, July 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118301352.

Anonymous:2018:EBf

[Ano18f]  Anonymous.  Editorial

board. Parallel Computing, 76(??):ii, August 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118301637.

Anonymous:2018:EBg

[Ano18g]  Anonymous.  Editorial

Board. Parallel Computing, 77(??):ii, September 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118302059.

Anonymous:2018:EBh

[Ano18h]  Anonymous.  Editorial

Board. Parallel Computing, 78(??):ii, October 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118302370.

Anonymous:2018:EBi

[Ano18i]  Anonymous.  Editorial

Board. Parallel Computing, 79(??):ii, November 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118302771.

Anonymous:2018:EBj

[Ano18j]  Anonymous.  Editorial

Board. Parallel Computing, 80(??):ii, December 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303661.

Anonymous:2018:FSI

[Ano18k]  Anonymous.  Foreword

for the special issue on the best papers from the EuroMPI 2016 conference. Parallel Computing, 72(??):42, February 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118300103.

Anonymous:2018:RA

[Ano18l] Anonymous. Reviewer acknowledgement 2017. Parallel Computing, 71(??):I– II, January 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117302016.

Anonymous:2019:D

[Ano19a] Anonymous. December 2019. Parallel Computing, 90(??):??, December 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2019:EBa

[Ano19b]  Anonymous.  Editorial

Board. Parallel Computing, 81(??):ii, January 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119300079.

Anonymous:2019:EBb

[Ano19c]  Anonymous.  Editorial

Board. Parallel Computing, 82(??):ii, ???? 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119300286.

Anonymous:2019:EBc

[Ano19d]  Anonymous.  Editorial

Board. Parallel Computing, 83(??):ii, April 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119300481.

Anonymous:2019:EBd

[Ano19e]  Anonymous.  Editorial

Board. Parallel Computing, 84(??):ii, May 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911930081X.

Anonymous:2019:EBe

[Ano19f]  Anonymous.  Editorial

Board. Parallel Computing, 85(??):ii, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301097.

Anonymous:2019:EBf

[Ano19g]  Anonymous.  Editorial

Board. Parallel Computing, 86(??):ii, August 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301218.

Anonymous:2019:EBg

[Ano19h]  Anonymous.  Editorial

Board. Parallel Computing, 87(??):ii, September 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301322.

Anonymous:2019:EBh

[Ano19i]  Anonymous.  Editorial

Board. Parallel Computing, 88(??):Article 102556, ???? 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301474.

Anonymous:2019:EBi

[Ano19j]  Anonymous.  Editorial

Board. Parallel Computing, 89(??):Article 102577, November 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301681.

Anonymous:2019:EBj

[Ano19k]  Anonymous.  Editorial

Board. Parallel Computing, 90(??):Article 102593, December 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911930184X.

Anonymous:2019:N

[Ano19l] Anonymous. November 2019. Parallel Computing, 89(??):??, November 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2019:PN

[Ano19m] Anonymous. Publisher’s note. Parallel Computing, 88(??):Article 102557, ???? 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301486.

Anonymous:2020:Aa

[Ano20a] Anonymous. April 2020. Parallel Computing, 92 (??):??, April 2020. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2020:Ab

[Ano20b] Anonymous. August 2020. Parallel Computing, 96 (??):??, August 2020. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2020:D

[Ano20c] Anonymous. December 2020. Parallel Computing, 100(??):??, December 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2020:EBa

[Ano20d]  Anonymous.  Editorial

Board. Parallel Computing, 91(??):Article 102603, March 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819119301942.

Anonymous:2020:EBb

[Ano20e]  Anonymous.  Editorial

Board. Parallel Computing, 92(??):Article 102612, April 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300053.

Anonymous:2020:EBc

[Ano20f]  Anonymous.  Editorial

Board. Parallel Computing, 93(??):Article 102628, May 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300211.

Anonymous:2020:EBd

[Ano20g]  Anonymous.  Editorial

Board. Parallel Computing, 96(??):Article 102654, August 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300478.

Anonymous:2020:EBe

[Ano20h]  Anonymous.  Editorial

Board. Parallel Computing, 97(??):Article 102676, September 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300673.

Anonymous:2020:EBf

[Ano20i]  Anonymous.  Editorial

Board. Parallel Computing, 98(??):Article 102690, October 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819120300806.

Anonymous:2020:EBg

[Ano20j]  Anonymous.  Editorial

Board. Parallel Computing, 99(??):Article 102714, November 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819120300995.

Anonymous:2020:EBh

[Ano20k]  Anonymous.  Editorial

Board. Parallel Computing, 100(??):Article 102728, December 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0167819120301125.

Anonymous:2020:Ma

[Ano20l] Anonymous. March 2020. Parallel Computing, 91 (??):??, March 2020. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2020:Mb

[Ano20m] Anonymous. May 2020.

Parallel Computing, 93 (??):??, May 2020. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2020:N

[Ano20n] Anonymous. November 2020. Parallel Computing, 99(??):??, November 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2020:O

[Ano20o]  Anonymous.  October

2020. Parallel Computing, 98(??):??, October 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2020:S

[Ano20p] Anonymous. September 2020. Parallel Computing, 97(??):??, September 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2021:A

[Ano21a] Anonymous. April 2021.

Parallel Computing, 101 (??):??, April 2021. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2021:D

[Ano21b] Anonymous. December 2021. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2021:EBa

[Ano21c]  Anonymous.  Editorial

Board. Parallel Computing, 94–95(??):Article 102644, June 2021. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819120300375.

Anonymous:2021:EBb

[Ano21d]  Anonymous.  Editorial

Board. Parallel Computing, 101(??):Article 102744, April 2021. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819121000041.

Anonymous:2021:EBc

[Ano21e]  Anonymous.  Editorial

Board. Parallel Computing, 102(??):Article 102764, May 2021. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0167819121000223.

Anonymous:2021:EBd

[Ano21f]  Anonymous.  Editorial

Board. Parallel Computing, 103(??):??, June 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000314.

Anonymous:2021:EBe

[Ano21g]  Anonymous.  Editorial

Board. Parallel Computing, 104–105(??):??, July 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000508.

Anonymous:2021:EBf

[Ano21h]  Anonymous.  Editorial

Board. Parallel Computing, 104–105(??):??, July 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000508.

Anonymous:2021:EBg

[Ano21i]  Anonymous.  Editorial

Board. Parallel Computing, 106(??):??, September 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000673.

Anonymous:2021:EBh

[Ano21j]  Anonymous.  Editorial

Board. Parallel Computing, 107(??):??, October 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000910.

Anonymous:2021:EBi

[Ano21k]  Anonymous.  Editorial

Board. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121001095.

Anonymous:2021:Jc

[Ano21l] Anonymous. July 2021. Parallel Computing, 104– 105(??):??, July 2021. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2021:Jd

[Ano21m] Anonymous. July 2021. Parallel Computing, 104– 105(??):??, July 2021. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2021:Ja

[Ano21n] Anonymous. June 2020.

Parallel Computing, 94–95 (??):??, June 2021. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2021:Jb

[Ano21o] Anonymous. June 2021.

Parallel Computing, 103 (??):??, June 2021. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2021:M

[Ano21p] Anonymous. May 2021.

Parallel Computing, 102 (??):??, May 2021. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2021:O

[Ano21q]  Anonymous.  October

2021. Parallel Computing, 107(??):??, October 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2021:S

[Ano21r] Anonymous. September 2021. Parallel Computing, 106(??):??, September 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2022:D

[Ano22a] Anonymous. December 2022. Parallel Computing, 114(??):??, December 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2022:EBa

[Ano22b]  Anonymous.  Editorial

Board. Parallel Computing, 109(??):??, March 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121001228.

Anonymous:2022:EBb

[Ano22c]  Anonymous.  Editorial

Board. Parallel Computing, 110(??):??, May 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000151.

Anonymous:2022:EBc

[Ano22d]  Anonymous.  Editorial

Board. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819122000321.

Anonymous:2022:EBd

[Ano22e]  Anonymous.  Editorial

Board. Parallel Computing, 112(??):??, September 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000540.

Anonymous:2022:EBe

[Ano22f]  Anonymous.  Editorial

Board. Parallel Computing, 113(??):??, October 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000679.

Anonymous:2022:EB

[Ano22g]  Anonymous.  Editorial

Board. Parallel Computing, 114(??):??, December 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000771.

Anonymous:2022:J

[Ano22h] Anonymous. July 2022.

Parallel Computing, 111 (??):??, July 2022. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2022:Ma

[Ano22i] Anonymous. March 2022. Parallel Computing, 109 (??):??, March 2022. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2022:Mb

[Ano22j] Anonymous. May 2022. Parallel Computing, 110 (??):??, May 2022. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2022:O

[Ano22k]  Anonymous.  October

2022. Parallel Computing, 113(??):??, October 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2022:S

[Ano22l] Anonymous. September 2022. Parallel Computing, 112(??):??, September 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2022:SIS

[Ano22m] Anonymous. Special issue of selected papers from EuroMPI/USA 2020. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121001162.

Anonymous:2023:EBa

[Ano23a]  Anonymous.  Editorial

Board. Parallel Computing, 115(??):??, February 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000194.

Anonymous:2023:EBb

[Ano23b]  Anonymous.  Editorial

Board. Parallel Computing, 116(??):??, July 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000352.

Anonymous:2023:EBc

[Ano23c]  Anonymous.  Editorial

Board. Parallel Computing, 117(??):??, September 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000534.

Anonymous:2023:EBd

[Ano23d]  Anonymous.  Editorial

Board. Parallel Computing, 118(??):??, November 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819123000637.

Anonymous:2023:F

[Ano23e]  Anonymous.  February

2023. Parallel Computing, 115(??):??, February 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2023:J

[Ano23f] Anonymous. July 2023. Parallel Computing, 116 (??):??, July 2023. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2023:N

[Ano23g] Anonymous. November 2023. Parallel Computing, 118(??):??, November 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2023:RA

[Ano23h]  Anonymous.  Reviewer

acknowledgment. Parallel Computing, 115(??):??, February 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000108.

Anonymous:2023:S

[Ano23i] Anonymous. September 2023. Parallel Computing, 117(??):??, September 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2024:EBa

[Ano24a]  Anonymous.  Editorial

Board. Parallel Computing, 119(??):??, February 2024. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819124000140.

Anonymous:2024:EBb

[Ano24b]  Anonymous.  Editorial

Board. Parallel Computing, 120(??):??, June 2024. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819124000279.

Anonymous:2024:F

[Ano24c]  Anonymous.  February

2024. Parallel Computing, 119(??):??, February 2024.

CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:2024:J

[Ano24d] Anonymous. June 2024. Parallel Computing, 120 (??):??, June 2024. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Adams:1989:APP

[AO89] Loyce M. Adams and Elizabeth G. Ong. Additive polynomial preconditioners for parallel computers. Parallel Computing, 9(3):333–345, February 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Aktilav:2022:PAP

[AÖ22] Busenur Aktilav and Isil Öz. Performance and accuracy predictions of approximation methods for shortest-path algorithms on GPUs. Parallel Computing, 112(??):??, September 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819122000400.

Aji:2016:MEA

[APBcF16] Ashwin M. Aji, Antonio J. Peña, Pavan Balaji, and Wu chun Feng. MultiCL: Enabling automatic scheduling for task-parallel workloads in OpenCL. Parallel Computing, 58(??):37–55, October 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300357.

Aluru:1992:RNG

[APG92] Srinivas Aluru, G. M.

Prabhu, and John Gustafson. A random number generator for parallel computers. Parallel Computing, 18(8):839–847, August 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Abad:2015:ILL

[APPG15] Pablo Abad, Pablo Prieto, Valentin Puente, and Jose-Angel Gregorio. Improving last level shared cache performance through mobile insertion policies (MIP). Parallel Computing, 49(??):13–27, November 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000800.

Arnaldi:1997:VMS

[APRP97] Bruno Arnaldi, Thierry Priol, Luc Renambot, and Xavier Pueyo. Visibility masks for solving complex radiosity computations on multiprocessors. Parallel Computing, 23(7):887–897, July 14, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/free/1997/
                                        23/7/arnaldi/fig4c.gif.

Akl:1993:NRS

[AQ93] S. G. Akl and Ke Qiu. A novel routing scheme on the star and pancake networks and its applications. Parallel Computing, 19(1): 95–101, January 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Antonopoulos:2018:PPR

[AQO18] Christos D. Antonopoulos and Enrique S. Quintana-Ortí. Parallel programming for resilience and energy efficiency. Parallel Computing, 74(??):1–2, ???? 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117301229.

Arafeh:2003:TDS

[Ara03] Bassel R. Arafeh. A task duplication scheme for resolving deadlocks in clustered DAGs. Parallel Computing, 29(6):795–820, June 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Arbenz:1992:DCA

[Arb92] Peter Arbenz. Divide and conquer algorithms for the bandsymmetric eigenvalue problem. Parallel Computing, 18(10):1105–1128, October 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Asenov:1994:SSI

[ARB94] A. Asenov, D. Reid, and J. R. Barker. Speed-up of scalable iterative linear solvers implemented on an array of transputers. Parallel Computing, 20(3):375–387, March 10, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=3&aid=832.

Asenov:1995:SSI

[ARB95] A. Asenov, D. Reid, and J. R. Barker. Speed-up of scalable iterative linear solvers implemented on an array of transputers. Parallel Computing, 21(4):669–682, April 1, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=4&aid=931.

Al-Rawi:2005:TMO

[ARCH05] Ghazi Al-Rawi, John Cioffi, and Mark Horowitz. On task mapping optimization for parallel decoding of low-density parity-check codes on message-passing architectures. Parallel Computing, 31(5):462–490, May 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Artemov:2019:PSA

[ARR19] Anton G. Artemov, Elias Rudberg, and Emanuel H. Rubensson. Parallelization and scalability analysis of inverse factorization using the chunks and tasks programming model. Parallel Computing, 89(??): Article 102548, November 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301395.

Arvind:1992:DCR

[Arv92] D. K. Arvind. On the detection of communicationrelated errors in concurrent programs. Parallel Computing, 18(12):1381– 1392, December 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Altekar:1993:PTA

[ARW93] S. D. Altekar, A. K. Ray, and B. R. Wienke. On the parallelization of a S_{n} transport algorithm on a CRAY Y MP. Parallel Computing, 19(7):823–834, July 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ancona:1997:TBI

[ARZ97] Fabio Ancona, Stefano Rovetta, and Rodolfo Zunino. Transputer-based implementation of distributed associative memories. Parallel Computing, 23(10):1479–1491, November 7, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=10&aid=1226.

Akl:1986:SSS

[AS86] S. G. Akl and H. Schmeck. Systolic sorting in a sequential input/output environment. Parallel Computing, 3(1):11–23, March 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Atiquzzaman:2000:PCC

[AS00] Mohammed Atiquzzaman and Pradip K. Srimani. Parallel computing on clusters of workstations. Parallel Computing, 26 (2–3):175–177, February 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/25/23/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/25/23/
article.pdf.

Acer:2016:IPS

[ASA16] Seher Acer, Oguz Selvitopi, and Cevdet Aykanat. Improving performance of sparse matrix dense matrix multiplication on large-scale parallel systems. Parallel Computing, 59(??):71–96, November 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819116301041.

Al-Sadi:2001:FTR

[ASDOK01] J. Al-Sadi, K. Day, and M. Ould-Khaoua. Faulttolerant routing in hypercubes using probability vectors. Parallel Computing, 27(10):1381–1399, September 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/40/32/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/40/32/
article.pdf.

Audet:1992:PIV

[ASH92] D. Audet, Y. Savaria, and J.-L. Houle. Performance improvements to VLSI parallel systems, using dynamic concatenation of processing resources. Parallel Computing, 18(2): 149–167, February 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Arbenz:2011:SIP

[ASSS11] Peter Arbenz, Yousef Saad, Ahmed Sameh, and Olaf Schenk. Special issue on Parallel Matrix Algorithms and Applications (PMAA’10). Parallel Computing, 37 (12):731–732, December 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001426.

Aref:1998:LMA

[AT98] Mostafa M. Aref and Mohammed A. Tayyib. Lana-Match algorithm: a parallel version of the Rete-Match algorithm. Parallel Computing, 24(5–6):763– 775, June 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/5-6/1280.pdf.

Alvanos:2016:USD

[ATA^{+}16] Michail Alvanos, Ettore Tiotto, José Nelson Amaral, Montse Farreras, and Xavier Martorell. Using shared-data localization to reduce the cost of inspector-execution in unified-parallel-c programs. Parallel Computing, 54(??):2–14, May 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300096.

Amritkar:2012:OPF

[ATL^{+}12] Amit Amritkar, Danesh Tafti, Rui Liu, Rick Kufrin, and Barbara Chapman. OpenMP parallelism for fluid and fluid-particulate systems. Parallel Computing, 38 (9):501–517, September 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000476.

Adamides:1989:SAC

[ATT89] E. Adamides, Ph. Tsalides, and A. Thanailakis. Synchronization of asynchronous concurrent processes using cellular automata. Parallel Computing, 11(2):163– 169, ???? 4, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Adamides:1992:HCA

[ATT92] E. D. Adamides, Ph.

Tsalides, and A. Thanailakis. Hierarchical cellular automata structures. Parallel Computing, 18(5):517– 524, May 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Asano:1988:SAC

[AU88] T. Asano and H. Umeo.

Systolic algorithms for computing the visibility polygon and triangulation of a polygonal region. Parallel Computing, 6(2):209– 216, February 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Aubanel:2011:STP

[Aub11] Eric Aubanel. Scheduling of tasks in the parareal algorithm. Parallel Computing, 37(3):172–182, March 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819110001419.

Asai:2015:ICP

[AV15] Ryo Asai and Andrey Vladimirov. Intel Cilk Plus for complex parallel algorithms: “Enormous Fast Fourier Transforms” (EFFT) library. Parallel Computing, 48(??):125– 142, October 2015. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0167819115000782.

Allouche:1997:LCA

[AvHL^{+}97] J.-P. Allouche, F. v. Haeseler, E. Lange, A. Petersen, and G. Skordev. Linear cellular automata and automatic sequences. Parallel Computing, 23 (11):1577–1592, December 1, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=11&aid=1232.

Ayala:2013:PIS

[AW13] Orlando Ayala and Lian-Ping Wang. Parallel implementation and scalability analysis of 3D Fast Fourier Transform using 2D domain decomposition. Parallel Computing, 39(1):58– 77, January 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000932.

Axelrod:1986:ESB

[Axe86] Tim S. Axelrod. Effects of synchronization barriers on multiprocessor performance. Parallel Computing, 3(2):129–140, May 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Amme:1998:DDA

[AZ98] Wolfram Amme and Eberhard Zehendner. Data dependence analysis in programs with pointers. Parallel Computing, 24 (3–4):505–525, May 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/3-4/1300.pdf.

Brightwell:2000:MPC

[B^{+}00] Ron Brightwell et al. Massively parallel computing using commodity components. Parallel Computing, 26(2–3):243–266, February 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/42/25/26/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/25/26/
                                        article.pdf.

Beynon:2001:DPV

[B^{+}01] Michael D. Beynon et al. Distributed processing of very large datasets with DataCutter. Parallel Computing, 27(11):1457–1478, October 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/47/41/31/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/41/31/
                                        article.pdf.

Beynon:2002:PLS

[B^{+}02] Michael Beynon et al. Processing large-scale multidimensional data in parallel and distributed environments. Parallel Computing, 28(5):827–859, May 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/57/35/
abstract.html.

Bhandarkar:1995:RMT

[BA95] Suchendra M. Bhandarkar and Hamid R. Arabnia. The REFINE multiprocessor — theoretical properties and algorithms. Parallel Computing, 21 (11):1783–1805, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=11&aid=1011.

Ben-Asher:2002:PCS

[BA02] Yosi Ben-Asher. The parallel client-server paradigm. Parallel Computing, 28(3): 503–523, March 2002. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/gej-ng/10/35/21/

                                        60/34/34/abstract.html;
                                         http://www.elsevier.
                                        nl/gej-ng/10/35/21/60/
                                        34/34/main.pdf.

Bader:2008:HPC

[BA08] David A. Bader and Srinivas Aluru. Highperformance computational biology. Parallel Computing, 34(11):613– 615, November 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Beck:2002:MUS

[BAB^{+}02] Micah Beck, Dorian Arnold, Alessandro Bassi, Fran Berman, Henri Casanova, Jack Dongarra, Terry Moore, Graziano Obertelli, James Plank, and Martin Swany. Middleware for the use of storage in communication. Parallel Computing, 28(12):1773–1787, December 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Baillie:1988:CSD

[Bai88] C. F. Baillie. Comparing shared and distributed memory computers. Parallel Computing, 8(1–3):101– 110, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bai:1999:CAP

[Bai99] Zhong-Zhi Bai. A class of asynchronous parallel multisplitting blockwise relaxation methods. Parallel Computing, 25(6):681–701, June 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/6/1407.pdf.

Bosch:2020:ARD

[BÁJG^{+}20] Jaume Bosch, Carlos Álvarez, Daniel Jiménez-González, Xavier Martorell, and Eduard Ayguadé. Asynchronous runtime with distributed manager for task-based programming models. Parallel Computing, 97(??): Article 102664, September 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300570.

Bakker:2000:SPB

[Bak00] J. A. Bakker. Semantic partitioning as a basis for parallel I/O in database management systems. Parallel Computing, 26(11):1491–1513, October 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/42/32/27/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/32/27/
                                        article.pdf.

Bader:2009:CDT

[BAK09] David A. Bader, Virat Agarwal, and Seunghwa Kang. Computing discrete transforms on the Cell Broadband Engine. Parallel Computing, 35(3):119– 137, March 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bader:2007:HPC

[BAMK07] David A. Bader, Virat Agarwal, Kamesh Madduri, and Seunghwa Kang. High performance combinatorial algorithm design on the Cell Broadband Engine processor. Parallel Computing, 33(10–11): 720–740, November 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Banks:1997:SPD

[Ban97] David C. Banks. Screenparallel determination of intersection curves. Parallel Computing, 23(7):953– 960, July 14, 1997. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cas/tree/store/parco/
free/1997/23/7/banks/
fig3c.tif.

Barada:1995:MMC

[Bar95] H. R. Barada. Modular matrix computations on multi-linear VLSI arrays. Parallel Computing, 21(8):1269–1284, August 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=8&aid=999.

Barth:1997:PMP

[Bar97] Dominique Barth. Parallel matrix product algorithm in the de Bruijn network using emulation of meshes of trees. Parallel Computing, 22(12): 1563–1578, February 21, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=12&aid=1115.

Benkahla:1998:PED

[BAR98] O. Benkahla, C. Aktouf, and C. Robach. Performance evaluation of distributed diagnosis algorithms in parallel systems. Parallel Computing, 24(8):1205–1222, August 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/8/1325.pdf.

Barlas:2012:CBO

[Bar12] Gerassimos Barlas. Clusterbased optimized parallel video transcoding. Parallel Computing, 38(4– 5):226–244, April/May 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000117.

Basu:1994:PAS

[Bas94] Amit J. Basu. A parallel algorithm for spectral solution of the threedimensional Navier–Stokes equations. Parallel Computing, 20(8):1191–1204, August 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=8&aid=874.

Buyukkececi:2013:POI

[BAS13] Ferit Büyükkeçeci, Omar Awile, and Ivo F. Sbalzarini. A portable OpenCL implementation of generic particle-mesh and meshparticle interpolation in 2D and 3D. Parallel Computing, 39(2):94–111, February 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000920.

Bruguera:1993:DPR

[BAZ93] J. D. Bruguera, E. Antelo, and E. L. Zapata. Design of a pipelined radix 4 CORDIC processor. Parallel Computing, 19(7):729– 744, July 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Booth:2021:NSS

[BB21] Joshua Dennis Booth and Gregory Bolet. An onnode scalable sparse incomplete LU factorization for a many-core iterative solver with Javelin. Parallel Computing, 94–95 (??):Article 102622, June 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300156.

Bala:1994:IEU

[BBB^{+}94] Vasanth Bala, Jehoshua Bruck, Raymond Bryant, Robert Cypher, Peter de Jong, Pablo Elustondo, D. Frye, Alex Ho, Ching-Tien Ho, Gail Irwin, Shlomo Kipnis, Richard Lawrence, and Marc Snir. The IBM External User Interface for scalable parallel systems. Parallel Computing, 20(4):445–462, March 31, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=4&aid=862.

Baboulin:2014:EDR

[BBB^{+}14] Marc Baboulin, Dulceneia Becker, George Bosilca, Anthony Danalis, and Jack Dongarra. An efficient distributed randomized algorithm for solving large dense symmetric indefinite linear systems. Parallel Computing, 40(7):213–223, July 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113001488.

Buurlage:2019:GPM

[BBB19] J. W. Buurlage, R. H. Bisseling, and K. J. Batenburg. A geometric partitioning method for distributed tomographic reconstruction. Parallel Computing, 81(??):104– 121, January 2019. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819118301856.

Bak:2022:OAE

[BBB^{+}22] Seonmyeong Bak, Colleen Bertoni, Swen Boehm, Reuben Budiardja, Barbara M. Chapman, Johannes Doerfert, Markus Eisenbach, Hal Finkel, Oscar Hernandez, Joseph Huber, Shintaro Iwasaki, Vivek Kale, Paul R. C. Kent, JaeHyuk Kwack, Meifeng Lin, Piotr Luszczek, Ye Luo, Buu Pham, Swaroop Pophale, Kiran Ravikumar, Vivek Sarkar, Thomas Scogland, Shilei Tian, and P. K. Yeung. OpenMP application experiences: Porting to accelerated nodes. Parallel Computing, 109(??):??, March 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121001009.

Beletska:2011:CGL

[BBC^{+}11] Anna Beletska, Włodzimierz Bielecki, Albert Cohen, Marek Palkowski, and Krzysztof Siedlecki. Coarsegrained loop parallelization: Iteration Space Slicing vs affine transformations. Parallel Computing, 37(8):479–497, August 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111000032.

Bosilca:2012:DGD

[BBD^{+}12] George Bosilca, Aurelien Bouteiller, Anthony Danalis, Thomas Herault, Pierre Lemarinier, and Jack Dongarra. DAGuE: a generic distributed DAG engine for High Performance Computing. Parallel Computing, 38(1–2): 37–51, January/February 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111001347.

Bouam:2021:CRA

[BBDN21] Mellila Bouam, Charles Bouillaguet, Claire Delaplace, and Camille Noûs. Computational records with aging hardware: Controlling half the output of SHA-256. Parallel Computing, 106(??):??, September 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000557.

Baker:1990:MIL

[BBK90] M. A. Baker, K. C. Bowler, and R. D. Kenway. MIMD implementations of linear solvers for oil reservoir simulation. Parallel Computing, 16(2–3):313–334, December 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Balzuweit:2016:LSI

[BBL^{+}16] Evan Balzuweit, David P.

Bunde, Vitus J. Leung, Austin Finley, and Alan C. S. Lee. Local search to improve coordinate-based task mapping. Parallel Computing, 51(??):67–78, January 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001441.

Bunn:2019:SCC

[BBL^{+}19] C. Bunn, H. Barclay, A. Lazarev, F. Yusuf, J. Fitch, J. Booth, K. Shivdikar, and D. Kaeli. Student cluster competition 2018, team northeastern university: Reproducing performance of a multiphysics simulations of the Tsunamigenic 2004 Sumatra Megathrust earthquake on the AMD EPYC 7551 architecture. Parallel Computing, 90(??): Article 102568, December 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301590.

Benner:2000:SAR

[BBQOQO00] Peter Benner, Ralph Byers, Enrique S. Quintana-Ortí, and Gregorio Quintana-Ortí. Solving algebraic Riccati equations on parallel computers using Newton’s method with exact line search. Parallel Computing, 26(10):1345–1368, August 15, 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/31/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/31/28/
article.pdf.

Bertsekas:1991:PSA

[BC91] D. P. Bertsekas and D. A. Castanon. Parallel synchronous and asynchronous implementations of the auction algorithm. Parallel Computing, 17 (6–7):707–732, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Berthou:1997:WAP

[BC97] Jean-Yves Berthou and Laurent Colombet. Which approach to parallelizing scientific codes — that is the question. Parallel Computing, 23 (1–2):165–179, April 16, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=1-2&aid=1154.

Bhandarkar:2004:PPM

[BC04] Suchendra M. Bhandarkar and Shankar R. Chandrasekaran. Parallel parsing of MPEG video on a shared-memory symmetric multiprocessor. Parallel Computing, 30(11):1233– 1276, November 2004. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Balaji:2019:SIM

[BC19a] Pavan Balaji and Marc Casas. Special issue on the message passing interface. Parallel Computing, 86(??):14–15, August 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911930095X.

Budiardja:2019:TGO

[BC19b] Reuben D. Budiardja and Christian Y. Cardall. Targeting GPUs with OpenMP directives on Summit: a simple and effective Fortran experience. Parallel Computing, 88(??):Article 102544, ???? 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301358.

Bekas:2008:AWI

[BCA08] C. Bekas, A. Curioni, and W. Andreoni. Atomic wavefunction initialization in ab initio molecular dynamics using distributed Lanczos. Parallel Computing, 34(6–8):441–450, July 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Benkrid:2002:TGF

[BCB02] K. Benkrid, D. Crookes, and A. Benkrid. Towards a general framework for FPGA based image processing using hardware skeletons. Parallel Computing, 28(7–8):1141–1154, August 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/58/37/
abstract.html.

Brandes:1997:HSIa

[BCC^{+}97a] T. Brandes, S. Chaumette, M. C. Counilh, J. Roman, A. Darte, F. Desprez, and J. C. Mignot. HP-FIT: a set of integrated tools for the parallelization of applications using High Performance Fortran. Part I: HPFIT and the TransTOOL environment. Parallel Computing, 23(1–2):71–87, April 16, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=1-2&aid=1148.

Brandes:1997:HSIb

[BCC^{+}97b] T. Brandes, S. Chaumette, M. C. Counilh, J. Roman, F. Desprez, and J. C. Mignot. HPFIT: a set of integrated tools for the parallelization of applications using High Performance Fortran. Part II: Data-structure visualization and HPF extensions for irregular problems. Parallel Computing, 23(1–2):89–105, April 16, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=1-2&aid=1149.

Bernardon:2007:AFV

[BCCS07] Fábio F. Bernardon, Steven P. Callahan, João L. D. Comba, and Cláudio T. Silva. An adaptive framework for visualizing unstructured grids with timevarying scalar fields. Parallel Computing, 33(6):391– 405, June 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Baranowski:2017:RPS

[BCE^{+}17] Marek Baranowski, Braden Caywood, Hannah Eyre, Janaan Lake, Kevin Parker, Kincaid Savoie, Hari Sundar, and Mary Hall. Reproducing ParConnect for SC16. Parallel Computing, 70(??):18–21, December 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117300923.

Bourbeau:2000:BBP

[BCG00] Benoît Bourbeau, Teodor Gabriel Crainic, and Bernard Gendron. Branch-and-bound parallelization strategies applied to a depot location and container fleet management problem. Parallel Computing, 26(1):27– 46, January 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/23/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/23/25/
article.pdf.

Balboni:1993:PST

[BCGS93] G. P. Balboni, G. P. Cabodi, S. Gai, and M. Sonza Reorda. A parallel system for test pattern generation. Parallel Computing, 19(2):177–185, February 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Berry:1991:SBC

[BCL91] M. Berry, G. Cybenko, and J. Larson. Scientific benchmark characterizations. Parallel Computing, 17(10–11):1173–1194, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Barton:1994:MPM

[BCM94] Eric Barton, James Cownie, and Moray McLaren. Message passing on the Meiko CS-2. Parallel Computing, 20(4):497–507, March 31, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=4&aid=866.

Bilbao-Castro:2007:POR

[BCMG^{+}07] J. R. Bilbao-Castro, A. Merino, I. García, J. M. Carazo, and J. J. Fernández. Parameter optimization in 3D reconstruction on a large scale grid. Parallel Computing, 33(4–5):250– 263, May 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Buenabad-Chavez:2003:VMD

[BCMSW03] Jorge Buenabad-Chávez, Henk L. Muller, Paul W. A. Stallard, and David H. D. Warren. Virtual memory on data diffusion architectures. Parallel Computing, 29(8):1021– 1052, August 2003. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Buenabad-Chavez:2004:DSD

[BCMSW04] Jorge Buenabad-Chávez, Henk L. Muller, Paul W. A. Stallard, and David H. D. Warren. The diffusion space of data diffusion architectures. Parallel Computing, 30(11):1169– 1193, November 2004. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Banicescu:2005:DIN

[BCPB05] Ioana Banicescu, Ricolindo L. Cariño, Jaderick P. Pabico, and Mahadevan Balasubramaniam. Design and implementation of a novel dynamic load balancing library for cluster computing. Parallel Computing, 31(7):736–756, July 2005. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bregier:2000:SLP

[BCR00] Frédéric Brégier, Marie-Christine Counilh, and Jean Roman. Scheduling loops with partial loopcarried dependencies. Parallel Computing, 26(13– 14):1789–1806, December 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/34/27/
abstract.html;   http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/34/27/
                                        article.pdf.

Benoit:2011:RAM

[BCRSR11] Anne Benoit, Henri Casanova, Veronika Rehn-Sonigo, and Yves Robert. Resource allocation for multiple concurrent in-network stream-processing applications. Parallel Computing, 37(8):331–348, August 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819110001298.

Bramas:2015:TDB

[BCS15] Berenger Bramas, Olivier Coulaud, and Guillaume Sylvand. Time-domain BEM for the wave equation on distributed-heterogeneous architectures: a blocking approach. Parallel Computing, 49(??):66–82, November 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115001088.

Bahi:2005:EAI

[BCVC05] Jacques M. Bahi, Sylvain Contassot-Vivier, and Raphaël Couturier. Evaluation of the asynchronous iterative algorithms in the context of distant heterogeneous clusters. Parallel Computing, 31(5):439– 461, May 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bouattane:2011:PMA

[BCYB11] Omar Bouattane, Bouchaib Cherradi, Mohamed Youssfi, and Mohamed O. Bensalah. Parallel c-means algorithm for image segmentation on a reconfigurable mesh computer. Parallel Computing, 37 (4–5):230–243, April/May 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911100024X.

Babolian:1989:PSF

[BD89a] E. Babolian and L. M.

Delves. Parallel solution of Fredholm integral equations. Parallel Computing, 12(1):95–106, October 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bonomo:1989:PIM

[BD89b] J. P. Bonomo and W. R.

Dyksen. Pipelined iterative methods for shared memory machines. Parallel Computing, 11(2):187– 199, ???? 4, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Blazewicz:1995:SCS

[BD95] J. Błażewicz and M. Drozdowski. Short Communication: Scheduling divisible jobs on hypercubes. Parallel Computing, 21 (12):1945–1956, December 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=12&aid=1025.

Berlinska:2010:HMR

[BD10] J. Berlińska and M. Drozdowski. Heuristics for multi-round divisible loads scheduling with limited memory. Parallel Computing, 36(4):199–211, April 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Berlinska:2018:CLB

[BD18] Joanna Berlińska and Maciej Drozdowski. Comparing load-balancing algorithms for MapReduce under Zipfian data skews. Parallel Computing, 72(??):14–28, February 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117302065.

Balewicz:2000:SPT

[BDF^{+}00] Jacek Bałewicz, Maciej Drozdowski, Piotr Formanowicz, Wiesław Kubiak, and Günter Schmidt. Scheduling preemptable tasks on parallel processors with limited availability. Parallel Computing, 26(9):1195–1211, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/30/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/30/30/
article.pdf.

Bekas:2010:SIP

[BDG^{+}10] Costas Bekas, Pasqua D’Ambra, Ananth Grama, Yousef Saad, and Petko Yanev. Special issue on parallel matrix algorithms and applications. Parallel Computing, 36(5–6):213– 214, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Borgeest:1995:TBP

[BDH95] Rolf Borgeest, Bernward Dimke, and Olav Hansen. A trace based performance evaluation tool for parallel real time systems. Parallel Computing, 21(4):551–564, April 1, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=4&aid=972.

Barros:1995:IMP

[BDI^{+}95] S. R. M. Barros, D. Dent, L. Isaksen, G. Robinson, G. Mozdzynski, and F. Wollenweber. The IFS model: a parallel production weather code. Parallel Computing, 21 (10):1621–1638, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=10&aid=1019.

Berry:1995:PAR

[BDK95] Michael W. Berry, Jack J. Dongarra, and Youngbae Kim. A parallel algorithm for the reduction of a nonsymmetric matrix to block upper-Hessenberg form. Parallel Computing, 21(8):1189–1211, August 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=8&aid=994.

Boldon:1996:MWD

[BDK96] Bruce Boldon, Narsingh Deo, and Nishit Kumar. Minimum-weight degree-constrained spanning tree problem: Heuristics and implementation on an SIMD parallel machine. Parallel Computing, 22(3):369–382, April 29, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=3&aid=1061.

Bampis:1998:OSG

[BDK98] E. Bampis, C. Delorme, and J.-C. König. Optimal schedules for d − D grid graphs with communication delays. Parallel Computing, 24(11):1653–1664, October 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/11/1324.pdf.

Boukerche:2006:ETD

[BDK06] Azzedine Boukerche, Caron Dzermajko, and Lu Kaiyuan. An enhancement towards dynamic grid-based DDM protocol for distributed simulation using multiple levels of data filtering. Parallel Computing, 32 (11–12):902–919, December 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Blazewicz:1999:DTS

[BDM99] Jacek Błażewicz, Maciej Drozdowski, and Mariusz Markiewicz. Divisible task scheduling — Concept and verification. Parallel Computing, 25(1):87–98, January 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/1/1380.pdf.

Benoit:2011:MWA

[BDNP11] Anne Benoit, Alexandru Dobrila, Jean-Marc Nicod, and Laurent Philippe. Mapping workflow applications with types on heterogeneous specialized platforms. Parallel Computing, 37(8):410–427, August 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819110001559.

Biswas:2017:HEC

[BDO17] Rupak Biswas, David Donofrio, and Leonid Oliker. High-end computing for next-generation scientific discovery. Parallel Computing, 64(??):1–2, May 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300492.

Bleck:1995:CDP

[BDOS95] Rainer Bleck, Sumner Dean, Matthew O’Keefe, and Aaron Sawdey. A comparison of data-parallel and message-passing versions of the Miami Isopycnic Coordinate Ocean Model (MICOM). Parallel Computing, 21(10): 1695–1720, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=10&aid=1022.

Bernaschi:2020:ABC

[BDP20] Massimo Bernaschi, Pasqua D’Ambra, and Dario Pasquini. AMG based on compatible weighted matching for GPUs. Parallel Computing, 92(??): Article 102599, April 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301905.

Bacci:1999:SHE

[BDPV99] B. Bacci, M. Danelutto, S. Pelagatti, and M. Vanneschi. SkIE: a heterogeneous environment for HPC applications. Parallel Computing, 25(13– 14):1827–1852, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/36/35/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/35/
article.pdf.

Boulet:1999:STH

[BDRV99] Pierre Boulet, Jack Dongarra, Yves Robert, and Frédéric Vivien. Static tiling for heterogeneous computing platforms. Parallel Computing, 25(5): 547–568, May 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cas/tree/store/parco/
sub/1999/25/5/1404.pdf.

Brewer:1988:TAA

[BDS88] O. Brewer, J. Dongarra, and D. Sorensen. Tools to aid in the analysis of memory access patterns for FORTRAN programs. Parallel Computing, 9(1):25–35, December 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Blazewicz:1994:SIM

[BDSd94] J. Błażewicz, M. Drozdowski, G. Schmidt, and D. de Werra. Scheduling independent multiprocessor tasks on a uniform kprocessor system. Parallel Computing, 20(1):15–28, January 16, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=1&aid=820.

Boulet:1998:LPA

[BDSV98] Pierre Boulet, Alain Darte, Georges-André Silber, and Frédéric Vivien. Loop parallelization algorithms: From parallelism extraction to code generation. Parallel Computing, 24 (3–4):421–444, May 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/3-4/1297.pdf.

Bekakos:1987:RFA

[BE87] M. P. Bekakos and D. J. Evans. A ‘rotating’ and ‘folding’ algorithm using a two-dimensional ‘systolic’ communication geometry. Parallel Computing, 4(2): 221–228, April 1987. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Berntsen:1988:PGA

[BE88] J. Berntsen and T. O. Espelid. A parallel global adaptive quadrature algorithm for hypercubes. Parallel Computing, 8(1–3): 313–323, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bekakos:1989:RPC

[BE89] M. P. Bekakos and D. J. Evans. Relative performance comparisons for the group explicit class of methods on MIMD, SIMD and pipelined vector computers. Parallel Computing, 10(3):357–364, May 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bekakos:1992:DAG

[BE92] M. P. Bekakos and D. J.

Evans. The double alternating group explicit method for nonlinear parabolic equations on MIMD parallel computers. Parallel Computing, 18 (8):879–895, August 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Barada:1993:MAR

[BE93a] H. Barada and A. El-Amawy. A methodology for algorithm regularization and mapping into time-optimal VLSI arrays. Parallel Computing, 19(1): 33–61, January 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bekakos:1993:PCO

[BE93b] M. P. Bekakos and D. J.

Evans. Parallel cyclic odd-even reduction algorithms for solving Toeplitz tridiagonal equations on MIMD computers. Parallel Computing, 19(5):545– 561, May 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Boffey:1993:IPC

[BE93c] T. B. Boffey and W. A. Essah. Implementing a parallel constrained ℓ_{1} approximation algorithm. Parallel Computing, 19(6):609– 620, June 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bahoshy:1995:GHE

[BE95] N. M. Bahoshy and D. J. Evans. A general harness for explicit parallel programming. Parallel Computing, 21(4):607–617, April 1, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=4&aid=955.

Boszormenyi:1997:MLH

[BE97] László Böszörményi and Karl-Heinz Eder. M3Set — a language for handling of distributed and persistent sets of objects. Parallel Computing, 22 (13):1913–1925, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=13&aid=1135.

Bebbington:1997:PIA

[Beb97] M. S. Bebbington. Parallel implementation of an aggregation/disaggregation method for evaluating quasi-stationary behavior in continuous-time Markov chains. Parallel Computing, 23(10): 1545–1559, November 7, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=10&aid=1225.

Beaumont:2016:ARC

[BEDdC16] Olivier Beaumont, Lionel Eyraud-Dubois, and Juan-Angel Lorenzo del Castillo. Analyzing real cluster data for formulating allocation algorithms in cloud platforms. Parallel Computing, 54(??):83– 96, May 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000939.

Bekakos:1995:NAF

[Bek95a] M. P. Bekakos. A notational approach to formulation of systolic array programs (short communication). Parallel Computing, 21(4):619–626, April 1, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=4&aid=956.

Bonniger:1995:CKP

[BEK95b] Tilmann Bönniger, Rüdiger Esser, and Dietrich Krekel. CM-5E, KSR2, Paragon XP/S: a comparative description of massively parallel computers. Parallel Computing, 21(2): 199–232, February 17, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=2&aid=941.

Benner:2011:MPA

[BEK^{+}11] Peter Benner, Pablo Ezzatti, Daniel Kressner, Enrique S. Quintana-Orti’, and Alfredo Remón. A mixed-precision algorithm for the solution of Lyapunov equations on hybrid CPU–GPU platforms. Parallel Computing, 37(8):439–450, August 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819110001560.

Benoit:2024:EPC

[Ben24]  Anne Benoit.  Edito-

rial for Parallel Computing. Parallel Computing, 119(??):??, February 2024. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819124000036.

Berntsen:1989:CEM

[Ber89] J. Berntsen. Communication efficient matrix multiplication on hypercubes. Parallel Computing, 12(3): 335–342, December 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bernholdt:2000:SCE

[Ber00] David E. Bernholdt. Scalability of correlated electronic structure calculations on parallel computers: a case study of the RI-MP2 method. Parallel Computing, 26(7–8):945– 963, July 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/29/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/29/30/
article.pdf.

Booth:2017:BPS

[BETR17] Joshua D. Booth, Nathan D. Ellingwood, Heidi K. Thornquist, and Sivasankaran Rajamanickam. Basker: Parallel sparse LU factorization utilizing hierarchical parallelism and data layouts. Parallel Computing, 68(??):17–31, October 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300868.

Bevan:1989:ERC

[Bev89] D. I. Bevan. An efficient reference counting solution to the distributed garbage collection problem. Parallel Computing, 9(2):179– 192, January 1989. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Balewicz:2000:NTS

[BEY00] Jacek Bałewicz, Klaus H. Ecker, and Tao Yang. New trends on scheduling in parallel and distributed systems. Parallel Computing, 26(9):1061–1063, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/30/23/
abstract.html;   http:

//www.elsevier.nl/gej-
ng/10/35/21/42/30/23/
article.pdf.

Bentley:1992:HCS

[BF92] M. Bentley and C. Froese Fischer. Hypercube conversion of serial codes for atomic structure calculations. Parallel Computing, 18(9):1023–1031, September 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Barlos:1995:LBM

[BF95] Fotis Barlos and Ophir Frieder. A load balanced multicomputer relational database system for highly skewed data. Parallel Computing, 21 (9):1451–1483, September 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=9&aid=1007.

Bisseling:2006:MSM

[BF06] Rob H. Bisseling and Ildikó Flesch. Mondriaan sparse matrix partitioning for attacking cryptosystems by a parallel block Lanczos algorithm — a case study. Parallel Computing, 32(7–8):551–567, September 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bernabe:2007:EIW

[BFG^{+}07] Gregorio Bernabé, Ricardo Fernández, Jose M. García, Manuel E. Acacio, and José González. An efficient implementation of a 3D wavelet transform based encoder on hyperthreading technology. Parallel Computing, 33(1):54– 72, February 2007. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bolten:2023:TGB

[BFH23] Matthias Bolten, Stephanie Friedhoff, and Jens Hahne. Task graph-based performance analysis of parallelin-time methods. Parallel Computing, 118(??):??, November 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781912300056X.

Basermann:2001:DMP

[BFL^{+}01] A. Basermann, J. Fingberg, G. Lonsdale, B. Maerten, and C. Walshaw. Dynamic multi-partitioning for parallel finite element applications. Parallel Computing, 27(7):869–881, June 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/31/24/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/31/24/
article.pdf.

Block:1990:BCS

[BFM90] U. Block, A. Frommer, and G. Mayer. Block colouring schemes for the SOR method on local memory parallel computers. Parallel Computing, 14(1):61– 75, May 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Buoni:1993:ADS

[BFR93] John J. Buoni, Paul A.

Farrell, and Arden Ruttan. Algorithms for LU decomposition on a shared memory multiprocessor. Parallel Computing, 19(8):925– 937, August 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Boton-Fernandez:2014:SAG

[BFVRC14] María Botón-Fernández, Miguel A. Vega-Rodríguez, and Francisco Prieto Castrillo. Self-adaptivity for grid applications. An efficient resources selection model based on evolutionary computation algorithms. Parallel Computing, 40(8):345–361, August 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000684.

Baker:2006:APA

[BFY06] A. H. Baker, R. D. Falgout, and U. M. Yang. An assumed partition algorithm for determining processor inter-communication. Parallel Computing, 32(5–6): 394–414, June 2006. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

BacchelliMontefusco:1991:DDM

[BG91a] L. Bacchelli Montefusco and C. Guerrini. A domain decomposition method for scattered data approximation on a distributed memory multiprocessor. Parallel Computing, 17(2–3): 253–263, June 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bader:1991:PTN

[BG91b] G. Bader and E. Gehrke. On the performance of transputer networks for solving linear systems of equations. Parallel Computing, 17(12):1397–1407, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Beraldi:1997:PAI

[BG97] P. Beraldi and F. Guerriero. Parallel asynchronous implementation of the ϵ-relaxation method for the linear minimum cost flow problem. Parallel Computing, 23(8):1021– 1044, July 25, 1997. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cgi-bin/cas/tree/
store/parco/cas_sub/
browse/browse.cgi?year=
1997&volume=23&issue=
8&aid=1161.

Brunett:1998:LSM

[BG98] Sharon Brunett and Thomas Gottschalk. A large-scale metacomputing framework for the ModSAF realtime simulation. Parallel Computing, 24(12– 13):1873–1900, November 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/12-13/1359.pdf.

Bekas:2001:CPP

[BG01] C. Bekas and E. Gallopoulos. Cobra: Parallel path following for computing the matrix pseudospectrum. Parallel Computing, 27(14):1879–1896, December 31, 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/44/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/44/29/
article.pdf.

Bekas:2002:PCP

[BG02] C. Bekas and E. Gallopoulos. Parallel computation of pseudospectra by fast descent. Parallel Computing, 28(2):223–242, February 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/33/31/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/33/31/
main.pdf.

Berten:2007:BSC

[BG07] Vandy Berten and Bruno Gaujal. Brokering strategies in computational grids using stochastic prediction models. Parallel Computing, 33(4–5):238–249, May 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bellas:2021:HHFa

[BG21a] Christos Bellas and Anastasios Gounaris. HySet: a hybrid framework for exact set similarity join using a GPU. Parallel Computing, 104–105(??): ??, July 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000442.

Bellas:2021:HHFb

[BG21b] Christos Bellas and Anastasios Gounaris. HySet: a hybrid framework for exact set similarity join using a GPU. Parallel Computing, 104–105(??): ??, July 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000442.

Buluc:2010:SPP

[BGB10] Aydın Buluç, John R.

Gilbert, and Ceren Budak. Solving path problems on the GPU. Parallel Computing, 36(5–6): 241–253, June 2010. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Barrows:1999:ESD

[BGG^{+}99] M. E. Barrows, D. E.

Gregory, L. Gao, A. L. Rosenberg, and P. R. Cohen. An empirical study of dynamic scheduling on rings of processors. Parallel Computing, 25(9):1063–1079, September 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/25/17/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/25/17/
article.pdf.

Brune:1998:SRS

[BGK^{+}98] Matthias Brune, Jorn Gehring, Axel Keller, Burkhard Monien, Friedhelm Ramme, and Alexander Reinefeld. Specifying resources and services in metacomputing environments. Parallel Computing, 24(12– 13):1751–1776, November 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/12-13/1353.pdf.

Brint:1988:CGM

[BGL^{+}88] A. T. Brint, V. J. Gillet, M. F. Lynch, P. Willett, G. A. Manson, and G. A. Wilson. Chemical graph matching using transputer networks. Parallel Computing, 8(1– 3):295–300, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Butler:2001:CIP

[BGL01] Ralph Butler, William Gropp, and Ewing Lusk. Components and interfaces of a process management system for parallel programs. Parallel Computing, 27(11):1417–1429, October 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/41/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/41/29/
article.pdf.

Blanco:2004:PPP

[BGL^{+}04] V. Blanco, J. A. González, C. León, C. Rodríguez, G. Rodríguez, and M. Printista. Predicting the performance of parallel programs. Parallel Computing, 30(3):337–356, March 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bortfeldt:2003:PTS

[BGM03] A. Bortfeldt, H. Gehring, and D. Mack. A parallel tabu search algorithm for solving the container loading problem. Parallel Computing, 29(5):641– 662, May 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Beraldi:2000:PAS

[BGMT00] P. Beraldi, L. Grandinetti, R. Musmanno, and C. Triki. Parallel algorithms to solve two-stage stochastic linear programs with robustness constraints. Parallel Computing, 26(13– 14):1889–1908, December 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/34/31/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/34/31/
article.pdf.

Bergadano:1997:SAI

[BGP^{+}97] F. Bergadano, A. Giallombardo, A. Puliafito, G. Ruffo, and L. Vita. Security agents for information retrieval in distributed systems. Parallel Computing, 22(13): 1719–1731, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.

cgi?year=1997&volume=
22&issue=13&aid=1124.

Brandes:2000:SCD

[BGR00] Thomas Brandes and Cécile Germain-Renaud. A schedule cache for data parallel unstructured computations. Parallel Computing, 26(13– 14):1807–1823, December 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/34/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/34/28/
article.pdf.

Bunnin:2000:DHP

[BGRD00] F. O. Bunnin, Y. Guo, Y. Ren, and J. Darlington. Design of high performance financial modelling environment. Parallel Computing, 26(5):601– 622, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/27/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/27/28/
article.pdf.

Bekas:2014:PMA

[BGSS14] Costas Bekas, Ananth Grama, Yousef Saad, and Olaf Schenk. Parallel matrix algorithms. Parallel Computing, 40(7):159–160, July 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000647.

Barve:1997:SRM

[BGV97] Rakesh D. Barve, Edward F. Grove, and Jeffrey Scott Vitter. Simple randomized mergesort on parallel disks. Parallel Computing, 23 (4–5):601–631, May 23, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=4-5&aid=1173.

Balos:2021:EGA

[BGWR21] Cody J. Balos, David J. Gardner, Carol S. Woodward, and Daniel R. Reynolds. Enabling GPU accelerated computing in the SUNDIALS time integration library. Parallel Computing, 108(??):??, December 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819121000831.

Bangay:1997:BSG

[BGWW97] Shaun Bangay, James Gain, Greg Watkins, and Kevan Watkins. Building the second generation of parallel/distributed virtual reality systems. Parallel Computing, 23 (7):991–1000, July 14, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=7&aid=1198.

Burg:1992:ICS

[BH92] H. C. Burg and J. Helin.

1991 International Conference on Supercomputing. Parallel Computing, 18(4): 467–??, April 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bowler:1999:PCQ

[BH99] Kenneth C. Bowler and Anthony J. G. Hey. Parallel computing and quantum chromodynamics. Parallel Computing, 25(13–14):2111–2134, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-

                                        ng/10/35/21/32/36/45/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/36/45/
                                        article.pdf.

Balaji:2014:SIP

[BH14] Pavan Balaji and Zhiyi Huang. Special issue on programming models and applications for multicores and manycores — Guest Editors’ introduction. Parallel Computing, 40(2):33–34, February 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113001476.

Ncir:2021:PSD

[BHB21] Chiheb-Eddine Ben Ncir, Abdallah Hamza, and Waad Bouaguel. Parallel and scalable Dunn index for the validation of big data clusters. Parallel Computing, 102 (??):Article 102751, May 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000119.

Barr:1989:MST

[BHK89] R. S. Barr, R. V. Helgaon, and J. L. Kennington. Minimal spanning trees: an empirical investigation of parallel algorithms. Parallel Computing, 12(1):45–52, October 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Biswas:2000:GPP

[BHK00] Rupak Biswas, Bruce Hendrickson, and George Karypis. Graph partitioning and parallel computing. Parallel Computing, 26(12):1515–1517, November 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/33/22/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/33/22/
article.pdf.

Beecroft:1994:MCI

[BHM94] Jon Beecroft, Mark Homewood, and Moray McLaren. Meiko CS-2 interconnect Elan-Elite design. Parallel Computing, 20(10– 11):1627–1638, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=918.

Bendtsen:1995:IDM

[BHM^{+}95] Claus Bendtsen, Per Christian Hansen, Kaj Madsen, Hans Bruun Nielsen, and Mustafa Pinar. Implementation of QR up- and downdating on a massively parallel computer. Parallel Computing, 21(1):49–61, January 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=1&aid=936.

Benoit:2009:CAF

[BHR09] Anne Benoit, Mourad Hakem, and Yves Robert. Contention awareness and fault-tolerant scheduling for precedence constrained tasks in heterogeneous systems. Parallel Computing, 35(2):83–108, February 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bohm:1999:PPI

[BHS99] A. P. Willem Böhm, Jeffrey P. Hammes, and Sumit S. Sur. On the performance of pure and impure parallel functional programs. Parallel Computing, 25(13– 14):1723–1740, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/36/32/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/32/
article.pdf.

Botta:2007:RBC

[BI07] N. Botta and C. Ionescu.

Relation-based computations in a monadic BSP model. Parallel Computing, 33(12):795–821, December 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bishop:1984:UVP

[Bis84] M. Bishop. The Ultracomputer as a vehicle for polymer simulations. Parallel Computing, 1(2):165–174, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bischof:1989:CSV

[Bis89] Christian H. Bischof. Computing the singular value decomposition on a distributed system of vector processors. Parallel Computing, 11(2):171–186, ???? 4, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Badr:2018:HLM

[BJ18] Mario Badr and Natalie Enright Jerger. A high-level model for exploring multi-core architectures. Parallel Computing, 80(??):23–35, December 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911830125X.

Biswas:2017:NPQ

[BJK^{+}17] Rupak Biswas, Zhang Jiang, Kostya Kechezhi, Sergey Knysh, Salvatore Mandrà, Bryan O’Gorman, Alejandro Perdomo-Ortiz, Andre Petukhov, John Realpe-Gómez, Eleanor Rieffel, Davide Venturelli, Fedir Vasko, and Zhihui Wang. A NASA perspective on quantum computing: Opportunities and challenges. Parallel Computing, 64(??):81–98, May 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116301326.

Bjorstad:1987:LSS

[Bjø87] Petter E. Bjørstad. A large scale, sparse, secondary storage, direct linear equation solver for structural analysis and its implementation on vector and parallel architectures. Parallel Computing, 5(1–2):3– 12, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Bohr:1989:PCS

[BJP^{+}89] H. Bohr, K. S. Jensen, T. Petersen, B. Rathjen, E. Mosekilde, and N.-H. Holstein-Rathlou. Parallel computer simulation of nearest-neighbour interaction in a system of nephrons. Parallel Computing, 12(1):113–120, October 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Boukerche:2004:AIB

[BJSN04] Azzedine Boukerche, Kathia Regina Lemos Jucá, João Bosco Sobral, and Mirela Sechi Moretti Annoni Notare. An artificial immune based intrusion detection model for computer and telecommunication systems. Parallel Computing, 30(5–6): 629–646, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bui:2016:ISD

[BJV^{+}16] Huy Bui, Eun-Sung Jung, Venkatram Vishwanath, Andrew Johnson, Jason Leigh, and Michael E. Papka. Improving sparse data movement performance using multiple paths on the Blue Gene/Q supercomputer. Parallel Computing, 51(??):3–16, January 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115001167.

Bonorden:2003:PUB

[BJvOR03] Olaf Bonorden, Ben Juurlink, Ingo von Otte, and Ingo Rieping. The Paderborn University BSP (PUB) library. Parallel Computing, 29(2):187–207, February 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Barros:1994:PGS

[BK94] Saulo R. M. Barros and Tuomo Kauranne. On the parallelization of global spectral weather models. Parallel Computing, 20 (9):1335–1356, September 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/

cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=9&aid=894.

Buchholz:1997:SRB

[BK97] Thomas Buchholz and Martin Kutrib. Some relations between massively parallel arrays. Parallel Computing, 23(11): 1643–1662, December 1, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=11&aid=1233.

Byun:2007:DDS

[BK07] Eun-Kyu Byun and Jin-Soo Kim. DynaGrid: a dynamic service deployment and resource migration framework for WSRF-compliant applications. Parallel Computing, 33(4–5):328–338, May 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bhardwaj:2024:FLB

[BKK24] Sanjay Bhardwaj, Da-Hye Kim, and Dong-Seong Kim. Federated learning based modulation classification for multipath channels. Parallel Computing, 120(??):??, June 2024. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819124000218.

Brewer:2019:SCC

[BKL^{+}19] Nicole Brewer, HyeJin Kim, Claudia Li, Heidi Anderson, Jessica Lanum, Jia Cheoh, Betsy Hillery, and Trinity Overmyer. Student cluster competition 2018, team Ada Six of Purdue University: Reproducing Extreme Scale Multi-Physics Simulations of Tsunamigenic 2004 Sumatra Megathrust Earthquake on Intel Skylake architecture. Parallel Computing, 90(??): Article 102565, December 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301565.

Bansal:2006:ITS

[BKS06] Savina Bansal, Padam Kumar, and Kuldip Singh. An improved two-step algorithm for task and data parallel scheduling in distributed memory machines. Parallel Computing, 32(10):759–774, November 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bazinska:2019:SCC

[BKS19] Julia Bazińska, Maciej Korpalski, and Maciej Szpindler. Student Cluster Competition 2018, Team University of Warsaw, University of Wroclaw, Warsaw University of Technology: Reproducing performance of a multiphysics simulations of the Tsunamigenic 2004 Sumatra megathrust earthquake on the Intel Skylake architecture. Parallel Computing, 90(??):Article 102567, December 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301589.

Bampis:1991:ICC

[BKT91] E. Bampis, J. C. Konig, and D. Trystram. Impact of communications on the complexity of the parallel Gaussian elimination. Parallel Computing, 17(1):55– 61, April 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Boissin:1993:PSA

[BL93a] Nicolas Boissin and Jean-Luc Lutton. A parallel simulated annealing algorithm. Parallel Computing, 19(8):859–872, August 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bossomaier:1993:PCH

[BL93b] Terry Bossomaier and Adrian Loeff. Parallel computation of the Hausdorff distance between images. Parallel Computing, 19(10):1129–1140, October 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Beazley:1994:MPM

[BL94a] D. M. Beazley and P. S. Lomdahl. Message-passing multi-cell molecular dynamics on the Connection Machine 5. Parallel Computing, 20(2): 173–195, February 24, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=2&aid=826.

Butler:1994:MMC

[BL94b] Ralph M. Butler and Ewing L. Lusk. Monitors, messages, and clusters: The p4 parallel programming system. Parallel Computing, 20(4):547–564, March 31, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=4&aid=864.

Barth:1999:SMS

[BL99] D. Barth and C. Laforest. Scattering and multiscattering in trees and meshes, with local routing and without buffering. Parallel Computing, 25(8):1035–1057, September 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/24/20/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/24/20/
article.pdf.

Brink:2021:EAPa

[BLCR21a] Stephanie Brink, Matthew Larsen, Hank Childs, and Barry Rountree. Evaluating adaptive and predictive power management strategies for optimizing visualization performance on supercomputers. Parallel Computing, 104–105(??): ??, July 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0167819121000399.

Brink:2021:EAPb

[BLCR21b] Stephanie Brink, Matthew Larsen, Hank Childs, and Barry Rountree. Evaluating adaptive and predictive power management strategies for optimizing visualization performance on supercomputers. Parallel Computing, 104–105(??): ??, July 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000399.

Bobrowicz:1984:VMC

[BLFT84] F. W. Bobrowicz, J. E. Lynch, K. J. Fisher, and J. E. Tabor. Vectorized Monte Carlo photon transport. Parallel Computing, 1(3–4):295–305, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Buttari:2009:CPT

[BLKD09] Alfredo Buttari, Julien Langou, Jakub Kurzak, and Jack Dongarra. A class of parallel tiled linear algebra algorithms for multicore architectures. Parallel Computing, 35(1):38– 53, January 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Blomvall:2003:MSP

[Blo03] Jörgen Blomvall. A multistage stochastic programming algorithm suitable for parallel computing. Parallel Computing, 29(4): 431–445, April 2003. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Beaumont:2003:SDW

[BLR03] Olivier Beaumont, Arnaud Legrand, and Yves Robert. Scheduling divisible workloads on heterogeneous platforms. Parallel Computing, 29(9):1121–1152, September 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Beaumont:2002:DLA

[BLRR02] Olivier Beaumont, Arnaud Legrand, Fabrice Rastello, and Yves Robert. Dense linear algebra kernels on heterogeneous platforms: Redistribution issues. Parallel Computing, 28(2):155–185, February 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/33/28/
abstract.html;   http:
//www.elsevier.nl/gej-

ng/10/35/21/60/33/28/
main.pdf.

Bielich:2022:LSG

[BLT^{+}22] Daniel Bielich, Julien Langou, Stephen Thomas, Kasia Świrydowicz, Ichitaro Yamazaki, and Erik G. Boman. Low-synch Gram– Schmidt with delayed reorthogonalization for Krylov solvers. Parallel Computing, 112(??):??, September 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000394.

Brugnano:1990:VSB

[BM90] L. Brugnano and M. Marrone. Vectorization of some block preconditioned conjugate gradient methods. Parallel Computing, 14(2):191–198, June 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bahi:1993:CMM

[BM93] M. Bahi and J. C. Miellou. Contractive mappings with maximum norms: comparison of constants of contraction and application to asynchronous iterations. Parallel Computing, 19(5):511–523, May 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bandini:2001:PPS

[BM01a] Stefania Bandini and Massimiliano Magagnini. Parallel processing simulation of dynamic properties of filled rubber compounds based on cellular automata. Parallel Computing, 27(5):643–661, April 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/29/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/29/30/
article.pdf.

Boleng:2001:LBP

[BM01b] Jeff Boleng and Manavendra Misra. Load balanced parallel QR decomposition on shared memory multiprocessors. Parallel Computing, 27(10):1321–1345, September 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/40/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/40/29/
article.pdf.

Biancardi:2002:EDP

[BM02] A. Biancardi and A. Mérigot.

Extending the data parallel paradigm with datadependent operators. Parallel Computing, 28(7–8): 995–1021, August 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.
                                        com/gej-ng/10/35/21/
                                        60/58/31/abstract.html.

Bader:2008:GTA

[BM08] David A. Bader and Kamesh Madduri. A graph-theoretic analysis of the human proteininteraction network using multicore parallel algorithms. Parallel Computing, 34(11):627–639, November 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Behara:2009:PFE

[BM09] Suresh Behara and Sanjay Mittal. Parallel finite element computation of incompressible flows. Parallel Computing, 35(4):195– 212, April 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Barbosa:2011:DSB

[BM11] Jorge G. Barbosa and Belmiro Moreira. Dynamic scheduling of a batch of parallel task jobs on heterogeneous clusters. Parallel Computing, 37(8):428– 438, August 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000020.

Barbosa:2006:CCH

[BMA06] Valmir C. Barbosa, Fernando M. N. Miranda, and Matheus C. M. Agostini. Cell-centric heuristics for the classification of cellular automata. Parallel Computing, 32(1):44–66, January 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bhandarkar:1998:PCC

[BMCA98] Suchendra M. Bhandarkar, Salem Machaka, Sridhar Chirravuri, and Jonathan Arnold. Parallel computing for chromosome reconstruction via ordering of DNA sequences. Parallel Computing, 24 (8):1177–1204, August 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/8/1330.pdf.

Buntinas:2007:IES

[BMG07] Darius Buntinas, Guillaume Mercier, and William Gropp. Implementation and evaluation of sharedmemory communication and synchronization operations in MPICH2 using the Nemesis communication subsystem. Parallel Computing, 33(9):634–644, September 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Badia:2015:SIB

[BMP15] Santiago Badia, Alberto F. Martín, and Javier Principe. On the scalability of inexact balancing domain decomposition by constraints with overlapped coarse/ fine corrections. Parallel Computing, 50(??):1–24, December 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001180.

Baillie:1997:RWM

[BMS97] C. Baillie, J. Michalakes, and R. Skålin. Regional weather modeling on parallel computers. Parallel Computing, 23(14): 2135–2142, December 17, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=14&aid=1262.

Bandini:2001:CAT

[BMS01a] S. Bandini, G. Mauri, and R. Serra. Cellular automata: From a theoretical parallel computational model to its application to complex systems. Parallel Computing, 27(5):539– 553, April 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/29/24/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/29/24/
article.pdf.

Bandini:2001:CAM

[BMS01b] Stefania Bandini, Giancarlo Mauri, and Roberto Serra. Cellular automata: From modeling to applications. Parallel Computing, 27(5):537–538, April 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/29/23/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/29/23/
article.pdf.

Bermond:1992:BWM

[BMT92] J.-C. Bermond, P. Michallon, and D. Trystram. Broadcasting in wraparound meshes with parallel monodirectional links. Parallel Computing, 18(6):639–648, June 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Basoglu:1998:CRM

[BMYK98] Chris Basoglu, Ravi Managuli, George York, and Yongmin Kim. Computing requirements of modern medical diagnostic ultrasound machines. Parallel Computing, 24(9– 10):1407–1431, September 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/9-10/1341.pdf.

Balsara:2001:HPS

[BN01] Dinshaw S. Balsara and Charles D. Norton. Highly parallel structured adaptive mesh refinement using parallel language-based approaches. Parallel Computing, 27(1–2):37–70, January 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/25/24/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/25/24/
                                        article.pdf.

Bromberger:2015:CHS

[BNK15] Michael Bromberger, Fabian Nowak, and Wolfgang Karl. Combined hardwaresoftware multi-parallel prefiltering on the Convey HC-1 for fast homology detection. Parallel Computing, 42(??):4–17, February 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001173.

Blagojevic:2007:RSD

[BNS^{+}07] Filip Blagojevic, Dimitrios S. Nikolopoulos, Alexandros Stamatakis, Christos D. Antonopoulos, and Matthew Curtis-Maury. Runtime scheduling of dynamic parallelism on accelerator-based multi-core systems. Parallel Computing, 33(10–11): 700–719, November 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bar-On:1991:ELT

[BO91] Ilan Bar-On. Efficient logarithmic time parallel algorithms for the Cholesky decomposition and Gram– Schmidt process. Parallel Computing, 17(4–5): 409–417, July 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Becka:2003:VBF

[BO03] Martin Bečka and Gabriel Okša. On variable blocking factor in a parallel dynamic block-Jacobi SVD algorithm. Parallel Computing, 29(9):1153–1174, September 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Boglaev:1992:EDL

[Bog92] Y. P. Boglaev. Exact dynamic load balancing of MIMD architectures with linear programming algorithms. Parallel Computing, 18(6):615–623, June 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Buluc:2015:SIG

[BOG15] Aydin Buluç, Leonid Oliker, and John Gilbert. Special issue “Graph analysis for scientific discovery”. Parallel Computing, 47(??):1–2, August 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000927.

Boley:1986:SGE

[Bol86]  Daniel Boley.  Solving

the generalized eigenvalue problem on a synchronous linear processor array. Parallel Computing, 3(2):153– 166, May 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bondeli:1991:DCP

[Bon91] S. Bondeli. Divide and conquer: a parallel algorithm for the solution of a tridiagonal linear system of equations. Parallel Computing, 17(4–5): 419–434, July 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Borrill:2009:HGF

[BOS^{+}09] Julian Borrill, Leonid Oliker, John Shalf, Hongzhang Shan, and Andrew Uselton. HPC global file system performance analysis using a scientificapplication derived benchmark. Parallel Computing, 35(6):358–373, June 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Becka:2002:DOP

[BOV02] M. Bečka, G. Okša, and M. Vajteršic. Dynamic ordering for a parallel block-Jacobi SVD algorithm. Parallel Computing, 28(2):243–262, February 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/33/32/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/60/33/32/
                                        main.pdf.

Biswas:2009:RTA

[BOV09] Rupak Biswas, Leonid Oliker, and Jeffrey Vetter. Revolutionary technologies for acceleration of emerging petascale applications. Parallel Computing, 35(3):117–118, March 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Becka:2010:IQP

[BOVG10] Martin Bečka, Gabriel Okša, Marián Vajteršic, and Laura Grigori. On iterative QR pre-processing in the parallel block-jacobi SVD algorithm. Parallel Computing, 36(5–6):297– 307, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Boxer:1992:FCR

[Box92] Laurence Boxer. Finding congruent regions in parallel. Parallel Computing, 18(7):807–810, July 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Brass:1986:TTD

[BP86] A. Brass and G. S. Pawley. Two- and threedimensional FFTs on highly parallel computers. Parallel Computing, 3(2): 167–184, May 1986. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Baillie:1989:CCD

[BP89] C. F. Baillie and G. S.

Pawley. A comparison of the CM with the DAP for lattice gauge theory. Parallel Computing, 12(2): 209–220, November 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Boreddy:1990:PTA

[BP90] J. Boreddy and A. Paulraj.

On the performance of transputer arrays for dense linear systems. Parallel Computing, 15(1–3):107– 117, September 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Beidas:1993:CAA

[BP93] Bassem F. Beidas and George P. Papavassilopoulos. Convergence analysis of asynchronous linear iterations with stochastic delays. Parallel Computing, 19(3):281–302, March 1993. CODEN PACOEJ.

ISSN 0167-8191 (print), 1872-7336 (electronic).

Beidas:1995:DAA

[BP95] Bassem F. Beidas and George P. Papavassilopoulos. Distributed asynchronous algorithms with stochastic delays for constrained optimization problems with conditions of time drift. Parallel Computing, 21(9):1431–1450, September 12, 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cgi-bin/cas/tree/
store/parco/cas_sub/
browse/browse.cgi?year=
1995&volume=21&issue=
9&aid=984.

Billard:1997:LBA

[BP97] Edward A. Billard and Joseph C. Pasquale. Load balancing to adjust for proximity in some network topologies. Parallel Computing, 22(14):2007–2023, March 24, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=14&aid=1139.

Bevilacqua:2000:PIR

[BP00] A. Bevilacqua and E. Loli Piccolomini. Parallel image restoration on parallel and distributed computers. Parallel Computing, 26(4):495–506, March 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/26/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/26/29/
article.pdf.

Brill:2002:PIB

[BP02] Stephen H. Brill and George F. Pinder. Parallel implementation of the Bi-CGSTAB method with block red-black Gauss– Seidel preconditioner applied to the Hermite collocation discretization of partial differential equations. Parallel Computing, 28(3):399–414, March 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/34/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/34/28/
main.pdf.

Brzenski:2021:ILGa

[BPC21a] Jared Brzenski, Christopher Paolini, and Jose E. Castillo. Improving the I/O of large geophysical models using PnetCDF and BeeGFS. Parallel Computing, 104–105(??): ??, July 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000429.

Brzenski:2021:ILGb

[BPC21b] Jared Brzenski, Christopher Paolini, and Jose E. Castillo. Improving the I/O of large geophysical models using PnetCDF and BeeGFS. Parallel Computing, 104–105(??): ??, July 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000429.

Balduccini:2005:IPE

[BPEL05] Marcello Balduccini, Enrico Pontelli, Omar Elkhatib, and Hung Le. Issues in parallel execution of nonmonotonic reasoning systems. Parallel Computing, 31(6):608–647, June 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bouhrour:2022:TLC

[BPJ22] Stephane Bouhrour, Thibaut Pepin, and Julien Jaeger. Towards leveraging collective performance with the support of MPI 4.0 features in MPC. Parallel Computing, 109(??):??, March 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121001034.

Bielecki:2012:FSS

[BPK12] Wlodzimierz Bielecki, Marek Palkowski, and Tomasz Klimek. Free scheduling for statement instances of parameterized arbitrarily nested affine loops. Parallel Computing, 38(9):518–532, September 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000336.

Bouziane:2010:ESC

[BPP10] Hinde Lilia Bouziane, Christian Pérez, and Thierry Priol. Extending software component models with the masterworker paradigm. Parallel Computing, 36(2–3): 86–103, February/March 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bergamaschi:2001:PPS

[BPS01] Luca Bergamaschi, Giorgio Pini, and Flavio Sartoretto. Parallel preconditioning of a sparse eigensolver. Parallel Computing, 27(7):963–976, June 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/47/31/30/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/31/30/
                                        article.pdf.

Buzbee:1984:CFL

[BR84] B. L. Buzbee and H. J. Raveché. Conference on forefronts of largescale computational problems. Parallel Computing, 1(3–4):307–315, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781918490245X.

Benaini:1989:EFS

[BR89] A. Benaini and Y. Robert. An even faster systolic array for matrix multiplication. Parallel Computing, 12(2):249–254, November 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Benaini:1990:SMS

[BR90] Abdelhamid Benaini and Yves Robert. Spacetimeminimal systolic arrays for Gaussian elimination and the algebraic path problem. Parallel Computing, 15(1–3):211–225, September 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Boeres:1999:VCM

[BR99] Cristina Boeres and Vinod E. F. Rebello. A versatile cost modelling approach for multicomputer task scheduling. Parallel Computing, 25(1):63–86, January 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/1/1379.pdf.

Brandes:1988:DDK

[Bra88] Thomas Brandes. Determination of dependencies in a knowledge-based parallelization tool. Parallel Computing, 8(1–3):111– 119, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the International Conference on Vector and Parallel Processors in Computational Science, III (Liverpool, 1987).

Bomans:1990:AGM

[BRH90] L. Bomans, D. Roose, and R. Hempel. The argonne/GMD macros in FORTRAN for portable parallel programming and their implementation on the Intel iPSC/2. Parallel Computing, 15(1–3): 119–132, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Briscolini:1995:PID

[Bri95] M. Briscolini. A parallel implementation of a 3-D pseudospectral based code on the IBM 9076 scalable POWER parallel system. Parallel Computing, 21(11):1849–1862, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=11&aid=1027.

Browne:1986:FFA

[Bro86] J. C. Browne. Framework for formulation and analysis of parallel computation structures. Parallel Computing, 3(1):1–9, March 1986. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Brooks:1987:BPM

[Bro87] E. D. Brooks, III. A butterfly processor-memory interconnection for a vector processing environment. Parallel Computing, 4(1):103–110, February 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Brooks:1988:IAC

[Bro88a] E. D. Brooks, III. The indirect k-ary n-cube for a vector processing environment. Parallel Computing, 6(3):339–348, March 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Brooks:1988:SMH

[Bro88b] E. D. Brooks, III. The shared memory hypercube. Parallel Computing, 6(2): 235–245, February 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Brochard:1989:ESP

[Bro89] L. Brochard. Efficiency of some parallel numerical algorithms on distributed systems. Parallel Computing, 12(1):21–44, October 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Brodal:1999:PQP

[Bro99]  G. S. Brodal.  Pri-

ority queues on parallel machines. Parallel Computing, 25(8): 987–1011, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/24/19/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/24/19/
article.pdf.

Benoit:2011:SWS

[BRRV11] Anne Benoit, Yves Robert, Arnold Rosenberg, and Frédéric Vivien. Static worksharing strategies for heterogeneous computers with unrecoverable interruptions. Parallel Computing, 37(8):365–378, August 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819110000372.

Basermann:1997:PCM

[BRS97] A. Basermann, B. Reichel, and C. Schelthoff. Preconditioned CG methods for sparse matrices on massively parallel machines. Parallel Computing, 23(3):381–398, May 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/

cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=3&aid=1163.

Brazile:2008:CVG

[BRS^{+}08] Jason Brazile, Rudolf Richter, Daniel Schläpfer, Michael E. Schaepman, and Klaus I. Itten. Cluster versus grid for operational generation of ATCOR’s MODTRAN-based look up tables. Parallel Computing, 34(1):32–46, January 2008. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bonisch:2017:HLH

[BRS^{+}17] Thomas Bönisch, Michael Resch, Thomas Schwitalla, Matthias Meinke, Volker Wulfmeyer, and Kirsten Warrach-Sagi. Hazel Hen — leading HPC technology and its impact on science in Germany and Europe. Parallel Computing, 64(??):3–11, May 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300194.

Brugnano:1991:PST

[Bru91] L. Brugnano. A parallel solver for tridiagonal linear systems for distributed memory parallel computers. Parallel Computing, 17(9):1017–1023, November 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Becker:2009:STS

[BRWL09] Daniel Becker, Rolf Rabenseifner, Felix Wolf, and John C. Linford. Scalable timestamp synchronization for event traces of message-passing applications. Parallel Computing, 35(12):595–607, December 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Brorsson:1996:CMS

[BS96] Mats Brorsson and Per Stenström. Characterising and modelling shared memory accesses in multiprocessor programs. Parallel Computing, 22(6): 869–893, September 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=6&aid=1076.

Bartha:1997:PSL

[BS97] Tamás Bartha and Endre Selényi. Probabilistic system-level fault diagnostic algorithms for multiprocessors. Parallel Computing, 22(13): 1807–1821, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=13&aid=1129.

Baker:2000:IQC

[BS00a] Jon Baker and Matt Shirel.

Ab initio quantum chemistry on PC-based parallel supercomputers. Parallel Computing, 26(7–8):1011– 1024, July 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/29/34/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/29/34/
article.pdf.

Burton:2000:MRP

[BS00b] F. Warren Burton and David J. Simpson. Memory requirements for parallel programs. Parallel Computing, 26(13– 14):1739–1763, December 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/34/25/
abstract.html;   http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/34/25/
                                        article.pdf.

Blikberg:2005:LBO

[BS05] R. Blikberg and T. Sørevik. Load balancing and OpenMP implementation of nested parallelism. Parallel Computing, 31(10–12):984–998, October/December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bogle:2022:PGC

[BSB^{+}22] Ian Bogle, George M. Slota, Erik G. Boman, Karen D. Devine, and Sivasankaran Rajamanickam. Parallel graph coloring algorithms for distributed GPU environments. Parallel Computing, 110(??):??, May 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819122000047.

Boeres:2011:EWB

[BSD11] Cristina Boeres, Idalmis Milián Sardiña, and Lúcia M. A. Drummond. An efficient weighted bi-objective scheduling algorithm for heterogeneous systems. Parallel Computing, 37 (8):349–364, August 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819110001407.

Babb:1988:PSD

[BSE88] R. G. Babb, II, L. Storc, and P. G. Eltgroth. Parallelization schemes for 2-D hydrodynamics codes using the independent time step method. Parallel Computing, 8(1–3):85–89, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Babb:1988:DPM

[BSH88] R. G. Babb, L. Storc, and R. Hiromoto. Developing a parallel Monte Carlo transport algorithm using largegrain data flow. Parallel Computing, 7(2):187–198, June 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Blochinger:2003:PPS

[BSK03] Wolfgang Blochinger, Carsten Sinz, and Wolfgang Küchlin. Parallel propositional satisfiability checking with distributed dynamic learning. Parallel Computing, 29(7):969–994, July 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Briggs:1988:FPS

[BT88] William L. Briggs and Thomas Turnbull. Fast Poisson solvers for MIMD computers. Parallel Computing, 6(3):265–274, March 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Benmaiza:1999:OAB

[BT99] Mohamed Benmaiza and Abderezak Touzene. Oneto-all broadcast algorithm for constant degree 4 Cayley graphs. Parallel Computing, 25(3):249–264, March 22, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/3/1385.pdf.

Boukerche:2001:LVG

[BT01] Azzedine Boukerche and Carl Tropper. Local versus global lookahead in conservative parallel simulations. Parallel Computing, 27(8):1033–1055, July 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/34/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/34/26/
article.pdf.

Boukaram:2018:BQS

[BTLK18] Wajih Halim Boukaram, George Turkiyyah, Hatem Ltaief, and David E. Keyes. Batched QR and SVD algorithms on GPUs with applications in hierarchical matrix compression. Parallel Computing, 74(??):19–33, ???? 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301461.

Broberg:2006:TAW

[BTZ06] James Broberg, Zahir Tari, and Panlop Zeephongsekul. Task assignment with work-conserving migration. Parallel Computing, 32(11–12):808–830, December 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Burrage:1990:ANI

[Bur90] K. Burrage. An adaptive numerical integration code for a chain of transputers. Parallel Computing, 16(2–3):305–312, December 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Burkowski:2004:PPA

[Bur04] Forbes J. Burkowski. Proximity and priority: applying a gene expression algorithm to the Traveling Salesperson Problem. Parallel Computing, 30(5–6): 803–816, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Busvine:1993:IRF

[Bus93] D. Busvine. Implementing recursive functions as processor farms. Parallel Computing, 19(10): 1141–1153, October 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Butel:1992:CVC

[But92] R. Butel. A Cray-2 versus CM-2 comparison using several polynomial benchmarks. Parallel Computing, 18(8):931–945, August 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Buzbee:1986:SV

[Buz86] B. L. Buzbee. A strategy for vectorization. Parallel Computing, 3(3):187– 192, July 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bassiliades:1996:HQE

[BV96] N. Bassiliades and I. Vlahavas. Hierarchical query execution in a parallel object-oriented database system. Parallel Computing, 22(7):1017–1048, October 1, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=7&aid=1082.

Barlas:2004:QLD

[BV04] Gerassimos Barlas and Bharadwaj Veeravalli. Quantized load distribution for tree and bus-connected processors. Parallel Computing, 30(7):841–865, July 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Balaji:2016:SIP

[BVC16] Pavan Balaji, Abhinav Vishnu, and Yong Chen. Special issue on parallel programming models and systems software for highend computing. Parallel Computing, 51(??):1– 2, January 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000053.

Balaji:2019:FSI

[BVC19] Pavan Balaji, Abhinav Vishnu, and Yong Chen.

Foreword to the special issue for the Workshop on Parallel Programming Models and Systems Software for High-End Computing (P2S2 2017). Parallel Computing, 83(??):1–2, April 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303223.

Barendregt:1989:LIL

[BVP^{+}89] H. P. Barendregt, M. C. J. D. Van Eekelen, M. J. Plasmeijer, J. R. W. Glauert, J. R. Kennaway, and M. R. Sleep. LEAN: an intermediate language based on graph rewriting. Parallel Computing, 9(2):163–177, January 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). PARLE: Conference on Parallel Architectures and Languages— Europe (Eindhoven, 1987).

Borstnik:2014:SMM

[BVWH14] Urban Borstnik, Joost VandeVondele, Valéry Weber, and Jürg Hutter. Sparse matrix multiplication: the distributed block-compressed sparse row library. Parallel Computing, 40(5–6):47–58, May 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000428.

Basermann:1992:PAD

[BW92] A. Basermann and P. Weidner. A parallel algorithm for determining all eigenvalues of large real symmetric tridiagonal matrices. Parallel Computing, 18(10):1129–1141, October 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Beckers:2001:PTC

[BW01] Andreas Beckers and Thomas Worsch. A perimeter-time CA for the queen bee problem. Parallel Computing, 27(5):555– 569, April 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/29/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/29/25/
article.pdf.

Blum:2000:AMP

[BWL00] Edward K. Blum, Xin Wang, and Patrick Leung. Architectures and message-passing algorithms for cluster computing: Design and performance.

Parallel Computing, 26 (2–3):313–332, February 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/25/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/25/29/
article.pdf.

Brenner:2008:BCP

[BWT^{+}08] P. Brenner, J. M. Wozniak, D. Thain, A. Striegel, J. W. Peng, and J. A. Izaguirre. Biomolecular committor probability calculation enabled by processing in network storage. Parallel Computing, 34(11): 652–660, November 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Basu:2017:CBC

[BWV^{+}17] Protonu Basu, Samuel Williams, Brian Van Straalen, Leonid Oliker, Phillip Colella, and Mary Hall. Compiler-based code generation and autotuning for geometric multigrid on GPU-accelerated supercomputers. Parallel Computing, 64(??):50–64, May 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819117300376.

Bessenrodt-Weberpals:1989:FVA

[BWW89] M. Bessenrodt-Weberpals and H. Weberpals. A fast vector algorithm for solving tridiagonal linear equations. Parallel Computing, 9(3):367–372, February 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Brown:1995:PAE

[BWZ95] John Brown, Jerzy Wasńiewski, and Zahari Zlatev. Practical aspects and experiences. running air pollution models on massively parallel machines. Parallel Computing, 21(6):971–991, June 12, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=6&aid=981.

Benchara:2021:NSD

[BY21] Fatéma Zahra Benchara and Mohamed Youssfi. A new scalable distributed k-means algorithm based on cloud micro-services for high-performance computing. Parallel Computing, 101(??):Article 102736, April 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120301186.

Baronio:1996:DDT

[BZ96] A. Baronio and F. Zama. A domain decomposition technique for spline image restoration on distributed memory systems. Parallel Computing, 22 (1):101–110, February 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=1&aid=1046.

Benkner:1999:CHP

[BZ99] Siegfried Benkner and Hans Zima. Compiling High Performance Fortran for distributed-memory architectures. Parallel Computing, 25(13– 14):1785–1825, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/36/34/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/34/
article.pdf.

Caron:2001:SSO

[C^{+}01a] E. Caron et al. Scilab to Scilab_{//}: The Ouragan project. Parallel Computing, 27(11):1497–1519, October 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/41/33/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/41/33/
article.pdf.

Christen:2001:SPA

[C^{+}01b] Peter Christen et al. Scalable parallel algorithms for surface fitting and data mining. Parallel Computing, 27(7):941–961, June 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/31/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/31/29/
article.pdf.

Chockalingam:1992:RHM

[CA92] T. Chockalingam and S. Arunkumar. A randomized heuristics for the mapping problem: The genetic approach. Parallel Computing, 18(10):1157–1165, October 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chase:1998:MDT

[CAA98] Craig Chase, Prakash Arunachalam, and Jacob Abraham. Memory distribution: Techniques and practice for CAD applications. Parallel Computing, 24(11):1597–1615, October 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/11/1329.pdf.

Carrion:2009:TBI

[CAC^{+}09] I. Marín Carrión, E. Arias Antúnez, M. M. Artigao Castillo, J. J. Águila Guerrero, and J. J. Miralles Canals. Thread-based implementations of the false nearest neighbors method. Parallel Computing, 35 (10–11):523–534, October/ November 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Carpen-Amarie:2017:EOC

[CAHT17] Alexandra Carpen-Amarie, Sascha Hunold, and Jesper Larsson Träff. On expected and observed communication performance with MPI derived datatypes. Parallel Computing, 69(??):98–117, November 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301217.

Calahan:1985:TGS

[Cal85] D. A. Calahan. Task granularity studies on a many-processor CRAY X-MP. Parallel Computing, 2(2):109–118, June 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Calvin:1996:IPF

[Cal96] C. Calvin. Implementation of parallel FFT algorithms on distributed memory machines with a minimum overhead of communication. Parallel Computing, 22(9):1255–1279, November 22, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=9&aid=1090.

Capdevila:1988:SDE

[Cap88] H. Capdevila. Solution of 2-D Euler equations with a parallel code. Parallel Computing, 7(3):451–460, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Carnevali:1988:TRS

[Car88a] P. Carnevali. Timing results of some internal sorting algorithms on the IBM 3090. Parallel Computing, 6(1):115–117, January 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Carver:1988:SMM

[Car88b] G. Carver. A spectral meteorological method on the ICL DAP. Parallel Computing, 8(1–3):121– 126, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Carriero:1998:ILN

[Car98] Nicholas Carriero. An implementation of Linda for a NUMA machine. Parallel Computing, 24 (7):1005–1021, July 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/7/1317.pdf.

Chakraborty:1991:NUT

[CARW91] A. Chakraborty, D. C. S. Allison, C. J. Ribbens, and L. T. Watson. Note on unit tangent vector computation for homotopy curve tracking on a hypercube. Parallel Computing, 17(12):1385–1395, December 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Canedo:2009:ECQ

[CAS09] Arquimedes Canedo, Ben A.

Abderazek, and Masahiro Sowa. Efficient compilation for queue size constrained queue processors. Parallel Computing, 35(4): 213–225, April 2009. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cichocki:1997:NNS

[CB97] A. Cichocki and A. Bargiela.

Neural networks for solving linear inequality systems. Parallel Computing, 22(11):1455–1475, January 26, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=11&aid=1116.

Correa:2009:POD

[CB09] Ricardo C. Corrêa and Valmir C. Barbosa. Partially ordered distributed computations on asynchronous point-to-point networks. Parallel Computing, 35(1):12–28, January 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Casas:2017:PIN

[CB17] Marc Casas and Greg Bronevetsky. Prediction of the impact of network switch utilization on application performance via active measurement. Parallel Computing, 67(??):38–56, September 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781911730090X.

Cirne:2007:EEE

[CBP^{+}07] Walfredo Cirne, Francisco Brasileiro, Daniel Paranhos, Luís Fabrício W. Góes, and William Voorsluys. On the efficacy, efficiency and emergent behavior of task replication in large distributed systems. Parallel Computing, 33(3):213– 234, April 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2013:SIP

[CBV13] Yong Chen, Pavan Balaji, and Abhinav Vishnu. Special issue on programming models, systems software, and tools for high-end computing. Parallel Computing, 39(12):751–752, December 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001208.

Chen:1991:SAP

[CC91] Gen Huey Chen and Maw Sheng Chern. Synthesis of algorithms on processor arrays. Parallel Computing, 17(2–3):279– 284, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chou:1992:BAS

[CC92] Hong Chich Chou and Chung Ping Chung. A bound analysis of scheduling instructions on pipelined processors with a maximal delay of one cycle. Parallel Computing, 18(4):393– 399, April 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cheng:1993:RAA

[CC93] Ren-Lianq Cheng and Chung-Ping Chung. Reaching approximate agreement on hypercube. Parallel Computing, 19(7):765– 775, July 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:1995:EAC

[CC95] Ling Chen and Henry Y. H. Chuang. An efficient algorithm for complete Euclidean distance transform on mesh-connected SIMD (short communication). Parallel Computing, 21(5):841–852, May 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=5&aid=969.

Calvin:1996:PEM

[CC96a] C. Calvin and L. Colombet. Performance evaluation and modeling of collective communications on Cray T3D. Parallel Computing, 22(10): 1413–1427, December 15, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=10&aid=1097.

Chandwani:1996:FAP

[CC96b] M. Chandwani and N. S. Chaudhari. Formulation and analysis of parallel context-free recognition and parsing on a PRAM model. Parallel Computing, 22(6): 845–868, September 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=6&aid=1067.

Cermele:1997:NUD

[CC97a] M. Cermele and M. Colajanni. Non-uniform and dynamic domain decompositions for hypercomputing. Parallel Computing, 23(6):699–720, June 20, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=6&aid=1179.

Chiola:1997:ILC

[CC97b] G. Chiola and G. Ciaccio.

Implementing a low cost, low latency parallel platform. Parallel Computing, 22(13):1703–1717, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.

                                        cgi?year=1997&volume=
                                        22&issue=13&aid=1123.

Chang:1998:ET

[CC98] Weng-Long Chang and Chih-Ping Chu. The extension of the I test. Parallel Computing, 24 (14):2101–2127, December 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/14/1368.pdf.

Clematis:1999:MPH

[CC99a] A. Clematis and A. Corana. Modeling performance of heterogeneous parallel computing systems. Parallel Computing, 25(9): 1131–1145, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/32/25/20/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/25/20/
                                        article.pdf.

Clementi:1999:EPL

[CC99b] Enrico Clementi and Giorgina Corongiu. Early parallelism with a loosely coupled array of processors: The ICAP experiment. Parallel Computing, 25(13–14):1583–1600, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/36/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/26/
article.pdf.

Chang:2000:ILT

[CC00a] Weng-Long Chang and Chih-Ping Chu. The infinity Lambda test: a multi-dimensional version of Banerjee infinity test. Parallel Computing, 26 (10):1275–1295, August 15, 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/31/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/31/25/
article.pdf.

Chatagny:2000:PMF

[CC00b] Rodolphe Chatagny and Bastien Chopard. A parallel model for the foreign exchange market. Parallel Computing, 26(5):587– 600, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/27/27/
abstract.html;   http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/27/27/
                                        article.pdf.

Chiola:2000:EPP

[CC00c] G. Chiola and G. Ciaccio. Efficient parallel processing on low-cost clusters with GAMMA active ports. Parallel Computing, 26(2–3):333–354, February 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/25/30/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/25/30/
                                        article.pdf.

Chang:2001:GDV

[CC01] Weng-Long Chang and Chih-Ping Chu. The generalized Direction Vector I test. Parallel Computing, 27(8):1117–1144, July 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/47/34/29/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/34/29/
                                        article.pdf.

Chang:2002:IFT

[CC02a] Huei-Huang Chang and Ge-Ming Chiu. An improved fault-tolerant routing algorithm in meshes with convex faults. Parallel Computing, 28(1):133–149, January 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/27/34/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/27/34/
00001682.pdf.

Chang:2002:EPA

[CC02b] Li-Chiu Chang and Fi-John Chang. An efficient parallel algorithm for LISSOM neural network. Parallel Computing, 28(11):1611–1633, November ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/61/31/
abstract.html.

Childs:2016:PVD

[CC16] Hank Childs and Franck Cappello. Preface: Visualization and data analytics for scientific discovery. Parallel Computing, 55 (??):1, July 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300308.

Casiccia:1992:SMC

[CCCP92] P. Casiccia, P. Castangia, S. Cincotti, and G. Parodi. Simulation of a molecular cellular array on a transputer-based parallel computer. Parallel Computing, 18(3):313–324, March 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cappuccio:2001:PIC

[CCEJ01] R. Cappuccio, G. Cattaneo, G. Erbacci, and U. Jocher. A parallel implementation of a cellular automata based model for coffee percolation. Parallel Computing, 27(5):685– 717, April 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/29/32/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/29/32/
                                        article.pdf.

Cave:2017:TGH

[CCG^{+}17] Vincent Cavé, Romain Clédat, Paul Griffin, Ankit More, Bala Seshasayee, Shekhar Borkar, Sanjay Chatterjee, Dave Dunning, and Joshua Fryman. Traleika Glacier: a hardware–software codesigned approach to exascale computing. Parallel Computing, 64(??):33– 49, May 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300169.

Camara:2014:ATN

[CCGG14] Jesús Cámara, Javier Cuenca, Luis-Pedro García, and Domingo Giménez. Auto-tuned nested parallelism: a way to reduce the execution time of scientific software in NUMA systems. Parallel Computing, 40(7):309–327, July 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000416.

Czappa:2021:DTP

[CCH^{+}21] Fabian Czappa, Alexandru Calotoiu, Thomas Höhl, Heiko Mantel, Toni Nguyen, and Felix Wolf. Design-time performance modeling of compositional parallel programs. Parallel Computing, 108(??):??, December 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000855.

Chen:1990:PAD

[CCJ90] Gen-Huey Chen, Maw-Sheng Chern, and Jin-Hwang Jang. Pipeline architectures for dynamic programming algorithms. Parallel Computing, 13(1): 111–117, January 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chang:2004:SOP

[CCL04] Rong-Guey Chang, Tyng-Ruey Chuang, and Jenq Kuen Lee. Support and optimization for parallel sparse programs with array intrinsics of Fortran 90. Parallel Computing, 30(4): 527–550, April 2004. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ciciani:1998:PED

[CCP98] B. Ciciani, M. Colajanni, and C. Paolucci. Performance evaluation of deterministic wormhole routing in k-ary n-cubes. Parallel Computing, 24 (14):2053–2075, December 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/14/1371.pdf.

Couch:2021:TPP

[CCP^{+}21] Sean M. Couch, Jared Carlson, Michael Pajkos, Brian W. O’Shea, Anshu Dubey, and Tom Klosterman. Towards performance portability in the Spark astrophysical magnetohydrodynamics solver in the Flash-X simulation framework. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000788.

Corana:1991:ECC

[CCRR91] A. Corana, A. Casaleggio, C. Rolando, and S. Ridella. Efficient computation of the correlation dimension from a time series on a LIW computer. Parallel Computing, 17(6–7): 809–820, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Carlisle:1987:AMS

[CCS87] H. Carlisle, A. Crawford, and S. Sheppard. ADA multitasking and the single source shortest path problem. Parallel Computing, 4(1):75–91, February 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chang:1994:PAK

[CCS94] Henry Ker-Chang Chang, Jonathan Jen-Rong Chen, and Shyong-Jian Shyu. A parallel algorithm for the knapsack problem using a generation and searching technique. Parallel Computing, 20(2): 233–243, February 24, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=2&aid=827.

Chan:2003:HLA

[CCS03] Fan Chan, Jiannong Cao, and Yudong Sun. Highlevel abstractions for messagepassing parallel programming. Parallel Computing, 29(11–12):1589–1621, November/December 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2018:OPD

[CCS^{+}18] Xuhao Chen, Cheng Chen, Jie Shen, Jianbin Fang, Tao Tang, Canqun Yang, and Zhiying Wang. Orchestrating parallel detection of strongly connected components on GPUs. Parallel Computing, 78(??):101–114, October 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301837.

Chang:2001:MDV

[CCW01] Weng-Long Chang, Chih-Ping Chu, and Jia-Hwa Wu. A multi-dimensional version of the I test. Parallel Computing, 27 (13):1783–1799, December 1, 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/43/33/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/43/33/
article.pdf.

Casanova:1998:UAB

[CD98] Henri Casanova and Jack Dongarra. Using agentbased software for scientific computing in the NetSolve system. Parallel Computing, 24(12– 13):1777–1790, November 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/12-13/1354.pdf.

Chu:2000:ACC

[CD00] John C. Chu and Patrick W.

Dowd. Adaptive cache coherence over a high bandwidth broadband mesh network. Parallel Computing, 26(2–3):285–311, February 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/25/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/25/28/
article.pdf.

Cruz:2016:LLA

[CDA^{+}16] Eduardo H. M. Cruz, Matthias Diener, Marco A. Z. Alves, Laércio L. Pilla, and Philippe O. A. Navaux. LAPT: a localityaware page table for thread and data mapping. Parallel Computing, 54(??):59– 71, May 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001556.

Chan:2008:CGP

[CDBL08] Albert Chan, Frank Dehne, Prosenjit Bose, and Markus Latzel. Coarse grained parallel algorithms for graph matching. Parallel Computing, 34(1):47–62, January 2008. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Clementi:1987:LSC

[CDC^{+}87] E. Clementi, J. Detrich, S. Chin, G. Corongiu, D. Folsom, D. Logan, R. Caltabiano, A. Carnevali, J. Helin, M. Russo, A. Gnudi, and P. Palamidese. Large-scale computations on a scalar, vector and parallel ‘supercomputer’. Parallel Computing, 5(1–2): 13–44, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Caromel:2007:PPC

[CdCM07] Denis Caromel, Alexandre di Costanzo, and Clément Mathieu. Peer-to-peer for computational grids: mixing clusters and desktop machines. Parallel Computing, 33(4–5):275– 288, May 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Campanini:1993:TBP

[CDDG93] R. Campanini, I. D’Antone, G. Di Caro, and G. Giusti. A transputer-based parallel expert diagnostic system. Parallel Computing, 19(6):685–692, June 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Collange:2015:NRP

[CDGI15] Sylvain Collange, David Defour, Stef Graillat, and Roman Iakymchuk. Numerical reproducibility for the parallel reduction on multi- and many-core architectures. Parallel Computing, 49(??):83–97, November 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001155.

Conforti:1996:PIA

[CDGM96] D. Conforti, L. De Luca, L. Grandinetti, and R. Musmanno. A parallel implementation of automatic differentiation for partially separable functions using PVM. Parallel Computing, 22(5):643–656, August 8, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=5&aid=1065.

Coulaud:2003:OKL

[CDH^{+}03] Olivier Coulaud, Michaël Dussere, Pascal Hénon, Erik Lefebvre, and Jean Roman. Optimization of a kinetic laser–plasma interaction code for large parallel systems. Parallel Computing, 29(9):1175–1189, September 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Couturier:2008:GLS

[CDJ08] Raphaël Couturier, Christophe Denis, and Fabienne Jézéquel. GREMLINS: a large sparse linear solver for grid environment. Parallel Computing, 34(6–8):380–391, July 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chervenak:2003:HPR

[CDK^{+}03] Ann Chervenak, Ewa Deelman, Carl Kesselman, Bill Allcock, Ian Foster, Veronika Nefedova, Jason Lee, Alex Sim, Arie Shoshani, Bob Drach, et al. High-performance remote access to climate simulation data: a challenge problem for data grid technologies. Parallel Computing, 29(10):1335–1356, October 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Carracciuolo:2006:TPC

[CDM06] L. Carracciuolo, L. D’Amore, and A. Murli. Towards a parallel component for imaging in PETSc programming environment: a case study in 3-D echocardiography. Parallel Computing, 32(1):67–83, January 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Corsaro:2010:PAL

[CDM^{+}10] S. Corsaro, P. L. De Angelis, Z. Marino, F. Perla, and P. Zanetti. On parallel asset-liability management in life insurance: a forward risk-neutral approach. Parallel Computing, 36(7):390–402, July 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cannataro:1995:PCA

[CDR^{+}95] M. Cannataro, S. Di Gregorio, R. Rongo, W. Spataro, G. Spezzano, and D. Talia. A parallel cellular automata environment on multicomputers for computational science. Parallel Computing, 21(5):803– 823 (or 803–824??), May 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=5&aid=965.

Calland:1997:PAO

[CDRV97] Pierre-Yves Calland, Alain Darte, Yves Robert, and Frederic Vivien. Plugging anti and output dependence removal techniques into loop parallelization algorithm. Parallel Computing, 23(1–2):251–266, April 16, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=1-2&aid=1159.

Colombo:2008:RTF

[CDV08] C. Colombo, A. Del Bimbo, and A. Valli. A real-time full body tracking and humanoid animation system. Parallel Computing, 34(12):718– 726, December 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Choi:1995:PMT

[CDW95] Jaeyoung Choi, Jack J.

Dongarra, and David W. Walker. Parallel matrix transpose algorithms on distributed memory concurrent computers. Parallel Computing, 21(9): 1387–1405, September 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=9&aid=995.

Ceron:1998:PID

[CDZ^{+}98] C. Ceron, J. Dopazo, E. L. Zapata, J. M. Carazo, and O. Trelles. Parallel implementation of DNAml program on message-passing architectures. Parallel Computing, 24(5–6):701– 716, June 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/5-6/1279.pdf.

Cheng:2021:ADQ

[CDZ^{+}21] Wen Cheng, Shijun Deng, Lingfang Zeng, Yang Wang, and André Brinkmann. AIOC^{2}: a deep Q-learning approach to autonomic I/O congestion control in Lustre. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000995.

Calinescu:1994:PSM

[CE94] R. Calinescu and D. J. Evans. A parallel simulation model for load balancing in clustered distributed systems. Parallel Computing, 20(1):77–91, January 16, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=1&aid=819.

Clune:1999:CAC

[CEM^{+}99] T. C. Clune, J. R. Elliott, M. S. Miesch, J. Toomre, and G. A. Glatzmaier. Computational aspects of a code to study rotating turbulent convection in spherical shells. Parallel Computing, 25(4):361–380, April 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/4/1401.pdf.

Cerin:2007:LSG

[Cér07] Christophe Cérin. Large scale grids. Parallel Computing, 33(4–5):235–237, May 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cosnard:1990:FRP

[CF90] Michel Cosnard and Pierre Fraigniaud. Finding the roots of a polynomial on an MIMD multicomputer. Parallel Computing, 15(1– 3):75–85, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:1991:EPT

[CF91] Wen Tsuen Chen and Ming Yi Fang. An efficient procedure for theorem proving in propositional logic on vector computers. Parallel Computing, 17(9): 983–995, November 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chan:1993:PAS

[CF93] I. W. Chan and D. K. Friesen. Parallel algorithm for segment visibility reporting. Parallel Computing, 19(9):973–978, September 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2004:MLT

[CF04] Ching-Wen Chen and Shih-Chang Fu. A minimal links traversed dynamic rerouting network. Parallel Computing, 30(7):883– 898, July 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:1999:NSP

[CFC99] S. W. Chen, C. Y. Fang, and K. E. Chang. Neural simulation of Petri nets. Parallel Computing, 25(2):183–207, February 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/2/1383.pdf.

Castro:2016:SWP

[CFD^{+}16] Márcio Castro, Emilio Francesquini, Fabrice Dupros, Hideo Aochi, Philippe O. A. Navaux, and Jean-François Méhaut. Seismic wave propagation simulations on low-power and performance-centric manycores. Parallel Computing, 54(??):108–120, May 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000417.

Cardellini:2017:CBL

[CFF17] Valeria Cardellini, Alessandro Fanfarillo, and Salvatore Filippone. Coarraybased load balancing on heterogeneous and manycore architectures. Parallel Computing, 68(??):45– 58, October 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300844.

Catalyurek:2012:GCA

[ÇFG^{+}12] Ümit V. Çatalyürek, John Feo, Assefaw H. Gebremedhin, Mahantesh Halappanavar, and Alex Pothen. Graph coloring algorithms for multicore and massively multithreaded architectures. Parallel Computing, 38 (10–11):576–594, October/ November 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819112000592.

Cosnard:1989:TLA

[CFH89] M. Cosnard, A. G. Ferreira, and H. Herbelin. The two list algorithm for the knapsack problem on a FPS T20. Parallel Computing, 9(3):385–388, February 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Feng:2007:HPC

[cFM07] Wu chun Feng and Dinesh Manocha. Highperformance computing using accelerators. Parallel Computing, 33(10–11): 645–647, November 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cattaneo:1997:TOD

[CFMM97] G. Cattaneo, E. Formenti, L. Margara, and G. Mauri. Transformations of the one-dimensional cellular automata rule space. Parallel Computing, 23(11): 1593–1611, December 1, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2003:OPP

[CFN03] Li Chen, Issei Fujishiro, and Kengo Nakajima. Optimizing parallel performance of unstructured volume rendering for the Earth Simulator. Parallel Computing, 29(3):355– 371, March 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2001:IGR

[CFS01] Ting Chen, Vladimir Filkov, and Steven S. Skiena. Identifying gene regulatory networks from experimental data. Parallel Computing, 27(1–2): 141–162, January 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/47/
25/28/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/47/
25/28/article.pdf.

Choi:2022:ACP

[CFW^{+}22] Jaemin Choi, Zane Fink, Sam White, Nitin Bhat, David F. Richards, and Laxmikant V. Kale. Accelerating communication for parallel programming models on GPU systems. Parallel Computing, 113(??):??, October 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000606.

Chu:1987:GEP

[CG87] Eleanor Chu and Alan George. Gaussian elimination with partial pivoting and load balancing on a multiprocessor. Parallel Computing, 5(1–2):65– 74, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Chronopoulos:1989:EIP

[CG89a] A. T. Chronopoulos and C. W. Gear. On the efficient implementation of preconditioned s-step conjugate gradient methods on multiprocessors with memory hierarchy. Parallel Computing, 11(1):37– 53, July 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chu:1989:FDM

[CG89b] Eleanor Chu and Alan George. QR factorization of a dense matrix on a shared-memory multiprocessor. Parallel Computing, 11(1):55–71, July 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chu:1992:BSM

[CG92] E. Chu and A. George. A balanced submatrix merging algorithm for multiprocessor architectures. Parallel Computing, 18(1):1– 10, January 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chalmers:1993:CMP

[CG93] A. G. Chalmers and S. Gregory. Constructing minimum path configurations for multiprocessor systems. Parallel Computing, 19(3):343–355, March 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Censor:2001:CAE

[CGG01] Yair Censor, Dan Gordon, and Rachel Gordon. Component averaging: an efficient iterative parallel algorithm for large and sparse unstructured problems. Parallel Computing, 27(6):777– 808, May 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/30/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/30/25/
article.pdf.

Cuenca:2004:AAT

[CGG04] Javier Cuenca, Domingo Giménez, and José González. Architecture of an automatically tuned linear algebra library. Parallel Computing, 30(2):187–210, February 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Caricato:2003:PTS

[CGGG03] Pierpaolo Caricato, Gianpaolo Ghiani, Antonio Grieco, and Emanuela Guerriero. Parallel tabu search for a pickup and delivery problem under track contention. Parallel Computing, 29(5):631– 639, May 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cojean:2019:RAT

[CGH^{+}19] T. Cojean, A. Guermouche, A. Hugo, R. Namyst, and P. A. Wacrenier. Resource aggregation for task-based Cholesky factorization on top of modern architectures. Parallel Computing, 83(??):73– 92, April 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303387.

Ceballos:2018:APV

[CGHBS18] Germán Ceballos, Thomas Grass, Andra Hugo, and David Black-Schaffer. Analyzing performance variation of task schedulers with TaskInsight. Parallel Computing, 75(??):11– 27, July 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300346.

Calore:2016:MPL

[CGK^{+}16] E. Calore, A. Gabbana, J. Kraus, E. Pellegrini, S. F. Schifano, and R. Tripiccione. Massively parallel lattice-Boltzmann codes on large GPU clusters. Parallel Computing, 58(??):1–24, October 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300825.

Conforti:1992:MEA

[CGM^{+}92] D. Conforti, L. Grandinetti, R. Musmanno, M. Cannataro, G. Spezzano, and D. Talia. A model of efficient asynchronous parallel algorithms on multicomputer systems. Parallel Computing, 18(1):31– 45, January 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cuenca:2005:HWD

[CGM05] Javier Cuenca, Domingo Giménez, and Juan-Pedro Martínez. Heuristics for work distribution of a homogeneous parallel dynamic programming scheme on heterogeneous systems. Parallel Computing, 31(7):711–735, July 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Corral:1999:PSP

[CGMM99] Cristina Corral, Isabel Giménez, José Marín, and José Mas. Parallel mstep preconditioners for the conjugate gradient method. Parallel Computing, 25(3):265–281, March 22, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1999/
                                        25/3/1387.pdf.

Carriero:1994:LAM

[CGMS94] Nicholas J. Carriero, David Gelernter, Timothy G. Mattson, and Andrew H. Sherman. The Linda alternative to message-passing systems. Parallel Computing, 20(4):633–655, March 31, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=4&aid=858.

Clint:2006:PMA

[CGNR06] Maurice Clint, Efstratios Gallopoulos, Esmond Ng, and Jean Roman. Parallel Matrix Algorithms and Applications (PMAA’04). Parallel Computing, 32(2): 113–114, February 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Correa:2003:PIA

[CGOP03] Ricardo C. Corrêa, Fernando C. Gomes, Carlos A. S. Oliveira, and Panos M. Pardalos. A parallel implementation of an asynchronous team to the point-to-point connection problem. Parallel Computing, 29(4):447–466, April 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chu:1993:PMI

[CGQ93] E. Chu, A. George, and D. Quesnel. Parallel matrix inversion on a subcube-grid. Parallel Computing, 19(3):243–256, March 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chu:1992:ANN

[CH92] Y. P. Chu and C. M. Hsieh. An artificial neural network model with modified perceptron algorithm. Parallel Computing, 18(9): 983–996, September 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cormen:1997:EEE

[CH97] Thomas H. Cormen and Melissa Hirschl. Early experiences in evaluating the parallel disk model with the ViC* implementation. Parallel Computing, 23(4–5):571–600, May 23, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=4-5&aid=1172.

Caprioli:1998:PQN

[CH98] Paul Caprioli and Mark H.

Holmes. A parallel quasi-Newton method for Gaussian data fitting. Parallel Computing, 24(11): 1635–1651, October 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/11/1332.pdf.

Chamberlain:1988:GCF

[Cha88] R. M. Chamberlain. Gray codes, fast Fourier transforms and hypercubes. Parallel Computing, 6(2): 225–233, February 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chadwick:1999:HPA

[Cha99] Edmund Chadwick. A hybrid parallel algorithm for the spectral transform method which uses functional parallelism. Parallel Computing, 25(4):345–360, April 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/4/1386.pdf.

Castain:2018:PPM

[CHBS18] Ralph H. Castain, Joshua Hursey, Aurelien Bouteiller, and David Solt. PMIx: Process management for exascale environments. Parallel Computing, 79 (??):9–29, November 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819118302424.

Cappello:2009:F

[CHD09] Franck Cappello, Thomas Herault, and Jack Dongarra. Foreword. Parallel Computing, 35(12):571, December 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:1996:PGH

[Che96] Lin Chen. Partitioning graphs into Hamiltonian ones. Parallel Computing, 22(4):607–618, June 11, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=4&aid=1058.

Chen:1998:EFC

[Che98a] Chung-Ming Chen. An efficient four-connected parallel system for PET image reconstruction. Parallel Computing, 24(9– 10):1499–1522, September 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/9-10/1345.pdf.

Chervenak:1998:CTS

[Che98b] Ann L. Chervenak. Challenges for tertiary storage in multimedia servers. Parallel Computing, 24 (1):157–176, March 10, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1998&volume=
24&issue=1&aid=1278.

Chedid:2008:OPT

[Che08] Fouad B. Chedid. An optimal parallelization of the two-list algorithm of cost O(2^{n/2}). Parallel Computing, 34(1):63–65, January 2008. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2001:GBS

[CHH^{+}01] Gen-Huey Chen, Shien-Ching Hwang, Hui-Ling Huang, Ming-Yang Su, and Dyi-Rong Duh. A general broadcasting scheme for recursive networks with complete connection. Parallel Computing, 27(9):1273–1278, August 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/35/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/35/28/
article.pdf.

Calkin:1994:PPP

[CHHW94] R. Calkin, R. Hempel, H.- C. Hoppe, and P. Wypior. Portable programming with the PAR-MACS message-passing library. Parallel Computing, 20(4):615–632, March 31, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=4&aid=859.

Clausing:1985:TAP

[CHLO85] J. A. Clausing, R. Hagstrom, E. L. Lusk, and R. A. Overbeek. A technique for achieving portability among multiprocessors: Implementation on the Lemur. Parallel Computing, 2(2):137–162, June 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:1990:DML

[CHLS90] Gen Huey Chen, Hong Fa Ho, Shieu Hong Lin, and Jang-Ping Sheu. Data mapping of linear programming on fixed-size hypercubes. Parallel Computing, 13(2):235–243, February 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Choi:1991:RVW

[Cho91] Y.-H. Choi. Reconfigurable VLSI/WSI multipipelines. Parallel Computing, 17(8):941–952, October 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Choi:1992:EDF

[Cho92] Y.-H. Choi. An easilydiagnosable fault-tolerant binary tree architecture (short communication). Parallel Computing, 18 (10):1185–1195, October 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Catalan:2018:EBB

[CHQORS18a] Sandra Catalán, José R.

Herrero, Enrique S. Quintana-Ortí, and Rafael Rodríguez-Sánchez. Energy balance between voltage-frequency scaling and resilience for linear algebra routines on low-power multicore architectures. Parallel Computing, 74(??):28–39, ???? 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300765.

Catalan:2018:SSF

[CHQORS18b] Sandra Catalán, José R. Herrero, Enrique S. Quintana-Ortí, and Rafael Rodríguez-Sánchez. Static scheduling of the LU factorization with look-ahead on asymmetric multicore processors. Parallel Computing, 76(??):18–27, August 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301194.

Christensen:1998:MVS

[Chr98] Gary E. Christensen. MIMD vs. SIMD parallel processing: a case study in 3D medical image registration. Parallel Computing, 24(9– 10):1369–1383, September 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/9-10/1339.pdf.

ChassindeKergommeaux:2000:PCI

[CHR00a] J. Chassin de Kergommeaux, P. J. Hatcher, and L. Rauchwerger. Parallel computing for irregular applications. Parallel Computing, 26(13– 14):1681–1684, December 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/34/22/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/34/22/
article.pdf.

Chretienne:2000:GBC

[Chr00b] Philippe Chrétienne. On Graham’s bound for cyclic scheduling. Parallel Computing, 26(9):1163–1174, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/30/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/30/28/
article.pdf.

Casanova:2018:CEM

[CHR18] Henri Casanova, Julien Herrmann, and Yves Robert. Computing the expected makespan of task graphs in the presence of silent errors. Parallel Computing, 75(??):41–60, July 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118300796.

Capelli:2019:IVC

[CHZ^{+}19] Ludovic A. R. Capelli, Zhenjiang Hu, Timothy A. K. Zakian, Nick Brown, and J. Mark Bull. iPregel: Vertex-centric programmability vs memory efficiency and performance, why choose? Parallel Computing, 86(??): 45–56, August 2019. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0167819118303788.

Creusillet:1998:IAF

[CI98] B. Creusillet and F. Irigoin. Interprocedural analyses of Fortran programs. Parallel Computing, 24(3–4):629– 648, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/3-4/1305.pdf.

Cielo:2021:VWL

[CIG^{+}21] Salvatore Cielo, Luigi Iapichino, Johannes Günther, Christoph Federrath, Elisabeth Mayer, and Markus Wiedemann. Visualizing the world’s largest turbulence simulation. Parallel Computing, 102(??): Article 102758, May 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S016781912100017X.

Costero:2017:RCT

[CIO^{+}17] Luis Costero, Francisco D.

Igual, Katzalin Olcoz, Sandra Catalán, Rafael Rodríguez-Sánchez, and Enrique S. Quintana-Ortí. Revisiting conventional task schedulers to exploit asymmetry in multi-core architectures for dense linear algebra operations. Parallel Computing, 68(??):59–76, October 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300856.

Cabibbo:1999:HPC

[CIS99] N. Cabibbo, Y. Iwasaki, and K. Schilling. High performance computing in lattice QCD. Parallel Computing, 25(10– 11):1197–1198, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/32/26/30/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/26/30/
                                        article.pdf.

Chen:1992:IPA

[CJ92] Gen Huey Chen and Jin Hwang Jang. An improved parallel algorithm for 0/1 knapsack problem. Parallel Computing, 18(7):811–821, July 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Crivelli:1995:CEC

[CJ95] S. Crivelli and E. R. Jessup. The cost of eigenvalue computation on distributed-memory MIMD multiprocessors. Parallel Computing, 21 (3):401–422, March 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=3&aid=966.

Chalmers:1997:PGV

[CJ97] A. Chalmers and F. W. Jansen. Parallel graphics and visualisation. Parallel Computing, 23(7):817– ??, July 14, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chandra:1993:SAT

[CJBK93] S. Chandra, M. Jain, A. Basu, and P. S. Kumar. Sorting algorithms on transputer arrays. Parallel Computing, 19(6):595– 607, June 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cho:2014:DCA

[CJLS14] Joong-Yeon Cho, Hyun-Wook Jin, Min Lee, and Karsten Schwan. Dynamic core affinity for highperformance file upload on Hadoop Distributed File System. Parallel Computing, 40(10):722–737, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000933.

Carnevali:1990:SMP

[CK90] P. Carnevali and M. Kindelan. A simplified model to predict the performance of FORTRAN vector loops on the IBM 3090/VF. Parallel Computing, 13(1):35– 46, January 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cotronis:2014:CCG

[CKLM14] Yiannis Cotronis, Elias Konstantinidis, Maria A. Louka, and Nikolaos M. Missirlis. A comparison of CPU and GPU implementations for solving the convection diffusion equation using the local modified SOR method. Parallel Computing, 40(7):173– 185, July 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000234.

Czech:1993:PAF

[CKM93] Zbigniew J. Czech, Marek Konopka, and Bohdan S. Majewski. Parallel algorithms for finding a suboptimal fundamental-cycle set in a graph. Parallel Computing, 19(9):961–971, September 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:1994:GMI

[CKM94] Lujuan Chen, E. V. Krishnamurthy, and Iain Macleod. Generalised matrix inversion and rank computation by successive matrix powering. Parallel Computing, 20(3):297–311, March 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=3&aid=834.

Chandy:1998:FSD

[CKRZ98] K. Mani Chandy, Joseph Kiniry, Adam Rifkin, and Daniel Zimmerman. A framework for structured distributed object computing. Parallel Computing, 24(12– 13):1901–1922, November 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/12-13/1360.pdf.

Correa:2003:CSF

[CKS03] Wagner T. Corrêa, James T.

Klosowski, and Cláudio T. Silva. Out-of-core sortfirst parallel rendering for cluster-based tiled displays. Parallel Computing, 29(3):325–338, March 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:1992:CFB

[CL92] Gen-Huey Chen and Wei-Wen Liang. Conflict-free broadcasting algorithms for graph traversals and their applications. Parallel Computing, 18(4):439– 448, April 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2002:PAP

[CL02] Ke Chen and Choi H. Lai. Parallel algorithms of the Purcell method for direct solution of linear systems. Parallel Computing, 28(9):1275–1291, September 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/59/31/
                                        abstract.html.

Chunlin:2005:DUB

[CL05] Li Chunlin and Li Layuan. A distributed utility-based two level market solution for optimal resource scheduling in computational grid. Parallel Computing, 31(3–4):332–351, March/April 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Clarke:2003:GFE

[Cla03] K. C. Clarke. Geocomputation’s future at the extremes: high performance computing and nanoclients. Parallel Computing, 29(10):1281–1295, October 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2014:PLS

[CLA14] Jie Chen, Tom L. H. Li, and Mihai Anitescu. A parallel linear solver for multilevel Toeplitz systems with possibly several righthand sides. Parallel Computing, 40(8):408–424, August 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000672.

Chan:2015:PMH

[CLA15] Siew Yin Chan, Teck Chaw Ling, and Eric Aubanel. Performance modeling for hierarchical graph partitioning in heterogeneous multi-core environment. Parallel Computing, 46(??):78–97, July 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000532.

Chamberlain:2008:VAD

[CLC08] Roger D. Chamberlain, Joseph M. Lancaster, and Ron K. Cytron. Visions for application development on hybrid computing systems. Parallel Computing, 34(4–5):201–216, May 2008. CODEN PACOEJ.

ISSN 0167-8191 (print), 1872-7336 (electronic).

Chang:2018:SCC

[CLC^{+}18] ChanJung Chang, YungChing Lin, YuHsuan Cheng, YuCheng Wang, LiYu Yu, TienChi Yang, and Jerry Chou. Student cluster competition 2017, team NTHU: Reproducing vectorization of the Tersoff multi-body potential on the Intel Skylake and Nvidia P100 architecture. Parallel Computing, 78(??):72–78, October 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300656.

Chen:2023:PNC

[CLHL23] Yidong Chen, Chen Li, Yonghong Hu, and Zhonghua Lu. A parallel non-convex approximation framework for risk parity portfolio design. Parallel Computing, 116(??):??, July 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000054.

Coetzee:2014:TUS

[CLJ14] P. Coetzee, M. Leeke, and S. Jarvis. Towards unified secure onand off-line analytics at scale. Parallel Computing, 40(10):738–753, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000842.

Carre:1999:HPS

[CLL99] G. Carré, S. Lanteri, and Mark Loriot. High performance simulations of compressible flows inside car engine geometries using the N3S-NATUR parallel solver. Parallel Computing, 25(12):1435–1458, November 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/35/24/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/35/24/
article.pdf.

Caffey:1993:PPL

[CLS93] H. Caffey, L. Z. Liao, and C. A. Shoemaker. Parallel processing of large scale discrete-time unconstrained differential dynamic programming. Parallel Computing, 19(9): 1003–1017, September 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Carpentieri:2016:UVM

[CLS^{+}16] Bruno Carpentieri, Jia Liao, Masha Sosonkina, Aldo Bonfiglioli, and Sven Baars. Using the VBARMS method in parallel computing. Parallel Computing, 57 (??):197–211, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116000351.

Chen:1998:ORF

[CLSM98] Jih-H. Chen, Shu-Yun Le, Bruce A. Shapiro, and Jacob V. Maizel. Optimization of an RNA folding algorithm for parallel architectures. Parallel Computing, 24(11):1617–1634, October 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/11/1331.pdf.

Cao:2019:OCR

[CLY^{+}19] Yanhua Cao, Li Lu, Jiadi Yu, Shiyou Qian, Yanmin Zhu, and Minglu Li. Online cost-rejection rate scheduling for resource requests in hybrid clouds. Parallel Computing, 81(??):85–103, January 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911830111X.

Corradi:2001:POA

[CLZ01] Antonio Corradi, Letizia Leonardi, and Franco Zambonelli. Parallel object allocation via userspecified directives: a case study in traffic simulation. Parallel Computing, 27(3):223–241, February 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/22/23/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/22/23/
article.pdf.

Casciola:1995:GPC

[CM95] G. Casciola and S. Morigi. Graphics in parallel computation for rendering 3D modelled scenes. Parallel Computing, 21 (8):1365–1382, August 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=8&aid=1002.

Chapman:2005:O

[CM05] Barbara M. Chapman and Federico Massaioli. OpenMP. Parallel Computing, 31(10–12):957–959, October/December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Corona:1988:SLE

[CMM^{+}88a] A. Corona, C. Martini, M. Morando, S. Ridella, and C. Rolando. Solving linear equation systems on vector computers with maximum efficiency. Parallel Computing, 8(1– 3):133–139, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Crookes:1988:APL

[CMM^{+}88b] D. Crookes, P. J. Morrow, P. Milligan, P. L. Kilpatrick, and N. S. Scott. An array processing language for transputer networks. Parallel Computing, 8(1–3):141–148, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Clematis:2003:HPC

[CMM03] Andrea Clematis, Mike Mineter, and Richard Marciano. High performance computing with geographical data. Parallel Computing, 29(10):1275–1279, October 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Clauss:1992:SSO

[CMP92] Ph. Clauss, C. Mongenet, and G. R. Perrin. Synthesis of size-optimal toroidal arrays for the algebraic path problem: a new contribution. Parallel Computing, 18(2):185–194, February 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chavarria-Miranda:2015:GTL

[CMPM^{+}15] Daniel Chavarría-Miranda, Ajay Panyala, Wenjing Ma, Adrian Prantl, and Sriram Krishnamoorthy. Global transformations for legacy parallel applications via structural analysis and rewriting. Parallel Computing, 43(??):1– 26, March 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000083.

Cosnard:1988:PGE

[CMRT88] M. Cosnard, M. Marrakchi, Y. Robert, and D. Trystram. Parallel Gaussian elimination on an MIMD computer. Parallel Computing, 6(3):275–296, March 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cesar:2006:MMW

[CMSL06] E. Cesar, A. Moreno, J. Sorribes, and E. Luque. Modeling master/worker applications for automatic performance tuning. Parallel Computing, 32(7–8): 568–589, September 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Colombet:1996:PMV

[CMT96] L. Colombet, Ph. Michallon, and D. Trystram. Parallel matrix-vector product on rings with a minimum of communications. Parallel Computing, 22(2):289–310, April 5, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=2&aid=1044.

Cahon:2004:BPR

[CMT04a] S. Cahon, N. Melab, and E.-G. Talbi. Building with ParadisEO reusable parallel and distributed evolutionary algorithms. Parallel Computing, 30(5–6):

677–697, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Calzarossa:2004:MTA

[CMT04b] Maria Calzarossa, Luisa Massari, and Daniele Tessera. A methodology towards automatic performance analysis of parallel applications. Parallel Computing, 30(2):211–223, February 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Clement:2006:DDS

[CMV^{+}06] F. Clément, V. Martin, A. Vodicka, R. Di Cosmo, and P. Weis. Domain decomposition and skeleton programming with OCamlP3l. Parallel Computing, 32(7–8):539–550, September 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cormen:1998:PCF

[CN98] Thomas H. Cormen and David M. Nicol. Performing out-of-core FFTs on parallel disk systems. Parallel Computing, 24(1):5–20, March 10, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/

                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1998&volume=
                                        24&issue=1&aid=1272.

Choudhary:1993:EHS

[CNK93] A. N. Choudhary, B. Narahari, and R. Krishnamurti. An efficient heuristic scheme for dynamic remapping of parallel computations (short communication). Parallel Computing, 19(6):621–632, June 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cole:2004:BSC

[Col04] Murray Cole. Bringing skeletons out of the closet: a pragmatic manifesto for skeletal parallel programming. Parallel Computing, 30(3):389–406, March 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Conroy:1989:NPC

[Con89] John M. Conroy. A note on the parallel Cholesky factorization of wide banded matrices. Parallel Computing, 10(2):239–246, April 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Conroy:1990:PND

[Con90] John M. Conroy. Parallel nested dissection. Parallel Computing, 16(2–3):

139–156, December 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cools:2019:AIM

[Coo19]  S. Cools.  Analyzing

and improving maximal attainable accuracy in the communication hiding pipelined BiCGStab method. Parallel Computing, 86(??):16–35, August 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118302576.

Corana:1999:PCC

[Cor99] Angelo Corana. Parallel computation of the correlation dimension from a time series. Parallel Computing, 25(6):639–666, June 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/6/1421.pdf.

Correa:2000:PAS

[Cor00] Ricardo C. Corrêa. A parallel approximation scheme for the multiprocessor scheduling problem. Parallel Computing, 26(1):47–72, January 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/42/23/26/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/23/26/
                                        article.pdf.

Chevalier:2008:PST

[CP08] C. Chevalier and F. Pellegrini. PT-Scotch: a tool for efficient parallel graph ordering. Parallel Computing, 34(6–8):318–331, July 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chandrasekaran:2017:SIT

[CP17] Sunita Chandrasekaran and Antonio J. Peña. Special issue on topics on heterogeneous computing. Parallel Computing, 68(??):1–2, October 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117301084.

Chandrasekaran:2018:SIA

[CP18] Sunita Chandrasekaran and Antonio J. Peña. Special issue on applications for the heterogeneous computing era 2017. Parallel Computing, 77(??):125–127, September 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301844.

Carretero:1997:PIM

[CPdM^{+}97] J. Carretero, F. Pérez, P. de Miguel, F. García, and L. Alonso. Performance increase mechanisms for parallel and distributed file systems. Parallel Computing, 23 (4–5):525–542, May 23, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=4-5&aid=1170.

Chen:2018:DMH

[CPR^{+}18] Chao Chen, Hadi Pouransari, Sivasankaran Rajamanickam, Erik G. Boman, and Eric Darve. A distributedmemory hierarchical solver for general sparse linear systems. Parallel Computing, 74(??):49–64, ???? 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117302077.

Clement:1997:APP

[CQ97] Mark J. Clement and Michael J. Quinn. Automated performance prediction for scalable parallel computing. Parallel Computing, 23(10): 1405–1420, November 7, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=10&aid=1224.

Cortellessa:2001:CRS

[CQ01] Vittorio Cortellessa and Francesco Quaglia. A checkpointing-recovery scheme for Time Warp parallel simulation. Parallel Computing, 27(9):1227–1252, August 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/35/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/35/26/
article.pdf.

Counilh:1990:EMP

[CR90] M. C. Counilh and J. Roman. Expression for massively parallel algorithmsdescription and illustrative example. Parallel Computing, 16(2–3):239–251, December 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Castells-Rufas:2023:GAL

[CR23] David Castells-Rufas. GPU acceleration of Levenshtein distance computation between long strings. Parallel Computing, 116(??): ??, July 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781912300025X.

Christiaens:2002:BNS

[CRD02] Mark Christiaens, Michiel Ronsse, and Koen De Bosschere. Bounding the number of segment histories during data race detection. Parallel Computing, 28(9):1221–1238, September 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/59/28/
abstract.html.

Cores:2016:ROM

[CRGM16] Iván Cores, Mónica Rodríguez, Patricia González, and María J. Martín. Reducing the overhead of an MPI application-level migration approach. Parallel Computing, 54(??):72– 82, May 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116000429.

Caminero:2013:PBR

[CRGR^{+}13] Agustín C. Caminero, Antonio Robles-Gómez, Salvador Ros, Roberto Hernández, and Llanos Tobarra. P2P-based resource discovery in dynamic grids allowing multi-attribute and range queries. Parallel Computing, 39(10): 615–637, October 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0167819113000951.

Crockett:1997:IPR

[Cro97] Thomas W. Crockett. An introduction to parallel rendering. Parallel Computing, 23(7):819–843, July 14, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=7&aid=1186.

Clint:1988:TCE

[CRS88] M. Clint, D. Roantree, and A. Stewart. Towards the construction of an eigenvalue engine. Parallel Computing, 8(1–3): 127–132, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the International Conference on Vector and Parallel Processors in Computational Science, III (Liverpool, 1987).

Cremonesi:1999:PEP

[CRSS99] Paolo Cremonesi, Emilia Rosti, Giuseppe Serazzi, and Evgenia Smirni. Performance evaluation of parallel systems. Parallel Computing, 25(13– 14):1677–1698, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/36/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/30/
article.pdf.

Cosnard:1989:ESD

[CRT89] M. Cosnard, Y. Robert, and B. Tourancheau. Evaluating speedups on distributed memory architectures. Parallel Computing, 10(2):247–253, April 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cheng:1987:VSB

[CS87] K. H. Cheng and S. Sahni. VLSI systems for band matrix multiplication. Parallel Computing, 4(3):239– 258, June 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Carlson:1989:ASE

[CS89a] David A. Carlson and Binay Sugla. Adapting shuffle-exchange like parallel processing organizations to work as systolic arrays. Parallel Computing, 11(1):93–106, July 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cheng:1989:VAB

[CS89b] Kam Hoi Cheng and S. Sahni. VLSI architectures for back substitution. Parallel Computing, 12(1):53–69, October 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cheng:1989:NVS

[CS89c] Kam-Hoi Cheng and Sartaj Sahni. A new VLSI system for adaptive recursive filtering. Parallel Computing, 10(1):109–115, March 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cap:1993:EPC

[CS93] C. H. Cap and V. Strumpen.

Efficient parallel computing in distributed workstation environments. Parallel Computing, 19(11): 1221–1234, November 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chan:1995:PCD

[CS95] Tony F. Chan and Jian Ping Shao. Parallel complexity of domain decomposition methods and optimal coarse grid size. Parallel Computing, 21 (7):1033–1049, July 11, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=7&aid=988.

Chronopoulos:1996:PIS

[CS96] A. T. Chronopoulos and C. D. Swanson. Parallel iterative S-step methods for unsymmetric linear systems. Parallel Computing, 22(5):623–641, August 8, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.

                                        cgi?year=1996&volume=
                                        22&issue=5&aid=1073.

Chen:1997:TFI

[CS97a] Yuh-Shyan Chen and Jang-Ping Sheu. Tolerating faults in injured hypercubes using maximal faultfree subcube-ring. Parallel Computing, 23(3):311–331, May 15, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=3&aid=1107.

Cho:1997:SVS

[CS97b] Jinsung Cho and Heonshik Shin. Scheduling video streams in a large-scale video-on-demand server. Parallel Computing, 23 (12):1743–1755, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=12&aid=1243.

Campos:2000:RCL

[CS00] Luis Miguel Campos and Isaac D. Scherson. Rate of change load balancing in distributed and parallel systems. Parallel Computing, 26(9):1213–1230, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/30/31/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/30/31/
article.pdf.

Chen:2019:EED

[CS19] Shaolong Chen and Miquel Angel Senar. Exploring efficient data parallelism for genome read mapping on multicore and manycore architectures. Parallel Computing, 87(??):11–24, September 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118302564.

ChassindeKergommeaux:2000:PIV

[CSB00] J. Chassin de Kergommeaux, B. Stein, and P. E. Bernard. Pajé, an interactive visualization tool for tuning multithreaded parallel applications. Parallel Computing, 26(10):1253–1274, August 15, 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/31/24/
abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/10/35/21/42/31/24/
                                        article.pdf.

Campbell:2003:SHA

[CSEK03] James F. Campbell, Gary Stiehr, Andreas T. Ernst, and Mohan Krishnamoorthy. Solving hub arc location problems on a cluster of workstations. Parallel Computing, 29(5):555– 574, May 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chiu:2011:DRA

[CSH^{+}11] Yung-Chang Chiu, Ce-Kuen Shieh, Tzu-Chi Huang, Tyng-Yeu Liang, and Kuo-Chih Chu. Data race avoidance and replay scheme for developing and debugging parallel programs on distributed shared memory systems. Parallel Computing, 37(1): 11–25, January 2011. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cibej:2005:CSD

[ČSR05] Uroš Čibej, Boštjan Slivnik, and Borut Robič. The complexity of static data replication in data grids. Parallel Computing, 31 (8–9):900–912, August/ September 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2002:EAC

[CST02] Shao Dong Chen, Hong Shen, and Rodney Topor. An efficient algorithm for constructing Hamiltonian paths in meshes. Parallel Computing, 28 (9):1293–1305, September 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/59/32/
abstract.html.

Cooper:2001:ESO

[CSW01] A. Cooper, M. Szularz, and J. Weston. External selective orthogonalization for the Lanczos algorithm in distributed memory environments. Parallel Computing, 27(7):913– 923, June 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/31/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/31/27/
article.pdf.

Cochrane:1988:SPN

[CT88] David L. Cochrane and Donald G. Truhlar. Strategies and performance norms for efficient utilization of vector pipeline computers as illustrated by the classical mechanical simulation of rotationally inelastic collisions. Parallel Computing, 6(1):63–85, January 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Conner:1994:SPH

[CT94] Michael Conner and Richard Tolimieri. Special purpose hardware for Discrete Fourier Transform implementation. Parallel Computing, 20(2): 215–232, February 24, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=2&aid=829.

Cojean:2022:GML

[CTA22] Terry Cojean, Yu-Hsiang Mike Tsai, and Hartwig Anzt. Ginkgo — a math library designed for platform portability. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819122000096.

Cannataro:2002:PDI

[CTS02] Mario Cannataro, Domenico Talia, and Pradip K. Srimani. Parallel data intensive computing in scientific and commercial applications. Parallel Computing, 28(5):673–704, May 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/57/28/
abstract.html.

Cosnard:1989:SGJ

[CTT89] Michel Cosnard, Maurice Tchuente, and Bernard Tourancheau. Systolic Gauss–Jordan elimination for dense linear systems. Parallel Computing, 10(1): 117–122, March 1989. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2014:ABS

[CTW14] Ray-Bing Chen, Yaohung M. Tsai, and Weichung Wang. Adaptive block size for dense QR factorization in hybrid CPU–GPU systems via statistical modeling. Parallel Computing, 40(5–6): 70–85, May 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000313.

Chavez:2018:ACR

[CTZ^{+}18] Gustavo Chávez, George Turkiyyah, Stefano Zampini, Hatem Ltaief, and David Keyes. Accelerated Cyclic Reduction: A distributedmemory fast solver for structured linear systems. Parallel Computing, 74(??):65–83, ???? 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117302041.

Caron:2004:PPF

[CU04] Eddy Caron and Gil Utard. On the performance of parallel factorization of out-of-core matrices. Parallel Computing, 30(3):357–375, March 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chang:1988:PHT

[CUSR88] H. Y. Chang, S. Utku, M. Salama, and D. Rapp. A parallel Householder tridiagonalization strategem using scattered square decomposition. Parallel Computing, 6(3):297–311, March 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chaudron:1998:FDP

[Cv98] Michel R. V. Chaudron and Arno C. N. van Duin. The formal derivation of parallel triangular system solvers using a coordination-based design method. Parallel Computing, 24(7):1023–1046, July 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/7/1318.pdf.

Coppola:2002:HPD

[CV02] Massimo Coppola and Marco Vanneschi. Highperformance data mining with skeleton-based structured parallel programming. Parallel Computing, 28(5):793–813, May 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/57/33/
abstract.html.

Cools:2017:CHP

[CV17] S. Cools and W. Vanroose.

The communication-hiding pipelined BiCGstab method for the parallel solution of large unsymmetric linear systems. Parallel Computing, 65(??):1–20, July 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0167819117300406.

Chronopoulos:1997:PAE

[CW97] Anthony Theodore Chronopoulos and Gang Wang. Practical aspects and experiences. parallel solution of a traffic flow simulation problem. Parallel Computing, 22(14):1965–1983, March 24, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        22&issue=14&aid=1121.

Chrabakh:2006:GSS

[CW06] Wahid Chrabakh and Rich Wolski. GridSAT: a system for solving satisfiability problems using a computational grid. Parallel Computing, 32(9):660–687, October 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Clint:1992:CTF

[CWB92] M. Clint, J. S. Weston, and C. W. Bleakney. A comparison of two Fortran dialects for expressing parallel solutions for a problem in linear algebra. Parallel Computing, 18(12):1325–1333, December 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2000:MCW

[CWC00] Tzung-Shi Chen, Nen-Chung Wang, and Chih-Ping Chu. Multicast communication in wormholerouted star graph interconnection networks. Parallel Computing, 26 (11):1459–1490, October 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/32/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/32/26/
article.pdf.

Cotting:2007:WSD

[CWDG07] D. Cotting, M. Waschbüsch, M. Duller, and M. Gross. WinSGL: synchronizing displays in parallel graphics using cost-effective software genlocking. Parallel Computing, 33(6):420– 437, June 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2009:OPD

[CWK09] Ching-Wen Chen, Chuan-Chi Weng, and Chang-Jung Ku. An overlapping and pipelining data transmission MAC protocol with multiple channels in ad hoc networks. Parallel Computing, 35(6):313– 330, June 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cai:2021:GAP

[CYDJ21] Linchao Cai, Junrong Yang, Shoubin Dong, and Zhenyu Jiang. GPU accelerated parallel reliabilityguided digital volume correlation with automatic seed selection based on 3D SIFT. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000727.

Che:2018:PSC

[CYXL18] Yonggang Che, Meifang Yang, Chuanfu Xu, and Yutong Lu. Petascale scramjet combustion simulation on the Tianhe-2 heterogeneous supercomputer. Parallel Computing, 77(??):101–117, September 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911830190X.

Chajakis:1991:SAI

[CZ91] E. D. Chajakis and S. A.

Zenios. Synchronous and asynchronous implementations of relaxation algorithms for nonlinear network optimization. Parallel Computing, 17(8):873– 894, October 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2012:ALF

[CZJS12] Yong Chen, Huaiyu Zhu, Hui Jin, and Xian-He Sun. Algorithm-level feedbackcontrolled adaptive data prefetcher: Accelerating data access for highperformance processors. Parallel Computing, 38 (10–11):533–551, October/ November 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000488.

Cai:1999:IAT

[CZTS99] W. Cai, K. Zhang, S. J.

Turner, and C. Sun. Interlock avoidance in transparent and dynamic parallel program instrumentation using logical clocks. Parallel Computing, 25(5):569– 591, May 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/5/1402.pdf.

daMotaTenorio:1990:TSN

[da 90] M. F. da Mota Tenorio. Topology synthesis networks: self-organization of structure and weight adjustment as a learning paradigm. Parallel Computing, 14(3):363–380, August 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Danelutto:2006:ASM

[DA06] Marco Danelutto and Marco Aldinucci. Algorithmic skeletons meeting grids. Parallel Computing, 32(7–8):449–462, September 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Delves:1994:DIP

[DAA94] L. M. Delves, C. A. Addison, and O. A. Aziz. The design and implementation of a portable parallel numerical library. Parallel Computing, 20(10– 11):1639–1651, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=10-11&aid=919.

Dubey:2011:PAM

[DAD11] Anshu Dubey, Katie Antypas, and Christopher Daley. Parallel algorithms for moving Lagrangian data on block structured Eulerian meshes. Parallel Computing, 37(2):101–113, February 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dag:2007:AIP

[Daḡ07]  Hasan Daḡ.  An ap-

proximate inverse preconditioner and its implementation for conjugate gradient method. Parallel Computing, 33(2):83–91, March 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dubey:2009:ECB

[DAG^{+}09] Anshu Dubey, Katie Antypas, Murali K. Ganapathy, Lynn B. Reid, Katherine Riley, Dan Sheeler, Andrew Siegel, and Klaus Weide. Extensible component-based architecture for FLASH, a massively parallel, multiphysics simulation code. Parallel Computing, 35 (10–11):512–522, October/ November 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Darte:2000:CLF

[Dar00]  Alain Darte.  On the

complexity of loop fusion. Parallel Computing, 26(9):1175–1193, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/42/30/29/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/30/29/
                                        article.pdf.

Davies:1988:BEM

[Dav88] A. J. Davies. The boundary element method on the ICL DAP. Parallel Computing, 8(1–3):335– 343, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Drummond:2003:RCM

[DB03] Lúcia M. A. Drummond and Valmir C. Barbosa. On reducing the complexity of matrix clocks. Parallel Computing, 29(7):895– 905, July 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Duro:2017:EEF

[DBI^{+}17] Francisco Rodrigo Duro, Javier Garcia Blas, Florin Isaila, Jesus Carretero, Justin M. Wozniak, and Rob Ross. Experimental evaluation of a flexible I/O architecture for accelerating workflow engines in ultrascale environments. Parallel Computing, 61(??):52–67, January 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116301053.

DiMartino:2001:PPP

[DBVS01] B. Di Martino, S. Briguglio, G. Vlad, and P. Sguazzero. Parallel PIC plasma simulation through particle decomposition techniques. Parallel Computing, 27(3):295–314, February 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/22/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/22/27/
article.pdf.

Dai:2019:VDB

[DBZ^{+}19] Dong Dai, Forrest Sheng Bao, Jiang Zhou, Xuanhua Shi, and Yong Chen. Vectorizing disks blocks for efficient storage system via deep learning. Parallel Computing, 82(??):75–90, ???? 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118300802.

Davidovic:2015:PLS

[DC15] Tatjana Davidović and Teodor Gabriel Crainic. Parallel Local Search to schedule communicating tasks on identical processors. Parallel Computing, 48(??):1–14, October 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000630.

deBiase:1989:VAA

[dCC89] G. de Biase, P. Ciucci, and M. Cottone. Vectorized algorithms for astronomical image processing. Parallel Computing, 10(3):339– 346, May 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DeGreef:1997:MSR

[DCD97] Eddy De Greef, Francky Catthoor, and Hugo De Man. Memory size reduction through storage order optimization for embedded parallel multimedia applications. Parallel Computing, 23(12):1811– 1837 (or 1811–1838??), December 15, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=12&aid=1247.

DeSutter:1998:USP

[DCDV98] Bjorn De Sutter, Mark Christiaens, Koen De Bosschere, and Jan Van Campenhout. On the use of subword parallelism in medical image processing. Parallel Computing, 24(9– 10):1537–1556, September 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/9-10/1347.pdf.

DantasDeMelo:1990:VMD

[DCG90] J. Dantas De Melo, J. L.

Calvet, and J. M. Garcia. Vectorization and multitasking of dynamic programming in control: experiments on a CRAY-2. Parallel Computing, 13 (3):261–269, March 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Debattista:2007:PSR

[DCG^{+}07] K. Debattista, A. Chalmers, R. Gillibrand, P. Longhurst, G. Mastoropoulou, and V. Sundstedt. Parallel selective rendering of highfidelity virtual environments. Parallel Computing, 33(6):361–376, June 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Diener:2015:CAP

[DCN^{+}15] Matthias Diener, Eduardo H. M. Cruz, Philippe O. A. Navaux, Anselm Busse, and Hans-Ulrich Heiß. Communicationaware process and thread mapping using online communication detection. Parallel Computing, 43(??): 43–63, March 2015. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819115000125.

Dai:2016:ATE

[DCR^{+}16] Dong Dai, Philip Carns, Robert B. Ross, John Jenkins, Nicholas Muirhead, and Yong Chen. An asynchronous traversal engine for graph-based rich metadata management. Parallel Computing, 58(??):140–156, October 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819116300576.

Dabah:2022:EPB

[DCY^{+}22] Adel Dabah, Ibrahim Chegrane, Saïd Yahiaoui, Ahcene Bendjoudi, and Nadia Nouali-Taboudjemat. Efficient parallel branchand-bound approaches for exact graph edit distance problem. Parallel Computing, 114(??):??, December 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000734.

Dave:1987:SMC

[DD87] Ameet K. Dave and Iain S.

Duff. Sparse matrix calculations on the CRAY-2. Parallel Computing, 5(1–2):55–64, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Damm:1989:SDC

[DD89] W. Damm and G. Dohmen.

Specifying distributed computer architectures in AADL. Parallel Computing, 9(2):193–211, January 1989. CODEN PACOEJ.

ISSN 0167-8191 (print), 1872-7336 (electronic).

DAmbra:2002:AEPa

[DDdS02] Pasqua D’Ambra, Marco Danelutto, and Daniela di Serafino. Advanced environments for parallel and distributed computing. Parallel Computing, 28(12):1635–1636, December 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DAmbra:2002:AEPb

[DDdSL02] Pasqua D’Ambra, Marco Danelutto, Daniela di Serafino, and Marco Lapegna. Advanced environments for parallel and distributed applications: a view of current status. Parallel Computing, 28(12):1637–1662, December 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dupros:2010:HPF

[DDF^{+}10] Fabrice Dupros, Florent De Martin, Evelyne Foerster, Dimitri Komatitsch, and Jean Roman. Highperformance finite-element simulations of seismic wave propagation in threedimensional nonlinear inelastic geological media. Parallel Computing, 36 (5–6):308–325, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Das:1990:PGA

[DDP90a] Sajal K. Das, Narsingh Deo, and Sushil Prasad. Parallel graph algorithms for hypercube computers. Parallel Computing, 13(2): 143–158, February 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Das:1990:TMS

[DDP90b] Sajal K. Das, Narsingh Deo, and Sushil Prasad. Two minimum spanning forest algorithms on fixedsize hypercube computers. Parallel Computing, 15(1–3):179–187, September 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

deBruijn:1987:EEP

[de 87] M. A. de Bruijn. EPS: an ‘elementary’ programming system for the Delft Parallel Processor. Parallel Computing, 5(3):323–337, November 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DeAgostino:1995:SCP

[De 95] Sergio De Agostino. Short communication: a parallel decoding algorithm for LZ2 data compression. Parallel Computing, 21(12):1957–1961, December 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=12&aid=1030.

deSturler:1996:PMK

[de 96] E. de Sturler. A performance model for Krylov subspace methods on mesh-based parallel computers. Parallel Computing, 22(1):57–74, February 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=1&aid=1036.

Daldorff:2009:PVM

[DE09] L. K. S. Daldorff and B. Eliasson. Parallelization of a Vlasov–Maxwell solver in four-dimensional phase space. Parallel Computing, 35(2):109–115, February 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Deak:1990:URN

[Deá90] István Deák. Uniform random number generators for parallel computers. Parallel Computing, 15(1–3): 155–164, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Decker:2000:VDS

[Dec00] Thomas Decker. Virtual data space — load balancing for irregular applications. Parallel Computing, 26(13–14):1825–1860, December 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/34/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/34/29/
article.pdf.

Dehn:1995:SSM

[DEGS95] T. Dehn, M. Eiermann, K. Giebermann, and V. Sperling. Structured sparse matrix-vector multiplication on massively parallel SIMD architectures. Parallel Computing, 21 (12):1867–1894, December 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=12&aid=1034.

Diaz:2011:POB

[DEH^{+}11] David Díaz, Francisco José Esteban, Pilar Hernández, Juan Antonio Caballero, Gabriel Dorado, and Sergio Gálvez. Parallelizing and optimizing a bioinformatics pairwise sequence alignment algorithm for many-core architecture. Parallel Computing, 37(4–5):244–259, April/May 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111000329.

DeGloria:1991:BMA

[DF91] A. De Gloria and P. Faraboschi. A Boltzmann Machine approach to code optimization. Parallel Computing, 17(9):969–982, November 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Drake:1995:GEP

[DF95a] J. Drake and I. Foster. Guest editorial: Parallel computing in climate and weather modeling. Parallel Computing, 21(10):1537– ??, November 29, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Drake:1995:ISI

[DF95b] J. Drake and I. Foster. Introduction to the special issue on parallel computing in climate and weather modeling. Parallel Computing, 21(10):1539–1544, November 29, 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Donatiello:1998:GCE

[DF98] Lorenzo Donatiello and Alessandro Fabbri. Generative coordination environments supporting parallel discrete event simulation. Parallel Computing, 24(7):1047–1080, July 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/7/1319.pdf.

deForcrand:1999:MM

[dF99] Philippe de Forcrand. The MultiBoson method. Parallel Computing, 25(10– 11):1341–1355, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/26/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/26/25/
article.pdf.

Diao:2000:ESD

[DF00] Shijun Diao and T. Fujiwara. Evaluation and strategy of different data parallel implementation methods of a stiff chemical non-equilibrium flow solver. Parallel Computing, 26(6):791–804, May 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/42/28/30/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/28/30/
                                        article.pdf.

Dongarra:2013:HQF

[DFH^{+}13] Jack Dongarra, Mathieu Faverge, Thomas Hérault, Mathias Jacquelin, Julien Langou, and Yves Robert. Hierarchical QR factorization algorithms for multi-core clusters. Parallel Computing, 39(4– 5):212–232, April/May 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000100.

Drake:1995:DPS

[DFM^{+}95] John Drake, Ian Foster, John Michalakes, Brian Toonen, and Patrick Worley. Design and performance of a scalable parallel community climate model. Parallel Computing, 21 (10):1571–1591, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=10&aid=1016.

Diekmann:1999:ESN

[DFM99] Ralf Diekmann, Andreas Frommer, and Burkhard Monien. Efficient schemes for nearest neighbor load balancing. Parallel Computing, 25(7):789–812, August 13, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1999&volume=
25&issue=7&aid=1410;
http://www.elsevier.
nl/gej-ng/10/35/21/32/
23/20/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/32/
23/20/article.pdf.

DeGloria:1993:CBM

[DFO93] A. De Gloria, P. Faraboschi, and M. Olivieri. Clustered Boltzmann Machines: Massively parallel architectures for constrained optimization problems. Parallel Computing, 19(2):163–175, February 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Diaz:2019:AOO

[DFP^{+}19] Jose Monsalve Diaz, Kyle Friedline, Swaroop Pophale, Oscar Hernandez, David E. Bernholdt, and Sunita Chandrasekaran. Analysis of OpenMP 4.5 offloading in implementations: Correctness and overhead. Parallel Computing, 89(??):Article 102546, November 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819119301371.

DeGloria:1992:DMP

[DFR92] A. De Gloria, P. Faraboschi, and S. Ridella. A dedicated massively parallel architecture for the Boltzmann machine. Parallel Computing, 18(1):57– 73, January 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dehne:1990:PBB

[DFRC90] F. Dehne, A. G. Ferreira, and A. Rau-Chaplin. Parallel branch and bound on fine-grained hypercube multiprocessors. Parallel Computing, 15(1–3):201– 209, September 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dehne:1994:MPK

[DFRC94] Frank Dehne, Afonso Ferreira, and Andrew Rau-Chaplin. A massively parallel knowledge-base server using a hypercube multiprocessor. Parallel Computing, 20(9): 1369–1382, September 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=9&aid=885.

Dongarra:1991:PLT

[DFRR91] J. Dongarra, M. Furtney, S. Reinhardt, and J. Russell. Parallel loops — a test suite for parallelizing compilers: description and example results. Parallel Computing, 17(10– 11):1247–1255, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DiSanto:2002:CBA

[DFRZ02] M. Di Santo, F. Frattolillo, W. Russo, and E. Zimeo. A componentbased approach to build a portable and flexible middleware for metacomputing. Parallel Computing, 28(12):1789–1810, December 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dehne:1990:NLB

[DG90] F. Dehne and M. Gastaldo. A note on the load balancing problem for coarse grained hypercube dictionary machines. Parallel Computing, 16(1):75– 79, November 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Draper:1994:MCM

[DG94] Jeffrey T. Draper and Joydeep Ghosh. The Mcache: a message-handling mechanism for multicomputer systems. Parallel Computing, 20(9): 1269–1288, September 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=9&aid=892.

DAmbra:1995:CBC

[DG95a] Pasqua D’Ambra and Giulio Giunta. Concurrent banded Cholesky factorization on workstation networks using PVM. Parallel Computing, 21(3):487–494, March 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=3&aid=945.

Desprez:1995:NSC

[DG95b] Frederic Desprez and Marc Garbey. Numerical simulation of a combustion problem on a Paragon machine. Parallel Computing, 21(3):495–508, March 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=3&aid=959.

deDoncker:1998:MIH

[dG98] Elise de Doncker and Ajay Gupta. Multivariate integration on hypercubic and mesh networks. Parallel Computing, 24(8):1223–1244, August 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/8/1333.pdf.

Drozdowski:1999:SDL

[DG99] Maciej Drozdowski and Włodzimierz Glazek. Scheduling divisible loads in a three-dimensional mesh of processors. Parallel Computing, 25(4):381–404, April 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1999/
                                        25/4/1396.pdf.

Dreher:2005:RPM

[DG05] Jürgen Dreher and Rainer Grauer. Racoon: a parallel mesh-adaptive framework for hyperbolic conservation laws. Parallel Computing, 31(8–9):913–932, August/ September 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dias:2015:TMS

[DG15] Sérgio E. D. Dias and Abel J. P. Gomes. Triangulating molecular surfaces over a LAN of GPU-enabled computers. Parallel Computing, 42(??):35–47, February 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114001203.

DeMarle:2005:MSI

[DGBP05] David E. DeMarle, Christiaan P. Gribble, Solomon Boulos, and Steven G. Parker. Memory sharing for interactive ray tracing on clusters. Parallel Computing, 31(2):221–242, February 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Durillo:2019:MOR

[DGKF19] Juan J. Durillo, Philipp Gschwandtner, Klaus Kofler, and Thomas Fahringer. Multi-objective regionaware optimization of parallel programs. Parallel Computing, 83(??):3–21, April 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300905.

Darema:1988:SPM

[DGNP88] F. Darema, D. A. George, V. A. Norton, and G. F. Pfister. A single-program-multiple-data computational model for EPEX/ FORTRAN. Parallel Computing, 7(1):11–24, April 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dumas:2016:RBP

[DGP^{+}16] Jean-Guillaume Dumas, Thierry Gautier, Clément Pernet, Jean-Louis Roch, and Ziad Sultan. Recursion based parallelization of exact dense linear algebra routines for Gaussian elimination. Parallel Computing, 57(??):235–249, September 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001222.

deRosa:1995:PTA

[dGR95] M. A. de Rosa, G. Giunta, and M. Rizzardi. Parallel Talbot’s algorithm for distributed memory machines. Parallel Computing, 21(5):783–801 (or 783– 802??), May 10, 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cgi-bin/cas/tree/
store/parco/cas_sub/
browse/browse.cgi?year=
1995&volume=21&issue=
5&aid=974.

Dongarra:1984:CPL

[DH84] J. J. Dongarra and R. E. Hiromoto. A collection of parallel linear equations routines for the Denelcor HEP. Parallel Computing, 1(2):133–142, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Detert:1991:CXM

[DH91] U. Detert and G. Hofemann. CRAY X-MP and Y-MP memory performance. Parallel Computing, 17(4–5):579–590, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DeRose:2007:DDP

[DHL07] César A. F. De Rose, Hans-Ulrich Heiss, and Barry Linnert. Distributed dynamic processor allocation for multicomputers. Parallel Computing, 33(3):145– 158, April 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dehne:1989:CCS

[DHS89] Frank Dehne, Anne-Lise Hassenklover, and Jörg-Rüdiger Sack. Computing the configuration space for a robot on a meshof-processors. Parallel Computing, 12(2):221–231, November 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Donaldson:2000:BCH

[DHS00] Stephen R. Donaldson, Jonathan M. D. Hill, and David B. Skillicorn. BSP clusters: High performance, reliable and very low cost. Parallel Computing, 26(2–3):199–242, February 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/25/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/25/25/
article.pdf.

Dongarra:1997:KCP

[DHW97] Jack J. Dongarra, Sven Hammarling, and David W. Walker. Key concepts for parallel out-of-core LU factorization. Parallel Computing, 23(1–2):49–70, April 16, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=1-2&aid=1147.

diSerafino:1997:PIM

[di 97] Daniela di Serafino. Parallel implementation of a multigrid multiblock Euler solver on distributed memory machines. Parallel Computing, 23(13): 2095–2113, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=13&aid=1229.

Dhamdhere:1997:DTD

[DIR97] D. M. Dhamdhere, Sridhar R. Iyer, and E. Kishore Kumar Reddy. Distributed termination detection for dynamic systems. Parallel Computing, 22(14): 2025–2045, March 24, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=14&aid=1119.

Dixit:1991:SB

[Dix91]  K. M. Dixit.  The

SPEC benchmarks. Parallel Computing, 17(10– 11):1195–1209, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dongarra:1987:SBS

[DJ87] Jack J. Dongarra and Lennart Johnsson. Solving banded systems on a parallel processor. Parallel Computing, 5(1–2):219– 246, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

DiBlas:2005:ONN

[DJH05] Andrea Di Blas, Arun Jagota, and Richard Hughey. Optimizing neural networks on SIMD parallel computers. Parallel Computing, 31(1):97–115, January 2005. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Deng:2001:PSB

[DK01] Yuefan Deng and Alex Korobka. The performance of a supercomputer built with commodity components. Parallel Computing, 27(1–2):91–108, January 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/47/25/26/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/25/26/
                                        article.pdf.

Dong:2004:DLP

[DK04] Suchuan Dong and George Em Karniadakis. Dual-level parallelism for high-order CFD methods. Parallel Computing, 30(1):1–20, January 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Daoudi:1992:IBE

[DL92] E. M. Daoudi and J. Lobry. Implementation of a boundary element method on distributed memory computers. Parallel Computing, 18(12):1317–1324, December 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Daoudi:1997:ESP

[DL97] El Mostafa Daoudi and Abdelhak Lakhouaja. Exploiting the symmetry in the parallelization of the Jacobi method. Parallel Computing, 23(1– 2):137–151, April 16, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=1-2&aid=1152.

Dolean:2004:PMM

[DL04] V. Dolean and S. Lanteri.

Parallel multigrid methods for the calculation of unsteady flows on unstructured grids: algorithmic aspects and parallel performances on clusters of PCs. Parallel Computing, 30(4):503–525, April 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Du:2005:NPA

[DL05] Z. Du and F. Lin. A novel parallelization approach for hierarchical clustering. Parallel Computing, 31(5): 523–527, May 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Du:2006:PPP

[DL06] Zhihua Du and Feng Lin. pNJTree: a parallel program for reconstruction of neighbor-joining tree and its application in ClustalW. Parallel Computing, 32(5–6):441–446, June 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dorta:2006:BSL

[DLdS06] Antonio Dorta, Pablo López, and Francisco de Sande. Basic skeletons in llc. Parallel Computing, 32(7–8):491–506, September 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DApuzzo:1997:SLB

[DLM97] Marco D’Apuzzo, Marco Lapegna, and Almerico Murli. Scalability and load balancing in adaptive algorithms for multidimensional integration. Parallel Computing, 23 (8):1199–1210, July 25, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=8&aid=1175.

Ditter:2017:RRT

[DLO17] Alexander Ditter, Jan Laukemann, and Benedikt Oehlrich. Reproducibility report: Team SegFAUlt @ SCC 2016. Parallel Computing, 70(??):41–45, December 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301205.

Dontje:1992:SAS

[DLPS92] T. Dontje, Th. Lippert, N. Petkov, and K. Schilling. Statistical analysis of simulationgenerated time series: Systolic vs. semi-systolic correlation on the connection machine. Parallel Computing, 18(5):575–588, May 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DeKeyser:1994:RTL

[DLR94] Johan De Keyser, Kurt Lust, and Dirk Roose. Run-time load balancing support for a parallel multiblock Euler/Navier– Stokes code with adaptive refinement on distributed memory computers. Parallel Computing, 20(8):1069–1088, August 10, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=8&aid=873.

Deng:2009:SPM

[DLS09] Lih-Yuan Deng, Huajiang Li, and Jyh-Jen Horng Shiau. Scalable parallel multiple recursive generators of large order. Parallel Computing, 35(1):29– 37, January 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819108001099.

Douglas:1990:BMP

[DM90] Craig C. Douglas and Willard L. Miranker. Beyond massive parallelism: numerical computation using associative tables. Parallel Computing, 16(1):1– 25, November 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dabdub:1997:PPA

[DM97] Donald Dabdub and Rajit Manohar. Performance and portability of an air quality model. Parallel Computing, 23 (14):2187–2200, December 17, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=14&aid=1266.

Dunn:2002:QFS

[DM02] Ian N. Dunn and Gerard G. L. Meyer. QR factorization for shared memory and message passing. Parallel Computing, 28 (11):1507–1530, November ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/61/26/
abstract.html.

DApuzzo:2003:PCI

[DM03] M. D’Apuzzo and M. Marino.

Parallel computational issues of an interior point method for solving large bound-constrained quadratic programming problems. Parallel Computing, 29(4):467–483, April 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DiMartino:2004:SOS

[DM04] V. Di Martino and M. Mililotti. Sub optimal scheduling in a grid using genetic algorithms. Parallel Computing, 30(5–6):553–565, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dubitzky:2004:GED

[DMG^{+}04] Werner Dubitzky, Damian McCourt, Mykola Galushka, Mathilde Romberg, and Bernd Schuller. Gridenabled data warehousing for molecular engineering. Parallel Computing, 30(9–10):1019–1035, September/October 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Deng:2006:AMS

[DMT06] Xiao Yan Deng, Greg Michaelson, and Phil Trinder. Autonomous mobility skeletons. Parallel Computing, 32(7–8): 463–478, September 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DuCroz:1988:ALB

[DMWW88] J. J. Du Croz, P. J. D.

Mayes, J. Wasniewski, and S. Wilson. Applications of Level 2 BLAS in the NAG library. Parallel Computing, 8(1–3):345–350, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Date:2016:GAH

[DN16] Ketan Date and Rakesh Nagi. GPU-accelerated Hungarian algorithms for the Linear Assignment Problem. Parallel Computing, 57(??):52–72, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911630045X.

deNardin:2021:REP

[dNdRRL^{+}21] Igor Fontana de Nardin, Rodrigo da Rosa Righi, Thiago Roberto Lima Lopes, Cristiano André da Costa, Heon Young Yeom, and Harald Köstler. On revisiting energy and performance in microservices applications: a cloud elasticity-driven approach. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121001010.

Dorostkar:2015:NCA

[DNL15] Ali Dorostkar, Maya Neytcheva, and Björn Lund. Numerical and computational aspects of some block-preconditioners for saddle point systems. Parallel Computing, 49 (??):164–178, November 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000897.

Dent:1988:MCM

[DO88] D. Dent and M. O’Neill. Microtasking as a complement to macrotasking. Parallel Computing, 8(1– 3):149–154, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dodd:1989:GMS

[Dod89] Nigel Dodd. Graph matching by stochastic optimisation applied to the implementation of multi layer perceptrons on transputer networks. Parallel Computing, 10(2):135–142, April 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dodd:1990:SAV

[Dod90] N. Dodd. Slow annealing versus multiple fast annealing runs — an empirical investigation. Parallel Computing, 16(2–3): 269–272, December 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dulger:2023:URP

[DOD23] Özcan Dülger, Halit Oguztüzün, and Mübeccel Demirekler. Uphill resampling for particle filter and its implementation on graphics processing unit. Parallel Computing, 115(??):??, February 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000837.

Dong:2005:LTP

[Don05] Tao Dong. A linear time pessimistic one-step diagnosis algorithm for hypercube multicomputer systems. Parallel Computing, 31(8–9):933–947, August/ September 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dowling:1990:OCP

[Dow90] Michael L. Dowling. Optimal code parallelization using unimodular transformations. Parallel Computing, 16(2–3):157–171, December 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dow:1995:TMV

[Dow95] Murray Dow. Transposing a matrix on a vector computer. Parallel Computing, 21(12):1997–2005, December 12, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=12&aid=1029.

DeMatteis:1990:CPR

[DP90a] A. De Matteis and S. Pagnutti. A class of parallel random number generators. Parallel Computing, 13(2):193–198, February 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

deMatteis:1990:LRC

[dP90b] A. de Matteis and S. Pagnutti. Long-range correlations in linear and nonlinear random number generators. Parallel Computing, 14(2):207–210, June 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DeMatteis:1995:CCP

[DP95] A. De Matteis and S. Pagnutti. Controlling correlations in parallel Monte Carlo. Parallel Computing, 21(1):73–84, January 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=1&aid=939.

Das:2000:PPQ

[DP00] Sajal K. Das and M. Cristina Pinotti. Parallel priority queues based on binomial heaps. Parallel Computing, 26(11):1411–1428, October 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/32/23/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/32/23/
article.pdf.

Dasu:2002:RMP

[DP02] Aravind Dasu and Sethuraman Panchanathan. Reconfigurable media processing. Parallel Computing, 28(7–8):1111–1139, August 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/58/36/
                                        abstract.html.

Diep:2019:TSS

[DPFT19] Thanh-Dang Diep, Kien Trung Pham, Karl Fürlinger, and Nam Thoai. A timestamping system to detect memory consistency errors in MPI one-sided applications. Parallel Computing, 86(??):36–44, August 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118303235.

Diekmann:2000:SOM

[DPSW00] Ralf Diekmann, Robert Preis, Frank Schlimbach, and Chris Walshaw. Shapeoptimized mesh partitioning and load balancing for parallel adaptive FEM. Parallel Computing, 26 (12):1555–1581, November 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/42/33/25/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/33/25/
                                        article.pdf.

Djamegni:2000:DSA

[DQRR00] Clémentin Tayou Djamégni, Patrice Quinton, Sanjay Rajopadhye, and Tanguy Risset. Derivation of systolic algorithms for the algebraic path problem by recurrence transformations. Parallel Computing, 26(11):1429–1445, October 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/32/24/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/32/24/
article.pdf.

DeKeyser:1993:LBD

[DR93] J. De Keyser and D. Roose.

Load balancing data parallel programs on distributed memory computers. Parallel Computing, 19(11):1199–1219, November 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Darte:1994:MUL

[DR94] Alain Darte and Yves Robert. Mapping uniform loop nests onto distributed memory architectures. Parallel Computing, 20(5):679–710, May 11, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/

                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=5&aid=831.

DeKeyser:1995:RTL

[DR95] J. De Keyser and D. Roose. Run-time load balancing techniques for a parallel unstructured multi-grid Euler solver with adaptive grid refinement. Parallel Computing, 21(2): 179–198, February 17, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=2&aid=952.

Dion:1996:MAL

[DR96] Michèle Dion and Yves Robert. Mapping affine loop nests. Parallel Computing, 22(10):1373–1397, December 15, 1996. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.
                                        com/cgi-bin/cas/tree/
                                        store/parco/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=22&issue=
                                        10&aid=1100.

Dumas:2002:PBA

[DR02] Jean-Guillaume Dumas and Jean-Louis Roch. On parallel block algorithms for exact triangularizations. Parallel Computing, 28(11):1531–1548, November ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/61/27/
abstract.html.

Durazzi:2003:NSS

[DR03] C. Durazzi and V. Ruggiero. Numerical solution of special linear and quadratic programs via a parallel interior-point method. Parallel Computing, 29(4):485–503, April 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Diegues:2015:STI

[DR15] Nuno Diegues and Paolo Romano. Self-tuning Intel Restricted Transactional Memory. Parallel Computing, 50(??):25–52, December 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001209.

Davina:2018:MCP

[DR18] A. Lamas Daviña and J. E. Roman. MPI-CUDA parallel linear solvers for block-tridiagonal matrices in the context of SLEPc’s eigensolvers. Parallel Computing, 74(??):118– 135, ???? 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301874.

Delaitre:1997:GTS

[DRJSW97] T. Delaitre, G. R. Ribeiro-Justo, F. Spies, and S. C. Winter. A graphical toolset for simulation modelling of parallel systems. Parallel Computing, 22 (13):1823–1836, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=13&aid=1130.

Diaz:2003:DIP

[DRST03] Manuel Díaz, Bartolomé Rubio, Enrique Soler, and José M. Troya. Domain interaction patterns to coordinate HPF tasks. Parallel Computing, 29(7):925– 951, July 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dongarra:1984:SPB

[DS84] Jack J. Dongarra and Ahmed H. Sameh. On some parallel banded system solvers. Parallel Computing, 1(3–4):223–235, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dongarra:1987:PED

[DS87] J. J. Dongarra and D. C.

Sorensen. A portable environment for developing parallel FORTRAN programs. Parallel Computing, 5(1–2):175–186, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Duller:1992:SVA

[DS92] A. W. G. Duller and R. Storer. Simulation and verification of associative processor arrays. Parallel Computing, 18(12):1403– 1414, December 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dhrif:1993:FAS

[DS93] H. Dhrif and D. Sarkar.

Fuzzy arithmetic on systolic arrays. Parallel Computing, 19(11):1283–1301, November 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dabdub:1996:PCA

[DS96] Donald Dabdub and John H.

Seinfeld. Parallel computation in atmospheric chemical modeling. Parallel Computing, 22(1): 111–130, February 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=1&aid=1042.

Dattilo:2003:SCL

[DS03] Giuseppe Dattilo and Giandomenico Spezzano. Simulation of a cellular landslide model with CAMELOT on high performance computers. Parallel Computing, 29(10): 1403–1418, October 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Duff:2005:SBB

[DS05] Iain S. Duff and Jennifer A. Scott. Stabilized bordered block diagonal forms for parallel sparse solvers. Parallel Computing, 31(3–4):275– 289, March/April 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DAmbrosio:2007:PEM

[DS07] D. D’Ambrosio and W. Spataro. Parallel evolutionary modelling of geological processes. Parallel Computing, 33(3):186–212, April 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dang:2013:CES

[DS13] Hoang-Vu Dang and Bertil Schmidt. CUDA-enabled sparse matrix-vector multiplication on GPUs using atomic operations. Parallel Computing, 39 (11):737–750, November 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001178.

Dongarra:1988:PMP

[DSCP88] J. J. Dongarra, D. C.

Sorensen, K. Connolly, and J. Patterson. Programming methodology and performance issues for advanced computer architectures. Parallel Computing, 8(1–3):41–58, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dang:2017:ECB

[DSG17] Hoang-Vu Dang, Marc Snir, and William Gropp. Eliminating contention bottlenecks in multithreaded MPI. Parallel Computing, 69(??):1–23, November 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117301187.

Dongarra:1986:ISC

[DSS86] J. J. Dongarra, A. H. Sameh, and D. C. Sorensen. Implementation of some concurrent algorithms for matrix factorization. Parallel Computing, 3(1):25– 34, March 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

daSilva:2009:ISB

[dSS09] Fabrício A. B. da Silva and Hermes Senger. Improving scalability of Bagof-Tasks applications running on master-slave platforms. Parallel Computing, 35(2):57–71, February 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dreier:2018:SSC

[DSSD18] Lisa Marie Dreier, Svilen Stefanov, David Schneller, and Alexander Ditter. SC17 student cluster competition, Team Technical University of Munich and Friedrich-Alexander University Erlangen–Nürnberg: Reproducing vectorization of the Tersoff multibody potential on the Intel Broadwell architecture. Parallel Computing, 78(??):79–83, October 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300644.

Dongarra:2001:CCG

[DST01] Jack Dongarra, Masaaki Shimasaki, and Bernard Tourancheau. Clusters and computational grids for scientific computing. Parallel Computing, 27 (11):1401–1402, October 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/41/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/41/27/
article.pdf.

Falco:2015:MTC

[DST15] Ivanoe De Falco, Umberto Scafuri, and Ernesto Tarantino. Mapping of time-consuming multitask applications on a cloud system by multiobjective Differential Evolution. Parallel Computing, 48(??):40–58, October 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000629.

Desprez:1995:BRR

[DT95] F. Desprez and B. Tourancheau. Basic routines for the rank-2k update: 2D torus vs. reconfigurable network. Parallel Computing, 21(3):353–372, March 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=3&aid=960.

Djordjevic:1996:HST

[DT96] Goran Lj. Djordjević and Milorad B. Tošić. A heuristic for scheduling task graphs with communication delays onto multiprocessors. Parallel Computing, 22(9): 1197–1214, November 22, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=9&aid=1092.

Dongarra:1997:WET

[DT97] J. Dongarra and B. Tourancheau. Workshop on environments and tools for parallel scientific computing. Parallel Computing, 23(1– 2):1–??, April 16, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Negro:2004:PPP

[dTNOR^{+}04] F. de Toro Negro, J. Ortega, E. Ros, S. Mota, B. Paechter, and J. M. Martín. PSFGA: Parallel processing and evolutionary computation for multiobjective optimisation. Parallel Computing, 30 (5–6):721–739, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Deveci:2018:MSM

[DTR18] Mehmet Deveci, Christian Trott, and Sivasankaran Rajamanickam. Multithreaded sparse matrix– matrix multiplication for many-core and GPU architectures. Parallel Computing, 78(??):33–46, October 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301923.

Dad:2021:SFD

[DTV21] Cherifa Dad, Jean-Philippe Tavella, and Stéphane Vialle. Synthesis and feedback on the distribution and parallelization of FMI-CS-based co-simulations with the DACCOSIM platform. Parallel Computing, 106(??):??, September 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000545.

Duff:1984:SE

[Duf84] Iain S. Duff. Supercomputers in Europe. Parallel Computing, 1(3–4): 321–324, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Duff:1986:PIM

[Duf86] Iain S. Duff. Parallel implementation of multifrontal schemes. Parallel Computing, 3(3):193–204, July 1986. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Drummond:2006:GED

[DUG^{+}06] Lúcia M. A. Drummond, Eduardo Uchoa, Alexandre D. Gonçalves, Juliana M. N. Silva, Marcelo C. P. Santos, and Maria Clícia S. de Castro. A grid-enabled distributed branch-and-bound algorithm with application on the Steiner Problem in graphs. Parallel Computing, 32(9):629– 642, October 2006. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dunigan:1991:PII

[Dun91] T. H. Dunigan. Performance of the Intel iPSC/ 860 and Ncube 6400 hypercubes. Parallel Computing, 17(10–11):1285–1302, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

dAcierno:1994:PRN

[dV94] Antonio d’Acierno and Roberto Vaccaro. On parallelizing recursive neural networks on coarsegrained parallel computers: a general algorithm. Parallel Computing, 20 (2):245–256, February 24, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=2&aid=823.

Dongarra:1992:RCF

[DvdG92] Jack J. Dongarra and Robert A. van de Geijn. Reduction to condensed form for the eigenvalue problem on distributed memory architectures. Parallel Computing, 18(9):

973–982, September 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Duff:1999:DTP

[DvdV99] Iain S. Duff and Henk A. van der Vorst. Developments and trends in the parallel solution of linear systems. Parallel Computing, 25(13–14):1931–1970, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/36/39/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/36/39/
                                        article.pdf.

DiazdeCerio:1998:MEC

[DVGG98] Luis Díaz de Cerio, Miguel Valero-García, and Antonio González. Method for exploiting communication/computation overlap in hypercubes. Parallel Computing, 24(2):221–245, February 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/2/1282.pdf.

Das:1990:SAH

[DVP90] S. R. Das, N. H. Vaidya, and L. M. Patnaik. A systolic algorithm for hidden surface removal. Parallel Computing, 15(1–3): 277–289, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Doi:1999:OSR

[DW99] Shun Doi and Takumi Washio. Ordering strategies and related techniques to overcome the trade-off between parallelism and convergence in incomplete factorizations. Parallel Computing, 25(13–14):1995–2014, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/36/41/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/41/
article.pdf.

Durand:2000:TAS

[DW00] M. D. Durand and Steve R.

White. Trading accuracy for speed in parallel simulated annealing with simultaneous moves. Parallel Computing, 26(1):135–150, January 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/23/30/
abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/10/35/21/42/23/30/
                                        article.pdf.

Dodson:1999:PIH

[DWB99] S. J. Dodson, S. P. Walker, and M. J. Bluck. Parallelisation issues for high speed time domain integral equation analysis. Parallel Computing, 25 (8):925–942, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/32/24/17/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/24/17/
                                        article.pdf.

Deng:2008:DSS

[DWH^{+}08] Yuhui Deng, Frank Wang, Na Helian, Sining Wu, and Chenhan Liao. Dynamic and scalable storage management architecture for Grid Oriented Storage devices. Parallel Computing, 34(1):17–31, January 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Du:2012:COT

[DWL^{+}12] Peng Du, Rick Weber, Piotr Luszczek, Stanimire Tomov, Gregory Peterson, and Jack Dongarra. From CUDA to OpenCL: Towards a performanceportable solution for multiplatform GPU programming. Parallel Computing, 38(8):391–407, August 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001335.

Dmitruk:2001:SPF

[DWM^{+}01] P. Dmitruk, L.-P. Wang, W. H. Matthaeus, R. Zhang, and D. Seckel. Scalable parallel FFT for spectral simulations on a Beowulf cluster. Parallel Computing, 27(14):1921–1936, December 31, 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/44/31/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/44/31/
article.pdf.

Dosanjh:2021:IEM

[DWS^{+}21] Matthew G. F. Dosanjh, Andrew Worley, Derek Schafer, Prema Soundararajan, Sheikh Ghafoor, Anthony Skjellum, Purushotham V. Bangalore, and Ryan E. Grant. Implementation and evaluation of MPI 4.0 partitioned communication libraries. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000752.

Deng:2023:PEL

[DWS^{+}23] Lih-Yuan Deng, Bryan R. Winter, Jyh-Jen Horng Shiau, Henry Horng-Shing Lu, Nirman Kumar, and Ching-Chi Yang. Parallelizable efficient large order multiple recursive generators. Parallel Computing, 117(??):??, September 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781912300042X.

Di:2012:PSG

[DWX^{+}12] Peng Di, Hui Wu, Jingling Xue, Feng Wang, and Canqun Yang. Parallelizing SOR for GPG-PUs using alternate loop tiling. Parallel Computing, 38(6–7):310–328, June/July 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781911200021X.

Das:1991:PPS

[DY91] Sajal K. Das and Cui-Qing Yang. Performance of parallel spanning tree algorithms on linear arrays of transputers and Unix systems. Parallel Computing, 17(4–5):527–551, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Doallo:1990:VSA

[DZ90] R. Doallo and E. L. Zapata. A VLSI systolic architecture for solving DBT-Transformed fuzzy clustering problems of arbitrary size. Parallel Computing, 13(3):321–335, March 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dymond:2001:DPC

[DZD01] Patrick Dymond, Jieliang Zhou, and Xiaotie Deng. A 2-D parallel convex hull algorithm with optimal communication phases. Parallel Computing, 27(3):243– 255, February 2001. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/47/
22/24/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/47/
22/24/article.pdf.

Dubey:1994:GPS

[DZG94] A. Dubey, M. Zubair, and C. E. Grosch. A general purpose subroutine for fast Fourier transform on a distributed memory parallel machine. Parallel Computing, 20 (12):1697–1710, November 28, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=12&aid=928.

Dong:2020:EGA

[DZLK20] Shi Dong, Pu Zhao, Xue Lin, and David Kaeli. Exploring GPU acceleration of deep neural networks using block circulant matrices. Parallel Computing, 100(??): Article 102701, December 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300909.

Deng:2013:ALP

[DZM^{+}13] Yuefan Deng, Peng Zhang, Carlos Marques, Reid Powell, and Li Zhang. Analysis of Linpack and power efficiencies of the world’s TOP500 supercomputers. Parallel Computing, 39(6–7):271–279, June/July 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000513.

Dzwinel:1991:SOM

[Dzw91] W. Dzwinel. The search for an optimal multiprocessor interconnection network. Parallel Computing, 17(1):95–100, April 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1995:SAD

[EA95] D. J. Evans and S. A.

Amin. Systolic algorithms for digital image filtering. Parallel Computing, 21(1):109–119, January 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=1&aid=930.

Evans:1990:SSP

[EAKT90] D. J. Evans, M. Adamopoulos, S. Kortesis, and K. Tsouros. Searching sets of properties with neural networks. Parallel Computing, 16(2–3):279–285, December 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

El-Amawy:1993:SSG

[EAR93] A. El-Amawy and R. Raja. Split sequence generation algorithms for efficient identification of operational subcubes in faulty hypercubes. Parallel Computing, 19(7):789– 805, July 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1988:SLS

[EB88] D. J. Evans and M. P. Bekakos. The solution of linear systems by the QIF algorithm on a wavefront array processor. Parallel Computing, 7(1):111–130, April 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1993:DLB

[EB93] D. J. Evans and W. U. N. Butt. Dynamic load balancing using task-transfer probabilities. Parallel Computing, 19(8):897–916, August 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1994:LBN

[EB94] D. J. Evans and W. U. N. Butt. Load balancing with network partitioning using host groups. Parallel Computing, 20(3):325–345, March 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=3&aid=836.

Evans:1998:BLS

[EB98] D. J. Evans and M. Barulli.

BSP linear solvers for dense matrices. Parallel Computing, 24(5–6):777– 795, June 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/5-6/1291.pdf.

Efe:1994:TPC

[EBSS94] Kemal Efe, P. K. Blackwell, W. Slough, and T. Shiau. Topological properties of the crossed cube architecture. Parallel Computing, 20(12): 1763–1775, November 28, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=12&aid=899.

Evans:1991:AGE

[EC91] D. J. Evans and S. Chikohora. The alternating group explicit (AGE) method on a transputer network. Parallel Computing, 17(6–7):833–843, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Edmonds:1993:DPS

[ECG93] P. Edmonds, E. Chu, and A. George. Dynamic programming on a sharedmemory multiprocessor. Parallel Computing, 19(1): 9–22, January 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Eckardt:1990:SPE

[Eck90] H. Eckardt. System performance and execution of scientific algorithms on the parallel computer parawell. Parallel Computing, 13(2):159–173, February 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Etinski:2012:PJS

[ECLV12] M. Etinski, J. Corbalan, J. Labarta, and M. Valero. Parallel job scheduling for power constrained HPC systems. Parallel Computing, 38(12):615–630, December 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000610.

Evans:1991:APA

[ED91] D. J. Evans and Wang Deren. An asynchronous parallel algorithm for solving a class of nonlinear simultaneous equations. Parallel Computing, 17 (2–3):165–180, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Eeckhout:2004:EST

[ED04] Lieven Eeckhout and Koen De Bosschere. Efficient simulation of trace samples on parallel machines. Parallel Computing, 30(3): 317–335, March 2004. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ellis:2019:OPA

[EEH^{+}19] J. Austin Ellis, Thomas M.

Evans, Steven P. Hamilton, C. T. Kelley, and Tara M. Pandya. Optimization of processor allocation for domain decomposed Monte Carlo calculations. Parallel Computing, 87(??):77–86, September 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301243.

Exum:1990:NDA

[EG90] M. R. Exum and J. L. Gaudiot. Network design and allocation considerations in the Hughes data-flow machine. Parallel Computing, 13(1):17– 34, January 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Erhard:1992:IPA

[EG92a] W. Erhard and A. Grefe. Improved parallel algorithms for the classification of electroencephalograms (EEGs) on the DAP510. Parallel Computing, 18(1): 109–115, January 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1992:IFT

[EG92b] D. J. Evans and M. Gusev. Implementation of folding transformations on linear VLSI processor arrays. Parallel Computing, 18(5):525–542, May 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1994:PAP

[EG94a] D. J. Evans and E. Galligani. A parallel additive preconditioner for conjugate gradient method for AX + XB = C. Parallel Computing, 20 (7):1055–1064, July 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=7&aid=879.

Evans:1994:NLS

[EG94b] D. J. Evans and M. Gusev. New linear systolic arrays for digital filters and convolution. Parallel Computing, 20(1):29–61, January 16, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=1&aid=822.

Evripidou:1995:IIO

[EG95] Paraskevas Evripidou and Jean-Luc Gaudiot. Incorporating input/output operations into dynamic data-flow graphs. Parallel Computing, 21(8): 1285–1311, August 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=8&aid=1000.

Ehold:2002:OLP

[EGKU02] Harald J. Ehold, Wilfried N. Gansterer, Dieter F. Kvasnicka, and Christoph W. Ueberhuber. Optimizing local performance in HPF. Parallel Computing, 28(3):415–432, March 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/34/29/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/60/34/29/
                                        main.pdf.

Edjlali:1999:IPP

[EGTD99] G. Edjlali, M. Garbey, and D. Tromeur-Dervout. Interoperability parallel programs approach to simulate 3D frontal polymerization processes. Parallel Computing, 25(9): 1161–1191, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/32/25/22/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/25/22/
                                        article.pdf.

El-Horbaty:1992:SAS

[EHE92] El-Sayed M. El-Horbaty and A. El-Din H. Mohamed. A synchronous algorithm for shortest paths on a tree machine. Parallel Computing, 18(1):103–107, January 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Eldredge:1997:HPP

[EHF^{+}97] Michael Eldredge, Thomas J. R. Hughes, Robert M. Ferencz, Steven M. Rifai, Arthur Raefsky, and Bruce Herndon. Highperformance parallel computing in industry. Parallel Computing, 23(9): 1217–1233, November 3, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=9&aid=1207.

Ewinger:1989:MMM

[EHHS89] W. Ewinger, O. Haan, E. Haupenthal, and C. Siemers. Modelling and measurement of memory access in SIEMENS VP supercomputers. Parallel Computing, 11(3):361–365, ???? 28, 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Edwards:1999:CFL

[EHN99] Robert G. Edwards, Urs M.

Heller, and Rajamani Narayanan. Chiral fermions on the lattice. Parallel Computing, 25(10– 11):1395–1407, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/26/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/26/29/
article.pdf.

Eickermann:1998:MGE

[EHR^{+}98] Th Eickermann, J. Henrichs, M. Resch, R. Stoy, and R. Volpel. Metacomputing in gigabit environments: networks, tools, and applications. Parallel Computing, 24(12– 13):1847–1872, November 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/12-13/1358.pdf.

Eisenstat:1989:CSP

[Eis89] Stanley C. Eisenstat. Comments on scheduling parallel iterative methods on multiprocessor systems. II. Parallel Computing, 11 (2):241–244, ???? 4, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1997:ACG

[EJLC97] E. W. Evans, S. P. Johnson, P. F. Leggett, and M. Cross. Automatic code generation of overlapped communications in a parallelisation tool. Parallel Computing, 23 (10):1493–1523, November 7, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=10&aid=1227.

Evans:2000:AEM

[EJLC00] E. W. Evans, S. P.

Johnson, P. F. Leggett, and M. Cross. Automatic and effective multidimensional parallelisation of structured mesh based codes. Parallel Computing, 26(6):677–703, May 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/28/23/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/28/23/
article.pdf.

Egecioglu:1992:PAG

[EK92] Ömer Eğecioğlu and Çetin K.

Koç. A parallel algorithm for generating discrete orthogonal polynomials. Parallel Computing, 18(6):649–659, June 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Everaars:1998:UCP

[EK98] Kees Everaars and Barry Koren. Using coordination to parallelize sparsegrid methods for 3-D CFD problems. Parallel Computing, 24(7):1081–1106, July 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/7/1320.pdf.

Eklund:2004:MPA

[Ekl04] Sven E. Eklund. A massively parallel architecture for distributed genetic algorithms. Parallel Computing, 30(5–6):647–676, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

ElBaz:1993:AIR

[El 93] D. El Baz. Asynchronous implementation of relaxation and gradient algorithms for convex network flow problems. Parallel Computing, 19(9):1019– 1028, September 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1990:SUG

[EL90] D. J. Evans and C. Li.

Successive underrelaxation (SUR) and generalised conjugate gradient (GCG) methods for hyperbolic difference equations on a parallel computer. Parallel Computing, 16(2–3):207– 220, December 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1991:MSV

[EL91] D. J. Evans and M. D.

Levin. A matrix-squaring variant of the power method on the DAP. Parallel Computing, 17(1):49– 54, April 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Elman:1995:ULA

[EL95] Howard C. Elman and Dennis K.-Y. Lee. Use of linear algebra kernels to build an efficient finite element solver. Parallel Computing, 21(1):161–173, January 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.

                                        cgi?year=1995&volume=
                                        21&issue=1&aid=896.

Engblom:2016:FMC

[EL16] Stefan Engblom and Dimitar Lukarski. Fast Matlab compatible sparse assembly on multicore computers. Parallel Computing, 56(??):1–17, August 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300102.

Evans:1985:TPA

[EM85] D. J. Evans and S. Mai. Two parallel algorithms for the convex hull problem in a two dimensional space. Parallel Computing, 2(4): 313–326, December 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1986:RIU

[EM86] D. J. Evans and G. M. Megson. Romberg integration using systolic arrays. Parallel Computing, 3(4):289–304, October 1986. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1987:CET

[EM87] D. J. Evans and G. M. Megson. Construction of extrapolation tables by systolic arrays for solving ordinary differential equations. Parallel Computing, 4(1):33–48, February 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1988:OPB

[EM88] D. J. Evans and K. Margaritis. Optical processing of banded matrix algorithms using outer product concepts. Parallel Computing, 6(1):119–125, January 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1990:SDE

[EM90] D. J. Evans and K. Margaritis. Systolic designs for eigenvalue-eigenvector computations using matrix powers. Parallel Computing, 14(1):77–87, May 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Emans:2010:PPA

[Ema10] Maximilian Emans. Performance of parallel AMGpreconditioners in CFDcodes for weakly compressible flows. Parallel Computing, 36(5–6):326–338, June 2010. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1989:SHP

[EMB89] D. J. Evans, K. Margaritis, and M. P. Bekakos. Systolic and holographic pyramidical soft-systolic designs for successive matrix powers. Parallel Computing, 9(3):373–384, February 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Emmen:1984:ISA

[Emm84]  Ad Emmen.  Inter-

national supercomputer applications symposium. Parallel Computing, 1 (3–4):317–319, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819184902618.

Estana:2018:HPM

[EMV^{+}18] Alejandro Estaña, Kevin Molloy, Marc Vaisset, Nathalie Sibille, Thierry Siméon, Pau Bernadó, and Juan Cortés. Hybrid parallelization of a multitree path search algorithm: Application to highlyflexible biomolecules. Parallel Computing, 77(??): 84–100, September 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0167819118301893.

Escaig:1991:ATM

[EO91] Y. Escaig and W. Oed.

Analysis tools for microand autotasking programs on CRAY multiprocessor systems. Parallel Computing, 17(12):1425–1433, December 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Estevez:2002:SPA

[EPMPU02] Pablo A. Estévez, Hélène Paugam-Moisy, Didier Puzenat, and Manuel Ugarte. A scalable parallel algorithm for training a hierarchical mixture of neural experts. Parallel Computing, 28(6):861– 891, June 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/56/27/
abstract.html.

Eisenhauer:1998:DHP

[EPS98] Greg Eisenhauer, Beth Plale, and Karsten Schwan. DataExchange: high performance communications in distributed laboratories. Parallel Computing, 24(12–13):1713–1733, November 1, 1998. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.
                                        com/cas/tree/store/parco/
                                        sub/1998/24/12-13/1351.
                                        pdf.

Evans:1989:NSF

[ER89] D. J. Evans and A. M. S. Rahma. The numerical solution of Fredholm integral equations on parallel computers. Parallel Computing, 10(2):193–205, April 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Embrechts:1995:MDCa

[ER95a] Hugo Embrechts and Dirk Roose. MIMD divideand-conquer algorithms for the distance transformation. part I: City block distance. Parallel Computing, 21(7):1051–1076, July 11, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=7&aid=992.

Embrechts:1995:MDCb

[ER95b] Hugo Embrechts and Dirk Roose. MIMD divide-and-conquer algorithms for the distance transformation. part II. chamfer 3-4 distance. Parallel Computing, 21(7):1077–1096, July 11, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=7&aid=993.

Eibl:2018:LPC

[ER18] Sebastian Eibl and Ulrich Rüde. A local parallel communication algorithm for polydisperse rigid body dynamics. Parallel Computing, 80(??):36–48, December 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300231.

Ercegovac:1988:HSA

[Erc88] M. D. Ercegovac. Heterogeneity in supercomputer architectures. Parallel Computing, 7(3):367– 372, September 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Eltgroth:1988:SIM

[ES88] Peter G. Eltgroth and Mark K. Seager. The subimplicit method: new multiprocessor algorithms for old implicit codes. Parallel Computing, 8(1–3): 155–163, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the International Conference on Vector and Parallel Processors in Computational Science, III (Liverpool, 1987).

Eskow:1989:MMP

[ES89a] E. Eskow and R. B. Schnabel. Mathematical modeling of a parallel global optimization algorithm. Parallel Computing, 12(3):315– 325, December 1989. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1989:PRK

[ES89b] D. J. Evans and B. B. Sanugi. A parallel Runge– Kutta integration method. Parallel Computing, 11(2): 245–251, ???? 4, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1989:PCV

[ES89c] David J. Evans and Ivan Stojmenović. On parallel computation of Voronoĭ diagrams. Parallel Computing, 12(1):121–125, October 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Emad:2006:TRI

[ES06] Nahid Emad and Ani Sedrakian. Toward the reusability for iterative linear algebra software in distributed environment. Parallel Computing, 32(3): 251–266, March 2006. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1988:CFP

[ESK88] D. J. Evans, Jian Ping Shao, and Li Shan Kang. The convergence factor of the parallel Schwarz overrelaxation method for linear systems. Parallel Computing, 6(3):313–324, March 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Elble:2010:GCK

[ESV10] Joseph M. Elble, Nikolaos V. Sahinidis, and Panagiotis Vouzis. GPU computing with Kaczmarz’s and other iterative algorithms for linear systems. Parallel Computing, 36(5–6):215–231, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Erhel:1991:EEP

[ETV91] J. Erhel, A. Traynard, and M. Vidrascu. An elementby-element preconditioned conjugate gradient method implemented on a vector computer. Parallel Computing, 17(9):1051– 1065, November 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Eijkhout:1989:PDA

[EV89] V. Eijkhout and P. Vassilevski. Positive definiteness aspects of vectorizable preconditioners. Parallel Computing, 10(1):93– 100, March 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1984:PSI

[Eva84] D. J. Evans. Parallel SOR iterative methods. Parallel Computing, 1(1):3–18, August 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1989:SDA

[Eva89] D. J. Evans. A systolic design for the Aitken extrapolation formula. Parallel Computing, 11(3):385–388, ???? 28, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1990:PSM

[Eva90] D. J. Evans. A parallel sorting-merging algorithm for tightly coupled multiprocessors. Parallel Computing, 14(1):111–121, May 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1991:PAM

[Eva91] D. J. Evans. The parallel AGE method for the elliptic problem in two dimensions. Parallel Computing, 17(8):925–940, October 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1992:SAD

[Eva92] D. J. Evans. A systolic array design for matrix system solution by the symmetric bordering method. Parallel Computing, 18(2): 195–205, February 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evripidou:2001:MDD

[Evr01] Paraskevas Evripidou. D^{3}- Machine: a decoupled data-driven multithreaded architecture with variable resolution support. Parallel Computing, 27 (9):1197–1225, August 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/35/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/35/25/
article.pdf.

Evans:1993:PDS

[EW93] D. J. Evans and C. R. Wan. Parallel direct solution for P -cyclic matrix systems. Parallel Computing, 19(1):79–93, January 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1986:PNS

[EY86] D. J. Evans and N. Y. Yousif. The parallel neighbor sort and 2-way merge algorithm. Parallel Computing, 3(1):85–90, March 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1990:IEB

[EY90] D. J. Evans and W. S. Yousif. The implementation of the explicit block iterative methods on the balance 8000 parallel computer. Parallel Computing, 16(1):81–97, November 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Evans:1994:SUT

[EY94] D. J. Evans and W. S. Yousif. The solution of unsymmetric tridiagonal Toeplitz systems by the strides reduction algorithm. Parallel Computing, 20(5):787–798, May 11, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=5&aid=842.

Elis:2020:QNG

[EYP^{+}20] Bengisu Elis, Dai Yang, Olga Pearce, Kathryn Mohror, and Martin Schulz. QMPI: a next generation MPI profiling interface for modern HPC platforms. Parallel Computing, 96 (??):Article 102635, August 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300284.

Ebner:1999:DFF

[EZ99] Ralf Ebner and Christoph Zenger. A distributed functional framework for recursive finite element simulations. Parallel Computing, 25(7):813–826, August 13, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1999&volume=
25&issue=7&aid=1411;
http://www.elsevier.

                                        nl/gej-ng/10/35/21/32/
                                        23/21/abstract.html;
                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/32/
                                        23/21/article.pdf.

Fragopoulou:1996:SSA

[FA96]  Paraskevi  Fragopoulou

and Selim G. Akl. Spanning subgraphs with applications to communication on the multidimensional torus network. Parallel Computing, 22 (7):991–1015, October 1, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=7&aid=1080.

Flaig:2011:SME

[FA11] Cyril Flaig and Peter Arbenz. A scalable memory efficient multigrid solver for micro-finite element analyses based on CT images. Parallel Computing, 37(12):846–854, December 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111001037.

Fabeiro:2016:WPP

[FAF16] Jorge F. Fabeiro, Diego Andrade, and Basilio B.

Fraguela. Writing a performance-portable matrix multiplication. Parallel Computing, 52(??):65– 77, February 2016. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819115001611.

Fahlman:1985:PPA

[Fah85]  S. E. Fahlman.  Par-

allel processing in artificial intelligence. Parallel Computing, 2(3):283–286, November 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Farrag:1994:TFE

[Far94] Abdel Aziz Farrag. Tolerating faulty edges in a multi-dimensional mesh. Parallel Computing, 20 (9):1289–1301, September 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=9&aid=891.

Farrag:1996:NAC

[Far96] Abdel Aziz Farrag. New algorithm for constructing fault-tolerant solutions of the circulant graph configuration. Parallel Computing, 22(9):1239–1253 (or 1239–1254??), November 22, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=9&aid=1081.

Ferreira:2002:DPL

[FAS02] Renato Ferreira, Gagan Agrawal, and Joel Saltz. Data parallel language and compiler support for data intensive applications. Parallel Computing, 28(5):725–748, May 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/57/30/
abstract.html.

Freeman:1991:APZ

[FB91] T. L. Freeman and M. K. Bane. Asynchronous polynomial zero-finding algorithms. Parallel Computing, 17(6–7):673–681, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Favre:2019:CET

[FB19] Jean M. Favre and Alexander Blass. A comparative evaluation of three volume rendering libraries for the visualization of sheared thermal convection. Parallel Computing, 88(??):Article 102543, ???? 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301280.

Fortmeier:2013:NME

[FBAB13] O. Fortmeier, H. M.

Bücker, B. O. Fagginger Auer, and R. H. Bisseling. A new metric enabling an exact hypergraph model for the communication volume in distributed-memory parallel applications. Parallel Computing, 39(8):319–335, August 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000690.

Fagg:2001:HFT

[FBD01] Graham E. Fagg, Antonin Bukovsky, and Jack J. Dongarra. HARNESS and fault tolerant MPI. Parallel Computing, 27 (11):1479–1495, October 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-

                                        ng/10/35/21/47/41/32/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/41/32/
                                        article.pdf.

Fraguela:2012:OTE

[FBG^{+}12] Basilio B. Fraguela, Ganesh Bikshandi, Jia Guo, María J. Garzarán, David Padua, and Christoph von Praun. Optimization techniques for efficient HTA programs. Parallel Computing, 38(9):465–484, September 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819112000312.

Finnemann:1988:SND

[FBMV88] H. Finnemann, J. Brehm, E. Michel, and J. Volkert. Solution of the neutron diffusion equation through multigrid methods implemented on a memorycoupled 25-processor system. Parallel Computing, 8(1–3):391–398, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Furumura:2005:PSS

[FC05] T. Furumura and L. Chen. Parallel simulation of strong ground motions during recent and historical damaging earthquakes in Tokyo, Japan. Parallel Computing, 31(2):149–165, February 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fonseca:2018:ONF

[FC18] Alcides Fonseca and Bruno Cabral. Overcoming the No Free Lunch Theorem in cut-off algorithms for fork–join programs. Parallel Computing, 76(??):42– 56, August 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301157.

Feng:2016:TLL

[FCC16] Daming Feng, Andrey N.

Chernikov, and Nikos P. Chrisochoides. Two-level locality-aware parallel Delaunay image-to-mesh conversion. Parallel Computing, 59(??):60–70, November 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000375.

Freitas:2003:CWA

[FCM03] Christopher J. Freitas, Derrick B. Coffin, and Richard L. Murphy. The characterization of a wide area network computation.

Parallel Computing, 29(7): 879–894, July 2003. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fujita:2019:EIM

[FCS^{+}19] Hajime Fujita, Chongxiao Cao, Sayantan Sur, Charles Archer, Erik Paulson, and Maria Garzaran. Efficient implementation of MPI-3 RMA over open-Fabrics interfaces. Parallel Computing, 87(??):1–10, September 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118303843.

Fu:2017:FEM

[FCZY17] Huansong Fu, Haiquan Chen, Yue Zhu, and Weikuan Yu. FARMS: Efficient MapReduce speculation for failure recovery in short jobs. Parallel Computing, 61(??):68–82, January 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116301077.

Fresse:2002:ARP

[FD02] Virginie Fresse and Olivier Deforges. ARIAL: Rapid Prototyping for Mixed and Parallel Platforms.

Parallel Computing, 28 (7–8):1179–1202, August 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/58/39/
abstract.html.

Fanfarillo:2018:NAC

[FD18] Alessandro Fanfarillo and Davide Del Vento. Notified access in coarraybased hydrodynamics applications on many-core architectures: Design and performance. Parallel Computing, 75(??):118– 129, July 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301005.

Fraguela:2004:CTP

[FDTZ04] B. B. Fraguela, R. Doallo, J. Touriño, and E. L. Zapata. A compiler tool to predict memory hierarchy performance of scientific codes. Parallel Computing, 30(2):225–248, February 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Feitelson:1991:DDW

[Fei91] D. G. Feitelson. Deadlock detection without wait-for graphs. Parallel Computing, 17(12):1377–1383, December 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Feo:1988:ACP

[Feo88] J. T. Feo. An analysis of the computational and parallel complexity of the Livermore loops. Parallel Computing, 7(2):163–185, June 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ferenczi:1997:PEO

[Fer97] Szabolcs Ferenczi. Parallel execution of objectoriented programs: Message handling strategies. Parallel Computing, 22 (13):1897–1912, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        22&issue=13&aid=1134.

Filin:1992:SAI

[FF92] Anatol G. Filin and Michael A. Frumkin. A systolic array for inversion of a finite Radon transform. Parallel Computing, 18(1):85–90, January 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fallah:2021:PBB

[FF21] Mohammad K. Fallah and Mahmood Fazlali. Parallel branch and bound algorithm for solving integer linear programming models derived from behavioral synthesis. Parallel Computing, 101(??): Article 102722, April 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120301071.

Fabbretti:1996:MSA

[FFLV96] G. Fabbretti, A. Farina, D. Laforenza, and F. Vinelli. Mapping the synthetic aperture radar signal processor on a distributed-memory MIMD architecture. Parallel Computing, 22(5):761– 784, August 8, 1996. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cgi-bin/cas/tree/
store/parco/cas_sub/
browse/browse.cgi?year=
1996&volume=22&issue=
5&aid=1064.

Fang:2018:BGM

[FFZ^{+}18] Minquan Fang, Jianbin Fang, Weimin Zhang, Haifang Zhou, Jianxing Liao, and Yuangang Wang. Benchmarking the GPU memory at the warp level. Parallel Computing, 71(??):23–41, January 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301825.

Fernandes:1989:NSS

[FG89] P. Fernandes and P. Girdinio. A new storage scheme for an efficient implementation of the sparse matrix-vector product. Parallel Computing, 12(3):327–333, December 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Franklin:1996:GMI

[FG96] Mark A. Franklin and Vasudha Govindan. A general matrix iterative model for dynamic load balancing. Parallel Computing, 22(7):969–989, October 1, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=7&aid=1077.

Florian:2001:APC

[FG01] Michael Florian and Michel Gendreau. Applications of parallel computing in transportation. Parallel Computing, 27 (12):1521–1522, November 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/42/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/42/27/
article.pdf.

Fanfarillo:2019:RCA

[FGBN19] Alessandro Fanfarillo, Sudip Kumar Garain, Dinshaw Balsara, and Daniel Nagle. Resilient computational applications using Coarray Fortran. Parallel Computing, 81(??):58–67, January 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303867.

Foster:1998:WAI

[FGG^{+}98] Ian Foster, Jonathan Geisler, William Gropp, Nicholas Karonis, Ewing Lusk, George Thiruvathukal, and Steven Tuecke. Wide-area implementation of the Message Passing Interface. Parallel Computing, 24(12– 13):1735–1749, November 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/12-13/1352.pdf.

Fleischmann:1994:PAP

[FGHB94] Georg Fleischmann, Matthias Gente, Fridolin Hofmann, and Gunter Bolch. Performance analysis of parallel programs based on model calculations. Parallel Computing, 20(10– 11):1583–1603, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=915.

Foyer:2023:SST

[FGP23] Clément Foyer, Brice Goglin, and Andrès Rubio Proaño. A survey of software techniques to emulate heterogeneous memory systems in high-performance computing. Parallel Computing, 116(??):??, July 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819123000297.

Flocchini:1997:CLE

[FGS97] Paola Flocchini, Frédéric Geurts, and Nicola Santoro. CA-like error propagation in fuzzy CA. Parallel Computing, 23 (11):1673–1682, December 1, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=11&aid=1235.

Frederickson:1984:PRT

[FHJ^{+}84] P. Frederickson, R. Hiromoto, T. L. Jordan, B. Smith, and T. Warnock. Pseudo-random trees in Monte Carlo. Parallel Computing, 1(2):175–180, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Feichtinger:2011:FPB

[FHK^{+}11] Christian Feichtinger, Johannes Habich, Harald Köstler, Georg Hager, Ulrich Rüde, and Gerhard Wellein. A flexible patchbased lattice Boltzmann parallelization approach for heterogeneous GPU– CPU clusters. Parallel Computing, 37(9):536–549, September 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111000342.

Feichtinger:2015:PMA

[FHK^{+}15] Christian Feichtinger, Johannes Habich, Harald Köstler, Ulrich Rüde, and Takayuki Aoki. Performance modeling and analysis of heterogeneous lattice Boltzmann simulations on CPU–GPU clusters. Parallel Computing, 46(??):1–13, July 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114001446.

Franke:1988:PBI

[FHKZ88] B. Franke, R. Harneit, A. Kern, and H. C. Zeidler. The pipeline bus: an interconnection network for multiprocessor systems. Parallel Computing, 7(3):403–412, September 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Federickson:1987:PMC

[FHL87] P. Federickson, R. Hiromoto, and J. Larson. A parallel Monte Carlo transport algorithm using a pseudo-random tree to guarantee reproducibility. Parallel Computing, 4(3): 281–290, June 1987. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Foresti:1993:PRO

[FHMS93] S. Foresti, S. Hassanzadeh, H. Murakami, and V. Sonnad. Parallel rapid operator for iterative finite element solvers on a shared memory machine. Parallel Computing, 19(1):1–7, January 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fiebach:1996:CBA

[Fie96]  Peter Fiebach.  Cyclic

block-algorithms for solving triangular systems on distributed-memory multiprocessors with mesh topology. Parallel Computing, 22(3):383–393, April 29, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=3&aid=1052.

Figueira:2006:OPN

[Fig06] Silvia M. Figueira. Optimal partitioning of nodes to space-sharing parallel tasks. Parallel Computing, 32(4):313–324, April 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fujiwara:1999:COP

[FIMF99] Akihiro Fujiwara, Michiko Inoue, Toshimitsu Masuzawa, and Hideo Fujiwara. A cost optimal parallel algorithm for weighted distance transforms. Parallel Computing, 25(4):405–416, April 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1999/
                                        25/4/1398.pdf.

Fischer:1990:ADP

[Fis90] Herbert Fischer. Automatic differentiation: parallel computation of function, gradient, and Hessian matrix. Parallel Computing, 13(1):101–110, January 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fischer:1991:SS

[Fis91] D. Fischer. On superlinear speedups. Parallel Computing, 17(6–7): 695–697, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Flick:2017:RPC

[FJPA17] Patrick Flick, Chirag Jain, Tony Pan, and Srinivas Aluru. Reprint of “A parallel connectivity algorithm for de Bruijn graphs in metagenomic applications”. Parallel Computing, 70(??):54–65, December 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301485.

Frederickson:1985:SCP

[FJS85] Paul O. Frederickson, Rondall E. Jones, and Brian T. Smith. Synchronization and control of parallel algorithms. Parallel Computing, 2(3):255–264, November 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Flatt:1989:PPP

[FK89] Horace P. Flatt and Ken Kennedy. Performance of parallel processors. Parallel Computing, 12(1):1– 20, October 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Flower:1994:EJM

[FK94] Jon Flower and Adam Kolawa. Express is not just a message passing system:

Current and future directions in Express. Parallel Computing, 20(4):597–614, March 31, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=4&aid=860.

Fan:1998:EBS

[FK98] Kuo-Pao Fan and Chung-Ta King. Efficient barrier synchronization in wormhole-routed mesh networks supporting turn model. Parallel Computing, 24(14):2077–2099, December 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/14/1367.pdf.

Fatoohi:2006:PEH

[FKK^{+}06] Rod Fatoohi, Ken Kardys, Sumy Koshy, Soundarya Sivaramakrishnan, and Jeffrey S. Vetter. Performance evaluation of highspeed interconnects using dense communication patterns. Parallel Computing, 32(11–12):794–807, December 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fischer:2022:NGA

[FKM^{+}22] Paul Fischer, Stefan Kerkemeier, Misun Min, Yu-Hsiang Lan, Malachi Phillips, Thilina Rathnayake, Elia Merzari, Ananias Tomboulides, Ali Karakus, Noel Chalmers, and Tim Warburton. NekRS, a GPU-accelerated spectral element Navier– Stokes solver. Parallel Computing, 114(??):??, December 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000710.

Fairbanks:2015:BCD

[FKPB15] James P. Fairbanks, Ramakrishnan Kannan, Haesun Park, and David A. Bader. Behavioral clusters in dynamic graphs. Parallel Computing, 47(??):38– 50, August 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000460.

Ferreira:2021:EMR

[FL21] Kurt B. Ferreira and Scott Levy. Evaluating MPI resource usage summary statistics. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000739.

Ferreira:2018:CMM

[FLPG18] Kurt B. Ferreira, Scott Levy, Kevin Pedretti, and Ryan E. Grant. Characterizing MPI matching via trace-based simulation. Parallel Computing, 77(??):57–83, September 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118301467.

Falgout:2021:PHH

[FLS^{+}21] Robert D. Falgout, Ruipeng Li, Björn Sjögreen, Lu Wang, and Ulrike Meier Yang. Porting hypre to heterogeneous computer architectures: Strategies and experiences. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000867.

Faber:1986:SSE

[FLW86] V. Faber, Olaf M. Lubeck, and Andrew B. White, Jr. Superlinear speedup of an efficient sequential algorithm is not possible. Parallel Computing, 3(3):259– 260, July 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Faber:1987:CPP

[FLW87] V. Faber, O. M. Lubeck, and A. B. White, Jr. Comments on the paper ‘Parallel efficiency can be greater than unity’. Parallel Computing, 4(2):209–210, April 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). See [Par86].

Fei:2016:SEF

[FLYL16] Xiongwei Fei, Kenli Li, Wangdong Yang, and Keqin Li. A secure and efficient file protecting system based on SHA3 and parallel AES. Parallel Computing, 52(??):106–132, February 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000028.

Fei:2021:AEE

[FLYL21] Xiongwei Fei, Kenli Li, Wangdong Yang, and Keqin Li. Analysis of energy efficiency of a parallel AES algorithm for CPU– GPU heterogeneous platforms. Parallel Computing, 94–95(??):Article 102621, June 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819120300144.

Francis:1988:SES

[FM88] R. Francis and I. Mathieson. Synchronised execution on shared memory multiprocessors. Parallel Computing, 8(1–3):165– 175, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ferretti:2015:GMS

[FM15] Marco Ferretti and Mirto Musci. Geometrical motifs search in proteins: a parallel approach. Parallel Computing, 42(??):60–74, February 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114001185.

Faraji:2017:EHC

[FMA17] Iman Faraji, Seyed H. Mirsadeghi, and Ahmad Afsahi. Exploiting heterogeneity of communication channels for efficient GPU selection on multi-GPU nodes. Parallel Computing, 68(??):3–16, October 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300911.

Ford:1998:RAI

[FMB98] Daniel A. Ford, Robert J. T. Morris, and Alan E. Bell. Redundant arrays of independent libraries (RAIL): the StarFish tertiary storage system. Parallel Computing, 24 (1):45–64, March 10, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1998&volume=
24&issue=1&aid=1274.

Fonlupt:1998:DPL

[FMD98] Cyril Fonlupt, Philippe Marquet, and Jean-Luc Dekeyser. Data-parallel load balancing strategies. Parallel Computing, 24 (11):1665–1684, October 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/11/1326.pdf.

Furmento:2002:IOC

[FMM^{+}02] Nathalie Furmento, Anthony Mayer, Stephen Mc-Gough, Steven Newhouse, Tony Field, and John Darlington. ICENI: Optimisation of component applications within a Grid environment. Parallel Computing, 28(12):1753–1772, December 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fusai:2010:OPM

[FMM10] Gianluca Fusai, Daniele Marazzina, and Marina Marena. Option pricing, maturity randomization and distributed computing. Parallel Computing, 36(7):403–414, July 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fraigniaud:1990:SRP

[FMR90] Pierre Fraigniaud, Serge Miguet, and Yves Robert. Scattering on a ring of processors. Parallel Computing, 13(3):377–383, March 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fisher:2001:PPW

[FMS01] M. S. Fisher, M. Mani, and D. Stookesberry. Parallel processing with the Wind CFD code at Boeing. Parallel Computing, 27(4):441–456, March 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/28/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/28/28/
article.pdf.

Folino:2006:MBC

[FMS^{+}06] Gianluigi Folino, Giuseppe Mendicino, Alfonso Senatore, Giandomenico Spezzano, and Salvatore Straface. A model based on cellular automata for the parallel simulation of 3D unsaturated flow. Parallel Computing, 32(5–6):357– 376, June 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fox:1987:MAH

[FOH87] G. C. Fox, S. W. Otto, and A. J. G. Hey. Matrix algorithms on a hypercube. I. matrix multiplication. Parallel Computing, 4(1):17–31, February 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Francis:1992:PPE

[FP92] R. S. Francis and L. J. H.

Pannan. A parallel partition for enhanced parallel QuickSort. Parallel Computing, 18(5):543–550, May 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fisette:1998:CPV

[FP98] P. Fisette and J. M. Péterkenne. Contribution to parallel and vector computation in multibody dynamics. Parallel Computing, 24(5–6):717–728, June 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/5-6/1313.pdf.

Francomano:1991:PEI

[FPT91] E. Francomano, A. Pecorella, and A. Tortorici Macaluso. Parallel experience on the inverse matrix computation. Parallel Computing, 17(8):907–912, October 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Floros:1995:ESE

[FR95] N. Floros and J. S. Reeve. Evaluation of a spectral element CFD code on parallel architectures. Parallel Computing, 21(7):1137–1150, July 11, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/

cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=7&aid=989.

Floros:1995:CED

[FRC^{+}95] N. Floros, J. S. Reeve, J. Clinckemaillie, S. Vlachoutsis, and G. Lonsdale. Comparative efficiencies of domain decompositions. Parallel Computing, 21(11):1823–1835, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=11&aid=1010.

Freeman:1989:CPZ

[Fre89] T. L. Freeman. Calculating polynomial zeros on a local memory parallel computer. Parallel Computing, 12(3): 351–358, December 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Forrest:1988:NNM

[FRS^{+}88] B. M. Forrest, D. Roweth, N. Stroud, D. J. Wallace, and G. V. Wilson. Neural network models. Parallel Computing, 8(1–3):71– 83, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Falcou:2006:QEC

[FSCL06] J. Falcou, J. Sérot, T. Chateau, and J. T. Lapresté. Quaff: efficient C++ design for parallel skeletons. Parallel Computing, 32(7–8): 604–615, September 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fan:2019:SAO

[FSG19] Xing Fan, Oliver Sinnen, and Nasser Giacaman. Supporting asynchronization in OpenMP for eventdriven programming. Parallel Computing, 82(??): 57–74, ???? 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118300838.

Farahabady:2006:CSF

[FSKF06] M. Hoseiny Farahabady, F. Safaei, A. Khonsari, and M. Fathy. Characterization of spatial fault patterns in interconnection networks. Parallel Computing, 32(11–12):886–901, December 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

FroeseFischer:1988:MCA

[FSY88] C. Froese Fischer, N. S. Scott, and J. Yoo. Multitasking the calculation of angular integrals on the CRAY-2 and CRAY X-MP. Parallel Computing, 8(1–3):385–390, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Funabiki:1993:PML

[FT93] N. Funabiki and Y. Takefuji. A parallel multilayer channel router on the HVH model. Parallel Computing, 19(1):63– 77, January 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fu:2020:IPP

[FTY^{+}20] Zhongming Fu, Zhuo Tang, Li Yang, Kenli Li, and Keqin Li. ImRP: a predictive partition method for data skew alleviation in spark streaming environment. Parallel Computing, 100(??):Article 102699, December 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300892.

Fu:2003:FTC

[Fu03] Jung-Sheng Fu. Faulttolerant cycle embedding in the hypercube. Parallel Computing, 29(6):821– 832, June 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fu:2007:CFT

[Fu07] Jung-Sheng Fu. Conditional fault-tolerant hamiltonicity of star graphs. Parallel Computing, 33(7– 8):488–496, August 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fang:2003:DGO

[FWL03] Weijian Fang, Cho-Li Wang, and Francis C. M. Lau. On the design of global object space for efficient multi-threading Java computing on clusters. Parallel Computing, 29(11–12):1563–1587, November/December 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Frieze:1989:AAP

[FYEHP89] A. M. Frieze, J. Yadegar, S. El-Horbaty, and D. Parkinson. Algorithms for assignment problems on an array processor. Parallel Computing, 11(2): 151–162, ???? 4, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fu:2018:SCC

[FYH^{+}18] Zhenxin Fu, Lei Yang, Wenbin Hou, Zhuohan Li, Yifan Wu, Yihua Cheng, Xiaolin Wang, and Yun Liang. Student Cluster Competition 2017, Team Peking University: Reproducing vectorization of the Tersoff multi-body potential on the Intel Broadwell architecture. Parallel Computing, 78(??):28– 32, October 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300693.

Goehner:2013:LFB

[GAA^{+}13] J. D. Goehner, D. C.

Arnold, D. H. Ahn, G. L. Lee, B. R. de Supinski, M. P. LeGendre, B. P. Miller, and M. Schulz. LIBI: a framework for bootstrapping extreme scale software systems. Parallel Computing, 39(3): 167–176, March 2013. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819112000774.

Gonzalez:2003:TAO

[GAMR03] D. González, F. Almeida, L. Moreno, and C. Rodríguez. Towards the automatic optimal mapping of pipeline algorithms. Parallel Computing, 29(2):241–254, February 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gao:1987:SCM

[Gao87] Guang R. Gao. A stability classification method and its application to pipelined solution of linear recurrences. Parallel Computing, 4(3):305–321, June 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gao:1990:EFG

[Gao90]  G. R. Gao.  Exploit-

ing fine-grain parallelism on dataflow architectures. Parallel Computing, 13(3): 309–320, March 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gutierrez:2000:API

[GAPZ00] E. Gutiérrez, R. Asenjo, O. Plata, and E. L. Zapata. Automatic parallelization of irregular applications. Parallel Computing, 26(13–14):1709–1738, December 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/34/24/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/34/24/
                                        article.pdf.

Goddeke:2015:FTF

[GAR15] Dominik Göddeke, Mirco Altenbernd, and Dirk Ribbrock. Fault-tolerant finite-element multigrid algorithms with hierarchically compressed asynchronous checkpointing. Parallel Computing, 49 (??):117–135, November 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001064.

Gonzalez-Alvarez:2018:SCP

[GÁVRRL18] David L. González-Álvarez, Miguel A. Vega-Rodríguez, and Álvaro Rubio-Largo. Searching for common patterns on protein sequences by means of a parallel hybrid honey-bee mating optimization algorithm. Parallel Computing, 76(??):1– 17, August 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300978.

Gutbrod:1996:SLG

[GAW96] F. Gutbrod, N. Attig, and M. Weber. The SU(2)- Lattice Gauge Theory simulation code on the Intel Paragon supercomputer. Parallel Computing, 22(3):443–463, April 29, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=3&aid=1057.

Green:2014:SDA

[GB14] Oded Green and Yitzhak Birk. Scheduling directives: Accelerating sharedmemory many-core processor execution. Parallel Computing, 40(2):90–106, February 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001464.

Galloway:1998:TOA

[GBH98] Robert L. Galloway, W. Andrew Bass, and Christopher E. Hockey. Taskoriented asymmetric multiprocessing for interactive image-guided surgery. Parallel Computing, 24(9– 10):1323–1343, September 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/9-10/1337.pdf.

Ghane:2019:PTP

[GCC19] Millad Ghane, Sunita Chandrasekaran, and Margaret S. Cheung. pointerchain: Tracing pointers to their roots — a case study in molecular dynamics simulations. Parallel Computing, 85(??):190–203, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118302618.

Gao:2021:TAS

[GCH21] Jiaquan Gao, Qi Chen, and Guixia He. A thread-adaptive sparse approximate inverse preconditioning algorithm on multi-GPUs. Parallel Computing, 101(??): Article 102724, April 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120301083.

Glatard:2011:MPJ

[GCP11] Tristan Glatard and Sorina Camarasu-Pop. A model of pilot-job resource provisioning on production grids. Parallel Computing, 37(10–11):684– 692, October/November 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111000366.

Geist:1990:FEE

[GD90] G. A. Geist and G. J. Davis. Finding eigenvalues and eigenvectors of unsymmetric matrices using a distributed-memory multiprocessor. Parallel Computing, 13(2):199–209, February 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Goumas:2006:MPC

[GDAK06] Georgios Goumas, Nikolaos Drosinos, Maria Athanasaki, and Nectarios Koziris. Message-passing code generation for non-rectangular tiling transformations. Parallel Computing, 32(10): 711–732, November 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gomez:2003:PTM

[GdCCS03] Susana Gómez, Nelson del Castillo, Longina Castellanos, and Julio Solano. The parallel tunneling method. Parallel Computing, 29(4):523–533, April 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gupta:1997:FNT

[GDMS97] Srabani Sen Gupta, Rajib K. Das, Krishnendu Mukhopadhyaya, and Bhabani P. Sinha. A family of network topologies with multiple loops and logarithmic diameter. Parallel Computing, 22(14):2047–2064, March 24, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=14&aid=1118.

Gusev:1992:VPA

[GE92] M. Gusev and D. J. Evans.

VLSI processor array IPS cells (short communication). Parallel Computing, 18(9):997–1007, September 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gusev:1993:NLS

[GE93] M. Gusev and D. J. Evans.

New linear systolic arrays for the string comparison algorithm. Parallel Computing, 19(10):1177–1193, October 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Garces-Erice:2011:ACR

[GE11] Luis Garcés-Erice. Admission control for a responsive distributed middleware using decision trees to model run-time parameters. Parallel Computing, 37(8):379–391, August 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819110001286.

Gentzsch:1984:NAC

[Gen84] W. Gentzsch. Numerical algorithms in computational fluid dynamics on vector computers. Parallel Computing, 1(1):19–33, August 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gentzsch:1987:FVS

[Gen87] W. Gentzsch. A fully vectorizable SOR variant. Parallel Computing, 4(3): 349–353, June 1987. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gerteisen:1994:PPR

[Ger94] Edgar A. Gerteisen. Preliminary performance results of the massive parallel Aircraft Euler Method. Parallel Computing, 20(10–11):1675–1683, November 3, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cgi-bin/cas/tree/
store/parco/cas_sub/
browse/browse.cgi?year=
1994&volume=20&issue=
10-11&aid=922.

Gerndt:1998:HLP

[Ger98] Michael Gerndt. Highlevel programming of massively parallel computers based on shared virtual memory. Parallel Computing, 24(3–4):383–400, May 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/3-4/1295.pdf.

Gerbessiotis:2004:AIP

[Ger04] Alexandros V. Gerbessiotis. Architecture independent parallel binomial tree option price valuations. Parallel Computing, 30(2):301–316, February 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gerbessiotis:2015:EBM

[Ger15] Alexandros V. Gerbessiotis. Extending the BSP model for multi-core and out-of-core computing:

MBSP. Parallel Computing, 41(??):90–102, January 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114001434.

Gonzalez-Escribano:2009:PIS

[GEvGCP09] Arturo González-Escribano, Arjan J. C. van Gemund, and Valentín Cardeñoso-Payo. Performance implications of synchronization structure in parallel programming. Parallel Computing, 35(8–9):455–474, August/September 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gary:1989:OPF

[GF89] J. Gary and L. Fosdick. An optimizing precompiler for finite-difference computations on a vector computer. Parallel Computing, 10(1):51–64, March 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ghoneim:2003:JPF

[GF03] Salma A. Ghoneim and Hossam M. A. Fahmy. Job preemption, fast subcube compaction, or waiting in hypercube systems? A selection methodology. Parallel Computing, 29(1):

111–134, January 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gonzalez:2013:FAB

[GF13] Carlos H. González and Basilio B. Fraguela. A framework for argumentbased task synchronization with automatic detection of dependencies. Parallel Computing, 39(9):475–489, September 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000562.

Gordon:2010:CCR

[GG10] Dan Gordon and Rachel Gordon. CARP–CG: a robust and efficient parallel solver for linear systems, applied to strongly convection dominated PDEs. Parallel Computing, 36(9): 495–515, September 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ghazimirsaeed:2019:DUD

[GGA19] S. Mahdieh Ghazimirsaeed, Ryan E. Grant, and Ahmad Afsahi. A dynamic, unified design for dedicated message matching engines for collective and point-to-point communications. Parallel Computing, 89(??):

Article 102547, November 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301383.

GracaRuano:1993:API

[GGFF93] M. Graca Ruano, D. F. Garcia Nocetti, P. J. Fish, and P. J. Fleming. Alternative parallel implementations of an AR-modified covariance spectral estimator for diagnostic ultrasonic blood flow studies. Parallel Computing, 19(4): 463–476, April 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gawande:2022:TSC

[GGH^{+}22] Nitin Gawande, Sayan Ghosh, Mahantesh Halappanavar, Antonino Tumeo, and Ananth Kalyanaraman. Towards scaling community detection on distributed-memory heterogeneous systems. Parallel Computing, 111(??): ??, July 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000060.

Geigl:1999:TDP

[GGL99] Max Geigl, Martin Griebl, and Christian Lengauer. Termination detection in parallel loop nests with while loops. Parallel Computing, 25(12):1489–1510, November 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/35/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/35/27/
article.pdf.

Guyetant:2005:CRC

[GGL^{+}05] Stéphane Guyetant, Mathieu Giraud, Ludovic L’Hours, Steven Derrien, Stéphane Rubini, Dominique Lavenier, and Frédéric Raimbault. Cluster of reconfigurable nodes for scanning large genomic banks. Parallel Computing, 31(1):73–96, January 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gonzales:2021:PFA

[GGL21] Ronald Gonzales, Yury Gryazin, and Yun Teck Lee. Parallel FFT algorithms for high-order approximations on threedimensional compact stencils. Parallel Computing, 103(??):??, June 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000168.

Gay:2011:YSP

[GGNY11] David Gay, Joel Galenson, Mayur Naik, and Kathy Yelick. Yada: Straightforward parallel programming. Parallel Computing, 37(9):592–609, September 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111000184.

Genaud:2004:LBS

[GGV04] Stéphane Genaud, Arnaud Giersch, and Frédéric Vivien. Load-balancing scatter operations for grid computing. Parallel Computing, 30(8):923–946, August 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gladwell:1989:VPO

[GH89] I. Gladwell and R. I. Hay. Vector- and parallelisation of ODE BVP codes. Parallel Computing, 12(3):343– 350, December 1989. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gilbert:1990:PSF

[GH90] John R. Gilbert and Hjálmtýr Hafsteinsson. Parallel symbolic factorization of sparse linear systems. Parallel Computing, 14(2):151–162, June 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gupta:1992:LBT

[GH92] Ajay K. Gupta and Susanne E. Hambrusch. Load balanced tree embeddings. Parallel Computing, 18(6): 595–614, June 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gregor:1998:CSF

[GH98] Jens Gregor and Dean A.

Huff. A computational study of the focus-of-attention EM-ML algorithm for PET reconstruction. Parallel Computing, 24(9–10): 1481–1497, September 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/9-10/1344.pdf.

Guo:2004:FPM

[GHC04] Minyi Guo, Michael (Shan-Hui) Ho, and Weng-Long Chang. Fast parallel molecular solution to the dominating-set problem on massively parallel bio-computing. Parallel Computing, 30(9–10): 1109–1125, September/ October 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ghezzi:1985:CPL

[Ghe85] C. Ghezzi. Concurrency in programming languages: a survey. Parallel Computing, 2(3):229–241, November 1985. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Govett:2003:SMS

[GHH^{+}03] M. Govett, L. Hart, T. Henderson, J. Middlecoff, and D. Schaffer. The Scalable Modeling System: directive-based code parallelization for distributed and shared memory computers. Parallel Computing, 29(8):995–1020, August 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Garrett:2018:SIS

[GHH18] C. Kris Garrett, Stephen Lien Harrell, and Michael A. Heroux. Special issue on SCC’17 reproducibility initiative. Parallel Computing, 79(??):48–49, November 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118302734.

Gutjahr:1997:TAC

[GHM97] W. J. Gutjahr, M. Hitz, and T. A. Mueck. Task assignment in Cayley interconnection topologies. Parallel Computing, 23 (10):1429–1460, November 7, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=10&aid=1228.

George:1987:SCF

[GHNL87] Alan George, Michael T.

Heath, Esmond Ng, and Joseph Liu. Symbolic Cholesky factorization on a local-memory multiprocessor. Parallel Computing, 5(1–2):85–95, July 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Giraud:2010:UML

[GHP10] L. Giraud, A. Haidar, and S. Pralet. Using multiple levels of parallelism to enhance the performance of domain decomposition solvers. Parallel Computing, 36(5–6):285–296, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Goscinski:2002:GET

[GHS02] A. Goscinski, M. Hobbs, and J. Silcock. GENE-SIS: an efficient, transparent and easy to use cluster operating system. Parallel Computing, 28(4):557– 606, April 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/39/28/
                                        abstract.html.

Giraud:2008:PSS

[GHW08] L. Giraud, A. Haidar, and L. T. Watson. Parallel scalability study of hybrid preconditioners in three dimensions. Parallel Computing, 34(6–8): 363–379, July 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gomez-Iglesias:2018:PET

[GICM18] Antonio Gómez-Iglesias and Miguel Cárdenas-Montes. Performance evaluation of the three-point angular correlation function. Parallel Computing, 76(??):28–41, August 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301236.

Gotz:2010:CMD

[GIF^{+}10] J. Götz, K. Iglberger, C. Feichtinger, S. Donath, and U. Rüde. Coupling multibody dynamics and computational fluid dynamics on 8192 processor cores. Parallel Computing, 36(2–3):142–151, February/March 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Giloi:1988:STM

[Gil88] W. K. Giloi. SUPRENUM:

a trendsetter in modern supercomputer development. Parallel Computing, 7(3): 283–296, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Giloi:1994:SSG

[Gil94a]  W. K. Giloi.  The

SUPRENUM supercomputer: Goals, achievements, and lessons learned. Parallel Computing, 20 (10–11):1407–1425, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=905.

Giloi:1994:PSA

[Gil94b] Wolfgang K. Giloi. Parallel supercomputer architectures and their programming models. Parallel Computing, 20(10– 11):1443–1470, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=907.

Ginsberg:1999:ICS

[Gin99]  Myron Ginsberg.  In-

fluences, challenges, and strategies for automotive HPC benchmarking and performance improvement. Parallel Computing, 25 (12):1459–1476, November 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/35/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/35/25/
article.pdf.

Girault:2002:ERM

[Gir02] Alain Girault. Elimination of redundant messages with a two-pass static analysis algorithm. Parallel Computing, 28(3):433–453, March 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/34/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/34/30/
main.pdf.

Guerreiro:2019:DAA

[GIRT19] João Guerreiro, Aleksandar Ilic, Nuno Roma, and Pedro Tomás. DVFSaware application classification to improve GPG-PUs energy efficiency. Parallel Computing, 83(??): 93–117, April 2019. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819118300243.

Glasgow:1989:EPA

[GJMM89] J. Glasgow, M. Jenkins, H. Meijer, and C. Mc-Crosky. Expressing parallel algorithms in Nial. Parallel Computing, 11(3): 331–347, ???? 28, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Goglin:2018:HTM

[GJMM18] Brice Goglin, Emmanuel Jeannot, Farouk Mansouri, and Guillaume Mercier. Hardware topology management in MPI applications through hierarchical communicators. Parallel Computing, 76(??):70–90, August 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118301480.

Gaujal:1997:HSS

[GJMMS97] Bruno Gaujal, Alain Jean-Marie, Philippe Mussi, and Gunther Siegel. High speed simulation of discrete event systems by mixing process oriented and equational approaches. Parallel Computing, 23 (1–2):219–233, April 16, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=1-2&aid=1157.

Gartel:1993:FRP

[GJS93a] U. Gartel, W. Joppich, and A. Schuller. First results with a parallelized 3D weather prediction code. Parallel Computing, 19 (12):1427–1429, December 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gartel:1993:PEW

[GJS93b] U. Gartel, W. Joppich, and A. Schuller. Parallelizing the ECMWF’s weather forecast program: the 2D case. Parallel Computing, 19(12):1413–1425, December 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Goles:1992:LBC

[GK92] Eric Goles and Marcos Kiwi. A lower bound on the computational complexity of the QR decomposition on a shared memory SIMD computer. Parallel Computing, 18(3): 345–354, March 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gatu:2003:PAC

[GK03] Cristian Gatu and Erricos J. Kontoghiorghes. Parallel algorithms for computing all possible subset regression models using the QR decomposition. Parallel Computing, 29(4):505–521, April 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Guralnik:2004:PTP

[GK04] Valerie Guralnik and George Karypis. Parallel treeprojection-based sequence mining algorithms. Parallel Computing, 30(4):443– 472, April 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Giuliani:2017:FCU

[GK17] Andrew Giuliani and Lilia Krivodonova. Face coloring in unstructured CFD codes. Parallel Computing, 63(??):17–37, April 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117300364.

Gowanlock:2019:HCG

[GK19] Michael Gowanlock and Ben Karsin. A hybrid CPU/GPU approach for optimizing sorting throughput. Parallel Computing, 85(??):45–55, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118302515.

Gohberg:1991:TAP

[GKAS91] I. Gohberg, I. Koltracht, A. Averbuch, and B. Shoham. Timing analysis of a parallel algorithm for Toeplitz matrices on a MIMD parallel machine. Parallel Computing, 17(4–5):563–577, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gropp:2001:HPP

[GKKS01] William D. Gropp, Dinesh K. Kaushik, David E. Keyes, and Barry F. Smith. High-performance parallel implicit CFD. Parallel Computing, 27(4): 337–362, March 2001. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/47/
28/24/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/47/
28/24/article.pdf.

Grama:1998:SPF

[GKS98] Ananth Grama, Vipin Kumar, and Ahmed Sameh. Scalable parallel formulations of the Barnes–Hut method for n-body simulations. Parallel Computing, 24(5–6):797–822, June 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/5-6/1288.pdf.

Gaidioz:2007:EHP

[GKS07] Benjamin Gaidioz, Birger Koblitz, and Nuno Santos. Exploring high performance distributed file storage using LDPC codes. Parallel Computing, 33 (4–5):264–274, May 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Galgon:2015:PIS

[GKT^{+}15] Martin Galgon, Lukas Krämer, Jonas Thies, Achim Basermann, and Bruno Lang. On the parallel iterative solution of linear systems arising in the FEAST algorithm for computing inner eigenvalues. Parallel Computing, 49(??):153–163, November 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000915.

Gutheil:1994:PPM

[GKV94] Inge Gutheil and Werner Krotz-Vogel. Performance of a parallel matrix multiplication routine on Intel iPSC/860. Parallel Computing, 20(7):953–974, July 12, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/

cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=7&aid=871.

Garbey:1990:MPC

[GL90] Marc Garbey and David Levine. Massively parallel computation of conservation laws. Parallel Computing, 16(2–3):293– 304, December 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gropp:1997:HPM

[GL97] William Gropp and Ewing Lusk. A highperformance MPI implementation on a sharedmemory vector supercomputer. Parallel Computing, 22(11):1513–1526, January 26, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=11&aid=1113.

Grosser:1999:EPR

[GL99] B. Großer and B. Lang.

Efficient parallel reduction to bidiagonal form. Parallel Computing, 25 (8):969–986, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/32/24/21/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/24/21/
                                        article.pdf.

Grover:2008:PSD

[GLD08] Radhika S. Grover, Qiang Li, and H.-P. Dommel. Performance study of data layout schemes for a SANbased video server. Parallel Computing, 34(12): 747–756, December 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gropp:1996:HPP

[GLDS96] William Gropp, Ewing Lusk, Nathan Doss, and Anthony Skjellum. Highperformance, portable implementation of the MPI Message Passing Interface Standard. Parallel Computing, 22(6): 789–828, September 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=6&aid=1075.

Gatani:2006:EDA

[GLG06] Luca Gatani, Giuseppe Lo Re, and Salvatore Gaglio. An efficient distributed algorithm for generating and updating multicast trees. Parallel Computing, 32 (11–12):777–793, December 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Garcia-Lopez:2003:PSS

[GLMBMPMV03] Félix García-López, Belén Melián-Batista, José A. Moreno-Pérez, and J. Marcos Moreno-Vega. Parallelization of the scatter search for the p-median problem. Parallel Computing, 29(5):575–589, May 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

George:1989:CRP

[GLN89] Alan George, Joseph W. H.

Liu, and Esmond Ng. Communication results for parallel sparse Cholesky factorization on a hypercube. Parallel Computing, 10(3):287–298, May 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Globisch:1995:APG

[Glo95a] Gerhard Globisch. On an automatically parallel generation technique for tetrahedral meshes. Parallel Computing, 21 (12):1979–1995, December 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=12&aid=1033.

Globisch:1995:PPM

[Glo95b] Gerhard Globisch. PARMESH — a parallel mesh generator. Parallel Computing, 21(3):509–524, March 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=3&aid=951.

Golubchik:1998:SAF

[GLP98] Leana Golubchik, John C. S. Lui, and Maria Papadopouli. Survey of approaches to fault tolerant design of VOD servers: techniques, analysis and comparison. Parallel Computing, 24(1):123–155, March 10, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1998&volume=
24&issue=1&aid=1277.

Gamba:2008:LMA

[GLP08] Paolo Gamba, Luca Lombardi, and Marco Porta.

Log-map analysis. Parallel Computing, 34(12): 757–764, December 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gusken:1999:LQT

[GLS99] Stephan Güsken, Thomas Lippert, and Klaus Schilling. Lattice QCD with two dynamical Wilson fermions on APE100 parallel systems. Parallel Computing, 25(10–11):1227–1242, September 1, 1999. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/32/
26/19/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/32/
26/19/article.pdf.

Gendreau:2001:DMP

[GLS01] Michel Gendreau, Gilbert Laporte, and Frédéric Semet. A dynamic model and parallel tabu search heuristic for realtime ambulance relocation. Parallel Computing, 27(12):1641–1653, November 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/42/32/
abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/10/35/21/47/42/32/
                                        article.pdf.

Guermouche:2003:IRM

[GLU03] Abdou Guermouche, Jean-Yves L’Excellent, and Gil Utard. Impact of reordering on the memory of a multifrontal solver. Parallel Computing, 29(9):1191– 1218, September 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Guerra:1989:SSA

[GM89] Concettina Guerra and Rami Melhem. Synthesis of systolic algorithm design. Parallel Computing, 12(2):195–207, November 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ghandeharizadeh:1998:DIS

[GM98] Shahram Ghandeharizadeh and Richard Muntz. Design and implementation of scalable continuous media servers. Parallel Computing, 24(1):91–122, March 10, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1998&volume=
                                        24&issue=1&aid=1276.

Guerriero:2003:CPR

[GM03] F. Guerriero and M. Mancini.

A cooperative parallel rollout algorithm for the sequential ordering problem. Parallel Computing, 29(5): 663–677, May 2003. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Green:2004:EMS

[GM04a] Mark L. Green and Russ Miller. Evolutionary molecular structure determination using gridenabled data mining. Parallel Computing, 30(9–10): 1057–1071, September/ October 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Green:2004:MSD

[GM04b] Mark L. Green and Russ Miller. Molecular structure determination on a computational and data grid. Parallel Computing, 30(9–10):1001–1017, September/October 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Govindaraju:2007:CEN

[GM07] Naga K. Govindaraju and Dinesh Manocha. Cacheefficient numerical algorithms using graphics hardware. Parallel Computing, 33(10–11):663–684, November 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Garcia:1990:PQI

[GMBZ90] I. Garcia, J. J. Merelo, J. D. Bruguera, and E. L. Zapata. Parallel quadrant interlocking factorization on hypercube computers. Parallel Computing, 15(1–3):87–100, September 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

George:2000:RTS

[GMF00] Alan D. George, Jeff Markwell, and Ryan Fogarty. Real-time sonar beamforming on highperformance distributed computers. Parallel Computing, 26(10):1231–1252, August 15, 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/31/23/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/31/23/
article.pdf.

Gmys:2016:GBB

[GMMT16] J. Gmys, M. Mezmaz, N. Melab, and D. Tuyttens. A GPU-based branch-and-bound algorithm using integer–vector– matrix data structure. Parallel Computing, 59 (??):119–139, November 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000387.

Guinand:2004:SAT

[GMS04] Frédéric Guinand, Aziz Moukrim, and Eric Sanlaville. Sensitivity analysis of tree scheduling on two machines with communication delays. Parallel Computing, 30(1):103–120, January 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gallivan:1996:SLN

[GMW96] K. A. Gallivan, B. A. Marsolf, and H. A. G. Wijshoff. Solving large nonsymmetric sparse linear systems using MCSPARSE. Parallel Computing, 22 (10):1291–1333, December 15, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=10&aid=1098.

Guo:2000:CFC

[GNY00] Minyi Guo, Ikuo Nakata, and Yoshiyuki Yamashita. Contention-free communication scheduling for array redistribution. Parallel Computing, 26(10): 1325–1343, August 15, 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/31/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/31/27/
article.pdf.

Glickfeld:1988:GSS

[GO88] B. W. Glickfeld and R. A. Overbeek. Geometric specification of scheduling constraints: a simplified approach to multiprocessing. Parallel Computing, 6(3): 325–337, March 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Godart:2000:PIT

[God00]  Cyril Godart.  Paral-

lel implementation of a two-factor Cheyette-beta model calibration. Parallel Computing, 26(5):569– 586, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/27/26/
abstract.html;   http:
//www.elsevier.nl/gej-

ng/10/35/21/42/27/26/
article.pdf.

Goglin:2011:HPM

[Gog11]  Brice Goglin.  High-

performance message-passing over generic Ethernet hardware with Open-MX. Parallel Computing, 37(2): 85–100, February 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Goldapp:1986:FSL

[Gol86] M. Goldapp. Fast scan-line conversion using vectorisation. Parallel Computing, 3(2):141–152, May 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Goldmann:1988:VMS

[Gol88] Mark Goldmann. Vectorisation of the multiple shooting method for the nonlinear boundary value problem in ordinary differential equations. Parallel Computing, 7(1):97– 110, April 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gorlatch:1997:GST

[Gor97]  Sergei Gorlatch.  N-

graphs: Scalable topology and design of balanced divide-and-conquer algorithms. Parallel Computing, 23(6):687–698, June 20, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=6&aid=1182.

Gottlieb:1989:PQC

[Got89] Israel Gottlieb. The partitioning of QSDF computation graphs. Parallel Computing, 9(3):347–358, February 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gajski:1985:CFM

[GP85] D. D. Gajski and J. K. Peir. Comparison of five multiprocessor systems. Parallel Computing, 2(3): 265–282, November 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gannon:1986:RSC

[GP86] D. Gannon and J. Panetta. Restructuring SIMPLE for the CHiP architecture. Parallel Computing, 3(4): 305–326, October 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gao:1990:NCA

[GP90] Feng Gao and Beresford N. Parlett. A note on communication analysis of parallel sparse Cholesky factorization on a hypercube. Parallel Computing, 16(1):59– 60, November 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gasieniec:1996:ABF

[GP96] Leszek Gaşieniec and Andrzej Pelc. Adaptive broadcasting with faulty nodes. Parallel Computing, 22(6): 903–912, September 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=6&aid=1074.

Gu:1998:NSS

[GP98] Qian-Ping Gu and Shietung Peng. Node-to-set and set-to-set cluster fault tolerant routing in hypercubes. Parallel Computing, 24(8):1245–1261, August 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/8/1327.pdf.

Gaudiot:1988:PGA

[GPC88] J. L. Gaudiot, J. I. Pi, and M. L. Campbell. Program graph allocation in distributed multicomputers. Parallel Computing, 7(2):227–247, June 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Geneva:2017:SIR

[GPLW17] Nicholas Geneva, Cheng Peng, Xiaoming Li, and Lian-Ping Wang. A scalable interface-resolved simulation of particleladen flow using the lattice Boltzmann method. Parallel Computing, 67 (??):20–37, September 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300960.

Gupta:1999:CTP

[GPPS99] Rajiv Gupta, Santosh Pande, Kleanthis Psarris, and Vivek Sarkar. Compilation techniques for parallel systems. Parallel Computing, 25(13– 14):1741–1783, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/36/33/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/33/
article.pdf.

Gendron:2003:PHH

[GPS03] Bernard Gendron, Jean-Yves Potvin, and Patrick Soriano. A parallel hybrid heuristic for the multicommodity capacitated location problem with balancing requirements. Parallel Computing, 29(5):591– 606, May 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Grigori:2008:PMA

[GPS^{+}08] Laura Grigori, Bernard Philippe, Ahmed Sameh, Damien Tromeur-Dervout, and Marian Vajtersic. Parallel matrix algorithms and applications. Parallel Computing, 34(6–8):293– 295, July 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Grinberg:2009:PPC

[GPSK09] L. Grinberg, D. Pekurovsky, S. J. Sherwin, and G. E. Karniadakis. Parallel performance of the coarse space linear vertex solver and low energy basis preconditioner for spectral/ hp elements. Parallel Computing, 35(5):284–304, May 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gothandaraman:2008:FAQ

[GPW^{+}08] Akila Gothandaraman, Gregory D. Peterson, G. L. Warren, Robert J. Hinde, and Robert J. Harrison. FPGA acceleration of a quantum Monte Carlo application. Parallel Computing, 34(4–5):278–291, May 2008. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gutierrez:2008:AML

[GPZ08] Eladio Gutiérrez, Oscar Plata, and Emilio L. Zapata. An analytical model of locality-based parallel irregular reductions. Parallel Computing, 34(3):133– 157, March 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Galligani:1989:SLS

[GR89] Ilio Galligani and Valeria Ruggiero. Solving large systems of linear ordinary differential equations on a vector computer. Parallel Computing, 9(3):359–365, February 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Geus:2001:TFP

[GR01] Roman Geus and Stefan Röllin. Towards a fast parallel sparse symmetric matrix-vector multiplication. Parallel Computing, 27(7):883–896, June 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/31/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/31/25/
article.pdf.

Gutknecht:2002:CIR

[GR02] Martin H. Gutknecht and Stefan Röllin. The Chebyshev iteration revisited. Parallel Computing, 28(2):263–283, February 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/33/33/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/33/33/
main.pdf.

Grassl:1991:PPA

[Gra91]  C. M. Grassl.  Paral-

lel performance of applications on supercomputers. Parallel Computing, 17(10–11):1257–1273, December 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Grant:2015:SCR

[GRBA15] Ryan E. Grant, Mohammad J. Rashti, Pavan Balaji, and Ahmad Afsahi. Scalable connectionless RDMA over unreliable datagrams. Parallel Computing, 48(??):15–39, October 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000617.

Greenbaum:1989:SCM

[Gre89] A. Greenbaum. Synchronization costs on multiprocessors. Parallel Computing, 10(1):3–14, March 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gropp:1987:SPL

[Gro87] William D. Gropp. Solving PDEs on loosely-coupled parallel processors. Parallel Computing, 5(1–2):165– 173, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Gropp:2019:UNS

[Gro19] William D. Gropp. Using node and socket information to implement MPI Cartesian topologies. Parallel Computing, 85(??):98–108, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303156.

Garmon:2022:RAT

[GRP22] Andrew Garmon, Vinay Ramakrishnaiah, and Danny Perez. Resource allocation for task-level speculative scientific applications: a proof of concept using Parallel Trajectory Splicing. Parallel Computing, 112(??):??, September 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000369.

Gallopoulos:1989:PBC

[GS89] E. Gallopoulos and Y. Saad.

A parallel block cyclic reduction algorithm for the fast solution of elliptic equations. Parallel Computing, 10(2):143–159, April 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gupta:1996:CEM

[GS96] Himanshu Gupta and P. Sadayappan. Communicationefficient matrix multiplication on hypercubes. Parallel Computing, 22 (1):75–99, February 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=1&aid=1037.

Gerbessiotis:2001:MBM

[GS01] Alexandros V. Gerbessiotis and Constantinos J. Siniolakis. Merging on the BSP model. Parallel Computing, 27(6):809– 822, May 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/30/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/30/26/
article.pdf.

Gonzalez:2004:CNN

[GS04] Teofilo F. Gonzalez and David Serena. n-cube network: node disjoint shortest paths for maximal distance pairs of vertices. Parallel Computing, 30(8):973–998, August 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Grelck:2006:MCA

[GS06] Clemens Grelck and Sven-Bodo Scholz. Merging compositions of array skeletons in SaC. Parallel Computing, 32(7–8): 507–522, September 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gilli:2010:DOP

[GS10] Manfred Gilli and Enrico Schumann. Distributed optimisation of a portfolio’s Omega. Parallel Computing, 36(7):381–389, July 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ghysels:2022:HPS

[GS22] Pieter Ghysels and Ryan Synk. High performance sparse multifrontal solvers on modern GPUs. Parallel Computing, 110(??):??, May 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000059.

Gu:2022:ODU

[GSC^{+}22] Rong Gu, Jun Shi, Xiaofei Chen, Zhaokang Wang, Yang Che, Kai Zhang, and Yihua Huang. Octopus-DF: Unified DataFramebased cross-platform data analytic system. Parallel Computing, 110(??):??, May 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121001186.

Gardner:2013:CCE

[GScFM13] Mark Gardner, Paul Sathre, Wu chun Feng, and Gabriel Martinez. Characterizing the challenges and evaluating the efficacy of a CUDA-to-OpenCL translator. Parallel Computing, 39(12):769–786, December 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001075.

Gajinov:2015:DBS

[GSE^{+}15] Vladimir Gajinov, Srdjan Stipić, Igor Erić, Osman S. Unsal, Eduard Ayguadé, and Adrian Cristal. DaSH: a benchmark suite for hybrid dataflow and shared memory programming models. Parallel Computing, 45(??):18–48, June 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819115000496.

Goddeke:2007:EWS

[GSMY^{+}07] Dominik Göddeke, Robert Strzodka, Jamaludin Mohd-Yusof, Patrick McCormick, Sven H. M. Buijssen, Matthias Grajewski, and Stefan Turek. Exploring weak scalability for FEM calculations on a GPUenhanced cluster. Parallel Computing, 33(10–11): 685–699, November 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gentzsch:1988:UPF

[GSZ88] W. Gentzsch, F. Szelenyi, and V. Zecca. Use of parallel FORTRAN for engineering problems on the IBM 3090 vector multiprocessor. Parallel Computing, 9(1):107–115, December 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gu:1990:SAP

[GT90] Qian Ping Gu and Tadao Takaoka. A sharper analysis of a parallel algorithm for the all pairs shortest path problem. Parallel Computing, 16(1):61– 67, November 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gusev:1992:CAM

[GT92] M. Gusev and J. Tasic. Comparative analysis of methods for broadcast elimination. Parallel Computing, 18(8):857–866, August 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gropp:2007:TSM

[GT07] William Gropp and Rajeev Thakur. Thread-safety in an MPI implementation: Requirements and analysis. Parallel Computing, 33(9):595–604, September 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gropp:2019:GEI

[GT19] William Gropp and Rajeev Thakur. Guest Editor’s introduction: Special issue on best papers from EuroMPI/USA 2017. Parallel Computing, 84(??):62, May 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819119300560.

Gates:2018:AST

[GTD18] Mark Gates, Stanimire Tomov, and Jack Dongarra. Accelerating the SVD two stage bidiagonal reduction and divide and conquer using GPUs. Parallel Computing, 74(??):3– 18, ???? 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301758.

Gong:2020:NMG

[GTW^{+}20] Dunwei Gong, Tian Tian, Jinxin Wang, Ying Du, and Zheng Li. A novel method of grouping target paths for parallel programs. Parallel Computing, 97(??):Article 102665, September 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300582.

Gupta:1999:GPM

[Gup99] R. Gupta. General physics motivations for numerical simulations of quantum field theory. Parallel Computing, 25(10– 11):1199–1215, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/26/17/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/26/17/
article.pdf.

Gurke:1988:ASE

[Gur88] R. Gurke. The approximate solution of the Euclidean traveling salesman problem on a CRAY X-MP. Parallel Computing, 8(1–3):177–183, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gusken:1999:SET

[Güs99] Stephan Güsken. Stochastic estimator techniques and their implementation on distributed parallel computers. Parallel Computing, 25(10– 11):1371–1381, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/32/26/27/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/26/27/
                                        article.pdf.

Gutheil:1988:SSS

[Gut88] I. Gutheil. SUPRENUM software for the symmetric eigenvalue problem. Parallel Computing, 7(3):419– 424, September 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Garbey:2001:PSU

[GV01] M. Garbey and Yu. V.

Vassilevski. A parallel solver for unsteady incompressible 3D Navier– Stokes equations. Parallel Computing, 27(4):363– 389, March 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/28/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/28/25/
article.pdf.

Gonzalez-Velez:2006:SAS

[GV06] Horacio González-Vélez.

Self-adaptive skeletal task farm for computational grids. Parallel Computing, 32(7–8):479–490, September 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Goh:2008:DPE

[GV08] Lee Kee Goh and Bharadwaj Veeravalli. Design and performance evaluation of combined first-fit task allocation and migration strategies in mesh multiprocessor systems. Parallel Computing, 34(9):508–520, September 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Giraud:2011:PPW

[GV11] Mathieu Giraud and Jean-Stéphane Varré. Parallel Position Weight Matrices algorithms. Parallel Computing, 37(8):466–478, August 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819110001389.

Ghysels:2014:HGS

[GV14] P. Ghysels and W. Vanroose. Hiding global synchronization latency in the preconditioned conjugate gradient algorithm. Parallel Computing, 40(7):224– 238, July 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000719.

Gu:2008:EDF

[GVH08] Yongfeng Gu, Tom Van-Court, and Martin C. Herbordt. Explicit design of FPGA-based coprocessors for short-range force computations in molecular dynamics simulations. Parallel Computing, 34(4–5): 261–277, May 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gonzalez:1987:DDE

[GW87] Ruth Gonzalez and Mary Fanett Wheeler. Domain decomposition for elliptic partial differential equations with Neumann boundary conditions. Parallel Computing, 5(1–2):257–263, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Geschiere:1995:ELG

[GW95] J. P. Geschiere and H. A. G.

Wijshoff. Exploiting large grain parallelism in a sparse direct linear system solver. Parallel Computing, 21(8):1339–1364, August 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=8&aid=1003.

Goscinski:2008:SCE

[GW08] Andrzej M. Goscinski and Adam K. L. Wong. A study of the concurrent execution of parallel and sequential applications on a non-dedicated cluster. Parallel Computing, 34(2): 69–91, February 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Giolmas:1999:ACM

[GWC^{+}99] Nicholas Giolmas, Daniel W. Watson, David M. Chelberg, Peter V. Henstock, June Ho Yi, and Howard Jay Siegel. Aspects of computational mode and data distribution for parallel range image segmentation. Parallel Computing, 25(5):499– 523, May 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1999/
                                        25/5/1399.pdf.

Gao:2005:PMV

[GWLS05] Jinzhu Gao, Chaoli Wang, Liya Li, and Han-Wei Shen. A parallel multiresolution volume rendering algorithm for large data visualization. Parallel Computing, 31(2):185–204, February 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Geimer:2009:STA

[GWWM09] Markus Geimer, Felix Wolf, Brian J. N. Wylie, and Bernd Mohr. A scalable tool architecture for diagnosing wait states in massively parallel applications. Parallel Computing, 35(7):375–388, July 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gan:1994:PAR

[GYH94] Qi Gan, Qing Yang, and Chen-Yi Hu. Parallel all-row preconditioned interval linear solver for nonlinear equations on multiprocessors. Parallel Computing, 20(9): 1249–1268, September 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=9&aid=890.

Gorb:2000:HPC

[GYL00] Leonid Gorb, Ilya Yanov, and Jerzy Leszczynski. High performance computing on the Cray T3E and IBM SP2 systems with the parallel version of GAUS-SIAN 94. Parallel Computing, 26(7–8):1043–1060, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/29/36/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/29/36/
article.pdf.

Griebel:1999:PMA

[GZ99] Michael Griebel and Gerhard Zumbusch. Parallel multigrid in an adaptive PDE solver based on hashing and space-filling curves. Parallel Computing, 25(7):827–843, August 13, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1999&volume=
                                        25&issue=7&aid=1412;
                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/32/
                                        23/22/abstract.html;
                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/32/
                                        23/22/article.pdf.

Gao:2023:WNO

[GZH^{+}23] Yunqi Gao, Zechao Zhang, Bing Hu, A-Long Jin, and Chunming Wu. OF-WFBP: a near-optimal communication mechanism for tensor fusion in distributed deep learning. Parallel Computing, 118(??):??, November 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819123000595.

Gao:2017:MGP

[GZHX17] Jiaquan Gao, Yuanshen Zhou, Guixia He, and Yifei Xia. A multi-GPU parallel optimization model for the preconditioned conjugate gradient algorithm. Parallel Computing, 63(??):1– 16, April 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300388.

Gualandris:2007:PAD

[GZTR07] Alessia Gualandris, Simon Portegies Zwart, and Alfredo Tirado-Ramos. Performance analysis of direct N. Parallel Computing, 33(3):159–173, April 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Heirich:1997:SMC

[HA97] Alan Heirich and James Arvo. Scalable Monte Carlo image synthesis. Parallel Computing, 23 (7):845–859, July 14, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/free/1997/
23/7/heirich/fig5_2.
jpg.

Hariharan:2005:EPA

[HA05] Bhanu Hariharan and Srinivas Aluru. Efficient parallel algorithms and software for compressed octrees with applications to hierarchical methods. Parallel Computing, 31(3– 4):311–331, March/April 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hallberg:2014:OIR

[HA14] Robert Hallberg and Alistair Adcroft. An orderinvariant real-to-integer conversion sum. Parallel Computing, 40(5–6):140– 143, May 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000507.

Haase:1998:PIC

[Haa98] G. Haase. Parallel incomplete Cholesky preconditioners based on the nonoverlapping data distribution. Parallel Computing, 24(11):1685–1703, October 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/11/1323.pdf.

Hack:1989:PGP

[Hac89] J. J. Hack. On the promise of general-purpose parallel computing. Parallel Computing, 10(3):261–275, May 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hagemann:1991:CSM

[Hag91]  J. Hagemann.  Com-

binatorial structures for multiprocessor-systems. Parallel Computing, 17(6–7): 699–706, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hoeflinger:2001:PSP

[HAJK01] Jay Hoeflinger, Prasad Alavilli, Thomas Jackson, and Bob Kuhn. Producing scalable performance with OpenMP: Experiments with two CFD applications. Parallel Computing, 27(4):391–413, March 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/28/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/28/26/
article.pdf.

Hasegawa:2021:TCA

[HAK^{+}21] Yuta Hasegawa, Takayuki Aoki, Hiromichi Kobayashi, Yasuhiro Idomura, and Naoyuki Onodera. Tree cutting approach for domain partitioning on forestof-octrees-based blockstructured static adaptive mesh refinement with lattice Boltzmann method. Parallel Computing, 108 (??):??, December 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819121000971.

Handler:1985:DCS

[Han85] W. Handler. Dynamic computer structures for manifold utilization. Parallel Computing, 2(1):15– 32, March 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hanzalek:1998:PAG

[Han98] Zdeněk Hanzálek. A parallel algorithm for gradient training of feedforward neural networks. Parallel Computing, 24(5–6):823– 839, June 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/5-6/1312.pdf.

Hasselbring:1990:CCL

[Has90] W. Hasselbring. CELIP: a Cellular Language for Image Processing. Parallel Computing, 14(1):99–109, May 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hascoet:2001:MAP

[Has01]  Laurent Hascoët.  A

method for automatic placement of communications in SPMD parallelisation. Parallel Computing, 27(13):1655–1664, December 1, 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/43/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/43/27/
article.pdf.

Hatzky:2006:DCP

[Hat06] R. Hatzky. Domain cloning for a particle-in-cell (PIC) code on a cluster of symmetric-multiprocessor (SMP) computers. Parallel Computing, 32(4):325– 330, April 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hsiung:1984:NSD

[HB84] C. C. Hsiung and W. Butscher.

A numerical seismic 3-D migration model for vector multiprocessors. Parallel Computing, 1(2):113–120, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Heermann:1990:PIM

[HB90] D. W. Heermann and A. N. Burkitt. Parallelization of the Ising model and its performance evaluation. Parallel Computing, 13(3):345–357, March 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hu:1994:NEP

[HB94] Y. F. Hu and R. J. Blake. Numerical experiences with partitioning of unstructured meshes. Parallel Computing, 20 (5):815–829, May 11, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=5&aid=837.

Hu:1999:IDA

[HB99] Y. F. Hu and R. J. Blake. An improved diffusion algorithm for dynamic load balancing. Parallel Computing, 25(4):417–444, April 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/4/1394.pdf.

Heinecke:2019:TOH

[HBC19] Alexander Heinecke, Alexander Breuer, and Yifeng Cui. Tensor-optimized hardware accelerates fused discontinuous Galerkin simulations. Parallel Computing, 89(??):Article 102550, November 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819119301413.

Herrmann:2016:ACR

[HBH^{+}16] Julien Herrmann, George Bosilca, Thomas Hérault, Loris Marchal, Yves Robert, and Jack Dongarra. Assessing the cost of redistribution followed by a computational kernel: Complexity and performance results. Parallel Computing, 52(??):22–41, February 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001192.

Hockney:1989:PCM

[HC89] R. W. Hockney and I. J.

Curington. f_{1/2}: a parameter to characterize memory and communication bottlenecks. Parallel Computing, 10(3):277– 286, May 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hockney:1992:CCI

[HC92a] R. W. Hockney and E. A. Carmona. Comparison of communications on the Intel iPSC/860 and Touchstone Delta (short communication). Parallel Computing, 18(9):1067–1072, September 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hung:1992:DAQ

[HC92b] Yung Chen Hung and Gen Huey Chen. Distributed algorithms for the quickest path problem. Parallel Computing, 18(7):823–834, July 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hu:1994:ERA

[HC94] Shu Hua Hu and Hsing Lung Chen. An effective routing algorithm in incomplete hypercubes. Parallel Computing, 20 (12):1721–1738, November 28, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.

elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=12&aid=926.

Hsieh:2004:PMC

[HC04] Sun-Yuan Hsieh and Chun-Hua Chen. Pancyclicity on Möbius cubes with maximal edge faults. Parallel Computing, 30(3):407– 421, March 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hutter:2005:DLP

[HC05] Jürg Hutter and Alessandro Curioni. Dual-level parallelism for ab initio molecular dynamics: Reaching teraflop performance with the CPMD code. Parallel Computing, 31(1):1–17, January 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hempel:1996:RAN

[HCH^{+}96] R. Hempel, R. Calkin, R. Hess, W. Joppich, C. W. Oosterlee, H. Ritzdorf, P. Wypior, W. Ziegler, N. Koike, T. Washio, and U. Keller. Real applications on the new parallel system NEC Cenju-3. Parallel Computing, 22(1):131–148, February 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=1&aid=1039.

Hawick:2003:DFP

[HCJ03] Kenneth A. Hawick, P. D. Coddington, and H. A. James. Distributed frameworks and parallel algorithms for processing largescale geographic data. Parallel Computing, 29(10): 1297–1333, October 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Horng:1994:PRA

[HCK94] M. S. Horng, D. J. Chen, and Kuo Lung Ku. Parallel routing algorithms for incomplete hypercube interconnection networks. Parallel Computing, 20 (12):1739–1761, November 28, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=12&aid=927.

Huang:2005:TME

[HCL05] Lei Huang, Barbara Chapman, and Zhenying Liu. Towards a more efficient implementation of OpenMP for clusters via translation to global arrays. Parallel Computing, 31(10–12):1114–1139, October/December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Heras:2001:MDL

[HCR01] D. B. Heras, J. C. Cabaleiro, and F. F. Rivera. Modeling data locality for the sparse matrixvector product using distance measures. Parallel Computing, 27(7):897– 912, June 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/31/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/31/26/
article.pdf.

He:2016:NEC

[HCT16] Jiangzhou He, Wenguang Chen, and Zhizhong Tang. NestedMP: Enabling cacheaware thread mapping for nested parallel shared memory applications. Parallel Computing, 51(??): 56–66, January 2016. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S016781911500143X.

Hsu:2005:FHA

[HCTH05] Hong-Chun Hsu, Liang-Chih Chiang, Jimmy J. M. Tan, and Lih-Hsing Hsu. Fault hamiltonicity of augmented cubes. Parallel Computing, 31(1):131–145, January 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hey:1997:RPP

[HDH97] Tony Hey, Alistair Dunlop, and Emilio Hernández. Realistic parallel performance estimation. Parallel Computing, 23(1–2):5–21, April 16, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=1-2&aid=1144.

Hatzopoulos:1988:CPS

[HE88] M. Hatzopoulos and D. J. Evans. Comments on the paper: “A short proof for the existence of the WZ-factorisation” [Parallel Comput. 4 (1987), no. 2, 229–232, MR 88j:65064a] by M. Kaps and M. Schlegl. Parallel Computing, 6(2):259, February 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). See [KS87].

Hebeker:1993:PC

[Heb93] F.-H Hebeker. Parallel CFD’93. Parallel Computing, 19(12):1431–??, December 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hu:1996:CPC

[HEB96] Y. F. Hu, D. R. Emerson, and R. J. Blake. The communication performance of the Cray T3D and its effect on iterative solvers. Parallel Computing, 22 (6):829–844, September 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=6&aid=1086.

Heen:1997:ECS

[Hee97] Olivier Heen. Efficient constant speed-up for one dimensional cellular automata calculators. Parallel Computing, 23(11): 1663–1671, December 1, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=11&aid=1236.

Hegland:1996:RCF

[Heg96] Markus Hegland. Real and complex fast Fourier transforms on the Fujitsu VPP 500. Parallel Computing, 22(4):539–553, June 11, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=4&aid=1066.

Hempel:1994:API

[Hem94] R. Hempel. Application programming interfaces for SUPRENUM. Parallel Computing, 20(10– 11):1519–1526, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=10-11&aid=910.

Hendrickson:1993:PFU

[Hen93] B. Hendrickson. Parallel QR factorization using the torus-wrap mapping. Parallel Computing, 19(11):1259–1271, November 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Herzog:1988:PEP

[Her88] U. Herzog. Performance evaluation principles for vector- and multiprocessor systems. Parallel Computing, 7(3):425–438, September 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hermenegildo:2000:PIP

[Her00] Manuel Hermenegildo. Parallelizing irregular and pointer-based computations automatically: Perspectives from logic and constraint programming. Parallel Computing, 26 (13–14):1685–1708, December 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/34/23/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/34/23/
article.pdf.

Hey:1991:GDM

[Hey91]  A. J. G. Hey.  The

genesis distributed memory benchmarks. Parallel Computing, 17(10– 11):1275–1283, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hey:1994:GEP

[Hey94] Tony Hey. The Genesis Esprit project — an overview. Parallel Computing, 20(10– 11):1605–1612, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=916.

Hussain:2020:GBP

[HF20] Md Maruf Hussain and Noriyuki Fujimoto. GPUbased parallel multi-objective particle swarm optimization for large swarms and high dimensional problems. Parallel Computing, 92(??):Article 102589, April 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301802.

Hidalgo:2010:PAB

[HFL^{+}10] J. Ignacio Hidalgo, Francisco Fernandez, Juan Lanchares, Erick Cantú-Paz, and Albert Zomaya. Parallel architectures and bioinspired algorithms. Parallel Computing, 36(10– 11):553–554, October/ November 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hursey:2012:AFA

[HG12] Joshua Hursey and Richard L.

Graham. Analyzing fault aware collective performance in a process fault tolerant MPI. Parallel Computing, 38(1–2): 15–25, January/February 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001414.

Heroux:2017:SIS

[HG17] Michael A. Heroux and C. Kristopher Garrett. Special issue on SC16 Student Cluster Competition Reproducibility Initiative. Parallel Computing, 70(??):3–4, December 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301643.

Hoefler:2007:OCG

[HGLR07] Torsten Hoefler, Peter Gottschling, Andrew Lumsdaine, and Wolfgang Rehm. Optimizing a conjugate gradient solver with nonblocking collective operations. Parallel Computing, 33(9):624–633, September 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Huang:2020:EAM

[HGM20] Yu Huang, Kai Gong, and Eric Mercer. An efficient algorithm for match pair approximation in message passing. Parallel Computing, 91(??):Article 102585, March 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819119301760.

Hruz:2010:PSM

[HGS10] Tomas Hruz, Stefan Geisseler, and Marcel Schöngens. Parallelism in simulation and modeling of scale-free complex networks. Parallel Computing, 36(8):469– 485, August 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hake:1989:LAS

[HH89] J.-F. Hake and W. Homberg. Linear algebra software on a vector computer. Parallel Computing, 10(1):65– 81, March 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hake:1991:IMO

[HH91] J.-Fr. Hake and W. Homberg.

The impact of memory organization on the performance of matrix calculations. Parallel Computing, 17(2–3):311–327, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hutter:2017:PRS

[HH17] Edward Hutter and Chung-Ting Huang. ParConnect: Results from the student cluster competition at SC16. Parallel Computing, 70(??):35–40, December 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301175.

Hussain:2015:AAM

[HHGA15] Tassadaq Hussain, Amna Haider, Shakaib A. Gursal, and Eduard Ayguadé. AMC: Advanced Multiaccelerator Controller. Parallel Computing, 41(??): 14–30, January 2015. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819114001264.

Hsu:1992:FPS

[HHH92] Su Chu Hsu, Hsien Fen Hsieh, and Shing Tsaan Huang. A fully-pipelined systolic algorithm for finding bridges on an undirected connected graph. Parallel Computing, 18(4): 377–391, April 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hupp:2016:GCS

[HHJP16] Philipp Hupp, Mario Heene, Riko Jacob, and Dirk Pflüger. Global communication schemes for the numerical solution of high-dimensional PDEs. Parallel Computing, 52(??):78–105, February 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001623.

Hambrusch:1995:COC

[HHK95] Susanne E. Hambrusch, Farooq Hameed, and Ashfaq A. Khokhar. Communication operations on coarse-grained mesh architectures. Parallel Computing, 21(5):731–751 (or 731–752??), May 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.

cgi?year=1995&volume=
21&issue=5&aid=976.

Hermanns:2019:MEI

[HHK^{+}19] Marc-André Hermanns, Nathan T. Hjelm, Michael Knobloch, Kathryn Mohror, and Martin Schulz. The MPI T events interface: an early evaluation and overview of the interface. Parallel Computing, 85(??):119–130, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303314.

Hidalgo-Herrero:2006:AIM

[HHOMR06] Mercedes Hidalgo-Herrero, Yolanda Ortega-Mallén, and Fernando Rubio. Analyzing the influence of mixed evaluation on the performance of Eden skeletons. Parallel Computing, 32(7–8):523–538, September 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Huang:2006:COP

[HHQ06] Chun-Hsi Huang, Xin He, and Min Qian. Communicationoptimal parallel parenthesis matching. Parallel Computing, 32(1):14–23, January 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Houstis:1987:PPC

[HHR87] Catherine E. Houstis, Elias N. Houstis, and John R. Rice. Partitioning PDE computations: methods and performance evaluation. Parallel Computing, 5(1–2):141– 163, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Hoshino:1988:MSP

[HHSM88] T. Hoshino, R. Hiromoto, S. Sekiguchi, and S. Majima. Mapping schemes of the particle-in-cell method implemented on the PAX computer. Parallel Computing, 9(1):53–75, December 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Holey:1992:IAP

[HI92] J. Andrew Holey and Oscar H. Ibarra. Iterative algorithms for the planar convex hull problem on mesh-connected arrays. Parallel Computing, 18(3): 281–296, March 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hill:1992:PLA

[Hil92] J. M. D. Hill. Parallel lexical analysis and parsing on the AMT distributed array processor. Parallel Computing, 18(6):699–714, June 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hioki:1996:CSL

[Hio96] S. Hioki. Construction of staples in lattice gauge theory on a parallel computer. Parallel Computing, 22(10):1335–1344, December 15, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=10&aid=1104.

Hiromoto:1986:SIP

[Hir86] Robert Hiromoto. Some issues in parallel processing as encountered on the Denelcor HEP. Parallel Computing, 3(2):111–127, May 1986. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hancu:1992:ERE

[HISS92] M. V. A. Hancu, K. Iwasaki, Y. Sato, and M. Sugie. Experimental results on the error detection capability of a concurrent test architecture for massivelyparallel computers. Parallel Computing, 18(10): 1079–1103, October 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Heise:1997:PSN

[HJ97a] Bodo Heise and Michael Jung. Parallel solvers for nonlinear elliptic problems based on domain decomposition ideas. Parallel Computing, 22(11):1527–1544, January 26, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=11&aid=1106.

Hess:1997:CPM

[HJ97b] R. Hess and W. Joppich. A comparison of parallel multigrid and a fast Fourier transform algorithm for the solution of the Helmholtz equation in numerical weather prediction. Parallel Computing, 22(11):1503–1512, January 26, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/

cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=11&aid=1109.

Hodgson:1997:DDP

[HJ97c] D. C. Hodgson and P. K.

Jimack. A domain decomposition preconditioner for a parallel finite element solver on distributed unstructured grids. Parallel Computing, 23(8):1157–1181, July 25, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=8&aid=1205.

Hagras:2005:HPL

[HJ05] T. Hagras and J. Janeček.

A high performance, low complexity algorithm for compile-time task scheduling in heterogeneous systems. Parallel Computing, 31(7):653–670, July 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hori:2021:ISM

[HJB^{+}21] Atsushi Hori, Emmanuel Jeannot, George Bosilca, Takahiro Ogura, Balazs Gerofi, Jie Yin, and Yutaka Ishikawa. An international survey on MPI users. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000983.

Halver:2023:UHG

[HJS23] Rene Halver, Christoph Junghans, and Godehard Sutmann. Using heterogeneous GPU nodes with a Cabana-based implementation of MPCD. Parallel Computing, 117(??):??, September 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781912300039X.

Huang:1990:DSS

[HK90a] Jau Hsiung Huang and Leonard Kleinrock. Distributed selectsort sorting algorithms on broadcast communication networks. Parallel Computing, 16(2–3):183–190, December 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Huang:1990:OPM

[HK90b] Jau-Hsiung Huang and Leonard Kleinrock. Optimal parallel merging and sorting algorithms using \sqrt{N} processors without memory contention. Parallel Computing, 14(1):89– 97, May 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hutchinson:1991:PAS

[HK91] D. Hutchinson and B. M. S.

Khalaf. Parallel algorithms for solving initial value problems: front broadening and embedded parallelism. Parallel Computing, 17(9):957–968, November 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hege:1992:PC

[HK92a] H.-C. Hege and R. Knecht.

Parallel computing 91. Parallel Computing, 18(4): 473–??, April 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Horton:1992:TPM

[HK92b] G. Horton and R. Knirsch.

A time-parallel multigridextrapolation method for parabolic partial differential equations. Parallel Computing, 18(1):21– 29, January 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Holvoet:1998:BSP

[HK98] Tom Holvoet and Thilo Kielmann. Behaviour specification of parallel active objects. Parallel Computing, 24(7):1107–1135, July 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/7/1321.pdf.

Hendrickson:2000:GPM

[HK00] Bruce Hendrickson and Tamara G. Kolda. Graph partitioning models for parallel computing. Parallel Computing, 26(12): 1519–1534, November 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/42/
33/23/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/42/
33/23/article.pdf.

Hanen:2002:MVS

[HK02a] Claire Hanen and Alix Munier Kordon. Minimizing the volume in scheduling an out-tree with communication delays and duplication. Parallel Computing, 28(11):1573–1585, November ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/61/29/
abstract.html.

Hsiao:2002:IED

[HK02b] Hung-Chang Hsiao and Chung-Ta King. Implementation and evaluation of directory hints in CC-NUMA multiprocessors. Parallel Computing, 28(1):107–132, January 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/27/33/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/27/33/
00001681.pdf.

Hofmann:2006:PGS

[HK06] Marc Hofmann and Erricos John Kontoghiorghes. Pipeline Givens sequences for computing the QR decomposition on a EREW PRAM. Parallel Computing, 32(3):222–230, March 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Haring:1997:DPS

[HKK97a] G. Haring, P. Kacsuk, and G. Kotsis. Distributed and parallel systems: Environments and tools. Parallel Computing, 22(13):1699–1701, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        22&issue=13&aid=1122.

Haring:1997:PWD

[HKK97b] G. Haring, P. Kacsuk, and G. Kotsis, editors. Proceedings of the 1996 Workshop on Distributed and Parallel Systems, DAPSYS, volume 22(13) of Parallel Computing. North-Hol-land, Amsterdam, The Netherlands, February 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Haase:2001:PMM

[HKL01] Gundolf Haase, Michael Kuhn, and Ulrich Langer. Parallel multigrid 3D Maxwell solvers. Parallel Computing, 27(6):761– 775, May 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/30/24/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/30/24/
                                        article.pdf.

Hossfeld:1989:MEA

[HKN89] F. Hossfeld, R. Knecht, and W. E. Nagel. Multitasking: experience with applications on a CRAY X-MP. Parallel Computing, 12(3):259–283, December 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Huckle:2010:EPI

[HKR^{+}10] T. Huckle, A. Kallischko, A. Roy, M. Sedlacek, and T. Weinzierl. An efficient parallel implementation of the MSPAI preconditioner. Parallel Computing, 36(5–6):273–284, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hamalainen:1997:MSL

[HKSK97] Timo Hamalainen, Harri Klapuri, Jukka Saarinen, and Kimmo Kaski. Mapping of SOM and LVQ algorithms on a tree shape parallel computer system. Parallel Computing, 23(3):271–289, May 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=3&aid=1178.

Hermanns:2013:SIP

[HKW13] Marc-André Hermanns, Sriram Krishnamoorthy, and Felix Wolf. A scalable infrastructure for the performance analysis of passive target synchronization. Parallel Computing, 39(3):132–145, March 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000762.

Harms:1988:EBT

[HL88] U. Harms and H. Luttermann. Experiences in benchmarking the three supercomputers CRAY-1M, CRAY-X/MP, FU-JITSU VP-200 compared with the CYBER 76. Parallel Computing, 6(3):373– 382, March 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hamdi:1997:DLB

[HL97] M. Hamdi and C. K. Lee. Dynamic load-balancing of image processing applications on clusters of workstations. Parallel Computing, 22(11):1477–1492, January 26, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=11&aid=1105.

Hwang:1999:CSG

[HL99] G.-H. Hwang and J. K.

Lee. Communication set generations with CSD calculus and expressionrewriting framework. Parallel Computing, 25(9): 1105–1130, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/25/19/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/25/19/
article.pdf.

Huang:2013:PEV

[HL13] Zhiyi Huang and Kai-Cheung Leung. Performance evaluation of View-Oriented Transactional Memory. Parallel Computing, 39(12):787–801, December 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000975.

Hammond:1995:IPI

[HLDS95] Steven W. Hammond, Richard D. Loft, John M. Dennis, and Richard K. Sato. Implementation and performance issues of a massively parallel atmospheric model. Parallel Computing, 21(10):

1593–1619, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=10&aid=1017.

Hiromoto:1984:EDH

[HLM84] Robert E. Hiromoto, Olaf M. Lubeck, and James Moore. Experiences with the Denelcor HEP. Parallel Computing, 1(3–4):197– 206, December 1984. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hawick:2010:PGC

[HLP10] K. A. Hawick, A. Leist, and D. P. Playne. Parallel graph component labelling with GPUs and CUDA. Parallel Computing, 36(12):655–678, December 2010. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Han:2011:GPP

[HLvHA11] Liangxiu Han, Chee Sun Liew, Jano van Hemert, and Malcolm Atkinson. A generic parallel processing model for facilitating data mining and integration. Parallel Computing, 37(3):157–171, March 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000202.

Hoppe:1986:PAF

[HM86] H.-C. Hoppe and H. Mühlenbein.

Parallel adaptive fullmultigrid methods on message-based multiprocessors. Parallel Computing, 3(4):269–287, October 1986. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hart:1989:AMA

[HM89a] L. Hart and S. McCormick.

Asynchronous multilevel adaptive methods for solving partial differential equations on multiprocessors: basic ideas. Parallel Computing, 12(2):131–144, November 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Henderson:1989:SPA

[HM89b] M. E. Henderson and W. L. Miranker. Synergy in parallel algorithms. Parallel Computing, 11(1):17– 35, July 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Horiguchi:1989:PAF

[HM89c] Susumu Horiguchi and Willard L. Miranker. A parallel algorithm for finding the maximum value. Parallel Computing, 10(1): 101–108, March 1989. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Huang:1997:AGR

[HM97] Yih Huang and Philip K. McKinley. Adaptive global reduction algorithm for wormhole-routed 2D meshes. Parallel Computing, 23(13):1909–1936, December 15, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=13&aid=1255.

Hovland:2001:PSC

[HM01] Paul D. Hovland and Lois C. McInnes. Parallel simulation of compressible flow using automatic differentiation and PETSc. Parallel Computing, 27(4):503–519, March 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/28/31/
abstract.html;   http:

//www.elsevier.nl/gej-
ng/10/35/21/47/28/31/
article.pdf.

Hameurlain:2002:CIM

[HM02] Abdelkader Hameurlain and Franck Morvan. CPU and incremental memory allocation in dynamic parallelization of SQL queries. Parallel Computing, 28(4):525–556, April 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/39/27/
abstract.html.

Hussain:2013:SRA

[HMH^{+}13] Hameed Hussain, Saif Ur Rehman Malik, Abdul Hameed, Samee Ullah Khan, Gage Bickler, Nasro Min-Allah, Muhammad Bilal Qureshi, Limin Zhang, Wang Yongji, Nasir Ghani, Joanna Kolodziej, Albert Y. Zomaya, Cheng-Zhong Xu, Pavan Balaji, Abhinav Vishnu, Fredric Pinel, Johnatan E. Pecero, Dzmitry Kliazovich, Pascal Bouvry, Hongxiang Li, Lizhe Wang, Dan Chenm, and Ammar Rayes. A survey on resource allocation in high performance distributed computing systems. Parallel Computing, 39(11):709–736, November 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781911300121X.

Herbein:2016:PCI

[HMP^{+}16] S. Herbein, S. McDaniel, N. Podhorszki, J. Logan, S. Klasky, and M. Taufer. Performance characterization of irregular I/O at the extreme scale. Parallel Computing, 51(??):17– 36, January 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115001386.

Hill:1998:BBP

[HMS^{+}98] Jonathan M. D. Hill, Bill McColl, Dan C. Stefanescu, Mark W. Goudreau, Kevin Lang, Satish B. Rao, Torsten Suel, Thanasis Tsantilas, and Rob H. Bisseling. BSPlib: The BSP programming library. Parallel Computing, 24 (14):1947–1980, December 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/14/1370.pdf.

Herbordt:2007:SPS

[HMS^{+}07] Martin C. Herbordt, Josh Model, Bharat Sukhwani, Yongfeng Gu, and Tom VanCourt. Single pass streaming BLAST on FP-GAs. Parallel Computing, 33(10–11):741–756, November 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Holmes:2019:PPE

[HMS^{+}19] Daniel J. Holmes, Bradley Morgan, Anthony Skjellum, Purushotham V. Bangalore, and Srinivas Sridharan. Planning for performance: Enhancing achievable performance for MPI through persistent collective operations. Parallel Computing, 81(??):32–57, January 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118302412.

Hanson:1993:PDP

[HMTX93] F. B. Hanson, J.-D. Mei, C. Tier, and H. Xu. PDAC: a data parallel algorithm for the performance analysis of closed queueing networks. Parallel Computing, 19 (12):1345–1358, December 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hockney:1985:MCU

[Hoc85a] R. W. Hockney. MIMD computing in the U.S.A.— 1984. Parallel Computing, 2(2):119–136, June 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hockney:1985:MCC

[Hoc85b] Roger W. Hockney. (r_{∞}, n_{1/2}, s_{1/2}) measurements on the 2-CPU CRAY X-MP. Parallel Computing, 2(1):1–14, March 1985. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hockney:1987:PCP

[Hoc87] R. W. Hockney. Parametrization of computer performance. Parallel Computing, 5(1–2):97–103, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hockney:1989:SCO

[Hoc89] R. Hockney. Synchronization and communication overheads on the LCAP multiple FPS-164 computer system. Parallel Computing, 9(3):279–290, February 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hockney:1991:PPB

[Hoc91] R. Hockney. Performance parameters and benchmarking of supercomputers. Parallel Computing, 17(10–11):1111–1130, December 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hockney:1994:CCM

[Hoc94] Roger W. Hockney. The communication challenge for MPP: Intel Paragon and Meiko CS-2. Parallel Computing, 20(3):389–398, March 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=3&aid=830.

Hoefler:2012:ENG

[Hoe12] Torsten Hoefler. Extensions for next-generation parallel programming models. Parallel Computing, 38(1–2):1, January/ February 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001542.

Holmstrom:1995:PAE

[Hol95] Mats Holmström. Practical aspects and experiences: Parallelizing the fast wavelet transform. Parallel Computing, 21 (11):1837–1848, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=11&aid=1026.

Hollingsworth:2010:E

[Hol10] Jeffrey K. Hollingsworth. Editorial. Parallel Computing, 36(1):1–2, January 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hollingsworth:2011:MAC

[Hol11] Jeffrey K. Hollingsworth. In memoriam: Angela C. Sodan, PhD (August 30, 1955–April 21, 2011). Parallel Computing, 37(8): 327, August 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111000627.

Hollingsworth:2017:E

[Hol17] Jeff Hollingsworth. Editorial. Parallel Computing, 70(??):1, December 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117301631.

Horiguchi:1991:HSS

[Hor91]  S. Horiguchi.  Hybrid

systolic sorters. Parallel Computing, 17(9):997– 1007, November 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Horng:1993:CDC

[Hor93a] S. J. Horng. Computing dominators on a cubeconnected machine. Parallel Computing, 19(7):713– 728, July 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Horton:1993:MLD

[Hor93b]  G. Horton.  A multi-

level diffusion method for dynamic load balancing. Parallel Computing, 19(2): 209–218, February 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hossfeld:1988:VS

[Hos88]  F. Hossfeld.  Vector-

supercomputers. Parallel Computing, 7(3):373–385, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Honjou:1991:PTS

[HOSS91] N. Honjou, K. Ohtsuki, M. Sekiya, and F. Sasaki. A parallelization technique for the speedup of configuration interaction computing. Parallel Computing, 17(2–3):297–310, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hotop:1989:NKF

[Hot89] Hans-Jürgen Hotop. New Kalman filter algorithms based on orthogonal transformations for serial and vector computers. Parallel Computing, 12(2):233–247, November 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Huang:1991:PFA

[HP91] Y. Huang and Y. Paker. A parallel FFT algorithm for transputer networks. Parallel Computing, 17(8): 895–906, October 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Higham:1994:PAC

[HP94] Nicholas J. Higham and Pythagoras Papadimitriou. A parallel algorithm for computing the polar decomposition. Parallel Computing, 20(8): 1161–1173, August 10, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/

cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=8&aid=882.

Heddaya:1997:CCA

[HP97] Abdelsalam Heddaya and Kihong Park. Congestion control for asynchronous parallel computing on workstation networks. Parallel Computing, 23(13):1855–1875, December 15, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=13&aid=1249.

Hadjidoukas:2005:OEM

[HP05] P. E. Hadjidoukas and T. S. Papatheodorou. OpenMP extensions for master-slave message passing computing. Parallel Computing, 31(10– 12):1155–1167, October/ December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Huang:2004:HPP

[HR04] Chun-Hsi Huang and Sanguthevar Rajasekaran. High-performance parallel bio-computing. Parallel Computing, 30(9–10):999– 1000, September/October 2004. CODEN PACOEJ.

ISSN 0167-8191 (print), 1872-7336 (electronic).

Henon:2002:PHP

[HRR02] P. Hénon, P. Ramet, and J. Roman. PaStiX: a high-performance parallel direct solver for sparse symmetric positive definite systems. Parallel Computing, 28(2):301–321, February 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/33/35/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/60/33/35/
                                        main.pdf.

Henon:2008:FAS

[HRR08a] Pascal Hénon, Pierre Ramet, and Jean Roman. On finding approximate supernodes for an efficient block-ILU(k) factorization. Parallel Computing, 34(6–8):345–362, July 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hunold:2008:CBB

[HRR08b] S. Hunold, T. Rauber, and G. Rünger. Combining building blocks for parallel multi-level matrix multiplication. Parallel Computing, 34(6–8):411– 426, July 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hernandez:2007:PAE

[HRT07] V. Hernandez, J. E. Roman, and A. Tomas. Parallel Arnoldi eigensolvers with enhanced scalability via global communications rearrangement. Parallel Computing, 33(7–8):521– 540, August 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hack:1995:CDN

[HRW^{+}95] James J. Hack, James M.

Rosinski, David L. Williamson, Byron A. Boville, and John E. Truesdale. Computational design of the NCAR community climate model. Parallel Computing, 21(10): 1545–1569, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=10&aid=1015.

Holt:1989:PTA

[HS89] C. Holt and A. Stewart. A parallel thinning algorithm with fine grain subtasking. Parallel Computing, 10(3):329–334, May 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hajj:1990:MPS

[HS90] Ibrahim N. Hajj and Stig Skelboe. A multilevel parallel solver for block tridiagonal and banded linear systems. Parallel Computing, 15(1–3):21–45, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hoel:2003:DPP

[HS03] Erik G. Hoel and Hanan Samet. Data-parallel polygonization. Parallel Computing, 29(10):1381–1401, October 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hartley:2012:IPA

[HSÇ12] Timothy D. R. Hartley, Erik Saule, and Ümit V. Çatalyürek. Improving performance of adaptive component-based dataflow middleware. Parallel Computing, 38(6–7):289–309, June/July 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819112000221.

Horelik:2014:MCD

[HSFS14] Nicholas Horelik, Andrew Siegel, Benoit Forget, and Kord Smith. Monte Carlo domain decomposition for robust nuclear reactor analysis. Parallel Computing, 40(10):646–660, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001240.

Hsieh:2006:FTC

[Hsi06] Sun-Yuan Hsieh. Faulttolerant cycle embedding in the hypercube with more both faulty vertices and faulty edges. Parallel Computing, 32(1):84– 91, January 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hyatt:1989:PAB

[HSN89] R. M. Hyatt, B. W. Suter, and H. L. Nelson. A parallel alpha/beta tree searching algorithm. Parallel Computing, 10(3):299–308, May 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hallock:2014:SRD

[HSR^{+}14] Michael J. Hallock, John E.

Stone, Elijah Roberts, Corey Fry, and Zaida Luthey-Schulten. Simulation of reaction diffusion processes over biologically relevant size and time scales using multi-GPU workstations. Parallel Computing, 40(5–6): 86–99, May 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000398.

Hammonds:2007:SCD

[HSS07] James S. Hammonds, Faisal Saied, and Mark A. Shannon. Solving coupled 3-D paraxial wave and thermal diffusion equations with mixedmode parallel computations. Parallel Computing, 33(1):43–53, February 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hardy:2009:MSE

[HSS09] David J. Hardy, John E. Stone, and Klaus Schulten. Multilevel summation of electrostatic potentials using graphics processing units. Parallel Computing, 35(3):164–177, March 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hornung:2017:APA

[HSSM17] Maximilian Hornung, Svilen Stefanov, David Schneller, and Sharru Møller. Analysis of the ParConnect algorithm ran on Intel Xeon Phi Knights Landing. Parallel Computing, 70(??):46–53, December 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301308.

Hoshino:1987:MCP

[HST87] T. Hoshino, T. Shirakawa, and K. Tsuboi. Meshconnected parallel computer PAX for scientific applications. Parallel Computing, 5(3):363–371, November 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hirano:2005:ODB

[HST05] Motonori Hirano, Mitsuhisa Sato, and Yoshio Tanaka. OpenGR: a directive-based grid programming environment. Parallel Computing, 31 (10–12):1140–1154, October/December 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hong:1992:PPL

[HT92] Tzung-Pei Hong and Shian-Shyong Tseng. Parallel perceptron learning on a single-channel broadcast communication model. Parallel Computing, 18(2):133–148, February 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hong:1994:OPP

[HT94] Tzung-Pei Hong and Shian-Shyong Tseng. An optimal parallel perceptron learning algorithm for a large training set. Parallel Computing, 20(3):347–352, March 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=3&aid=850.

Houle:1998:DET

[HT98] Michael E. Houle and Gavin Turner. Dimensionexchange token distribution on the mesh and the torus. Parallel Computing, 24(2):247–265, February 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/2/1283.pdf.

Han:2000:ECR

[HT00] Hwansoo Han and Chau-Wen Tseng. Efficient compiler and run-time support for parallel irregular reductions. Parallel Computing, 26(13–14):1861–1887, December 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/34/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/34/30/
article.pdf.

Hribar:2001:IPS

[HTB01] Michelle R. Hribar, Valerie E. Taylor, and David E. Boyce. Implementing parallel shortest path for parallel transportation applications. Parallel Computing, 27 (12):1537–1568, November 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/42/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/42/29/
article.pdf.

Hurley:1993:TMU

[Hur93] S. Hurley. Taskgraph mapping using a genetic algorithm: a comparison of fitness functions. Parallel Computing, 19(11):1313– 1317, November 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Habata:2004:HSE

[HUYK04] Shinichi Habata, Kazuhiko Umezawa, Mitsuo Yokokawa, and Shigemune Kitawaki. Hardware system of the Earth Simulator. Parallel Computing, 30(12):1287– 1313, December 2004. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hamidouche:2016:CAO

[HVA^{+}16] Khaled Hamidouche, Akshay Venkatesh, Ammar Ahmad Awan, Hari Subramoni, Ching-Hsiang Chu, and Dhabaleswar K. Panda. CUDA-aware OpenSHMEM: Extensions and designs for high performance OpenSHMEM on GPU clusters. Parallel Computing, 58(??):27– 36, October 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300345.

Hijma:2012:GSS

[HvNJB12] Pieter Hijma, Rob V. van Nieuwpoort, Ceriel J. H. Jacobs, and Henri E. Bal. Generating synchronization statements in divideand-conquer programs. Parallel Computing, 38 (1–2):75–89, January/ February 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001384.

Heydorn:1991:OPA

[HW91] S. Heydorn and P. Weidner. Optimization and performance analysis of thinning algorithms on parallel computers. Parallel Computing, 17(1):17– 27, April 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Huang:1994:ARS

[HW94] Tien-Yu Huang and Jean-Lien C. Wu. Alternate resolution strategy in multistage interconnection networks. Parallel Computing, 20(6):887–896, June 10, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=6&aid=851.

He:2002:EPI

[HW02] Fusen He and Jie Wu. An efficient parallel implementation of the Everglades Landscape Fire Model using checkpointing. Parallel Computing, 28(1):65– 82, January 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/27/31/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/60/27/31/
                                        00001683.pdf.

Hwang:2002:PGC

[Hwa02] Yuan-Shin Hwang. Parallelizing graph construction operations in programs with cyclic graphs. Parallel Computing, 28 (9):1307–1328, September 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/59/33/
                                        abstract.html.

Hwang:2004:EAC

[Hwa04] Gwan-Hwan Hwang. An efficient algorithm for communication set generation of data parallel programs with block-cyclic distribution. Parallel Computing, 30(4):473–501, April 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hiromoto:1992:PAI

[HWB92] R. Hiromoto, B. R. Wienke, and R. G. Brickner. The performance of asynchronous iteration schemes applied to the linearized Boltzmann transport equation. Parallel Computing, 18(3):241–268, March 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Huang:2022:FCI

[HWL^{+}22] Yan Huang, Qingbin Wang, Minghao Lv, Xingguang Song, Jinkai Feng, Xuli Tan, Ziyan Huang, and Chuyuan Zhou. Fast calculation of isostatic compensation correction using the GPU-parallel prism method. Parallel Computing, 113(??):??, October 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000618.

Hsu:1992:PAF

[HWW92] Lih-Hsing Hsu, Peng Fei Wang, and Chu Tao Wu. Parallel algorithms for finding the most vital edge with respect to minimum spanning tree. Parallel Computing, 18(10):1143– 1155, October 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Huang:1989:GTC

[HWYL89] Tsung Chuan Huang, Jhing-Fa Wang, Chu Sing Yang, and Jau-Yien Lee. Graph theoretic characterization and reliability of the generalized Boolean ncube network. Parallel Computing, 12(3):375–385, December 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hao:2014:XFT

[HXCZ14] Zhijun Hao, Chenning Xie, Haibo Chen, and Binyu Zang. X10-FT: Transparent fault tolerance for AP-GAS language and runtime. Parallel Computing, 40(2):136–156, February 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001440.

Huang:2013:EMC

[HXW^{+}13] Libo Huang, Nong Xiao, Zhiying Wang, Yongwen Wang, and Mingche Lai. Efficient multimedia coprocessor with enhanced SIMD engines for exploiting ILP and DLP. Parallel Computing, 39(10):586–602, October 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819113000823.

Hoffmann:1993:PED

[HZ93] K.-H. Hoffmann and J. Zou.

Parallel efficiency of domain decomposition methods. Parallel Computing, 19(12):1375–1391, December 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

He:2019:SCC

[HZY^{+}19] Jiaao He, Chenggang Zhao, Jiping Yu, Xinjian Yu, Liyan Zheng, Chenyao Lou, Shizhi Tang, Wentao Han, and Jidong Zhai. Student Cluster Competition 2018, Team Tsinghua University: Reproducing performance of multiphysics simulations of the Tsunamigenic 2004 Sumatra megathrust earthquake on the Intel Skylake Architecture. Parallel Computing, 90(??):Article 102570, December 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301619.

Inda:2001:SEP

[IB01] Márcia A. Inda and Rob H.

Bisseling. A simple and efficient parallel FFT algorithm using the BSP model. Parallel Computing, 27(14):1847–1878, December 31, 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/44/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/44/28/
article.pdf.

Ibanez:2016:HMT

[IDS16] Dan Ibanez, Ian Dunn, and Mark S. Shephard. Hybrid MPI-thread parallelization of adaptive mesh operations. Parallel Computing, 52(??):133–143, February 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000041.

Iwashita:2012:LST

[IHM^{+}12] Takeshi Iwashita, Yu Hirotani, Takeshi Mifune, Toshio Murayama, and Hideki Ohtani. Large-scale time-harmonic electromagnetic field analysis using a multigrid solver on a distributed memory parallel computer. Parallel Computing, 38(9):485–500, September 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819112000464.

Ierotheou:1996:CAP

[IJCL96] C. S. Ierotheou, S. P.

Johnson, M. Cross, and P. F. Leggett. Computer aided parallelisation tools (CAPTools) — conceptual overview and performance on the parallelisation of structured mesh codes. Parallel Computing, 22(2):163–195, April 5, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=2&aid=1055.

Ierotheou:2005:GOC

[IJM^{+}05] C. S. Ierotheou, H. Jin, G. Matthews, S. P. Johnson, and R. Hood. Generating OpenMP code using an interactive parallelization environment. Parallel Computing, 31(10– 12):999–1012, October/ December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Iverson:2015:ECC

[IKK15] J. Iverson, C. Kamath, and G. Karypis. Evaluation of connected-component labeling algorithms for distributed-memory systems. Parallel Computing, 44(??):53–68, May 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000435.

Inoue:1988:AMV

[IM88] A. Inoue and A. Maeda. The architecture of a multi-vector processor system, VVP. Parallel Computing, 8(1–3):185–193, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Iserte:2018:DAI

[IMQO^{+}18] Sergio Iserte, Rafael Mayo, Enrique S. Quintana-Ortí, Vicenç Beltran, and Antonio J. Peña. DMR API: Improving cluster productivity by turning applications into malleable. Parallel Computing, 78(??):54– 66, October 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118302229.

Ino:2005:DDP

[IOH05] Fumihiko Ino, Kanrou Ooyama, and Kenichi Hagihara. A data distributed parallel algorithm for nonrigid image registration. Parallel Computing, 31(1): 19–43, January 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ibarra:1993:EPG

[IR93] Louis Ibarra and Dana Richards. Efficient parallel graph algorithms based on open ear decomposition. Parallel Computing, 19(8):873–886, August 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Iliopoulos:2000:OPA

[IR00] Costas S. Iliopoulos and James F. Reid. Optimal parallel analysis and decomposition of partially occluded strings. Parallel Computing, 26(4):483– 494, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/26/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/26/28/
article.pdf.

Imakura:2018:BSC

[IS18] Akira Imakura and Tetsuya Sakurai. Block SS-CAA: A complex momentbased parallel nonlinear eigensolver using the block communication-avoiding Arnoldi procedure. Parallel Computing, 74(??):34– 48, ???? 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117301886.

Isard:2003:DRI

[ISH03] M. Isard, M. Shand, and A. Heirich. Distributed rendering of interactive soft shadows. Parallel Computing, 29(3):311–323, March 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Iwamoto:2003:RMP

[ISO^{+}03] Yoshiyuki Iwamoto, Koichi Suga, Kanemitsu Ootsu, Takashi Yokota, and Takanobu Baba. Receiving message prediction method. Parallel Computing, 29(11– 12):1509–1538, November/ December 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Isgro:2008:DGA

[IT08] Francesco Isgrò and Domenico Tegolo. A distributed genetic algorithm for restoration of vertical line scratches. Parallel Computing, 34 (12):727–734, December 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Itoh:1987:TFL

[IU87] M. Itoh and K. Uchida.

Trends in Fujitsu large scale computer technology. Parallel Computing, 5(1–2):105–115, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Itakura:2004:SHP

[IUY^{+}04] K. Itakura, A. Uno, M. Yokokawa, T. Ishihara, and Y. Kaneda. Scalability of hybrid programming for a CFD code on the Earth Simulator. Parallel Computing, 30(12):1329–1343, December 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Iwabuchi:2022:MPM

[IYV^{+}22] Keita Iwabuchi, Karim Youssef, Kaushik Velusamy, Maya Gokhale, and Roger Pearce. Metall: a persistent memory allocator for data-centric analytics. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000114.

Islam:2020:MUH

[IZS^{+}20] Nusrat Sharmin Islam, Gengbin Zheng, Sayantan Sur, Akhil Langer, and Maria Garzaran. Minimizing the usage of hardware counters for collective communication using triggered operations. Parallel Computing, 98 (??):Article 102636, October 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300296.

Javadi:2006:PMA

[JAA06] Bahman Javadi, Mohammad K. Akbari, and Jemal H. Abawajy. A performance model for analysis of heterogeneous multicluster systems. Parallel Computing, 32(11–12): 831–851, December 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Judd:2018:PEP

[JAH^{+}18] Patrick Judd, Jorge Albericio, Tayler Hetherington, Tor Aamodt, Natalie Enright Jerger, Raquel Urtasun, and Andreas Moshovos. Proteus: Exploiting precision variability in deep neural networks. Parallel Computing, 74(??):40–51, ???? 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300741.

Janssen:1987:NSS

[Jan87] R. Janssen. A note on superlinear speedup. Parallel Computing, 4(2):211– 213, April 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jana:2006:PIP

[Jan06] Prasanta K. Jana. Polynomial interpolation and polynomial root finding on OTIS-mesh. Parallel Computing, 32(4):301–312, April 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jacq:2004:GBT

[JBC^{+}04] N. Jacq, C. Blanchet, C. Combet, E. Cornillot, L. Duret, K. Kurata, H. Nakamura, T. Silvestre, and V. Breton. Grid as a bioinformatic tool. Parallel Computing, 30(9–10):1093–1107, September/October 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jacq:2007:VSL

[JBC^{+}07] Nicolas Jacq, Vincent Breton, Hsin-Yen Chen, Li-Yung Ho, Martin Hofmann, Vinod Kasam, Hurng-Chun Lee, Yannick Legré, Simon C. Lin, Astrid Maaß, Emmanuel Medernach, Ivan Merelli, Luciano Milanesi, Giulio Rastelli, Matthieu Reichstadt, Jean Salzemann, Horst Schwichtenberg, Ying-Ta Wu, and Marc Zimmermann. Virtual screening on large scale grids. Parallel Computing, 33(4–5):289–301, May 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jann:2014:TIO

[JBWE14] Joefon Jann, R. Sarma Burugula, Ching-Farn E. Wu, and Kaoutar El Maghraoui. Towards an immortal operating system in virtual environments. Parallel Computing, 40(9):526–535, October 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000374.

Jain:1994:EPR

[JC94] Abhay Jain and N. S. Chaudhari. Efficient parallel recognition of context-free languages. Parallel Computing, 20 (9):1303–1321 (or 1303– 1322??), September 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=9&aid=888.

Jiang:2024:ELL

[JCC^{+}24] Qingcai Jiang, Zhenwei Cao, Xinhui Cui, Lingyun Wan, Xinming Qin, Huanqi Cao, Hong An, Junshi Chen, Jie Liu, Wei Hu, and Jinlong Yang. Extending the limit of LR-TDDFT on two different approaches: Numerical algorithms and new Sunway heterogeneous supercomputer. Parallel Computing, 120(??):??, June 2024. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819124000231.

Johnson:1996:ESI

[JCE96] S. P. Johnson, M. Cross, and M. G. Everett. Exploitation of symbolic information in interprocedural dependence analysis. Parallel Computing, 22(2):197–226, April 5, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=2&aid=1053.

Jeong:1992:PER

[JCL92] Chang-Sung Jeong, Jung-Ju Choi, and Der Tsai Lee. Parallel enclosing rectangle on SIMD machines. Parallel Computing, 18(2): 221–229, February 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jie:2007:SIS

[JCWP07] Wei Jie, Wentong Cai, Lizhe Wang, and Rob Procter. A secure information service for monitoring large scale grids. Parallel Computing, 33(7– 8):572–591, August 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ju:2002:PMC

[JDG02] Lili Ju, Qiang Du, and Max Gunzburger. Probabilistic methods for centroidal Voronoi tessellations and their parallel implementations. Parallel Computing, 28(10): 1477–1500, October ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/60/31/
abstract.html.

Johnson:1996:CDT

[JDH96] Theodore Johnson, Timothy A. Davis, and Steven M. Hadfield. A concurrent dynamic task graph. Parallel Computing, 22(2):327–333, April 5, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=2&aid=1040.

Jeong:1991:IPA

[Jeo91a] Chang-Sung Jeong. An improved parallel algorithm for constructing Voronoĭ diagram on a mesh-connected computer. Parallel Computing, 17 (4–5):505–514, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jeong:1991:PVD

[Jeo91b] Chang-Sung Jeong. Parallel Voronoĭ diagram in L_{1}(L_{∞}) metric on a meshconnected computer. Parallel Computing, 17(2–3): 241–252, June 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jesshope:1988:TSO

[Jes88] C. Jesshope. Transputers and switches as objects in OCCAM. Parallel Computing, 8(1–3):19–30, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jakobs:1993:SAP

[JG93] A. Jakobs and R. W. Gerling. Scaling aspects for the performance of parallel algorithms. Parallel Computing, 19(9):1063–1073, September 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jordan:2020:AFA

[JGT^{+}20] Herbert Jordan, Philipp Gschwandtner, Peter Thoman, Peter Zangerl, Alexander Hirsch, Thomas Fahringer, Thomas Heller, and Dietmar Fey. The allscale framework architecture. Parallel Computing, 99 (??):Article 102648, November 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819120300417.

Jackson:1997:SYE

[JH97] David J. Jackson and Chris W. Humphres. A simple yet effective load balancing extension to the PVM software system. Parallel Computing, 22(12):1647–1660, February 21, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=12&aid=1112.

Jiang:2022:OSC

[JHD^{+}22] Jiazhi Jiang, Dan Huang, Jiangsu Du, Yutong Lu, and Xiangke Liao. Optimizing small channel 3D convolution on GPU with tensor core. Parallel Computing, 113(??):??, October 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000473.

Jarvis:2002:PAI

[JHSV02] S. A. Jarvis, J. M. D.

Hill, C. J. Siniolakis, and V. P. Vasilev. Portable and architecture independent parallel performance tuning using BSP. Parallel Computing, 28 (11):1587–1609, November ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/61/30/
                                        abstract.html.

Jia:2009:PCM

[Jia09] Bin Jia. Process cooperation in multiple message broadcast. Parallel Computing, 35(12):572– 580, December 2009. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Johnson:1996:APC

[JIC96] S. P. Johnson, C. S. Ierotheou, and M. Cross. Automatic parallel code generation for message passing on distributed memory systems. Parallel Computing, 22(2):227–258, April 5, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=2&aid=1056.

Junqueira-Junior:2020:SCT

[JJAP^{+}20] Carlos Junqueira-Junior, João Luiz F. Azevedo, Jairo Panetta, William R. Wolf, and Sami Yamouni. On the scalability of CFD tool for supersonic jet flow configurations.

Parallel Computing, 93 (??):Article 102620, May 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300132.

Jin:2011:HPC

[JJM^{+}11] Haoqiang Jin, Dennis Jespersen, Piyush Mehrotra, Rupak Biswas, Lei Huang, and Barbara Chapman. High performance computing using MPI and OpenMP on multi-core parallel systems. Parallel Computing, 37(9):562–575, September 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000159.

Jeong:1991:FPS

[JK91] Chang-Sung Jeong and Myung-Ho Kim. Fast parallel simulated annealing for traveling salesman problem on SIMD machines with linear interconnections. Parallel Computing, 17(2–3):221–228, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Johnsson:1992:CTF

[JK92] S. L. Johnsson and R. L.

Krawitz. Cooley–Tukey FFT on the Connection Machine. Parallel Computing, 18(11):1201–1221, November 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jones:1995:PCD

[JKH95] Philip W. Jones, Christopher L. Kerr, and Richard S. Hemler. Practical considerations in development of a parallel SKYHI general circulation model. Parallel Computing, 21 (10):1677–1694, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=10&aid=1021.

Jeun:2008:OPB

[JKHK08] Woo-Chul Jeun, Yang-Suk Kee, Soonhoi Ha, and Changdon Kee. Overcoming performance bottlenecks in using OpenMP on SMP clusters. Parallel Computing, 34(10):570– 592, October 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jang:2022:PMV

[JLLC22] Seo Jin Jang, Wei Liu, Wei Li, and Yong Beom Cho. Parallel multiview HEVC for heterogeneously embedded cluster system. Parallel Computing, 112(??):??, September 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000448.

Jacquelin:2018:PDM

[JLY18] Mathias Jacquelin, Lin Lin, and Chao Yang. PSelInv — a distributed memory parallel algorithm for selected inversion: The non-symmetric case. Parallel Computing, 74(??): 84–98, ???? 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301941.

Joven:2013:IPM

[JMA^{+}13] J. Joven, A. Marongiu, F. Angiolini, L. Benini, and G. De Micheli. An integrated, programming model-driven framework for NoC–QoS support in cluster-based embedded many-cores. Parallel Computing, 39(10):549–566, October 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0167819113000720.

Jean-Marie:1998:AAP

[JMLBL98] Alain Jean-Marie, Sophie Lefebvre-Barbaroux, and Zhen Liu. An analytical approach to the performance evaluation of master–slave computational models. Parallel Computing, 24(5–6):841– 862, June 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/5-6/1290.pdf.

Joubert:2019:PAC

[JNC^{+}19] Wayne Joubert, James Nance, Sharlee Climer, Deborah Weighill, and Daniel Jacobson. Parallel accelerated Custom Correlation Coefficient calculations for genomics applications. Parallel Computing, 84(??):15–23, May 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118301431.

Joubert:2018:PAV

[JNWJ18] Wayne Joubert, James Nance, Deborah Weighill, and Daniel Jacobson. Parallel accelerated vector similarity calculations for genomics applications. Parallel Computing, 75(??):130–145, July 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911830084X.

Johnsson:1993:MCT

[Joh93] S. L. Johnsson. Minimizing the communication time for matrix multiplication on multiprocessors. Parallel Computing, 19(11):1235–1257, November 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jocksch:2021:OAC

[JOL^{+}21] Andreas Jocksch, Noé Ohana, Emmanuel Lanti, Eirini Koutsaniti, Vasileios Karakasis, and Laurent Villard. An optimisation of allreduce communication in message-passing systems. Parallel Computing, 107(??):??, October 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000612.

Jonker:2008:DBP

[JON08] P. P. Jonker, J. G. E. Olk, and C. Nicolescu. Distributed bucket processing: a paradigm embedded in a framework for the parallel processing of pixel sets. Parallel Computing, 34(12):735–746, December 2008. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jordan:1986:SPA

[Jor86] Harry F. Jordan. Structuring parallel algorithms in an MIMD, shared memory environment. Parallel Computing, 3(2):93– 110, May 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jordan:1988:PLC

[Jor88] H. F. Jordan. Programming language concepts for multiprocessors. Parallel Computing, 8(1–3):31–40, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jou:1989:LRB

[Jou89] I.-C. Chang Jou. Linear rotation based algorithm and systolic architecture for solving linear system equations. Parallel Computing, 11(3):367–379, ???? 28, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Joubert:1997:PIW

[Jou97] G. R. Joubert, editor. Proceedings of the 1996 International Workshop on Environments and Tools for Parallel Scientific Computing, volume 23(1–2) of Parallel Computing. North-Holland, Amsterdam, The Netherlands, April 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Joubert:2002:E

[Jou02] Gerhard R. Joubert. Editorial. Parallel Computing, 28(1):1–2, January 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/27/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/27/27/
00001675.pdf.

Joubert:2004:EN

[Jou04] Gerhard R. Joubert. Editorial note. Parallel Computing, 30(8):917–918, August 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jones:1994:SIS

[JP94] Mark T. Jones and Paul E.

Plassmann. Scalable iterative solution of sparse linear systems. Parallel Computing, 20(5):753–773, May 11, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=5&aid=844.

JavadiAbhari:2015:SSC

[JPK^{+}15] Ali JavadiAbhari, Shruti Patil, Daniel Kudrow, Jeff Heckey, Alexey Lvov, Frederic T. Chong, and Margaret Martonosi. ScaffCC: Scalable compilation and analysis of quantum programs. Parallel Computing, 45(??):2–17, June 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001422.

Jacobsen:2013:MLP

[JS13] Dana A. Jacobsen and Inanc Senocak. Multi-level parallelism for incompressible flow computations on GPU clusters. Parallel Computing, 39(1):1–20, January 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000804.

Jadav:1997:BDA

[JSC97] Divyesh Jadav, Chutimet Srinilta, and Alok Choudhary. Batching and dynamic allocation techniques for increasing the stream capacity of an on-demand media server. Parallel Computing, 23 (12):1727–1742, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=12&aid=1242.

Jevremovic:2011:PNA

[JTS^{+}11] Dimitrije Jevremović, Cong T.

Trinh, Friedrich Srienc, Carlos P. Sosa, and Daniel Boley. Parallelization of Nullspace Algorithm for the computation of metabolic pathways. Parallel Computing, 37(6–7):261–278, June/ July 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000378.

Jager:2009:EPH

[JW09] Gerold Jäger and Clemens Wagner. Efficient parallelizations of Hermite and Smith normal form algorithms. Parallel Computing, 35(6):345–357, June 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jiang:2013:ERA

[JWS13] Guiyuan Jiang, Jigang Wu, and Jizhou Sun. Efficient reconfiguration algorithms for communicationaware three-dimensional processor arrays. Parallel Computing, 39(9):490–503, September 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000495.

Kortas:1996:PPM

[KA96] Samuel Kortas and Philippe Angot. A practical and portable model of programming for iterative solvers on distributed memory machines. Parallel Computing, 22(4):487–512, June 11, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=4&aid=1068.

Koehn:2019:DSE

[KA19] Thaddeus Koehn and Peter Athanas. Data staging for efficient high throughput stream processing. Parallel Computing, 90 (??):Article 102566, December 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301577.

Kurzak:2009:OMM

[KAD09] Jakub Kurzak, Wesley Alvaro, and Jack Dongarra. Optimizing matrix multiplication for a shortvector SIMD architecture — CELL processor. Parallel Computing, 35(3):138– 150, March 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kak:1988:TLM

[Kak88]  S. C. Kak.  A two-

layered mesh array for matrix multiplication. Parallel Computing, 6(3):383– 385, March 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kalamboukis:1990:STE

[Kal90] T. Z. Kalamboukis. The symmetric tridiagonal eigenvalue problem on a transputer network. Parallel Computing, 15(1–3):101– 106, September 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kalamboukis:1992:PAD

[Kal92] T. Z. Kalamboukis. A parallel algorithm for the dense symmetric eigenvalue problem on a transputer array. Parallel Computing, 18(2):207–212, February 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kamowitz:1987:SME

[Kam87] D. Kamowitz. SOR and MGR(ν) experiments on the Crystal multicomputer. Parallel Computing, 4(2):117–142, April 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kammer:1988:SVF

[Kam88] H. Kammer. The SUPRENUM vector floating-point unit. Parallel Computing, 7(3): 315–323, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kumbhare:2020:DPM

[KAM^{+}20] Nirmal Kumbhare, Ali Akoglu, Aniruddha Marathe, Salim Hariri, and Ghaleb Abdulla. Dynamic power management for valueoriented schedulers in power-constrained HPC system. Parallel Computing, 99(??):Article 102686, November 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300764.

Kanada:1993:MVP

[Kan93] Y. Kanada. A method of vector processing for shared symbolic data. Parallel Computing, 19(10): 1155–1175, October 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kanawati:1997:LRD

[Kan97] Rushed Kanawati. LI-CRA: a replicated-data management algorithm for distributed synchronous group-ware applications. Parallel Computing, 22 (13):1733–1746, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=13&aid=1125.

Kao:2008:PIR

[Kao08] O. Kao. On parallel image retrieval with dynamically extracted features. Parallel Computing, 34(12): 700–709, December 2008.

CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kascic:1984:VDC

[Kas84] M. J. Kascic, Jr. Vorton dynamics: a case study of developing a fluid dynamics model for a vector processor. Parallel Computing, 1(1):35–44, August 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kasabov:1985:MSM

[Kas85] Nikola K. Kasabov. A method for SIMD/MIMD functionally reconfigurable multimicroprocessor systems design and parallel data exchange algorithms. Parallel Computing, 2(1): 73–78, March 1985. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Katona:1986:LMC

[Kat86] E. Katona. A lattice model for cellular (systolic) algorithms. Parallel Computing, 3(3):251–258, July 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Katsinis:2001:PAS

[Kat01] Constantine Katsinis. Performance analysis of the simultaneous optical multiprocessor exchange bus. Parallel Computing, 27(8):

1079–1115, July 2001. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/47/
34/28/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/47/
34/28/article.pdf.

Kato:2003:KPG

[Kat03] Toshi Kato. “Kilauea”— -parallel global illumination renderer. Parallel Computing, 29(3):289–310, March 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kauranne:1994:SGW

[Kau94] Tuomo Kauranne. Summary of GENESIS work at the European Centre for Medium-range Weather Forecasts (ECMWF). Parallel Computing, 20(10– 11):1685–1688, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=923.

Kalra:1985:PAT

[KB85] N. C. Kalra and P. C. P.

Bhatt. Parallel algorithms for tree traversals. Parallel Computing, 2(2):163– 171, June 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kindratenko:2008:RSS

[KB08] Volodymyr Kindratenko and Duncan Buell. Reconfigurable Systems Summer Institute 2007. Parallel Computing, 34(4–5):199– 200, May 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Knigge:2020:IEA

[KB20] Timon E. Knigge and Rob H. Bisseling. An improved exact algorithm and an NP-completeness proof for sparse matrix bipartitioning. Parallel Computing, 96(??): Article 102640, August 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819120300338.

Kashko:1988:PMR

[KBBC88] A. Kashko, H. Buxton, B. F. Buxton, and D. A. Castelow. Parallel matching and reconstruction algorithms in computer vision. Parallel Computing, 8(1–3):3–17, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kumar:1993:GPC

[KBD93] M. Kumar, Y. Baransky, and M. Denneau. The GF11 parallel computer. Parallel Computing, 19(12):1393–1412, December 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kielmann:2001:NPA

[KBG^{+}01] Thilo Kielmann, Henri E.

Bal, Sergei Gorlatch, Kees Verstoep, and Rutger F. H. Hofman. Network performance-aware collective communication for clustered wide-area systems. Parallel Computing, 27(11):1431–1456, October 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/41/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/41/30/
article.pdf.

Kremer:1988:ATT

[KBGZ88] U. Kremer, H.-J. Bast, M. Gerndt, and H. P. Zima. Advanced tools and techniques for automatic parallelization. Parallel Computing, 7(3):387–393, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kim:1991:CLL

[KC91] S. K. Kim and A. T. Chronopoulos. A class of Lanczos-like algorithms implemented on parallel computers. Parallel Computing, 17(6–7):763–778, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Koc:1993:SAI

[KC93a]  Çetin Kaya Koç and Pe-

ter Cappello. Systolic arrays for integer Chinese remaindering. Parallel Computing, 19(11):1303–1311, November 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kontoghiorghes:1993:PRD

[KC93b] E. J. Kontoghiorghes and M. R. B. Clarke. Parallel reorthogonalization of the QR decomposition after deleting columns (short communication). Parallel Computing, 19(6):703–707, June 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kose:1997:PEP

[KC97] Cemal Köse and Alan Chalmers. Profiling for efficient parallel volume visualisation. Parallel Computing, 23(7):943–952, July 14, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=7&aid=1194.

Kuchen:2006:E

[KC06] Herbert Kuchen and Murray Cole. Editorial. Parallel Computing, 32(7–8): 447–448, September 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kacsuk:1997:GDD

[KCD^{+}97] Péter Kacsuk, José C.

Cunha, Gábor Dózsa, João Lourenço, Tibor Fadgyas, and Tiago Antão. A graphical development and debugging environment for parallel programs. Parallel Computing, 22(13): 1747–1770, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=13&aid=1126.

Koehler:2008:PAC

[KCG08] Seth Koehler, John Curreri, and Alan D. George. Performance analysis challenges and framework for high-performance reconfigurable computing. Parallel Computing, 34(4–5): 217–230, May 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kontoghiorghes:1999:RLS

[KCN99] E. J. Kontoghiorghes, M. Clint, and H.-H. Naegeli. Recursive leastsquares using a hybrid Householder algorithm on massively parallel SIMD systems. Parallel Computing, 25(9):1147–1159, September 1, 1999. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/32/
25/21/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/32/
25/21/article.pdf.

Kandemir:1998:CTC

[KCRB98] M. Kandemir, A. Choudhary, J. Ramanujam, and R. Bordawekar. Compilation techniques for outof-core parallel computations. Parallel Computing, 24(3–4):597–628, May 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.

elsevier.com/cas/tree/
store/parco/sub/1998/
24/3-4/1304.pdf.

Kapenga:1988:PAT

[Kd88] J. A. Kapenga and E. de Doncker. A parallelization of adaptive task partitioning algorithms. Parallel Computing, 7(2):211– 225, June 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kechadi:1997:ASO

[KD97] M.-Tahar Kechadi and J.- Luc Dekeyser. Analysis and simulation of an outof-order execution model in vector multiprocessor systems. Parallel Computing, 23(13):1963–1986, December 15, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=13&aid=1253.

Krotkiewski:2010:PSS

[KD10] M. Krotkiewski and M. Dabrowski.

Parallel symmetric sparse matrix-vector product on scalar multi-core CPUs. Parallel Computing, 36(4): 181–198, April 2010. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Krotkiewski:2013:ESC

[KD13] Marcin Krotkiewski and Marcin Dabrowski. Efficient 3D stencil computations using CUDA. Parallel Computing, 39 (10):533–548, October 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911300094X.

Kolberg:2013:MMS

[KdBMA^{+}13] Wagner Kolberg, Pedro de B. Marcos, Julio C. S. Anjos, Alexandre K. S. Miyazaki, Claudio R. Geyer, and Luciana B. Arantes. MRSG — a MapReduce simulator over SimGrid. Parallel Computing, 39(4–5):233–244, April/May 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000215.

Kavas:2001:UMP

[KEEF01] Avi Kavas, David Er-El, and Dror G. Feitelson. Using multicast to pre-load jobs on the ParPar cluster. Parallel Computing, 27(3):315–327, February 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/22/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/22/28/
article.pdf.

Kentridge:1990:NNL

[Ken90] R. W. Kentridge. Neural networks for learning in the real world: representation, reinforcement and dynamics. Parallel Computing, 14(3):405–414, August 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kennedy:1999:HMC

[Ken99] A. D. Kennedy. The Hybrid Monte Carlo algorithm on parallel computers. Parallel Computing, 25(10–11):1311–1339, September 1, 1999. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/32/
26/24/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/32/
26/24/article.pdf.

Kanevsky:1995:ECP

[KF95] Arkady Kanevsky and Chao Feng. On the embedding of cycles in pancake graphs. Parallel Computing, 21(6):923–936, June 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=6&aid=962.

Karp:1987:IPJ

[KG87] Alan H. Karp and John Greenstadt. An improved parallel Jacobi method for diagonalizing a symmetric matrix. Parallel Computing, 5(3):281–294, November 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kohout:2003:HPC

[KG03] James Kohout and Alan D. George. A high-performance communication service for parallel computing on distributed DSP systems. Parallel Computing, 29(7): 851–878, July 2003. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Khelghatdoust:2021:SLL

[KG21] Mansour Khelghatdoust and Vincent Gramoli. A scalable and low latency probe-based scheduler for data analytics frameworks. Parallel Computing, 103(??):??, June 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000120.

Kahou:2008:PAB

[KGS08] Guy Antoine Atenekeng Kahou, Laura Grigori, and Masha Sosonkina. A partitioning algorithm for block-diagonal matrices with overlap. Parallel Computing, 34(6–8):332– 344, July 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kranzlmuller:1997:DME

[KGV97] D. Kranzlmüller, S. Grabner, and J. Volkert. Debugging with the MAD environment. Parallel Computing, 23(1–2):199–217, April 16, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=1-2&aid=1156.

Khalaf:1992:PAI

[KH92] B. M. S. Khalaf and D. Hutchinson. Parallel algorithms for initial value problems: parallel shooting. Parallel Computing, 18(6):661–673, June 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kao:1995:OAC

[KH95] Tzong Wann Kao and Shi Jinn Horng. Optimal algorithms for computing articulation points and some related problems on a circular-arc graph (short communication). Parallel Computing, 21(6):953–969, June 12, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=6&aid=986.

Kim:1997:FTW

[KH97] Seong-Pyo Kim and Taisook Han. Fault-tolerant wormhole routing in mesh with overlapped solid fault regions. Parallel Computing, 23(13):1937–1962, December 15, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=13&aid=1251.

Khan:2012:SHS

[Kha12] Minhaj Ahmad Khan. Scheduling for heterogeneous systems using constrained critical paths.

Parallel Computing, 38 (4–5):175–193, April/May 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000105.

Khan:1992:OHO

[KHC92] B. Khan, L. Hayes, and A. P. Cracknell. The optimisation of higher order resampling methods in a multiprocessor environment. Parallel Computing, 18(12):1335–1347, December 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kruzik:2020:CSF

[KHHC20] Jakub Kruzik, David Horak, Vaclav Hapla, and Martin Cermak. Comparison of selected FETI coarse space projector implementation strategies. Parallel Computing, 93 (??):Article 102608, May 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300016.

Komeiji:2001:PMD

[KHN01] Yuto Komeiji, Makoto Haraguchi, and Umpei Nagashima. Parallel molecular dynamics simulation of a protein. Parallel Computing, 27(8):977–987, July 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/47/34/23/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/34/23/
                                        article.pdf.

Kishimoto:2005:OCH

[KI05] Yoshinori Kishimoto and Shuichi Ichikawa. Optimizing the configuration of a heterogeneous cluster with multiprocessing and execution-time estimation. Parallel Computing, 31(7):691–710, July 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kiehl:1991:VIO

[Kie91] M. Kiehl. A vector implementation of an ODE code for multi-point-boundary-value problems. Parallel Computing, 17(2–3):347– 352, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kiehl:1994:PMS

[Kie94] M. Kiehl. Parallel multiple shooting for the solution of initial value problems. Parallel Computing, 20(3):275–295, March 10, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=3&aid=833.

Kim:1990:PAP

[Kim90] Sung Kwon Kim. Parallel algorithms for planar dominance counting. Parallel Computing, 15(1–3): 241–246, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kim:1997:RRP

[Kim97] Sung Kwon Kim. Rectangulating rectilinear polygons in parallel. Parallel Computing, 23(3):349–367, May 15, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=3&aid=1142.

Kim:1998:CTR

[Kim98] Sung Kwon Kim. Constanttime RMESH algorithms for the range minima and co-minima problems. Parallel Computing, 24 (5–6):965–977, June 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/5-6/1307.pdf.

Kwak:2007:TRI

[KJ07] Jong Wook Kwak and Chu Shik Jhon. Torus Ring: improving performance of interconnection network by modifying hierarchical ring. Parallel Computing, 33(1):2–20, February 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Koibuchi:2005:PSA

[KJA05] Michihiro Koibuchi, Akiya Jouraku, and Hideharu Amano. Path selection algorithm: the strategy for designing deterministic routing from alternative paths. Parallel Computing, 31(1):117–130, January 2005. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Khaldi:2015:PBR

[KJA15] Dounia Khaldi, Pierre Jouvelot, and Corinne Ancourt. Parallelizing with BDSC, a resourceconstrained scheduling algorithm for shared and distributed memory systems. Parallel Computing, 41(??):66–89, January 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001410.

Kohlhoff:1992:PES

[KK92] S. Kohlhoff and J. Krone.

Performance evaluation of SUPRENUM for the LIN-PACK benchmark (short communication). Parallel Computing, 18(2):231–238, February 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Krishnamurthy:1994:AMP

[KK94] E. V. Krishnamurthy and Vikram Krishnamurthy. An ANN model perceptron algorithm using generalized matrix inversion. Parallel Computing, 20(5):799–806, May 11, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=5&aid=835.

Katagiri:2001:EIP

[KK01] Takahiro Katagiri and Yasumasa Kanada. An efficient implementation of parallel eigenvalue computation for massively parallel processing. Parallel Computing, 27(14): 1831–1845, December 31, 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/44/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/44/27/
article.pdf.

Kang:2003:TDB

[KK03] Oh-Han Kang and Si-Gwan Kim. A task duplication based scheduling algorithm for shared memory multiprocessors. Parallel Computing, 29(1):161– 166, January 2003. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Karlsson:2011:PTS

[KK11] L. Karlsson and B. Kågström. Parallel two-stage reduction to Hessenberg form using dynamic scheduling on shared-memory architectures. Parallel Computing, 37(12):771–782, December 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000482.

Kumar:1992:PAE

[KKB92] P. Sreenivasa Kumar, M. Kishore Kumar, and A. Basu. A parallel algorithm for elimination tree computation and symbolic factorization. Parallel Computing, 18(8):849– 856, August 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kumar:1993:PAS

[KKB93] P. Sreenivasa Kumar, M. K. Kumar, and A. Basu. Parallel algorithms for sparse triangular system solution. Parallel Computing, 19(2):187–196, February 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kee:2004:MMM

[KKH04] Yang-Suk Kee, Jin-Soo Kim, and Soonhoi Ha. Memory management for multi-threaded software DSM systems. Parallel Computing, 30(1):121–138, January 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Katagiri:2006:APE

[KKHY06a] Takahiro Katagiri, Kenji Kise, Hiroki Honda, and Toshitsugu Yuba. AB-CLib DRSSED: a parallel eigensolver with an autotuning facility. Parallel Computing, 32(3):231–250, March 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Katagiri:2006:ADS

[KKHY06b] Takahiro Katagiri, Kenji Kise, Hiroki Honda, and Toshitsugu Yuba. AB-CLibscript: a directive to support specification of an auto-tuning facility for numerical software. Parallel Computing, 32(1):92–112, January 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Koutsovasilis:2018:AEA

[KKK^{+}18] Panos Koutsovasilis, Christos Kalogirou, Christos Konstantas, Manolis Maroudas, Michalis Spyrou, and Christos D. Antonopoulos. AcHEe: Evaluating approximate computing and heterogeneity for energy efficiency. Parallel Computing, 74(??):52–67, ???? 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300285.

Krishnamurthy:1990:SAT

[KKSS90] E. V. Krishnamurthy, M. Kunde, M. Schimmler, and H. Schröder. Systolic algorithm for tensor products of matrices: implementation and applications. Parallel Computing, 13(3):301–308, March 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kalinowski:2000:LSG

[KKT00] Tomasz Kalinowski, Iskander Kort, and Denis Trystram. List scheduling of general task graphs under LogP. Parallel Computing, 26(9):1109–1128, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/30/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/30/26/
article.pdf.

Karlsson:2016:PAT

[KKU16] Lars Karlsson, Daniel Kressner, and André Uschmajew. Parallel algorithms for tensor completion in the CP format. Parallel Computing, 57(??):222–234, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0167819115001210.

Karlsson:2014:FGB

[KKW14] L. Karlsson, B. Kågström, and E. Wadbro. Finegrained bulge-chasing kernels for strongly scalable parallel QR algorithms. Parallel Computing, 40(7):271–288, July 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000465.

Kohout:2005:PDT

[KKŽ05] Josef Kohout, Ivana Kolingerová, and Jiří Žára. Parallel Delaunay triangulation in E^{2} and E^{3} for computers with shared memory. Parallel Computing, 31(5):491–522, May 2005. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kim:1990:PAS

[KL90a] Hyoung Joong Kim and Jang Gyu Lee. A parallel algorithm solving a tridiagonal Toeplitz linear system. Parallel Computing, 13(3):289–294, March 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kindermann:1990:INN

[KL90b] J. Kindermann and A. Linden. Inversion of neural networks by gradient descent. Parallel Computing, 14(3):277–286, August 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Keh:2000:FTE

[KL00] Huan-Chao Keh and Jen-Chih Lin. On faulttolerant embedding of Hamiltonian cycles, linear arrays and rings in a Flexible Hypercube. Parallel Computing, 26(6):769– 781, May 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/28/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/28/28/
article.pdf.

Lee:1995:PAE

[kLH95] Chi kin Lee and Mounir Hamdi. Practical aspects and experiences: Parallel image processing applications on a network of workstations. Parallel Computing, 21(1):137–160, January 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/

                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=1&aid=934.

Kueng:2009:EPV

[KLL^{+}09] Tz-Liang Kueng, Cheng-Kuan Lin, Tyne Liang, Jimmy J. M. Tan, and Lih-Hsing Hsu. Embedding paths of variable lengths into hypercubes with conditional link-faults. Parallel Computing, 35(8–9):441–454, August/September 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Knecht:1990:PQD

[KLN90] S. Knecht, E. Laermann, and W. E. Nagel. Parallelizing QCD with dynamical fermions on a Cray multiprocessor system. Parallel Computing, 15(1–3):3–20, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kerbyson:2011:AWF

[KLP11] Darren J. Kerbyson, Michael Lang, and Scott Pakin. Adapting wave-front algorithms to efficiently utilize systems with deep communication hierarchies. Parallel Computing, 37 (9):550–561, September 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000226.

Kalinov:2005:HC

[KLR05] Alexey Kalinov, Alexey Lastovetsky, and Yves Robert. Heterogeneous computing. Parallel Computing, 31(7):649–652, July 2005. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kunde:1988:ISA

[KLS^{+}88] Manfred Kunde, Hans-Werner Lang, Manfred Schimmler, Hartmut Schmeck, and Heiko Schröder. The instruction systolic array and its relation to other models of parallel computers. Parallel Computing, 7 (1):25–39, April 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kolp:1988:PES

[KM88] O. Kolp and H. Mierendorff. Performance estimations for SUPRENUM systems. Parallel Computing, 7(3):357–366, September 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kramer:1989:MSM

[KM89] O. Krämer and H. Mühlenbein. Mapping strategies in message-based multiprocessor systems. Parallel Computing, 9(2):213–225, January 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). PARLE: Conference on Parallel Architectures and Languages— Europe (Eindhoven, 1987).

Kalantzis:2018:SID

[KMB^{+}18] Vassilis Kalantzis, A. Cristiano I. Malossi, Costas Bekas, Alessandro Curioni, Efstratios Gallopoulos, and Yousef Saad. A scalable iterative dense linear system solver for multiple right-hand sides in data analytics. Parallel Computing, 74(??):136– 153, ???? 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117302119.

Kus:2019:OEE

[KMK^{+}19] P. Kůs, A. Marek, S. S. Köcher, H.-H. Kowalski, C. Carbogno, Ch. Scheurer, K. Reuter, M. Scheffler, and H. Lederer. Optimizations of the eigensolvers in the ELPA library. Parallel Computing, 85(??):167–177, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911830334X.

Karavanic:1997:IVP

[KMLM97] Karen L. Karavanic, Jussi Myllymaki, Miron Livny, and Barton P. Miller. Integrated visualization of parallel program performance data. Parallel Computing, 23(1–2):181–198, April 16, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=1-2&aid=1155.

Kampe:1988:PCC

[KN88] F. C. Kampe and T. M.

Nguyen. Performance comparison of the Cray-2 and Cray X-MP on a class of seismic data processing algorithms. Parallel Computing, 7(1):41–53, April 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Konda:2009:NAS

[KN09] Taro Konda and Yoshimasa Nakamura. A new algorithm for singular value decomposition and its parallelization. Parallel Computing, 35(6):331–344, June 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kacsuk:1997:TML

[KNP97] Péter Kacsuk, Zsolt Németh, and Zsolt Puskás. Tools for mapping, load balancing and monitoring in the LOGFLOW parallel Prolog project. Parallel Computing, 22(13): 1853–1881, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=13&aid=1132.

Kontoghiorghes:2000:PCE

[KNR00] Erricos John Kontoghiorghes, Anna Nagurney, and Berç Rustem. Parallel computing in economics, finance and decision-making. Parallel Computing, 26(5): 507–509, March 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/42/
27/23/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/42/
27/23/article.pdf.

Kumar:2008:IMC

[KNTG08] V. Santhosh Kumar, R. Nanjundiah, M. J. Thazhuthaveetil, and R. Govindarajan. Impact of message compression on the scalability of an atmospheric modeling application on clusters. Parallel Computing, 34(1):1– 16, January 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Koc:1997:PAM

[Koç97]  Ç. K. Koç. Parallel p-

adic method for solving linear systems of equations. Parallel Computing, 23(13):2067–2074, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=13&aid=1220.

Kogge:1985:FBC

[Kog85] P. M. Kogge. Functionbased computing and parallelism: a review. Parallel Computing, 2(3):243–253, November 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kohring:1995:DLB

[Koh95] G. A. Kohring. Dynamic load balancing for parallelized particle simulations on MIMD computers. Parallel Computing, 21(4):683–693, April 1, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=4&aid=978.

Kolp:1994:PEP

[Kol94] Otto Kolp. Performance estimation for a parallel system with a hierarchical switch network. Parallel Computing, 20(10– 11):1613–1626, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=917.

Kontoghiorghes:2002:GGA

[Kon02] Erricos John Kontoghiorghes. Greedy Givens algorithms for computing the rank-k updating of the QR decomposition. Parallel Computing, 28(9):1257–1273, September 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-

ng/10/35/21/60/59/30/
abstract.html.

Konrad:2011:TCD

[Kon11] Christian Konrad. Twoconstraint domain decomposition with Space Filling Curves. Parallel Computing, 37(4–5):203–216, April/May 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000251.

Kotsis:1997:SAW

[Kot97] Gabriele Kotsis. A systematic approach for workload modeling for parallel processing systems. Parallel Computing, 22 (13):1771–1787, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=13&aid=1127.

Krishnamurthy:1993:DPE

[KP93] E. V. Krishnamurthy and Chen Pin. Data parallel evaluation-interpolation algorithm for polynomial matrix inversion. Parallel Computing, 19(5):577– 589, May 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Khardon:1996:PSC

[KP96] Roni Khardon and Shlomit S. Pinter. Partitioning and scheduling to counteract overhead. Parallel Computing, 22(4):555–593, June 11, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=4&aid=1051.

Kindratenko:2012:AAH

[KP12] Volodymyr Kindratenko and Gregory D. Peterson. Application accelerators in HPC — editorial introduction. Parallel Computing, 38(8):343, August 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000324.

Krogh:1997:PSR

[KPH97] Michael Krogh, James Painter, and Charles Hansen. Parallel sphere rendering. Parallel Computing, 23(7):961–974, July 14, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.

elsevier.com/cas/tree/
store/parco/free/1997/
23/7/krogh/fig2.jpg.

Klockner:2012:PPS

[KPL^{+}12] Andreas Klöckner, Nicolas Pinto, Yunsup Lee, Bryan Catanzaro, Paul Ivanov, and Ahmed Fasih. Py-CUDA and PyOpenCL: a scripting-based approach to GPU run-time code generation. Parallel Computing, 38(3):157–174, March 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001281.

Krechel:1990:PVA

[KPS90] Arno Krechel, Hans-Joachim Plum, and Klaus Stüben. Parallelization and vectorization aspects of the solution of tridiagonal linear systems. Parallel Computing, 14(1):31–49, May 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kadalbajoo:1997:PGE

[KR97] Mohan K. Kadalbajoo and A. Appaji Rao. Parallel group explicit method for two-dimensional parabolic equations. Parallel Computing, 23(6):649–666, June 20, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=6&aid=1137.

Komasa:2000:SQM

[KR00] Jacek Komasa and Jacek Rychlewski. Solving quantum-mechanical problems on parallel systems. Parallel Computing, 26(7–8):999–1009, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/42/29/33/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/29/33/
                                        article.pdf.

Karbowski:2010:ACB

[KR10] Andrzej Karbowski and Maciej Remiszewski. Assessment of the Cell Broadband Engine architecture as a platform to solve closed-loop optimal control problems. Parallel Computing, 36(4):169–180, April 2010. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kratz:1984:VFE

[Kra84]  M. Kratz.  Vectorized

finite-element stiffness generation: tuning the Noor-Lambiotte algorithm. Parallel Computing, 1(2):121– 132, December 1984. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Krentel:2013:LTF

[Kre13] Mark W. Krentel. Libmonitor: a tool for firstparty monitoring. Parallel Computing, 39(3):114– 119, March 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000798.

Khan:2011:EDC

[KRN^{+}11] M. M. Khan, A. D. Rast, J. Navaridas, X. Jin, L. A. Plana, M. Luján, S. Temple, C. Patterson, D. Richards, J. V. Woods, J. Miguel-Alonso, and S. B. Furber. Eventdriven configuration of a neural network CMP system over an homogeneous interconnect fabric. Parallel Computing, 37(8):392– 409, August 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000160.

Kumar:2021:ODC

[KRS^{+}21] Bipin Kumar, Matt Rehme, Neethi Suresh, Nihanth Cherukuru, Stanislaw Jaroszynski, Samual Li, Scott Pearse, Tim Scheitlin, Suryachandra A. Rao, and Ravi S. Nanjundiah. Optimization of DNS code and visualization of entrainment and mixing phenomena at cloud edges. Parallel Computing, 107(??): ??, October 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000600.

Kruyt:1997:CGM

[Kru97] N. P. Kruyt. A conjugate gradient method for the spectral partitioning of graphs. Parallel Computing, 22(11):1493–1502, January 26, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=11&aid=1110.

Katz:1988:MSC

[KRW88] S. Katz, W. A. Ray, and G. Walder. Multiprocessor software for the CYBER-PLUS high performance system. Parallel Computing, 8(1–3):231–244, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kaps:1987:SPE

[KS87] Michael Kaps and Michael Schlegl. A short proof for the existence of the W Z-factorisation. Parallel Computing, 4(2):229–232, April 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). See comments [HE88].

Kamath:1989:PMS

[KS89] Chandrika Kamath and Ahmed Sameh. A projection method for solving nonsymmetric linear systems on multiprocessors. Parallel Computing, 9(3): 291–312, February 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kockler:1991:PSV

[KS91a] N. Kockler and M. Simon. Parallel singular value decomposition with cyclic storing. Parallel Computing, 17(1):39– 47, April 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Krishnamurthy:1991:SAM

[KS91b] E. V. Krishnamurthy and H. Schröder. Systolic algorithm for multivariable approximation using tensor products of basis functions. Parallel Computing, 17(4–5):483–492, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kumar:1995:DPG

[KS95] K. G. Kumar and D. B. Skillicorn. Data parallel geometric operations on lists. Parallel Computing, 21(3):447–459, March 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=3&aid=970.

Krizanc:1999:BSP

[KS99] Danny Krizanc and Anton Saarimaki. Bulk synchronous parallel: practical experience with a model for parallel computing. Parallel Computing, 25(2):159–181, February 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/2/1382.pdf.

Krechel:2001:PAM

[KS01] Arnold Krechel and Klaus Stüben. Parallel algebraic multigrid based on subdomain blocking. Parallel Computing, 27(8):1009– 1031, July 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/34/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/34/25/
article.pdf.

Kechadi:2005:DTS

[KS05] M-Tahar Kechadi and Ilias K. Savvas. Dynamic task scheduling for irregular network topologies. Parallel Computing, 31(7): 757–776, July 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kulla:2007:SPS

[KS07] Fabian Kulla and Peter Sanders. Scalable parallel suffix array construction. Parallel Computing, 33(9): 605–612, September 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kohler:2016:GAC

[KS16] Martin Köhler and Jens Saak. On GPU acceleration of common solvers for (quasi-) triangular generalized Lyapunov equations. Parallel Computing, 57(??):212–221, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300436.

Kleis:1994:ERE

[KSM^{+}94] U. Kleis, J. M. Singer, I. Morgenstern, Th. Hußlein, and H.-G. Matuttis. Experiences with re-engineering and parallelizing a high-T_{c} superconductivity code. Parallel Computing, 20 (3):399–407, March 10, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=3&aid=848.

Kim:1997:LDS

[KSP97] Jong-Uk Kim, Kyu-Hyun Shim, and Kyu Ho Park. A link-disjoint subcube for processor allocation in hypercube computers. Parallel Computing, 22 (12):1579–1595, February 21, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=12&aid=1108.

Korosec:2004:SMPa

[KŠR04a] Peter Korošec, Jurij Šilc, and Borut Robič. Solving the mesh-partitioning problem with an antcolony algorithm. Parallel Computing, 30(5–6): 785–801, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). See [KŠR04b].

Korosec:2004:SMPb

[KŠR04b] Peter Korošec, Jurij Šilc, and Borut Robič. “Solving the mesh-partitioning problem with an antcolony algorithm” [Parallel Computing 30 (2004) 785–801]. Parallel Computing, 30(8):919–921, August 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). See [KŠR04a].

Konchady:1998:IPE

[KSS98] Manu Konchady, Arun Sood, and Paul S. Schopf. Implementation and performance evaluation of a parallel ocean model. Parallel Computing, 24 (2):181–203, February 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/2/1250.pdf.

Kumar:2006:PAA

[KSS06] P. Rajesh Kumar, K. Sridharan, and S. Srinivasan. A parallel algorithm, architecture and FPGA realization for landmark determination and map construction in a planar unknown environment. Parallel Computing, 32(3): 205–221, March 2006. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kontoghiorghes:2002:SIP

[KST02] Erricos John Kontoghiorghes, Ahmed Sameh, and Denis Trystram. Special issue on parallel matrix algorithms and applications. Parallel Computing, 28(2):151–153, February 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/33/27/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/60/33/27/
                                        main.pdf.

Kalro:1997:PCU

[KT97] V. Kalro and T. Tezduyar. Parallel 3D computation of unsteady flows around circular cylinders. Parallel Computing, 23 (9):1235–1248, November 3, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=9&aid=1208.

Kessler:1999:LLS

[KT99] Christoph W. Keßler and Jesper Larsson Träff. Language and library support for practical PRAM programming. Parallel Computing, 25(2):105–135, February 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/2/1369.pdf.

Keppens:2000:UHP

[KT00] R. Keppens and G. Tóth.

Using high performance Fortran for magnetohydrodynamic simulations. Parallel Computing, 26(6): 705–722, May 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/42/
28/24/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/42/
28/24/article.pdf.

Kang:2019:SAM

[KTAB^{+}19] Qiao Kang, Jesper Larsson Träff, Reda Al-Bahrani, Ankit Agrawal, Alok Choudhary, and Wei keng Liao. Scalable algorithms for MPI intergroup Allgather and Allgatherv. Parallel Computing, 85(??):220–230, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911830320X.

Kodama:2014:SRM

[KTN^{+}14] Chihiro Kodama, Masaaki Terai, Akira T. Noda, Yohei Yamada, Masaki Satoh, Tatsuya Seiki, Shin ichi Iga, Hisashi Yashiro, Hirofumi Tomita, and Kazuo Minami. Scalable rank-mapping algorithm for an icosahedral grid system on the massive parallel computer with a 3-D torus network. Parallel Computing, 40(8):362– 373, August 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000659.

Krommer:1993:AAA

[KU93] A. R. Krommer and C. W. Ueberhuber. Architecture adaptive algorithms. Parallel Computing, 19(4): 409–435, April 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kutti:1985:TPP

[Kut85] S. Kutti. Taxonomy of parallel processing and definitions. Parallel Computing, 2(4):353–359, December 1985. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kutil:2002:AZI

[Kut02] Rade Kutil. Approaches to zerotree image and video coding on MIMD architectures. Parallel Computing, 28(7–8):1095–1109, August 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/58/35/
abstract.html.

Kuznetsov:1998:ORD

[Kuz98] S. V. Kuznetsov. Orthogonal reduction of dense matrices to bidiagonal form on computers with distributed memory architectures. Parallel Computing, 24(2):305–313, February 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/2/1285.pdf.

Kim:1999:INO

[KV99] Sunil Kim and Alexander V. Veidenbaum. Interconnection network organization and its impact on performance and cost in shared memory multiprocessors. Parallel Computing, 25(3):283–309, March 22, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/3/1388.pdf.

Kutrib:1997:PIW

[KVW97a] M. Kutrib, R. Vollmar, and T. Worsch, editors. Proceedings of the 1996 IFIP Workshop on Cellular Automata, volume 23(11) of Parallel Computing. North-Holland, Amsterdam, The Netherlands, November 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kutrib:1997:ISI

[KVW97b] M. Kutrib, R. Vollmar, and Th. Worsch. Introduction to the special issue on cellular automata. Parallel Computing, 23 (11):1567–1576, December 1, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/

cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=11&aid=1231.

Kamath:1990:ITP

[KW90] C. Kamath and S. Weeratunga. Implementation of two projection methods on a shared memory multiprocessor: DEC VAX 6240. Parallel Computing, 16(2–3):375–382, December 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kim:1994:TPS

[KY94] Dongseung Kim and Byung-Guoen Yi. A two-pass scheduling algorithm for parallel programs. Parallel Computing, 20(6):869–885, June 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=6&aid=849.

Kasahara:1998:DLC

[KY98] Hironori Kasahara and Akimasa Yoshida. A data-localization compilation scheme using partialstatic task assignment for Fortran coarse-grain parallel processing. Parallel Computing, 24(3–4):579– 596, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/3-4/1303.pdf.

Kumar:2001:PUI

[KYLH01] B. V. Rathish Kumar, T. Yamaguchi, H. Liu, and R. Himeno. A parallel 3D unsteady incompressible flow solver on VPP700. Parallel Computing, 27(13):1687–1713, December 1, 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/47/43/29/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/43/29/
                                        article.pdf.

Kabeer:2023:DSD

[KYS23] Muhammad Kabeer, Ibrahim Yusuf, and Nasir Ahmad Sufi. Distributed software defined networkbased fog to fog collaboration scheme. Parallel Computing, 117(??):??, September 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819123000467.

Koning:1997:VVS

[KZV97] Anton H. J. Koning, Karel J. Zuiderveld, and Max A. Viergever. Volume visualization on shared memory architectures. Parallel Computing, 23(7): 915–925, July 14, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cgi-bin/cas/tree/
store/parco/cas_sub/
browse/browse.cgi?year=
1997&volume=23&issue=
7&aid=1192.

Laforenza:2002:GPS

[Laf02] D. Laforenza. Grid programming: some indications where we are headed. Parallel Computing, 28 (12):1733–1752, December 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lahlou:2000:AAS

[Lah00] Chams Lahlou. Approximation algorithms for scheduling with a limited number of communications. Parallel Computing, 26(9):1129–1162, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/30/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/30/27/
article.pdf.

Langguth:2014:PPR

[LAHM14] J. Langguth, A. Azad, M. Halappanavar, and F. Manne. On parallel push-relabel based algorithms for bipartite maximum matching. Parallel Computing, 40(7):289–308, July 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000349.

Lange:1990:SHN

[Lan90] T. E. Lange. Simulation of heterogeneous neural networks on serial and parallel machines. Parallel Computing, 14(3):287–303, August 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lang:1996:PRB

[Lan96a]  Bruno Lang.  Paral-

lel reduction of banded matrices to bidiagonal form. Parallel Computing, 22(1):1–18, February 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=1&aid=1043.

Lanteri:1996:PSC

[Lan96b] S. Lanteri. Parallel solutions of compressible flows using overlapping and nonoverlapping mesh partitioning strategies. Parallel Computing, 22(7):943–968, October 1, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=7&aid=1087.

Lang:1999:EES

[Lan99]  Bruno Lang.  Efficient

eigenvalue and singular value computations on shared memory machines. Parallel Computing, 25 (7):845–860, August 13, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1999&volume=
25&issue=7&aid=1413;
http://www.elsevier.
nl/gej-ng/10/35/21/32/
23/18/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/32/
23/18/article.pdf.

Leach:1990:ACP

[LAS90] Ronald J. Leach, O. Michael Atogi, and Razeyah R.

Stephen. The actual complexity of parallel evaluation of low degree polynomials. Parallel Computing, 13(1):73–83, January 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lastovetsky:2002:APC

[Las02] Alexey Lastovetsky. Adaptive parallel computing on heterogeneous networks with mpC. Parallel Computing, 28(10):1369–1407, October ??, 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/60/28/
                                        abstract.html.

Laursen:1993:SAP

[Lau93]  P. S. Laursen.  Sim-

ple approaches to parallel Branch and Bound. Parallel Computing, 19(2):143– 152, February 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lane:2023:HSM

[LB23] Phillip Allen Lane and Joshua Dennis Booth. Heterogeneous sparse matrix– vector multiplication via compressed sparse row format. Parallel Computing, 115(??):??, February 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000030.

Lockhart:2023:CPN

[LBGO23] Shelby Lockhart, Amanda Bienz, William D. Gropp, and Luke N. Olson. Characterizing the performance of node-aware strategies for irregular point-to-point communication on heterogeneous architectures. Parallel Computing, 116(??):??, July 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000273.

Lorenz:2007:DTD

[LBH07] Mario Lorenz, Guido Brunnett, and Marcel Heinz. Driving tiled displays with an extended Chromium system based on stream cached multicast communication. Parallel Computing, 33(6):438–466, June 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lakamsani:1995:PAE

[LBL95] Vamsee Lakamsani, Laxmi N.

Bhuyan, and D. Scott Linthicum. Practical aspects and experiences. mapping molecular dynamics computations on to hypercubes. Parallel Computing, 21(6):993–1013, June 12, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=6&aid=985.

Li:1990:SPM

[LBWR90] J. Li, A. Brass, D. J. Ward, and B. Robson. A study of parallel molecular dynamics algorithms for N-body simulations on a transputer system. Parallel Computing, 14(2):211– 222, June 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lin:1989:PRT

[LC89] Ferng-Ching Lin and R. Charng. Pin reduction through variable duplications and substitutions in a data dependence graph. Parallel Computing, 10(2):231–238, April 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lin:1990:COP

[LC90] Ferng-Ching Lin and Kuo Liang Chung. A cost-optimal parallel tridiagonal system solver. Parallel Computing, 15(1–3):189–199, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lou:1997:PTL

[LC97] Der-Chyuan Lou and Chin-Chen Chang. A parallel two-list algorithm for the knapsack problem. Parallel Computing, 22(14):1985–1996, March 24, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=14&aid=1136.

Lin:1999:MCC

[LC99a] W.-Y. Lin and C.-L. Chen.

Minimum communication cost reordering for parallel sparse Cholesky factorization. Parallel Computing, 25(8):943–967, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/24/18/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/24/18/
article.pdf.

Liu:1999:ORL

[LC99b] Zhiyong Liu and David W.

Cheung. Oblivious routing for LC permutations on hypercubes. Parallel Computing, 25(4):445–460, April 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/4/1400.pdf.

Lee:2002:GCS

[LC02] PeiZong Lee and Wen-Yao Chen. Generating communication sets of array assignment statements for block-cyclic distribution on distributed memory parallel computers. Parallel Computing, 28(9):1329–1368, September 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/59/34/
abstract.html.

Leidel:2017:HSC

[LC17] John D. Leidel and Yong Chen. HMC-Sim-2.0: a co-design infrastructure for exploring custom memory cube operations. Parallel Computing, 68(??):77–88, October 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300984.

Levine:1991:CSA

[LCD91] D. Levine, D. Callahan, and J. Dongarra. A comparative study of automatic vectorizing compilers. Parallel Computing, 17(10–11):1223–1244, December 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lake:2018:SCC

[LCE^{+}18] Janaan Lake, Qixiang Chao, Hannah Eyre, Emerson Ford, Kevin Parker, and Kincaid Savoie. Student Cluster Competition 2017, Team University of Utah: Reproducing vectorization of the Tersoff multibody potential on the Intel Broadwell and Intel Skylake platforms. Parallel Computing, 79(??):1–8, November 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911830070X.

Liang:2019:CCA

[LCLA19] Weihao Liang, Yong Chen, Jialin Liu, and Hong An. CARS: a contentionaware scheduler for efficient resource management of HPC storage systems. Parallel Computing, 87(??):25–34, September 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781911830382X.

Liu:2000:NAF

[LCLL00] J. G. Liu, F. H. Y. Chan, F. K. Lam, and H. F. Li. A new approach to fast calculation of moments of 3-D gray level images. Parallel Computing, 26(6):805– 815, May 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/28/31/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/28/31/
                                        article.pdf.

Lee:1998:EMM

[LD98] Kangwoo Lee and Michel Dubois. Empirical models of miss rates. Parallel Computing, 24(2):205–219, February 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/2/1259.pdf.

Lakshmivarahan:1999:RTH

[LD99] S. Lakshmivarahan and Sudarshan K. Dhall. Ring, torus and hypercube architectures/algorithms for parallel computing. Parallel Computing, 25(13– 14):1877–1906, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/36/37/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/37/
article.pdf.

Li:2016:PSA

[LDK16] Lu Li, Usman Dastgeer, and Christoph Kessler. Pruning strategies in adaptive off-line tuning for optimized composition of components on heterogeneous systems. Parallel Computing, 51(??):37–45, January 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001179.

Levin:1991:IMD

[LE91] M. D. Levin and D. J.

Evans. The inversion of matrices by the doublebordering algorithm on MIMD computers. Parallel Computing, 17(4–5):591– 602, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lecot:1989:AGL

[Lec89] C. Lecot. An algorithm for generating low discrepancy sequences on vector computers. Parallel Computing, 11(1):113–116, July 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lee:1995:TCP

[Lee95] PeiZong Z. Lee. Techniques for compiling programs on distributed memory multicomputers. Parallel Computing, 21(12): 1895–1923, December 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=12&aid=1031.

Lee:1997:EIP

[Lee97] Tong-Yee Lee. Exploitation of image parallelism for ray tracing 3D scenes on 2D mesh multicomputers. Parallel Computing, 23(13):1993–2015, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.

cgi?year=1997&volume=
23&issue=13&aid=1257.

Lefevre:1997:PPT

[Lef97] Laurent Lefèvre. Parallel programming on top of DSM system. an experimental study. Parallel Computing, 23(1– 2):235–249, April 16, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=1-2&aid=1158.

Liu:2014:GTA

[LEH14] Yu Liu, Kento Emoto, and Zhenjiang Hu. A Generate-Test-Aggregate parallel programming library for systematic parallel programming. Parallel Computing, 40(2):116–135, February 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001403.

Lepage:1999:IDL

[Lep99] G. Peter Lepage. Improved discretizations for lattice QCD. Parallel Computing, 25(10– 11):1383–1393, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/32/26/28/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/26/28/
                                        article.pdf.

Leszczynski:2000:CC

[Les00] Jerzy Leszczynski. Computational chemistry. Parallel Computing, 26(7–8): 817–818, July 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/42/
                                        29/23/abstract.html;
                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/42/
                                        29/23/article.pdf.

Leuze:1989:ISO

[Leu89] Michael R. Leuze. Independent set orderings for parallel matrix factorization by Gaussian elimination. Parallel Computing, 10(2):177–191, April 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Leung:1993:LFI

[Leu93] Y.-W. Leung. On-line fault identification in multistage interconnection networks. Parallel Computing, 19(6): 693–702, June 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Levin:1990:FVQ

[Lev90] S. A. Levin. A fully vectorized quicksort. Parallel Computing, 16(2–3): 369–373, December 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lubeck:1988:MPH

[LF88] O. M. Lubeck and V. Faber.

Modeling the performance of hypercubes: a case study using the particlein-cell application. Parallel Computing, 9(1):37–52, December 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Li:1989:PCA

[LF89] Xiaobo Li and Zhi Xi Fang. Parallel clustering algorithms. Parallel Computing, 11(3):275–290, ???? 28, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lefebvre:1998:ASM

[LF98] Vincent Lefebvre and Paul Feautrier. Automatic storage management for parallel programs. Parallel Computing, 24(3–4):649– 671, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/3-4/1306.pdf.

Lowenthal:2000:AIP

[LF00] David K. Lowenthal and Vincent W. Freeh. Architectureindependent parallelism for both shared- and distributed-memory machines using the Filaments package. Parallel Computing, 26(10):1297–1323, August 15, 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/31/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/31/26/
article.pdf.

Lim:2011:ATC

[LFL11] Min Yeol Lim, Vincent W. Freeh, and David K. Lowenthal. Adaptive, transparent CPU scaling algorithms leveraging inter-node MPI communication regions. Parallel Computing, 37(10– 11):667–683, October/ November 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000871.

Li:2014:HAC

[LFS14] Yong Li, Dan Feng, and Zhan Shi. Heterogeneousaware cache partitioning: Improving the fairness of shared storage cache. Parallel Computing, 40(10):710–721, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000945.

Levy:2019:USE

[LFS^{+}19] Scott Levy, Kurt B. Ferreira, Whit Schonbein, Ryan E. Grant, and Matthew G. F. Dosanjh. Using simulation to examine the effect of MPI message matching costs on application performance. Parallel Computing, 84(??):63–74, May 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303272.

Letherwood:2001:GVM

[LG01] Michael D. Letherwood and David D. Gunter. Ground vehicle modeling and simulation of military vehicles using high performance computing. Parallel Computing, 27(1–2):109–140, January 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/47/25/27/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/25/27/
                                        article.pdf.

Li:2003:VCE

[LG03] Hon F. Li and Gabriel Girard. View consistencies and exact implementations. Parallel Computing, 29(1):37–67, January 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ltaief:2009:PAA

[LG09] Hatem Ltaief and Marc Garbey. A parallel Aitkenadditive Schwarz waveform relaxation suitable for the grid. Parallel Computing, 35(7):416–428, July 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lawson:2021:PPT

[LG21] John Lawson and Mehdi Goli. Performance portability through machine learning guided kernel selection in SYCL libraries. Parallel Computing, 107(??):??, October 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000624.

Labarta:1997:ASP

[LGC97] Jesus Labarta, Sergi Girona, and Toni Cortes. Analyzing scheduling policies using Dimemas. Parallel Computing, 23(1–2):23–34, April 16, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=1-2&aid=1145.

Lima:2015:DAS

[LGD^{+}15] João V. F. Lima, Thierry Gautier, Vincent Danjean, Bruno Raffin, and Nicolas Maillard. Design and analysis of scheduling strategies for multi-CPU and multi-GPU architectures. Parallel Computing, 44(??):37–52, May 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000459.

Lingrand:2009:MLP

[LGM09] Diane Lingrand, Tristan Glatard, and Johan Montagnat. Modeling the latency on production grids with respect to the execution context. Parallel Computing, 35 (10–11):493–511, October/ November 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lopez-Gomez:2019:ESP

[LGMdRA^{+}19] Javier López-Gómez, Javier Fernández Muñoz, David del Rio Astorga, Manuel F. Dolz, and J. Daniel Garcia. Exploring stream parallel patterns in distributed MPI environments. Parallel Computing, 84(??):24–36, May 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118303442.

Lehmann:1988:MDR

[LH88] L. Lehmann and F. Hopfl.

A model of distributed recovery for the SUPRENUM multiprocessor. Parallel Computing, 7(3):395–401, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Loh:2000:JCN

[LH00] Peter K. K. Loh and Wen Jing Hsu. The Josephus cube: a novel interconnection network. Parallel Computing, 26(4):427– 453, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/26/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/26/25/
article.pdf.

Lin:2002:RBT

[LH02] Jen-Chih Lin and Nan-Chen Hsien. Reconfiguring binary tree structures in a faulty supercube with unbounded expansion. Parallel Computing, 28(3):471– 483, March 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/34/32/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/34/32/
main.pdf.

Lee:2004:TCG

[LH04a] Chain-Wu Lee and Chun-Hsi Huang. Toward cooperative genomic knowledge inference. Parallel Computing, 30(9–10): 1127–1135, September/ October 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Loh:2004:FTR

[LH04b] Peter K. K. Loh and W. J. Hsu. Fault-tolerant routing for complete Josephus Cubes. Parallel Computing, 30(9–10):1151–1167, September/October 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Li:2023:MLP

[LHG^{+}23] Kuan Li, Kang He, Stef Graillat, Hao Jiang, Tongxiang Gu, and Jie Liu. Multi-level parallel multi-layer block reproducible summation algorithm. Parallel Computing, 115(??):??, February 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819123000029.

Lee:1996:PCE

[LHK^{+}96] Yu-Hua Lee, Shi-Jinn Horng, Tzong-Wann Kao, Ferng-Shi Jaung, Yuung-Jih Chen, and Horng-Ren Tsai. Parallel computation of exact Euclidean distance transform. Parallel Computing, 22(2):311–325, April 5, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/

cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=2&aid=1045.

Lu:2015:PHS

[LHK15] Hao Lu, Mahantesh Halappanavar, and Ananth Kalyanaraman. Parallel heuristics for scalable community detection. Parallel Computing, 47(??):19– 37, August 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000472.

Liu:2018:SCC

[LHL18] Siyuan Liu, Meiru Hao, and Bu-Sung Lee. Student Cluster Competition 2017, team Nanyang Technological University: Reproducing vectorization of the Tersoff multibody potential on the Intel Broadwell architecture. Parallel Computing, 77(??):118–124, September 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300607.

Fevre:2019:CPR

[LHR^{+}19] Valentin Le Fèvre, Thomas Herault, Yves Robert, Aurelien Bouteiller, Atsushi Hori, George Bosilca, and Jack Dongarra. Comparing the performance of rigid, moldable and gridshaped applications on failure-prone HPC platforms. Parallel Computing, 85(??):1–12, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118302230.

Lam:2013:DFP

[LHS13] Michael O. Lam, Jeffrey K. Hollingsworth, and G. W. Stewart. Dynamic floating-point cancellation detection. Parallel Computing, 39(3):146– 155, March 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000622.

Li:1997:PIO

[LHZ97] Wei Li, Xiaohu Huang, and Nanning Zheng. Parallel implementing OpenGL on PVM. Parallel Computing, 23(12):1839–1850, December 15, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.

cgi?year=1997&volume=
23&issue=12&aid=1248.

Liang:2020:AMD

[LHZ^{+}20] Jianguo Liang, Rong Hua, Hao Zhang, Wenqiang Zhu, and You Fu. Accelerated molecular dynamics simulation of silicon crystals on TaihuLight using OpenACC. Parallel Computing, 99(??): Article 102667, November 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300600.

Liang:2022:OAC

[LHZ^{+}22] Jianguo Liang, Rong Hua, Wenqiang Zhu, Yuxi Ye, You Fu, and Hao Zhang. OpenACC + Athread collaborative optimization of silicon-crystal application on Sunway Taihu-Light. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000023.

Li:1989:PIR

[Li89] Tao Li. Parallel implementation of rule-based expert systems for interactive applications. Parallel Computing, 10(3):309– 318, May 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Li:1990:SCF

[Li90] Liwu Li. Systolic computation with fault diagnosis. Parallel Computing, 14(2):235–243, June 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Li:1991:NNC

[Li91] Xiaobo Li. Nearest neighbor classification on two types of SIMD machines. Parallel Computing, 17 (4–5):381–407, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Li:1994:MCC

[Li94]  Jian-Jin Li.  Multi-

scattering on the Cube-Connected Cycles. Parallel Computing, 20(3):313–324, March 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=3&aid=839.

Li:1997:BVG

[Li97] Guangye Li. A block variant of the GMRES method on massively parallel processors. Parallel Computing, 23(8):1005–1019, July 25, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=8&aid=1162.

Lilja:1994:MAC

[Lil94] David J. Lilja. A multiprocessor architecture combining fine-grained and coarse-grained parallelism strategies. Parallel Computing, 20(5):729–751, May 11, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=5&aid=854.

Lin:1990:PGP

[Lin90]  Chau-Jy Lin.  Parallel

generation of permutations on systolic arrays. Parallel Computing, 15(1–3): 267–276, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lin:1991:ASA

[Lin91a]  Yen-Chun Lin.  Array

size anomaly of problemsize independent systolic arrays for matrix-vector multiplication. Parallel Computing, 17(4–5):515– 522, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lin:1991:FBT

[Lin91b] Yen-Chun Lin. An FPbased tool for the synthesis of regular array algorithms. Parallel Computing, 17(4–5):457–470, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lin:1993:PTD

[Lin93]  Yi-Bing Lin.  Paral-

lel trace-driven simulation for packet loss in finitebuffered voice multiplexers. Parallel Computing, 19(2):219–228, February 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lin:1994:NSA

[Lin94]  Yen Chun Lin.  New

systolic arrays for the longest common subsequence problem^{+}. Parallel Computing, 20(9): 1323–1334, September 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.

cgi?year=1994&volume=
20&issue=9&aid=878.

Lin:2001:UGM

[Lin01] H. X. Lin. A unifying graph model for designing parallel algorithms for tridiagonal systems. Parallel Computing, 27(7):925– 939, June 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/31/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/31/28/
article.pdf.

Lin:2008:PMA

[Lin08] Wei-Ming Lin. Performance modeling and analysis of correlated parallel computations. Parallel Computing, 34(9):521–538, September 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lippert:1999:PSP

[Lip99a]  Th. Lippert.  Paral-

lel SSOR preconditioning for lattice QCD. Parallel Computing, 25(10– 11):1357–1370, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/26/26/

                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/26/26/
                                        article.pdf.

Lippert:1999:HSA

[Lip99b] Thomas Lippert. Hypersystolic algorithms for Nbody computations and parallel level-3 BLAS libraries. Parallel Computing, 25(7):877–891, August 13, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1999&volume=
                                        25&issue=7&aid=1415;
                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/32/
                                        23/24/abstract.html;
                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/32/
                                        23/24/article.pdf.

Liu:1986:CMT

[Liu86] J. W. H. Liu. Computational models and task scheduling for parallel sparse Cholesky factorization. Parallel Computing, 3 (4):327–342, October 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Liu:1989:RSM

[Liu89] Joseph W. H. Liu. Reordering sparse matrices for parallel elimination.

Parallel Computing, 11(1): 73–91, July 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Liu:1998:WCA

[Liu98] Zhen Liu. Worst-case analysis of scheduling heuristics of parallel systems. Parallel Computing, 24 (5–6):863–891, June 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/5-6/1314.pdf.

Lakshmivarahan:1993:SIN

[LJD93] S. Lakshmivarahan, Jung Sing Jwo, and S. K. Dhall. Symmetry in interconnection networks based on Cayley graphs of permutation groups: a survey. Parallel Computing, 19(4):361– 407, April 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Liu:2000:ESR

[LK00] Yung-Lin Liu and Chung-Ta King. EXPLORER: Supporting run-time parallelization of DOACROSS loops on general networks of workstations. Parallel Computing, 26 (2–3):355–375, February 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/42/25/31/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/25/31/
                                        article.pdf.

Luisier:2010:NST

[LK10] Mathieu Luisier and Gerhard Klimeck. Numerical strategies towards petascale simulations of nanoelectronics devices. Parallel Computing, 36(2–3): 117–128, February/March 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

LaSalle:2014:MBD

[LK14] Dominique LaSalle and George Karypis. MPI for big data: New tricks for an old dog. Parallel Computing, 40(10):754–767, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000830.

Leon:2016:POI

[LKGD16] Edgar A. León, Ian Karlin, Ryan E. Grant, and Matthew Dosanjh. Program optimizations: the interplay between power, performance, and energy. Parallel Computing, 58(??):56–75, October 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300369.

Lee:2003:TSU

[LKHL03] Heejo Lee, Jong Kim, Sung Je Hong, and Sunggu Lee. Task scheduling using a block dependency DAG for block-oriented sparse Cholesky factorization. Parallel Computing, 29(1):135–159, January 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Liu:2024:GBH

[LKP24] Xi Liu, Gizem Kayar, and Ken Perlin. A GPUbased hydrodynamic simulator with boid interactions. Parallel Computing, 119(??):??, February 2024. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000686.

Legendi:1988:MM

[LKTZ88] T. Legendi, E. Katona, J. Toth, and A. Zsoter. Megacell machine. Parallel Computing, 8(1–3):195– 199, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lai:1988:FEU

[LL88] C. H. Lai and H. M. Liddell. Finite elements using long vectors of the DAP. Parallel Computing, 8(1– 3):351–361, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lin:1990:PSC

[LL90] Yen Chun Lin and Ferng-Ching Lin. Parallel sorting with cooperating heaps in a linear array of processors. Parallel Computing, 16(2–3):273–278, December 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lim:1998:MPM

[LL98] Amy W. Lim and Monica S. Lam. Maximizing parallelism and minimizing synchronization with affine partitions. Parallel Computing, 24(3–4):445– 475, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/3-4/1298.pdf.

Li:2019:HPI

[LLD19] Shengguo Li, Jie Liu, and Yunfei Du. A high performance implementation of Zolo-SVD algorithm on distributed memory systems. Parallel Computing, 86(??):57–65, August 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301807.

Liu:2023:LCA

[LLFZ23] Zhexu Liu, Shaofeng Liu, Zhiyong Fan, and Zhen Zhao. Low consumption automatic discovery protocol for DDS-based largescale distributed parallel computing. Parallel Computing, 118(??):??, November 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000583.

Lin:2010:TAC

[LLL10] Yi-Neng Lin, Ying-Dar Lin, and Yuan-Cheng Lai. Thread allocation in CMPbased multithreaded network processors. Parallel Computing, 36(2–3): 104–116, February/March 2010. CODEN PACOEJ.

ISSN 0167-8191 (print), 1872-7336 (electronic).

Laszloffy:2000:SDM

[LLP00] Andras Laszloffy, Jingping Long, and Abani K. Patra. Simple data management, scheduling and solution strategies for managing the irregularities in parallel adaptive hp finite element simulations. Parallel Computing, 26 (13–14):1765–1788, December 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/34/26/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/34/26/
                                        article.pdf.

Lagana:2006:SSO

[LLPV06] Demetrio Laganá, Pasquale Legato, Ornella Pisacane, and Francesca Vocaturo. Solving simulation optimization problems on grid computing systems. Parallel Computing, 32(9):688– 700, October 2006. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Li:1993:VPS

[LLS^{+}93] Xiaobo Li, Paul Lu, Jonathan Schaeffer, John Shillington, Pok Sze Wong, and Hanmao Shi. On the versatility of parallel sorting by regular sampling. Parallel Computing, 19(10):1079–1103, October 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lambert:2021:OOFa

[LLVM21a] Jacob Lambert, Seyong Lee, Jeffrey S. Vetter, and Allen D. Malony. Optimization with the OpenACC-to-FPGA framework on the Arria 10 and Stratix 10 FP-GAs. Parallel Computing, 104–105(??):??, July 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000417.

Lambert:2021:OOFb

[LLVM21b] Jacob Lambert, Seyong Lee, Jeffrey S. Vetter, and Allen D. Malony. Optimization with the OpenACC-to-FPGA framework on the arria 10 and stratix 10 FP-GAs. Parallel Computing, 104–105(??):??, July 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000417.

Li:2015:COP

[LLW^{+}15] Kenli Li, Jing Liu, Lanjun Wan, Shu Yin, and Keqin Li. A cost-optimal parallel algorithm for the 0– 1 knapsack problem and its performance on multicore CPU and GPU implementations. Parallel Computing, 43(??):27–42, March 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000113.

Li:2024:POA

[LLW^{+}24] Jianjiang Li, Lin Li, Qingwei Wang, Wei Xue, Jiabi Liang, and Jinliang Shi. Parallel optimization and application of unstructured sparse triangular solver on new generation of Sunway architecture. Parallel Computing, 120(??): ??, June 2024. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819124000188.

Lau:2018:SCC

[LLX^{+}18] Ka Cheong Jason Lau, Yuxuan Li, Lei Xie, Qian Xie, Beichen Li, Yu Chen, Guanyu Feng, Jiping Yu, Xinjian Yu, Miao Wang, Wentao Han, and Jidong Zhai. Student cluster competition 2017, team Tsinghua University: Reproducing vectorization of the Tersoff multi-body potential on the Intel Skylake and NVIDIA Volta architectures. Parallel Computing, 78(??):47–53, October 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300668.

Li:2000:PCQ

[LM00] Jenny X. Li and Gary L.

Mullen. Parallel computing of a quasi-Monte Carlo algorithm for valuing derivatives. Parallel Computing, 26(5):641– 653, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/27/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/27/30/
article.pdf.

Laporte:2003:PCL

[LM03] Gilbert Laporte and Roberto Musmanno. Parallel computing in logistics. Parallel Computing, 29(5):553– 554, May 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Linderoth:2006:OGO

[LM06] Jeff Linderoth and Roberto Musmanno. Optimization on grids — optimization for grids. Parallel Computing, 32(9):627–628, October 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Liang:2005:IPV

[LMC05] Kevin Liang, Patricia Monger, and Huge Couchman. Interactive parallel visualization of large particle datasets. Parallel Computing, 31(2):243–260, February 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Li:2009:DPI

[LMG09] Yaohang Li, Michael Mascagni, and Andrey Gorin. A decentralized parallel implementation for parallel tempering algorithm. Parallel Computing, 35(5):269–283, May 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lucka:2008:AMS

[LMH08] Maria Lucka, Igor Melichercik, and Ladislav Halada. Application of multistage stochastic programs solved in parallel in portfolio management. Parallel Computing, 34(6–8):469– 485, July 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Leggett:1996:IUK

[LMJC96] P. F. Leggett, A. T. J.

Marsh, S. P. Johnson, and M. Cross. Integrating user knowledge with information from parallelisation tools to facilitate the automatic generation of efficient parallel FORTRAN code. Parallel Computing, 22(2):259–288, April 5, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=2&aid=1054.

Luthi:1997:PBD

[LMKH97] J. Lüthi, S. Majumdar, G. Kotsis, and G. Haring. Performance bounds for distributed systems with workload variabilities and uncertainties. Parallel Computing, 22 (13):1789–1806, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.

                                        cgi?year=1997&volume=
                                        22&issue=13&aid=1128.

Louter-Nool:1987:BLA

[LN87] M. Louter-Nool. Basic linear algebra subprograms (BLAS) on the CDC Cyber 205. Parallel Computing, 4(2):143–165, April 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Luna:2006:OUG

[LNA06] F. Luna, A. J. Nebro, and E. Alba. Observations in using Grid-enabled technologies for solving multi-objective optimization problems. Parallel Computing, 32(5–6):377– 393, June 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Liu:2019:MMT

[LND^{+}19] Zhuo Liu, Amit Kumar Nath, Xiaoning Ding, Huansong Fu, Md. Muhib Khan, and Weikuan Yu. Multivariate modeling and two-level scheduling of analytic queries. Parallel Computing, 85(??):66–78, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118303557.

Lee:2013:SRN

[LNLK13] Junghee Lee, Chrysostomos Nicopoulos, Hyung Gyu Lee, and Jongman Kim. Sharded Router: a novel on-chip router architecture employing bandwidth sharding and stealing. Parallel Computing, 39(9):372–388, September 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000483.

Lanthier:2003:PIG

[LNS03] Mark Lanthier, Doron Nussbaum, and Jörg-Rüdiger Sack. Parallel implementation of geometric shortest path algorithms. Parallel Computing, 29(10):1445–1479, October 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lin:1992:FCO

[LO92] R. Lin and S. Olariu. A fast cost-optimal parallel algorithm for the lowest common ancestor problem. Parallel Computing, 18(5): 511–516, May 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Loh:1996:AIS

[Loh96] Peter K. K. Loh. Artificial intelligence search techniques as fault-tolerant routing strategies. Parallel Computing, 22(8): 1127–1147, October 28, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=8&aid=1084.

Loucif:1999:AFA

[LOKM99] S. Loucif, M. Ould-Khaoua, and L. M. Mackenzie. Analysis of fully adaptive wormhole routing in tori. Parallel Computing, 25 (12):1477–1487, November 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/35/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/35/26/
article.pdf.

Lopez:1993:MBB

[Lop93]  L. Lopez.  Methods

based on boundary value techniques for solving parabolic equations on parallel computers. Parallel Computing, 19(9):979– 991, September 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Llorens:2016:DSC

[LOST16] M. Llorens, J. Oliver, J. Silva, and S. Tamarit. Dynamic slicing of concurrent specification languages. Parallel Computing, 53(??):1–22, April 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000363.

Lopez:1992:PMN

[LP92] L. Lopez and T. Politi.

Parallel methods in the numerical treatment of population dynamic models. Parallel Computing, 18(7): 767–777, July 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lopez:1997:UFP

[LPAZ97] J. Lopez, O. Plata, F. Arguello, and E. L. Zapata. Unified framework for the parallelization of divide and conquer based tridiagonal systems. Parallel Computing, 23(6):667–686, June 20, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=6&aid=1180.

Laurent:1998:PIR

[LPCA98] C. Laurent, F. Peyrin, J-M Chassery, and M. Amiel. Parallel image reconstruction on MIMD computers for three- dimensional cone-beam tomography. Parallel Computing, 24(9– 10):1461–1479, September 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/9-10/1343.pdf.

Li:2000:SGT

[LPH00] Keqin Li, Yi Pan, and Mounir Hamdi. Solving graph theory problems using reconfigurable pipelined optical buses. Parallel Computing, 26(6): 723–735, May 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/42/
28/25/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/42/
28/25/article.pdf.

Langguth:2011:PAB

[LPM11] Johannes Langguth, Md. Mostofa Ali Patwary, and Fredrik Manne. Parallel algorithms for bipartite matching problems on distributed memory computers. Parallel Computing, 37(12):820–845, December 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001311.

Llorente:2001:PMS

[LPMD01] Ignacio M. Llorente, Manuel Prieto-Matías, and Boris Diskin. A parallel multigrid solver for 3D convection and convectiondiffusion problems. Parallel Computing, 27(13): 1715–1741, December 1, 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/43/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/43/30/
article.pdf.

Larriba-Pey:1996:RGT

[LPNJR96] Josep-Lluis Larriba-Pey, Juan J. Navarro, Angel Jorba, and Oriol Roig. Review of general and Toeplitz vector bidiagonal solvers. Parallel Computing, 22(8):1091–1125 (or 1091–1126??), October 28, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=8&aid=1079.

Lee:2020:AIR

[LPNV20] JunKyu Lee, Gregory D. Peterson, Dimitrios S. Nikolopoulos, and Hans Vandierendonck. AIR: Iterative refinement acceleration using arbitrary dynamic precision. Parallel Computing, 97(??): Article 102663, September 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819120300569.

Lippert:2001:HSM

[LPPS01] Th. Lippert, N. Petkov, P. Palazzari, and K. Schilling. Hyper-systolic matrix multiplication. Parallel Computing, 27(6):737–759, May 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/47/30/23/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/30/23/
                                        article.pdf.

Lu:1992:APC

[LQ92] Mi Lu and Xiangzhen Qiao. Applying parallel computer systems to solve symmetric tridiagonal eigenvalue problems. Parallel Computing, 18 (12):1301–1315, December 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lootsma:1988:SAP

[LR88] F. A. Lootsma and K. M.

Ragsdell. State-of-the-art in parallel nonlinear optimization. Parallel Computing, 6(2):133–155, February 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lozier:1989:SPC

[LR89] D. W. Lozier and R. G.

Rehm. Some performance comparisons for a fluid dynamics code. Parallel Computing, 11(3):305–320, ???? 28, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lenhardt:1999:KSM

[LR99] Ingrid Lenhardt and Thomas Rottner. Krylov subspace methods for structural finite element analysis. Parallel Computing, 25(7):861–875, August 13, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1999&volume=
25&issue=7&aid=1414;
http://www.elsevier.
nl/gej-ng/10/35/21/32/
23/23/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/32/
23/23/article.pdf.

Lastovetsky:2004:PAH

[LR04] Alexey Lastovetsky and Ravi Reddy. On performance analysis of heterogeneous parallel algorithms. Parallel Computing, 30(11):1195–1216, November 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lastovetsky:2007:DDD

[LR07] Alexey Lastovetsky and Ravi Reddy. Data distribution for dense factorization on computers with memory heterogeneity. Parallel Computing, 33(12): 757–779, December 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lou:2016:PSA

[LR16] Zhihao Lou and John Reinitz. Parallel simulated annealing using an adaptive resampling interval. Parallel Computing, 53(??):23–31, April 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000430.

Lambert:1997:PDM

[LRH97] Michael A. Lambert, Garry H.

Rodrigue, and Dennis W. Hewett. Parallel DS-DADI method for solution of the steady state diffusion equation. Parallel Computing, 23(13): 2041–2065, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=13&aid=1223.

Lenders:1990:PSD

[LS90] P. Lenders and H. Schroder.

A programmable systolic device for image processing based on mathematical morphology. Parallel Computing, 13(3):337–344, March 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Loots:1992:PTP

[LS92] W. Loots and T. H. C.

Smith. A parallel three phase sorting procedure for a k-dimensional hypercube and a transputer implementation. Parallel Computing, 18(3):335–344, March 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lonsdale:1993:MEC

[LS93] G. Lonsdale and A. Schuller. Multigrid efficiency for complex flow simulations on distributed memory machines. Parallel Computing, 19(1):23–32, January 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Liang:1997:FMV

[LS97] Weifa Liang and Xiaojun Shen. Finding the k most vital edges in the minimum spanning tree problem. Parallel Computing, 23(13):1889–1907, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=13&aid=1256.

Latifi:1998:WBS

[LS98] Shahram Latifi and Pradip K. Srimani. Wormhole broadcast in star graph networks. Parallel Computing, 24(8):1263–1276, August 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/8/1334.pdf.

Lega:1995:PAS

[LSA^{+}95] E. Lega, H. Scholl, J.- M. Alimi, A. Bijaoui, and P. Bury. A parallel algorithm for structure detection based on wavelet and segmentation analysis. Parallel Computing, 21(2):265–285, February 17, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=2&aid=942.

Li:2015:CEC

[LSC^{+}15] Jian Li, Sen Su, Xiang Cheng, Meina Song, Liyu Ma, and Jie Wang. Cost-efficient coordinated scheduling for leasing cloud resources on hybrid workloads. Parallel Computing, 44(??):1– 17, May 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0167819115000332.

Lyzinski:2015:SCD

[LSF^{+}15] Vince Lyzinski, Daniel L. Sussman, Donniell E. Fishkind, Henry Pao, Li Chen, Joshua T. Vogelstein, Youngser Park, and Carey E. Priebe. Spectral clustering for divideand-conquer graph matching. Parallel Computing, 47(??):70–87, August 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000484.

Liang:2002:STW

[LSL02] Tyng-Yeu Liang, Ce-Kuen Shieh, and Jun-Qi Li. Selecting threads for workload migration in software distributed shared memory systems. Parallel Computing, 28(6):893– 913, June 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/56/28/
                                        abstract.html.

LExcellent:2014:SSM

[LSL14] Jean-Yves L’Excellent and Wissam M. Sid-Lakhdar. A study of shared-memory parallelism in a multifrontal solver. Parallel Computing, 40(3–4):34–46, March 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000246.

Lippert:1992:QPC

[LSP92] T. Lippert, K. Schilling, and N. Petkov. Quark propagator on the Connection Machine. Parallel Computing, 18(12):1291– 1299, December 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Linden:1988:PMS

[LSS88] J. Linden, B. Steckel, and K. Stuben. Parallel multigrid solution of the Navier– Stokes equations on general 2D domains. Parallel Computing, 7(3):461–475, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Laha:2005:PAI

[LSS05] Arijit Laha, Amitava Sen, and Bhabani P. Sinha. Parallel algorithms for identifying convex and non-convex basis polygons in an image. Parallel Computing, 31(3–4):290–310, March/April 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Louka:1990:TMI

[LT90] Basile Louka and Maurice Tchuente. Triangular matrix inversion on systolic arrays. Parallel Computing, 14(2):223–228, June 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lee:1994:SDG

[LT94] Wei Ping Lee and Jong Chuang Tsay. A systolic design for generating permutations in lexicographic order. Parallel Computing, 20(5):775–785, May 11, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=5&aid=843.

Li:2009:HPH

[LT09] Tien-Yien Li and Chih-Hsiung Tsai. HOM4PS-2.0para: Parallelization of HOM4PS-2.0 for solving polynomial systems. Parallel Computing, 35(4): 226–238, April 2009. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Loghin:2019:TEE

[LT19] Dumitrel Loghin and Yong Meng Teo. The time and energy efficiency of modern multicore systems. Parallel Computing, 86(??):1–13, August 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303740.

Laguna:2023:FIT

[LTG23] Ignacio Laguna, Anh Tran, and Ganesh Gopalakrishnan. Finding inputs that trigger floatingpoint exceptions in heterogeneous computing via Bayesian optimization. Parallel Computing, 117 (??):103042:1–103042:13, September 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000480.

Liu:2024:IFB

[LTS^{+}24] Ke Liu, Haonan Tong, Zhongxiang Sun, Zhixin Ren, Guangkui Huang, Hongyin Zhu, Luyang Liu, Qunyang Lin, and Chuang Zhang. Integrating FPGA-based hardware acceleration with relational databases. Parallel Computing, 119(??):??, February 2024. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819124000024.

Llorente:1996:SAA

[LTV96] Ignacio Martín Llorente, Francisco Tirado, and Luis Vázquez. Some aspects about the scalability of scientific applications on parallel architectures. Parallel Computing, 22(9):1169–1195, November 22, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=9&aid=1089.

Luccio:2001:NSP

[Luc01] Alfredo U. Luccio. Numerical simulation of particle accelerators. Parallel Computing, 27(1–2): 163–177, January 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/47/
                                        25/29/abstract.html;
                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/47/
                                        25/29/article.pdf.

Lundin:1998:CVR

[Lun98] L. K. Lundin. Computing the velocity of a rotating flow. Parallel Computing, 24(14):2021–2034, December 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/14/1364.pdf.

Lotti:1992:AVP

[LV92] G. Lotti and M. Vajtersic.

The application of VLSI Poisson solvers to the biharmonic problem. Parallel Computing, 18(1):11– 19, January 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Liu:2015:SSS

[LV15] Weifeng Liu and Brian Vinter. Speculative segmented sum for sparse matrix-vector multiplication on heterogeneous processors. Parallel Computing, 49(??):179–193, November 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000770.

Lewis:2016:VMD

[LVC16] Christopher Lewis, Miguel Valenciano, and Charles Cornwell. Visualizations of molecular dynamics simulations of high-performance polycrystalline structural ceramics. Parallel Computing, 55(??):35–42, July 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001374.

Li:2011:CLT

[LW11] Bo Li and Koichi Wada. Communication latency tolerant parallel algorithm for particle swarm optimization. Parallel Computing, 37(1):1–10, January 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Liu:2015:SPG

[LW15] Yan Y. Liu and Shaowen Wang. A scalable parallel genetic algorithm for the Generalized Assignment Problem. Parallel Computing, 46(??):98–119, July 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000519.

Li:2006:QAR

[LWG06] Hon F. Li, Zunce Wei, and Dhrubajyoti Goswami. Quasi-atomic recovery for distributed agents. Parallel Computing, 32(10): 733–758, November 2006.

CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lu:2017:HSA

[LWJ^{+}17] Weiming Lu, Yaoguang Wang, Jingyuan Jiang, Jian Liu, Yapeng Shen, and Baogang Wei. Hybrid storage architecture and efficient MapReduce processing for unstructured data. Parallel Computing, 69(??):63–77, November 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301230.

Liang:2002:GLS

[LWS02] Y. Liang, J. Weston, and M. Szularz. Generalized least-squares polynomial preconditioners for symmetric indefinite linear equations. Parallel Computing, 28(2):323–341, February 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/33/36/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/33/36/
main.pdf.

Long:2021:CSC

[LWW^{+}21] Xin Long, Jigang Wu, Yalan Wu, Long Chen, and Yidong Li. Context switch cost aware joint task merging and scheduling for deep learning applications. Parallel Computing, 102 (??):Article 102753, May 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000132.

Li:2024:PPD

[LWZL24] Shushan Li, Meng Wang, Hong Zhang, and Yao Liu. Program partitioning and deadlock analysis for MPI based on logical clocks. Parallel Computing, 119(??):??, February 2024. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000674.

Lin:2000:LST

[LX00] W.-M. Lin and W. Xie. Load-skewing task assignment to minimize communication conflicts on network of workstations. Parallel Computing, 26 (2–3):179–197, February 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/25/24/

abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/25/24/
article.pdf.

Lin:2018:ESM

[LXC^{+}18] James Lin, Zhigeng Xu, Linjin Cai, Akira Nukada, and Satoshi Matsuoka. Evaluating the SW26010 many-core processor with a micro-benchmark suite for performance optimizations. Parallel Computing, 77(??):128–143, September 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301820.

Liu:2023:HPM

[LXD^{+}23] Zeshi Liu, Zhen Xie, Wenqian Dong, Mengting Yuan, Haihang You, and Dong Li. A heterogeneous processing-in-memory approach to accelerate quantum chemistry simulation. Parallel Computing, 116(??):??, July 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000236.

Liu:2022:OCN

[LXL^{+}22] Zhong Liu, Xin Xiao, Chen Li, Sheng Ma, and Deng Rangyu. Optimizing convolutional neural networks on multi-core vector accelerator. Parallel Computing, 112(??):??, September 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000424.

Liu:1993:CFR

[LY93] Zhiyong Liu and Jia-Huai You. Conflict-free routing for BPC-permutations on synchronous hypercubes. Parallel Computing, 19(3): 323–342, March 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lin:2019:SCC

[LYC^{+}19] ShaoFu Lin, ChiChen Yang, YuHsuan Cheng, KengJui Hsu, HungHsin Chen, YuanChing Lin, and Jerry Chou. Student Cluster Competition 2018, team NTHU: Reproducing performance of multiphysics simulations of the tsunamigenic 2004 Sumatra megathrust earthquake on the Intel Skylake architecture. Parallel Computing, 90(??):Article 102569, December 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819119301607.

Lee:2022:CSP

[LyHW^{+}22] Sunwoo Lee, Kai yuan Hou, Kewei Wang, Saba Sehrish, Marc Paterno, James Kowalkowski, Quincey Koziol, Robert B. Ross, Ankit Agrawal, Alok Choudhary, and Wei keng Liao. A case study on parallel HDF5 dataset concatenation for high energy physics data analysis. Parallel Computing, 110(??): ??, May 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121001174.

Lowe:2000:SBT

[LZ00] Welf Löwe and Wolf Zimmermann. Scheduling balanced task-graphs to LogP-machines. Parallel Computing, 26(9):1083– 1108, July 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/30/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/30/25/
article.pdf.

Long:2015:PBD

[LZCT15] Saiqin Long, Yuelong Zhao, Wei Chen, and Yuanbin Tang. A predictionbased dynamic file assignment strategy for parallel file systems. Parallel Computing, 41(??):1–13, January 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001276.

Morales:2000:PDP

[M^{+}00] D. G. Morales et al. Parallel dynamic programming and automata theory. Parallel Computing, 26(1):113–134, January 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/23/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/23/29/
article.pdf.

Mutlu:2022:SSS

[MA22] Gizen Mutlu and Çigdem Inan Aci. SVM-SMO-SGD: a hybrid-parallel support vector machine algorithm using sequential minimal optimization with stochastic gradient descent. Parallel Computing, 113(??): ??, October 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819122000485.

Myers:2021:PWG

[MAA^{+}21] A. Myers, A. Almgren, L. D. Amorim, J. Bell, L. Fedeli, L. Ge, K. Gott, D. P. Grote, M. Hogan, A. Huebl, R. Jambunathan, R. Lehe, C. Ng, M. Rowan, O. Shapoval, M. Thévenet, J.-L. Vay, H. Vincenti, E. Yang, N. Zaïm, W. Zhang, Y. Zhao, and E. Zoni. Porting WarpX to GPUaccelerated platforms. Parallel Computing, 108(??): ??, December 2021. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819121000818.

MacDonald:2002:PFP

[MAB^{+}02] S. MacDonald, J. Anvik, S. Bromling, J. Schaeffer, D. Szafron, and K. Tan. From patterns to frameworks to parallel programs. Parallel Computing, 28 (12):1663–1683, December 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mahmoudi:2017:CCT

[MAB17] Ramzi Mahmoudi, Mohamed Akil, and Mohamed Hédi Bedoui. Concurrent computation of topological watershed on shared memory parallel machines. Parallel Computing, 69(??):78–97, November 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301242.

Mills:2021:TPP

[MAB^{+}21] Richard Tran Mills, Mark F. Adams, Satish Balay, Jed Brown, Alp Dener, Matthew Knepley, Scott E. Kruger, Hannah Morgan, Todd Munson, Karl Rupp, Barry F. Smith, Stefano Zampini, Hong Zhang, and Junchao Zhang. Toward performance-portable PETSc for GPU-based exascale systems. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781912100079X.

Moreira:2016:DBL

[MAD^{+}16] Francis B. Moreira, Marco A. Z. Alves, Matthias Diener, Philippe O. A. Navaux, and Israel Koren. A dynamic block-level execution profiler. Parallel Computing, 54(??):15–28, May 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000405.

Marzulo:2014:CPP

[MAFC14] Leandro A. J. Marzulo, Tiago A. O. Alves, Felipe M. G. França, and Vítor Santos Costa. Couillard: Parallel programming via coarse-grained Data-flow Compilation. Parallel Computing, 40 (10):661–680, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001252.

Maheshwari:1996:IGL

[Mah96] Piyush Maheshwari. Improving granularity and locality of data in multiprocessor execution of functional programs. Parallel Computing, 22(10): 1359–1372, December 15, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=10&aid=1101.

Masliah:2019:AOT

[MAH^{+}19] I. Masliah, A. Abdelfattah, A. Haidar, S. Tomov, M. Baboulin, J. Falcou, and J. Dongarra. Algorithms and optimization techniques for high-performance matrix– matrix multiplications of very small matrices. Parallel Computing, 81(??):1– 21, January 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301091.

Mariani:2017:CTP

[MAJD17] Giovanni Mariani, Andreea Anghel, Rik Jongerius, and Gero Dittmann. Classification of thread profiles for scaling application behavior. Parallel Computing, 66(??):1– 21, August 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300418.

Makino:1994:LFR

[Mak94a]  Jun Makino.  Lagged-

Fibonacci random number generators on parallel computers. Parallel Computing, 20(9): 1357–1367, September 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/

cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=9&aid=893.

Makino:1994:SNP

[Mak94b] Takenori Makino. Shiftnet and power shift-net for parallel processor systems. Parallel Computing, 20(7):1027–1039, July 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=7&aid=877.

Malard:2002:PRM

[Mal02] Joël M. Malard. Parallel restricted maximum likelihood estimation for linear models with a dense exogenous matrix. Parallel Computing, 28(2):343–353, February 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/33/37/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/33/37/
main.pdf.

Meredith:2009:APG

[MAM^{+}09] Jeremy S. Meredith, Gonzalo Alvarez, Thomas A. Maier, Thomas C. Schulthess, and Jeffrey S. Vetter. Accuracy and performance of graphics processors: a quantum Monte Carlo application case study. Parallel Computing, 35(3):151– 163, March 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Manke:2001:PCA

[Man01a] J. W. Manke. Parallel computing in aerospace. Parallel Computing, 27(4): 329–336, March 2001. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/47/
28/23/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/47/
28/23/article.pdf.

Manoharan:2001:ETD

[Man01b] Sathiamoorthy Manoharan. Effect of task duplication on the assignment of dependency graphs. Parallel Computing, 27(3):257–268, February 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/22/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/22/25/
article.pdf.

Marrakchi:1992:OPS

[Mar92] M. Marrakchi. Optimal parallel scheduling for the 2-steps graph with constant task cost. Parallel Computing, 18(2):169–176, February 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Margaritis:1995:SIA

[Mar95] K. G. Margaritis. On the systolic implementation of associative memory artificial neural networks. Parallel Computing, 21(5):825–840, May 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=5&aid=949.

Martone:2014:EMU

[Mar14] Michele Martone. Efficient multithreaded untransposed, transposed or symmetric sparse matrixvector multiplication with the Recursive Sparse Blocks format. Parallel Computing, 40(7):251–270, July 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819114000386.

Mascagni:1998:PLC

[Mas98] Michael Mascagni. Parallel linear congruential generators with prime moduli. Parallel Computing, 24(5–6):923–936, June 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/5-6/1287.pdf.

Mehta:2006:MSG

[MAS06] Paras Mehta, José Nelson Amaral, and Duane Szafron. Is MPI suitable for a generative designpattern system? Parallel Computing, 32(7–8): 616–626, September 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Matstoms:1995:PSF

[Mat95] Pontus Matstoms. Parallel sparse QR factorization on shared memory architectures. Parallel Computing, 21(3):473–486, March 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=3&aid=958.

Mawhinney:1999:TQC

[Maw99] Robert D. Mawhinney.

The 1 Teraflops QCDSP computer. Parallel Computing, 25(10–11):1281– 1296, September 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/32/
26/22/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/32/
26/22/article.pdf.

Miura:1988:TGP

[MB88] K. Miura and R. G. Babb, II. Tradeoffs in granularity and parallelization for a Monte Carlo shower simulation code. Parallel Computing, 8(1–3):91–100, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Murshed:2000:AOA

[MB00] M. Manzur Murshed and Richard P. Brent. Adaptive AT 2 optimal algorithms on reconfigurable meshes. Parallel Computing, 26(11):1447–1458, October 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/32/25/
abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/10/35/21/42/32/25/
                                        article.pdf.

Marcus:2018:SCC

[MBB^{+}18] Z. Marcus, J. Booth, C. Bunn, M. Leger, S. Hance, T. Sweeney, C. McCardwell, and D. Kaeli. Student cluster competition 2017, team Northeastern University: Reproducing vectorization of the Tersoff multi-body potential on the NVIDIA V100. Parallel Computing, 78(??):67–71, October 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118300619.

Megson:1992:SDA

[MBC92] G. M. Megson, O. Brudaru, and D. Comish. Systolic designs for Aitken’s root finding method. Parallel Computing, 18(4): 415–429, April 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Morel:1997:CSE

[MBC97] E. Morel, J. Briat, and J. Chassin de Kergommeaux. Cuts and sideeffects in distributed memory OR-parallel Prolog. Parallel Computing, 22 (13):1883–1896, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=13&aid=1133.

Miller:2012:USA

[MBK12] Phil Miller, Aaron Becker, and Laxmikant Kalé. Using shared arrays in message-driven parallel programs. Parallel Computing, 38(1–2):66–74, January/February 2012. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819111001360.

Mascagni:2004:PLC

[MC04] Michael Mascagni and Hongmei Chi. Parallel linear congruential generators with Sophie–Germain moduli. Parallel Computing, 30(11):1217–1231, November 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Meng:2009:BDT

[MC09] X. Meng and V. Chaudhary. Boosting data throughput for sequence database similarity searches on FPGAs using an adaptive buffering scheme. Parallel Computing, 35(1):1– 11, January 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

McBryan:1988:CMP

[McB88a] O. A. McBryan. The connection machine: PDE solution on 65536 processors. Parallel Computing, 9(1):1–24, December 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

McBryan:1988:NAP

[McB88b] O. A. McBryan. New architectures: performance highlights and new algorithms. Parallel Computing, 7(3):477–499, September 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

McBryan:1994:OMP

[McB94a] Oliver A. McBryan. An overview of message passing environments. Parallel Computing, 20(4):417– 443 (or 417–444??), March 31, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=4&aid=869.

McBryan:1994:SGP

[McB94b] Oliver A. McBryan. The SUPRENUM and GEN-ESIS projects. Parallel Computing, 20(10– 11):1389–1396, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=903.

McBryan:1994:SPP

[McB94c] Oliver A. McBryan. SUPRENUM:

Perspectives and performance. Parallel Computing, 20(10–11):1427–1442, November 3, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cgi-bin/cas/tree/
store/parco/cas_sub/
browse/browse.cgi?year=
1994&volume=20&issue=
10-11&aid=906.

Massaioli:2005:OPA

[MCB05] Federico Massaioli, Filippo Castiglione, and Massimo Bernaschi. OpenMP parallelization of agent-based models. Parallel Computing, 31(10–12):1066–1081, October/December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Marongiu:2016:CNE

[MCB16] Andrea Marongiu, Alessandro Capotondi, and Luca Benini. Controlling NUMA effects in embedded manycore applications with lightweight nested parallelism support. Parallel Computing, 59(??):24–42, November 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116000442.

McCrosky:1989:RPA

[McC89] C. McCrosky. Realizing the parallelism of arraybased computation. Parallel Computing, 10(1):29– 43, March 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Massie:2004:GDM

[MCC04] Matthew L. Massie, Brent N. Chun, and David E. Culler. The ganglia distributed monitoring system: design, implementation, and experience. Parallel Computing, 30(7): 817–840, July 2004. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Moga:1998:PWT

[MCG98] Alina N. Moga, Bogdan Cramariuc, and Moncef Gabbouj. Parallel watershed transformation algorithms for image segmentation. Parallel Computing, 24(14):1981–2001, December 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/14/1362.pdf.

Moreno:2012:LBH

[MCG^{+}12] A. Moreno, E. Cesar, A. Guevara, J. Sorribes, and T. Margalef. Load balancing in homogeneous pipeline based applications. Parallel Computing, 38(3):125–139, March 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001566.

Morales:2001:ETD

[MCM01] F. Jiménez Morales, J. P.

Crutchfield, and M. Mitchell. Evolving two-dimensional cellular automata to perform density classification: a report on work in progress. Parallel Computing, 27(5):571–585, April 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/29/26/

                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/29/26/
                                        article.pdf.

Milanez:2014:TSM

[MCP^{+}14] Teo Milanez, Sylvain Collange, Fernando Magno Quintão Pereira, Wagner Meira, Jr., and Renato Ferreira. Thread scheduling and memory coalescing for dynamic vectorization of SPMD workloads. Parallel Computing, 40(9):548–558, October 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000362.

Ma:2024:PFE

[MCY^{+}24] Kaihao Ma, Zhenkun Cai, Xiao Yan, Yang Zhang, Zhi Liu, Yihui Feng, Chao Li, Wei Lin, and James Cheng. PPS: Fair and efficient black-box scheduling for multi-tenant GPU clusters. Parallel Computing, 120(??):??, June 2024. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819124000206.

McKerrell:1988:MCS

[MD88] A. McKerrell and L. M. Delves. Monte Carlo simulation of neutron diffusion on SIMD architectures. Parallel Computing, 8(1–3):363–370, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mahapatra:2004:AQE

[MD04] Nihar R. Mahapatra and Shantanu Dutt. Adaptive Quality Equalizing: Highperformance load balancing for parallel branchand-bound across applications and computing systems. Parallel Computing, 30(7):867–881, July 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Murli:2008:HPE

[MDC^{+}08] A. Murli, L. D’Amore, L. Carracciuolo, M. Ceccarelli, and L. Antonelli. High performance edgepreserving regularization in 3D SPECT imaging. Parallel Computing, 34(2): 115–132, February 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mai:1984:PAE

[ME84] Shao-Wen Mai and D. J.

Evans. A parallel algorithm for the enumeration of the spanning trees of a graph. Parallel Computing, 1(3–4):275–286, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Megson:1989:AFT

[ME89] G. M. Megson and D. J. Evans. Algorithmic fault tolerance for matrix operations on triangular arrays. Parallel Computing, 10(2):207–219, April 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Margaritis:1990:SDB

[ME90a] K. Margaritis and D. J. Evans. Systolic designs for Bernoulli’s method. Parallel Computing, 15(1–3): 227–240, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Megson:1990:OSD

[ME90b] G. M. Megson and D. J. Evans. An orthogonal systolic design for the assignment problem. Parallel Computing, 16(2–3): 253–267, December 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Megson:1990:SAG

[ME90c] G. M. Megson and D. J. Evans. Systolic arrays for group explicit methods for solving first order hyperbolic equations. Parallel Computing, 16(2–3): 191–205, December 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Margaritis:1992:SIN

[ME92a] K. G. Margaritis and D. J.

Evans. Systolic implementation of neural networks for searching sets of properties. Parallel Computing, 18(3):325–334, March 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Megson:1992:MSL

[ME92b] G. M. Megson and D. J.

Evans. More on systolic line drawing. Parallel Computing, 18(3):355– 358, March 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Morris:1992:MDP

[ME92c] D. Morris and D. G. Evans.

Modelling distributed and parallel computer systems. Parallel Computing, 18(7): 793–806, July 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mantharam:1993:NJS

[ME93a] M. Mantharam and P. J.

Eberlein. New Jacobisets for parallel computations. Parallel Computing, 19(4):437–454, April 1993.

CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mantharam:1993:BRA

[ME93b] Mythili Mantharam and P. J. Eberlein. Block recursive algorithm to generate Jacobi-sets. Parallel Computing, 19(5):481– 496, May 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Megson:1990:RAR

[Meg90a] G. M. Megson. Rank annihilation on a ring of processors. Parallel Computing, 13(1):85–94, January 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Megson:1990:SHM

[Meg90b] G. M. Megson. A systolic helix for matrix triangularisation with partial pivoting. Parallel Computing, 14(2):199–206, June 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mehrmann:1993:DCM

[Meh93] Volker Mehrmann. Divide and conquer methods for block tridiagonal systems. Parallel Computing, 19(3): 257–279, March 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Meier:1985:PPM

[Mei85] U. Meier. A parallel partition method for solving banded systems of linear equations. Parallel Computing, 2(1):33–43, March 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Meier:1986:TPS

[Mei86] U. Meier. Two parallel SOR variants of the Schwarz alternating procedure. Parallel Computing, 3(3):205–215, July 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Melhem:1987:PGJ

[Mel87] Rami Melhem. Parallel Gauss–Jordan elimination for the solution of dense linear systems. Parallel Computing, 4(3):339–343, June 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Melhem:1988:PSL

[Mel88] Rami Melhem. Parallel solution of linear systems with striped sparse matrices. Parallel Computing, 6(2):165–184, February 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Meurant:1987:MCG

[Meu87] Gérard Meurant. Multitasking the conjugate gradient method on the CRAY X-MP/48. Parallel Computing, 5(3):267–280, November 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Meribout:2016:NSM

[MF16] Mahmoud Meribout and Ahmad Firadus. A new systolic multiprocessor architecture for real-time soft tomography algorithms. Parallel Computing, 52(??):144–155, February 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911600003X.

Moreton-Fernandez:2017:TAD

[MFGEL17] Ana Moreton-Fernandez, Arturo Gonzalez-Escribano, and Diego R. Llanos. A technique to automatically determine ad-hoc communication patterns at runtime. Parallel Computing, 69(??):45–62, November 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301254.

Matejka:2019:CPC

[MFS^{+}19] Joel Matějka, Björn Forsberg, Michal Sojka, Přemysl Šůcha, Luca Benini, Andrea Marongiu, and Zdeněk Hanzálek. Combining PREM compilation and static scheduling for highperformance and predictable MPSoC execution. Parallel Computing, 85(??):27–44, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301789.

Mahfoud:1995:PRS

[MG95] Samir W. Mahfoud and David E. Goldberg. Parallel recombinative simulated annealing: a genetic algorithm. Parallel Computing, 21(1):1–28, January 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=1&aid=937.

Mastoras:2019:LBL

[MG19] Aristeidis Mastoras and Thomas R. Gross. Loadbalancing for load-imbalanced fine-grained linear pipelines. Parallel Computing, 85 (??):178–189, July 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0167819118301546.

Marin:2010:SAP

[MGCB^{+}10] Mauricio Marin, Veronica Gil-Costa, Carolina Bonacic, Ricardo Baeza-Yates, and Isaac D. Scherson. Sync/Async parallel search for the efficient design and construction of Web search engines. Parallel Computing, 36(4):153– 168, April 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Muhlenbein:1987:NSM

[MGSK87] H. Mühlenbein, M. Gorges-Schleuter, and O. Kramer. New solutions to the mapping problem of parallel systems: the evolution approach. Parallel Computing, 4(3):269–279, June 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Muhlenbein:1988:EAC

[MGSK88] H. Mühlenbein, M. Gorges-Schleuter, and O. Kramer. Evolution algorithms in combinatorial optimization. Parallel Computing, 7 (1):65–85, April 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mudalige:2013:DIP

[MGT^{+}13] G. R. Mudalige, M. B.

Giles, J. Thiyagalingam, I. Z. Reguly, C. Bertolli, P. H. J. Kelly, and A. E. Trefethen. Design and initial performance of a high-level unstructured mesh framework on heterogeneous parallel systems. Parallel Computing, 39(11):669–692, November 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001166.

Mair:2019:MUD

[MHE19] Jason Mair, Zhiyi Huang, and David Eyers. Manila: Using a densely populated PMC-space for power modelling within largescale systems. Parallel Computing, 82(??):37–56, ???? 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301443.

Mitra:1997:DPA

[MHH97] G. Mitra, I. Hai, and M. T.

Hajian. A distributed processing algorithm for solving integer programs using a cluster of workstations. Parallel Computing, 23(6):733–753, June 20, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=6&aid=1174.

Markus:1997:NQC

[MHK97] Mario Markus, Tomas Hahn, and Ingo Kusch. A novel quantification of cellular automata. Parallel Computing, 23(11): 1635–1642, December 1, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=11&aid=1237.

Montero:2006:BHT

[MHL06] R. S. Montero, E. Huedo, and I. M. Llorente. Benchmarking of high throughput computing applications on Grids. Parallel Computing, 32(4):267–279, April 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Matsuzaki:2006:PSM

[MHT06] Kiminori Matsuzaki, Zhenjiang Hu, and Masato Takeichi. Parallel skeletons for manipulating general trees. Parallel Computing, 32(7–8):590–603, September 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

McCormick:2007:SDP

[MIA^{+}07] Patrick McCormick, Jeff Inman, James Ahrens, Jamaludin Mohd-Yusof, Greg Roth, and Sharen Cummins. Scout: a dataparallel programming language for graphics processors. Parallel Computing, 33(10–11):648–662, November 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Michielse:1990:PAR

[Mic90] P. Michielse. Parallel adaptive reservoir simulation. Parallel Computing, 13(3): 359–368, March 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Michalakes:1997:MSP

[Mic97] J. Michalakes. MM90: a scalable parallel implementation of the Penn State/ NCAR Mesoscale Model (MM5). Parallel Computing, 23(14):2173–2186, December 17, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=14&aid=1265.

Madduri:2011:GPC

[MII^{+}11] Kamesh Madduri, Eun-Jin Im, Khaled Z. Ibrahim, Samuel Williams, Stéphane Ethier, and Leonid Oliker. Gyrokinetic particle-in-cell optimization on emerging multi- and manycore platforms. Parallel Computing, 37(9):501–520, September 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000147.

Milyukova:2001:PAF

[Mil01] O. Yu. Milyukova. Parallel approximate factorization method for solving discrete elliptic equations. Parallel Computing, 27(10):1365–1379, September 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/40/31/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/40/31/
article.pdf.

Missirlis:1987:SPI

[Mis87] Nikolaos M. Missirlis.

Scheduling parallel iterative methods on multiprocessor systems. Parallel Computing, 5(3):295–302, November 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Misic:1998:UBM

[Miš98] Jelena Mišić. Unicastbased multicast algorithm in wormhole-routed star graph interconnection networks. Parallel Computing, 24(2):267–286, February 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/2/1286.pdf.

Mathur:1994:MMA

[MJ94] Kapil K. Mathur and S. Lennart Johnsson. Multiplication of matrices of arbitrary shape on a data parallel computer. Parallel Computing, 20(7):919–951, July 12, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=7&aid=870.

Mahjoub:1995:RPS

[MJ95] Zaher Mahjoub and Mohamed Jemni. Restructuring and parallelizing a static conditional loop. Parallel Computing, 21 (2):339–347, February 17, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=2&aid=943.

Mierendorff:1999:EPM

[MJ99] Hermann Mierendorff and Wolfgang Joppich. Empirical performance modeling for parallel weather prediction codes. Parallel Computing, 25(13– 14):2135–2148, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/36/46/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/46/
article.pdf.

Miller:1997:REU

[MK97] Ethan L. Miller and Randy H. Katz. RAMA: An easy-to-use, highperformance parallel file system. Parallel Computing, 23(4–5):419–446, May 23, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=4-5&aid=1166.

Mohror:2012:TPS

[MK12] Kathryn Mohror and Karen L. Karavanic. Trace profiling: Scalable event tracing on high-end parallel systems. Parallel Computing, 38(4–5):194–225, April/May 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001852.

Murthy:1992:SAR

[MKC92] V. K. Murthy, E. V. Krishnamurthy, and Pin Chen. Systolic algorithm for rational interpolation and Padé approximation. Parallel Computing, 18(1):75– 83, January 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mehta:2021:PFMa

[MKJC21a] Chahak Mehta, Amarnath Karthi, Vishrut Jetly, and Bhaskar Chaudhury. Parallel fast multipole method accelerated FFT on HPC clusters. Parallel Computing, 104–105(??):??, July 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000405.

Mehta:2021:PFMb

[MKJC21b] Chahak Mehta, Amarnath Karthi, Vishrut Jetly, and Bhaskar Chaudhury. Parallel fast multipole method accelerated FFT on HPC clusters. Parallel Computing, 104–105(??):??, July 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000405.

Mermer:2003:EFI

[MKK03] Coskun Mermer, Donglok Kim, and Yongmin Kim. Efficient 2D FFT implementation on mediaprocessors. Parallel Computing, 29(6):691–709, June 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Machovec:2019:UBR

[MKK^{+}19] Dylan Machovec, Bhavesh Khemka, Nirmal Kumbhare, Sudeep Pasricha, Anthony A. Maciejewski, Howard Jay Siegel, Ali Akoglu, Gregory A.

Koenig, Salim Hariri, Cihan Tunc, Michael Wright, Marcia Hilton, Rajendra Rambharos, Christopher Blandin, Farah Fargo, Ahmed Louri, and Neena Imam. Utility-based resource management in an oversubscribed energyconstrained heterogeneous environment executing parallel applications. Parallel Computing, 83(??): 48–72, April 2019. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819117301862.

Muhlenbein:1988:MPE

[MKL^{+}88] H. Mühlenbein, O. Kramer, F. Limburger, M. Mevenkamp, and S. Streitz. MUP-PET: a programming environment for message-based multiprocessors. Parallel Computing, 8(1–3):201– 221, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Manke:2001:PPT

[MKL^{+}01] Joseph W. Manke, G. David Kerlick, David Levine, Subhankar Banerjee, and Eric Dillon. Parallel performance of two applications in the Boeing high performance computing benchmark suite. Parallel Computing, 27(4):457– 475, March 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/28/29/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/28/29/
                                        article.pdf.

Mahjoub:1991:PAR

[MKS91] Z. Mahjoub and F. Karoui-Sahtout. Parallel algorithms for redundant precedence relations elimination in task systems. Parallel Computing, 17 (4–5):471–481, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Marco:2000:TLP

[ML00] N. Marco and S. Lanteri. A two-level parallelization strategy for Genetic Algorithms applied to optimum shape design. Parallel Computing, 26(4):377– 397, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/26/23/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/26/23/
                                        article.pdf.

Manin:2020:CTT

[ML20] Valeriy Manin and Bruno Lang. Cannon-type triangular matrix multiplication for the reduction of generalized HPD eigenproblems to standard form. Parallel Computing, 91(??):Article 102597, March 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301887.

Manin:2023:EPR

[ML23] Valeriy Manin and Bruno Lang. Efficient parallel reduction of bandwidth for symmetric matrices. Parallel Computing, 115(??):??, February 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000042.

MirHassani:2000:CSC

[MLM^{+}00] S. A. MirHassani, C. Lucas, G. Mitra, E. Messina, and C. A. Poojari. Computational solution of capacity planning models under uncertainty. Parallel Computing, 26(5):511– 538, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/27/24/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/27/24/
                                        article.pdf.

Ma:2007:PEF

[MLX07] Meijie Ma, Guizhen Liu, and Jun-Ming Xu. Panconnectivity and edgefault-tolerant pancyclicity of augmented cubes. Parallel Computing, 33(1):36– 42, February 2007. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ma:1994:IMC

[mM94] Chang ming Ma. Implementation of a Monte Carlo code on a parallel computer system. Parallel Computing, 20 (7):991–1005, July 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=7&aid=883.

Makino:1995:PFS

[MM95] Jun Makino and Osamu Miyamura. Parallelized feedback shift register generators of pseudorandom numbers. Parallel Computing, 21(6):1015–1028, June 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=6&aid=987.

Mahapatra:1996:MNN

[MM96a] Rabi N. Mahapatra and Sudipta Mahapatra. Mapping of neural network models onto two-dimensional processor arrays. Parallel Computing, 22(10): 1345–1357, December 15, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=10&aid=1099.

Min:1996:MAI

[MM96b] Dugki Min and Matt W.

Mutka. A model for analyzing interactions in 2-D mesh wormhole-routed multicomputers. Parallel Computing, 22(5):675–699, August 8, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=5&aid=1062.

Maros:2000:ISS

[MM00] I. Maros and G. Mitra. Investigating the sparse simplex algorithm on a distributed memory multiprocessor. Parallel Computing, 26(1):151–170, January 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/42/23/31/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/23/31/
                                        article.pdf.

Mahapatra:1997:PFB

[MMC97] S. Mahapatra, R. N. Mahapatra, and B. N. Chatterji. A parallel formulation of back-propagation learning on distributed memory multiprocessors. Parallel Computing, 22 (12):1661–1675, February 21, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        22&issue=12&aid=1102.

Mohamed:2013:MMM

[MMM13] Hisham Mohamed and Stéphane Marchand-Maillet. MRO-MPI: MapReduce overlapping using MPI and an optimized data exchange policy. Parallel Computing, 39(12):851– 866, December 2013. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819113001026.

Marzolla:2007:PPS

[MMO07] Moreno Marzolla, Matteo Mordacchini, and Salvatore Orlando. Peer-to-peer systems for discovering resources in a dynamic grid. Parallel Computing, 33(4–5):339–358, May 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Moreland:2021:MDC

[MMP^{+}21] Kenneth Moreland, Robert Maynard, David Pugmire, Abhishek Yenpure, Allison Vacanti, Matthew Larsen, and Hank Childs. Minimizing development costs for efficient manycore visualization using MCD^{3}. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781912100082X.

Milovanovic:1990:OAG

[MMS90] Igor Ž. Milovanović, Emina I. Milovanović, and Mile K. Stojčev. An optimal algorithm for Gaussian elimination of band matrices on an MIMD computer. Parallel Computing, 15(1–3):133–145, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Meyerhenke:2009:GPD

[MMS09] Henning Meyerhenke, Burkhard Monien, and Stefan Schamberger. Graph partitioning and disturbed diffusion. Parallel Computing, 35(10–11):544–569, October/November 2009. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mariotti:2022:BMC

[MMSS22] Mirko Mariotti, Daniel Magalotti, Daniele Spiga, and Loriano Storchi. The BondMachine, a moldable computer architecture. Parallel Computing, 109(??):??, March 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121001150.

Melab:2006:PCM

[MMT06] N. Melab, M. Mezmaz, and E.-G. Talbi. Parallel cooperative meta-heuristics on the computational grid.: a case study: the biobjective Flow-Shop problem. Parallel Computing, 32(9):643–659, October 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mezmaz:2007:ELB

[MMT07] M. Mezmaz, N. Melab, and E.-G. Talbi. An efficient load balancing strategy for grid-based branch and bound algorithm. Parallel Computing, 33(4–5): 302–313, May 2007. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Matar:2018:OND

[MMTU18] Hassan Salehe Matar, Erdal Mutlu, Serdar Tasiran, and Didem Unat. Output nondeterminism detection for programming models combining dataflow with shared memory. Parallel Computing, 71(??):42–57, January 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911730193X.

mongaMade:2001:PIF

[mMvdV01] Mardochée Magolu monga Made and Henk A. van der Vorst. Parallel incomplete factorizations with pseudo-overlapped subdomains. Parallel Computing, 27(8):989–1008, July 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/47/34/24/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/34/24/
                                        article.pdf.

Misra:1993:EVI

[MNP93] M. Misra, D. Nassimi, and V. K. Prasanna. Efficient VLSI implementation of iterative solutions to sparse linear systems. Parallel Computing, 19(5):525–544, May 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Moller-Nielsen:1987:PHP

[MNS87] P. Moller-Nielsen and J. Staunstrup. Problemheap: a paradigm for multiprocessor algorithms. Parallel Computing, 4(1): 63–74, February 1987. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Malek:1989:CBM

[MO89] M. Malek and E. Opper. The cylindrical banyan multicomputer: a reconfigurable systolic architecture. Parallel Computing, 10(3):319–327, May 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Murty:1999:EPA

[MO99] Ravi Murty and Daniel Okunbor. Efficient parallel algorithms for molecular dynamics simulations. Parallel Computing, 25(3):217–230, March 22, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/3/1390.pdf.

Morillo:2004:CSE

[MOF04] P. Morillo, J. M. Orduña, and M. Fernández. A comparison study of evolutive algorithms for solving the partitioning problem in distributed virtual environment systems. Parallel Computing, 30(5–6): 585–610, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Moller:1999:SIM

[Möl99] Reinhard Möller. A systolic implementation of the MLEM reconstruction algorithm for positron emission tomography images. Parallel Computing, 25(7):905–920, August 13, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1999&volume=
25&issue=7&aid=1417;
http://www.elsevier.
nl/gej-ng/10/35/21/32/
23/19/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/32/
23/19/article.pdf.

Moore:2004:APG

[Moo04] Michelle Moore. An accurate parallel genetic algorithm to schedule tasks on a cluster. Parallel Computing, 30(5–6):567– 583, May/June 2004. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Moncrieff:1995:PP

[MOW95] D. Moncrieff, R. E. Overill, and S. Wilson. αcritical for parallel processors. Parallel Computing, 21 (3):467–471, March 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=3&aid=971.

Moncrieff:1996:HCM

[MOW96] D. Moncrieff, R. E. Overill, and S. Wilson. Heterogeneous computing machines and Amdahl’s law. Parallel Computing, 22(3):407–413, April 29, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=3&aid=1050.

Modi:1987:IBS

[MP87] J. Modi and R. Prager.

Implementation of bubble sort and the odd-even transposition sort on a rack of transputers. Parallel Computing, 4(3):345– 348, June 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mahadevan:1992:PEB

[MP92] I. Mahadevan and L. M.

Patnaik. Performance evaluation of bidirectional associative memory on a transputer-based parallel system. Parallel Computing, 18(4):401–413, April 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Michalewicz:1995:PSE

[MP95] Marek T. Michalewicz and Mark Priebatsch. Perfect scaling of the electronic structure problem on a SIMD architecture. Parallel Computing, 21(5):853–870, May 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=5&aid=963.

Mezher:2002:PCP

[MP02] Dany Mezher and Bernard Philippe. Parallel computation of pseudospectra of large sparse matrices. Parallel Computing, 28(2):199–221, February 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/33/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/33/30/
main.pdf.

Merigot:2008:PPIa

[MP08a] Alain Merigot and Alfredo Petrosino. Parallel processing for image and video processing. Parallel Computing, 34(12):693, December 2008. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Merigot:2008:PPIb

[MP08b] Alain Merigot and Alfredo Petrosino. Parallel processing for image and video processing: Issues and challenges. Parallel Computing, 34(12): 694–699, December 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mendiratta:2011:TSA

[MP11] Karan Mendiratta and Eric Polizzi. A threaded SPIKE algorithm for solving general banded systems. Parallel Computing, 37(12):733–741, December 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911100130X.

Monien:2000:QML

[MPD00] Burkhard Monien, Robert Preis, and Ralf Diekmann. Quality matching and local improvement for multilevel graph-partitioning. Parallel Computing, 26 (12):1609–1634, November 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/33/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/33/27/
article.pdf.

Muir:2003:PPM

[MPJ03] P. H. Muir, R. N. Pancer, and K. R. Jackson. PMIRKDC: a parallel mono-implicit Runge–Kutta code with defect control for boundary value ODEs. Parallel Computing, 29(6):711– 741, June 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mariani:2013:AAS

[MPZS13] Giovanni Mariani, Gianluca Palermo, Vittorio Zaccaria, and Cristina Silvano. ARTE: an Application-specific Run-Time managEment framework for multi-cores based on queuing models. Parallel Computing, 39(9):504– 519, September 2013. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S016781911300046X.

McCormick:1989:AMA

[MQ89] S. McCormick and D. Quinlan. Asynchronous multilevel adaptive methods for solving partial differential equations on multiprocessors: performance results. Parallel Computing, 12(2): 145–156, November 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Moore:1997:EDD

[MQ97] Jason A. Moore and Michael J. Quinn. Enhancing disk-directed I/O for fine-grained redistribution of file data. Parallel Computing, 23(4–5):477–499, May 23, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=4-5&aid=1168.

Modi:1985:AIS

[MR85] J. J. Modi and J. S. Rollett. An algorithm for inverse square-roots. Parallel Computing, 2(1):69– 71, March 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Modi:1986:SPE

[MR86] J. J. Modi and J. S. Rollett. Some problems of exploiting a pipeline processor. Parallel Computing, 3(3):263–265, July 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Marrakchi:1989:OAG

[MR89] Mounir Marrakchi and Yves Robert. Optimal algorithms for Gaussian elimination on an MIMD computer. Parallel Computing, 12(2):183–194, November 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Martin:2014:LTP

[MRBQO14] Alberto F. Martín, Ruymán Reyes, Rosa M. Badia, and Enrique S. Quintana-Ortí. Leveraging task-parallelism in message-passing dense matrix factorizations using SMPSs. Parallel Computing, 40(5–6):113–128, May 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000441.

Moscinski:1989:TRS

[MRJ89] J. Moscinski, Z. A. Rycerz, and P. W. M. Jacobs. Timing results of some internal sorting algorithms on the ETA 10-P. Parallel Computing, 11(1):117–119, July 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mininni:2011:HMO

[MRRP11] Pablo D. Mininni, Duane Rosenberg, Raghu Reddy, and Annick Pouquet. A hybrid MPI– OpenMP scheme for scalable parallel pseudospectral computations for fluid turbulence. Parallel Computing, 37(6–7):316–326, June/July 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000512.

Montagne:1994:MOG

[MRSB94] E. Montagne, M. Rukoz, R. Surós, and F. Breant. Modeling optimal granularity when adapting systolic algorithms to transputer based supercomputers. Parallel Computing, 20(5):807–814, May 11, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=5&aid=840.

Menezes:1993:APE

[MRT93] B. L. Menezes, I. L. M. Ricarte, and R. Thurimella. Analysis of pipelined external sorting on a reconfigurable message-passing multicomputer. Parallel Computing, 19(8):839–858, August 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Maimaitijiang:2008:PMI

[MRW^{+}08] Yasheng Maimaitijiang, Mohammed Ali Roula, Stuart Watson, Ralf Patz, Robert J. Williams, and Huw Griffiths. Parallelization methods for implementation of a magnetic induction tomography forward model in symmetric multiprocessor systems. Parallel Computing, 34(9): 497–507, September 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Murthy:1989:SAP

[MS89] V. K. Murthy and H. Schröder. Systolic arrays for parallel matrix g-inversion and finding Petri net invariants. Parallel Computing, 11(3):349–359, ???? 28, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Morandi:1990:PAI

[MS90] R. Morandi and F. Sgallari. Parallel algorithms for the iterative solution of sparse least-squares problems. Parallel Computing, 13(3):271–280, March 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Muller:1991:MPT

[MS91] S. M. Muller and D. Scheerer. A method to parallelize tridiagonal solvers. Parallel Computing, 17(2–3): 181–188, June 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Maheshwari:1998:ECA

[MS98] Piyush Maheshwari and Hong Shen. An efficient clustering algorithm for partitioning parallel programs. Parallel Computing, 24(5–6):893–909, June 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/5-6/1281.pdf.

Mierendorff:1999:AMG

[MS99a] Hermann Mierendorff and Helmut Schwamborn. Automatic model generation for performance estimation of parallel programs. Parallel Computing, 25(6):667–680, June 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/6/1418.pdf.

Mohring:1999:SSP

[MS99b] Rolf H. Möhring and Markus W. Schäffter. Scheduling series–parallel orders subject to 0/1communication delays. Parallel Computing, 25 (1):23–40, January 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1999/
                                        25/1/1376.pdf.

Manolakos:2000:SSP

[MS00] Elias S. Manolakos and Haris M. Stellakis. Systematic synthesis of parallel architectures for the computation of higher order cumulants. Parallel Computing, 26(5):655– 676, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/27/31/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/27/31/
                                        article.pdf.

Mahawar:2003:PIM

[MS03a] Hemant Mahawar and Vivek Sarin. Parallel iterative methods for dense linear systems in inductance extraction. Parallel Computing, 29(9):1219–1235, September 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

McCombs:2003:PMI

[MS03b] James R. McCombs and Andreas Stathopoulos. Parallel, multigrain iterative solvers for hiding network latencies on MPPs and networks of clusters. Parallel Computing, 29 (9):1237–1259, September 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mascagni:2004:PPM

[MS04a] Michael Mascagni and Ashok Srinivasan. Parameterizing parallel multiplicative lagged-Fibonacci generators. Parallel Computing, 30(7):899–916, July 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mihajlovic:2004:EPS

[MS04b] Milan D. Mihajlović and David J. Silvester. Efficient parallel solvers for the biharmonic equation. Parallel Computing, 30(1):35– 55, January 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Murai:2020:DEE

[MS20] Hitoshi Murai and Mitsuhisa Sato. Design and evaluation of efficient global data movement in partitioned global address space. Parallel Computing, 96(??):Article 102624, August 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781912030017X.

Marowka:2023:EAH

[MS23] Ami Marowka and Przemysław Stpiczyński. Editorial on advances in high performance programming. Parallel Computing, 117(??):??, September 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819123000431.

Mahini:2009:RPT

[MSA09] Hamid Mahini and Hamid Sarbazi-Azad. Resource placement in three-dimensional tori. Parallel Computing, 35(10–11):535–543, October/November 2009. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Muhlenbein:1991:PGA

[MSB91] H. Mühlenbein, M. Schomisch, and J. Born. The parallel genetic algorithm as function optimizer. Parallel Computing, 17(6–7): 619–632, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mohd-Saman:1993:ISB

[MSE93] M. Y. Mohd-Saman and D. J. Evans. Investigation of a set of Bernstein Tests for the detection of loop parallelization. Parallel Computing, 19(2):197–207, February 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mohd-Saman:1995:IPA

[MSE95] M. Y. Mohd-Saman and D. J. Evans. Interprocedural analysis for parallel computing. Parallel Computing, 21(2): 315–338, February 17, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=2&aid=948.

Muller:2007:ALB

[MSE07] C. Müller, M. Strengert, and T. Ertl. Adaptive load balancing for raycasting of non-uniformly bricked volumes. Parallel Computing, 33(6):406–419, June 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Matheou:2019:TDD

[MSE19] George Matheou, Vassos Soteriou, and Paraskevas Evripidou. Toward datadriven architectural support in improving the performance of future HPC architectures. Parallel Computing, 86(??):82–106, August 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118301121.

Meneses:2014:EPR

[MSK14] Esteban Meneses, Osman Sarood, and Laxmikant V. Kalé. Energy profile of rollback-recovery strategies in high performance computing. Parallel Computing, 40(9):536–547, October 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000350.

Marrocu:1998:PPM

[MSM98] M. Marrocu, R. Scardovelli, and P. Malguzzi. Parallelization and performance of a meteorological limited area model. Parallel Computing, 24 (5–6):911–922, June 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/5-6/1309.pdf.

Martin:2015:EPM

[MSMC15a] Gonzalo Martín, David E.

Singh, Maria-Cristina Marinescu, and Jesús Carretero. Enhancing the performance of malleable MPI applications by using performance-aware dynamic reconfiguration. Parallel Computing, 46 (??):60–77, July 2015. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819115000642.

Martin:2015:TEL

[MSMC15b] Gonzalo Martín, David E.

Singh, Maria-Cristina Marinescu, and Jesús Carretero. Towards efficient large scale epidemiological simulations in Epi-Graph. Parallel Computing, 42(??):88–102, February 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911400115X.

Montanola-Sales:2016:APC

[MSOCG^{+}16] Cristina Montañola-Sales, Bhakti S. S. Onggo, Josep Casanovas-Garcia, Jose María Cela-Espín, and Adriana Kaplan-Marcusán. Approaching parallel computing to simulating population dynamics in demography. Parallel Computing, 59(??):151–170, November 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300618.

Matrone:1993:LPC

[MSP93] A. Matrone, P. Schiano, and V. Puoti. LINDA and PVM: a comparison between two environments for parallel programming. Parallel Computing, 19 (8):949–957, August 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Midorikawa:2005:PSS

[MSS^{+}05] Takashi Midorikawa, Daisuke Shiraishi, Masayoshi Shigeno, Yasuki Tanabe, Toshihiro Hanawa, and Hideharu Amano. The performance of SNAIL-2 a (S2SS-MIN connected multiprocessor with cache coherent mechanism). Parallel Computing, 31(3–4):352–370, March/April 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Marchal:2019:SIP

[MSS19] Loris Marchal, Erik Saule, and Oliver Sinnen. Special issue proposal for the Parallel Computing journal: HeteroPar 2016 and HCW 2016 workshops. Parallel Computing, 83(??):47, April 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119300857.

Miao:2010:POM

[MSSC10] Qiankun Miao, Guangzhong Sun, Jiulong Shan, and Guoliang Chen. Parallelization and optimization of Mfold on shared memory system. Parallel Computing, 36(9):487–494, September 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mierendorff:1994:PMG

[MST94] Hermann Mierendorff, Helmut Schwamborn, and Maurizio Tazza. Performance modelling of grid problems — a case study on the SUPRENUM system. Parallel Computing, 20(10–11):1527–1546, November 3, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cgi-bin/cas/tree/
store/parco/cas_sub/
browse/browse.cgi?year=
1994&volume=20&issue=
10-11&aid=911.

Muller:2015:PSA

[MSV15] Eike Hermann Müller, Robert Scheichl, and Eero Vainikko. Petascale solvers for anisotropic PDEs in atmospheric modelling on GPU clusters. Parallel Computing, 50(??):53–69, December 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115001362.

Moncrieff:1991:PPU

[MSW91] D. Moncrieff, V. R. Saunders, and S. Wilson. Parallel processing using macrotasking in a multi-job environment on a CRAY Y-MP computer. Parallel Computing, 17(6–7): 733–750, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mirin:1998:EFT

[MSW98] A. A. Mirin, D. E. Shumaker, and M. F. Wehner. Efficient filtering techniques for finite-difference atmospheric general circulation models on parallel processors. Parallel Computing, 24(5–6):729– 740, June 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/

store/parco/sub/1998/
24/5-6/1289.pdf.

McInnes:2014:HKN

[MSZM14] Lois Curfman McInnes, Barry Smith, Hong Zhang, and Richard Tran Mills. Hierarchical Krylov and nested Krylov methods for extreme-scale computing. Parallel Computing, 40(1):17–31, January 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001221.

Martin:1989:CAR

[MT89] A. R. Martin and J. V.

Tucker. The concurrent assignment representation of synchronous systems. Parallel Computing, 9(2):227–256, January 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). PARLE: Conference on Parallel Architectures and Languages— Europe (Eindhoven, 1987).

Manoharan:1990:GBS

[MT90] Sathiamoorthy Manoharan and Nigel P. Topham. A general bound on schedule length for independent tasks. Parallel Computing, 16(1):69–73, November 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Manoharan:1991:ADG

[MT91] S. Manoharan and P. Thanisch. Assigning dependency graphs onto processor networks. Parallel Computing, 17(1): 63–73, April 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Manoharan:1995:AAS

[MT95] Sathiamoorthy Manoharan and Nigel P. Topham. An assessment of assignment schemes for dependency graphs. Parallel Computing, 21(1):85–107, January 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=1&aid=932.

Matsumoto:1997:PCD

[MT97] Y. Matsumoto and T. Tokumasu. Parallel computing of diatomic molecular rarefied gas flows. Parallel Computing, 23 (9):1249–1260, November 3, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/

cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=9&aid=1209.

Melab:2000:PAC

[MT00] N. Melab and E.-G.

Talbi. Parallel adaptive computing on metasystems including NOWs. Parallel Computing, 26 (2–3):267–284, February 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/25/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/25/27/
article.pdf.

Migdalas:2003:NOP

[MTK03a] A. Migdalas, G. Toraldo, and V. Kumar. Nonlinear optimization and parallel computing. Parallel Computing, 29(4):375–391, April 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Migdalas:2003:PCN

[MTK03b] A. Migdalas, G. Toraldo, and V. Kumar. Parallel computing in numerical optimization. Parallel Computing, 29(4):373, April 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Morris:1990:EPS

[MTPE90] D. Morris, C. J. Theaker, R. Phillips, and D. G. Evans. An experimental parallel system (EPS). Parallel Computing, 15 (1–3):247–259, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mastroianni:2008:DIS

[MTV08] Carlo Mastroianni, Domenico Talia, and Oreste Verta. Designing an information system for Grids: Comparing hierarchical, decentralized P2P and super-peer models. Parallel Computing, 34(10):593–611, October 2008. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mohr:2007:SPE

[MTW07] Bernd Mohr, Jesper Larsson Träff, and Joachim Worringen. Selected papers from EuroPVM/MPI 2006. Parallel Computing, 33(9):593–594, September 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Muhlenbein:1990:LML

[Müh90] H. Mühlenbein. Limitations of multi-layer perceptron networks-steps towards genetic neural networks. Parallel Computing, 14(3):249–260, August 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Munier:1999:AAS

[Mun99] Alix Munier. Approximation algorithms for scheduling trees with general communication delays. Parallel Computing, 25(1):41–48, January 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/1/1377.pdf.

McBryan:1987:MVO

[MV87a] Oliver A. McBryan and Eric F. Van de Velde. Matrix and vector operations on hypercube parallel processors. Parallel Computing, 5(1–2):117– 125, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Mehrotra:1987:BLP

[MV87b] Piyush Mehrotra and John Van Rosendale. The BLAZE language: a parallel language for scientific programming. Parallel Computing, 5(3):339–361, November 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Michielse:1988:DTW

[Mv88] Peter H. Michielse and Henk A. van der Vorst. Data transport in Wang’s partition method. Parallel Computing, 7(1):87–95, April 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mackens:1999:GMP

[MV99] Wolfgang Mackens and Heinrich Voss. General masters in parallel condensation of eigenvalue problems. Parallel Computing, 25(7):893–903, August 13, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1999&volume=
                                        25&issue=7&aid=1416;
                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/32/
                                        23/25/abstract.html;
                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/32/
                                        23/25/article.pdf.

Marchal:2011:E

[MV11] Loris Marchal and Frédéric Vivien. Editorial. Parallel Computing, 37(10–11): 693, October/November 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001098.

Malakar:2017:DMO

[MV17] Preeti Malakar and Venkatram Vishwanath. Data movement optimizations for independent MPI I/O on the Blue Gene/Q. Parallel Computing, 61(??): 35–51, January 2017. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S016781911630062X.

Madougou:2016:LGP

[MVdLvN16] Souley Madougou, Ana Varbanescu, Cees de Laat, and Rob van Nieuwpoort. The landscape of GPGPU performance modeling tools. Parallel Computing, 56(??):18–33, August 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300114.

Mehrotra:1998:HPF

[MVZ98] Piyush Mehrotra, John Van Rosendale, and Hans Zima. High Performance Fortran: History, status and future. Parallel Computing, 24(3–4):325–354, May 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/3-4/1293.pdf.

Muller-Wichards:1988:PEA

[MW88] Dieter Müller-Wichards. Performance estimates for applications: an algebraic framework. Parallel Computing, 9(1):77–106, December 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Muller-Wichards:1991:PSS

[MW91] Dieter Müller-Wichards. Problem size scaling in the presence of parallel overhead. Parallel Computing, 17(12):1361–1376, December 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Miller:1994:IDS

[MW94] J. J. H. Miller and S. Wang. On the implementation of a 3-D semiconductor device simulator on distributedmemory MIMD/SIMD machines. Parallel Computing, 20(10–11):1689– 1691, November 3, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cgi-bin/cas/tree/
store/parco/cas_sub/
browse/browse.cgi?year=
1994&volume=20&issue=
10-11&aid=924.

Maurer:2011:PBL

[MW11] Daniel Maurer and Christian Wieners. A parallel block LU decomposition method for distributed finite element matrices. Parallel Computing, 37(12):742–758, December 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000548.

Mattor:1995:AST

[MWH95] Nathan Mattor, Timothy J. Williams, and Dennis W. Hewett. Algorithm for solving tridiagonal matrix problems in parallel. Parallel Computing, 21(11):1769–1782, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.

                                        cgi?year=1995&volume=
                                        21&issue=11&aid=1012.

Maeyama:2015:ISS

[MWI^{+}15] Shinya Maeyama, Tomohiko Watanabe, Yasuhiro Idomura, Motoki Nakata, Masanori Nunami, and Akihiro Ishizawa. Improved strong scaling of a spectral/finite difference gyrokinetic code for multi-scale plasma turbulence. Parallel Computing, 49(??):1–12, November 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000794.

Muller-Wichards:1995:SAA

[MWR95] Dieter Müller-Wichards and Wolfgang Rönsch. Scalability of algorithms: an analytic approach. Parallel Computing, 21 (6):937–952, June 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=6&aid=979.

Mo:2007:RRC

[MX07] Zeyao Mo and Xiaowen Xu. Relaxed RS0 or CLJP coarsening strategy for parallel AMG. Parallel Computing, 33(3):174– 185, April 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Muniz:1995:PLB

[MZ95] F. J. Muniz and E. J.

Zaluska. Parallel loadbalancing: an extension to the gradient model. Parallel Computing, 21 (2):287–301, February 17, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=2&aid=933.

Mehrotra:2001:HPF

[MZ01] Piyush Mehrotra and Hans Zima. High Performance Fortran for aerospace applications. Parallel Computing, 27(4):477–501, March 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/28/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/28/30/
article.pdf.

Miller:2004:LSS

[MZ04] John H. Miller and Fang Zheng. Large-scale simulations of cellular signaling processes. Parallel Computing, 30(9–10): 1137–1149, September/ October 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Newns:2000:MDS

[N^{+}00] Dennis M. Newns et al. Molecular dynamics study of structure and gating of low molecular weight ion channels. Parallel Computing, 26(7–8):965–976, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/42/29/31/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/29/31/
                                        article.pdf.

Nagy:2003:MFP

[NA03] Naya Nagy and Selim G. Akl. The maximum flow problem: a real-time approach. Parallel Computing, 29(6):767–794, June 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ng:2008:PAE

[NA08] Kok Fu Ng and Norhashidah Hj. Mohd Ali. Performance analysis of explicit group parallel algorithms for distributed memory multicomputer. Parallel Computing, 34(6–8):427–440, July 2008. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Navarro:2014:CSD

[NAC^{+}14] Angeles Navarro, Rafael Asenjo, Francisco Corbera, Antonio J. Dios, and Emilio L. Zapata. A case study of different task implementations for multioutput stages in nontrivial parallel pipeline applications. Parallel Computing, 40(8):374–393, August 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000635.

Nagel:1988:UMC

[Nag88] W. E. Nagel. Using multiple CPUs for problem solving: experiences in multitasking on the CRAY X-MP/48. Parallel Computing, 8(1–3):223–230, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nagel:1990:EAC

[Nag90] W. E. Nagel. Exploiting autotasking on a CRAY Y-MP: an improved software interface to multitasking.

Parallel Computing, 13(2): 225–233, February 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nakajima:2005:PIS

[Nak05] Kengo Nakajima. Parallel iterative solvers for finiteelement methods using an OpenMP/MPI hybrid programming model on the Earth Simulator. Parallel Computing, 31(10– 12):1048–1065, October/ December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Natarajan:1990:PAG

[Nat90] Ramesh Natarajan. A parallel algorithm for the generalized symmetric eigenvalue problem on a hybrid multiprocessor. Parallel Computing, 14(2):129– 150, June 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nagpal:2012:CGE

[NB12] Rahul Nagpal and Anasua Bhowmik. Criticality guided energy aware speculation for speculative multithreaded processors. Parallel Computing, 38(6–7):329–341, June/July 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000191.

Norris:2002:PCP

[NBB^{+}02] Boyana Norris, Satish Balay, Steven Benson, Lori Freitag, Paul Hovland, Lois McInnes, and Barry Smith. Parallel components for PDEs and optimization: some issues and experiences. Parallel Computing, 28(12):1811–1831, December 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nakano:1997:ACC

[NC97] Aiichiro Nakano and Timothy Campbell. Adaptive curvilinear-coordinate approach to dynamic load balancing of parallel multiresolution molecular dynamics. Parallel Computing, 23(10): 1461–1478, November 7, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=10&aid=1218.

Nkonga:2002:GPM

[NC02] B. Nkonga and P. Charrier. Generalized parcel method for dispersed spray and message passing strategy on unstructured meshes. Parallel Computing, 28(3):369–398, March 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/34/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/34/27/
main.pdf.

Nicastro:1995:OMS

[ND95] L. Nicastro and N. D’Amico. An optimized mass storage FFT for vector computers. Parallel Computing, 21(3):423–432, March 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=3&aid=950.

Nielsen:2017:HPA

[ND17] Eric J. Nielsen and Boris Diskin. High-performance aerodynamic computations for aerospace applications. Parallel Computing, 64(??):20–32, May 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819117300182.

Natu:2020:GAL

[NDN20] Shardul Natu, Ketan Date, and Rakesh Nagi. GPUaccelerated Lagrangian heuristic for multidimensional assignment problems with decomposable costs. Parallel Computing, 97(??):Article 102666, September 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300594.

Narasimhamurthy:2019:SPS

[NDW^{+}19] Sai Narasimhamurthy, Nikita Danilov, Sining Wu, Ganesan Umanesan, Stefano Markidis, Sergio Rivas-Gomez, Ivy Bo Peng, Erwin Laure, Dirk Pleiter, and Shaun de Witt. SAGE: Percipient storage for exascale data centric computing. Parallel Computing, 83(??):22–33, April 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300772.

Noulard:2001:KRP

[NE01] Eric Noulard and Nahid Emad. A key for reusable parallel linear algebra software. Parallel Computing, 27(10):1299–1319, September 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/40/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/40/28/
article.pdf.

Nelson:1993:HMM

[Nel93] P. A. Nelson. Hypercube matrix multiplication. Parallel Computing, 19(7):777–788, July 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Navaridas:2013:SFT

[NFG^{+}13] Javier Navaridas, Steve Furber, Jim Garside, Xin Jin, Mukaram Khan, David Lester, Mikel Luján, José Miguel-Alonso, Eustace Painkras, Cameron Patterson, Luis A. Plana, Alexander Rast, Dominic Richards, Yebin Shi, Steve Temple, Jian Wu, and Shufan Yang. SpiN-Naker: Fault tolerance in a power- and area- constrained large-scale neuromimetic architecture. Parallel Computing, 39 (11):693–708, November 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001051.

Ng:1993:SSC

[Ng93] E. Ng. Supernodal symbolic Cholesky factorization on a local-memory multiprocessor. Parallel Computing, 19(2):153–162, February 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nam:2019:SIS

[NGM19] Hai Ah Nam, Elsa Gonsiorowski, and Scott Michael. Special issue on the SC’18 Student Cluster Competition Reproducibility Initiative. Parallel Computing, 90(??):Article 102572, December 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301632.

Numrich:2011:SSP

[NH11] Robert W. Numrich and Michael A. Heroux. Selfsimilarity of parallel machines. Parallel Computing, 37(2):69–84, February 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nagashima:1995:ESL

[NHS^{+}95] Umpei Nagashima, Sachiko Hyugaji, Satoshi Sekiguchi, Mitsuhisa Sato, and Haruo Hosoya. An experience with super-linear speedup achieved by parallel computing on a workstation cluster: Parallel calculation of density of states of large scale cyclic polyacenes. Parallel Computing, 21 (9):1491–1504, September 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=9&aid=1005.

Nicol:1995:NBS

[Nic95] David M. Nicol. Noncommittal barrier synchronization. Parallel Computing, 21(4):529–549, April 1, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=4&aid=967.

Nicolescu:2002:DTP

[NJ02] Cristina Nicolescu and Pieter Jonker. A data and task parallel image processing environment. Parallel Computing, 28(7–8):945–965, August 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/58/29/
abstract.html.

Nieuwejaar:1997:GPF

[NK97] Nils Nieuwejaar and David Kotz. The Galley parallel file system. Parallel Computing, 23(4–5):447–476, May 23, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=4-5&aid=1167.

Nagel:1991:BPP

[NL91] W. E. Nagel and M. A.

Linn. Benchmarking parallel programs in a multiprogramming environment: the PAR-Bench system. Parallel Computing, 17(10–11):1303–1321, December 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nguyen:1998:IAL

[NL98] Trung N. Nguyen and Zhiyuan Lib. Interprocedural analysis for loop scheduling and data allocation. Parallel Computing, 24(3–4):477–504, May 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/3-4/1299.pdf.

Najjar:1999:ADC

[NLG99] Walid A. Najjar, Edward A. Lee, and Guang R. Gao. Advances in the dataflow computational model. Parallel Computing, 25(13–14):1907–1929, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/36/38/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/38/
article.pdf.

Navaridas:2015:SEM

[NLP^{+}15] Javier Navaridas, Mikel Luján, Luis A. Plana, Steve Temple, and Steve B. Furber. SpiNNaker: Enhanced multicast routing. Parallel Computing, 45(??):49–66, June 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819115000095.

Nasri:2001:OPR

[NM01] Wahid Nasri and Zaher Mahjoub. Optimal parallelization of a recursive algorithm for triangular matrix inversion on MIMD computers. Parallel Computing, 27(13): 1767–1782, December 1, 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/43/32/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/43/32/
article.pdf.

Nagasaka:2019:POM

[NMAB19] Yusuke Nagasaka, Satoshi Matsuoka, Ariful Azad, and Aydin Buluç. Performance optimization, modeling and analysis of sparse matrix-matrix products on multi-core and manycore processors. Parallel Computing, 90(??): Article 102545, December 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911930136X.

Navaridas:2010:RCT

[NMARD10] Javier Navaridas, Jose Miguel-Alonso, Francisco Javier Ridruejo, and Wolfgang Denzel. Reducing complexity in tree-like computer interconnection networks. Parallel Computing, 36(2–3):71–85, February/March 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Norouzi:2024:FDD

[NMI^{+}24] Mohammad Norouzi, Nicolas Morew, Qamar Ilias, Lukas Rothenberger, Ali Jannesari, and Felix Wolf. Fast data-dependence profiling through prior static analysis. Parallel Computing, 119(??):??, February 2024. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819124000012.

Nejad:2013:HSP

[NMMG13] Ashkan Beyranvand Nejad, Anca Molnos, Matias Escudero Martinez, and Kees Goossens. A hardware/software platform for QoS bridging over multi-chip NoC-based systems. Parallel Computing, 39(9):424–441, September 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819113000550.

Nanayakkara:1993:PIR

[NMW93] A. Nanayakkara, D. Moncrieff, and S. Wilson. Performance of IBM RISC System/6000 workstation clusters in a quantum chemical application. Parallel Computing, 19(9): 1053–1062, September 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nakazawa:1999:CPM

[NNB^{+}99] Kisaburo Nakazawa, Hiroshi Nakamura, Taisuke Boku, Ikuo Nakata, and Yoshiyuki Yamashita. CP-PACS: a massively parallel processor at the University of Tsukuba. Parallel Computing, 25(13– 14):1635–1661, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/36/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/28/
article.pdf.

Netti:2022:ODA

[NOG^{+}22] Alessio Netti, Michael Ott, Carla Guillen, Daniele Tafani, and Martin Schulz. Operational data analytics in practice: Experiences from design to deployment in production HPC environments. Parallel Computing, 113(??):??, October 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781912200045X.

Neusius:1992:EDT

[NOS92] C. Neusius, J. Olszewski, and D. Scheerer. An efficient distributed thinning algorithm. Parallel Computing, 18(1):47–55, January 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nota:1995:EPD

[Not95] Y. Nota. An efficient parallel discrete PDE solver. Parallel Computing, 21 (11):1725–1748, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=11&aid=1009.

Nool:2005:PLS

[NP05] Margreet Nool and Michael M. J. Proot. A parallel least-squares spectral element solver for incompressible flow problems on unstructured grids. Parallel Computing, 31(5):414– 438, May 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nagel:2001:PIT

[NR01] Kai Nagel and Marcus Rickert. Parallel implementation of the TRAN-SIMS micro-simulation. Parallel Computing, 27 (12):1611–1639, November 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/42/31/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/42/31/
article.pdf.

Nguyen:2020:DTL

[NR20] Cuong M. Nguyen and Philip J. Rhodes. Delaunay triangulation of large-scale datasets using two-level parallelism. Parallel Computing, 98 (??):Article 102672, October 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300624.

Nobes:2000:CCF

[NRN00] Ross H. Nobes, Alistair P. Rendell, and Jarek Nieplocha. Computational chemistry on Fujitsu vector-parallel processors: Hardware and programming environment. Parallel Computing, 26(7–8): 869–886, July 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/42/
29/27/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/42/
29/27/article.pdf.

Nagel:1994:MTM

[NS94] K. Nagel and A. Schleicher. Microscopic traffic modeling on parallel high performance computers. Parallel Computing, 20(1):125–146, January 16, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=1&aid=812.

Nesheiwat:2002:IDS

[NS02] Jeffrey Nesheiwat and Boleslaw K. Szymanski. Instrumentation database system for performance analysis of parallel scientific applications. Parallel Computing, 28(10): 1409–1449, October ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/60/29/
abstract.html.

Nagpal:2011:CAP

[NS11] Rahul Nagpal and Y. N.

Srikant. Compiler-assisted power optimization for clustered VLIW architectures. Parallel Computing, 37(1):42–59, January 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nakamura:2022:STD

[NSF^{+}22] Tetsuro Nakamura, Shogo Saito, Kei Fujimoto, Masashi Kaneko, and Akinori Shiraga. Spatial- and timedivision multiplexing in CNN accelerator. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000254.

Nakao:2021:GOA

[NSH^{+}21] Masahiro Nakao, Maaki Sakai, Yoshiko Hanada, Hitoshi Murai, and Mitsuhisa Sato. Graph optimization algorithm for low-latency interconnection networks. Parallel Computing, 106(??):??, September 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000569.

Nukada:2023:ECR

[NSM23] Akira Nukada, Taichiro Suzuki, and Satoshi Matsuoka. Efficient checkpoint/restart of CUDA applications. Parallel Computing, 116(??):??, July 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819123000248.

Novakovic:2015:BPH

[NSS15] Vedran Novaković, Sanja Singer, and Sasa Singer. Blocking and parallelization of the Hari–Zimmermann variant of the Falk– Langemeyer algorithm for the generalized SVD. Parallel Computing, 49 (??):136–152, November 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000903.

Nakao:2022:GOA

[NTHY22] Masahiro Nakao, Masaki Tsukamoto, Yoshiko Hanada, and Keiji Yamamoto.

Graph optimization algorithm using symmetry and host bias for low-latency indirect network. Parallel Computing, 114(??):??, December 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000722.

Norcen:2005:HPJ

[NU05] Roland Norcen and Andreas Uhl. High performance JPEG 2000 and MPEG-4 VTC on SMPs using OpenMP. Parallel Computing, 31(10– 12):1082–1098, October/ December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Numrich:2005:PNA

[Num05] Robert W. Numrich. Parallel numerical algorithms based on tensor notation and Co-Array Fortran syntax. Parallel Computing, 31(6):588–607, June 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Neumann:1996:PAA

[NW96] Ingmar Neumann and Wolfgang Wilhelmi. A parallel algorithm for achieving the Smith normal form of an integer matrix. Parallel Computing, 22(10):1399–1412, December 15, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=10&aid=1091.

Nakata:2006:PPD

[NYFK06] Kazuhide Nakata, Makoto Yamashita, Katsuki Fujisawa, and Masakazu Kojima. A parallel primaldual interior-point method for semidefinite programs using positive definite matrix completion. Parallel Computing, 32(1):24– 43, January 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nielsen:1992:DSN

[NZ92] S. S. Nielsen and S. A. Zenios. Data structures for network algorithms on massively parallel architectures. Parallel Computing, 18(9):1033–1052, September 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nielsen:1997:SPB

[NZ97] Soren S. Nielsen and Stavros A. Zenios. Scalable parallel Benders decomposition for stochastic linear programming.

Parallel Computing, 23 (8):1069–1088, July 25, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=8&aid=1202.

Nishtala:2011:TCC

[NZHY11] Rajesh Nishtala, Yili Zheng, Paul H. Hargrove, and Katherine A. Yelick. Tuning collective communication for Partitioned Global Address Space programming models. Parallel Computing, 37(9):576–591, September 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000536.

Osiakwan:1991:PCM

[OA91] Constantine N. K. Osiakwan and Selim G. Akl. Parallel computation of matchings in trees. Parallel Computing, 17(6–7): 643–656, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Oh:2008:LSI

[OA08] Myeongsoo Oh and Kiyoharu Aizawa. Largescale image sensing by a group of smart image sensors. Parallel Computing, 34(12):710–717, December 2008. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

OLeary:2016:CIB

[OAJ^{+}16] Patrick O’Leary, James Ahrens, Sébastien Jourdain, Scott Wittenburg, David H. Rogers, and Mark Petersen. Cinema image-based in situ analysis and visualization of MPAS-ocean simulations. Parallel Computing, 55(??):43–48, July 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115001349.

Oberaigner:1985:PAR

[Obe85] Wilhelm Oberaigner. Parallel algorithms for rounding exact evaluation of sums of products. Parallel Computing, 2(2):173– 182, June 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Oliker:2000:PTM

[OBG00] Leonid Oliker, Rupak Biswas, and Harold N. Gabow. Parallel tetrahedral mesh adaptation with dynamic load balancing. Parallel Computing, 26(12):1583–1608, November 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/33/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/33/26/
article.pdf.

Ongaro:2007:PMF

[OCE^{+}07] T. Esposti Ongaro, C. Cavazzoni, G. Erbacci, A. Neri, and M. V. Salvetti. A parallel multiphase flow code for the 3D simulation of explosive volcanic eruptions. Parallel Computing, 33(7– 8):541–560, August 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Oussaidene:1997:PGP

[OCPT97] Mouloud Oussaidène, Bastien Chopard, Olivier V. Pictet, and Marco Tomassini. Parallel genetic programming and its application to trading model induction. Parallel Computing, 23(8):1183–1198, July 25, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.

                                        cgi?year=1997&volume=
                                        23&issue=8&aid=1203.

Ostrovsky:2001:CAP

[OCSBY01] B. Ostrovsky, G. Crooks, M. A. Smith, and Y. Bar-Yam. Cellular automata for polymer simulation with application to polymer melts and polymer collapse including implications for protein folding. Parallel Computing, 27(5):613–641, April 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/47/29/29/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/29/29/
                                        article.pdf.

Oed:1992:CMC

[Oed92] W. Oed. Cray Y-MP C90: System features and early benchmark results (short communication). Parallel Computing, 18(8):947–954, August 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Orobitg:2015:HPC

[OGC^{+}15] Miquel Orobitg, Fernando Guirado, Fernando Cores, Jordi Llados, and Cedric Notredame. High performance computing improvements on bioinformatics consistency-based multiple sequence alignment tools. Parallel Computing, 42(??):18–34, February 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001215.

Ohbuchi:1985:OPP

[Ohb85] R. Ohbuchi. Overview of parallel processing research in Japan. Parallel Computing, 2(3):219–228, November 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ostromsky:1998:CGP

[OHZ98] Tz. Ostromsky, P. C.

Hansen, and Z. Zlatev. A coarse-grained parallel QR-factorization algorithm for sparse least squares problems. Parallel Computing, 24(5–6):937– 964, June 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/5-6/1311.pdf.

Ovcharenko:2012:NCP

[OID^{+}12] Aleksandr Ovcharenko, Daniel Ibanez, Fabien Delalondre, Onkar Sahni, Kenneth E. Jansen, Christopher D. Carothers, and Mark S. Shephard. Neighborhood communication paradigm to increase scalability in large-scale dynamic scientific applications. Parallel Computing, 38(3):140–156, March 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001554.

Okitsu:2010:HPC

[OIH10] Yusuke Okitsu, Fumihiko Ino, and Kenichi Hagihara. High-performance cone beam reconstruction using CUDA compatible GPUs. Parallel Computing, 36(2–3): 129–141, February/March 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ohira:2021:SVQ

[OIK21] Ryoma Ohira, Md. Saiful Islam, and Humayun Kayesh. Speedup vs. quality: Asynchronous and cluster-based distributed adaptive genetic algorithms for ordered problems. Parallel Computing, 103(??):??, June 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819121000156.

Otstott:2017:TAS

[OILZ17] Douglas Otstott, Latchesar Ionkov, Michael Lang, and Ming Zhao. TCASM: an asynchronous shared memory interface for highperformance application composition. Parallel Computing, 63(??):61–78, April 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300157.

Olsson:1990:DIE

[OJ90] Pelle Olsson and S. Lennart Johnsson. A dataparallel implementation of an explicit method for the threedimensional compressible Navier–Stokes equations. Parallel Computing, 14(1): 1–30, May 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Oden:2022:ICA

[OK22] Lena Oden and Jörg Keller. Improving cryptanalytic applications with stochastic runtimes on GPUs and multicores. Parallel Computing, 112 (??):??, September 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0167819122000412.

Oden:2016:AGC

[OKF16] Lena Oden, Benjamin Klenk, and Holger Fröning. Analyzing GPU-controlled communication with dynamic parallelism in terms of performance and energy. Parallel Computing, 57(??):125–134, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300011.

Okamoto:1992:MDT

[OKSY92] K. Okamoto, Y. Kodama, S. Sakai, and Y. Yamaguchi. Methodologies in development and testing of the dataflow machine EM-4. Parallel Computing, 18 (8):901–912, August 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Obrecht:2013:SLB

[OKTR13] Christian Obrecht, Frédéric Kuznik, Bernard Tourancheau, and Jean-Jacques Roux. Scalable lattice Boltzmann solvers for CUDA GPU clusters. Parallel Computing, 39(6–7):259–270, June/July 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000458.

Oed:1986:MMS

[OL86] W. Oed and O. Lange.

Modelling, measurement, and simulation of memory interference in the Cray X-MP. Parallel Computing, 3 (4):343–358, October 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

OLeary:1987:PIB

[O’L87] Dianne P. O’Leary. Parallel implementation of the block conjugate gradient algorithm. Parallel Computing, 5(1–2):127– 139, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Oliveira:1996:PMM

[Oli96] S. Oliveira. Parallel multigrid methods for transport equations: the anisotropic case. Parallel Computing, 22(4):513–537, June 11, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=4&aid=1063.

Olson:1995:PAH

[Ols95] Clark F. Olson. Parallel algorithms for hierarchical clustering. Parallel Computing, 21(8):1313–1325, August 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=8&aid=996.

Oldehoeft:1990:MAI

[OM90] R. R. Oldehoeft and J. R. McGraw. Mixed applicative and imperative programs. Parallel Computing, 13(2):175–191, February 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Oliker:2011:EPP

[ONB11] Leonid Oliker, Rajesh Nishtala, and Rupak Biswas. Emerging programming paradigms for large-scale scientific computing. Parallel Computing, 37(9):499–500, September 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000883.

Otoo:2013:CED

[ONOK13] E. J. Otoo, Gideon Nimako, and Daniel Ohene-Kwofie. Chunked extendible dense arrays for scientific data storage. Parallel Computing, 39 (12):802–818, December 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000987.

Ortega:1988:FFMb

[OR88] J. M. Ortega and C. H.

Romine. The ijk forms of factorization methods. II. parallel systems. Parallel Computing, 7(2):149– 162, June 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Orii:2010:MEP

[Ori10] Shigeo Orii. Metrics for evaluation of parallel efficiency toward highly parallel processing. Parallel Computing, 36(1):16– 25, January 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

OReilly:2010:HSU

[ORM^{+}10] Una-May O’Reilly, Eric Robinson, Sanjeev Mohindra, Julie Mullen, and Nadya Bliss. Hogs and slackers: Using operations balance in a genetic algorithm to optimize sparse algebra computation on distributed architectures. Parallel Computing, 36 (10–11):635–644, October/ November 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Obermayer:1990:LSS

[ORS90] K. Obermayer, H. Ritter, and K. Schulten. Large-scale simulations of self-organizing neural networks on parallel computers: application to biological modelling. Parallel Computing, 14(3):381–404, August 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ortega:1988:FFMa

[Ort88]  J. M. Ortega.  The

ijk forms of factorization methods. I. vector computers. Parallel Computing, 7(2):135–147, June 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ostermark:1994:PIV

[ÖS94] Ralf Östermark and Martin Saarinen. Parallel implementation of a VAR-MAX algorithm. Parallel Computing, 20(12): 1711–1720, November 28, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=12&aid=898.

Orr:2021:OTS

[OS21] Michael Orr and Oliver Sinnen. Optimal task scheduling for partially heterogeneous systems. Parallel Computing, 107 (??):??, October 2021. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819121000636.

Olariu:1993:ARM

[OSZ93] Stephan Olariu, James L.

Schwing, and Jingyuan Zhang. Applications of reconfigurable meshes to constant-time computations. Parallel Computing, 19(2):229–237, February 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ohmaki:1989:TPE

[OTI^{+}89] K. Ohmaki, S. Tomura, K. Inoue, T. Ito, K. Ito, and K. Torii. TERM: a parallel executable graph reduction machine for equational language. Parallel Computing, 11(1):1– 16, July 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Oksa:2006:EPP

[OV06] Gabriel Okša and Marián Vajteršic. Efficient preprocessing in the parallel block-jacobi SVD algorithm. Parallel Computing, 32(2):166–176, February 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

OLeary:1990:PFH

[OW90] D. P. O’Leary and P. Whitman. Parallel QR factorization by Householder and modified Gram–Schmidt algorithms. Parallel Computing, 16(1):99–112, November 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Olariu:1991:EPA

[OW91] S. Olariu and Z. Wen. An efficient parallel algorithm for multiselection. Parallel Computing, 17(6–7):

689–693, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Overill:1994:PPA

[OW94] Richard E. Overill and Stephen Wilson. Performance of parallel algorithms for the evaluation of power series. Parallel Computing, 20(8): 1205–1213, August 10, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=8&aid=881.

Overill:1997:DPE

[OW97] R. E. Overill and S. Wilson. Data parallel evaluation of univariate polynomials by the Knuth-Eve algorithm. Parallel Computing, 23(13):2115–2127, December 15, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=13&aid=1254.

Okten:2010:PBR

[ÖW10] Giray Ökten and Matthew Willyard. Parameterization based on randomized quasi-Monte Carlo methods. Parallel Computing, 36(7):415–422, July 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Orf:2016:VSL

[OWW16] Leigh Orf, Robert Wilhelmson, and Louis Wicker. Visualization of a simulated long-track EF5 tornado embedded within a supercell thunderstorm. Parallel Computing, 55 (??):28–34, July 2016. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0167819115001465.

Olariu:1991:FOA

[OWZ91] Stephan Olariu, Zhaofang Wen, and Wei Xiong Zhang. A faster optimal algorithm for the measure problem. Parallel Computing, 17(6–7):683–687, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Oyanagi:1999:DSJ

[Oya99] Yoshio Oyanagi. Development of supercomputers in Japan: Hardware and software. Parallel Computing, 25(13–14):1545–1567, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/36/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/25/
article.pdf.

Oliver:2008:IFA

[OYS08] Tim Oliver, Leow Yuan Yeow, and Bertil Schmidt. Integrating FPGA acceleration into HMMer. Parallel Computing, 34(11): 681–691, November 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Owczarz:2002:PMC

[OZ02] Wojciech Owczarz and Zahari Zlatev. Parallel matrix computations in air pollution modelling. Parallel Computing, 28(2):355–368, February 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/33/38/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/33/38/
main.pdf.

Pissaloux:2002:VAA

[PAD02] Edwige Pissaloux, Franck Amiot, and Tharam Dillon. A vision-application adaptable computer concept and its implementation in FreeTIV computer. Parallel Computing, 28(7–8):1203–1219, August 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/58/40/
abstract.html.

Paglieri:1997:PCS

[PAF^{+}97] L. Paglieri, D. Ambrosi, L. Formaggia, A. Quarteroni, and A. L. Scheinine. Parallel computation for shallow water flow: a domain decomposition approach. Parallel Computing, 23(9): 1261–1277, November 3, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=9&aid=1210.

Plata:2005:PID

[PAG^{+}05] Oscar Plata, Rafael Asenjo, Eladio Gutiérrez, Francisco Corbera, Angeles Navarro, and Emilio L. Zapata. On the parallelization of irregular and dynamic programs. Parallel Computing, 31(6):544– 562, June 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pancake:2001:PTT

[Pan01] Cherri M. Pancake. Performance tools for today’s HPC: Are we addressing the right issues? Parallel Computing, 27(11):1403–1415, October 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/41/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/41/28/
article.pdf.

Papadopoulos:1998:DPS

[Pap98] George A. Papadopoulos. Distributed and parallel systems engineering in MANIFOLD. Parallel Computing, 24 (7):1137–1160, July 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/7/1322.pdf.

Parkinson:1986:PEC

[Par86] D. Parkinson. Parallel efficiency can be greater than unity. Parallel Computing, 3(3):261–262, July 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). See comments [FLW87].

Parberry:1987:SPS

[Par87a] I. Parberry. Some practical simulations of impractical parallel computers. Parallel Computing, 4 (1):93–101, February 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Parkinson:1987:OAU

[Par87b] D. Parkinson. Organisational aspects of using parallel computers. Parallel Computing, 5(1–2):75–83, July 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Park:1991:PAU

[Par91] H. Park. A parallel algorithm for the unbalanced orthogonal Procrustes problem. Parallel Computing, 17(8):913–923, October 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Passoni:1999:WPC

[PATC99] Giuseppe Passoni, Giancarlo Alfonsi, Giovanni Tula, and Umberto Cardu. A wavenumber parallel computational code for the numerical integration of the Navier–Stokes equations. Parallel Computing, 25(5):593–611, May 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/5/1395.pdf.

Polymenakos:1994:PSP

[PB94] L. C. Polymenakos and D. P. Bertsekas. Parallel shortest path auction algorithms. Parallel Computing, 20(9): 1221–1247, September 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=9&aid=895.

Pang:1998:SBD

[PB98] Yuan-Ping Pang and Stephen Brimijoin. Supercomputingbased dimeric analog approach for drug optimization. Parallel Computing, 24(9–10):1557–1566, September 1, 1998. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cas/tree/store/parco/
sub/1998/24/9-10/1348.
pdf.

Petschow:2011:MSS

[PB11] M. Petschow and P. Bientinesi. MR^{3}-SMP: a symmetric tridiagonal eigensolver for multi-core architectures. Parallel Computing, 37(12):795–805, December 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001323.

Pena:2016:DOP

[PB16] Antonio J. Peña and Pavan Balaji. A data-oriented profiler to assist in data partitioning and distribution for heterogeneous memory in HPC. Parallel Computing, 51(??):46– 55, January 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001350.

Palkowski:2023:NBS

[PB23] Marek Palkowski and Wlodzimierz Bielecki. NPDP benchmark suite for the evaluation of the effectiveness of automatic optimizing compilers. Parallel Computing, 116(??):??, July 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000224.

Pawley:1989:BBU

[PBTC89] G. S. Pawley, C. F. Baillie, E. Tenenbaum, and W. Celmaster. The BBN Butterfly used to simulate a molecular liquid. Parallel Computing, 11(3):321–329, ???? 28, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Passoni:2001:AIP

[PCA01] Giuseppe Passoni, Paolo Cremonesi, and Giancarlo Alfonsi. Analysis and implementation of a parallelization strategy on a Navier–Stokes solver for shear flow simulations. Parallel Computing, 27 (13):1665–1685, December 1, 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/43/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/43/28/
article.pdf.

Proano:2019:EMS

[PCC19] Julio Proaño, Carmen Carrión, and Blanca Caminero. Empirical modeling and simulation of an heterogeneous cloud computing environment. Parallel Computing, 83(??):118– 134, April 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117301850.

Pastrana-Cruz:2023:LSC

[PCL23] Andres Pastrana-Cruz and Manuel Lafond. A lightweight semi-centralized strategy for the massive parallelization of branching algorithms. Parallel Computing, 116(??):??, July 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819123000303.

Pickartz:2018:RLA

[PCLM18] Simon Pickartz, Carsten Clauss, Stefan Lankes, and Antonello Monti. Revisiting locality-awareness in view of dynamically changing topologies. Parallel Computing, 77(??):1–18, September 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118301455.

Piernas-Canovas:2010:IEA

[PCN10] Juan Piernas-Canovas and Jarek Nieplocha. Implementation and evaluation of active storage in modern parallel file systems. Parallel Computing, 36(1):26– 47, January 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Plimpton:2011:MML

[PD11] Steven J. Plimpton and Karen D. Devine. MapReduce in MPI for large-scale graph algorithms. Parallel Computing, 37(9):610–632, September 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000172.

Pearce:2019:EUO

[Pea19] Olga Pearce. Exploring utilization options of heterogeneous architectures for multi-physics simulations. Parallel Computing, 87(??):35–45, September 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303855.

Peinze:1988:SPS

[Pei88] K. Peinze. The SUPRENUM preprototype: status and experiences. Parallel Computing, 7(3):297–313, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pellegrini:1997:GPB

[Pel97] François Pellegrini. Graph partitioning based methods and tools for scientific computing. Parallel Computing, 23(1–2):153–164, April 16, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=1-2&aid=1153.

Prugger:2016:EPG

[PEO16] Martina Prugger, Lukas Einkemmer, and Alexander Ostermann. Evaluation of the partitioned global address space (PGAS) model for an inviscid Euler solver. Parallel Computing, 60(??):22–40, December 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116301193.

Predari:2017:CIP

[PER17] Maria Predari, Aurélien Esnard, and Jean Roman. Comparison of initial partitioning methods for multilevel direct k-way graph partitioning with fixed vertices. Parallel Computing, 66(??):22–39, August 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300510.

Peters:1984:PPA

[Pet84] F. J. Peters. Parallel pivoting algorithms for sparse symmetric matrices. Parallel Computing, 1(1):99– 110, August 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Peters:1991:SMV

[Pet91] A. Peters. Sparse matrix vector multiplication techniques on the IBM 3090 VF. Parallel Computing, 17(12):1409–1424, December 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Peters:1992:P

[Pet92] F. J. Peters. Preface. Parallel Computing, 18(12): 1289–??, December 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Petrova:1997:PIF

[Pet97] Svetozara Petrova. Parallel implementation of fast elliptic solver. Parallel Computing, 23(8):1113–1128, July 25, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=8&aid=1204.

Pais:2004:UHP

[PFS^{+}04] V. A. Pais, N. Fournier, M. A. Sutton, K. J. Weston, and U. Dragosits. Using High Performance Fortran to parallelise a multilayer atmospheric transport model. Parallel Computing, 30(1):21–33, January 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Peise:2015:HPS

[PFTB15] Elmar Peise, Diego Fabregat-Traver, and Paolo Bientinesi. High performance solutions for bigdata GWAS. Parallel Computing, 42(??):75–87, February 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001161.

Paprzyck:1991:SAB

[PG91] Marcin Paprzyck and Ian Gladwell. Solving almost block diagonal systems on parallel computers. Parallel Computing, 17(2–3):

133–153, June 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Paprzycki:1993:PCA

[PG93] M. Paprzycki and I. Gladwell. A parallel chopping algorithm for ODE boundary value problems. Parallel Computing, 19(6):651– 666, June 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pereira:2016:SHT

[PGAA16] Marcio Machado Pereira, Matthew Gaudet, J. Nelson Amaral, and Guido Araujo. Study of hardware transactional memory characteristics and serialization policies on Haswell. Parallel Computing, 54(??):46–58, May 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001568.

Pjesivac-Grbovic:2007:MCA

[PGBF^{+}07] Jelena Pješivac-Grbović, George Bosilca, Graham E. Fagg, Thara Angskun, and Jack J. Dongarra. MPI collective algorithm selection and quadtree encoding. Parallel Computing, 33(9):613– 623, September 2007. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Poshtkohi:2011:DGB

[PGG11] Alireza Poshtkohi and M. B. Ghaznavi-Ghoushchi. DotDFS: a grid-based high-throughput file transfer system. Parallel Computing, 37(2):114–136, February 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819110001572.

Penaranda:2017:XBH

[PGGL17] Roberto Peñaranda, Crispín Gómez, María Engracia Gómez, and Pedro López. XOR-based HoL-blocking reduction routing mechanisms for direct networks. Parallel Computing, 67(??):57–74, September 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911730087X.

Palyvos-Giannas:2019:GFG

[PGGP19] Dimitris Palyvos-Giannas, Vincenzo Gulisano, and Marina Papatriantafilou. GeneaLog: Fine-grained data streaming provenance in cyber-physical systems. Parallel Computing, 89(??):Article 102552, November 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301437.

Perry:2000:POF

[PGK^{+}00] S. C. Perry, R. H. Grimwood, D. J. Kerbyson, E. Papaefstathiou, and G. R. Nudd. Performance optimization of financial option calculations. Parallel Computing, 26(5):623– 639, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/27/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/27/29/
article.pdf.

Peng:2018:CPB

[PGK^{+}18] Ivy Bo Peng, Roberto Gioiosa, Gokcen Kestor, Jeffrey S. Vetter, Pietro Cicotti, Erwin Laure, and Stefano Markidis. Characterizing the performance benefit of hybrid memory system for HPC applications. Parallel Computing, 76(??):57–69, August 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118301224.

Prikopa:2016:PIR

[PGW16] Karl E. Prikopa, Wilfried N. Gansterer, and Elias Wimmer. Parallel iterative refinement linear least squares solvers based on all-reduce operations. Parallel Computing, 57(??):167–184, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300473.

Pichel:2005:POI

[PHCR05] J. C. Pichel, D. B. Heras, J. C. Cabaleiro, and F. F. Rivera. Performance optimization of irregular codes based on the combination of reordering and blocking techniques. Parallel Computing, 31(8–9):858–876, August/September 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Popescu:2016:PSS

[PHD16] Radu Popescu, Michael A. Heroux, and Simone Deparis. Parallel subdomain solver strategies for the algebraic additive Schwarz preconditioner. Parallel Computing, 57(??):137–153, September 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300448.

Phillips:1991:PBL

[Phi91] C. Phillips. The performance of the BLAS and LAPACK on a shared memory scalar multiprocessor. Parallel Computing, 17(6–7):751–761, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Protze:2022:MDT

[PHM^{+}22] Joachim Protze, Marc-André Hermanns, Matthias S. Müller, Van Man Nguyen, Julien Jaeger, Emmanuelle Saillard, Patrick Carribault, and Denis Barthou. MPI detach — towards automatic asynchronous local completion. Parallel Computing, 109(??):??, March 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121001022.

Pierce:1994:NMP

[Pie94] Paul Pierce. The NX message passing interface. Parallel Computing, 20(4):463–480, March 31, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=4&aid=863.

Perdikaris:2016:VMF

[PIG^{+}16] Paris Perdikaris, Joseph A. Insley, Leopold Grinberg, Yue Yu, Michael E. Papka, and George Em. Karniadakis. Visualizing multiphysics, fluid-structure interaction phenomena in intracranial aneurysms. Parallel Computing, 55(??):9– 16, July 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781911500160X.

Pini:1991:PAP

[Pin91] G. Pini. A parallel algorithm for the partial eigensolution of sparse symmetric matrices on the CRAY Y-MP. Parallel Computing, 17(4–5):553–561, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pirozzi:1993:FNS

[Pir93] Maria Antonietta Pirozzi. The fast numerical solution of mildly nonlinear elliptic boundary value problems on multiprocessors. Parallel Computing, 19(10):1117–1128, October 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pirozzi:1996:FNM

[Pir96] Maria Antonietta Pirozzi.

A fast numerical method for mildly nonlinear parabolic initial boundary value problems. II: The parallel implementation on the Intel Touchstone Delta system. Parallel Computing, 22(9):1281–1285, November 22, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=9&aid=1095.

Piskoulijski:1992:EAP

[Pis92] Pl. Iv. Piskoulijski. Error analysis of parallel algorithm for the solution of a tridiagonal Toeplitz linear system of equations. Parallel Computing, 18(4):431– 438, April 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Patel:1984:PPR

[PJ84] Nisheeth R. Patel and Harry F. Jordan. A parallelized point rowwise successive over-relaxation method on a multiprocessor. Parallel Computing, 1(3–4):207–222, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pavese:2000:CMD

[PJV00] Marc Pavese, Soonmin Jang, and Gregory A. Voth. Centroid molecular dynamics: a quantum dynamics method suitable for the parallel computer. Parallel Computing, 26(7–8):1025–1041, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/29/35/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/29/35/
article.pdf.

Pronold:2022:RBT

[PJW^{+}22] J. Pronold, J. Jordan, B. J. N. Wylie, I. Kitayama, M. Diesmann, and S. Kunkel. Routing brain traffic through the von Neumann bottleneck: Efficient cache usage in spiking neural network simulation code on general purpose computers. Parallel Computing, 113(??):??, October 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000461.

Park:2005:SOA

[PK05] Inho Park and Seon Wook Kim. Study of OpenMP applications on the InfiniBandbased software distributed shared-memory system. Parallel Computing, 31 (10–12):1099–1113, October/December 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Porto:2000:PEP

[PKR00] Stella C. S. Porto, João Paulo F. W. Kitajima, and Celso C. Ribeiro. Performance evaluation of a parallel tabu search task scheduling algorithm. Parallel Computing, 26(1):73– 90, January 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/23/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/23/27/
article.pdf.

Piecuch:2000:PMR

[PL00] Piotr Piecuch and Joseph I.

Landman. Parallelization of multi-reference coupledcluster method. Parallel Computing, 26(7–8):913– 943, July 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/29/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/29/29/
article.pdf.

Pakzad:1997:ICN

[PLP97] M. Pakzad, J. L. Lloyd, and C. Phillips. Independent columns: a new parallel ILU preconditioner for the PCG method. Parallel Computing, 23(6):637– 647 (or 637–648??), June 20, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=6&aid=1184.

Park:2002:ECL

[PLY02] Taesoon Park, Inseon Lee, and Heon Y. Yeom. An efficient causal logging scheme for recoverable distributed shared memory systems. Parallel Computing, 28(11): 1549–1572, November ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-

ng/10/35/21/60/61/28/
abstract.html.

Papaspyropoulos:1989:PDE

[PM89] George T. Papaspyropoulos and D. G. Maritsas. Parallel discrete event simulation with SIMULA. Parallel Computing, 12(3): 359–373, December 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pfenning:1995:OCP

[PM95]  Jörg-Thomas  Pfenning

and Christoph Moll. Optimized communication patterns on workstation clusters. Parallel Computing, 21(3):373–388, March 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=3&aid=964.

Papadopoulos:2003:PBL

[PM03] Apostolos Papadopoulos and Yannis Manolopoulos. Parallel bulk-loading of spatial data. Parallel Computing, 29(10):1419–1444, October 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pascual:2014:AAM

[PMAL14] Jose A. Pascual, Jose Miguel-Alonso, and Jose A. Lozano. Application-aware metrics for partition selection in cube-shaped topologies. Parallel Computing, 40(5–6):129–139, May 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000490.

Prieto:2003:PMS

[PMLT03] Manuel Prieto, Ruben S. Montero, Ignacio M. Llorente, and Francisco Tirado. A parallel multigrid solver for viscous flows on anisotropic structured grids. Parallel Computing, 29(7):907– 923, July 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Perez-Miguel:2010:PED

[PMMAM10] Carlos Pérez-Miguel, Jose Miguel-Alonso, and Alexander Mendiburu. Porting estimation of distribution algorithms to the Cell Broadband Engine. Parallel Computing, 36 (10–11):618–634, October/ November 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pande:2013:FLC

[PMS^{+}13] Sandeep Pande, Fearghal Morgan, Gerard Smit, Tom Bruintjes, Jochem Rutgers, Brian McGinley, Seamus Cawley, Jim Harkin, and Liam McDaid. Fixed latency on-chip interconnect for hardware spiking neural network architectures. Parallel Computing, 39(9):357–371, September 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000549.

Penke:2020:HPS

[PMV^{+}20] Carolin Penke, Andreas Marek, Christian Vorwerk, Claudia Draxl, and Peter Benner. High performance solution of skew-symmetric eigenvalue problems with applications in solving the Bethe–Salpeter eigenvalue problem. Parallel Computing, 96(??):Article 102639, August 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300326.

Putman:2016:GPA

[PODd16] William M. Putman, Lesley Ott, Anton Darmenov, and Arlindo daSilva. A global perspective of atmospheric carbon dioxide concentrations. Parallel Computing, 55(??):2–8, July 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300023.

Peredo:2014:THP

[POHS14] Oscar Peredo, Julián M. Ortiz, José R. Herrero, and Cristóbal Samaniego. Tuning and hybrid parallelization of a genetic-based multi-point statistics simulation code. Parallel Computing, 40(5–6):144–158, May 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000489.

Poplawski:1988:MRG

[Pop88] D. A. Poplawski. Mapping rings and grids onto the FPS T-Series hypercube. Parallel Computing, 7(1): 1–10, April 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Prakash:1998:DCS

[PP98] Ravi Prakash and Dhabaleswar K. Panda. Designing communication strategies for heterogeneous parallel systems. Parallel Computing, 24 (14):2035–2052, December 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/14/1366.pdf.

Pagano:1993:PIA

[PPZ93] F. Pagano, G. Parodi, and R. Zunino. Parallel implementation of associative memories for image classification. Parallel Computing, 19(6):667–684, June 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Page:1990:DFS

[PR90] R. M. R. Page and S. F.

Reddaway. The DAP as a filestore search engine. Parallel Computing, 13(3): 369–376, March 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Plastino:2003:DSA

[PRR03] A. Plastino, C. C. Ribeiro, and N. Rodriguez. Developing SPMD applications with load balancing. Parallel Computing, 29(6):743– 766, June 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pena:2014:CEC

[PRS^{+}14] Antonio J. Peña, Carlos Reaño, Federico Silla, Rafael Mayo, Enrique S. Quintana-Ortí, and José Duato. A complete and efficient CUDA-sharing solution for HPC clusters. Parallel Computing, 40(10):574–588, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001227.

Patrick:1987:IDP

[PRV87] Merrell L. Patrick, Daniel A. Reed, and Robert G. Voigt. The impact of domain partitioning on the performance of a shared memory multiprocessor. Parallel Computing, 5(1– 2):211–217, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Pineau:2008:IHM

[PRV08] Jean-François Pineau, Yves Robert, and Frédéric Vivien. The impact of heterogeneity on masterslave scheduling. Parallel Computing, 34(3):158– 176, March 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pogue:1988:SCD

[PRW88] C. A. Pogue, E. M. Rasmussen, and P. Willett. Searching and clustering of databases using the ICL distributed array processor. Parallel Computing, 8(1–3):399–407, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Prylli:1997:CEM

[Pry97] Loïc Prylli. The CAP-DYN environment and its message-passing library implementation. Parallel Computing, 23(1– 2):107–120, April 16, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=1-2&aid=1150.

Pettorossi:1984:HOC

[PS84] Alberto Pettorossi and Andrzej Skowron. Higherorder communications for concurrent programming. Parallel Computing, 1 (3–4):331–336, December 1984. CODEN PACOEJ.

ISSN 0167-8191 (print), 1872-7336 (electronic).

Papatheodorou:1989:PAA

[PS89a] Theodore S. Papatheodorou and Yiannis G. Saridakis. Parallel algorithms and architectures for multisplitting iterative methods. Parallel Computing, 12(2): 171–182, November 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Petkov:1989:BLS

[PS89b] N. Petkov and F. Sloboda. A bit-level systolic array for digital contour smoothing. Parallel Computing, 12(3):301–313, December 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Prasad:1991:DNP

[PS91] V. V. R. Prasad and C. Siva Ram Murthy. Downloading node programs/data into hypercubes. Parallel Computing, 17(6–7):633–642, September 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pflug:2000:SPO

[PŚ00] G. Ch. Pflug and A. Świȩtanowski. Selected parallel optimization methods for financial management under uncertainty. Parallel Computing, 26(1):3–25, January 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/23/24/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/23/24/
article.pdf.

Polizzi:2006:PHB

[PS06] Eric Polizzi and Ahmed H.

Sameh. A parallel hybrid banded system solver: the SPIKE algorithm. Parallel Computing, 32(2):177–194, February 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pena:2020:GES

[PS20] Antonio J. Pena and Min Si. Guest editorial: Special issue on applications and system software for hybrid exascale systems. Parallel Computing, 91(??):Article 102583, March 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301747.

Psarris:2002:PAT

[Psa02] Kleanthis Psarris. Program analysis techniques for transforming programs for parallel execution. Parallel Computing, 28(3):455–469, March 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/34/31/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/34/31/
main.pdf.

Prades:2015:DND

[PSF^{+}15] Javier Prades, Federico Silla, Holger Fröning, Mondrian Nüssle, and José Duato. On the design of a new dynamic credit-based end-to-end flow control mechanism for HPC clusters. Parallel Computing, 46(??):32–59, July 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000502.

Patzold:1994:PAP

[PSS94] Ortwin Pätzold, Anton Schüller, and Horst Schwichtenberg. Parallel applications and performance measurements on SUPRENUM. Parallel Computing, 20(10– 11):1571–1582, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=914.

Park:2001:ASP

[PSS01] Daeyeon Park, Byeong Hag Seong, and Rafael H. Saavedra. Adaptive software prefetching in scalable multiprocessors using cache information. Parallel Computing, 27 (9):1173–1195, August 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/35/24/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/35/24/
article.pdf.

Paszynski:2018:CTD

[PSW18] Maciej Paszyński, Leszek Siwik, and Maciej Woźniak. Concurrency of threedimensional refined isogeometric analysis. Parallel Computing, 80(??):1–22, December 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300875.

Pasini:2020:PSD

[PTGF20] Massimiliano Lupo Pasini, Bruno Turcksin, Wenjun Ge, and Jean-Luc Fattebert. A parallel strategy for density functional theory computations on accelerated nodes. Parallel Computing, 100(??): Article 102703, December 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819120300910.

Piggott:2004:EPU

[PTK04] David Piggott, Conor Teljeur, and Alan Kelly. Exploring the potential for using the grid to support health impact assessment modelling. Parallel Computing, 30(9–10): 1073–1091, September/ October 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pratas:2012:FGP

[PTS^{+}12] Frederico Pratas, Pedro Trancoso, Leonel Sousa, Alexandros Stamatakis, Guochun Shi, and Volodymyr Kindratenko. Fine-grain parallelism using multi-core, Cell/BE, and GPU systems. Parallel Computing, 38(8):365– 390, August 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001049.

Parsons:1997:POP

[PUSS97] Ian Parsons, Ron Unrau, Jonathan Schaeffer, and Duane Szafron. PI/ OT: Parallel I/O templates. Parallel Computing, 23(4–5):543–570, May 23, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=4-5&aid=1171.

Patronas:2023:ASN

[PVBR23] G. Patronas, N. Vlassopoulos, Ph. Bellos, and D. Reisis. Accelerating the scheduling of the network resources of the nextgeneration optical data centers. Parallel Computing, 115(??):??, February 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000825.

Pekkila:2022:SCH

[PVK^{+}22] Johannes Pekkilä, Miikka S. Väisälä, Maarit J.

Käpylä, Matthias Rheinhardt, and Oskar Lappi. Scalable communication for high-order stencil computations using CUDAaware MPI. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000102.

Plaza:2008:ECP

[PVP08] Antonio Plaza, David Valencia, and Javier Plaza. An experimental comparison of parallel algorithms for hyperspectral analysis using heterogeneous and homogeneous networks of workstations. Parallel Computing, 34(2):92–114, February 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Parkinson:1984:CAG

[PW84] D. Parkinson and M. Wunderlich. A compact algorithm for Gaussian elimination over GF(2) implemented on highly parallel computers. Parallel Computing, 1(1):65–73, August 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pogue:1987:UTS

[PW87] C. A. Pogue and P. Willett. Use of text signatures for document retrieval in a highly parallel environment. Parallel Computing, 4(3):259–268, June 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pelz:1989:MLE

[PW89] Wolfgang Pelz and Layne T.

Watson. Message length effects for solving polynomial systems on a hypercube. Parallel Computing, 10(2):161–176, April 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Phisutthangkoon:2022:OAT

[PW22] Nuntipat Phisutthangkoon and Jeeraporn Werapun. Optimal ATAPE task scheduling on reconfigurable and partitionable hierarchical hypercube networks. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000266.

Pua:2000:MPD

[PWM00] C. S. Pua, M. H. Williams, and D. H. Marwick. Modelling parallel databases with process algebra. Parallel Computing, 26(13– 14):1909–1924, December 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/34/32/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/34/32/
article.pdf.

Peng:2003:SIM

[PWY03] Liang Peng, Weng-Fai Wong, and Chung-Kwong Yuen. SilkRoad II: mixed paradigm cluster computing with RC dag consistency. Parallel Computing, 29(8):1091–1115, August 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Park:2000:ACC

[PY00] Taesoon Park and Heon Y. Yeom. Application controlled checkpointing coordination for fault-tolerant distributed computing systems. Parallel Computing, 26(4):467–482, March 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/26/27/
abstract.html;   http:
//www.elsevier.nl/gej-

ng/10/35/21/42/26/27/
article.pdf.

Pasini:2021:SAOa

[PYLE21a] Massimiliano Lupo Pasini, Junqi Yin, Ying Wai Li, and Markus Eisenbach. A scalable algorithm for the optimization of neural network architectures. Parallel Computing, 104–105(??): ??, July 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000430.

Pasini:2021:SAOb

[PYLE21b] Massimiliano Lupo Pasini, Junqi Yin, Ying Wai Li, and Markus Eisenbach. A scalable algorithm for the optimization of neural network architectures. Parallel Computing, 104–105(??): ??, July 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000430.

Peng:2011:PEA

[PYZC11] Lizhi Peng, Bo Yang, Lei Zhang, and Yuehui Chen. A parallel evolving algorithm for flexible neural tree. Parallel Computing, 37(10– 11):653–666, October/ November 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000615.

Pena:1994:FES

[PZE94] T. F. Pena, E. L. Zapata, and D. J. Evans. Finite element simulation of semiconductor devices on multiprocessor computers. Parallel Computing, 20(8):1129–1159, August 10, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=8&aid=886.

Quaglia:2002:PAA

[QCC02] F. Quaglia, B. Ciciani, and M. Colajanni. Performance analysis of adaptive wormhole routing in a twodimensional torus. Parallel Computing, 28(3):485– 501, March 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/34/33/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/34/33/
main.pdf.

Quan:2019:PCA

[QH19] Zixi Quan and Volker Haarslev. A parallel computing architecture for high-performance OWL reasoning. Parallel Computing, 83(??):34–46, April 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911830142X.

Qin:2006:NFT

[QJ06] Xiao Qin and Hong Jiang.

A novel fault-tolerant scheduling algorithm for precedence constrained tasks in real-time heterogeneous systems. Parallel Computing, 32(5–6):331– 356, June 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Qu:2016:CHT

[QP16] Yun R. Qu and Viktor K. Prasanna. Compact hash tables for decisiontrees. Parallel Computing, 54(??):121–127, May 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911500157X.

Quatember:1992:CCS

[Qua92] B. Quatember. Concept of a crossbar switch for large-scale multiple processor systems in the field of process control. Parallel Computing, 18(12):1415– 1431, December 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Quinn:1988:PSA

[Qui88]  M. J. Quinn.  Paral-

lel sorting algorithms for tightly coupled multiprocessors. Parallel Computing, 6(3):349–357, March 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Qararyah:2021:CGPa

[QWD^{+}21a] Fareed Qararyah, Mohamed Wahib, Doga Dikbayir, Mehmet Esat Belviranli, and Didem Unat. A computational-graph partitioning method for training memory-constrained DNNs. Parallel Computing, 104–105(??):??, July 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000454.

Qararyah:2021:CGPb

[QWD^{+}21b] Fareed Qararyah, Mohamed Wahib, Doga Dikbayir, Mehmet Esat Belviranli, and Didem Unat. A computational-graph partitioning method for training memory-constrained DNNs. Parallel Computing, 104–105(??):??, July 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000454.

Rendell:2000:CCF

[R^{+}00] Alistair P. Rendell et al.

Computational chemistry on Fujitsu vector-parallel processors: Development and performance of applications software. Parallel Computing, 26(7–8):887– 911, July 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/29/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/29/28/
article.pdf.

Ribizel:2020:PSG

[RA20] Tobias Ribizel and Hartwig Anzt. Parallel selection on GPUs. Parallel Computing, 91(??):Article 102588, March 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0167819119301796.

Raghavan:1997:POU

[Rag97] Padma Raghavan. Parallel ordering using edge contraction. Parallel Computing, 23(8):1045–1067, July 25, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=8&aid=1176.

Ralphs:2003:PBC

[Ral03] T. K. Ralphs. Parallel branch and cut for capacitated vehicle routing. Parallel Computing, 29(5): 607–629, May 2003. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ramesh:1999:ICR

[Ram99] S. Ramesh. Implementation of communicating reactive processes. Parallel Computing, 25(6):703–727, June 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1999/
                                        25/6/1405.pdf.

Ramadan:2007:TDM

[Ram07] Omar Ramadan. Three dimensional MPI parallel implementation of the PML algorithm for truncating finite-difference timedomain Grids. Parallel Computing, 33(2):109–115, March 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rantakokko:1997:SPV

[Ran97] Jarmo Rantakokko. Strategies for parallel variational data assimilation. Parallel Computing, 23 (13):2017–2039, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=13&aid=1252.

Rantakokko:2000:PSS

[Ran00] J. Rantakokko. Partitioning strategies for structured multiblock grids. Parallel Computing, 26 (12):1661–1680, November 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/33/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/33/29/
article.pdf.

Rao:1997:EUW

[Rao97] S. Chandra Sekhara Rao. Existence and uniqueness of WZ factorization. Parallel Computing, 23(8):1129–1139, July 25, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=8&aid=1200.

Rapuano:1999:QPA

[Rap99] F. Rapuano. Quenched physics on APE computers. Parallel Computing, 25(10–11):1217–1226, September 1, 1999. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/32/
                                        26/18/abstract.html;
                                        http://www.elsevier.
                                        nl/gej-ng/10/35/21/32/
                                        26/18/article.pdf.

Rauchwerger:1998:RTP

[Rau98] Lawrence Rauchwerger. Run-time parallelization: Its time has come. Parallel Computing, 24(3–4):527– 556, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/

store/parco/sub/1998/
24/3-4/1301.pdf.

Reinhard:2003:PGV

[RB03] Erik Reinhard and Dirk Bartz. Parallel graphics and visualisation. Parallel Computing, 29(3):285– 288, March 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ramos:2014:RPH

[RB14] Luiz Ramos and Ricardo Bianchini. Robust performance in hybrid-memory cooperative caches. Parallel Computing, 40(9):514– 525, October 2014. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819114000477.

Ruefenacht:2017:GRD

[RBB17] Martin Ruefenacht, Mark Bull, and Stephen Booth. Generalisation of recursive doubling for AllReduce: Now with simulation. Parallel Computing, 69(??):24–44, November 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117301199.

Ramdane:2022:BNP

[RBB^{+}22] Yassine Ramdane, Omar Boussaid, Doulkifli Boukraà, Nadia Kabachi, and Fadila Bentayeb. Building a novel physical design of a distributed big data warehouse over a Hadoop cluster to enhance OLAP cube query performance. Parallel Computing, 111(??): ??, July 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000230.

Rottmann:1997:SDS

[RBL97] Valentin Rottmann, Petra Berenbrink, and Reinhard Luling. Simple distributed scheduling policy for parallel interactive continuous media servers. Parallel Computing, 23 (12):1757–1776, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=12&aid=1244.

Riha:2016:IEC

[RBM^{+}16] Lubomír Ríha, Tomás Brzobohatý, Alexandros Markopoulos, Marta Jarosová, Tomás Kozubek, David Horák, and Václav Hapla. Implementation of the efficient communication layer for the highly parallel total FETI and hybrid total FETI solvers. Parallel Computing, 57 (??):154–166, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300333.

Reisner:2020:SLP

[RBMO20] Andrew Reisner, Markus Berndt, J. David Moulton, and Luke N. Olson. Scalable line and plane relaxation in a parallel structured multigrid solver. Parallel Computing, 100 (??):Article 102705, December 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300922.

Rumley:2017:OIE

[RBP^{+}17] Sébastien Rumley, Meisam Bahadori, Robert Polster, Simon D. Hammond, David M. Calhoun, Ke Wen, Arun Rodrigues, and Keren Bergman. Optical interconnects for extreme scale computing systems. Parallel Computing, 64(??):65–80, May 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117300170.

Rouhipour:2010:FBI

[RBS10] Marjan Rouhipour, Peter J. Bentley, and Hooman Shayani. Fast bio-inspired computation using a GPUbased systemic computer. Parallel Computing, 36 (10–11):591–617, October/ November 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Roth:2014:GEI

[RC14] Philip C. Roth and Yong Chen. Guest editors’ introduction to the special issue on “DISCS-2013”. Parallel Computing, 40(10):681, December 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781911400129X.

Rito:2015:ATS

[RC15] Hugo Rito and João Cachopo. Adaptive transaction scheduling for mixed transactional workloads. Parallel Computing, 41 (??):31–49, January 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819114001288.

Roth:2017:SID

[RC17] Philip C. Roth and R. Shane Canon. Special issue on data-intensive scalable computing systems. Parallel Computing, 61(??):1– 2, January 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300017.

Riakiotakis:2011:DDL

[RCAP11] Ioannis Riakiotakis, Florina M. Ciorba, Theodore Andronikos, and George Papakonstantinou. Distributed dynamic load balancing for pipelined computations on heterogeneous systems. Parallel Computing, 37(10– 11):713–729, October/ November 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000196.

Rissland:2007:EFC

[RD07] Peter Rissland and Yuefan Deng. Electrostatic force computation for biomolecules on supercomputers with torus networks. Parallel Computing, 33(2):116–123, March 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rivera:1990:GEP

[RDB^{+}90] F. F. Rivera, R. Doallo, J. D. Bruguera, E. L. Zapata, and R. Peskin. Gaussian elimination with pivoting on hypercubes. Parallel Computing, 14(1):51– 60, May 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Reinefeld:2013:APS

[RDS13] Alexander Reinefeld, Robert Döbbelin, and Thorsten Schütt. Analyzing the performance of SMP memory allocators with iterative MapReduce applications. Parallel Computing, 39(12):879–889, December 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113001191.

Ribeiro:1992:POR

[RE92] C. Ribeiro and D. El Baz. A parallel optimal routing algorithm. Parallel Computing, 18(12):1393–1402, December 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Radi:1998:APT

[RE98] Bouchaib Radi and Jean-François Estrade. Adaptive parallelization techniques in global weather models. Parallel Computing, 24(8):1167–1175, August 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/8/1328.pdf.

Reale:1990:TSM

[Rea90] F. Reale. A tridiagonal solver for massively parallel computer systems. Parallel Computing, 16(2–3): 361–368, December 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Reed:2007:CU

[Ree07] Dan Reed. Changes and updates. Parallel Computing, 33(1):1, February 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rego:2001:NEC

[Reg01] César Rego. Node-ejection chains for the vehicle routing problem: Sequential and parallel algorithms. Parallel Computing, 27(3):201–222, February 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/22/22/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/22/22/
article.pdf.

Reuter:1988:STS

[Reu88] R. Reuter. Solving tridiagonal systems of linear equations on the IBM 3090 VF. Parallel Computing, 8(1–3):371–376, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Reuter:1999:MPE

[Reu99] Andreas Reuter. Methods for parallel execution of complex database queries. Parallel Computing, 25(13–14):2177–2188, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/36/48/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/48/
article.pdf.

Rajopadhye:1990:SSA

[RF90] Sanjay V. Rajopadhye and Richard M. Fujimoto. Synthesizing systolic arrays from recurrence equations. Parallel Computing, 14(2): 163–189, June 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rothberg:1992:PIH

[RG92] E. Rothberg and A. Gupta.

Parallel ICCG on a hierarchical memory multiprocessor — addressing the triangular solve bottleneck. Parallel Computing, 18(7):719–741, July 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ro:2009:CEM

[RG09] Won W. Ro and Jean-Luc Gaudiot. A complexityeffective microprocessor design with decoupled dispatch queues and prefetching. Parallel Computing, 35(5):255–268, May 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Robles-Gomez:2011:ENM

[RGBC11] Antonio Robles-Gómez, Aurelio Bermúdez, and Rafael Casado. Efficient network management applied to source routed networks. Parallel Computing, 37(3):137– 156, March 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000056.

Robles-Gomez:2009:DDM

[RGBCS09] Antonio Robles-Gómez, Aurelio Bermúdez, Rafael Casado, and Åshild Grønstad Solheim. A dynamic distributed mechanism for reconfiguring highperformance networks. Parallel Computing, 35(5): 305–312, May 2009. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rico-Gallego:2015:ILM

[RGDM15] Juan-Antonio Rico-Gallego and Juan-Carlos Díaz-Martín. τ-Lop: Modeling performance of shared memory MPI. Parallel Computing, 46(??):14–31, July 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000447.

Rountree:2013:PHD

[RGdS^{+}13] Barry Rountree, Todd Gamblin, Bronis R. de Supinski, Martin Schulz, David K. Lowenthal, Guy Cobb, and Henry Tufo. Parallelizing heavyweight debugging tools with mpiecho. Parallel Computing, 39(3):156– 166, March 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000907.

Rivas-Gomez:2018:MWS

[RGGP^{+}18] Sergio Rivas-Gomez, Roberto Gioiosa, Ivy Bo Peng, Gokcen Kestor, Sai Narasimhamurthy, Erwin Laure, and Stefano Markidis. MPI windows on storage for HPC applications. Parallel Computing, 77(??):38–56, September 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301571.

Radulovic:2017:LLI

[RGT17] Milan B. Radulović, Sylvain Girbal, and Milo V. Tomasević. Low-level implementation of the SISC protocol for thread-level speculation on a multicore architecture. Parallel Computing, 67(??):1–19, September 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300972.

Reeve:1999:EPV

[RH99] J. S. Reeve and M. Heath.

An efficient parallel version of the householder-QL matrix diagonalisation algorithm. Parallel Computing, 25(3):311–319, March 22, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/3/1392.pdf.

Rutar:2012:DCT

[RH12] Nick Rutar and Jeffrey K. Hollingsworth. Data centric techniques for mapping performance data to program variables. Parallel Computing, 38(1–2): 2–14, January/February 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001372.

Rutar:2013:STN

[RH13] Nick Rutar and Jeffrey K. Hollingsworth. Software techniques for negating skid and approximating cache miss measurements. Parallel Computing, 39(3):120–131, March 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000786.

Roberts:1988:EPP

[RHM^{+}88] J. B. G. Roberts, J. G.

Harp, B. C. Merrifield, K. J. Palmer, P. Simpson, J. S. Ward, and H. C. Webber. Evaluating parallel processors for realtime applications. Parallel Computing, 8(1–3):245– 254, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Reitz:2023:LBL

[RHWF23] Lukas Reitz, Kai Hardenbicker, Tobias Werner, and Claudia Fohry. Lifelinebased load balancing schemes for Asynchronous Many-Task runtimes in clusters. Parallel Computing, 116(??):??, July 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000261.

Ribeiro:1984:PEV

[Rib84] Celso Ribeiro. Performance evaluation of vector implementations of combinatorial algorithms. Parallel Computing, 1(3–4): 287–294, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rucinski:2010:IMT

[RIB10] M. Ruciński, D. Izzo, and F. Biscani. On the impact of the migration topology on the Island Model. Parallel Computing, 36 (10–11):555–571, October/ November 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Risset:1990:IGE

[Ris90] Tanguy Risset. Implementing Gaussian elimination on a matrix-matrix multiplication systolic array. Parallel Computing, 16(2–3):351–359, December 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rizzi:1985:VCF

[Riz85] Arthur Rizzi. Vector coding the finite-volume procedure for the CYBER 205. Parallel Computing, 2(4):295–312, December 1985. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Reinhard:1997:RLS

[RJ97] Erik Reinhard and Frederik W. Jansen. Rendering large scenes using parallel ray tracing. Parallel Computing, 23(7):873–885, July 14, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=7&aid=1189.

Roman:2010:FEC

[RKMJ10] Jose E. Roman, Matthias Kammerer, Florian Merz, and Frank Jenko. Fast eigenvalue calculations in a massively parallel plasma turbulence code. Parallel Computing, 36(5–6):339– 358, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Robic:1992:AOD

[RKS92] B. Robic, P. Kolbezen, and J. Silc. Area optimization of dataflow-graph mappings. Parallel Computing, 18(3):297–311, March 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rippl:2019:PEC

[RLH19] Michael Rippl, Bruno Lang, and Thomas Huckle. Parallel eigenvalue computation for banded generalized eigenvalue problems. Parallel Computing, 88(??):Article 102542, ???? 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301279.

Rabhi:1991:DCP

[RM91] F. A. Rabhi and G. A. Manson. Divide-and-conquer and parallel graph reduction. Parallel Computing, 17(2–3):189–205, June 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rao:1997:DCP

[RM97] P. S. Rao and G. Mouney. Data communication in parallel block predictorcorrector methods for solving ODE’s. Parallel Computing, 23(13): 1877–1888, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=13&aid=1260.

Risco-Martin:2010:PEA

[RMACG10] José L. Risco-Martín, David Atienza, J. Manuel Colmenar, and Oscar Garnica. A parallel evolutionary algorithm to optimize dynamic memory managers in embedded systems. Parallel Computing, 36(10–11):572–590, October/November 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Richard:2005:CCD

[RMRN05] Bruno Richard, Nicolas Maillard, César A. F. De Rose, and Reynaldo Novaes. The I-Cluster Cloud: distributed management of idle resources for intense computing. Parallel Computing, 31(8–9):813–838, August/September 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ramesh:2018:MPE

[RMS^{+}18a] Srinivasan Ramesh, Aurèle Mahéo, Sameer Shende, Allen D. Malony, Hari Subramoni, Amit Ruhela, and Dhabaleswar K. (DK) Panda. MPI performance engineering with the MPI tool interface: the integration of MVAPICH and TAU. Parallel Computing, 77(??):19–37, September 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301479.

Rizzi:2018:EIR

[RMS^{+}18b] F. Rizzi, K. Morris, K. Sargsyan, P. Mycek, C. Safta, O. Le Maître, O. M. Knio, and B. J. Debusschere. Exploring the interplay of resilience and energy consumption for a task-based partial differential equations preconditioner. Parallel Computing, 74(??): 16–27, ???? 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300753.

Romine:1988:PST

[RO88] Charles H. Romine and James M. Ortega. Parallel solution of triangular systems of equations. Parallel Computing, 6(1):109–114, January 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Robb:1999:VBC

[Rob99] Richard A. Robb. Visualization in biomedical computing. Parallel Computing, 25(13–14):2067–2110, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/36/44/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/44/
article.pdf.

Robson:2000:SMP

[Rob00] Barry Robson. Simplified models of protein folding exploiting the Lagrange radius of gyration of the hydrophobic component. Parallel Computing, 26(7–8):977–998, July 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/29/32/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/29/32/
article.pdf.

Rodrigue:1985:IOI

[Rod85] Garry Rodrigue. Inner/ outer iterative methods and numerical Schwarz algorithms. Parallel Computing, 2(3):205–218, November 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ronsch:1984:SAU

[Ron84] W. Ronsch. Stability aspects in using parallel algorithms. Parallel Computing, 1(1):75–98, August 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rothberg:1995:ASS

[Rot95] Edward Rothberg. Alternatives for solving sparse triangular systems on distributed-memory multiprocessors. Parallel Computing, 21(7):1121–1136, July 11, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=7&aid=982.

Romero:2001:DTP

[ROZ01] L. F. Romero, E. M. Ortigosa, and E. L. Zapata. Data-task parallelism for the VMEC program. Parallel Computing, 27(10):1347–1364, September 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/40/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/40/30/
article.pdf.

Reed:1985:PIS

[RP85] Daniel A. Reed and Merrell L. Patrick. Parallel, iterative solution of sparse linear systems: models and architectures. Parallel Computing, 2(1):45–67, March 1985. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

RadicatidiBrozolo:1989:PCG

[RR89] G. Radicati di Brozolo and Y. Robert. Parallel conjugate gradient-like algorithms for solving sparse nonsymmetric linear systems on a vector multiprocessor. Parallel Computing, 11(2):223–239, ???? 4, 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Renaud:1997:FMP

[RR97] Christophe Renaud and François Rousselle. Fast massively parallel progressive radiosity on the MP-1. Parallel Computing, 23(7):899–913, July 14, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=7&aid=1191.

Ribeiro:2007:EPC

[RR07] Celso C. Ribeiro and Isabel Rosseti. Efficient parallel cooperative implementations of GRASP heuristics. Parallel Computing, 33(1):21–35, February 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rubensson:2014:CTP

[RR14] Emanuel H. Rubensson and Elias Rudberg. Chunks and tasks: a programming model for parallelization of dynamic algorithms. Parallel Computing, 40(7):328–343, July 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911300118X.

Rubensson:2016:LAP

[RR16] Emanuel H. Rubensson and Elias Rudberg. Localityaware parallel block-sparse matrix-matrix multiplication using the Chunks and Tasks programming model. Parallel Computing, 57(??):87–106, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300606.

Robey:2011:SNC

[RRA11] Robert W. Robey, Jonathan M. Robey, and Rob Aulwes. In search of numerical consistency in parallel programming. Parallel Computing, 37(4–5):217–229, April/May 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000238.

Rastello:2003:OTS

[RRP03] Fabrice Rastello, Amit Rao, and Santosh Pande. Optimal task scheduling at run time to exploit intratile parallelism. Parallel Computing, 29(2):209–239, February 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Radicati:1988:DLS

[RRS88] G. Radicati, Y. Robert, and P. Sguazzero. Dense linear systems FORTRAN solvers on the IBM 3090 vector multiprocessor. Parallel Computing, 8(1–3): 377–384, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rodriguez:2000:NPM

[RRS^{+}00] C. Rodríguez, J. L. Roda, F. Sande, D. G. Morales, and F. Almeida. A new parallel model for the analysis of asynchronous algorithms. Parallel Computing, 26(6):753–767, May 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/28/27/

                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/28/27/
                                        article.pdf.

Ronsch:1987:TRS

[RS87] W. Ronsch and H. Strauss. Timing results of some internal sorting algorithms on vector computers. Parallel Computing, 4(1):49– 61, February 1987. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ronsch:1988:LAP

[RS88] W. Ronsch and H. Strauss. A linear algebra package for a local memory multiprocessor: problems, proposals and solutions. Parallel Computing, 7(3):413– 418, September 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Robert:1992:RCS

[RS92] Y. Robert and S. W. Song. Revisiting cycle shrinking. Parallel Computing, 18(5): 481–496, May 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rao:2008:PSL

[RS08] S. Chandra Sekhara Rao and Sarita. Parallel solution of large symmetric tridiagonal linear systems. Parallel Computing, 34(3):

177–197, March 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Raffin:2005:PGV

[RSB05] Bruno Raffin, Han-Wei Shen, and Dirk Bartz. Parallel graphics and visualization. Parallel Computing, 31(2):147–148, February 2005. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ruhela:2019:EDM

[RSC^{+}19] Amit Ruhela, Hari Subramoni, Sourav Chakraborty, Mohammadreza Bayatpour, Pouya Kousha, and Dhabaleswar K. (DK) Panda. Efficient design for MPI asynchronous progress without dedicated resources. Parallel Computing, 85(??):13–26, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303302.

Rennich:2016:ASC

[RSD16] Steven C. Rennich, Darko Stosic, and Timothy A. Davis. Accelerating sparse Cholesky factorization on GPUs. Parallel Computing, 59(??):140–150, November 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911630059X.

Rathe:1999:CSS

[RSK99] U. W. Rathe, P. Sanders, and P. L. Knight. A case study in scalability: An ADI method for the two-dimensional timedependent Dirac equation. Parallel Computing, 25(5):525–533, May 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/5/1403.pdf.

Ramachandran:1996:SSK

[RSRM96] Umakishore Ramachandran, Gautam Shah, S. Ravikumar, and Jeyakumar Muthukumarasamy. Scalability study of the KSR-1. Parallel Computing, 22(5):739–759, August 8, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=5&aid=1072.

Ritzdorf:1994:EPM

[RSSS94] Hubert Ritzdorf, Anton Schüller, Barbara A. Steckel, and Klaus Stüben. L_{i}SS — an environment for the parallel multigrid solution of partial differential equations on general 2D domains. Parallel Computing, 20(10– 11):1559–1570, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=913.

Reed:2005:E

[RST05] Daniel A. Reed, Mitsuhisa Sato, and Denis Trystram. Editorial. Parallel Computing, 31(5):413, May 2005. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Robert:2011:PCS

[RST11] Yves Robert, Leonel Sousa, and Denis Trystram. Parallel computing — special issue. Parallel Computing, 37(8):329–330, August 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000330.

Robert:1988:CSP

[RT88] Yves Robert and Denis Trystram. Comments on scheduling parallel iterative methods on multiprocessor systems. Parallel Computing, 7(2):253–255, June 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rothe:1995:FPC

[RV95] Kai Rothe and Heinrich Voss. A fully parallel condensation method for generalized eigenvalue problems on distributed memory computers. Parallel Computing, 21(6):907–921, June 12, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=6&aid=980.

Robic:1996:ISM

[RV96] Borut Robič and Boštjan Vilfan. Improved schemes for mapping arbitrary algorithms onto processor meshes. Parallel Computing, 22(5):701–724, August 8, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/

cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=5&aid=1070.

Rul:2010:PBT

[RVD10] Sean Rul, Hans Vandierendonck, and Koen De Bosschere. A profile-based tool for finding pipeline parallelism in sequential programs. Parallel Computing, 36(9):531–551, September 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Reijns:2005:PET

[RvG05] G. L. Reijns and A. J. C.

van Gemund. Predicting the execution times of parallel-independent programs using Pearson distributions. Parallel Computing, 31(8–9):877–899, August/September 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Royo:2001:IOS

[RVGG01] Dolors Royo, Miguel Valero-García, and Antonio González. Implementing the onesided Jacobi method on a 2D/3D mesh multicomputer. Parallel Computing, 27(9):1253–1271, August 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-

ng/10/35/21/47/35/27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/35/27/
article.pdf.

Ruppelt:1989:ATH

[RW89] Th. Ruppelt and G. Wirtz. Automatic transformation of high-level objectoriented specifications into parallel programs. Parallel Computing, 10(1):15– 28, March 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rastogi:1996:PAL

[RW96] Sanjeev R. Rastogi and Norman J. Wagner. A parallel algorithm for Lees-Edwards boundary conditions. Parallel Computing, 22(6):895–901, September 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=6&aid=1078.

Roerdink:1998:DPT

[RW98] Jos B. T. M. Roerdink and Michel A. Westenberg. Data-parallel tomographic reconstruction: a comparison of filtered backprojection and direct Fourier reconstruction. Parallel Computing, 24(14):2129–2142, December 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/14/1372.pdf.

Ray:1997:PIF

[RWT97] S. E. Ray, G. P. Wren, and T. E. Tezduyar. Parallel implementations of a finite element formulation for fluid-structure interactions in interior flows. Parallel Computing, 23 (9):1279–1292, November 3, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=9&aid=1211.

Romero:1995:DDS

[RZ95] L. F. Romero and E. L.

Zapata. Data distributions for sparse matrix vector multiplication. Parallel Computing, 21(4):583–605, April 1, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=4&aid=953.

Sarje:2013:APC

[SA13] Abhinav Sarje and Srinivas Aluru. All-pairs computations on many-core graphics processors. Parallel Computing, 39(2):79–93, February 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000094.

Selvitopi:2016:RLC

[SA16] Oguz Selvitopi and Cevdet Aykanat. Reducing latency cost in 2D sparse matrix partitioning models. Parallel Computing, 57(??):1–24, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300138.

Spychalski:2018:MLM

[SA18] Przemyslaw Spychalski and Ryszard Arendt. Machine learning in multiagent systems using associative arrays. Parallel Computing, 75(??):88–99, July 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118300814.

Sabbatel:1997:HSE

[Sab97] Gilles Berger Sabbatel.

Hardware solutions for efficient distributed computing on ATM networks. Parallel Computing, 23(1–2):35–48, April 16, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=1-2&aid=1146.

Schenk:2021:GEV

[SAGV21] Olaf Schenk, Peter Arbenz, Luc Giraud, and Wim Vanroose. Guest editorial: Virtual special issue on parallel matrix algorithms and applications (PMAA’18). Parallel Computing, 102 (??):Article 102720, May 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120301058.

Sanders:1999:ANN

[San99]  P. Sanders.  Analysis

of nearest neighbor load balancing algorithms for random loads. Parallel Computing, 25(8): 1013–1033, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/24/22/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/24/22/
article.pdf.

Sanders:2002:RSR

[San02]  P. Sanders.  Reconcil-

ing simplicity and realism in parallel disk models. Parallel Computing, 28(5):705–723, May 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/57/29/
abstract.html.

Sarbazi-Azad:2001:CDH

[SAOKM01] H. Sarbazi-Azad, M. Ould-Khaoua, and L. M. Mackenzie. Communication delay in hypercubes in the presence of bit-reversal traffic. Parallel Computing, 27(13):1801–1816, December 1, 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/43/34/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/43/34/
article.pdf.

Sarbazi-Azad:2006:PEC

[SAOKZ06] H. Sarbazi-Azad, M. Ould-Khaoua, and A. Y. Zomaya. Performance evaluation of communication networks for parallel and distributed systems. Parallel Computing, 32(11–12):775–776, December 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sato:2004:ESR

[Sat04] Tetsuya Sato. The Earth Simulator: roles and impacts. Parallel Computing, 30(12):1279–1286, December 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Seldner:1988:PPS

[SAWH88] D. Seldner, M. Alef, T. Westermann, and E. Halter. Parallel particle simulation in high voltage diodes (algorithms and concepts for implementation on SUPRENUM). Parallel Computing, 7(3): 445–449, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Stark:1992:LDO

[SB92a] S. Stark and A. N. Beris.

LU decomposition optimized for a parallel computer with a hierarchical distributed memory.

Parallel Computing, 18(9): 959–971, September 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Suraweera:1992:PCO

[SB92b] F. Suraweera and P. Bhattacharya. A parallel costoptimal algorithm to compute the supremum of max-min powers. Parallel Computing, 18(5):551– 556, May 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sawley:1994:CSU

[SB94a] M. L. Sawley and C. M. Bergman. A comparative study of the use of the data-parallel approach for compressible flow calculations. Parallel Computing, 20(3):363–373, March 10, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=3&aid=841.

Seguel:1994:FDR

[SB94b] Jaime Seguel and Dorothy Bollman. Fast digitreversal algorithms on a shared-memory machine. Parallel Computing, 20 (1):93–99, January 16, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=1&aid=817.

Shu:1995:PPP

[SB95] Chang Shu and Hilary Buxton. Parallel path planning on the distributed array processor. Parallel Computing, 21 (11):1749–1767, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=11&aid=1024.

Savari:1996:FTA

[SB96] S. A. Savari and D. P.

Bertsekas. Finite termination of asynchronous iterative algorithms. Parallel Computing, 22(1):39– 56, February 20, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cgi-bin/cas/tree/
store/parco/cas_sub/
browse/browse.cgi?year=
1996&volume=22&issue=
1&aid=1038.

Skaalin:1997:IPP

[SB97] R. Skålin and D. Bjørge. Implementation and performance of a parallel version of the HIRLAM limited area atmospheric model. Parallel Computing, 23(14):2161–2172, December 17, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=14&aid=1264.

Scheetz:1998:GDS

[SBM^{+}98] Todd E. Scheetz, Terry A. Braun, Kyle J. Munn, Edwin M. Stone, Val C. Sheffield, and Thomas L. Casavant. GenoMap: a distributed system for unifying genotyping and genetic linkage analysis. Parallel Computing, 24(9– 10):1567–1592, September 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/9-10/1349.pdf.

Spies:2022:THE

[SBM^{+}22] Lukas Spies, Amanda Bienz, David Moulton, Luke Olson, and Andrew Reisner. Tausch: a halo exchange library for large heterogeneous computing systems using MPI, OpenCL, and CUDA. Parallel Computing, 114(??):??, December 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000631.

Sweet:1991:FTD

[SBO^{+}91] Roland A. Sweet, William L.

Briggs, Suely Oliveira, Jules L. Porsche, and Tom Turnbull. FFTs and three-dimensional Poisson solvers for hypercubes. Parallel Computing, 17 (2–3):121–131, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sandrieser:2012:UEP

[SBP12] Martin Sandrieser, Siegfried Benkner, and Sabri Pllana. Using explicit platform descriptions to support programming of heterogeneous many-core systems. Parallel Computing, 38(1–2):52–65, January/ February 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001396.

Seredynski:2004:CAC

[SBZ04]  Franciszek  Seredynski,

Pascal Bouvry, and Albert Y. Zomaya. Cellular automata computations and secret key cryptography. Parallel Computing, 30(5–6):753–766, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shiau:1992:AEM

[SC92] Yuh-Horng Shiau and Chung-Ping Chung. Adoptability and effectiveness of microcode compaction algorithms in superscalar processing. Parallel Computing, 18(5):497–510, May 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sridhar:1995:HPP

[SC95] R. Sridhar and N. Chandrasekharan. Highly parallelizable problems on sorted intervals. Parallel Computing, 21(3):433–446, March 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=3&aid=968.

Shen:2003:DMS

[SC03] Xiaohui Shen and Alok Choudhary. A distributed multi-storage I/O system for data intensive scientific computing. Parallel Computing, 29(11– 12):1623–1643, November/ December 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Srinivasan:2005:LTT

[SC05] A. Srinivasan and N. Chandra. Latency tolerance through parallelization of time in scientific applications. Parallel Computing, 31(7):777–796, July 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shterenlikht:2019:MVF

[SC19] Anton Shterenlikht and Luis Cebamanos. MPI vs Fortran coarrays beyond 100k cores: 3D cellular automata. Parallel Computing, 84(??):37–49, May 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303181.

Sanchez-Curto:2008:FPI

[SCCP08] J. Sánchez-Curto and P. Chamorro-Posada. On a faster parallel implementation of the split-step Fourier method. Parallel Computing, 34(9):539–549, September 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sanchez-Curto:2014:EPI

[SCCPM14] J. Sánchez-Curto, P. Chamorro-Posada, and G. S. McDonald. Efficient parallel implementation of the nonparaxial beam propagation method. Parallel Computing, 40(8):394–407, August 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000660.

Sharp:1992:PAD

[SCD92] D. Sharp, M. Cripps, and J. Darlington. Parallelarchitecture-directed program transformation. Parallel Computing, 18(12): 1363–1380, December 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schwandt:1987:IAM

[Sch87] Hartmut Schwandt. An interval arithmetic method for the solution of nonlinear systems of equations on a vector computer. Parallel Computing, 4(3):323– 337, June 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schaffler:1988:CPU

[Sch88a] G. Schaffler. Connecting PEACE to UNIX. Parallel Computing, 7(3):335–339, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schroder:1988:PDS

[Sch88b] W. Schroder. PEACE: the distributed SUPRENUM operating system. Parallel Computing, 7(3):325–333, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schimmler:1991:PSO

[Sch91a] Manfred Schimmler. Parallel strong orientation on a mesh connected computer. Parallel Computing, 17(6–7):657–664, September 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schwandt:1991:MAP

[Sch91b] H. Schwandt. Memory access problems in block cyclic reduction on vector computers. Parallel Computing, 17(2–3):329– 346, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schuller:1992:PPS

[Sch92] A. Schuller. Parallelizing particle simulations based on the Boltzmann equation. Parallel Computing, 18(3):269–279, March 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schnabel:1995:VLO

[Sch95] Robert B. Schnabel. A view of the limitations, opportunities, and challenges in parallel nonlinear optimization. Parallel Computing, 21(6):875–905, June 12, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=6&aid=983.

Shi:2023:NYS

[SCL^{+}23] Jinliang Shi, Dewu Chen, Jiabi Liang, Lin Li, Yue Lin, and Jianjiang Li. New YARN sharing GPU based on graphics memory granularity scheduling. Parallel Computing, 117(??):??, September 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819123000443.

Shock:1998:DEH

[SCM^{+}98] Carter T. Shock, Chialin Chang, Bongki Moon, Anurag Acharya, Larry Davis, Joel Saltz, and Alan Sussman. Design and evaluation of a highperformance earth science database. Parallel Computing, 24(1):65–89, March 10, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1998&volume=
24&issue=1&aid=1275.

Swirydowicz:2022:LSP

[ŚDJ^{+}22] Kasia Świrydowicz, Eric Darve, Wesley Jones, Jonathan Maack, Shaked Regev, Michael A. Saunders, Stephen J. Thomas, and Slaven Peleš. Linear solvers for power grid optimization problems: a review of GPU-accelerated linear solvers. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121001125.

Strohmaier:1999:MHP

[SDMS99] Erich Strohmaier, Jack J. Dongarra, Hans W. Meuer, and Horst D. Simon. The marketplace of highperformance computing. Parallel Computing, 25 (13–14):1517–1544, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/36/24/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/36/24/
                                        article.pdf.

Strohmaier:2005:RTM

[SDMS05] Erich Strohmaier, Jack J. Dongarra, Hans W. Meuer, and Horst D. Simon. Recent trends in the marketplace of high performance computing. Parallel Computing, 31(3–4):261– 273, March/April 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schmidt:2012:ISB

[SDMS12] Andrew G. Schmidt, Siddhartha Datta, Ashwin A. Mendon, and Ron Sass. Investigation into scaling I/O bound streaming applications productively with an all-FPGA cluster. Parallel Computing, 38(8):344–364, August 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001712.

Sips:1998:ALE

[SDv98] Henk J. Sips, Will Denissen, and Kees van Reeuwijk. Analysis of local enumeration and storage schemes in HPF. Parallel Computing, 24(3–4):355–382, May 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/3-4/1294.pdf.

Stojmenovic:1987:CTP

[SE87] I. Stojmenovic and D. J.

Evans. Comments on two parallel algorithms for the planar convex hull problem. Parallel Computing, 5(3):373–375, November 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Spaletta:1993:PRD

[SE93] G. Spaletta and D. J.

Evans. The Parallel Recursive Decoupling algorithm for solving tridiagonal linear systems. Parallel Computing, 19(5):563– 576, May 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sumiyoshi:1998:PPS

[SE98] K. Sumiyoshi and T. Ebisuzaki. Performance of parallel solution of a blocktridiagonal linear system on Fujitsu VPP500. Parallel Computing, 24 (2):287–304, February 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/2/1284.pdf.

Seager:1986:PCG

[Sea86] M. K. Seager. Parallelizing conjugate gradient for the Cray X-MP. Parallel Computing, 3(1):35–47, March 1986. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Saha:2008:PLD

[SEAH^{+}08] Proshanta Saha, Esam El-Araby, Miaoqing Huang, Mohamed Taher, Sergio Lopez-Buedo, Tarek El-Ghazawi, Chang Shu, Kris Gaj, Alan Michalski, and Duncan Buell. Portable library development for reconfigurable computing systems: a case study. Parallel Computing, 34(4–5): 245–260, May 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sela:1995:WFP

[Sel95]  J. G. Sela.  Weather

forecasting on parallel architectures. Parallel Computing, 21(10): 1639–1654, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=10&aid=1018.

Senechaud:1991:MIB

[Sen91] P. Senechaud. A MIMD implementation of the Buchberger algorithm for Boolean polynomials. Parallel Computing, 17(1):29– 37, April 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sadayappan:1990:CPA

[SER90] P. Sadayappan, F. Ercal, and J. Ramanujam. Cluster partitioning approaches to mapping parallel programs onto a hypercube. Parallel Computing, 13(1):1–16, January 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Seutter:1985:CCP

[Seu85] F. Seutter. CEPROL: a cellular programming language. Parallel Computing, 2(4):327–333, December 1985. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shih:1991:ANN

[SF91] Pao-Hsu Shih and Wu-Shung Feng. An application of neural networks on channel routing problem. Parallel Computing, 17(2–3):229–240, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shephard:1997:PAA

[SFB^{+}97] M. S. Shephard, J. E. Flaherty, C. L. Bottasso, H. L. de Cougny, C. Ozturan, and M. L. Simone. Parallel automatic adaptive analysis. Parallel Computing, 23(9):1327–1347, November 3, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=9&aid=1214.

Shen:2013:ACE

[SFSV13] Jie Shen, Jianbin Fang, Henk Sips, and Ana Lucia Varbanescu. An application-centric evaluation of OpenCL on multicore CPUs. Parallel Computing, 39(12):834–850, December 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001014.

Sun:1991:TBP

[SG91] Xian-He Sun and J. L.

Gustafson. Toward a better parallel performance metric. Parallel Computing, 17(10–11):1093–1109, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sunderam:1999:HPD

[SG99] V. S. Sunderam and G. A.

Geist. Heterogeneous parallel and distributed computing. Parallel Computing, 25(13–14):1699–1721, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/36/31/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/31/
article.pdf.

Schenk:2002:TLD

[SG02a] Olaf Schenk and Klaus Gärtner. Two-level dynamic scheduling in PAR-DISO: Improved scalability on shared memory multiprocessing systems. Parallel Computing, 28(2):187–197, February 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/33/29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/33/29/
main.pdf.

Serot:2002:SPI

[SG02b] Jocelyn Sérot and Dominique Ginhac. Skeletons for parallel image processing: an overview of the SKIPPER project. Parallel Computing, 28 (12):1685–1708, December 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sitaraman:2016:BCR

[SG16] Hariswaran Sitaraman and Ray Grout. Balancing conflicting requirements for grid and particle decomposition in continuum-Lagrangian solvers. Parallel Computing, 52(??):1– 21, February 2016. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0167819115001428.

Sunderam:1994:PCC

[SGDM94] V. S. Sunderam, G. A.

Geist, J. Dongarra, and R. Manchek. The PVM concurrent computing system: Evolution, experiences, and trends. Parallel Computing, 20(4):531–545, March 31, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=4&aid=861.

Sofranac:2022:ADP

[SGP22] Boro Sofranac, Ambros Gleixner, and Sebastian Pokutta. Accelerating domain propagation: an efficient GPU-parallel algorithm over sparse matrices. Parallel Computing, 109(??):??, March 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121001149.

Sekharan:1995:LBM

[SGS95] Chandra N. Sekharan, Vineet Goel, and R. Sridhar. Load balancing methods for ray tracing and binary tree computing using PVM. Parallel Computing, 21(12):1963–1978, December 12, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=12&aid=1028.

Snelling:1988:CSL

[SH88] D. F. Snelling and G.-R. Hoffmann. A comparative study of libraries for parallel processing. Parallel Computing, 8(1–3):255– 266, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Samad:1990:HOH

[SH90] T. Samad and P. Harper. High-order Hopfield and Tank optimization networks. Parallel Computing, 16(2–3):287–292, December 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schonauer:1991:PES

[SH91] W. Schonauer and H. Hafner. Performance estimates for supercomputers: the responsibilities of the manufacturer and of the user. Parallel Computing, 17 (10–11):1131–1149, December 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Swarztrauber:2001:COF

[SH01] Paul N. Swarztrauber and Steven W. Hammond. A comparison of optimal FFTs on torus and hypercube multicomputers. Parallel Computing, 27(6):847–859, May 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/30/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/30/28/
article.pdf.

Shahrabi:2006:PCR

[Sha06] A. Shahrabi. Performance comparison of routing algorithms in wormholeswitched networks. Parallel Computing, 32(11–12): 870–885, December 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Si:2019:IWP

[SHB19] Min Si, Zhiyi Huang, and Pavan Balaji. International workshop on programming models and applications for multicores and manycores (PMAM 2018). Parallel Computing, 88(??):Article 102541, ???? 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301267.

Su:2000:NDP

[SHCD00] Ming-Yang Su, Hui-Ling Huang, Gen-Huey Chen, and Dyi-Rong Duh. Nodedisjoint paths in incomplete WK-recursive networks. Parallel Computing, 26(13–14):1925–1944, December 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/34/33/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/34/33/
article.pdf.

Stacy:1998:HPC

[SHCR98] Mahlon Stacy, Dennis Hanson, Jon Camp, and Richard A. Robb. High performance computing in biomedical imaging research. Parallel Computing, 24(9–10):1287–1321, September 1, 1998. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cas/tree/store/parco/
sub/1998/24/9-10/1336.
pdf.

Shen:1986:FTP

[SHCS86] John P. Shen, John P.

Hayes, Luigi Ciminiera, and Angelo Serra. Faulttolerance and performance analysis of beta-networks. Parallel Computing, 3(3): 231–249, July 1986. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shen:1992:IUS

[She92] Hong Shen. Improved universal k-selection in hypercubes. Parallel Computing, 18(2):177–184, February 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shen:1993:HPI

[She93] Hong Shen. A high performance interconnection network for multiprocessor systems. Parallel Computing, 19(9):993–1001, September 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shen:1995:EPB

[She95] Hong Shen. An efficient permutation-based parallel algorithm for rangejoin in hypercubes. Parallel Computing, 21(2): 303–313, February 17, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=2&aid=935.

Shearer:1996:COF

[She96] M. M. Shearer. Computational optimization of finite difference methods on the CM5. Parallel Computing, 22(3):465–481, April 29, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=3&aid=1060.

Shen:1997:OPM

[She97]  Hong Shen.  Optimal

parallel multiselection on EREW PRAM. Parallel Computing, 23(13): 1987–1992, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=13&aid=1258.

Shadid:1997:EPC

[SHH^{+}97] John Shadid, Scott Hutchinson, Gary Hennigan, Harry Moffat, Karen Devine, and A. G. Salinger. Efficient parallel computation of unstructured finite element reacting flow solutions. Parallel Computing, 23(9):1307–1325, November 3, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=9&aid=1213.

Shih:2001:WRT

[Shi01]  Jau-Der Shih.  Worm-

hole routing for torus networks with faults. Parallel Computing, 27(13): 1817–1829, December 1, 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/43/35/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/43/35/
article.pdf.

Schnorr:2012:HAM

[SHN12] Lucas Mello Schnorr, Guillaume Huard, and Philippe Olivier Alexandre Navaux. A hierarchical aggregation model to achieve visualization scalability in the analysis of parallel applications. Parallel Computing, 38(3):91–110, March 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111001578.

Shoji:2017:LLD

[Sho17] Fumiyoshi Shoji. Lessons learned from development and operation of the K computer. Parallel Computing, 64(??):12–19, May 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117300273.

Salomon:2005:MPA

[SHPA05] M. Salomon, F. Heitz, G.-R. Perrin, and J.- P. Armspach. A massively parallel approach to deformable matching of 3D medical images via stochastic differential equations. Parallel Computing, 31(1):45–71, January 2005. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schimpf:1998:RCM

[SHRN98] Paul Schimpf, Jens Haueisen, Ceon Ramon, and Hannes Nowak. Realistic computer modelling of electric and magnetic fields of human head and torso.

Parallel Computing, 24(9– 10):1433–1460, September 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/9-10/1342.pdf.

Sandroos:2013:MGS

[SHvAP13] A. Sandroos, I. Honkonen, S. von Alfthan, and M. Palmroth. Multi-GPU simulations of Vlasov’s equation using vlasiator. Parallel Computing, 39(8):306–318, August 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000574.

Shyu:1990:PAF

[Shy90] Chii Huah Shyu. A parallel algorithm for finding a maximum weight clique of an interval graph. Parallel Computing, 13(2):253–256, February 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sugimoto:2014:ICL

[SIH14] Yuki Sugimoto, Fumihiko Ino, and Kenichi Hagihara. Improving cache locality for GPU-based volume rendering. Parallel Computing, 40(5–6):59–69, May 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781911400043X.

Silvester:1988:OFE

[Sil88] D. J. Silvester. Optimising finite element matrix calculations using the general technique of element vectorisation. Parallel Computing, 6(2):157– 164, February 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Simmen:1991:CBA

[Sim91] M. Simmen. Comments on broadcast algorithms for two-dimensional grids. Parallel Computing, 17(1): 109–112, April 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Singer:1990:IAN

[Sin90] A. Singer. Implementations of artificial neural networks on the connection machine. Parallel Computing, 14(3):305–315, August 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Staunstrup:1985:PDM

[SJJ85] J. Staunstrup, J. O. Jespersen, and O. V. Johansen. Physical datarepresentation in a multiprocessor database machine. Parallel Computing, 2(4): 335–343, December 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shao:1987:APM

[SK87] Jian Ping Shao and Li Shan Kang. An asynchronous parallel mixed algorithm for linear and nonlinear equations. Parallel Computing, 5(3):313–321, November 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schroder:1991:SCC

[SK91] H. Schröder and E. V. Krishnamurthy. Systolic computation of characteristic polynomials of Hessenberg matrices. Parallel Computing, 17(2–3):273–277, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schattler:1997:PDM

[SK97a] Ulrich Schättler and Elisabeth Krenzien. Parallel ‘Deutschland-Modell’ — a message-passing version for distributed memory computers. Parallel Computing, 23(14):

2215–2226, December 17, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=14&aid=1268.

Shinjo:1997:LPF

[SK97b] Yasushi Shinjo and Yasushi Kiyoki. A lightweight process facility supporting meta-level programming. Parallel Computing, 22(11):1429–1454, January 26, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=11&aid=1096.

Schwarz:2019:SEC

[SK19] Angelika Schwarz and Lars Karlsson. Scalable eigenvector computation for the non-symmetric eigenvalue problem. Parallel Computing, 85(??):131– 140, July 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118302692.

Sheu:1990:GSA

[SKC90] Jang-Ping Sheu, Nan-Ling Kuo, and Gen-Huey Chen. Graph search algorithms and maximum bipartite matching algorithm on the hypercube network model. Parallel Computing, 13(2): 245–251, February 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schade:2022:TES

[SKE^{+}22] Robert Schade, Tobias Kenter, Hossam Elgabarty, Michael Lass, Ole Schütt, Alfio Lazzaro, Hans Pabst, Stephan Mohr, Jürg Hutter, Thomas D. Kühne, and Christian Plessl. Towards electronic structurebased ab-initio molecular dynamics simulations with hundreds of millions of atoms. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000242.

Seinstra:2002:SAU

[SKG02] F. J. Seinstra, D. Koelma, and J. M. Geusebroek. A software architecture for user transparent parallel image processing. Parallel Computing, 28(7–8):967– 993, August 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/58/30/
                                        abstract.html.

Scherl:2012:ESA

[SKH^{+}12] Holger Scherl, Markus Kowarschik, Hannes G. Hofmann, Benjamin Keck, and Joachim Hornegger. Evaluation of state-of-the-art hardware architectures for fast cone-beam CT reconstruction. Parallel Computing, 38(3):111–124, March 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111001359.

Skillicorn:2002:PFS

[Ski02] D. B. Skillicorn. Parallel frequent set counting. Parallel Computing, 28(5):815–825, May 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/57/34/
                                        abstract.html.

Sano:2004:DCS

[SKN04] Kentaro Sano, Yusuke Kobayashi, and Tadao Nakamura. Differential coding scheme for efficient parallel image composition on a PC cluster system. Parallel Computing, 30(2): 285–299, February 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sakai:1995:RIC

[SKS^{+}95] Shuichi Sakai, Yuetsu Kodama, Mitsuhisa Sato, Andrew Shaw, Hiroshi Matsuoka, Hideo Hirono, Kazuaki Okamoto, and Takashi Yokota. Reduced interprocessor-communication architecture and its implementation on EM-4. Parallel Computing, 21(5):753– 769 (or 753–770??), May 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=5&aid=975.

Sachdeva:2008:EVC

[SKST08] Vipin Sachdeva, Michael Kistler, Evan Speight, and Tzy-Hwa Kathy Tzeng. Exploring the viability of the Cell Broadband Engine for bioinformatics applications. Parallel Computing, 34(11):616–626, November 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sakai:1993:DIC

[SKY93] S. Sakai, Y. Kodama, and Y. Yamaguchi. Design and implementation of a circular omega network in the EM-4. Parallel Computing, 19(2):125–142, February 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shyu:1990:SSC

[SL90a] S. J. Shyu and R. C. T. Lee. Solving the set cover problem on a supercomputer. Parallel Computing, 13(3):295–300, March 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shyu:1990:VPP

[SL90b] S. J. Shyu and R. C. T. Lee. The vectorization of the partition problem. Parallel Computing, 16 (2–3):343–350, December 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schulze:2003:PPS

[SL03] Jürgen P. Schulze and Ulrich Lang. The parallelized perspective shearwarp algorithm for volume rendering. Parallel Computing, 29(3):339–354, March 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Strout:2016:ACG

[SLC^{+}16] Michelle Mills Strout, Alan LaMielle, Larry Carter, Jeanne Ferrante, Barbara Kreaseck, and Catherine Olschanowsky. An approach for code generation in the Sparse Polyhedral Framework. Parallel Computing, 53(??):32– 57, April 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116000557.

Shi:2022:EES

[SLG^{+}22] Wenhu Shi, Hongjian Li, Junzhe Guan, Hang Zeng, and Rafe Misskat jahan. Energy-efficient scheduling algorithms based on task clustering in heterogeneous spark clusters. Parallel Computing, 112(??):??, September 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000436.

Servat:2013:FPP

[SLH^{+}13a] Harald Servat, Germán Llort, Kevin Huck, Judit Giménez, and Jesús Labarta. Framework for a productive performance optimization. Parallel Computing, 39(8):336–353, August 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000707.

Su:2013:CET

[SLH^{+}13b] Sen Su, Jian Li, Qingjia Huang, Xiao Huang, Kai Shuang, and Jie Wang. Cost-efficient task scheduling for executing large programs in the cloud. Parallel Computing, 39 (4–5):177–188, April/May 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000355.

Servat:2018:UMA

[SLH^{+}18] Harald Servat, Jesús Labarta, Hans-Christian Hoppe, Judit Giménez, and Antonio J. Peña. Understanding memory access patterns using the BSC performance tools. Parallel Computing, 78(??):1– 14, October 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118301911.

Sloboda:1991:PMC

[Slo91] Fridrich Sloboda. A projection method of the Cimmino type for linear algebraic systems. Parallel Computing, 17(4–5):435– 442, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sener:2021:MMC

[SLS^{+}21] Melih Sener, Stuart Levy, John E. Stone, AJ Christensen, Barry Isralewitz, Robert Patterson, Kalina Borkiewicz, Jeffrey Carpenter, C. Neil Hunter, Zaida Luthey-Schulten, and Donna Cox. Multiscale modeling and cinematic visualization of photosynthetic energy conversion processes from electronic to cell scales. Parallel Computing, 102 (??):Article 102698, May 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300880.

Sheu:1989:SFK

[SlWC89] Jang-Ping Sheu, Chun lien Wu, and Gen-Huey Chen. Selection of the first k largest processes in hypercubes. Parallel Computing, 11(3):381–384, ???? 28, 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shih:1990:PAF

[SLY90] Z. C. Shih, R. C. T. Lee, and S. N. Yang. A parallel algorithm for finding congruent regions. Parallel Computing, 13(2):135–142, February 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Suros:1987:OSN

[SM87] R. Suros and E. Montagne. Optimizing systolic networks by fitting diagonals. Parallel Computing, 4(2):167–174, April 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Stojmenovic:1988:OPA

[SM88] I. Stojmenovic and M. Miyakawa. An optimal parallel algorithm for solving the maximal elements problem in the plane. Parallel Computing, 7(2):249–251, June 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Smieja:1990:GML

[SM90] F. J. Smieja and H. Mühlenbein. The geometry of multilayer perceptron solutions. Parallel Computing, 14 (3):261–275, August 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Selvakumar:1991:EAM

[SM91] S. Selvakumar and C. Siva Ram Murthy. An efficient algorithm for mapping VLSI circuit simulation programs onto multiprocessors. Parallel Computing, 17(9):1009–1016, November 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Saougkos:2013:SAR

[SM13] Dimitris Saougkos and George Manis. Self adaptive run time scheduling for the automatic parallelization of loops with the C2μTC/SL compiler. Parallel Computing, 39(10):603–614, October 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000835.

Slota:2015:PCC

[SM15] George M. Slota and Kamesh Madduri. Parallel color-coding. Parallel Computing, 47(??):51–69, August 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000423.

Srinivasan:2003:TPR

[SMC03] Ashok Srinivasan, Michael Mascagni, and David Ceperley. Testing parallel random number generators. Parallel Computing, 29(1):69–94, January 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sanjuan:2016:ADD

[SMC16] Gemma Sanjuan, Tomàs Margalef, and Ana Cortés. Applying domain decomposition to wind field calculation. Parallel Computing, 57(??):185–196, September 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300461.

Sikora:2015:ORC

[SMJ15] Anna Sikora, Tomàs Margalef, and Josep Jorba. Online root-cause performance analysis of parallel applications. Parallel Computing, 48(??):81–107, October 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000769.

Schroder:1991:SAP

[SMK91] H. Schroder, V. K. Murthy, and E. V. Krishnamurthy. Systolic algorithm for polynomial interpolation and related problems. Parallel Computing, 17(4–5): 493–503, July 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schwarz:2020:RPE

[SMK20] Angelika Schwarz, Carl Christian Kjelgaard Mikkelsen, and Lars Karlsson. Robust parallel eigenvector computation for the non-symmetric eigenvalue problem. Parallel Computing, 100(??):Article 102707, December 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819120300934.

Situ:2014:PLE

[SML^{+}14] Yingchong Situ, Chandra S. Martha, Matthew E. Louis, Zhiyuan Li, Ahmed H. Sameh, Gregory A. Blaisdell, and Anastasios S. Lyrintzis. Petascale large eddy simulation of jet engine noise based on the truncated SPIKE algorithm. Parallel Computing, 40(9):496–511, October 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000829.

Stojcev:1990:AMC

[SMM90] M. K. Stojčev, E. I. Milovanović, and I. Ž. Milovanović. An algorithm for multiplication of concatenated matrices. Parallel Computing, 13(2):211–223, February 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shylo:2011:RSO

[SMP11] Oleg V. Shylo, Timothy Middelkoop, and Panos M. Pardalos. Restart strategies in optimization: parallel and serial cases. Parallel Computing, 37(1):60– 68, January 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sano:2004:EPP

[SMT^{+}04] Kentaro Sano, Shintaro Momose, Hiroyuki Takizawa, Hiroaki Kobayashi, and Tadao Nakamura. Efficient parallel processing of competitive learning algorithms. Parallel Computing, 30(12):1361–1383, December 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Stuben:1996:IPC

[SMTT96] Klaus Stüben, Hermann Mierendorff, Clemens-August Thole, and Owen Thomas. Industrial parallel computing with real codes. Parallel Computing, 22(5):725–737, August 8, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=5&aid=1071.

Strengert:2005:LVV

[SMW^{+}05] M. Strengert, M. Magallón, D. Weiskopf, Stefan Guthe, and T. Ertl. Large volume visualization of compressed timedependent datasets on GPU clusters. Parallel Computing, 31(2):205–219, February 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Saltz:1988:TDR

[SN88] Joel H. Saltz and Vijay K.

Naik. Towards developing robust algorithms for solving partial differential equations on MIMD machines. Parallel Computing, 6(1):19–44, January 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schnorr:2014:BSP

[SN14] Lucas Mello Schnorr and Philippe Olivier Alexandre Navaux. Best of SBAC– PAD 2012. Parallel Computing, 40(9):512–513, October 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000623.

Snelling:1988:SFP

[Sne88] D. F. Snelling. Standard FORTRAN 77 as a parallel language. Parallel Computing, 8(1–3):409– 414, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Satoh:2006:ATL

[SNK06] Makoto Satoh, Kiyoshi Negishi, and Atsushi Kobayashi. Analysis of two-level data mapping in an HPF compiler for distributed-memory machines. Parallel Computing, 32(4):280–300, April 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Suda:1999:HPP

[SNO99] Reiji Suda, Akira Nishida, and Yoshio Oyanagi. A high performance parallelization scheme for the Hessenberg double shift QR algorithm. Parallel Computing, 25(6):729–744, June 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/6/1427.pdf.

Sakata:1997:PEW

[SNS^{+}97] Satoko Sakata, Umpei Nagashima, Mitsuhisa Sato, Satoshi Sekiguchi, and Haruo Hosoya. Performance evaluation of a workstation cluster, TMC CM-5, and Intel Paragon/ XP using a parallel homology analysis program. Parallel Computing, 22 (12):1677–1693, February 21, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=12&aid=1117.

Sak:2007:PCA

[SÖB07] Halis Sak, Süleyman Özekici, and İlkay Boduroğlu. Parallel computing in Asian option pricing. Parallel Computing, 33(2):92–108, March 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sodan:2002:AMA

[Sod02] Angela C. Sodan. Applications on a multithreaded architecture: a case study with EARTH-MANNA. Parallel Computing, 28(1):3–33, January 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/27/28/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/60/27/28/
                                        00001684.pdf.

Schreiber:1994:NEP

[SOH94] Thomas Schreiber, Peter Otto, and Fridolin Hofmann. A new efficient parallelization strategy for the QR algorithm. Parallel Computing, 20(1):63– 75 (or 63–76??), January 16, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=1&aid=815.

Solchenbach:1988:GAD

[Sol88] K. Solchenbach. Grid applications on distributed memory architectures: implementation and evaluation. Parallel Computing, 7(3):341–356, September 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sonnenschein:1986:ELC

[Son86] M. Sonnenschein. An extension of the language C for concurrent programming. Parallel Computing, 3(1):59–71, March 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Song:1992:DTE

[Son92] Jianjian Song. A distributedtermination experiment on a mesh-connected array of processors. Parallel Computing, 18(7):779–791, July 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Song:1994:PAI

[Son94] Jianjian Song. A partially asynchronous and iterative algorithm for distributed load balancing. Parallel Computing, 20(6):853–868, June 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=6&aid=853.

Sorensen:1984:BVP

[Sor84] D. C. Sorensen. Buffering for vector performance on a pipelined MIMD machine. Parallel Computing, 1(2):143–164, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Satofuka:1997:PCS

[SOS97] N. Satofuka, M. Obata, and T. Suzuki. Parallel computation of super-/ hypersonic flows on workstation network and Transputer arrays. Parallel Computing, 23(9): 1293–1305, November 3, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=9&aid=1212.

Struckmeier:1993:ESM

[SP93] J. Struckmeier and F. J. Pfreundt. On the efficiency of simulation methods for the Boltzmann equation on parallel computers. Parallel Computing, 19(1):103– 119, January 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schroder-Preikschat:1994:PSB

[SP94] Wolfgang Schröder-Preikschat.

PEACE — a software backplane for parallel computing. Parallel Computing, 20(10–11): 1471–1485, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=908.

Sharma:1997:EIB

[SP97] Neeraj K. Sharma and Madhusudhana R. Pinnu. An efficient implementation of bypass queue under bursty traffic. Parallel Computing, 23(6):777–781, June 20, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=6&aid=1183.

Schmid:2022:EFS

[SPCB22] Rafael F. Schmid, Flávia Pisani, Edson N. Cáceres, and Edson Borin. An evaluation of fast segmented sorting implementations on GPUs. Parallel Computing, 110(??):??, May 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121001289.

Smith:2018:HFO

[SPK18] Shaden Smith, Jongsoo Park, and George Karypis. HPC formulations of optimization algorithms for tensor completion. Parallel Computing, 74(??):99– 117, ???? 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117301849.

Sky:2023:FSM

[SPMB23] Adam Sky, César Polindara, Ingo Muench, and Carolin Birk. A flexible sparse matrix data format and parallel algorithms for the assembly of finite element matrices on shared memory systems. Parallel Computing, 117(??):??, September 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819123000455.

Shi:2015:PCE

[SPW^{+}15] Xiaohua Shi, Fredrick Park, Lina Wang, Jack Xin, and Yingyong Qi.

Parallelization of a colorentropy preprocessed Chan– Vese model for face contour detection on multicore CPU and GPU. Parallel Computing, 49(??): 28–49, November 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819115001052.

Slone:1997:EBR

[SR97] Dale M. Slone and Garry H.

Rodrigue. Efficient biased random bit generation for parallel lattice gas simulations. Parallel Computing, 22(12):1597–1620, February 21, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=12&aid=1111.

Sudarsan:2010:DPS

[SR10] Rajesh Sudarsan and Calvin J. Ribbens. Design and performance of a scheduling framework for resizable parallel applications. Parallel Computing, 36(1):48–64, January 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Selim:2014:CLI

[SR14] Mohammad Reza Selim and Mohammed Ziaur Rahman. Carrying on the legacy of imperative languages in the future parallel computing era. Parallel Computing, 40(3–4):1– 33, March 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000222.

Speck:2017:TFT

[SR17] Robert Speck and Daniel Ruprecht. Toward faulttolerant parallel-in-time integration with PFASST. Parallel Computing, 62 (??):20–37, February 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0167819116301338.

Santo:2007:PGV

[SRH07] Luis Paulo Santo, Bruno Raffin, and Alan Heirich. Parallel graphics and visualization. Parallel Computing, 33(6):359–360, June 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schneider:2021:NIA

[SRK^{+}21] Evan Schneider, Brant Robertson, Alexander Kuhn, Christopher Lux, and Marc Nienhaus. NVIDIA IndeX accelerated computing for visualizing Cholla’s galactic winds. Parallel Computing, 107(??):??, October 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000594.

Sultana:2019:FRB

[SRS^{+}19] Nawrin Sultana, Martin Rüfenacht, Anthony Skjellum, Ignacio Laguna, and Kathryn Mohror. Failure recovery for bulk synchronous applications with MPI stages. Parallel Computing, 84(??):1–14, May 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118303260.

Srikant:1987:NPA

[SS87] Y. N. Srikant and P. Shankar.

A new parallel algorithm for parsing arithmetic infix expressions. Parallel Computing, 4(3):291–304, June 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schonauer:1988:FBB

[SS88] Willi Schönauer and Eric Schnepf. FIDISOL: a “black box” solver for partial differential equations. Parallel Computing, 6(2): 185–193, February 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Saad:1989:DCP

[SS89a] Youcef Saad and Martin H. Schultz. Data communication in parallel architectures. Parallel Computing, 11(2):131–150, ???? 4, 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schimmler:1989:SSM

[SS89b] Manfred Schimmler and Heiko Schröder. A simple systolic method to find all bridges of an undirected graph. Parallel Computing, 12(1):107–111, October 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Stewart:1990:PMF

[SS90] A. Stewart and G. J. Shaw. A parallel multigrid FAS scheme for transputer networks. Parallel Computing, 16(2–3):335–342, December 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schroder:1991:PCI

[SS91a] H. Schroder and P. Strazdins.

Program compression on the instruction systolic array. Parallel Computing, 17(2–3):207–219, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Storoy:1991:NOS

[SS91b] S. Storoy and T. Sorevik.

A note on an orthogonal systolic design for the assignment problem. Parallel Computing, 17(4–5): 523–525, July 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Strongin:1992:GMO

[SS92] Roman G. Strongin and Yaroslav D. Sergeyev. Global multidimensional optimization on parallel computer. Parallel Computing, 18(11):1259–1273, November 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Selvakumar:1994:STA

[SS94a] S. Selvakumar and C. Siva Ram Murthy. Static task allocation of concurrent programs for distributed computing systems with processor and resource heterogeneity. Parallel Computing, 20(6):835–851, June 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=6&aid=852.

Suman:1994:CSG

[SS94b] H. Suman and K. Schilling. A comparative study of gauge fixing procedures on the connection machines CM2 and CM5. Parallel Computing, 20(7):975–990, July 12, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=7&aid=872.

Saikia:1995:OPC

[SS95] Dilip K. Saikia and Ranjan K. Sen. Order preserving communication on a star network. Parallel Computing, 21(5):771–782, May 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=5&aid=957.

Shi:2001:RSF

[SS01] Wei Shi and Pradip K.

Srimani. A regular scalable fault tolerant interconnection network for distributed processing. Parallel Computing, 27(14): 1897–1919, December 31, 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/44/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/44/30/
article.pdf.

Sameh:2002:PAI

[SS02] Ahmed H. Sameh and Vivek Sarin. Parallel algorithms for indefinite linear systems. Parallel Computing, 28(2):285–299, February 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/33/34/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/33/34/
main.pdf.

Sinnen:2004:LSE

[SS04] Oliver Sinnen and Leonel Sousa. List scheduling: extension for contention awareness and evaluation of node priorities for heterogeneous cluster architectures. Parallel Computing, 30(1):81–101, January 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Swarztrauber:1991:BFA

[SSB^{+}91] Paul N. Swarztrauber, Roland A. Sweet, William L. Briggs, Van Emden Henson, and James Otto. Bluestein’s FFT for arbitrary N on the hypercube. Parallel Computing, 17(6–7):607–617, September 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sathe:2012:ABW

[SSB12] Madan Sathe, Olaf Schenk, and Helmar Burkhart. An auction-based weighted matching implementation on massively parallel architectures. Parallel Computing, 38(12):595–614, December 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819112000750.

Starzec:2019:DAC

[SSBT19] Mateusz Starzec, Grazyna Starzec, Aleksander Byrski, and Wojciech Turek. Distributed ant colony optimization based on actor model. Parallel Computing, 90(??):Article 102573, December 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301644.

Skjellum:1994:DEZ

[SSD^{+}94] Anthony Skjellum, Steven G.

Smith, Nathan E. Doss, Alvin P. Leung, and Manfred Morari. The design and evolution of Zipcode. Parallel Computing, 20(4):565–596, March 31, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=4&aid=857.

Sips:2001:LSA

[SSD01] Henk J. Sips, Ruud Sommerhalder, and Erik D’Hollander. Linear systems and associated problems. Parallel Computing, 27(7):867–868, June 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/31/23/
abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/10/35/21/47/31/23/
                                        article.pdf.

Sosa:2000:IQC

[SSGF00] C. P. Sosa, G. Scalmani, R. Gomperts, and M. J. Frisch. Ab initio quantum chemistry on a cc-NUMA architecture using openMP. III. Parallel Computing, 26(7–8):843– 856, July 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/29/25/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/29/25/
                                        article.pdf.

Schwabe:1997:EAB

[SSH97] Eric J. Schwabe, Ian M. Sutherland, and Bruce K. Holmer. Evaluating approximately balanced parity-declustered data layouts for disk arrays. Parallel Computing, 23 (4–5):501–523, May 23, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=4-5&aid=1169.

Sariyuce:2015:ICC

[SSKÇ15] Ahmet Erdem Sariyüce, Erik Saule, Kamer Kaya, and Ümit V. Çatalyürek. Incremental closeness centrality in distributed memory. Parallel Computing, 47(??):3–18, August 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000101.

Sim:2003:MSH

[SSL03] Leo Chin Sim, Heiko Schroder, and Graham Leedham. MIMD–SIMD hybrid system—-towards a new low cost parallel system. Parallel Computing, 29(1):21–36, January 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schreiber:2019:EIP

[SSL19] Martin Schreiber, Nathanaël Schaeffer, and Richard Loft. Exponential integrators with parallelin-time rational approximations for the shallowwater equations on the rotating sphere. Parallel Computing, 85(??):56–65, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300620.

Schoneveld:2000:FDL

[SSLK00] Arjen Schoneveld, Peter M. A. Sloot, Martin Lees, and Erwan Karyadi. A framework for dynamic load balancing: a case study on explosive containment simulation. Parallel Computing, 26(6):737– 751, May 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/28/26/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/28/26/
                                        article.pdf.

Sousa:2004:PSB

[SSN04] Tiago Sousa, Arlindo Silva, and Ana Neves. Particle swarm-based data mining algorithms for classification tasks. Parallel Computing, 30(5–6):767– 783, May/June 2004. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schuchart:2021:CBC

[SSN^{+}21] Joseph Schuchart, Philipp Samfass, Christoph Niethammer, José Gracia, and George Bosilca. Callback-based completion notification using MPI continuations. Parallel Computing, 106(??):??, September 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000466.

Sheng:2014:CIE

[SSO^{+}14] Weihua Sheng, Stefan Schürmans, Maximilian Odendahl, Mark Bertsch, Vitaliy Volevach, Rainer Leupers, and Gerd Ascheid. A compiler infrastructure for embedded heterogeneous MP-SoCs. Parallel Computing, 40(2):51–68, February 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001452.

Shan:2003:MPS

[SSOB03] Hongzhang Shan, Jaswinder P.

Singh, Leonid Oliker, and Rupak Biswas. Message passing and shared address space parallelism on an SMP cluster. Parallel Computing, 29(2):167–186, February 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sheehan:1998:LSC

[SSP^{+}98] Timothy J. Sheehan, William A.

Shelton, Thomas J. Pratt, Philip M. Papadopoulos, Philip LoCascio, and Thomas H. Dunigan.

The locally self-consistent multiple scattering code in a geographically distributed linked MPP environment. Parallel Computing, 24(12–13):1827–1846, November 1, 1998. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cas/tree/store/parco/
sub/1998/24/12-13/1357.
pdf.

Schreiber:1992:PMD

[SSS92] H. Schreiber, O. Steinhauser, and P. Schuster. Parallel molecular dynamics of biomolecules. Parallel Computing, 18(5):557– 573, May 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sen:1999:PBD

[SSS99] Vikramaditya Sen, Mrinal K. Sen, and Paul L. Stoffa. PVM based 3-D Kirchhoff depth migration using dynamically computed travel-times: an application in seismic data processing. Parallel Computing, 25(3):231–248, March 22, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/3/1389.pdf.

Shivle:2005:MSM

[SSS^{+}05] S. Shivle, P. Sugavanam, H. J. Siegel, A. A. Maciejewski, T. Banka, K. Chindam, S. Dussinger, A. Kutruff, P. Penumarthy, P. Pichumani, P. Satyasekaran, D. Sendek, J. Smith, J. Sousa, J. Sridharan, and J. Velazco. Mapping subtasks with multiple versions on an ad hoc grid. Parallel Computing, 31(7):671–690, July 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Smith:2009:RRA

[SSS^{+}09] Jay Smith, Vladimir Shestak, Howard Jay Siegel, Suzy Price, Larry Teklits, and Prasanna Sugavanam. Robust resource allocation in a cluster based imaging system. Parallel Computing, 35(7):389–400, July 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sambavaram:2003:MBP

[SSSG03] Sreekanth R. Sambavaram, Vivek Sarin, Ahmed Sameh, and Ananth Grama. Multipolebased preconditioners for large sparse linear systems. Parallel Computing, 29(9):1261–1273, September 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sanders:2009:TTA

[SST09] Peter Sanders, Jochen Speck, and Jesper Larsson Träff. Two-tree algorithms for full bandwidth broadcast, reduction and scan. Parallel Computing, 35(12):581–594, December 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Stone:2016:ADV

[SSV^{+}16] John E. Stone, Melih Sener, Kirby L. Vandivort, Angela Barragan, Abhishek Singharoy, Ivan Teo, João V. Ribeiro, Barry Isralewitz, Bo Liu, Boon Chong Goh, James C. Phillips, Craig MacGregor-Chatwin, Matthew P. Johnson, Lena F. Kourkoutis, C. Neil Hunter, and Klaus Schulten. Atomic detail visualization of photosynthetic membranes with GPU-accelerated ray tracing. Parallel Computing, 55(??):17–27, July 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115001593.

Sanches:2002:SCC

[SSY02] Carlos Alberto Alonso Sanches, Nei Yoshihiro Soma, and Horacio Hideki Yanasse. Short communication: Comments on parallel algorithms for the knapsack problem. Parallel Computing, 28 (10):1501–1505, October ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/60/32/
abstract.html.

Sanches:2010:OOP

[SSY10] Carlos Alberto Alonso Sanches, Nei Yoshihiro Soma, and Horacio Hideki Yanasse. Observations on optimal parallelizations of two-list. Parallel Computing, 36(1):65–67, January 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Solchenbach:1988:SSE

[ST88] K. Solchenbach and U. Trottenberg. SUPRENUM: system essentials and grid applications. Parallel Computing, 7(3):265–281, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Smith:1995:PIC

[ST95] T. H. C. Smith and G. L.

Thompson. A parallel implementation of the column subtraction algorithm. Parallel Computing, 21(1):63–71, January 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=1&aid=938.

Salinger:2002:OBG

[ST02] Petr Salinger and Pavel Tvrdík. Optimal broadcasting and gossiping in one-port meshes of trees with distance-insensitive routing. Parallel Computing, 28(4):627–647, April 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/39/30/
abstract.html.

Shirvani:2021:NHH

[ST21] Mirsaeid Hosseini Shirvani and Reza Noorian Talouki. A novel hybrid heuristicbased list scheduling algorithm in heterogeneous cloud computing environment for makespan optimization. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000764.

Stewart:1987:PIA

[Ste87] G. W. Stewart. A parallel implementation of the QR-algorithm. Parallel Computing, 5(1–2):187– 196, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Steffen:1988:IRC

[Ste88] B. Steffen. Implementation of a resonant cavity package on MIMD computers. Parallel Computing, 7(1): 55–63, April 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Stewart:1990:CMC

[Ste90] G. W. Stewart. Communication and matrix computations on large message passing systems. Parallel Computing, 16(1):27– 40, November 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Stewart:1994:UUD

[Ste94] G. W. Stewart. Updating URV decompositions in parallel. Parallel Computing, 20(2):151–172, February 24, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=2&aid=828.

Steinfadt:2013:FGP

[Ste13] Shannon Steinfadt. Finegrained parallel implementations for SWAMP+ Smith–Waterman alignment. Parallel Computing, 39(12):819–833, December 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001002.

Simic:2023:BDB

[STE23] Srdan Daniel Simić, Nikola Tanković, and Darko Etinger. Big data BPMN workflow resource optimization in the cloud. Parallel Computing, 117(??): ??, September 2023. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819123000315.

Surridge:1996:PAE

[STKA96] M. Surridge, D. J. Tildesley, Y. C. Kong, and D. B. Adolf. Practical aspects and experiences.

A parallel molecular dynamics simulation code for dialkyl cationic surfactants. Parallel Computing, 22(8):1053–1071, October 28, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=8&aid=1093.

Storoy:1989:HAP

[Sto89] S. Storoy. Holistic algorithms: a paradigm for multiprocessor programming. Parallel Computing, 10(2):221–229, April 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Stpiczynski:1992:PCF

[Stp92] Przemysław Stpiczyński.

Parallel Cholesky factorization on orthogonal multiprocessors. Parallel Computing, 18(2):213–219, February 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Stpiczynski:1993:EAT

[Stp93] Przemysław Stpiczyński.

Error analysis of two parallel algorithms for solving linear recurrence systems. Parallel Computing, 19(8):917–923, August 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sala:2019:IBN

[STP^{+}19] Kevin Sala, Xavier Teruel, Josep M. Perez, Antonio J. Peña, Vicenç Beltran, and Jesus Labarta. Integrating blocking and non-blocking MPI primitives with task-based programming models. Parallel Computing, 85(??):153– 166, July 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118303326.

Strakos:1987:EOA

[Str87] Z. Strakos. Effectivity and optimizing of algorithms and programs on the host-computer/array-processor system. Parallel Computing, 4(2):189–207, April 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Strauss:1992:PC

[Str92]  H. Strauss.  Paral-

lel CFD’92. Parallel Computing, 18(9):1073–??, September 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Straubhaar:2008:PPC

[Str08] Julien Straubhaar. Parallel preconditioners for the conjugate gradient algorithm using Gram– Schmidt and least squares methods. Parallel Computing, 34(10):551–569, October 2008. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Solchenbach:1994:GAS

[STT94] Karl Solchenbach, Clemens-August Thole, and Ulrich Trottenberg. GENESIS application software. Parallel Computing, 20(10– 11):1669–1673, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=921.

Sugimoto:1999:GPC

[Sug99] D. Sugimoto. GRAPE:

a parallel computer dedicated to astrophysical many-body problems. Parallel Computing, 25(13– 14):1663–1676, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/36/29/

                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/36/29/
                                        article.pdf.

Sun:1995:AAP

[Sun95] Xian-He Sun. Application and accuracy of the parallel diagonal dominant algorithm. Parallel Computing, 21(8):1241–1267, August 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=8&aid=997.

Sun:1997:PSS

[Sun97a] Chunguang Sun. Parallel solution of sparse linear least squares problems on distributed-memory multiprocessors. Parallel Computing, 23(13): 2075–2093, December 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=13&aid=1222.

Sunderam:1997:HNC

[Sun97b] Vaidy Sunderam. Heterogeneous network computing: The next generation. Parallel Computing, 23(1–2):121–135, April 16, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=1-2&aid=1151.

Surkov:2010:POP

[Sur10] Vladimir Surkov. Parallel option pricing with Fourier space time-stepping method on graphics processing units. Parallel Computing, 36(7):372–380, July 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sutner:1997:LCA

[Sut97] Klaus Sutner. Linear cellular automata and Fischer automata. Parallel Computing, 23(11): 1613–1634, December 1, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=11&aid=1238.

Schmidt-Voigt:1994:EPC

[SV94] M. Schmidt-Voigt. Efficient parallel communication with the nCUBE 2S processor. Parallel Computing, 20(4):509–530, March 31, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=4&aid=867.

Singh:1997:IPA

[SV97] Ajit Singh and Vincent Van Dongen. An integrated performance analysis tool for SPMD data-parallel programs. Parallel Computing, 23 (8):1089–1112, July 25, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=8&aid=1201.

Scherson:2007:SAR

[SVC07] Isaac D. Scherson, Daniel S. Valencia, and Enrique Cauich. Service address routing: a networkembedded resource management layer for cluster computing. Parallel Computing, 33(7–8):561–571, August 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Si:2019:PPM

[SVC19] Min Si, Abhinav Vishnu, and Yong Chen. Parallel programming models and systems software for high-end computing (P2S2 2018). Parallel Computing, 89(??):Article 102549, November 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301401.

Serra:2001:CGN

[SVS01] Roberto Serra, Marco Villani, and Anna Salvemini. Continuous genetic networks. Parallel Computing, 27(5):663– 683, April 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/47/29/31/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/29/31/
article.pdf.

Serra:2002:ECG

[SVS02] Roberto Serra, Marco Villani, and Anna Salvemini. Erratum to “Continuous genetic networks” [Parallel Comput. 27(5) (2001) 663–683]. Parallel Computing, 28(4):667, April 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-

                                        ng/10/35/21/60/39/32/
                                        abstract.html;   http:
                                        //www.elsevier.com/PII/
                                        S0167819100000831.

Solis-Vasquez:2022:BPI

[SVTSM^{+}22] Leonardo Solis-Vasquez, Andreas F. Tillack, Diogo Santos-Martins, Andreas Koch, Scott LeGrand, and Stefano Forli. Benchmarking the performance of irregular computations in AutoDock–GPU molecular docking. Parallel Computing, 109(??):??, March 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121001046.

Saunders:1991:SPC

[SW91] V. R. Saunders and S. Wilson. “scavenger” programming for the CRAY X-MP computer (short communication). Parallel Computing, 17(9):1025–1034, November 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sui:2000:FTR

[SW00] Pao-Hwa Sui and Sheng-De Wang. A faulttolerant routing algorithm for wormhole routed meshes. Parallel Computing, 26(4):455–465, March 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/26/26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/26/26/
article.pdf.

Sun:2003:SIS

[SW03] Yudong Sun and Cho-Li Wang. Solving irregularly structured problems based on distributed object model. Parallel Computing, 29(11– 12):1539–1562, November/ December 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Swarztrauber:1984:FAV

[Swa84] P. N. Swarztrauber. FFT algorithms for vector computers. Parallel Computing, 1(1):45–63, August 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Swarztrauber:1987:MF

[Swa87] Paul N. Swarztrauber.

Multiprocessor FFTs. Parallel Computing, 5(1–2): 197–210, July 1987. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Szularz:1999:ERL

[SWC99] M. Szularz, J. Weston, and M. Clint. Explicitly restarted Lanczos algorithms in an MPP environment. Parallel Computing, 25(5):613–631, May 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1999/
                                        25/5/1397.pdf.

Sha:2019:TCE

[SWCBQ19] Shi Sha, Wujie Wen, Gustavo A. Chaparro-Baquero, and Gang Quan. Thermal-constrained energy efficient real-time scheduling on multi-core platforms. Parallel Computing, 85(??):231–242, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118300280.

Sosonkina:2013:APC

[SWR^{+}13] Masha Sosonkina, Layne T. Watson, Nicholas R. Radcliffe, Rafael T. Haftka, and Michael W. Trosset. Adjusting process count on demand for petascale global optimization. Parallel Computing, 39(1):21– 35, January 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000890.

Sullivan:2018:SCC

[SWR^{+}18] James Sullivan, Collin Weir, Austin Reichert, R. Todd Evans, W. Cyrus Proctor, and Nicolas Thorne. Student cluster competition 2017, Team University of Texas at Austin/Texas State University: Reproducing vectorization of the Tersoff multi-body potential on the Intel Skylake and NVIDIA V100 architectures. Parallel Computing, 79(??):30–35, November 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300681.

Spee:1992:ECB

[SWSG92] P. Spee, W. F. Wong, M. Sato, and E. Goto. Evaluation of the continuation bit in the Cyclic Pipeline Computer. Parallel Computing, 18(12): 1349–1361, December 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schwichtenberg:1999:AMM

[SWW99] H. Schwichtenberg, G. Winter, and H. Wallmeier. Acceleration of molecular mechanic simulation by parallelization and fast multipole techniques. Parallel Computing, 25(5):535– 546, May 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1999/
                                        25/5/1406.pdf.

Shi:2017:CST

[SWYM17] Lizhen Shi, Zhong Wang, Weikuan Yu, and Xiandong Meng. A case study of tuning MapReduce for efficient bioinformatics in the cloud. Parallel Computing, 61(??):83–95, January 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116301065.

Sathye:1997:PWM

[SXBD97] A. Sathye, M. Xue, G. Bassett, and K. Droegemeier. Parallel weather modeling with the advanced regional prediction system. Parallel Computing, 23 (14):2243–2256, December 17, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=14&aid=1270.

Simpson:1987:OEM

[SY87] R. B. Simpson and A. Yazici.

An organization of the extrapolation method for vector processing. Parallel Computing, 4(2):175–188, April 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shioya:2021:BRB

[SY21] Akemi Shioya and Yusaku Yamamoto. Block redblack MILU(0) preconditioner with relaxation on GPU. Parallel Computing, 103(??):??, June 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000181.

Sahin:2007:FDA

[SYAU07] Ferat Sahin, M. Çetin Yavuz, Ziya Arnavut, and Önder Uluyol. Fault diagnosis for airplane engines using Bayesian networks and distributed particle swarm optimization.

Parallel Computing, 33(2): 124–143, March 2007. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sykora:1984:VSS

[Sýk84] Ondrej Sýkora. VLSI systems for some problems of computational geometry. Parallel Computing, 1(3–4):337–342, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shen:2013:GEI

[SYP13] Fangyang Shen, Mei Yang, and Maurizio Palesi. Guest editors’ introduction to the special issue on “Novel On-Chip Parallel Architectures and Software Support”. Parallel Computing, 39(9):355–356, September 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000847.

Sun:2022:QAD

[SYY^{+}22] Qingxiao Sun, Liu Yi, Hailong Yang, Mingzhen Li, Zhongzhi Luan, and Depei Qian. QoS-aware dynamic resource allocation with improved utilization and energy efficiency on GPU. Parallel Computing, 113(??):??, October 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000503.

Shen:2002:PTL

[SZ02] Chi Shen and Jun Zhang.

Parallel two level block ILU preconditioning techniques for solving large sparse linear systems. Parallel Computing, 28(10): 1451–1475, October ??, 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/60/30/
abstract.html.

Shimasaki:2004:ES

[SZ04] Masaaki Shimasaki and Hans P. Zima. The Earth Simulator. Parallel Computing, 30(12):1277–1278, December 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shen:2022:TPT

[SZS22] Zhongyu Shen, Jilin Zhang, and Tomohiro Suzuki. Task-parallel tiled direct solver for dense symmetric indefinite systems. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000072.

Turan:2014:LSM

[TA14] Erhan Turan and Peter Arbenz. Large scale micro finite element analysis of 3D bone poroelasticity. Parallel Computing, 40(7):239– 250, July 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001063.

Tournier:2019:MTI

[TAB^{+}19] P.-H. Tournier, I. Aliferis, M. Bonazzoli, M. de Buhan, M. Darbas, V. Dolean, F. Hecht, P. Jolivet, I. El Kanfoud, C. Migliaccio, F. Nataf, Ch. Pichot, and S. Semenov. Microwave tomographic imaging of cerebrovascular accidents by using high-performance computing. Parallel Computing, 85(??):88–97, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301042.

Taft:2001:AGP

[Taf01] James R. Taft. Achieving 60 GFLOP/s on the production CFD code OVERFLOW-MLP. Parallel Computing, 27(4): 521–536, March 2001. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/47/
28/32/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/47/
28/32/article.pdf.

Taillard:1991:RTS

[Tai91] E. Taillard. Robust taboo search for the quadratic assignment problem. Parallel Computing, 17(4–5): 443–455, July 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Takahashi:2003:PDF

[Tak03] Daisuke Takahashi. A parallel 1-D FFT algorithm for the Hitachi SR8000. Parallel Computing, 29(6): 679–690, June 2003. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Takesue:2006:PCB

[Tak06] Masaru Takesue. The psicube: a bus-based cubetype clustering network for high-performance on-chip systems. Parallel Computing, 32(11–12):852–869, December 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Takahashi:2010:PIM

[Tak10] Daisuke Takahashi. Parallel implementation of multiple-precision arithmetic and 2, 576, 980, 370, 000 decimal digits of π calculation. Parallel Computing, 36(8):439–448, August 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Takahashi:2018:CQH

[Tak18] Daisuke Takahashi. Computation of the 100 quadrillionth hexadecimal digit of π on a cluster of Intel Xeon Phi processors. Parallel Computing, 75(??):1–10, July 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118300334.

Tappe:1984:APC

[Tap84a] J. Tappe. Algorithms for pipeline control. Parallel Computing, 1(2):185–188, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tappe:1984:MAL

[Tap84b] J. Tappe. The minimal average latency of multiconfigurable pipelines. Parallel Computing, 1(2):181– 183, December 1984. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Teuliere:2003:PPF

[TB03] V. Teulière and Olivier Brun. Parallelisation of the particle filtering technique and application to Doppler-bearing tracking of maneuvering sources. Parallel Computing, 29(8): 1069–1090, August 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Thomaszewski:2007:PBS

[TB07] Bernhard Thomaszewski and Wolfgang Blochinger. Physically based simulation of cloth on distributed memory architectures. Parallel Computing, 33(6):377–390, June 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Timcheck:2022:RQI

[TB22] Stephen Timcheck and Jeremy Buhler. Reducing queuing impact in streaming applications with irregular dataflow. Parallel Computing, 109(??):??, March 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819121001058.

Taufer:2016:SIC

[TBM16a] Michela Taufer, Pavan Balaji, and Satoshi Matsuoka. Special issue on cluster computing. Parallel Computing, 58(??):25– 26, October 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300837.

Tsugane:2016:HVP

[TBM^{+}16b] Keisuke Tsugane, Taisuke Boku, Hitoshi Murai, Mitsuhisa Sato, William Tang, and Bei Wang. Hybridview programming of nuclear fusion simulation code in the PGAS parallel programming language XcalableMP. Parallel Computing, 57(??):37–51, September 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300552.

Thompson:1992:PAM

[TCL92] C. P. Thompson, W. R. Cowell, and G. K. Leaf. On the parallelization of an adaptive multigrid algorithm for a class of flow problems. Parallel Computing, 18(4):449–466, April 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Toulouse:2004:SBC

[TCS04] Michel Toulouse, Teodor Gabriel Crainic, and Brunilde Sansó. Systemic behavior of cooperative search algorithms. Parallel Computing, 30(1):57–79, January 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Toulouse:2000:GOP

[TCT00] Michel Toulouse, Teodor Gabriel Crainic, and K. Thulasiraman. Global optimization properties of parallel cooperative search algorithms: a simulation study. Parallel Computing, 26(1):91–112, January 2000. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/42/23/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/23/28/
article.pdf.

Tomov:2010:TDL

[TDB10] Stanimire Tomov, Jack Dongarra, and Marc Baboulin. Towards dense linear algebra for hybrid GPU accelerated manycore systems. Parallel Computing, 36(5–6):232–240, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tavakoli:2019:CSS

[TDC19] Neda Tavakoli, Dong Dai, and Yong Chen. Clientside straggler-aware I/O scheduler for object-based parallel file systems. Parallel Computing, 82(??):3– 18, ???? 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118302138.

Tan:2018:UVQ

[TDG^{+}18] Li Tan, Nathan De-Bardeleben, Qiang Guan, Sean Blanchard, and Michael Lang. Using virtualization to quantify power conservation via near-threshold voltage reduction for inherently resilient applications. Parallel Computing, 74(??):3– 15, ???? 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117300996.

Tseng:2003:DST

[TDW03] Y. Tseng, R. F. De-Mara, and P. J. Wilder.

Distributed-sum termination detection supporting multithreaded execution. Parallel Computing, 29(7): 953–968, July 2003. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tomov:2004:ART

[TDW^{+}04] N. Tomov, E. Dempster, M. H. Williams, A. Burger, H. Taylor, P. J. B. King, and P. Broughton. Analytical response time estimation in parallel relational database systems. Parallel Computing, 30(2):249–283, February 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tout:1992:PFC

[TE92] K. R. Tout and D. J.

Evans. Parallel forward chaining technique with dynamic scheduling, for rule-based expert systems. Parallel Computing, 18 (8):913–930, August 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Temperton:1988:IPF

[Tem88] Clive Temperton. Implementation of a prime factor FFT algorithm on CRAY-1. Parallel Computing, 6 (1):99–108, January 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Temperton:1989:FMC

[Tem89]  C. Temperton.  Fur-

ther measurements of (r_{∞}, n_{1/2}) on the CRAY-1 and CRAY X-MP. Parallel Computing, 11(1):107– 111, July 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Terekhov:2010:PDA

[Ter10] Andrew V. Terekhov. Parallel Dichotomy Algorithm for solving tridiagonal system of linear equations with multiple right-hand sides. Parallel Computing, 36(8):423–438, August 2010. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Terekhov:2013:FPA

[Ter13] Andrew V. Terekhov. A fast parallel algorithm for solving block-tridiagonal systems of linear equations including the domain decomposition method. Parallel Computing, 39 (6–7):245–258, June/July 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000367.

Triantafillou:1998:OSO

[TF98] Peter Triantafillou and Christos Faloutsos. Overlay striping and optimal parallel I/O for modern applications. Parallel Computing, 24(1):21–43, March 10, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1998&volume=
24&issue=1&aid=1273.

Tremblay:2001:TSP

[TF01] N. Tremblay and M. Florian. Temporal shortest paths: Parallel computing implementations. Parallel Computing, 27 (12):1569–1609, November 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/42/30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/42/30/
article.pdf.

Topkar:1991:DRH

[TFS91] V. Topkar, O. Frieder, and A. K. Sood. Duplicate removal on hypercube engines: an experimental analysis. Parallel Computing, 17(8):845–871, October 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tumeo:2016:SIT

[TFV16] Antonino Tumeo, John Feo, and Oreste Villa. Special issue on Theory and Practice of Irregular Applications (TaPIA). Parallel Computing, 59(??):21–23, November 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116301089.

Thakur:2009:TSE

[TG09] Rajeev Thakur and William Gropp. Test suite for evaluating performance of multithreaded MPI communication. Parallel Computing, 35(12):608–617, December 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tasic:1992:SIP

[TGE92] J. Tasic, M. Gusev, and D. J. Evans. Systolic implementation of preconditioned conjugate gradient method in adaptive transversal filters. Parallel Computing, 18(9):1053– 1065, September 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Thakur:2002:ONA

[TGL02] Rajeev Thakur, William Gropp, and Ewing Lusk. Optimizing noncontiguous accesses in MPI-IO. Parallel Computing, 28(1):83–105, January 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/27/32/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/60/27/32/
00001686.pdf.

Tsay:1989:GFE

[TH89] Jong-Chuang Tsay and Yodung-Chang Hou. Generating function and equivalent transformation for systolic arrays. Parallel Computing, 10(3):347–356, May 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Traff:2020:SIS

[TH20] Jesper Larsson Träff and Torsten Hoefler. Special issue: Selected papers from EuroMPI 2019. Parallel Computing, 99(??): Article 102695, November 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0167819120300855.

Talbi:1998:PAT

[THG98] E. G. Talbi, Z. Hafidi, and J-M. Geib. A parallel adaptive tabu search approach. Parallel Computing, 24(14):2003–2019, December 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/14/1363.pdf.

Tian:2005:CEN

[THH^{+}05] Xinmin Tian, Jay P. Hoeflinger, Grant Haab, Yen-Kuang Chen, Milind Girkar, and Sanjiv Shah. A compiler for exploiting nested parallelism in OpenMP programs. Parallel Computing, 31 (10–12):960–983, October/ December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tsai:1999:FDM

[THK^{+}99] Horng-Ren Tsai, Shi-Jinn Horng, Tzong-Wann Kao, Shung-Shing Lee, and Shun-Shan Tsai. Fundamental data movement operations and its applications on a hyperbus broadcast network. Parallel Computing, 25 (2):137–157, February 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/2/1381.pdf.

Totoni:2014:SAP

[THK14] Ehsan Totoni, Michael T.

Heath, and Laxmikant V. Kale. Structure-adaptive parallel solution of sparse triangular linear systems. Parallel Computing, 40(9): 454–470, October 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819114000799.

Trefftz:1995:SES

[THM^{+}95] C. Trefftz, C. C. Huang, P. K. McKinley, T.-Y. Li, and Z. Zeng. A scalable eigenvalue solver for symmetric tridiagonal matrices. Parallel Computing, 21(8):1213–1240, August 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=8&aid=998.

Traff:2021:MCC

[THMH21] Jesper Larsson Träff, Sascha Hunold, Guillaume Mercier, and Daniel J. Holmes. MPI collective communication through a single set of interfaces: a case for orthogonality. Parallel Computing, 107(??):??, October 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000740.

Thune:1990:PSE

[Thu90] Michael Thuné. A partitioning strategy for explicit difference methods. Parallel Computing, 15 (1–3):147–154, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Thune:1991:SPC

[Thu91] Michael Thuné. Straightforward partitioning of composite grids for explicit difference methods. Parallel Computing, 17(6–7): 665–672, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Thune:1992:PPC

[Thu92]  M. Thune.  The par-

titioning problem for a class of data parallel algorithms. Parallel Computing, 18(8):867–878, August 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Thulasiram:2010:P

[Thu10] Ruppa K. Thulasiram.

Preface. Parallel Computing, 36(7):371, July 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Trobec:1997:LDM

[TJ97] Roman Trobec and Izidor Jerebic. Local diagnosis in massively parallel systems. Parallel Computing, 23(6):721–731, June 20, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=6&aid=1181.

Trobec:1993:PAM

[TJJ93] R. Trobec, I. Jerebic, and D. Janezic. Parallel algorithm for molecular dynamics integration. Parallel Computing, 19(9):1029– 1039, September 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tan:2014:SPE

[TKC^{+}14] Li Tan, Shashank Kothapalli, Longxiang Chen, Omar Hussaini, Ryan Bissiri, and Zizhong Chen. A survey of power and energy efficient techniques for high performance numerical linear algebra operations. Parallel Computing, 40(10):559–573, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001112.

Tezduyar:1997:PCM

[TKG97] T. Tezduyar, V. Kalro, and W. Garrard. Parallel computational methods for 3D simulation of a parafoil with prescribed shape changes. Parallel Computing, 23 (9):1349–1363, November 3, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=9&aid=1215.

Tamura:1985:FVS

[TKI85] Hiroshi Tamura, Sachio Kamiya, and Takahiro Ishigai. FACOM VP-100/ 200: supercomputers with ease of use. Parallel Computing, 2(2):87–107, June 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tangpongprasit:2005:TLB

[TKK^{+}05] Sanya Tangpongprasit, Takahiro Katagiri, Kenji Kise, Hiroki Honda, and Toshitsugu Yuba. A timeto-live based reservation algorithm on fully decentralized resource discovery in Grid computing. Parallel Computing, 31(6):529– 543, June 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tsay:1990:SDG

[TL90] J. C. Tsay and C. J. Lin.

A systolic design for generating combinations in lexicographic order. Parallel Computing, 13(1):119–125, January 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tsay:1994:OPA

[TL94] Jong-Chuang Tsay and Wei-Ping Lee. An optimal parallel algorithm for generating permutations in minimal change order. Parallel Computing, 20(3):353–361, March 10, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=3&aid=845.

Tremolet:1996:PAV

[TL96] Y. Trémolet and F.-X. Le Dimet. Parallel algorithms for variational data assimilation and coupling models. Parallel Computing, 22(5):657–674, August 8, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=5&aid=1069.

Tong:2021:APS

[TLC^{+}21] Qianqian Tong, Guannan Liang, Xingyu Cai, Chunjiang Zhu, and Jinbo Bi. Asynchronous parallel stochastic quasi-Newton methods. Parallel Computing, 101(??): Article 102721, April 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781912030106X.

Trotter:2023:TPU

[TLC23] James D. Trotter, Johannes Langguth, and Xing Cai. Targeting performance and userfriendliness: GPU-accelerated finite element computation with automated code generation in FEniCS. Parallel Computing, 118(??):??, November 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000571.

Taufer:2008:RVL

[TLS^{+}08] Michela Taufer, Ming-Ying Leung, Thamar Solorio, Abel Licon, David Mireles, Roberto Araiza, and Kyle L. Johnson. RNAVLab: a virtual laboratory for studying RNA secondary structures based on grid computing technology. Parallel Computing, 34(11):661–680, November 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tran:2016:BPA

[TLS16] Tuan Tu Tran, Yongchao Liu, and Bertil Schmidt. Bit-parallel approximate pattern matching: Kepler GPU versus Xeon Phi. Parallel Computing, 54(??):128–138, May 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001477.

Theoharis:1990:IMM

[TM90] T. Theoharis and J. J.

Modi. Implementation of matrix multiplication on the T-RACK. Parallel Computing, 14(2):229–233, June 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tucker:1994:CAM

[TM94] Lewis W. Tucker and Alan Mainwaring. CMMD: Active messages on the CM-5. Parallel Computing, 20(4):481–496, March 31, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=4&aid=856.

Touati:2009:PRS

[TM09] Sid-Ahmed-Ali Touati and Zsolt Mathe. Periodic register saturation in innermost loops. Parallel Computing, 35(4):239–254, April 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tarkov:2002:MAF

[TMCC02] Mikhail S. Tarkov, Youngsong Mun, Jaeyoung Choi, and Hyung-Il Choi. Mapping adaptive fuzzy Kohonen clustering network onto distributed image processing system.

Parallel Computing, 28 (9):1239–1256, September 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/59/29/
abstract.html.

Thomas:1997:MPI

[TMD^{+}97] S. J. Thomas, A. V.

Malevsky, M. Desgagne, R. Benoit, P. Pellerin, and M. Valin. Massively parallel implementation of the mesoscale compressible community model. Parallel Computing, 23(14):2143–2160, December 17, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=14&aid=1263.

Tauer:2013:MRL

[TN13] Gregory Tauer and Rakesh Nagi. A map-reduce Lagrangian heuristic for multidimensional assignment problems with decomposable costs. Parallel Computing, 39(11):653–668, November 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S016781911300104X.

Tomov:2010:ARU

[TND10] Stanimire Tomov, Rajib Nath, and Jack Dongarra. Accelerating the reduction to upper Hessenberg, tridiagonal, and bidiagonal forms through hybrid GPU-based computing. Parallel Computing, 36(12):645–654, December 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tchendji:2020:ECB

[TNZM20] Vianney Kengne Tchendji, Armel Nkonjoh Ngomade, Jerry Lacmou Zeutouo, and Jean Frédéric Myoupo. Efficient CGMbased parallel algorithms for the longest common subsequence problem with multiple substringexclusion constraints. Parallel Computing, 91(??): Article 102598, March 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819119301899.

Troya:1989:SPB

[TO89] J. M. Troya and M. Ortega. A study of parallel branch-and-bound algorithms with best-bound-first search. Parallel Computing, 11(1):121–126, July 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tollenaere:1991:SMN

[TO91] T. Tollenaere and G. A.

Orban. Simulating modular neural networks on message-passing multiprocessors. Parallel Computing, 17(4–5):361–379, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tezduyar:1999:MPC

[TO99] Tayfun Tezduyar and Yasuo Osawa. Methods for parallel computation of complex flow problems. Parallel Computing, 25(13–14):2039–2066, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/36/43/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/43/
article.pdf.

Tolman:2002:DMC

[Tol02] Hendrik L. Tolman. Distributedmemory concepts in the wave model WAVEWATCH III. Parallel Computing, 28(1):35–52, January 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/27/29/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/60/27/29/
                                        00001687.pdf.

Terao:2020:CAT

[TOO20] Takeshi Terao, Katsuhisa Ozaki, and Takeshi Ogita. LU-Cholesky QR algorithms for thin QR decomposition. Parallel Computing, 92(??):Article 102571, April 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819119301620.

Torralba:1991:SAA

[Tor91] A. Torralba. A systolic array with applications to image processing and wirerouting in VLSI circuits. Parallel Computing, 17(1): 85–93, April 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Touzene:2002:EDS

[Tou02] Abderezak Touzene. Edgesdisjoint spanning trees on the binary wrapped butterfly network with applications to fault tolerance. Parallel Computing, 28(4):649–666, April 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/39/31/
abstract.html.

Teodoro:2013:EIW

[TPK^{+}13] George Teodoro, Tony Pan, Tahsin M. Kurc, Jun Kong, Lee A. D. Cooper, and Joel H. Saltz. Efficient irregular wavefront propagation algorithms on hybrid CPU–GPU machines. Parallel Computing, 39(4–5):189–211, April/May 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000343.

Teodoro:2014:RTD

[TPK^{+}14] George Teodoro, Tony Pan, Tahsin Kurc, Jun Kong, Lee Cooper, Scott Klasky, and Joel Saltz. Region templates: Data representation and management for high-throughput image analysis. Parallel Computing, 40(10):589– 610, December 2014. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819114001148.

Traff:1995:ECT

[Trä95] Jesper Larsson Träff. An experimental comparison of two distributed singlesource shortest path algorithms. Parallel Computing, 21(9):1505–1532, September 12, 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.
                                        com/cgi-bin/cas/tree/
                                        store/parco/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=21&issue=
                                        9&aid=1004.

Traff:2012:AUE

[Trä12] Jesper Larsson Träff. Alternative, uniformly expressive and more scalable interfaces for collective communication in MPI. Parallel Computing, 38(1–2):26–36, January/ February 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111001402.

Traff:2018:PDL

[Trä18] Jesper Larsson Träff. Practical, distributed, low overhead algorithms for irregular gather and scatter collectives. Parallel Computing, 75(??):100– 117, July 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301017.

Treleaven:1985:CDD

[Tre85] P. C. Treleaven. Controldriven, data-driven and demand-driven computer architecture. Parallel Computing, 2(3):287–288, November 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Treleaven:1988:PAO

[Tre88] P. C. Treleaven. Parallel architecture overview. Parallel Computing, 8(1– 3):59–70, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tripiccione:1999:A

[Tri99] R. Tripiccione. APEmille.

Parallel Computing, 25 (10–11):1297–1309, September 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/26/23/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/26/23/
article.pdf.

Tang:2013:TBS

[TRLD13] Wei Tang, Dongxu Ren, Zhiling Lan, and Narayan Desai. Toward balanced and sustainable job scheduling for production supercomputers. Parallel Computing, 39 (12):753–768, December 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000999.

Trottenberg:1994:SRS

[Tro94] Ulrich Trottenberg. Some remarks on the SUPRENUM project. Parallel Computing, 20(10–11):1397–1406, November 3, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cgi-bin/cas/tree/
store/parco/cas_sub/
browse/browse.cgi?year=
1994&volume=20&issue=
10-11&aid=904.

Trobec:2000:TDR

[Tro00]  Roman Trobec.  Two-

dimensional regular dmeshes. Parallel Computing, 26(13–14):1945–1953, December 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/34/34/
abstract.html;   http:
//www.elsevier.nl/gej-

ng/10/35/21/42/34/34/
article.pdf.

Tomas:2019:DLA

[TRSC^{+}19] Andrés E. Tomás, Rafael Rodríguez-Sánchez, Sandra Catalán, Rocío Carratalá-Sáez, and Enrique S. Quintana-Ortí. Dynamic look-ahead in the reduction to band form for the singular value decomposition. Parallel Computing, 81(??):22–31, January 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301534.

Tirado-Ramos:2004:IAH

[TRSHB04] Alfredo Tirado-Ramos, Peter M. A. Sloot, Alfons G. Hoekstra, and Marian Bubak. An integrative approach to highperformance biomedical problem solving environments on the Grid. Parallel Computing, 30(9–10): 1037–1055, September/ October 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tanqueray:1988:DSS

[TS88] D. A. Tanqueray and D. F.

Snelling. A distributed self-scheduler for partially ordered tasks. Parallel Computing, 8(1–3):267– 273, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Thole:1999:ISP

[TS99] Clemens-August Thole and Klaus Stüben. Industrial simulation on parallel computers. Parallel Computing, 25(13– 14):2015–2037, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/36/42/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/42/
article.pdf.

Talia:2002:GEP

[TS02] Domenico Talia and Pradip K. Srimani. Guest editorial: Parallel data-intensive algorithms and applications. Parallel Computing, 28(5):669–671, May 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/57/27/
abstract.html.

Trapnell:2009:ODI

[TS09] Cole Trapnell and Michael C. Schatz. Optimizing data intensive GPGPU computations for DNA sequence alignment. Parallel Computing, 35(8–9):429–440, August/September 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tramm:2021:IRR

[TS21] John R. Tramm and Andrew R. Siegel. Immortal rays: Rethinking random ray neutron transport on GPU architectures. Parallel Computing, 108(??):??, December 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000806.

Teixido:2014:MBI

[TSCS14] Ivan Teixidó, Francesc Sebé, Josep Conde, and Francesc Solsona. MPIbased implementation of an enhanced algorithm to solve the LPN problem in a memory-constrained environment. Parallel Computing, 40(5–6):100–112, May 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000453.

Takizawa:2021:OLO

[TSEE21] Hiroyuki Takizawa, Shinji Shiotsuki, Naoki Ebata, and Ryusuke Egawa. OpenCLlike offloading with metaprogramming for SX-aurora TSUBASA. Parallel Computing, 102(??):Article 102754, May 2021. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819121000144.

Touheed:2000:CSD

[TSJB00] N. Touheed, P. Selwood, P. K. Jimack, and M. Berzins. A comparison of some dynamic loadbalancing algorithms for a parallel adaptive flow solver. Parallel Computing, 26(12):1535–1554, November 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/33/24/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/33/24/
article.pdf.

Tan:2017:SCC

[TSLL17] Ying Hao Tan, Yiyang Shao, Siyuan Liu, and Bu-Sung Lee. Student cluster competition: ParConnect reproducibility task report. Parallel Computing, 70(??):11–17, December 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300959.

Trdlicka:2000:ECA

[TT00] Jan Trdlička and Pavel Tvrdík. Embedding complete k-ary trees into ksquare 2D meshes with optimal edge congestion. Parallel Computing, 26(6): 783–790, May 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/42/
28/29/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/42/
28/29/article.pdf.

Tanno:1995:PFA

[TTH95] Kuninobu Tanno, Toshihiro Taketa, and Susumu Horiguchi. Parallel FFT algorithms using radix 4 butterfly computation on an eight-neighbor processor array. Parallel Computing, 21(1):121–136, January 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=1&aid=900.

Teng:2005:HRD

[TTH05] Yuan-Hsiang Teng, Jimmy J. M. Tan, and Lih-Hsing Hsu. Honeycomb rectangular disks. Parallel Computing, 31(3–4):371– 388, March/April 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tiwari:2009:TPA

[TTH09] Ananta Tiwari, Vahid Tabatabaee, and Jeffrey K. Hollingsworth. Tuning parallel applications in parallel. Parallel Computing, 35(8–9):475–492, August/September 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tchendji:2022:CGM

[TTO^{+}22] Vianney Kengne Tchendji, Hermann Bogning Tepiele, Mathias Akong Onabid, Jean Frédéric Myoupo, and Jerry Lacmou Zeutouo. A coarse-grained multicomputer parallel algorithm for the sequential substring constrained longest common subsequence problem. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781912200028X.

Takeda:1992:PSM

[TTT^{+}92] T. Takeda, K. Tani, T. Tsunematsu, Y. Kishimoto, G. I. Kurita, S. Matsushita, and T. Nakata. Plasma simulator METIS for tokamak confinement and heating studies. Parallel Computing, 18(7):743– 765, July 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

tenCate:1996:PEP

[tV96] H. H. ten Cate and E. A. H. Vollebregt. On the portability and efficiency of parallel algorithms and software. Parallel Computing, 22(8):1149–1163, October 28, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=8&aid=1094.

Talnikar:2019:TLC

[TW19] Chaitanya Talnikar and Qiqi Wang. A twolevel computational graph method for the adjoint of a finite volume based compressible unsteady flow solver. Parallel Computing, 81(??):68–84, January 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118300310.

Tang:2000:GET

[TX00] Peiyi Tang and Jingling Xue. Generating efficient tiled code for distributed memory machines. Parallel Computing, 26(11):1369–1410, October 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/32/22/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/32/22/
                                        article.pdf.

Tervola:1991:PJA

[TY91a] P. Tervola and W. Yeung. Parallel Jacobi algorithm for matrix diagonalisation on transputer networks. Parallel Computing, 17(2–3):155–163, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tien:1991:HST

[TY91b] Jenn Yang Tien and Wei Pang Yang. Hierarchical spanning trees and distributing on incomplete hypercubes. Parallel Computing, 17(12):1343–1360, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Terasawa:1995:PEM

[TYKA95] Takuya Terasawa, Ou Yamamoto, Tomohiro Kudoh, and Hideharu Amano. A performance evaluation of the multiprocessor testbed ATTEMPT-0. Parallel Computing, 21(5):701–730, May 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=5&aid=977.

Teng:2022:TLB

[TYLL22] Fei Teng, Lei Yu, Xiao Liu, and Pei Lai. Tight lower bound on power consumption for scheduling realtime periodic tasks in corelevel DVFS systems. Parallel Computing, 110(??): ??, May 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000011.

Tyrtyshnikov:1990:NAD

[Tyr90] Evgenij E. Tyrtyshnikov.

New approaches to deriving parallel algorithms. Parallel Computing, 15 (1–3):261–265, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tan:2013:MCG

[TYSF13] Jingweijia Tan, Yang Yi, Fangyang Shen, and Xin Fu. Modeling and characterizing GPGPU reliability in the presence of soft errors. Parallel Computing, 39(9):520–532, September 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000021.

Uchida:1999:PLP

[UAR^{+}99] Shunichi Uchida, Akira Aiba, Kazuaki Rokusawa, Takashi Chikayama, and Ryuzo Hasegawa. The parallel logic programming system in the FGCS project and its future directions. Parallel Computing, 25(13–14):1601–1633, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/36/27/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/36/27/
                                        article.pdf.

Ucar:2010:MPI

[UÇA10] Bora Uçar, Ümit V.

Çatalyürek, and Cevdet Aykanat. A matrix partitioning interface to Pa-ToH in MATLAB. Parallel Computing, 36(5–6):254– 272, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Uresin:1989:SCC

[ÜD89] Aydin Üresin and Michel Dubois. Sufficient conditions for the convergence of asynchronous iterations. Parallel Computing, 10(1): 83–92, March 1989. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Uhl:1996:WPB

[Uhl96] Andreas Uhl. Wavelet packet best basis selection on moderate parallel MIMD architectures. Parallel Computing, 22 (1):149–158, February 20, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=1&aid=1035.

Ukawa:1999:LQR

[Uka99]  Akira Ukawa.  Lattice

QCD results from the CP-PACS computer. Parallel Computing, 25(10– 11):1257–1280, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/32/26/21/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/32/26/21/
                                        article.pdf.

Umeo:1989:DTO

[Ume89] Hiroshi Umeo. A design of time-optimum and register-number-minimum systolic convolvers. Parallel Computing, 12(3):285– 299, December 1989. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Umeo:2001:LTR

[Ume01] Hiroshi Umeo. Linear-time recognition of connectivity of binary images on 1-bit inter-cell communication cellular automaton. Parallel Computing, 27(5):587– 599, April 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/29/27/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/29/27/
                                        article.pdf.

Umland:1994:PSR

[Uml94] Thomas Umland. Parallel sorting revisited. Parallel Computing, 20(1):115–124, January 16, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=1&aid=813.

Umeo:1987:DPI

[UN87] H. Umeo and I. Nakatsuka. A design of pipelineinterval-optimum systolic stack. Parallel Computing, 4(2):215–219, April 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Utting:2014:JLP

[UWC14] Mark Utting, Min-Hsien Weng, and John G. Cleary. The JStar language philosophy. Parallel Computing, 40(2):35–50, February 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001427.

Uhl:2002:GEP

[UZ02] Andreas Uhl and Peter Zinterhof. Guest editorial: Parallel computing in image and video processing. Parallel Computing, 28(7–8):941–943, August 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/gej-
                                        ng/10/35/21/60/58/28/
                                        abstract.html.

Vu:2014:NPM

[VA14] Lan Vu and Gita Alaghband. Novel parallel method for association rule mining on multicore shared memory systems. Parallel Computing, 40(10):768–785, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114001124.

Vajtersic:1984:PMP

[Vaj84] Marian Vajtersic. Parallel marching Poisson solvers. Parallel Computing, 1(3–4):325–330, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

vanderVorst:1986:PFI

[van86] H. A. van der Vorst. The performance of FOR-TRAN implementations for preconditioned conjugate gradients on vector computers. Parallel Computing, 3(1):49–58, March 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

vanderVorst:1987:APS

[van87a] Henk A. van der Vorst.

Analysis of a parallel solution method for tridiagonal linear systems. Parallel Computing, 5(3):303–311, November 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

vanderVorst:1987:LTB

[van87b] Henk A. van der Vorst.

Large tridiagonal and block tridiagonal linear systems on vector and parallel computers. Parallel Computing, 5(1–2):45–54, July 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

VandeVelde:1990:DRC

[Van90] Eric F. Van de Velde. Data redistribution and concurrency. Parallel Computing, 16(2–3):125–138, December 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

vanderSteen:1991:BEG

[van91] A. J. van der Steen. The benchmark of the EuroBen group. Parallel Computing, 17(10–11):1211–1221, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Vanneschi:2002:PMA

[Van02] Marco Vanneschi. The programming model of ASSIST, an environment for parallel and distributed portable applications. Parallel Computing, 28(12):1709–1732, December 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Varvarigos:1992:CAI

[VB92] Emmanouel A. Varvarigos and Dimitri P. Bertsekas. Communication algorithms for isotropic tasks in hypercubes and wraparound meshes. Parallel Computing, 18(11): 1233–1257, November 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Varvarigos:1995:TBM

[VB95] Emmanouel A. Varvarigos and Dimitri P. Bertsekas. Transposition of banded matrices in hypercubes: a nearly isotropic task. Parallel Computing, 21(2):243–264, February 17, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=2&aid=897.

Vishnu:2019:GEI

[VBC19] Abhinav Vishnu, Pavan Balaji, and Yong Chen. Guest Editor’s introduction: P2S2: SI 2016. Parallel Computing, 82(??):1– 2, ???? 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119300249.

Vasseur:2015:IDM

[VBS^{+}15] Romain Vasseur, Stéphanie Baud, Luiz Angelo Steffenel, Xavier Vigouroux, Laurent Martiny, Michaël Krajecki, and Manuel Dauchez. Inverse docking method for new proteins targets identification: a parallel approach. Parallel Computing, 42(??):48–59, February 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001197.

vanSwaaij:1990:DAA

[vCd90] M. F. X. B. van Swaaij, F. V. M. Catthoor, and H. J. de Man. Deriving ASIC architectures for the Hough transform. Parallel Computing, 16(1):113–121, November 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Vydyanathan:2011:OLT

[VCK^{+}11] Naga Vydyanathan, Umit Catalyurek, Tahsin Kurc, Ponnuswamy Sadayappan, and Joel Saltz. Optimizing latency and throughput of application workflows on clusters. Parallel Computing, 37(10– 11):694–712, October/ November 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819110000815.

Vegh:2018:IEM

[Vég18] János Végh. Introducing the explicitly manyprocessor approach. Parallel Computing, 75(??):28– 40, July 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118300577.

Verriet:1999:SIO

[Ver99] Jacques Verriet. Scheduling interval-ordered tasks with non-uniform deadlines subject to nonzero communication delays. Parallel Computing, 25(1):3–21, January 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/1/1375.pdf.

Verriet:2000:SOH

[Ver00] Jacques Verriet. Scheduling outtrees of height one in the LogP model. Parallel Computing, 26(9):1065– 1082, July 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/42/30/24/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/42/30/24/
article.pdf.

Vazquez:2012:ATS

[VFG12] Francisco Vázquez, José Jesús Fernández, and Ester M. Garzón. Automatic tuning of the sparse matrix vector product on GPUs based on the ELLR-T approach. Parallel Computing, 38(8):408–420, August 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001050.

Vega-Gisbert:2016:DIJ

[VGRS16] Oscar Vega-Gisbert, Jose E. Roman, and Jeffrey M. Squyres. Design and implementation of Java bindings in Open MPI. Parallel Computing, 59(??):1–20, November 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819116300758.

Vikas:2014:MGA

[VGS14] Vikas, Nasser Giacaman, and Oliver Sinnen. Multiprocessing with GUI-awareness using OpenMP-like directives in Java. Parallel Computing, 40(2):69–89, February 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113001439.

Varman:1990:PMA

[VIHD90] Peter J. Varman, Balakrishna R. Iyer, Donald J. Haderle, and Stephen M. Dunn. Parallel merging: algorithm and implementation results. Parallel Computing, 15(1–3): 165–177, September 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Violard:2004:SFA

[Vio04] Eric Violard. A semantic framework to address data locality in data parallel languages. Parallel Computing, 30(1):139–161, January 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Varoglu:2011:AST

[VJ11] Sevin Varoglu and Stephen Jenks. Architectural support for thread communications in multi-core processors. Parallel Computing, 37(1):26–41, January 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Verschoor:2012:APE

[VJ12] Mickeal Verschoor and Andrei C. Jalba. Analysis and performance estimation of the conjugate gradient method on multiple GPUs. Parallel Computing, 38 (10–11):552–575, October/ November 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000609.

Vlahavas:1992:PPR

[VK92] I. Vlahavas and P. Kefalas.

A parallel Prolog resolution based on multiple unifications. Parallel Computing, 18(11):1275–1283, November 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Verma:2017:HMO

[VK17] Amandeep Verma and Sakshi Kaushal. A hybrid multi-objective Particle Swarm Optimization for scientific workflow scheduling. Parallel Computing, 62(??):1–19, February 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117300145.

Vlahavas:1999:OAP

[VKH99] I. Vlahavas, P. Kefalas, and C. Halatsis. OASys: an AND/OR parallel logic programming system. Parallel Computing, 25(3):321–336, March 22, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1999/
25/3/1384.pdf.

Venetis:2015:DTS

[VKS^{+}15] I. E. Venetis, A. Kouris, A. Sobczyk, E. Gallopoulos, and A. H. Sameh. A direct tridiagonal solver based on Givens rotations for GPU architectures. Parallel Computing, 49(??):101–116, November 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000605.

Verkaik:2005:CNP

[VL05] J. Verkaik and H. X. Lin.

A class of novel parallel algorithms for the solution of tridiagonal systems. Parallel Computing, 31(6):563– 587, June 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Veltman:1990:MSC

[VLL90] B. Veltman, B. J. Lageweg, and J. K. Lenstra. Multiprocessor scheduling with communication delays. Parallel Computing, 16 (2–3):173–182, December 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Valero-Lara:2019:MTS

[VLSPL19] Pedro Valero-Lara, Raül Sirvent, Antonio J. Peña, and Jesús Labarta. MPI + OpenMP tasking scalability for multi-morphology simulations of the human brain. Parallel Computing, 84(??):50–61, May 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781911830317X.

vanHalderen:1998:HRM

[vOS^{+}98] A. W. van Halderen, B. J. Overeinder, P. M. A. Sloot, R. van Dantzig, D. H. J. Epema, and M. Livny. Hierarchical resource management in the Polder Metacomputing Initiative. Parallel Computing, 24(12– 13):1807–1825, November 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1998/
                                        24/12-13/1356.pdf.

Violard:1992:PLR

[VP92] E. Violard and G.-R. Perrin. PEI: a language and its refinement calculus for parallel programming. Parallel Computing, 18(10):1167–1184, October 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

VanHuffel:1994:PTB

[VP94] Sabine Van Huffel and Haesun Park. Parallel triand bi-diagonalization of bordered bidiagonal matrices. Parallel Computing, 20(8):1107–1128, August 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=8&aid=884.

Vaidyanathan:1995:SCB

[VP95] Ramachandran Vaidyanathan and Anand Padmanabhan. Short communication: Bus-based networks for fan-in and uniform hypercube algorithms. Parallel Computing, 21(11): 1807–1821, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=11&aid=1023.

VanDriessche:1995:ISB

[VR95] R. Van Driessche and D. Roose. An improved spectral bisection algorithm and its application to dynamic load balancing. Parallel Computing, 21(1):29–48, January 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.

                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=1&aid=925.

Vega-Rodriguez:2015:PBV

[VRGÁ15] Miguel A. Vega-Rodríguez and David L. González-Álvarez. Parallelism in bioinformatics: a view from different parallelismbased technologies. Parallel Computing, 42(??):1–3, February 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115000204.

Vollebregt:2003:LSC

[VRL03] E. A. H. Vollebregt, M. R. T. Roest, and J. W. M. Lander. Large scale computing at Rijkswaterstaat. Parallel Computing, 29(1):1–20, January 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

vonHanxleden:1992:CDP

[vS92] R. von Hanxleden and L. R. Scott. Correctness and determinism of Parallel Monte Carlo Processes. Parallel Computing, 18(2): 121–132, February 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Vaidyanathan:2003:DSS

[VTmL03] Ramachandran Vaidyanathan, Jerry L. Trahan, and Chun ming Lu. Degree of scalability: scalable reconfigurable mesh algorithms for multiple addition and matrix–vector multiplication. Parallel Computing, 29(1):95–109, January 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

vanderPas:1991:PMU

[vv91] R. J. van der Pas and J. M. van Kats. Parallelism in a multi-user environment. Parallel Computing, 17(2–3):285–296, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Vanneschi:2007:DDA

[VV07] M. Vanneschi and L. Veraldi. Dynamicity in distributed applications: issues, problems and the ASSIST approach. Parallel Computing, 33(12): 822–845, December 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

VanderVlugt:1990:INS

[VvBv90] F. F. Van der Vlugt, D. A. van Delft, A. F. Bakker, and T. H. van der Meer. The implementation of a 3D Navier–Stokes algorithm on an algorithm oriented processor. Parallel Computing, 15(1–3):47– 60, September 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Vuik:1998:PIP

[VvW98] C. Vuik, R. R. P. van Nooyen, and P. Wesseling. Parallelism in ILUpreconditioned GMRES. Parallel Computing, 24 (14):1927–1946, December 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/14/1361.pdf.

vanWagensveld:2019:EMS

[vWWM^{+}19] Remko van Wagensveld, Tobias Wägemann, Ralph Mader, Ramin Tavakoli Kolagari, and Ulrich Margull. Evaluation and modeling of the supercore parallelization pattern in automotive real-time systems. Parallel Computing, 81(??):122–130, January 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301595.

Womble:1999:MPC

[W^{+}99] David E. Womble et al. Massively parallel computing: A Sandia perspective. Parallel Computing, 25(13–14):1853–1876, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.nl/gej-
ng/10/35/21/32/36/36/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/36/
article.pdf.

Wang:2003:QAD

[WA03] Shaowen Wang and Marc P.

Armstrong. A quadtree approach to domain decomposition for spatial interpolation in Grid computing environments. Parallel Computing, 29(10): 1481–1504, October 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Womble:1989:IIM

[WAB89] David E. Womble, Richard C.

Allen, Jr., and Lorraine S. Baca. Invariant imbedding and the method of lines for parallel computers. Parallel Computing, 11(3):263–273, ???? 28, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wagner:1989:PSA

[Wag89] Robert A. Wagner. Parallel solution of arbitrarily sparse linear systems.

Parallel Computing, 9(3): 313–331, February 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wait:1988:PPF

[Wai88] R. Wait. Partitioning and preconditioning of finite element matrices on the DAP. Parallel Computing, 8(1–3):275–284, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wakatani:2004:PSA

[Wak04] Akiyoshi Wakatani. A parallel and scalable algorithm for ADI method with pre-propagation and message vectorization. Parallel Computing, 30(12): 1345–1359, December 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Walker:1994:DSM

[Wal94a] David W. Walker. The design of a standard message passing interface for distributed memory concurrent computers. Parallel Computing, 20(4):657–673, March 31, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.

cgi?year=1994&volume=
20&issue=4&aid=865. See

erratum [Wal94b].

Walker:1994:EDS

[Wal94b] David W. Walker. Erratum to: “The design of a standard message passing interface for distributed memory concurrent computers”. Parallel Computing, 20(8):1215, August 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). See [Wal94a].

Walker:2001:SEC

[Wal01] Reginald L. Walker. Search engine case study: searching the Web using genetic programming and MPI. Parallel Computing, 27(1–2):71–89, January 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/25/25/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/25/25/
article.pdf.

Wang:2008:LTA

[Wan08] Dajin Wang. A linear-time algorithm for computing collision-free path on reconfigurable mesh. Parallel Computing, 34(9):487–496, September 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wang:2009:NTA

[Wan09] Yuh-Rau Wang. A novel O(1) time algorithm for 3D block-based medial axis transform by peeling corner shells. Parallel Computing, 35(2):72–82, February 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Watanabe:1987:APN

[Wat87] T. Watanabe. Architecture and performance of NEC supercomputer SX system. Parallel Computing, 5(1–2):247–255, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Watts:2000:PAC

[Wat00] John D. Watts. Parallel algorithms for coupledcluster methods. Parallel Computing, 26(7–8):857– 867, July 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/29/26/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/29/26/
                                        article.pdf.

Watanabe:2020:DPM

[WAT20] Seiya Watanabe, Takayuki Aoki, and Tomohiro Takaki. A domain partitioning method using a multiphase-field model for blockbased AMR applications. Parallel Computing, 97 (??):Article 102647, September 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300405.

Wait:1988:OBM

[WB88] R. Wait and N. G. Brown.

Overlapping block methods for solving tridiagonal systems on transputer arrays. Parallel Computing, 8(1–3):325–333, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the International Conference on Vector and Parallel Processors in Computational Science, III (Liverpool, 1987).

Watlington:1997:SPM

[WB97] John A. Watlington and V. Michael Bove, Jr. A system for parallel media processing. Parallel Computing, 23(12):1793–1809, December 15, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=12&aid=1246.

Wedemann:1999:EDT

[WBD99a] R. S. Wedemann, V. C. Barbosa, and R. Donangelo. Erratum to “Defeasible time-stepping” [Parallel Computing 25 (4) (April 1999) pp. 461–489]. Parallel Computing, 25(6): 767, June 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/cas/
                                        tree/store/parco/sub/
                                        1999/25/6/1425.pdf. See

[WBD99b].

Wedemann:1999:DTS

[WBD99b] Roseli S. Wedemann, Valmir C. Barbosa, and Raul Donangelo. Defeasible time-stepping. Parallel Computing, 25(4):461–489, April 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/sub/1999/
                                        25/4/1393.pdf. See erra-

tum [WBD99a].

Williams:2017:SSC

[WBN^{+}17] G. R. Williams, G. P. Behm, T. Nguyen, A. Esparza, V. G. Haka, A. Ramos, B. Wright, J. C. Otto, C. P. Paolini, and M. P. Thomas. SC16 student cluster competition challenge: Investigating the reproducibility of results for the ParConnect application. Parallel Computing, 70(??):27–34, December 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911730100X.

Wang:1997:DPP

[WBPM97] Xin Wang, Edward K.

Blum, D. Stott Parker, and Daniel Massey. The dance party problem and its application to collective communication in computer networks. Parallel Computing, 23(8):1141–1156, July 25, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=8&aid=1199.

Wapperom:2006:NTS

[WBS06] P. Wapperom, A. N. Beris, and M. A. Straka. A new transpose split method for three-dimensional FFTs: performance on an Origin2000 and Alphaserver cluster. Parallel Computing, 32(1):1–13, January 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Weston:1990:TAP

[WC90] J. S. Weston and M. Clint. Two algorithms for the parallel computation of eigenvalues and eigenvectors of large symmetric matrices using the ICL DAP. Parallel Computing, 13(3):281–288, March 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Waring:1991:PGS

[WC91] L. C. Waring and M. Clint. Parallel Gram–Schmidt orthogonalisation on a network of transputers. Parallel Computing, 17(9):1043– 1050, November 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wang:1994:DSM

[WC94] Hsiao-Hsi Wang and Ruei-Chuan Chang. A distributed shared memory system with self-adjusting coherence scheme. Parallel Computing, 20(7):1007– 1025, July 12, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.
                                        com/cgi-bin/cas/tree/
                                        store/parco/cas_sub/

browse/browse.cgi?year=
1994&volume=20&issue=
7&aid=876.

Walshaw:2000:POA

[WC00] C. Walshaw and M. Cross.

Parallel optimisation algorithms for multilevel mesh partitioning. Parallel Computing, 26(12):1635– 1660, November 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
nl/gej-ng/10/35/21/42/
33/28/abstract.html;
http://www.elsevier.
nl/gej-ng/10/35/21/42/
33/28/article.pdf.

Wang:2015:PPC

[WC15] Haifeng Wang and Yunpeng Cao. Predicting power consumption of GPUs with fuzzy wavelet neural networks. Parallel Computing, 44(??):18–36, May 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115000320.

Wang:2022:LAC

[WCC^{+}22] Ziheng Wang, Heng Chen, Weiling Cai, Xiaoshe Dong, and Xingjun Zhang. C-Lop: Accurate contentionbased modeling of MPI concurrent communication. Parallel Computing, 111(??):??, July 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000278.

Wang:2011:CCI

[WCKM11] Chenqi Wang, Neil Cafferkey, James Kennedy, and John P. Morrison. CG3DR: Coordination of icosahedral virus reconstruction using Condensed Graphs. Parallel Computing, 37(8):451–465, August 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819110001390.

Weaver:2014:TDS

[WCM^{+}14] Jesse Weaver, Vito Giovanni Castellana, Alessandro Morari, Antonino Tumeo, Sumit Purohit, Alan Chappell, David Haglin, Oreste Villa, Sutanay Choudhury, Karen Schuchardt, and John Feo. Toward a data scalable solution for facilitating discovery of science resources. Parallel Computing, 40(10):682–696, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001100.

Wheat:1991:APM

[WE91a] M. Wheat and D. J.

Evans. Asynchronous parallel merging. Parallel Computing, 17(9):1035– 1041, November 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wheat:1991:MSD

[WE91b] M. Wheat and D. J.

Evans. Maintenance of shared data structures on tightly coupled multiprocessors. Parallel Computing, 17(1):101–107, April 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wheat:1992:EPS

[WE92] M. Wheat and D. J. Evans.

An efficient parallel sorting algorithm for shared memory multiprocessors. Parallel Computing, 18(1):91– 102, January 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wan:1993:SAA

[WE93] C. R. Wan and D. J.

Evans. A systolic array architecture for linear and inverse matrix systems. Parallel Computing, 19(3):303–321, March 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wan:1994:SAA

[WE94] C. R. Wan and D. J. Evans. A systolic array architecture for QR decomposition of block structured sparse systems. Parallel Computing, 20(6):903–914, June 10, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=6&aid=838.

Westermann:1997:DVV

[WE97] Rüdiger Westermann and Thomas Ertl. Distributed volume visualization: a step towards integrated data analysis and image synthesis. Parallel Computing, 23(7):927–941, July 14, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=7&aid=1193.

Weberpals:1990:AAI

[Web90] H. Weberpals. Architectural approach to the IBM 3090E vector performance.

Parallel Computing, 13(1): 47–59, January 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Weicker:1991:DLS

[Wei91] R. P. Weicker. A detailed look at some popular benchmarks. Parallel Computing, 17(10– 11):1153–1172, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wei:1993:PAI

[Wei93]  Jiawang Wei.  Parallel

asynchronous iterations of least fixed points. Parallel Computing, 19(8):887– 895, August 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Weimar:1997:CAR

[Wei97] Jörg R. Weimar. Cellular automata for reaction– diffusion systems. Parallel Computing, 23(11): 1699–1715, December 1, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=11&aid=1239.

Weimar:2001:CMM

[Wei01] Jörg R. Weimar. Coupling microscopic and macroscopic cellular automata. Parallel Computing, 27(5):601–611, April 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/29/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/29/28/
article.pdf.

Welch:1989:EDL

[Wel89] P. H. Welch. Emulating digital logic using transputer networks (very high parallelism=simplicity=performance). Parallel Computing, 9(2): 257–272, January 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wen:1995:FPA

[Wen95]  Zhaofang Wen.  Fast

parallel algorithms for the maximum sum problem. Parallel Computing, 21(3):461–466, March 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=3&aid=929.

Wang:2011:CPC

[WFW^{+}11] Hao Wang, Xudong Fu, Guangqian Wang, Tiejian Li, and Jie Gao. A common parallel computing framework for modeling hydrological processes of river basins. Parallel Computing, 37(6–7):302–315, June/July 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000500.

Womble:1997:PI

[WG97] David E. Womble and David S. Greenberg. Parallel I/O: an introduction. Parallel Computing, 23(4–5):403–417, May 23, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=4-5&aid=1165.

Wiegand:1991:DIR

[WH91] F. Wiegand and B. S.

Hoyle. Development and implementation of realtime ultrasound process tomography using a transputer network. Parallel Computing, 17(6–7):791– 807, September 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wong:1993:PDS

[WH93] Yung-Chang Wong and Shu-Yuen Hwang. On parallelizing the Dempster-Shafer method using transputer network. Parallel Computing, 19(7):807–822, July 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wu:1994:NBC

[WH94] Jean-Lien C. Wu and T.-Y. Huang. A new bus contention scheme in S/NET with dynamic priority. Parallel Computing, 20(7):1041–1054, July 12, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=7&aid=875.

Wabnig:1997:PTP

[WH97] H. Wabnig and G. Haring. PAPS — a testbed for performance prediction of parallel applications. Parallel Computing, 22(13):1837–1851, February 28, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.

elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=13&aid=1131.

Weise:2015:RMF

[WH15] S. Weise and C. Hasse.

Reducing the memory footprint in Large Eddy Simulations of reactive flows. Parallel Computing, 49(??):50–65, November 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819115001076.

Wang:2022:SCF

[WH22] Xinyuan Wang and Hejiao Huang. SGPM: a coroutine framework for transaction processing. Parallel Computing, 114(??):??, December 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819122000709.

Wu:2019:SCC

[WHL19] Bian Wu, Weiliang Heng, and Bu-Sung Lee. Student Cluster Competition 2018, Team Nanyang Technological University: Reproducing performance of a Multi-Physics Simulations of the Tsunamigenic 2004 Sumatra Megathrust Earthquake on the Intel Skylake architecture. Parallel Computing, 90(??): Article 102564, December 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819119301553.

Wu:1991:PSO

[WHW91] C.-H. Wu, R. E. Hodges, and C. J. Wang. Parallelizing the self-organizing feature map on multiprocessor systems. Parallel Computing, 17(6–7):821– 832, September 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wilde:2011:SLD

[WHW^{+}11] Michael Wilde, Mihael Hategan, Justin M. Wozniak, Ben Clifford, Daniel S. Katz, and Ian Foster. Swift: a language for distributed parallel scripting. Parallel Computing, 37(9):633–652, September 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111000524.

Williams:1985:TCC

[Wil85] S. A. Williams. The transformation of collections of communicating sequential processes that represent pipeline configurations. Parallel Computing, 2(4):345–351, December 1985. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Williams:1988:FLH

[Wil88]  R. Williams.  Free-

Lagrange hydrodynamics with a distributed-memory parallel processor. Parallel Computing, 7(3):439–443, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wittenbrink:1998:EPW

[Wit98] Craig M. Wittenbrink. Extensions to permutation warping for parallel volume rendering. Parallel Computing, 24(9– 10):1385–1406, September 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/9-10/1340.pdf.

Warfield:1998:HPC

[WJK98] Simon K. Warfield, Ferenc A. Jolesz, and Ron Kikinis. A high performance computing approach to the registration of medical imaging data. Parallel Computing, 24(9– 10):1345–1368, September 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/9-10/1338.pdf.

Wu:1998:EOI

[WL98] Joe Shang-Chieh Wu and Ying-Dar Lin. An efficient and orderly implementation of bypass queue under bursty traffic. Parallel Computing, 24(14):2143–2148, December 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/14/1365.pdf.

Wan:2013:HPC

[WL13a] Xiaoliang Wan and Guang Lin. Hybrid parallel computing of minimum action method. Parallel Computing, 39(10):638–651, October 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000963.

Wang:2013:DDD

[WL13b] Yang Wang and Paul Lu. DDS: a deadlock detection-based scheduling algorithm for workflow computations in HPC systems with storage constraints. Parallel Computing, 39(8):291–305, August 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000501.

White:2014:DOI

[WL14] Andrew White and Soo-Young Lee. Derivation of optimal input parameters for minimizing execution time of matrixbased computations on a GPU. Parallel Computing, 40(10):628–645, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001136.

Wang:2002:MSP

[WLCG02] P. Wang, Karen Y. Liu, Tom Cwik, and Robert Green. MODTRAN on supercomputers and parallel computers. Parallel Computing, 28(1):53–64, January 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/27/30/

                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/60/27/30/
                                        00001685.pdf.

Wang:1997:SSC

[WLK97] Jiahong Wang, Jie Li, and Hisao Kameda. Simulation studies on concurrency control in parallel transaction processing systems. Parallel Computing, 23(6):755–775, June 20, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=6&aid=1177.

Wolfe:2018:ODM

[WLK^{+}18] Michael Wolfe, Seyong Lee, Jungwon Kim, Xiaonan Tian, Rengan Xu, Barbara Chapman, and Sunita Chandrasekaran. The OpenACC data model: Preliminary study on its major challenges and implementations. Parallel Computing, 78(??):15–27, October 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819118302175.

Wimberly:1996:PTP

[WLN^{+}96] Frank C. Wimberly, Michael H.

Lambert, Nicholas A. Nystrom, Alex Ropelewski, and William Young. Porting third-party applications packages to the Cray T3D: Programming issues and scalability results. Parallel Computing, 22(8):1073–1089, October 28, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=8&aid=1088.

Wang:2013:ARR

[WLYJ13] Xiaohang Wang, Peng Liu, Mei Yang, and Yingtao Jiang. Avoiding request-request type message-dependent deadlocks in networks-on-chips. Parallel Computing, 39 (9):408–423, September 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000689.

Wu:2023:OSA

[WLZ^{+}23] Ou Wu, Shanshan Li, He Zhang, Liwen Liu, Haoming Li, Yanze Wang, and Ziyi Zhang. An optimal scheduling algorithm considering the transactions worst-case delay for multi-channel hyperledger fabric network. Parallel Computing, 117(??):??, September 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819123000479.

Wallcraft:1997:NLO

[WM97] Alan J. Wallcraft and Daniel R. Moore. The NRL layered ocean model. Parallel Computing, 23 (14):2227–2242, December 17, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=14&aid=1269.

Walker:1997:NCF

[WMCU97] Edward Walker, Gary Morgan, Bruce Cass, and Zygmunt Ulanowski. A note on compiling FOR-TRAN loop kernels onto a dataflow architecture. Parallel Computing, 22 (11):1545–1557, January 26, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.

cgi?year=1997&volume=
22&issue=11&aid=1103.

Wehner:1995:PDM

[WME^{+}95] M. F. Wehner, A. A. Mirin, P. G. Eltgroth, W. P. Dannevik, C. R. Mechoso, J. D. Farrara, and J. A. Spahr. Performance of a distributed memory finite difference atmospheric general circulation model. Parallel Computing, 21 (10):1655–1675, November 29, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=10&aid=1020.

Wang:1991:PNP

[WN91] C.-J. Wang and V. P. Nelson. Petri net performance modeling of a modified mesh-connected parallel computer. Parallel Computing, 17(1):75– 84, April 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wahle:2001:CAT

[WNES01] J. Wahle, L. Neubert, J. Esser, and M. Schreckenberg. A cellular automaton traffic flow model for online simulation of traffic. Parallel Computing, 27(5):719–735, April 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/47/29/33/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/29/33/
                                        article.pdf.

Wismuller:1996:IDP

[WOKH96] Roland Wismüller, Michael Oberhuber, Johann Krammer, and Olav Hansen. Interactive debugging and performance analysis of massively parallel applications. Parallel Computing, 22(3):415–442, April 29, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=3&aid=1049.

Wolfstahl:1989:MPP

[Wol89] Y. Wolfstahl. Mapping parallel programs to multiprocessors: a dynamic approach. Parallel Computing, 10(1):45–50, March 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wong:1988:SLE

[Won88] Yau Shu Wong. Solving large elliptic difference equations on CYBER 205. Parallel Computing, 6(2): 195–207, February 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Worlton:1991:TTP

[Wor91] J. Worlton. Toward a taxonomy of performance metrics. Parallel Computing, 17(10–11):1073–1092, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Worley:1992:EMR

[Wor92] P. H. Worley. The effect of multiprocessor radius on scaling. Parallel Computing, 18(4):361–376, April 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Worsch:1997:PTM

[Wor97] Thomas Worsch. On parallel Turing machines with multi-head control units. Parallel Computing, 23 (11):1683–1697, December 1, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=11&aid=1240.

Williams:2009:OSM

[WOV^{+}09] Samuel Williams, Leonid Oliker, Richard Vuduc, John Shalf, Katherine Yelick, and James Demmel. Optimization of sparse matrix-vector multiplication on emerging multicore platforms. Parallel Computing, 35(3): 178–194, March 2009. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wagner:1988:SMA

[WP88] Robert A. Wagner and Merrell L. Patrick. A sparse matrix algorithm on the Boolean vector machine. Parallel Computing, 6(3):359–371, March 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wu:2019:DPA

[WP19] Xinzhe Wu and Serge G. Petiton. A distributed and parallel asynchronous unite and conquer method to solve large scale non-Hermitian linear systems with multiple right-hand sides. Parallel Computing, 89(??):Article 102551, November 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301425.

Whaley:2001:AEO

[WPD01] R. Clint Whaley, Antoine Petitet, and Jack J. Dongarra. Automated empirical optimizations of software and the ATLAS project. Parallel Computing, 27(1–2):3–35, January 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/25/23/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/25/23/
article.pdf.

Wirthlin:2008:OCC

[WPS^{+}08] M. Wirthlin, D. Poznanovic, P. Sundararajan, A. Coppola, D. Pellerin, W. Najjar, R. Bruce, M. Babst, O. Pritchard, P. Palazzari, and G. Kuzmanov. OpenFPGA Core-Lib core library interoperability effort. Parallel Computing, 34(4–5):231– 244, May 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wang:1997:PAE

[WRB97] Xiaodong Wang, Vwani P.

Roychowdhury, and Pratheep Balasingam. Practical aspects and experiences. scalable massively parallel algorithms for computational nanoelectronics. Parallel Computing, 22(14):1931–1963, March 24, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        22&issue=14&aid=1120.

Wright:1990:SDT

[Wri90] Stephen J. Wright. Solution of discrete-time optimal control problems on parallel computers. Parallel Computing, 16(2–3): 221–237, December 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wright:1991:PAD

[Wri91] K. Wright. Parallel algorithms for QR decomposition on a shared memory multiprocessor. Parallel Computing, 17(6–7): 779–790, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wyrzykowski:2012:MDA

[WRS12] Roman Wyrzykowski, Krzysztof Rojek, and Lukasz Szustak. Model-driven adaptation of double-precision matrix multiplication to the Cell processor architecture. Parallel Computing, 38(4–5):260–276, April/May 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001086.

Wu:1994:HAT

[WS94] Shen Shen Wu and David Sweeting. Heuristic algorithms for task assignment and scheduling in a processor network. Parallel Computing, 20(1):1–14, January 16, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=1&aid=824.

Williams:1998:HPA

[WS98] Roy Williams and Bruce Sears. A high-performance active digital library. Parallel Computing, 24(12– 13):1791–1806, November 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/
24/12-13/1355.pdf.

Wensch:2004:PSA

[WS04] Jörg Wensch and Ben Sommeijer. Parallel simulation of axon growth in the nervous system. Parallel Computing, 30(2):163– 186, February 2004. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wernsing:2012:ECP

[WS12] John R. Wernsing and Greg Stitt. Elastic computing: a portable optimization framework for hybrid computers. Parallel Computing, 38(8):438–464, August 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819112000300.

Whitley:1990:GAN

[WSB90] D. Whitley, T. Starkweather, and C. Bogart. Genetic algorithms and neural networks: optimizing connections and connectivity. Parallel Computing, 14(3):347–361, August 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wasserman:1988:PMA

[WSL88] H. J. Wasserman, M. L. Simmons, and O. M. Lubeck. The performance of minisupercomputers: Alliant FX/8, Convex C-1, and SCS-40. Parallel Computing, 8(1–3):285– 293, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wyrzykowski:2014:PME

[WSR14] Roman Wyrzykowski, Lukasz Szustak, and Krzysztof Rojek. Parallelization of 2D MPDATA EULAG algorithm on hybrid architectures with GPU accelerators. Parallel Computing, 40(8):425–447, August 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114000520.

Wu:2023:PSP

[WTLW23] Guoqing Wu, Hongyun Tian, Guo Lu, and Wei Wang. ParVoro++: a scalable parallel algorithm for constructing 3D Voronoi tessellations based on kdtree decomposition. Parallel Computing, 115(??):??, February 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000017.

Weinzierl:2016:TPG

[WVHR16] T. Weinzierl, B. Verleye, P. Henri, and D. Roose. Two particle-in-grid realisations on spacetrees. Parallel Computing, 52(??): 42–64, February 2016. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0167819115001635.

Wilson:2016:TUD

[WvR16] Lucas A. Wilson and Jeffery von Ronne. A taskuncoordinated distributed dataflow model for scalable high performance parallel program execution. Parallel Computing, 51(??):79– 87, January 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115001453.

Wang:1990:SPC

[WW90] Chien Min Wang and Sheng-De Wang. Structured partitioning of concurrent programs for execution on multiprocessors. Parallel Computing, 16(1): 41–57, November 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wang:1992:HSE

[WW92] Chien-Min Wang and Sheng-De Wang. A hybrid scheme for efficiently executing nested loops on multiprocessors. Parallel Computing, 18(6):625–637, June 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wakatani:1995:OAR

[WW95] Akiyoshi Wakatani and Michael Wolfe. Optimization of array redistribution for distributed memory multicomputers. Parallel Computing, 21 (9):1485–1490, September 12, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=9&aid=1006.

Wu:2022:RAS

[WWL^{+}22] Yalan Wu, Jigang Wu, Peng Liu, Yinhe Han, and Thambipillai Srikanthan. Reconfiguration algorithms for synchronous communication on switch based degradable arrays. Parallel Computing, 111 (??):??, July 2022. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819122000084.

Wang:2005:DPO

[WWR05] Z. G. Wang, Y. S. Wong, and M. Rahman. Development of a parallel optimization method based on genetic simulated annealing algorithm. Parallel Computing, 31(8–9):839– 857, August/September 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wong:2001:PGA

[WWT01] S. C. Wong, C. K. Wong, and C. O. Tong. A parallelized genetic algorithm for the calibration of Lowry model. Parallel Computing, 27 (12):1523–1536, November 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/47/42/28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/42/28/
article.pdf.

Wang:2018:IAD

[WWZ^{+}18] Junxiong Wang, Hongzhi Wang, Chenxu Zhao, Jianzhong Li, and Hong Gao. Iteration acceleration for distributed learning systems. Parallel Computing, 72(??):29–41, February 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819118300024.

Wang:2018:DMN

[WXL^{+}18] Hongzhi Wang, Feng Xiong, Jianing Li, Shengfei Shi, Jianzhong Li, and Hong Gao. Data management on new processors: A survey. Parallel Computing, 72(??):1–13, February 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819117302053.

Wei:2011:HMP

[WY11] Fangzhou Wei and Ali E.

Yilmaz. A hybrid message passing/shared memory parallelization of the adaptive integral method for multi-core clusters. Parallel Computing, 37 (6–7):279–301, June/July 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111000354.

Williams-Young:2021:APP

[WYBdJ^{+}21] David B. Williams-Young, Abhishek Bagusetty, Wibe A. de Jong, Douglas Doerfler, Hubertus J. J. van Dam, Álvaro Vázquez-Mayagoitia, Theresa L. Windus, and Chao Yang.

Achieving performance portability in Gaussian basis set density functional theory on accelerator based architectures in NWChemEx. Parallel Computing, 108(??):??, December 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000776.

Wang:2016:GFD

[WYT^{+}16] Chen Wang, Ce Yu, Shanjiang Tang, Jian Xiao, Jizhou Sun, and Xiangfei Meng. A general and fast distributed system for large-scale dynamic programming applications. Parallel Computing, 60(??):1–21, December 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781911630103X.

Wang:2022:MER

[WYX^{+}22] Hao Wang, Ce Yu, Jian Xiao, Shanjiang Tang, Yu Lu, Hao Fu, Bo Kang, Gang Zheng, and Chenzhou Cui. A method for efficient radio astronomical data gridding on multicore vector processor. Parallel Computing, 113(??): ??, October 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781912200062X.

Witbrock:1990:IBL

[WZ90] M. Witbrock and M. Zagha. An implementation of backpropagation learning on GF11, a large SIMD parallel computer. Parallel Computing, 14(3):329– 346, August 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wang:2014:CGS

[WZCG14] Zhenning Wang, Long Zheng, Quan Chen, and Minyi Guo. CPU + GPU scheduling with asymptotic profiling. Parallel Computing, 40(2):107–115, February 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001415.

Wang:2009:DKM

[WZL09] Yang Wang, Ming Zhu, and Hua Li. A distributed Key Message algorithm to optimize the communication in clusters. Parallel Computing, 35(7):401–415, July 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wang:2014:AHT

[WZS^{+}14] Yizhuo Wang, Yang Zhang, Yan Su, Xiaojun Wang, Xu Chen, Weixing Ji, and Feng Shi. An adaptive and hierarchical task scheduling scheme for multi-core clusters. Parallel Computing, 40(10):611–627, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001239.

Wallace:2016:APP

[WZV^{+}16] Sean Wallace, Zhou Zhou, Venkatram Vishwanath, Susan Coghlan, John Tramm, Zhiling Lan, and Michael E. Papka. Application power profiling on IBM Blue Gene/ Q. Parallel Computing, 57(??):73–86, September 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300485.

Xue:2002:TMT

[XC02] Jingling Xue and Wentong Cai. Time-minimal tiling when rise is larger than zero. Parallel Computing, 28(6):915–939, June 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/56/29/
abstract.html.

Xie:2017:AFA

[XCR17] Wei Xie, Yong Chen, and Philip C. Roth. ASA-FTL: an adaptive separation aware flash translation layer for solid state drives. Parallel Computing, 61(??):3–17, January 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116301181.

Xiang:2005:FTR

[XCS05] Dong Xiang, Ai Chen, and Jiaguang Sun. Faulttolerant routing and multicasting in hypercubes using a partial path setup. Parallel Computing, 31(3–4):389–411, March/ April 2005. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Xia:2010:FGP

[XDZL10] Fei Xia, Yong Dou, Dan Zhou, and Xin Li. Fine-grained parallel RNA secondary structure prediction using SCFGs on FPGA. Parallel Computing, 36(9):516–530, September 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Xu:1996:EPM

[XH96] Zhiwei Xu and Kai Hwang. Early prediction of MPP performance: The SP2, T3D, and Paragon experiences. Parallel Computing, 22(7):917–942, October 1, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=
                                        22&issue=7&aid=1085.

Xu:2022:PPM

[XKL^{+}22] Tianshi Xu, Vassilis Kalantzis, Ruipeng Li, Yuanzhe Xi, Geoffrey Dillon, and Yousef Saad. parGeMSLR: a parallel multilevel Schur complement low-rank preconditioning and solution package for general sparse matrices. Parallel Computing, 113(??):??, October 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819122000497.

Xuan:2017:ABD

[XLS^{+}17] Pengfei Xuan, Walter B.

Ligon, Pradip K. Srimani, Rong Ge, and Feng Luo. Accelerating big data analytics on HPC clusters using two-level storage. Parallel Computing, 61(??):18–34, January 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300631.

Xiao:2020:PON

[XLY^{+}20] Jian Xiao, Min Long, Ce Yu, Xin Zhou, and Li Ji. Performance optimization of non-equilibrium ionization simulations from MapReduce and GPU acceleration. Parallel Computing, 98(??): Article 102682, October 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300727.

Xu:2019:FNL

[XRW^{+}19] Liyang Xu, Xiaoguang Ren, Qian Wang, Xinhai Xu, and Xuejun Yang. Full-neighbor-list based numerical reproducibility method for parallel molecular dynamics simulations. Parallel Computing, 85(??):109– 118, July 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819119300754.

Xian:2011:MGP

[XT11] Wang Xian and Aoki Takayuki. Multi-GPU performance of incompressible flow computation by lattice Boltzmann method on GPU cluster. Parallel Computing, 37(9):521–535, September 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819111000214.

Xu:2007:BTS

[Xu07] J. Xu. Benchmarks on tera-scalable models for DNS of turbulent channel flow. Parallel Computing, 33(12):780–794, December 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Xue:1994:ANU

[Xue94] Jingling Xue. Automating non-unimodular loop transformations for massive parallelism. Parallel Computing, 20(5):711–728, May 11, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.

elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=5&aid=847.

Xue:1996:TNL

[Xue96] Jingling Xue. Transformations of nested loops with non-convex iteration spaces. Parallel Computing, 22(3):339–368, April 29, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=3&aid=1048.

Xue:1997:UTN

[Xue97] Jingling Xue. Unimodular transformations of non-perfectly nested loops. Parallel Computing, 22 (12):1621–1645, February 21, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
22&issue=12&aid=1114.

Xiaoguang:2015:GDL

[XXQ^{+}15] Ren Xiaoguang, Xu Xinhai, Wang Qian, Chen Juan, Wang Miao, and Yang Xuejun. GS-DMR: Low-overhead soft error detection scheme for stencil-based computation. Parallel Computing, 41(??):50–65, January 2015. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819114001409.

Xun:2021:HDP

[XZYQ21] Yaling Xun, Jifu Zhang, Haifeng Yang, and Xiao Qin. HBPFP-DC: a parallel frequent itemset mining using Spark. Parallel Computing, 101 (??):Article 102738, April 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120301198.

Yamamoto:1999:PES

[Y^{+}99] J. Yamamoto et al. Performance evaluation of SNAIL: a multiprocessor based on the simple serial synchronized multistage interconnection network architecture. Parallel Computing, 25(9): 1081–1103, September 1, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/25/18/

abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/25/18/
article.pdf.

Yang:2002:ABD

[Y^{+}02] Yanyan Yang et al. Agent based data management in digital libraries. Parallel Computing, 28(5):773–792, May 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/57/32/
abstract.html.

Yazdanpanah:2024:ALP

[YA24]  Fahimeh  Yazdanpanah

and Mohammad Alaei. An approach for lowpower heterogeneous parallel implementation of ALC-PSO algorithm using OmpSs and CUDA. Parallel Computing, 120(??): ??, June 2024. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781912400022X.

Yalamov:1997:SPA

[Yal97] Plamen Y. Yalamov. Stability of a partitioning algorithm for bidiagonal systems. Parallel Computing, 23(3):333–348, May 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=3&aid=1160.

Yang:2004:PFE

[Yan04] Xin-She Yang. Pattern formation in enzyme inhibition and cooperativity with parallel cellular automata. Parallel Computing, 30(5–6):741–751, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yasar:2001:NIM

[Yaş01] O. Yaşar. A new ignition model for sparkignited engine simulations. Parallel Computing, 27(1–2):179–200, January 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/47/25/30/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/47/25/30/
                                        article.pdf.

Yzelman:2011:TDC

[YB11] A. N. Yzelman and Rob H. Bisseling. Two-dimensional cache-oblivious sparse matrixvector multiplication. Parallel Computing, 37(12):

806–819, December 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819111001062.

Yu:2005:WMR

[YBM05] Han Yu, Xin Bai, and Dan C. Marinescu. Workflow management and resource discovery for an intelligent grid. Parallel Computing, 31(7):797–811, July 2005. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yamazaki:2019:PAO

[YCBD19] Ichitaro Yamazaki, Edmond Chow, Aurelien Bouteiller, and Jack Dongarra. Performance of asynchronous optimized Schwarz with one-sided communication. Parallel Computing, 86(??):66–81, August 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301261.

Yang:2007:DEP

[YCS07] Chao-Tung Yang, Kuan-Wei Cheng, and Wen-Chung Shih. On development of an efficient parallel loop self-scheduling for grid computing environments. Parallel Computing, 33(7–8):467–487, August 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yeckel:1997:PCI

[YD97] Andrew Yeckel and Jeffrey J. Derby. Parallel computation of incompressible flows in materials processing: Numerical experiments in diagonal preconditioning. Parallel Computing, 23 (9):1379–1400, November 3, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=9&aid=1217.

Yoo:2001:EPM

[YD01] Byung S. Yoo and Chita R. Das. Efficient processor management schemes for mesh-connected multicomputers. Parallel Computing, 27(8):1057–1078, July 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/34/27/
abstract.html;   http:
//www.elsevier.nl/gej-

ng/10/35/21/47/34/27/
article.pdf.

Yasar:2007:TPC

[YD07] Osman Yaşar and Hasan Daḡ. Trends in parallel computing. Parallel Computing, 33(2):81–82, March 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yasar:2001:NTH

[YDTS01] O. Yaşar, Y. Deng, R. E.

Tuzun, and D. Saltz. New trends in high performance computing. Parallel Computing, 27(1–2):1–2, January 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/47/25/22/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/25/22/
article.pdf.

Yalamov:1994:FSM

[YE94] P. Yalamov and D. J.

Evans. On the forward stability of a modified ‘stride of 3’ reduction method. Parallel Computing, 20(8):1175–1190, August 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.

cgi?year=1994&volume=
20&issue=8&aid=887.

Yalamov:1995:MFM

[YE95] P. Yalamov and D. J. Evans. The W Z matrix factorisation method. Parallel Computing, 21 (7):1111–1120, July 11, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=7&aid=954.

Yoon:1997:IPR

[YEC97] Hyeon-Ju Yoon, Seongbae Eun, and Jung Wan Cho. Image parallel ray tracing using static load balancing and data prefetching. Parallel Computing, 23(7):861–872, July 14, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=7&aid=1188.

Yamashita:2003:SSP

[YFK03] M. Yamashita, K. Fujisawa, and M. Kojima. SD-PARA: SemiDefinite Programming Algorithm paR-Allel version. Parallel Computing, 29(8):1053– 1067, August 2003. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yang:1993:LSC

[YG93] T. Yang and A. Gerasoulis. List scheduling with and without communication delays. Parallel Computing, 19(12):1321–1344, December 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Young:2019:MCE

[YHE^{+}19] Jeffrey S. Young, Eric Hein, Srinivas Eswar, Patrick Lavin, Jiajia Li, Jason Riedy, Richard Vuduc, and Tom Conte. A microbenchmark characterization of the Emu Chick. Parallel Computing, 87(??):60–69, September 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118302503.

Yu:2021:PPB

[YHZ21] Zhixing Yu, Kejing He, and Xiuhong Zou. PEAB: a pool-based distributed evolutionary algorithm model with buffer. Parallel Computing, 106(??):??, September 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121000582.

Yuba:1987:JNP

[YK87] T. Yuba and H. Kashiwagi. The Japanese national project for new generation supercomputing systems. Parallel Computing, 4(1):1–16, February 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yanev:2006:EAE

[YK06] Petko Yanev and Erricos John Kontoghiorghes. Efficient algorithms for estimating the general linear model. Parallel Computing, 32(2):195–204, February 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yanev:2008:PAD

[YK08] Petko I. Yanev and Erricos J. Kontoghiorghes. Parallel algorithms for downdating the least squares estimator of the regression model. Parallel Computing, 34(6–8):451–468, July 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yang:1986:MDA

[YL86] C. B. Yang and R. C. T.

Lee. The mapping of 2-D array processors to 1-D array processors. Parallel Computing, 3(3):217– 229, July 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yu:2019:MEM

[YL19] Ting Yu and Mengchi Liu.

A memory efficient maximal clique enumeration method for sparse graphs with a parallel implementation. Parallel Computing, 87(??):46–59, September 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301297.

Yirang:2018:DDM

[YLCT18] Yuan Yirang, Chang Luo, Li Changfeng, and Sun Tongjun. Domain decomposition modified with characteristic mixed finite element of compressible oil-water seepage displacement and its numerical analysis. Parallel Computing, 79(??):36–47, November 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0167819118302643.

Yong:1995:AEA

[YLE95] Liu Yong, Kang Lishan, and D. J. Evans. The annealing evolution algorithm as function optimizer. Parallel Computing, 21(3):389–400, March 10, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=
                                        21&issue=3&aid=944.

Yang:2017:PRR

[YLF^{+}17] Lei Yang, Yilong Li, Zhenxin Fu, Zhuohan Li, Wenbin Hou, Haoze Wu, Xiaolin Wang, and Yun Liang. ParConnect reproducibility report. Parallel Computing, 70(??):22–26, December 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117300935.

Yu:2023:APR

[YLT^{+}23] Bin Yu, Xu Lu, Cong Tian, Meng Wang, Chu Chen, Ming Lei, and Zhenhua Duan. Adaptively parallel runtime verification based on distributed network for temporal properties. Parallel Computing, 117(??):??, September 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000406.

Yu:2013:EPS

[YLW^{+}13] Jiyang Yu, Peng Liu, Weidong Wang, Chunming Huang, Jie Yang, Yingtao Jiang, and Qingdong Yao. An efficient protocol with synchronization accelerator for multi-processor embedded systems. Parallel Computing, 39(9):461–474, September 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113000525.

Yu:2005:SMP

[YM05] Hongfeng Yu and Kwan-Liu Ma. A study of I/O methods for parallel visualization of large-scale data. Parallel Computing, 31(2):167–183, February 2005. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yavits:2014:ECS

[YMG14] L. Yavits, A. Morad, and R. Ginosar. The effect of communication and synchronization on Amdahl’s law in multicore systems. Parallel Computing, 40(1):1–16, January 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819113001324.

Yalamanchili:2010:AHB

[YMJT10] Pavan Yalamanchili, Sumod Mohan, Rommel Jalasutram, and Tarek Taha. Acceleration of hierarchical Bayesian network based cortical models on multicore architectures. Parallel Computing, 36(8):449– 468, August 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yoritaka:2019:PGM

[YMY^{+}19] Hiroshi Yoritaka, Ken Matsui, Masahiro Yasugi, Tasuku Hiraishi, and Seiji Umatani. Probabilistic guards: a mechanism for increasing the granularity of work-stealing programs. Parallel Computing, 82(??):19–36, ???? 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301881.

Yoshida:2024:AIC

[YMYH24] Kohei Yoshida, Shinobu Miwa, Hayato Yamaki, and Hiroki Honda. Analyzing the impact of CUDA versions on GPU applications. Parallel Computing, 120(??):??, June 2024. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S016781912400019X.

Yagawa:1997:LSF

[YNO97] Genki Yagawa, Yasushi Nakabayashi, and Hiroshi Okuda. Large-scale finite element fluid analysis by massively parallel processors. Parallel Computing, 23(9):1365–1377, November 3, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=9&aid=1216.

Yoshida:2017:CCE

[YNY^{+}17] Ryuji Yoshida, Seiya Nishizawa, Hisashi Yashiro, Sachiho A. Adachi, Yousuke Sato, Tsuyoshi Yamaura, and Hirofumi Tomita. CONeP: a cost-effective online nesting procedure for regional atmospheric models. Parallel Computing, 65(??):21–31, July 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S016781911730039X.

Yamanaka:2008:PAA

[YORO08] N. Yamanaka, T. Ogita, S. M. Rump, and S. Oishi. A parallel algorithm for accurate dot product. Parallel Computing, 34(6–8): 392–410, July 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yang:2003:PDS

[YPG03] Laurence T. Yang, Yi Pan, and Minyi Guo. Parallel and distributed scientific and engineering computing. Parallel Computing, 29(11–12): 1505–1508, November/ December 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yang:2012:CSC

[YPL12] Depeng Yang, Gregory. D. Peterson, and Husheng Li. Compressed sensing and Cholesky decomposition on FPGAs and GPUs. Parallel Computing, 38(8):421–437, August 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819112000129.

Yan:2018:SSP

[YR18] Beichuan Yan and Richard A.

Regueiro. Superlinear speedup phenomenon in parallel 3D Discrete Element Method (DEM) simulations of complexshaped particles. Parallel Computing, 75(??):61–87, July 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300826.

Yan:1996:APP

[YS96] Jerry C. Yan and Sekhar R.

Sarukkai. Analyzing parallel program performance using normalized performance indices and trace transformation techniques. Parallel Computing, 22 (9):1215–1237, November 22, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=9&aid=1083.

Yanagawa:2004:SSE

[YS04] Takashi Yanagawa and Kenji Suehiro. Software system of the Earth Simulator. Parallel Computing, 30(12):1315–1327, December 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yuan:1997:LBP

[YSBM97] X. Yuan, C. Salisbury, D. Balsara, and R. Melhem. Load balancing package on distributed memory systems and its application to particle-particle particle-mesh (P3M) methods. Parallel Computing, 23(10):1525–1544, November 7, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=10&aid=1221.

Yang:1995:AWR

[YTCS95] C. S. Yang, Y. M. Tsai, S. L. Chi, and Shepherd S. B. Shi. Adaptive wormhole routing in k-ary ncubes. Parallel Computing, 21(12):1925–1943, December 12, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.

cgi?year=1995&volume=
21&issue=12&aid=1032.

Yang:1997:UKB

[YTCS97] Chao-Tung Yang, Shian-Shyong Tseng, Cheng-Der Chuang, and Wen-Chung Shih. Using knowledgebased techniques on loop parallelization for parallelizing compilers. Parallel Computing, 23(3):291–309, May 15, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=
23&issue=3&aid=1141.

Yang:2007:PCO

[YTCW07] Jinn-Shyong Yang, Shyue-Ming Tang, Jou-Ming Chang, and Yue-Li Wang. Parallel construction of optimal independent spanning trees on hypercubes. Parallel Computing, 33(1): 73–79, February 2007. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yamazaki:2013:PCP

[YTSI13] Ichitaro Yamazaki, Hiroto Tadano, Tetsuya Sakurai, and Tsutomu Ikegami. Performance comparison of parallel eigensolvers based on a contour integral method and a Lanczos method. Parallel Computing, 39(6–7):280–290, June/July 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819112000233.

Yuan:1994:EFT

[Yua94] Shyan-Ming Yuan. An efficient fault-tolerant decentralized commit protocol. Parallel Computing, 20(1):101–114, January 16, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=1&aid=818.

Yuen:1997:PPC

[Yue97]  C. K. Yuen.  Parallel

programming — a critique. Parallel Computing, 23(3):369–380, May 15, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=
                                        23&issue=3&aid=1143.

Yu:2011:CGH

[YWP11] Chenhan D. Yu, Weichung Wang, and Dan’l Pierce. A CPU–GPU hybrid approach for the unsymmetric multifrontal method. Parallel Computing, 37(12):759–770, December 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819111001293.

Yamauchi:2007:PVC

[YX07] Hiroshi Yamauchi and Dongyan Xu. Portable virtual cycle accounting for large-scale distributed cycle sharing systems. Parallel Computing, 33(4–5): 314–327, May 2007. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yan:2021:QQT

[YXQ^{+}21] Baicheng Yan, Limin Xiao, Guangjun Qin, Zhang Yang, Bin Dong, Haonan Yu, and Hongyu Wu. QTMS: a quadratic time complexity topologyaware process mapping method for large-scale parallel applications on shared HPC system. Parallel Computing, 94–95 (??):Article 102637, June 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300302.

Yu:2023:SBP

[YZB^{+}23] Lei Yu, Tianqi Zhong, Peng Bi, Lan Wang, and Fei Teng. Segment based power-efficient scheduling for real-time DAG tasks on edge devices. Parallel Computing, 116(??):??, July 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819123000285.

Yin:2014:SNS

[YZWcF14] Jiangling Yin, Junyao Zhang, Jun Wang, and Wu chun Feng. SDAFT: a novel scalable data access framework for parallel BLAST. Parallel Computing, 40(10):697–709, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819114000957.

Zhou:2021:MAG

[ZAA^{+}21] Keren Zhou, Laksono Adhianto, Jonathon Anderson, Aaron Cherian, Dejan Grubisic, Mark Krentel, Yumeng Liu, Xiaozhu Meng, and John Mellor-Crummey. Measurement and analysis of GPUaccelerated applications with HPCToolkit. Parallel Computing, 108(??):??, December 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819121000843.

Zambonelli:1999:EBL

[Zam99] Franco Zambonelli. Exploiting biased load information in direct-neighbour load balancing policies. Parallel Computing, 25(6): 745–766, June 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.
com/cas/tree/store/parco/
sub/1999/25/6/1422.pdf.

Zima:1994:SVF

[ZBC94] Hans P. Zima, Peter Brezany, and Barbara M. Chapman. SUPERB and Vienna Fortran. Parallel Computing, 20(10– 11):1487–1517, November 3, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=10-11&aid=909.

Zima:1988:STS

[ZBG88] H. P. Zima, H.-J. Bast, and M. Gerndt. SUPERB: a tool for semi-automatic MIMD/SIMD parallelization. Parallel Computing, 6(1):1–18, January 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zhong:2022:ULV

[ZCBD22] Dong Zhong, Qinglei Cao, George Bosilca, and Jack Dongarra. Using long vector extensions for MPI reductions. Parallel Computing, 109(??):??, March 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819121001137.

Zanghirati:2000:CTI

[ZCPT00] G. Zanghirati, F. Cocco, G. Paruolo, and F. Taddei. A Cray T3E implementation of a parallel stochastic dynamic assets and liabilities management model. Parallel Computing, 26(5):539– 567, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/27/25/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/27/25/
                                        article.pdf.

Zhou:2020:TTE

[ZCT^{+}20] Yi Zhou, Yuanqi Chen, Shubbhi Taneja, Ajit Chavan, Xiao Qin, and Jifu Zhang. ThermoBench: a thermal efficiency benchmark for clusters in data centers. Parallel Computing, 98(??):Article 102671, October 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300636.

Zhongzhi:1995:MAP

[ZDE95] Bai Zhongzhi, Wang Deren, and D. J. Evans. Models of asynchronous parallel matrix multisplitting relaxed iterations. Parallel Computing, 21(4):565–582, April 1, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=4&aid=946.

Zhang:1992:PSC

[ZE92] G. Zhang and H. C. Elman. Parallel sparse Cholesky factorization on a shared memory multiprocessor. Parallel Computing, 18(9):1009–1022, September 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zhang:2017:ESA

[ZEC^{+}17] Boyu Zhang, Trilce Estrada, Pietro Cicotti, Pavan Balaji, and Michela Taufer. Enabling scalable and accurate clustering of distributed ligand geometries on supercomputers. Parallel Computing, 63(??):38– 60, April 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819117300261.

Zomaya:2004:PNI

[ZEgT04] Albert Y. Zomaya, Fikret Ercal, and El ghazali Talbi. Parallel and nature-inspired computational paradigms and applications. Parallel Computing, 30(5–6):551–552, May/June 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zenios:1999:HPC

[Zen99] Stavros A. Zenios. Highperformance computing in finance: The last 10 years and the next. Parallel Computing, 25(13– 14):2149–2175, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.elsevier.nl/gej-
                                        ng/10/35/21/32/36/47/

abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/47/
article.pdf.

Zerovnik:1990:PVH

[Zer90] J. Zerovnik. A parallel variant of a heuristical algorithm for graph colouring. Parallel Computing, 13(1):95–100, January 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zmijewski:1988:PAS

[ZG88] Earl Zmijewski and John R.

Gilbert. A parallel algorithm for sparse symbolic Cholesky factorization on a multiprocessor. Parallel Computing, 7(2):199–210, June 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zhu:2016:PMC

[ZG16] Yao Zhu and David F.

Gleich. A parallel mincut algorithm using iteratively reweighted least squares targeting at problems with floating-point edge weights. Parallel Computing, 59(??):43–59, November 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0167819116000545.

Zubair:1992:VPA

[ZGG92] M. Zubair, S. N. Gupta, and C. E. Grosch. A variable precision approach to speedup iterative schemes on fine grained parallel machines (short communication). Parallel Computing, 18(11):1223–1231, November 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zhu:2019:DES

[ZGL^{+}19] Guanghui Zhu, Chen Guo, Le Lu, Zhi Huang, Chunfeng Yuan, Rong Gu, and Yihua Huang. DGST: Efficient and scalable suffix tree construction on distributed data-parallel platforms. Parallel Computing, 87(??):87–102, September 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819119301255.

Zhou:2020:CHM

[ZGZS20] Huan Zhou, José Gracia, Naweiluo Zhou, and Ralf Schneider. Collectives in hybrid MPI+MPI code: Design, practice and performance. Parallel Computing, 99(??):

Article 102669, November 2020. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819120300612.

Zhang:1991:APD

[Zha91] Hong Zhang. On the accuracy of the parallel diagonal dominant algorithm. Parallel Computing, 17(2–3):265–272, June 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zhang:1995:POR

[Zha95] Xiaodong Zhang. Parallelizing an oil refining simulation: Numerical methods, implementations and experience. Parallel Computing, 21(4):627–647, April 1, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=4&aid=961.

Zhang:2023:EES

[ZHC^{+}23] Hao Zhang, Zhiyi Huang, Yawen Chen, Jianguo Liang, and Xiran Gao. ESA: an efficient sequence alignment algorithm for biological database search on Sunway TaihuLight. Parallel Computing, 117(??):??, September 2023. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819123000492.

Zhou:1993:TSW

[Zho93] H. B. Zhou. Two-stage m-way graph partitioning. Parallel Computing, 19(12):1359–1373, December 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zhu:1993:FRL

[Zhu93] Jianping Zhu. QR factorization for the regularized least squares problem on hypercubes. Parallel Computing, 19(8):939–948, August 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zheng:2023:OMP

[ZJD^{+}23] Jiang Zheng, Jiazhi Jiang, Jiangsu Du, Dan Huang, and Yutong Lu. Optimizing massively parallel sparse matrix computing on ARM many-core processor. Parallel Computing, 117(??):??, September 2023. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819123000418.

Zhang:2018:IRP

[ZJDW18] Xuechen Zhang, Song Jiang, Alseny Diallo, and Lei Wang. IR+: Removing parallel I/O interference of MPI programs via data replication over heterogeneous storage devices. Parallel Computing, 76(??):91–105, August 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118300140.

Zerovnik:1992:PVH

[ZK92] J. Zerovnik and M. Kaufman. A parallel variant of a heuristical algorithm for graph coloring — corrigendum (short communication). Parallel Computing, 18(8):897–900, August 1992. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zhu:2004:KMS

[ZKCL04] Ming Zhu, Constantine Katsinis, Wentong Cai, and Bu-Sung Lee. Key messaging on SOME-bus clusters. Parallel Computing, 30(8):947–971, August 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zhang:1994:ASC

[ZL94] Bao Lin Zhang and Wen Zhi Li. On Alternating Segment Crank–Nicolson scheme (short communication). Parallel Computing, 20(6):897–902, June 10, 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/
                                        cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=
                                        20&issue=6&aid=846.

Zhou:2004:MPM

[ZL04] Jipeng Zhou and Francis C. M. Lau. Multi-phase minimal fault-tolerant wormhole routing in meshes. Parallel Computing, 30(3): 423–442, March 2004. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zlatev:1988:TSM

[Zla88]  Zahari Zlatev.  Treat-

ment of some mathematical models describing longrange transport of air pollutants on vector processors. Parallel Computing, 6(1):87–98, January 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zhang:2015:DDP

[ZLD15] Peng Zhang, Ling Liu, and Yuefan Deng. A datadriven paradigm for mapping problems. Parallel Computing, 48(??):108– 124, October 2015. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0167819115000757.

Zhang:1993:SAD

[ZLJ93] C. N. Zhang, H. F. Li, and R. Jayakumar. A systematic approach for designing concurrent errordetecting systolic arrays using redundancy. Parallel Computing, 19(7):745– 764, July 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zaidi:1998:IEM

[ZLM98] Habib Zaidi, Claire Labbé, and Christian Morel. Implementation of an environment for Monte Carlo simulation of fully 3-D positron tomography on a high-performance parallel platform. Parallel Computing, 24(9– 10):1523–1536, September 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

                                        store/parco/sub/1998/
                                        24/9-10/1346.pdf.

Zhang:2016:TSS

[ZLPF16] Ziming Zhang, Michael Lang, Scott Pakin, and Song Fu. TracSim: Simulating and scheduling trapped power capacity to maximize machine room throughput. Parallel Computing, 57(??):108–124, September 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819115001489.

Zafari:2019:DEP

[ZLT19] Afshin Zafari, Elisabeth Larsson, and Martin Tillenius. DuctTeip: an efficient programming model for distributed task-based parallel computing. Parallel Computing, 90(??): Article 102582, December 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819119301735.

Zenios:1988:DAC

[ZM88a] Stavros A. Zenios and John M. Mulvey. A distributed algorithm for convex network optimization problems. Parallel Computing, 6(1):45–56, January 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zubair:1988:ESA

[ZM88b] M. Zubair and B. B.

Maden. Efficient systolic algorithm for finding bridges in a connected graph. Parallel Computing, 6(1):57–61, January 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zhang:1994:UPB

[ZM94] Hong Zhang and William F.

Moss. Using parallel banded linear system solvers in generalized eigenvalue problems. Parallel Computing, 20(8):1089–1105, August 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=
20&issue=8&aid=889.

Ziavras:1996:DBR

[ZM96] Sotirios G. Ziavras and Arup Mukherjee. Data broadcasting and reduction, prefix computation, and sorting on reduced hypercube parallel computers. Parallel Computing, 22(4):595–606, June 11, 1996. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1996&volume=
22&issue=4&aid=1059.

Zhang:1990:BAG

[ZMMW90] Xiru Zhang, M. McKenna, J. P. Mesirov, and D. L. Waltz. The backpropagation algorithm on grid and hypercube architectures. Parallel Computing, 14(3):317–327, August 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zomaya:1995:PPR

[Zom95] Albert Y. Zomaya. Parallel processing for robot dynamics computations. Parallel Computing, 21 (4):649–668, April 1, 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/
cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=
21&issue=4&aid=973.

Zeng:2016:AMT

[ZP16] Jiaan Zeng and Beth Plale. Argus: a multitenancy NoSQL store with workload-aware resource reservation. Parallel Computing, 58(??):76–89, October 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300588.

Zapata:1999:DPS

[ZPAT99] E. L. Zapata, O. Plata, R. Asenjo, and G. P. Trabado. Data-parallel support for numerical irregular problems. Parallel Computing, 25(13– 14):1971–1994, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-
ng/10/35/21/32/36/40/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/32/36/40/
article.pdf.

Zapatam:1989:PFC

[ZRPI89] E. L. Zapatam, F. F.

Rivera, O. G. Plata, and M. A. Ismail. Parallel fuzzy clustering on fixed size hypercube SIMD computers. Parallel Computing, 11(3):291–303, ???? 28, 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zhao:2013:RWS

[ZRS13] Laiping Zhao, Yizhi Ren, and Kouichi Sakurai. Reliable workflow scheduling with less resource redundancy. Parallel Computing, 39(10):567–585, October 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819113000732.

Zha:2021:IPS

[ZS21] Benbo Zha and Hong Shen. Improved probabilistic I/O scheduling for limitedsize burst-buffers deployed HPC. Parallel Computing, 101(??):Article 102708, April 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0167819120300946.

Zockler:1997:PLI

[ZSH97] Malte Zöckler, Detlev Stalling, and Hans-Christian Hege. Parallel line integral convolution. Parallel Computing, 23(7):975–989, July 14, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/parco/free/1997/
                                        23/7/zockler/fig9.jpg.

Zawada:2002:CHC

[ZSI02] A. C. Zawada, N. L. Seed, and P. A. Ivey. Continuous and high coverage self-testing of dynamically re-configurable systems. Parallel Computing, 28(7–8):1155–1178, August 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/gej-
ng/10/35/21/60/58/38/
abstract.html.

Zubair:1990:OSA

[Zub90] M. Zubair. An optimal speedup algorithm for the measure problem. Parallel Computing, 13(1):61– 71, January 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zlatev:1988:CSP

[ZVWS88] Z. Zlatev, Phuong Vu, J. Wasniewski, and K. Schaumburg. Computations with symmetric, positive definite and band matrices on a parallel vector processor. Parallel Computing, 8(1– 3):301–312, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zhou:2019:VVA

[ZWJ^{+}19] Fang Zhou, Song Wu, Youchuang Jia, Xiang Gao, Hai Jin, Xiaofei Liao, and Pingpeng Yuan. VAIL: a victim-aware cache policy to improve NVM lifetime for hybrid memory system. Parallel Computing, 87(??):70–76, September 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819118301765.

Zeng:2016:SGP

[ZY16] Jianping Zeng and Hongfeng Yu. A study of graph partitioning schemes for parallel graph community detection. Parallel Computing, 58(??):131–139, October 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300400.

Zhou:2016:ABA

[ZYZ^{+}16] Zhou Zhou, Xu Yang, Dongfang Zhao, Paul Rich, Wei Tang, Jia Wang, and Zhiling Lan. I/O-aware bandwidth allocation for petascale computing systems. Parallel Computing, 58(??):107–116, October 2016. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0167819116300382.

Zanghirati:2003:PSL

[ZZ03] G. Zanghirati and L. Zanni.

A parallel solver for large quadratic programs in training support vector machines. Parallel Computing, 29(4):535–551, April 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).